A Complete Bibliography of Publications in Numerische Mathematik (1959–1999)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 3.42

Title word cross-reference

(−t^{2})dt [Luk61]. (0, 1) [Els84a]. (0, 1) [Sch87a]. (1 + z/n)^{n} [vI84]. (A − λB)x = 0 [Sch74b]. (k, l) [Gut89b]. (λ^{2} − λA − B)x = 0 [Lin76]. (λ^{2}I − λA − B)x = 0 [Ric74b]. (λ^{2}λA − B)x = 0 [Lin76]. (m, s) [LA89]. (m > 0.5) [LMW84]. (n, n) [Lub89b]. (∂^{2}z/∂x^{2})(∂^{2}z/∂y^{2}) − ((∂^{2}z/∂x∂y))^{2} = f [OP88]. 1 [ASW93, Chr70a, Chr70b, Jeg68]. 1 + 2 [Wer79, Wer82]. 1/t [And88]. 1 ≤ n ≤ 3 [Koh87, Koh88]. 2 [CFS97, Mal99]. 2^{k} [BL81]. 2m [Guo88, Kos85]. 3 [Jay93, Str66]. 3 × 3 [Bat90]. 4k + 1 [VC92a]. << [Leb70]. >> [Leb70]. A [Alb85, Hai82, Hai86, PW93]. A(Θ) [Sav93]. A_{0} [Jel79]. AB [Kub84]. AX + Y B = C [Zie85]. Ax = b [Har80].

Ax = λBx [Rod73, MW68b]. AX ≈ B [PW93]. B [CS94, DKS90a, DKS90b, FHL86a, FHL86b, HS81, Hun86, vV83, Cro79, Hai86, HS90, SU89, Sch87c]. β [Sto84]. C [Nør75, Lor92]. C(λ) [Eck76]. C[0, 1] [Sab78]. C^{0} [Ley86]. C^{1} [Gme87, KK97, Zha89]. C^{2} [Cle92, SS90, SH95]. C^{∘} [Bak82]. C^{ρ} [LLS95]. C_{0}(a, b) [Pal76]. C_{0}[a, b] [Pal76]. C(Ω) [Ben86]. C^{∈} [AL93]. J [Hom95]. cos(z) [Kap96]. CS [Ste82, Van85]. D [vV83, Wil62e]. d = 3r + 1 [AS90]. D^{m} [AY93, LA88]. e^{−z} [SSV76a, SSV76b, SVN76]. e^{z} [SV75, SV76a, SV78, VC94]. E_{n} [Wil62e]. ℓ_{1} [SW87]. ℓ_{∞} [BL78]. ϵ [Bre71, Bre78, Cor77, Pyl67]. exp(−z) [vI84]. exp(z) [Hai82]. f [Ans73]. f^{(n)}(0) [Str73b].

G [Med73, Gra68]. GT H [O’C96]. H [Lyn63, Bea86a, Mes95, BD81, GB86a, GB86b, GB86c, Guo88, GC96, GS98c, Li90, SS96, Bea79, FS92]. h − p [GB86a, GB86b, GB86c]. H_{0}^{1} [Hav84]. H^{2} [Auz88]. H^{p} [AB84]. H_{p} [Sik82b, Ste78a]. hp [HMS99, SS97, SSS99]. hZ^{s} [RS95]. IIIC [Sch87c]. J [Ves76b, DH93b]. k [BL81, GNV86, Jac85, LP95, NV83]. K(1/b_{n}) [FJ72, Fie78]. K(a_{n}/1) [BJ85b, TW80]. KCx = λx [BRS71]. k × k [LP95]. k → ∞ [LP95]. L [AD74, Gro89, Luc70, Pre71, EG94]. L(λ)u + ϵR(λ, u) = 0 [DC74]. L^{1} [SO87]. L^{2} [Ber73, Cas85, DDW75, Vui90]. L^{∞} [CD85, Fet87, GN87, Koh87, Koh88, Noc88, SO87, SF86, Wan89]. L^{p} [ES90]. L^{q} [DDW75]. L_{1} [BY66, Car72b, Str82]. L_{2} [BS79, LTT92]. L_{∞} [BY66, Hav84, KJ90]. L_{p} [Fis81, Lam70, Wol79]. L_{p}(0, t_{i}) [SW96]. λ [GK90, CNV88]. lim a_{n} = 0 [BJ85b]. LR [RS63, DT71, MW68c, PW70, Ves99]. LU [FNP82, Mes95, O’C96, Heg91, VC81]. M [Bea86a, Bod68, Kel72, Rob66, Sch73a, Bea79, Bod68, VC81]. M/G/1 [Mei97]. N [Hek87, LM70, Bro65b, ELR92, GR70b, Kle81b, Rut60, Str66, Str68, Mar84]. ∇ · a∇u = f [Gun64a]. n → ∞ [Lub89b]. P [Cas81a, Dix82, ENR90, Kel72, Mar62, Zen86, AK99, BD81, BS89, GB86a, GB86b, GB86c, Guo88, GC96, GS98c, HP94, Heu98, Heu01, Hig92, Jen91, LV89, Li90, LK78, MP96a, Neu61, OD95, Pav94a, Pav94b, PHP90, Sar86, Sar90, SS96, Tee64, Vog83a, Vog83b, Gue86]. P1 [Osw92]. P_{0} [Wil71]. P_{1} [Won93]. P_{k} [Gue86]. P ≤ A ≤ Q [Har80]. p ≤ b ≤ q [Har80]. Q_{1} [AKP95]. Q_{k′} [Gue86]. qd [FP94, Wuy73]. QL [BMRW68, Dub70, MW68a]. QR [BMRW68, RB68, Bat90, BBdH86, BGBM89, CMR86, Leb70, MRW70, MPW70, Par65a, PW70, Sac72, Ste99]. R [BS88, AS90]. R^{2} [GH92]. R^{3} [Néd86, SS97, Néd80, Néd82]. R^{k} [GM82].

R^{n} [Ste75a, Sus64, Gou77, Koh87, Koh88, LA88, Rua88, Sch95a]. R_{3} [Wen68, Wen68]. ρ [Bre71, Osa96]. S [Ver77b, For68b]. DQR [Uhl97]. Σ^{2} [Tim65]. sin(z) [Kap96]. T [WB69]. θ [Jac84, BJ86, CG88a, Liu95]. U(a, b, z) [Tem83]. u(x) [Sch84]. u′′(x) [Sch84]. u_{t} = urr + \frac{2}{r}u_{r} [Eis67b]. u ∈ W_{2}^{m} [LMW84]. V [Reu88b]. ε [Bre75a, Cuy82]. W^{m,2} [Kos85]. ℘ [Jar94]. ℘(z|e_{1}, e_{2}, e_{3}) [Tod90]. X [SBH84]. x = 0 [Mid75]. y′′ + (λ + 2λ1cosχ + 2λ2cos2χ)y = 0 [Sto67]. y′′ = f(x, y) [Hai77]. Z [KvdD77]. z^{2} ∫_{0}^{z} [Luk61].

-adic [Dix82]. -algorithm

[Bre75a, Bre78, Cor77, Cuy82, Kub84, Osa96, Pyl67, Wuy73]. -algorithme [Bre75a]. -algorithmes [Bre71]. -algorithms [Bre71]. -Approximation [Lam70, WB69, LTT92]. -approximations [Str82]. -body [Mar84]. -class [Sto84]. -collocation-like [Ley86]. -condition [Mar62]. -Convergence [Ans73, DKS90a, Sch87c, DKS90b, FHL86a, FHL86b, Hun86]. -convergent [HS90]. -cubic [KK97]. -curve [EG94]. -cycle [Reu88b]. -cyclic [HP94, LV89, PHP90, Sar86, Sar90, Tee64, ENR90]. -d-elasticity [CFS97]. -dimensional [For68b, Str66, Hek87, LM70]. -dimensions [GR70b]. -error [Koh88, Kos85, Fet87, GN87, Koh87, Noc88, Vui90, Wan89]. -factorizations [Mes95]. -Fehlerabschätzungen [Kos85]. -FEM [SSS99]. -fractions [Gra68, Lor92]. -function [Jar94]. -Galerkin [SS97]. -generating [Med73]. -generator [LK78]. -groups [Neu61]. -Gruppen [Neu61]. -interpolants [AL93]. -Konvergenz [Ans73]. -line [LP95]. -matrice [Rob66]. -matrices [Lyn63, VC81, Mes95]. -matrix [GK90, Rob66, KvdD77]. -Matrizen [Bod68]. -methods [Alb85, BJ86, CG88a, Jac84, Liu95]. -minimal [Wil62e]. -monosplines [Gro89].

-norm [Hig92, Mal99, CD85]. -operators [Bea79, Bea86a]. -order [BS88]. -periodic [Jac85]. -polynomials [Nør75]. -projection [DDW75]. -projections [Hav84]. -Projektionen [Hav84]. -regularity [Auz88]. -Schritt-quadratische [Kle81b]. -series [CS94]. -sphere [Str68]. -Spline-Interpolierenden [Sch73a]. -splines [AY93, LA88, LA89, SU89, SS90, AD74, Gme87, Luc70, Pre71, SBH84]. -splittings [FS92]. -Stabilität [Hav84]. -stabilité [Cro79]. -Stability [Ans73, Cro79, Hai86, Jel79, Hav84, HS81, Ver77b, Zen86, vV83]. -stable [Hai82, Sav93, Cas81a]. -step [GNV86, Gut89b, Kle81b, NV83]. -Stetigkeit [Ans73]. -Symmetric [DH93b, Ves76b]. -ter [Kos85, Rut60]. -th [Rut60]. -version [AK99, BS89, GB86a, Heu98, HMS99, Heu01, MP96a, OD95, Pav94a, Pav94b, Vog83a, Vog83b].

1 [Ber60]. 11/27/1978 [BGHS79]. 1333 [Lod70]. 1978/1979 [Bea86a]. 1994 [FV94].

2 [Hal60a]. 22.8.1970 [GB71]. 25.5.1980 [GB81]. 25th [Ano83b]. 26 [SV77]. 2D [CHS96, Roh98].

31 [Bea86a]. 37 [JN82a]. 396 [Heu01]. 3D [AS97b, Per99].

4/21/1909 [BGHS79].

5 [Aro63a]. 50 [Koh88]. 52 [MC88a]. 58 [CS92, Lub92b].

60 [AKK^{+}68, BBG^{+}60]. 68 [vWMPK69].

79 [Heu01].

81e [Bea86a]. 82g [JN82a]. 88i [Koh88]. 88k [MC88a].

91m [CS92]. 92d [Lub92b].

A-posteriori [Sch96, Fae98]. A-priori [Gek78]. A-stable [Mat76]. A.D.I. [GM68]. Abbildung [Men74, Weg78]. Abbildungen [Bod68, Wei80]. Abbrechen [Lod70]. Abel [BJ86, BN81, Egg84, Egg88, GK66]. Abel-type [Egg84, GK66]. Abhängigkeit [Lam70, Nit62]. Ableitsungfreie [SV70]. Ableitungen [Kan73]. ableitungsfreie [Hae63, Häm65]. ableitungsfreien [Eng77]. ableitungsfreies [Kre71]. above [Mar87]. ABS [AGS87]. Abschätzung [Bra71b, Ehr69, Häm66, Sch70a]. Abschätzungen [Bit68, Bit71]. Abscissas [NR68, Ask72]. absence [Car72a]. Absolute [BSW61, BH61a, Kra86, vdGK86, Cas82]. Absorbing [HR95, HR87, TV98, Wag83a]. Abstiegsverfahren [Sch78c]. abstract [DC99, GO95, Pal99, Sav93, Sre69, Sre71]. Abszissen [NR68]. abzählbaren [Red67]. Accelerated [Han91b, Pao98, Ber84, ID86, PS84]. Accelerating [Cuy83a, Ste69a, TW80, Wil87, Gab84, Sch66a]. Acceleration [Alb62, Bra65, Lou79, Mat90, Bra70, Bre78, Chu76, DGB80, Gus85, Ho90b, JW88, KS92, Nev90a, NW84, Ort86, Osa96, Sch68b, Wal91, DGB80]. accelerations [Nie91]. account [ABF99]. Accuracy [FH95, JN86, dFSS97, ADK91, BS99b, CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, Dum98, Egg86, FGM67, GS95a, GM68, Had69b, JN82b, Kel65, Kre59, Noc89, Raz69, Ste61, Ste90b, Str86, Ura70, ZL94, Zlá70]. Accurate [FP94, Str64, VD91, Deá80, LRZ95, Mat96]. achieve [MO68]. ACM [PS59]. acoustic [GJ89]. active [CCCD93]. acyclic [KvdD77]. Adams [Ans63, Ans63, Jac82]. adapted [FF87, Mel99]. adaption [WW84]. Adaptive [BD98, BS99b, LN99, Man78b, VL92, WW97, WhY88, BV84, BB82b, BSW88, CGR94, CFS97, Dör96, GW94, GW95a, Grü97, GB86c, HP75, JM87, Jer85, SW84, Ver77a]. Addendum [BF90, Dej66a, EP76]. addition [CH86]. Additive [Cai91a, Pav94a, BD97, Bre96, FS99, GO94b, GO95, Hac92, LM99]. ADI [GT67, LT98a, LR93]. adic [Dix82]. adjoint [BG95, Ber59a, HR94, Jer70]. adjusted [DHP99b]. adjustment [Dia80]. Admissible [EM87]. advantage [Del73]. advection [DHP99b, JK91b]. aerodynamics [EP86]. against [Heb88]. aggregation [MS94]. ahead [BZS99, FZ93, Fre94, GH95]. Aid [Bit68, Bit71, Ess60, Sau63, Zur64]. Airfoil [ML90]. Aitken [Müh78, Nie91]. Aitken-algorithm [Müh78]. Algebra [WR71, DC99]. Algebraic [AV89, ELV91, LL90, TV98, VV89, Wei92a, vdS70a, vdS70b, AGS87, AM94b, ASW93, Cam93, DKS90a, DKS90b, DHZ87, EMN88, FL91, Gau73, Hai78, Han88, HS90, Jay93, Kem68, KM98, Las92, Lub89a, Not98, RS86, RRS89, RDM62, Roc88, Sch73a, SG78, Sri94, gS97, gS98b, TK64, WW68, fX96]. algebraically [Hai82]. algebraische [Sch73a]. Algebras [Gas71b, Ser99a]. algèbres [Gas71b]. ALGOL [AKK^{+}68, BBG^{+}63, PS59, Wyn64, vWMPK69, BBG^{+}60]. Algorithm [AL85, Bar91a, BW72b, BD68, DH89, Hua90, JP90, KS73, Lau67, MW68a, MW68c, MPW70, Rut63b, Smi91, Tha90, VH89, AO77b, Arn80, BDG97, BX96, BL70b, Bat90, BL78, Ber91, BHK^{+}85, BW72a, BD99, Brä81b, Bre75a, Bre78, Bre80, BRS92, BF66b, BGBM89, CR94, Car72a, Cha98, CVV88, Chu76, Chu79, Cla76, CW77, Cor77, Cuy82, Cuy83b, DT71, DH97, Dör96, Dub70, Dum98, Dzi91, EA88, EP95, ESR82, Fis81, FZ93, Fre94, GG87, GT87, GK94, GP67, Gre82, Gre79, Gre75, Gre80, Hag87, HL64a, HL65, Hee68, Heg94, Hou71, Hou73, ID86, Jer85, KC93, KZ94, KP81, KP90, Kub84, KT65, LLS98, Leb70, LR89b, LRZ95, MRW70, MvdS76, Mei66, MS99, NS95a, NS83b, NSS86, O’C96, Osa96, Owe77, Paa71, PE72]. algorithm [Par65a, Par67, Pei85, PB87, Pfl98, Pir82, Pyl67, Qia88, Ren80, SD72, Sac72, Sch85a, Sch98, SS90, See82, She77, Shr91, Sid82, Slo80, Stu89, Uhl97, VB92b, Van79, Van81, VN93, Ves93, Wil97, Wol79, Wuy73, Wyn59d, Wyn64, YK88, xY95, Zah77, vG80, Müh78, Müh79]. algorithme [Bre75a, Car72a, Lau67]. Algorithmen [Sto61]. Algorithmes [Del80, El 84, Tao84, Bre71, El 82, RM72]. Algorithmic [vWMPK69, AKK^{+}68, BBG^{+}60, BBG^{+}63, Müh81, PS59, CR89]. Algorithmique [CR89]. Algorithms [BY66, BL80, BE73, BMRW68, Bre70, Cai91a, CMW63, DM85, FM68, Hig87, JS91, LV74a, Lan64a, Ley90, Mat93a, MM67, Met65a, Mil65a, MT89, AS83, AM94a, AY83, Ari98, Ban78, BS77, Bre71, BZS99, BD72, CL61, El 82, El 84, EKL91, FP94, GA87, GK85, GL99, GO95, GH95, HJR88a, HJ90, Kac84, KjS95, Kin92, KvdD77, LMR95, Mey75, PE97, PB81, Ral65, RM72, jS99, Sik82b, Sre71, Ste65, Ste99, Sto61, Sto77, Sun98, Tsa79, Wer78, Wit89, Wyn59c, Wyn60, vGS84, Del80, Met65b]. Algorithmus [Arn80, KS73, Wyn60, Rut63b, See82]. allgemeine [Wet60, vW77a]. allgemeinen [Döh82, Haa60, Mül68]. allgemeiner [Sch71a]. Allowing [SS79]. alloy [BB97]. alloys [NS91]. Almost [AY93, Gut89b, SS95]. along [TS91]. Alpha [GP85]. Alston [Ano83a, Ano86, CV94, JHW77, H. 73, Wil78]. Alternanten [Pit75]. alternants [Bri64, Pit75]. Alternating [Alb62, BS85, CW79a, Dou61, Dou62a, Dou62b, GM66a, Lee61, Bad89, BS97a, Bia91, DP63, Gun64a, Had71, Mat93a, Mat93b, Mil65b, MF64, Pea62, Weg89]. alternating-direction [Gun64a]. Alternation [NS83a]. Alternative [MC84, CGT94, Lam80, Sán75]. alternierenden [Alb62]. always [SH95]. Alwin [Sau67]. among [Rob66, Wyn66]. analog [Bra66, Raz69]. analogs [Mil65b]. analogues [BH62, BD72, Sus64]. Analyse [BN90, FM92a, FM92b, Pau81b, Maî68]. Analysis [Ano83b, AL85, BO80, Bar91a, Ber91, BL97, Bou90, Bre70, CMQ82, CS86, CC95, DGL99, Edd93, EP94, EW92, GMP95, GHR97, HR89, Hig87, HT95, JS95, JO74, Maî68, Man83, O’C93, Ort63, PW93, Pit88, Sau78, Slo81, Wil60, YM86, Zou99, dBP77, Agu89, And71, AM88, AF89, BJ86, BC93, Bao96, BN90, BH92a, BL93, BJZ88, BP92, BL86b, BLR89, BS97c, Bri83, BHK94, Cam70, Car99, CL94, Chi91, CCK95, CE92, Cro78, DN89, DT71, DNvW84, DMJ86, Gau72, GT87, GR79, GMS95, GT82a, GB86a, GB86b, GS93b, Gut87, Hac92, HL97, Hai99, HR87, Han92, HNS95, Han80, Häu86, HUY95, HMSW99, iS91, Jac84, JT76, KS99, Lan66, LP78, Liu95, LA97, Man81a, May86, Mie79, Mil77, NV83, PE97, Pau81b, Ped96, PLCC92]. analysis [Per99, PZ99, Pit92, PS96b, Pot82, Rei82, Reu95, RKL89, Rok72, Sar86, Sau73, Sch85a, Sch81b, Sch88a, Sim98, Slo76, Sta97, Ste95, Stu85a, Stu85b, Sug97, gS93a, gS96b, gS98a, Tsa79, Tsu99, Vog83a, Woź78, fX96, Zah77, Zen97, Zha92b, tMS86, Ano84, FM92a, FM92b]. Analytic [BL74, BGBMN91, Häm66, Mor80, BLW74, BS68, Cha71, GS95a, IK77, KWW85, Kre71, Las92, LMPY85, MSW91, MSW92, Not93a, Pau77, Ste71, Wil97, Wri92, YK88]. analytical [Nie73b]. analyticity [SW99b]. analytische [Häm66]. analytischen [Nie73b]. analytischer [BS68, Kre71, Pau77]. andere [Wyn60]. aneinandergesetzte [Her74]. Anfangs [vFvH73]. Anfangs-Randwertprobleme [vFvH73]. Anfangsrandwertaufgaben [GK80]. Anfangswertaufgaben [Hof75, Mar75, Ans64, Ans67, Ans69].

Anfangswertprobleme [Sau63].

Anfangswertproblemen [Kre59, Mül82b]. Anfangswertproblems [AU61]. Angenäherte [Fil64]. angenäherten [Zur64]. angle [Lou86, Žen95]. Anisotropic [AL96, Hac89, HN92, NMC89, SSS99, Sie96, Ste93a]. Anmerkung [Jet84]. Anmerkungen [Smo90]. Anniversary [Ano83b]. Announcemen [Ano83a]. Announcement [Ano86, Ano98a, Ano98b]. Announcing [Ano83b, Ano84]. annular [CM70, PPSS97]. Anwendung [Com59, Mar65, Sch66b]. Anwendungen [Had69a, Mar75, Rei71a, Sch77a, Wit78c, Ale71, BH74, Hof70, Kos86, Kul68]. Anwendungsbereich [Sch70b]. any [Man78a]. Anzahl [Hüb82]. appearance [Chi99]. appearing [Bri64]. Appendix [Sch95a]. applicable [DPRW93]. Application [AABM98, BH61b, CR94, Hac89, Kat60, LV74a, Pau81a, Pau81b, RR90, Sch70b, TK64, Vog83b, AK99, BG76b, Bit70, BHK94, Cha73b, Deu74, Deu79, FG77, Fra76, Fra77, Huc98, Le 82, LA97, Mar75, McC75, MS85, MI82, Nie80, RRS89, SL86, SU89, Tee64, Tou92, Tsu99, Zen84a, Ale71, Bre72, Mei66, BH74]. Applications [BS77, BM87, CD90, Kul68, Laf75, LL90, Tan71, Van94, vGS84, BJ89, Bar93, BL76, BS97a, BSV68, BH71b, CW71, DML83, Dah85, Kos86, Krä89, Müh78, Müh79, NN87, Pir82, Pyl67, Rei71a, Rob71, Wit78c, Wit90, Žen95]. Applied [Kes79, And71, BS96, BEHH75, CL75, DP72, DN89, Fis61, Gol81, GP78, GLC89, Hac77b, Han97, Hoc79, LV89, NS95b, NS96, PV72, PR94, SSVH87, Sar86, Sar90, Sar96, Ste93a, Wes75]. appliquée [CL75, Kes79]. Approach [ADdR89, KM86, AO93, AG91, BB82b, CDV99, CKSS97, Chi99, DC99, DNvW84, FF97, Fro88b, Gün98, Har62, Krä89, Kul71, MO78, MD77, Mey75, Müh81, Rum96, Slo88, SD94, Sug97, VB92a]. approaches [CZ99]. approximant [Ben86, JT76].

Approximants [DD84b, DD84a, Cha84, Cuy82, Cuy83b, CV84b, DMGVO96, SV75, SV76a, SV78, VC94]. Approximate [AU61, Leh61, NR68, OP64, Zur64, Ama72, Ban87, BR81, BS92b, CH60, CW78, DPRW93, DM96, EN88, Ell79, EP95, GP79, HR94, Hop81, KL83, Leh67, LE81, Mag92, Mit78, Not98, Owe77, PLCC92, Per66, RS96, Sch79a, Sch87d, gS97, gS98b, Šve70, Tho75]. approximating [Dör90, Gue94, Mon92, PS92a, Wil97]. Approximation [Alb60, Arn82, Aum63, BL86a, BO90, BD82, BDT88, Ber82, Bla77a, BL81, BL67, BS68, Cha73b, Com85, Con84, Cro77, DN86, Ehr69, Ess60, Fag82, Fil64, Gai70, Gme87, Gri75b, Gri75c, Gri75a, GJ89, HSS96, Her70, Hof69, Hof70, IK89, KK74, KK76, Koc95, Kol78, Lam70, Lus96, Mit73, MRMW90, Mül68, Nat70, PPSS97, Rap77, Rap84, Sch70a, Sch95a, SS65, SB90, Sik61, Str72, Tay69a, VV84, Wag83a, Wer62, WB69, Wet70, Wol78a, Zur64, AGS98, AK99, AO76, AO77a, AO77b, AD73, BS92a, Bac70, BDS80, Ban92, BS92b, BE86, BE87, BS91, BC91, BL94, BB95, BW95, BB97, BBG98, BS99a, BS79, BL78, BMM87, Boe64, Bra71a, BRR80, BRR81a, BRR81b, Bri64, BD72, CMQ84, CC91, Car72b, CD73, Cha97]. approximation [Cha71, CG59, CL61, CL62, CFH68, CDR98, CD85, Cro80, DD98, Dob89, DGP93, Dun67, Dun68, Elb91, EK99, Fei87, FN98, Fet87, Fis81, Fis93, FR86, FGM87, FH90, Gar65, Gek74, GP79, Gol63, GR86, Gue94, GQ98, Gün94, GHR97, HS97, HS79, HS80, Hay65, Hin96, HT95, Hwa99, Ish79, JK91a, Jer70, JO80, Joh69, Kik81, Kiw84, Kre59, Kri64, Kut71, Lam80, Lau67, LPDT70, LMGR94, Law64, LMW84, Le 80b, Le 82, LR89a, LTT92, Li97, LMS77, MP95a, MPP99, Mei66, Mei67, Men74, Mey79, MM74, MP92, Mor94, MSW88, NS95a, NS81, Nie73a, Nit70, Nør75, NSS86, Opf82, PB87, Per99, Pit73, Pom68, Prö91, Qun91, Ral65, RSSV97, Rau86, RL81, Rei82, Reu88a, Reu88b, Rie91, Riv62, Rou87, RO70, Row71]. approximation [Saa86, Sac88, SA84, Sch88b, Sch78c, Seo95, Sik59, Soc88, Spa94, SW87, Ste66c, Sti60, SF86, TT87, Tor97, Tre81, Tsa77, VB92b, VŽ95, Ver87a, Ver87b, Ver91a, Vui90, WSB67, Wer86, Wid77, Wol79, Zha93, ZL94, vG82, Cha73b, LA88, LA89, Nat70, Sch71a, Sik59, Cro77]. Approximationen [Bla77b, Het75, Mer92, Bro65a, Cro76, Het74, MW60a, MW60b, PS66, Sti59]. Approximations [Cho90, Jou71a, Kal74, MW60a, MW60b, Moo68, PS66, Rob71, Sti59, AKK74, AF88a, AG88, And88, Aro63a, Aro63b, BBFS90, BL70a, BY66, Bey78, Bla77b, Bra69, Bro65a, CH71, CGH76, CS66, CS68, Dan66, Des61, DM85, DD74, FL70, FN72, Fun88, Gal89, Gau84, GM86, Hai82, Hel87, Het75, HMSW99, HMW86, Ish84a, IU82, Jam70, Jou71b, Kel65, KJ90, Kos85, LN99, Las92, MQ81, Mer92, Mül79a, Mül79b, Mül80, NS91, Nit62, NS83a, OW69, Por78, Pru75, Red88, RG88, Ric78, Ros75, SVN76, Sav93, SHH85, SS98, Spe71, Ste78a, Ste99, Str82, SP94, Vei60, Voi74, Won84, Yse83, Zou97, vI84, vV76, vPS96, Rob71, Cro76, For72, Het74]. Approximationsaufgaben [Sch78c, Wet70]. Approximationseigenschaften [NR68]. Approximationsproblemen [Aum63]. approximative [Str66]. Approximierenden [Nit62]. äquidistanten [Sch80a]. Äquivalenzsatzes [Ans66]. Arbeit [Ber60, Dir69, Els69, Had72a]. arbitrary [AR93, Bad89, Dol81, FQ99, Gme87, GH73, LG76a, LG76b, LL85, LZ80, Mar84, Mül79a, Mül79b, Mül80, Ves76a, Ves79a]. Arch [Zha92b, BD82, HC92, Red88, Zer94, Zha95]. arches [Ish81, Ish85]. area [ES92]. argument [YK88]. arguments [Sal64].

Arise [Nab92, FR86]. Arising [Cia90, BS88, CHS96, CGT94, EM91, KN84, OD95, Saa86, Šve70]. Arithmetic [LV74a, Lor92, Nic66b, Pic72, FM68, GH72, Ish85, Kup71, MM67, Met65a, Met65b, MH72]. arithmetic-geometric [MH72]. Arithmetik [Kup71, Nic66b]. Arithmétique [LV74a, Pic72]. Arnold [Dur91]. Arnold-Falk [Dur91]. Arnoldi [Sad93b, Sad93a, Sim98, SV93]. Arnoldi-Chebyshev [Sad93b]. Arnoldi-Faber [SV93]. arrays [Wim74]. ART [EKL91, Yse83]. artificial [HB97]. ascent [Buc62]. Aspect [Nic74]. aspects [Bro68, GS77, Rut69]. associated [Gut89a, RM72, Sab78, Skr98, SF86, SW80, Utr83]. Associative [Mai67]. associée [Mai67]. associèes [Sab78]. associés [RM72]. assumption [Kil75]. assumptions [Yam86b, Sch75d]. Asymptotic [AF89, BM96, Bar68, BLR86, BS66a, CS98, DVZ98, Din83, Gri75a, HL84, HW84, Hoc79, Hof67, Jac84, Kim96, Luk93, Mik70, Pin75, Sha84, ST90, Ste65, SP94, Wal91, Wan89, Zen97, BYI86, CG88a, DMR91, DR92, Fie87, For68b, Gug97, Ish84b, IU82, JW88, NW73, Wes75, Mer92, Ort86, Sch66a, Sch68b, Wol78a]. Asymptotically [Ess60, KV72, MM74, KW99, Köh95, Wil97, Zer94]. Asymptotics [Car92, ELR92, Pet93, VC94]. asymptotisch [Ess60]. Asymptotische [Gri75a, Fie87, Mer92, Sch66a, Sch68b, Wol78a]. asymptotischen [NW73]. Asymptotisches [Ort86]. asynchrones [El 82, El 84]. Asynchronous [FS94, BMPS99, El 82, EKN92, El 84]. attaining [VC92a]. attenuated [Nat79]. Attenuation [Gau72, Gut87, LTT92, Yan94]. attracting [KL90]. attraction [Bra77]. Attractive [HV85]. attractors [DH97]. Aufgaben [Boh67]. Auflösung [Alb61a]. Aufspaltungen [Tro76]. auftretenden [Wer72]. Augmented [ADdR89, BL97, Gov90, KP91, Sch81b, Sch81c]. Ausdehnung [Rut63a]. Ausgangsvektoren [Sch64a]. ausgewählte [Mar65]. Ausgleichsproblems [Sau84]. Ausgleichung [Läu61]. Ausnahmemengen [Red67]. Aussage [Hav84, Pit75]. Auswahlproblem [NO68]. Auswertungsverfahren [Rei75]. Author [Ano83c]. autoadjoints [Cha73b]. automatic [Chr70a, Chr70b, CC60, DD78, Gao89]. Automorphism [FN68]. Automorphismengruppe [FN68]. Autour [Rob76]. aux [AABM98, Cha73b, Del73, Kes79, Pau81a, Pau81b]. Available [Dek71]. avantage [Del73]. average [GNP89, Nov86, WW84]. averaging [Kir88, KN84]. avoids [AD64]. Award [Ano83a, Ano86, Ano98a, Ano98b, JHW77, H. 73, Wil78]. Ax=b [HW78, Owe77]. axiomatic [Kul71]. Axisymmetric [RBS82].

Backward [CI94, Hai99, Pal99, gS93a, gS99, dBP77, vdHtR81, Cas80b, Dei89, Gek78, Gek82, HL97, Lu97, LM98]. backwards [Höh82]. Bad [Tyr94]. Bairstow [FK87, IM94, Sal61, SZ61]. Balancing [ByS99, PR69, Läu61]. ball [Tao84]. Banach [Dör70, Ade69, Dör70, Gal89, LW82, Mey84b, Mil80, Pie72, Söd86, Yam87]. Banach-Räumen [Ade69, Dör70, Mey84b]. Band [MW65, MW67, ALM97, All70, All73, ER67, MRW70, Sch68c]. bang [GS77]. bang-bang-principle [GS77]. Bank [DR92]. Bank-Weiser [DR92]. Bareiss [Fre94]. barrier [CGT94, Jar94, JKR88]. Bartky [Bul69a]. Bartky-transformation [Bul69a]. Barycentric [Ber89b, Ber89c, Hen79]. Base [Laf75]. Based [Dag90, BS97b, CGR94, Car91, Cha99, DP88a, FHLM88, Gau61, Hac92, Jar94, Jer85, Jer89, Kan98, Kil75, Kir88, KP90, Mat90, Par88, Ser99b, Slo88, Smi65, Woh99, Wyn62, Yse90]. Basen [Hal92]. Bases [Sab78, CP94a, Hal92, Hwa99, Laf75, Sch95a, Ste98, Zie97, Sab78]. basic [BY93, Goë68, GK72, Uhl97]. basin [VG88]. Basis [BDY88, BB92, BHK^{+}85, IS96, JS91, LW99, Osw92, Pot89, Ste97]. Bateman [JS86]. Bauer [wEC86, DZ75, NS64, Rut69, Šve70, Zen68]. BDF [CGG90]. BDF-Formulas [CGG90]. Be [Bus68, BS98b]. beam [BD82, Cha97, Li90, Zha92b]. beams [LMGR94]. Behandlung [Bra60, Cro77, Kad67, Nat69b, Saw82, WS59, Web79, Wen68]. behavior [Din83, Fie87, Ish84b, Ort86]. behaviour [Nev77, Sto97]. beim [BS60, Bra77, Col61, Nit76]. beliebige [Dol81]. beliebigem [Mül79a, Mül79b, Mül80]. Bellman [FF94, Grü97, Hop86]. below [Mar87]. BEM [CFS97, GH92, Kim96, SS97, SS88]. Bemerkung [Els69]. Bemerkungen [ES73, BS60, Bul64, Dir69, For68a, För82, Had72a, Had72b, Red67, Rut60, Sch64b, Wal67, Zur63]. bending [Com85, Dur91, Joh73, PS96b]. bent [DN86]. BEPS [ELV91]. Berechnung [Döh82, Dre77, FN68, GL65, Ker66, Kil78, Koß81, Nic66a, PK64, Sau78, SS66, Sch61a, SG62a, SG62b, SEG62, Sch80b, Sto67, Wag76, Wag80, Woh63, WL79]. Berechnungskomplexität [Str73a]. Bergman [PW86]. Berichtigung [Met65b, Ber60, Hal60a]. Bernoulli [Mig76]. Bernstein [Ess60, Not90, Not93b, Sik59, Sik61]. Bernstein-Polynomen [Ess60]. Bernstein-Polynomen. [Sik61]. Berstein [Ess60]. beschränkte [Sac73]. Bessel [Car92, Elb91, Lun85]. Best [Ehr69, Ess60, KV72, Kri64, Lau67, Rei84a, Sau78, Sto62, AD73, BY66, Bla77b, Boe64, Bri64, Cro76, El 79, Gar65, Het74, Het75, JS70, Köh95, Med71, Men74, Mir71, NS83a, OW69, Owe77, Riv62, Sal71, SA84, Str82, Vei60, Wid77]. Bestapproximationsfehlers [Ehr69]. beste [Ess60, Het74, Het75]. bestehen [Wyn60]. bester [Cro76]. bestimmte [Bra70]. Bestimmung [Dir69, For68a, PS82, Haa60, Men74, Neu61, Rut66a]. Bettis [Del73]. Between [AR85, BS77, DKS90a, DKS90b, Gue86, Hod67, Lam70, Sto77, VV89, Wei92a]. Beweis [See82]. bezüglich [KS73, Nit70]. bi [DRW81]. bi-Lipschitz [DRW81]. biconjugate [BC93]. bidiagonal [LRZ95]. bidimensional [FMMS99]. bie [ST73, STR76]. Bifurcation [RR90, Ber82, Boh80, BRR81b, DC74, Gov90, Had80, Lan77, LO98, Rau86, Sey83, Web79, Mar70b, Sch77a]. Biharmonic [Bre75b, BL92, Bra66, Had71, Kes84, KS98, Lan87, LBF98, Par59, Pei85, PB87, Rua88, Woo68, Zha89]. biharmonique [Bre75b]. biharmonischen [Lan87]. bilinear [BN84, GMS95]. bilinear-constant [BN84]. bilipschitzienne [DRW81]. Binärmodellen [Com59]. Binary [Com59]. binomial [Fra62a, Fra62b, Fra63, Hen75]. Biorthogonal [Ros95, Zie97]. Birkhoff [Boj93, Müh81]. Birthday [Cia98, FV94]. Bisection [MBW67, GNP89, Nie95, Sik82a, ST90, Ber84, ESR82, Kea79, VI86, Wil62a]. Bivariate [LL90, Nie73a, NR95, AS90, Man72, Mel72, Xu94, vG82]. Björck [Hig87]. Björck-Pereyra [Hig87]. black [Ale82a, Bra92, Par88]. blending [CGH76, GH73]. blending-function [GH73]. Block [BMPS99, EP94, Elf80, HV64, KS99, MS98, Pao95, CR94, Els84b, EM91, Guo99, Jia98, LP95, LA97, Mag92, Man90, Ros64, Sad93b, Swe91, Sad93a]. block-approximate [Mag92]. Block-Arnoldi [Sad93a]. Block-iterative [Elf80]. block-scaling [Els84b]. blocking [Wen68]. bodies [Heb88, LV84]. body [Mar84]. Bogoliubov [Wet60, Wet60].

Boltzmann [DN89]. Boolean [KR82]. booleschen [KR82]. Bordered [Gov90]. bordering [GW95b]. born [GB71, GB81]. borné [LA88]. bornes [Cha79]. both [BS98b]. boule [Tao84]. Bound [Ber59b, Ber60, Due59, Leh60, Mit81, Uhl62, BB95, CNV88, MT89, Rei65, Sto64a, VC92a, Vit75, Wei95, You71, tR91]. Bound-constrained [Mit81]. boundaries [EJSS97, Fag82, Hin96, MSF91, Noc88, Tho71, VG88, vdH72]. Boundary [AM94b, BH92a, Ber59b, Cho90, CS59, KS71, Leh60, Leh61, MP96b, Pit79, Sch94b, SB90, Tre62, Wer69, vFvH73, Ain98, AM78, Alt80, AS97b, Asc85, AM88, AHS84, AKS88, BO80, BV84, BS89, BL92, Bak82, BL86a, BYI86, BE86, Ber73, BH61b, BH65, Ber59a, Bey78, Bla72, BS96, BF70, Bos71, Bra66, BCS96, Bro85, CDV99, CH71, CKL98, CS90, CS92, CK82, CŠ79, CM62, CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, Com85, CS98, Cry73, DDM92, DP72, DPR76, DH80, DD73a, DD73b, DD74, Dut99, ES90, EG95, EJSS97, Ess77, EN80, Fae98, Fag81b, FN98, För67, Fra76, Fra77, GS74, Gek74, GK80, GeN80b, GOS99, Gri70a, GKR86, GN66, GS98c, Gus73, HN89, HHM76, HR87, HR95, Han91a, Han94, HB97]. boundary [HS83, Heb88, Hei89a, Heu98, HMS99, Heu01, Hof67, Ish84a, IU82, IJ87, Jam70, Jer70, KM67, Kea68, KV85, Kos85, KS83, Kre90, LSW85, Lan87, LN99, Lau98, Ley86, Li86, Li87, LS92, Lub94, LR73, MW99, Mey78, Mey84a, Moc74, Nat75a, NY83, Nit69b, NW76, OA75, Osb67, Osb68, Pao87, PLCC92, Per68, PS75, PPV69, Pet70, Rat93, Red73, RS76, Ren79, Reu88a, Run70, RW88, RS72, Rus74, Sán75, SK96, Şch73b, Sch85b, SS88, SSS99, Sch66b, ST73, Sch75c, STR76, Sch84, Sch71b, Sch72b, SW92, Ser99b, Sha68, SW98, Spa93, Spr76, Sta77a, Ste68, SO86, Tro76, TIK67, TV98, Ura67, VŽ95, VV84, Ver87a, Ver91a, VH93, Vol86, Vos75, Wag83a, Wen68, WhY88, Wit78b, Wit78c, Won92a, Won92b, Wri94, XS92, Zlá69, Zou97]. boundary [vPS96]. boundary- [BH61b, Gri70a]. boundary-flux [PLCC92]. Boundary-Value [Cho90, BH65, Jam70, TIK67, Wen68]. boundary/eigenvalues [Nat75a]. bounded [BM84, CD73, HP92, LB97, LA88, MS90, Mil67, Nev90b, Wil97, YK88]. boundedness [DML83, Spi86]. Bounds [Alb61a, Alb62, Boh69, BS60, BS63, BS70, Bra61, BH71b, Des63, DH93b, GI88, Hae63, Häm66, HR94, Hen62, Kna71, Leh61, Loi67, OP64, OD95, Söd86, Ste66a, Bab71, BJ85b, BJ89, BS92b, BM96, BG76b, BL94, BG95, Bjö88, Bla72, Bod68, BF99, BS66a, BEN93, CR89, Cha79, CK76, Coo71, CW78, DP72, Die96, Eng77, Fie78, För67, For72, Gar65, GF73, Gra68, GK72, Gün94, Gut72, HS79, Har91, IK77, Kam71, Kin74, Kos86, Küp76b, LW82, Leh67, Lin76, Lyn63, MV69, McA68, Mei67, MVW79, NV80b, NW76, Not93a, O’C96, PV72, PP80, Ric74b, Rom85, RO70, Sch81a, Sch66a, Sch83, Sch61d, Sch61e, Sch75c, Sch84, Sha68, Sha84, Spe78, Spe79, Spi76, Spr76, Spr80, Ste66b, Ste63, Sti88, Sto68, Str73b]. bounds [gS83, gS97, gS98b, SF86, SW80, Wid77, Yam80, Yam82, Yam86a, Yam86b, Zen72, Bit68, Bit71, Nat69a, Nat70, Nat71, Tre62]. Boussinesq [OSS90]. box [JS91, RS91]. Bracketing [CS59]. Branch [Sey83, Sey79b, Sey79a]. Branches [BRR80, Spr80]. Braß [FP90]. breakdown [BRS92]. breakdown-free [BRS92]. Broken [MW60b, Wer62]. Brouwer [SD67, SD67]. Brown [Jel79, Fro88a, JK78]. Broyden [Sto84]. Bruijn [CNV88, tR91]. Bruijn-Newman [CNV88, tR91]. buckling [Ish79]. Burgers [MQ81].

Cahn [BB95, CE92, EFM89]. calcul [Gas71a, Par73, Pau81a, Pau81b, Tao84]. calculate [Sto83]. Calculating [RvdS90, Stu89, Wag76, WL79, Wag80, Wag83b, dHH87]. Calculation [MBW67, Cul86, GL65, GK83, JS70, KE82, Lod70, MW67, Mor78, Mor99, Sch61a, SG62b, SEG62, Wil62a, Wil62b, Woh63, AB87, Ber84, BG71, Bul65a, Bul65b, Bul67, Bul69b, Cle66, Coh60, Cry67, Döh82, HvdR63, Kil78, Lyn68, Mid75, MH72, PR69, Par73, Pau81a, Pau81b, Gas71a]. calculations [CdM70, DB77, FDD92, PLCC92, TIK67]. Calculs [Rom85, CR89]. calculus [Lub88a, Lub88b]. Can [Bus68, WW84]. cannot [BS98b]. Canonical [CS94, MP92]. capacitance [Dry82, Dry84, Ern96, Shi78]. capillary [SDK91]. Car [BS98a, BS99b]. caractérisation [Con80, Lau67]. Cardinal [Rei89b, RS91, Ber89b, Ber89c, Gün94, JS91]. Carlo [CdM70, CP70, Gir89, Han64, Ksc68, Oga89, TK64, TIK67]. Carlson [Tod88]. carreau [BN90, BN90, BL93]. Cartesian [Ber72]. cascade [Fei87, GL99, LLS98]. cascadic [BD96, BD99]. Case [MT90, AS87, AF89, BNZ72b, BNZ72a, CT93, CC95, FMMS99, Fro88a, Gar76, Han80, Kio79, Las92, MS78, Sid82, gS95b, Tay69a, dFSS97]. Cases [Mül68, Rut63b, Sch75b]. Cauchy [Cha92, CR74, CM89, DP85, DD99, Die96, EP75, EP76, Gol62, Huc98, Hun72, Hun73, Lyn85, MRS95, Mon84, Mon87, PE72, Sav93, Tom84, Voi74, Wes75]. Cauchy-type [Huc98]. cavitational [Agu89]. CELF [SS84]. cell [ELV91, Gir76]. cell-centered [ELV91]. centered [ELV91]. Centers [LH66]. central [LT98b, Str73b]. centres [LH66]. centroid [GG71]. Certain [Gas65, HB59, HB60, Lig90, Sto62, All73, Ans64, Aro63a, Aro63b, Ber59a, Bra70, BF66b, DP85, Dör90, Ein71, EP75, EP76, EV97, För67, Gal89, Gri67, Hac77b, Hal60a, Hal60b, Hel68, Kar68, Lev89, Mis84, Mon84, Mon87, Müh85, Nic75, PE72, Pon66, Riv62, RW65, Run70, SV76b, SV77, Sch74a, Sei63, gS93a, Wil87, Wyn59c, Wyn60, Wyn63]. certains [Gas65]. Chains [O’C93, SZ91b, Bar93, Mei97, Ste93b]. champ [Tho90]. change [Chr70b, Chr70a]. Changes [SS79]. chaotic [VG88]. Characteristic [BH60a, HB59, Law64, SS66, Sch61a, SEG62, AR68, Ans63, BH61a, Hou68, Ste61, Sto67, gS93a, Wag76, WL79, Wag80, dO70, Ber91]. Characteristics [Sau63, Bro85, BEHH75, DH80, DHP99b, TS91]. Characterization [CD90, RS90, Tan74, Con80, Hai82, HS90, Man72, Sto64a, Utr85, WW68]. Characterizations [Gri67, Mat80, Hai86, Mes95]. Characterizing [Lau67]. Characters [Com59, Dix67]. Charaktere [Com59]. Charakteristikenverfahren [Ans63]. charakteristischen [SS66, Sch61a, SEG62, Sto67, Wag76, WL79, Wag80]. charge [Gün98]. charge-oriented [Gün98]. Chebyshev [Bac70, BL70a, Bla77a, Boe64, Bra71a, Bra82, Bri64, Bro65a, BL67, BS68, CR94, CGR94, CL62, CS66, CS68, CFH68, Cro76, Cro77, Des61, Dör70, Dun67, Dun68, Fil64, För82, FR86, GM77, GO94a, GV61a, GV61b, GO88, GeN80a, yGL93, Hof69, Hof70, Lig90, MQ81, MW60a, MW60b, Mül68, Ném65, Nit62, Ral65, RO70, Row71, Sad93b, Sal75, SK73, Sti59, TT87, Tre81, Ura67, Vei60, VHS90, Wer62, WSB67, Woź77b, Zie85, Zur64]. Chebyshev-type [För82, GM77]. Chebyshevian [Lyc72]. checkerboard [Loi68]. choice [DB77, Dia80, nJyH99, Kal98, LS70]. choix [Dia80]. Cholesky [EP94, Kiw94]. choosing [Luk93]. Chord [Boh80, DK83]. Christoffel [HR79]. Cimmino [ID86]. circle [CGR90, MRS95]. circuits [Gün98]. circulant [Hem98]. Circular [GH72, BDT88, Ish81, Ish85, VC92a, Zha95]. cisaillement [DN86]. clamped [Kut71]. Class [ABS84, Cia90, Fre90, Kin73, Maî68, Nab92, Wei64, Ari98, AL86a, BWY90, BH71a, Bor94, BH71b, CR93, CK82, CP94b, Coh60, Cro76, DB77, FR86, GLL91, He94, Jar94, Jeg68, Kem72, Kiw84, Kle65, LLS95, Mis84, Pet93, Pom71, Reu88a, Rit80, SP72, SU89, Slo76, Sta81, Sto84, Ves79a, Ves79b, Wah75b, WA72, Xu94, Yse97a]. classe [Maî68]. Classes [Fel61, Gri67, SG78, SWB91]. classical [vK66a, LRS91, MPP99]. classifying [Rya67]. Clenshaw [Bra73, Fil64, Bra73, ES88, Fil64, Imh63, RJ72, SS78, SS80, Smi65]. close [HH76, Ye95]. closed [Bey87, PS92b]. Clough [BDT88]. Clough-Johnson [BDT88]. clustering [HN89]. Co [BW93]. Co-volume [BW93]. coarse [CSZ96, Sar97]. coarsening [BX96]. coefficient [BL86a, KP91, SW99a]. Coefficients [Bra65, OP64, Uhl92, Ban87, DSW96, Eis67a, GM68, Hen75, Jov89, KW96, Lam70, Las92, Lin93, Mar87, PR86, Sar97, Sch82, SS65, Sto83, Žen90, Str73a]. coerciveness [Gri70b]. Coherence [WB69, HS91]. Collection [GL65]. collinear [Ari98]. collision [DLD94]. Collocation [CP76, Jay93, Lie90, Oli80, Rus74, AW85, BS95, Bia91, BD97, BN81, Cha92, CM92, Egg84, Egg86, EG95, EJSS97, Fer97, GS95a, Han88, Hei89a, JY93, Kad67, Kau89, KP95, KKL97, Lau98, Ley86, LBF98, MP96b, NW81, RS76, RF94, RS72, SV90, Sar88, Sar89, Sch90, Sch87a, Sch91, Slo88, Sun98, vV76]. column [MO68, Ste84]. columns [PW93]. combination [PZ99, Wer68]. Combinations [Bau65b]. Combined [CMQ84, NV96, BD81, CMQ82, Gir76, Li86, Li87, Tau76, VŽ95]. combustion [Edd93]. Comment [Dir69]. Comments [HS69, Pyt02]. committee [PS59]. common [Akr88, Gün75]. commutators [Ise84]. Compact [BS96, AS87, Car72b, Fri70, LPDT70, Pie70, Rap77, Yam68]. Compactly [DGM88, RS95]. Compactness [FŽ88]. Companion [TT94, Ste71]. comparative [CdM70]. comparatively [Rai69]. Comparing [DLO^{+}77]. Comparison [Aum63, BH90, Dic87, Fro88a, Gre79, GLC89, Gro89, MS90, SS62, Bou90, BEN93, Kle66, MP96a, MN85, RKL89, Son93, Str84, Tou92, Wyn62, Zen72]. Comparisons [CV84a, Els89]. Compatibility [OP64]. compatible [Joh77, Man74]. compiles [Kel72]. complement [Bre99, KW99, Man90, WKW97]. complementarity [ACP99, Kan98, KZ94, Sch87b, Smo90, SW99b]. complementary [Dan70]. Complete [GL65, HS86, Ker66, Con80, Kal74, MH72, Mor78, Con80]. Completion [EHM94]. Completions [Häm65, Cam93]. Complex [Bar65, BMPW66, Fre90, MW68c, Mue66, PW70, AJS90, AL76, Cha84, Ebe70, FR88, Gar76, GS96a, MS72, MS73, MH72, Mor78, Mor99, Opf82, Saa86, Shr91, Tho90, VW79, Zen84a]. complexe [Tho90]. complexes [Cha84]. complexité [CR89, Rom85]. complexity [CR89, Jer89, Rom85, Str73a]. complicated [HS97]. component [BB97, Bou90]. Componentwise [Ale86]. Composed [Mai67]. composée [Mai67]. Composite [Bre85, HS97, Sta77a, Alb78, BC93, Rya67, Sta80]. composition [Müh85]. compound [BNvW89, Dic87, För93]. compressed [NMC89]. compressible [LV84, Yse97b, Yse99a]. compressibles [LV84]. Computable [Sti88]. Computation [BH71a, BGBMN91, DV99, FN68, GK90, GM94, Ker66, Nic66a, Nod71, PK64, SS66, SG62a, Wil60, BP92, CD79, DH97, Dix67, DGL99, Eck76, Fro88b, GK72, Had80, HM86, JO72, Kea79, Koß81, Mor80, NMC89, Ren80, Rhe88, RP85, Sch80b, Sey79b, Sey79a, Ste99, Sto67, Str73b, Sty81, Tem83, VD91, Web81, vD73, Ale71]. Computational [BD73, BLLN95, CKP95, RR90, Rut69, Lam80, LMS77, MP95b, Raz79, Str73a, Yam66, Sch61d]. Computationally [Dan67]. Computations [BW72b, BW72a, GW94, GW95a, Kul71, Pot89]. Compute [Bar91a, BHK^{+}85, EJ81]. computed [FH95, Sch79a, Yam80, Yam82]. Computer [LV74a, CC60]. computerized [Lou86, LR89c]. Computers [Kar61, Nic66b, Lin62, Nie89]. Computing [Can71, Car79, GZ95, MSW91, Ste75a, Ste82, Van85, Zha96, Akr88, BG95, BMX98, Bra77, CVV88, DM85, Dre77, EA88, For68a, Jen91, Kun74, Lev89, Li95, Lue79, MSW92, PS82, RW65, Sad93b, SKL88, Ste76b, Tao84, vG80]. concave [SS90]. concentration [Bea86b]. concept [Wil71]. concepts [MNS74]. Concerning [Har80, FP90, Hav84, Med71, SV76b, SV77]. concurrent [VL92]. condensation [JM87]. Condition [AR85, Dem87, EHM94, GI88, Geu82, Kat60, MP96a, SZ91a, vdS69, vdS70a, BH60b, Bre99, Buc64, Car72a, Dör90, Dur91, FN98, Gau73, He97, Kim96, KS83, Mar62, Mil67, OD95, Osb67, Osb68, RS86, Saa86, SSS86, Sau84, cS84, Sik84, Smi67, gS92, gS95b, Šve70, Tor91, Van82, VV84, Ver87a, Ver91a, Won93, Wyn59d, Žen95]. conditional [Pie70, Pie72, Uhl97]. Conditioned [Gau75b, Uhl92, BRZRS98, Deu74, Ruh70, Wil59a, Wil59b, Wil72]. Conditioning [LA97, Ric65, CGR90, Dör90]. Conditions [Bre72, MW60b, Sha82, Wet70, BS89, Ber73, CSV68a, CSV68c, DD73a, GW86, Hai81, HR87, HR95, HB97, Het74, Het75, JW88, KK76, MW99, Mey78, NW79, Rok72, Run70, Sau84, TV98, Wag83a, Wil64, Won92a, Won92b, Zou97, Vos75]. conduction [Die87, Raz69]. cone [Rob72]. cones [WW68]. Conference [Ano83b, Ano84]. Confluent [BE73, GP71, Gau62, Gau63, Gau78, Tem83, Bre75a]. confluente [Bre75a]. Conformal [Opf80, Sym69, Weg89, DHP99a, Ell79, Gai72, Gut81, HP81, HP83, KP90, Men74, Men88, Opf79, PW86, PS92a, PPSS97, Šve70, Sym66, Weg78, Weg84]. conforming [GP83, Sar97, Ver91b]. Congratulations [FV94]. congruential [EHG89, Mar70a, Nie85]. Conjecture [PC83]. Conjugate [Dan67, Fre90, Not93c, Ste73a, vdSvdV86, AS83, AL86b, Axe87, AC94, FR86, Gin63, Gre79, NBE97, PB87, Shi78, Slo80, Ste75b]. conjugates [Hen79]. connected [DHP99a, Ell79, HP83, Sym69, Vit75]. connecting [Mar62]. connection [FR86, Hai82, MNS74, Sto93]. connections [VV89]. connexion [Dij59]. Conquer [Cup81, GR90, KM86, BDG97, BG94]. Conservation [ESS92, CL96a, CM80, KNR95, LS95, LT98b, MO78, Roh98, dFSS97]. conserving [LG76a, LG76b, Mar84]. Consistency [PR94, Wat68, MNS74, Yse83]. Consistent [BV68, NP78, Nic74, SHH85, Bro68, Elf80, NF69]. consistently [Sar96, Tay69b]. constant [BN84, CNV88, JY93, Lin93, MS92, tR91]. Constants [Sik61, Die96, HM86, Rei84a]. Constrained [Bar91b, GvM91, SS79, Cro77, EA88, FH87, FL91, Krä89, McC75, Mit81, MT89, Rei97, Won84]. Constraint [CM95, Com85, Gan81, Het75, Kiw87, Utr85]. Constraints [Hof69, Hof70, BS92b, Dun88, Häm65, Mav77, Rou87]. Constructing [Sta77b, Ghi68, Opf79, Pyl67]. Construction [FM92b, Goë68, HP76a, Köh95, MS67, AL68, Cha84, CH89, Gal89, Ric74b, Sch71a, Ste86, Sty81, Sch73a]. Constructive [Hai82, BHH88, HJR88b, HS70, MS72]. contained [Fri70]. containing [LR98]. contiguous [Her74]. continuation [Bos71, DPR76, Deu79, Gov90, Rei82, VL92, VC92b, Wam94, Nie73b]. Continued [BW72b, Gut89a, Hay65, RDM62, Arn80, BJ85b, BJ89, BW72a, FJ72, Fie78, Fra62a, Fra62b, Fra63, Jac85, JW88, JTW89, NW84, RW85b, Sto83, TW80, Van81, Wyn59a, Wyn59b, Wyn63]. continuity [CDL96, DM88, Rei84a, Bla77b]. Continuous [BF90, Ess60, GW95b, Kau89, BL67, DP88a, DD74, Her70, NY83, PR86, SS65, Stu76, Tou92, Ver88]. contour [Wal89]. contours [CCCD93]. Contrôle [LV74a]. Contraction [Bre89, Bra81a, CM62, KS71]. Contractivity [Len89, Spi83, Ver84]. Contribution [Geu82]. Control [Kna85, LV74a, Arn82, Arn84, BS99b, Deu79, Deu83, GHR97, He97, Hoc81, HT95, KR79, KS89, OO88, Ren85, RRS89, Rou87, Tau76, Zhe89]. controlled [Neu61]. controls [HT95]. Convection [Cai91a, BBFS90, BB92, EK99, EGH99, GS93b, Hem98, JMT97, KR97, Reu95, SO89, TS91, Ver98a]. Convection-Diffusion [Cai91a, BBFS90, BB92, EK99, GS93b, Hem98, JMT97, KR97, Reu95, TS91, Ver98a]. Convenient [Dan67]. Convergence [Alb62, Ans73, AR85, Bat90, Bre78, BD68, Bus77, CGK92, CH60, CGH93, CS76, Dan67, DD95, DG98, DMJ86, DH93b, EM91, EKL91, EKN92, EGH99, GG87, GK94, GP79, GO88, GL99, Gri75c, HL98, ID86, Jac79, JW88, Kik81, KNR95, Ksc68, LRS91, Las92, LLS98, LZ80, Lin75, LS95, MT96, Mat93b, Mis84, Mon84, MH72, NC93, NN97, NV80a, Not93c, Par65a, RS90, Ral66, Red88, Reu88a, Reu88b, RKL89, SS84, SSVH87, Sch85a, Sch93, Sch81b, Sch95b, Spi66, Ste94b, Sül88, Tho71, Van94, VHS90, Wat68, Weg84, Ypm84, Zou97, vdSvdV86, vPS90, AGS87, Ale76, AV76, Ale82a, Ale82b, Ans64, AY93, AW85, AL86b, BD77, BS77, BL70a, Ber72, Bra70, Bra69, BdP72, Bre72, BP66, CGPM98, CG88b, CL75, CK82, CZ98]. convergence [CC85, CDR98, CM89, CMO91, Cro78, Cuy83a, hDxY99, DD75, DK83, Dej66a, Dej66b, DKS90a, DKS90b, DGB80, DMGVO96, DP63, Eis67a, El 82, El 84, EP75, EP76, EGS94, Eng72, FB89, Fer89, FK87, FHL86a, FHL86b, Gab84, GT67, GT84, Gek74, Geo79, vG72, Gri80, GT82a, Gri70b, Gri70a, Gus85, HNS95, Han97, Har82, Har91, Häu86, Hav84, HMS99, Hun86, Jam70, Jer89, JO80, Joh73, Jov89, vK66a, vK66b, Kio79, Kle81b, KL83, Lau98, LMW84, LP95, LE81, Lor80, Lou79, Lub91, Lub92b, Lub89b, MNS74, Man72, Mar69, MS94, Mat90, May85, Mei97, Mer91, Mik70, MK80, Mül82b, NBE97, Nat75b, Nic75, Nit70, Opf82, OSS90, Ort86, PW86, Par67, Par62b, Pea62, PS84, PV72, Pow94, Ptá76, Pyt02, RG88, RH93, Rit80]. convergence [RS88, SSV76a, SSV76b, SVN76, SP72, Sar88, Sar96, Sch66a, Sch68b, SS88, Sch87c, Sch61c, Sch74a, Sch75d, Sco73, Ser99b, cS84, Shi78, SS78, Spe75, Sta81, Ste78a, Ste61, Ste75b, Sto77, Sto84, Stu76, SO87, TW80, Tro90, Uhl97, Utr85, Wag76, Wal91, WFS95, Wer78, Wil87, Wil62d, Wit90, Wol79, Yam87, xY95, tMS86, AV86, Ans67, Ans69, EH70, ES73, Hac92, Hei89a, Pal77, Wit77, vFvH73, Gri75b]. Convergent [Ale73, CL96b, NP78, NS91, Ber89a, BC78, Dol81, Gal89, HS90, Lei71, Man70, Nii84, PS82, Sch79b, SO89, SS95, Ves76a, Ves76b, VW79, Wyn62]. convergentes [Del80]. converges [Nie95]. Converging [Wyn59a, Wyn59b, Wyn63]. Convex [Att68, MS99, Ari98, Auz88, CG59, EM87, FP90, Gre75, Har78, Kiw87, Krä89, LPDT70, MN68, Owe77, Ped96, SS90, SH95, Sch71b, Ste93a, Stu89, WW68]. convex-concave [SS90]. convexe [Att68]. convexity [Cle92, SKL88, Zen68]. convexity-preserving [Cle92, SKL88]. Convolution [Lub88a, Lub88b, Roa71]. convolutions [Egg92, ST89]. cooling [Raz69]. coordinate [Sch74b]. coordinates [Sch95a]. Coordinex [PC83]. coques [Tro90]. corner [LdD93]. corners [BR72, BR88, Heb88, Kre90, Qun91, Ste93a, Vas88, ZG78]. corps [LV84]. correct [CW79b]. Corrected [BS95]. Correcting [Dan66, MO78]. Correction [AP85, AP86, Aro63a, Par67, Pic72, Auz87, Dou62b, DM96, Fra76, Fra77, FHL86a, FHL86b, Fre60, Hai78, Moo88, Oli80, Sch90, Ste78b, Ber60, Hal60a, Pic72]. corrections [BZ85, Böh81, CR80, Mar90, Per66, Per67, Per68]. corrector [FH87]. correlated [Blu61]. correlations [DP88b, DA71, EHG89]. Corrigendum [CS68]. cosine [Bul67, JY93]. Cotes [Ask72, Bra78, GR70b]. Coulomb [Tau76]. Countable [Red67]. counterexample [She77]. Counterexamples [JKR88, DNvW84]. counting [GS96a]. coupled [BW95, BS88, GH92, KS92, Mor62, Pao87, Roh98, vD73]. Coupling [QLV91, BCS96, CFS97, GS98c, de 95]. Courant [SSS86]. Courant-Friedrichs-Lewy [SSS86]. Covolume [HN92]. Cowell [SB69]. crimes [Žen89]. Criteria [All70, Ans64, Ari98, Eng72, För93]. Criterion [Nit68, Cro78, Mar69, Sch68a]. critical [GR86, MVW79]. cross [CW79b, Gir89, HS91, Luk93]. cross-validation [CW79b, Gir89, Luk93]. Crouzeix [Cha99]. crystalline [GK94]. Cubature [Gue86, LR98, BNvW89, CH86, CH89, El 79, Gue88, HP76a, Möl76, Sch78a, VC92a, VH93, Xu94, GG86]. cubatures [SW92]. cubes [GT78, NR96]. Cubic [Dag90, Cle92, Cul86, DP88a, DP82, Gme87, KP95, KK97, KKL97, OO88, RH93, Sak97, SS90]. Current [BL90, HT95]. Curtis [Bra73, Fil64, Imh63, RJ72, SS78, SS80, Smi65, Bra73]. curvature [CL96b, DD95, GK94, NV96]. curve [Chu83, EG94, GA87]. curve-tracing [Chu83]. curved [BE87, Cha97, EJSS97, GH73]. curves [AW85, Bey87, DM88, PS82, PS92b, Sak97, Sar89, SW99a, Sto93]. curvilinear [Sta77b]. cusps [RP85]. cycle [Reu88b, vG80]. cycles [HV85, vG82]. Cyclic [KS73, Zaw91, Dol81, ENR90, Fer89, HP94, vK66b, Kje61, LV89, PHP90, Rei89a, RH93, Sar86, Sar90, Tee64, Wil62d]. cyclically [EGS94, HV64]. cycling [AD64]. Cylindrical [vFvH73, BDT88, Pit92].

D [Ale82b, JBC92, CFS97, Wyn59d]. D. [Ale82b, Lin61]. DAE [Bar91b]. DAE’s [Lam91, ACR94, CM95, CG95]. Dai [Pyt02]. d’ajustement [Dia80, LA88, LA89]. d’Algorithmes [LV74a]. damped [HR89]. damping [Tau76]. Dans [LV74a, Att77, BK96, Des61, DN86, Dia80, DRW81, Gou77, Lau67, LH66, Néd82, SA84, Tho90]. DAP [MP85]. d’application [Bre72]. d’approximants [Cha84]. d’approximation [AG88, LA88, LA89, Tor97]. Darmstadt [Sau67]. Darstellung [Lin62, Mar73]. Data [Cho90, SS79, AB74, AY93, Blu61, Cas85, CEL98, CW79b, Cul86, DDM92, DM85, Dut99, Gir89, GH84, Hel87, HSY91, HdH85, KO98, Lou86, LR89c, Mil77, Ran84, SL86, SS90, ST91, Sco73, SW93, Utr85, VL92, Wah75a]. Davidson [Sad93a]. Davis [Mei86]. Davis-Kahan-Weinberger [Mei86]. DCR2 [BL78]. decay [KK97]. Decimal [Lod70]. decision [AD64]. Decomposable [Her74]. Decomposition [Bre70, BKP76, BN78, BGBMN91, EP94, GR70a, MW65, MPW65, Smi91, Aus71, BW99, BS97b, Bör89, ByS99, Bre99, CGK92, CKL98, CVV88, CMR86, Fun88, Gas63, Hac89, Hac92, Kim98a, Kim98b, Krä89, Lou86, MQ89, NN97, O’C96, PS92a, QLV91, SP72, SHH85, Ste82, Wri92, d’H93a, Aus71, SP72]. decomposition-dualization [Krä89]. Decompositions [CI94, FNP82, KS73, FH95, GW95b, Ros95, Van85, Wag97a, Wag97b]. découpe [RM72]. decoupled [NS95a]. Defect [Auz87, Moo88, Böh81, Fra76, Fra77, FHL86a, FHL86b, Hai78, Kna85, Mül79a, Mül79b, Mül80, Sch90, Ste78b]. defect-correction [Fra76, Fra77]. Defekt [Mül79a, Mül79b, Mül80]. Deferred [CR80, Per66, Per67, Per68]. deficient [Bra82, LSY87, VC92b]. Defined [Att68, OW69, PS82]. definierter [PS82]. définies [Att68]. Definite [AR85, EHM94, Kna71, MW65, MPW65, MPW66, Nic74, BE87, Huc98, Mat96, Mis99, gS83, Ves93]. definiteness [All70, DE94, FP66, Wet70]. definiter [Kna71]. Definitheitsbedingungen [Wet70]. Definition [dJ77]. deflating [gS96b]. Deflation [BG71]. deformations [LV84, LV84]. degenerate [BBG98, BW93, MP95a, Slo76]. degenerate-kernel [Slo76]. Degeneration [Des61]. Dégénérescence [Des61]. degré [LL85]. degree [AS90, CW79b, DV97, ELR92, Fri70, vG72, Gün75, HP76a, Kea79, LL85, OO98, Rei84a, Sar88, Sau97, Sik59, Sri94, Ste75a, Str66, Str67, Str68, Sty79, Sty81, VC92a]. degree-computation [Kea79]. degrees [Bea86b]. delay [BZ85, BJZ88, BMR96, CG88a, CŠ79, Gug97, GH99, Had80, iS91, IK89, Kot98, OP81, QYK99, Tor91, Ver92, Zen86, Zen97]. delay-differential [BJZ88]. d’éléments [Con84, Tro90]. demi [DRW81]. demi-espace [DRW81]. denominator [Lam70]. denominators [vI84]. Dense [HO90a, CGC95]. density [HT95, Lus96]. dependence [MVW79]. Dependent [BB90, Arn81, BO80, BS97b, CZ99, Dör96, Fag82, FN72, GS93b, Lan64b, Pal99, PB81, Sch99, Sch86]. depending [DL97, MS78, Wol78a]. d’équations [Del73, Tho90]. d’équilibre [LV84]. derivation [OO88, TV98]. Derivative [BD72, Hae63, BLLN95, Kio79, Kre71, Mat92]. Derivative-Free [Hae63, Kio79, Kre71]. derivatives [Brä81b, CMO91, FG73, GW86, GM94, Häm65, Kan73, KM91, LMW84, LM69, LR98, Mon87, Voi79, Eng77]. derived [SG78, Yam86a]. deriving [HJR88b]. d’erreur [AABM98, Cha79, FM92b]. Descent [Bol84, Rus84, Sch78c]. Design [Lod70, PS96b, Cas81b, GMP95, GM94, SD94]. detecting [BK87, Neg90]. Detection [GS96a, LV74b, Ren85, LV74b]. Deterioration [HC92]. determinant [Mar62]. Determinantal [Hom95, Wag80]. Determinantenmethode [Wag80]. Determination [Bra61, KO98, DA71, ESR82, GH72, KM98, Vas88, Vei60, Wei80, Gai72, Haa60, Men74, Neu61, Rut66a]. determinations [Kem68]. determined [Par62b]. Determining [Dir69, EG94, Gre75, HR79, VN93, Wer79]. deux [Cha84]. development [Mar73]. deviation [RO70]. device [CC95]. d’évolution [Cro80, RW85a]. d’extrapolation [Bre72]. Dezimalzahlen [Lod70]. DFT’s [Hek87, ST89]. d’Hermite [Gou77]. Diagonal [Por66, SW62, MP96a, Man90, Peñ98, Ros64]. diagonalising [Ves76b]. diagonalization [MP85]. diagonalizing [Ves79a]. Diagonally [FNP82, VN93]. diagrams [Had80]. died [GB71, GB81]. dielectric [BP92]. Difference [AKK74, AU61, Ans69, Jou71a, Kre63, Mir71, Rut63b, Sus64, Tör63, Uhl62, vFvH73, Akr92, Aro63a, Aro63b, Asc85, Auz87, Bey78, BS96, BH62, Bra66, Bra69, BEN93, Cam70, CK82, CT93, CL96b, Dej66a, Dej66b, Del73, Dob89, Dou61, Dou62b, Eis67a, FGM67, Föß89, Gai72, Gek74, GK80, GW95a, Gor70, GM66b, Gri70b, Gri70a, Gri85, Gun64b, Had69b, Hel68, Her90, HMW86, IJ87, JBC92, Jam70, JK91b, Jou71b, Jov89, KM67, Kel65, KKL97, Koß81, Kre59, Kut71, LMW84, Lee61, LRT64, May65, McA68, Mer67, Nat75a, Nii84, Nii85, OSS90, Pao85, Pao98, Par59, Par61, Par62a, Par62b, Rei65, ST73, STR76, Sei63, Spi71, Sta77a, Sta80, Ste61, Ste63, Str64, SO87, SO89, SS95, Sun68, Tau76, TW68, Tho71, Van79, Wid66]. difference [Yse83, YM86, ZG78, Zen84b, Sau63, See82]. difference-differential [Zen84b]. differences [Cry73, GJ89, Lar81, Loc80, Mey79, Müh79, Sal64, Str73b, Del73]. different [SP72]. différentes [SP72]. differentiability [Rok72]. Differentiable [Ehr69, GK64, Aus71, ASS88, BL67, BS68, For78, Nit62]. différentiables [Aus71]. Differential [AU61, CKP95, Kre63, Nat70, Nat71, Red67, Rut60, SS66, Sch61a, SG62a, SG62b, SEG62, Sto70, Tör63, Uhl62, Wal68, Wri92, AKK74, Alb60, AU62, Alb78, AD93, AM94b, And69, Ans63, Arn79, ASW93, Aro63a, Aro63b, Axe77, BD83, Ban87, Bel83, BZ85, BJZ88, BSV68, BMR96, BS66b, BEHH75, CG88a, CGPM98, Cam93, Cas78, Cas83, CP76, CS76, Cha73b, CL75, CT91, CŠ79, Chr70a, Chr70b, CG67, Coo71, CW78, Cry73, Del73, DD98, DHZ87, DE94, DE96, DLO^{+}77, DH80, DPRW93, Ein71, ESS92, Ess77, FG73, FP94, För67, FL91, Gar96, Gau61, GS74, GF73, GP65, GHR65, GeN80a, Gug97, GH99, Hac77b, Hai77, Hai79, Hai81, HL88, Han94, Han88, Hil83, HW87, HL99, Hor75, HI71, HL60, iS91, Ise84, Jac79]. differential [Jac82, Jac84, Jay93, KR79, Kel65, Kem72, Koc95, Koß81, Kot98, Kre59, KM98, Küp76a, Küp76b, Küp78, LS66, Lan77, Lee61, Leh67, Lin75, Liu95, Lub82, Lub89a, Lub91, Lub92b, Lyc72, Man78a, Mar75, Mat76, Mey70, MM74, Mül79a, Mül79b, Mül80, Nat69a, Nat75a, OP81, OW69, Par65b, PS75, Pfl98, QYK99, Raz79, Red77, RRS89, RW85a, Roc88, Ros75, SSVH87, Saw82, Sch87a, Sch95a, Sch61d, Sch61e, Sey79b, Sey83, Sha67, Squ69, Sta80, Sti88, Sto67, Tan92, Tho90, Tor91, Tro71, Tur78, Ura70, Voi74, Voi79, Wag76, WL79, Wag80, Wag83b, Web79, Wer68, WH86, Wet60, Zad76, Zen84b, Zen86, Zen97, Zha93]. differential-algebraic [AM94b, ASW93, DHZ87, FL91, Han88, Jay93, KM98, Lub89a, RRS89]. differential-equations [Alb60].

Differentialgleichung

[SS66, Sch61a, SEG62, Sto67, WL79, SG62a, Uhl62, Wag76, Wag80]. Differentialgleichungen [AU61, Arn79, Mar75, Mül79a, Mül79b, Mül80, Nat69a, Nat70, Red77, Rut60, Sch61d, Sch61e, Tro71, WH86, Ans63, Ess77, För67, GS74, Koß81, Kre59, Kre63, Leh67, Nat75a, Red77, Tör63, Wal68, Web79, Wet60]. Differentialgleichungssysteme [Saw82]. Differentialoperatoren [Küp78, Küp76a, Nat71]. Differentials [Leg66]. Differentialungleichungen [Red67, Sto70]. Differentiated [NR68]. Differentiation [LM66, Sch63, AB74, Cas80b, Cha61, DA86, Eng72, Gek78, Gek82, KW95, Kra63, Pal76, Pal77, Sal60, Sal63, Sch64b, Tsu72, Wei92b, vdHtR81, Sch64b]. Differentiations [Pal76]. Differentiations- [Pal76]. Differentiationsformeln [Pal77]. Differentiationsverfahren [Eng72]. différentiel [Hil83]. différentielle [Hai77, RW85a]. différentielles [CL75, Del73, Tho90]. differentiels [Cha73b, HW87, Leg66]. Differenzapproximation [Kre59]. Differenzen [Loc80, Rut63b, See82]. Differenzenapproximationen [Yse83]. DIfferenzengleichungen [ST73, STR76, Gri85, Gri70a]. Differenzenmethoden [Gai72]. Differenzenoperatoren [Gri70b]. Differenzenschemata [Gor70]. Differenzenverfahren [AU61, Ans69, Bey78, GK80, Nat75a, Uhl62, Koß81, Rei65, Sau63]. Differenzenverfahrens [vFvH73]. differenzierbare [BL67]. differenzierbaren [Nit62]. differenzierbarer [BS68, Ehr69, For78]. differenzierter [NR68]. Differenzmethoden [Kre63]. Differenzverfahren [Tör63, Tau76]. difficulties [MP95b]. diffraction [Bao96].

Diffusion [Cai91a, Jou71a, ASS88, BBFS90, BB92, BS91, BW95, CC95, EK99, EGH99, Gor70, GLC89, GS93b, HR95, Hem98, JMT97, Jou71b, KR97, Kre68, KP91, Lub92a, LO98, Nii90, Pao85, Pao98, Pir82, Ran84, Reu95, Soc88, SO89, SS95, TL91, TS91, Ver98a, Ver98b, YM86]. diffusion-convection [SO89]. diffusion-type [GLC89]. Diffusionsgleichungen [Gor70]. digit [Deá80]. digital [Gün98]. digitalen [Lin62]. digraph [vG80]. dilation [Mei86]. dimension [AS90, BL81, CP76, Fun88, GK94, Nak85, Seg99]. Dimensional [Jou71a, KS73, LB97, AF88a, AFW93, Ale86, BV84, BF70, Bou90, BRR80, BRR81a, BRR81b, CH71, CKSS97, CSV67, CC85, CH89, DP85, EP86, For68b, For85, FH90, GB86a, GB86b, GB86c, Gün75, yGL93, Hal60a, Hal60b, Han94, Han80, HS83, Hek87, Jam70, JM78, Jou71b, LSW85, LM70, MT78, Mit73, Mon84, NV80a, NR96, Oga89, Rau86, RK84, Rhe88, Rua85, Sac88, Seo95, Sha70, Shi79, Söd86, SV79, Str66, Ver84]. dimensionalen [KS73]. Dimensions [Smi91, BP85, Cha99, DSW96, GR70b, GC96, Han91a, Heu98, Heu01, HH99, KZ74, KNR95, MBL90, MMS92, Pow94, Zwa73]. DIN [Lod70, Lod70]. d’intégration [GG86]. d’interpolation [Che79, Gou77, LA88, LA89, SA84]. Diophantine [Gre80, HL64a, HL65, RL66]. diophantisches [RL66]. Direct [ABS84, BKP79, JO74, LRT64, Mer67, SZ91b, Tha90, Bjö88, LE81, SS88, Sun95]. direction [Bia91, Dou61, Dou62a, Dou62b, DP63, GM66a, Gun64a, Had71, Lee61, MF64, Pea62, Ste73a]. directions [For68b]. directly [HSY91]. Directory [Ano99]. Dirichlet [BS89, BE86, Bör89, BH62, Bra69, Cas85, Dry82, DRW81, GH92, HMW86, KS98, KS83, Mey79, SB90, Shi79, WZ95, Wil61, Won92a, Won92b, ZG78]. disc [BP98, Sun95]. discontinuities [CS76]. discontinuity [DJ69]. Discontinuous [BS92b, SW96, DSW96, FG73, Man78a, Sar97, Ste90b, Žen90]. discrepancy [Pla96, Whi77]. Discrete [AO76, AO77a, AO77b, AD74, Böh81, Cho90, CV70, Des61, DP82, FF94, Gau84, Hal77, Ple76, Sti59, WB69, Ain98, AGS98, ADK91, Auz88, BY66, BM87, Boh80, CRZ99, CZ99, Dob89, Fis81, GLM89, GLM91, Grü97, Hüb82, JO80, Kil75, Kos86, Lin93, LA88, Lor77, Mar90, NV80a, NY83, PS75, Pom68, Stu76, gS98b, SSBB88, TV98, VB92b, Wer62, Wil61, Wol79, For72, Gri75b, Gri75c, Gri75a]. discrete-ordinates [NV80a]. discrete-time [gS98b]. discrètes [Des61, LA88]. discretisation [Jeg68]. Discretization [Arn81, Bru78, Li90, MSF91, Mül82a, Akr92, AD93, DMJ86, Fae98, FR84, Gar96, GKR86, Hac77b, Hof67, Lin75, LS92, Lub94, LO98, Sch94b, Ste65, Ste78b, Web94, Yse89]. Discretizations [KP95, Wat68, Ess77, EN80, GO94b, JN81, JN82a, JN82b, OP88, Ste68, VW92, Ver91b, Woh99]. discretized [Krä89, Lub88a, Lub88b]. discs [DZ75, JS70, Zen84a]. disk [BK96, Tre81]. diskete [Wer62]. diskrete [For72, Kos86, Sti59, Gri75b, Gri75c, Gri75a]. diskreten [Hüb82]. diskreter [Lor77, NY83, WB69]. Diskretisierung [Jeg68]. Diskretisierungen [Ess77, EN80]. Diskretisierungsverfahren [GKR86]. Disks [MV68a, MV68b, Bau68, Med71, Med73, Por66, Ste67, Tod65, Var64]. Dismantlings [Kul68]. displacement [RS68]. disque [BK96]. dissection [CR89, GT87, Rom85]. dissections [CR89]. Distance [Dem87, He97, Mal99]. distances [Ric74b]. distinct [vK66a]. distributed [Arn82]. distribution [AL86a, CDV99, JTW89, Loi68, NW73, Sim78, VL92, Yam75, Yam76]. distributions [Gau84, GM86, LW99, Won93]. divergence [Axe77, FK87, Sar90, Vog83b, WH86]. divergent [Nin66]. Divergenzaussagen [WH86]. Divide [Cup81, GR90, KM86, BDG97, BG94]. Divide-and-Conquer [KM86, BG94]. Divided [Sal64, Lar81, Müh79, Loc80]. Dividierte [Loc80]. divisor [BZ82]. divisors [Akr88, Wil84]. do [EN88]. Domain [Bör90, Fun88, Kim98a, Smi91, d’H93a, BD98, BE87, Bör89, ByS99, Bre99, CGK92, CKL98, CGC95, DG86, Fis93, HB97, Kim98b, LB97, LA88, Lor80, MQ89, NN97, PS92a, QLV91, YK88, Küp76a, Küp78]. domaine [LA88]. Domains [Sar90, Ain98, Auz88, BR72, BG76b, Bou90, DG98, DZ71, Ell79, EJSS97, GH73, Gri80, GS98c, HS97, Hof67, HP83, KS98, Kre90, Pal99, Qun91, Sar88, Ste93a, Sym67, Sym69, YN93]. dominance [Peñ98]. Dominant [FNP82, Mat80, VN93, Yam66]. dominated [Mat80]. dont [BN90]. d’opérateurs [Cha79]. Doppelschichtpotentialen [Wen68]. d’ordre [LA89, SA84]. d’Orlicz [Rob71]. Double [Spe71, Leb70, Lus96, Rat93, Sug97, Wen68, vG82, Ish85]. doubly [HP83, Sym69]. doubly-connected [HP83, Sym69]. Douglas [Gun64b, MF64]. Downdating [EP94]. Drazin [SFR93]. drift [CC95]. drift-diffusion [CC95]. driven [NV96]. drop [Edd93]. Dual [FF97, GL65, Neu60, GS74, Kiw94, Kle66, SS90, SS98, Neu61, VV98]. dualen [GS74]. Dualitätssatz [Sto64b]. Dualitätssätzen [Bit68, Bit71]. Duality [Bit68, Bit71, CS88, Sto63, Meh67, Sto64b]. dualization [Krä89]. Dualmaschine [Neu60, Neu61]. Dualmaschinen [GL65]. duals [Chu76]. Duffing [vD73]. d’un [Cha73b, DRW81, FM92a, FM92b, Hil83, LV75, Maî68, Pau81b, Rob66]. d’Une [LV74b, CL71, Lau67, Mai67, Pic72, Rob66]. dünnen [Hal92]. duoplex [KT65]. duoplex-algorithm [KT65]. duplication [Car79]. durch [Ehr69, Küp78, Nat70, Nat77, Sch78b, WL79, Wyn60]. dynamic [BS97c, Hod67, Man83, Tur78]. dynamical [BK98, Che97, KjS95, KL90]. dynamics [Dör90, SFR93, VG88].

each [KK97]. Eberlein [Har82].

economical [Ste76a]. écoulements [BN90]. edge [Per99]. edges [Heb88, LRT93]. Editor [FV94]. Editorial [Cia98]. Eduard [BGHS79]. Effect [BO90, JN94, Ban92, BR88, NV96, SK96, VŽ95]. Effective [Had80, CD79, GP67]. effects [BS92a, Nit76]. effektiven [KR82]. Effektivität [Saw82]. efficiency [Hal60a, Hal60b, Oli71, Saw82]. Efficient [Ban78, BS79, FGS95, Fer97, MP85, Sto83, CVV88, Kea79, KR82, LRZ95, Mit80, Mit81, Sch81c, Ste99, Tsa77, Wam94, dHH87]. Eigenanalysis [EO99]. eigenfunctions [Küp76b, PV72, Sti88]. eigenpairs [Sad93b]. Eigenproblem [BF66a, BNS78, Cup81, EB68, MW68b, BRS71, Ebe70, RS68, Roa71, Ruh70, Ves76a, Wil72, dV73]. Eigenproblems [GR90, BDG97, Faß97, Jia98, MS95a]. eigenreduction [Ves93]. Eigenschaften [Eck68, Sch73a, Kle81a, Sto70]. eigenspaces [gS83]. Eigensystems [Mue66]. eigentuple [BC78, BG78]. eigentuple-eigenvector [BG78]. Eigenvalue [BO90, Bra61, CI94, Had68, Had69a, Hav77, Leh63, Leh66, Mor74, Nat70, Nod71, Ort63, SZ91a, AP85, AL86a, Ban92, BK81, BG94, CDR98, Cry67, Dei89, Döh82, Fis93, FG89, GK90, Geo79, Gri70a, Gri75b, Gri75c, Gri75a, Haa60, Ish81, Ish85, Kol78, LR89b, Lin76, Mat96, Mer91, Mer92, MSW88, Mül82a, Osb69, Paa71, PV72, Pon66, Qun91, Ran79, Ric74b, Ric74a, Rod73, Sad93a, Smi67, Spe75, Spr80, gS92, gS95b, gS99, SW80, Utr83, VŽ95, Wet60, Wol78a, CSV68b]. eigenvalue-problems [Gri75a]. eigenvalue/eigenvector [Pon66].

Eigenvalues [MBW67, BD73, Eck68, Kna71, KM86, MV68a, MV68b, NR68, Rut66a, Sch70a, SG62a, SG62b, SZ91a, Sto70, gS91, AP86, BG95, BMX98, Ber84, DL97, DZ71, Döh82, Enn95, ESR82, Haa60, HUY95, HVW70, Kel65, Kel72, vK66a, Kut71, Lan64b, Lin76, Mal99, MO64, MV80, Nat69a, Nat75a, PR69, Par73, Pól68, Pru75, RW65, Sch79a, Shr91, Spe78, Spe79, Sti88, Tod65, VD91, Var64, VW79, Wag83b, Wei92b, Wei63, Wil62a, Wol78a, Yam68, Yam80, Yam82, Ye95, Zen72, Žit89, Mar90]. Eigenvector [PW84, BC78, BG78, Pon66, SW80]. eigenvector-eigentuple [BC78]. Eigenvectors [BD73, MW67, PW70, CH60, KL83, PR69, Shr91, Ste69a, Wil62b, Wol78a, Yam80, Yam82]. Eigenvektoren [Wol78a]. Eigenwertaufgaben [Lin76, Nat75a, Ric74b, Wet60, Had68, Had69a, Mer91, Mer92, Wol78a]. Eigenwertbestimmung [Bra61]. Eigenwerte [Bra71b, Döh82, Haa60, Kna71, Lin76, Nat69a, Nat70, Rut66a, Sch70a, SG62a, SG62b, Wol78a]. Eigenwerteingrenzungen [Leh66]. Eigenwerteinschliessungen [Leh63]. Eigenwerteproblemen [Gri75a]. Eigenwertnäherungen [Sch79a]. Eigenwertproblem [Hav77]. Eigenwertproblemen [Gri75b, Gri75c]. Eigenwertproblems [Gri70a, Döh82, Haa60]. eighth [Cur70]. Eindeutigkeit [Bro65a, BL67, Mül68, Pet70]. einem [Smo90]. Einfach [Wen68]. Einfach- [Wen68]. einfachen [Zhe89]. Einfluß [Nit76]. Eingabelung [Mar70b]. eingeschränktem [Zet70]. einige [Kul68, Eck68, Gri85, Had69a]. einiger [Cro77, Fie87, Saw82, Sch70b, Sik61]. Eins [Kle81a, Kle81b]. Einschließen [CS59]. einschließende [GKR86]. Einschließung [Hof75, Sch66a, Sch68b, Spr76, Sch78b].

Einschließungen [Nat69a].

Einschließungsaussagen [Küp78]. Einschließungsenaussagen [Küp76a]. Einschließungssaätze [Wer69]. Einschließungssatz [Wet60]. Einzelschrittverfahren [Due59]. Einzugsbereiche [Bra77]. Elastic [Cia90, AT79, AR93, DG86, Han92, NMC89, RSSV97, RG88, Vas88, d’H93a]. elastic-plastic [Han92, RG88]. elasticity [AS97b, ADG84, AF88b, BS92a, BD82, CFS97, CD98, FF97, FHLM88, Han94, JM78, Le 82, Mor89, Neg90, Rei91, Ste86, Ste88, SS96]. elastodynamics [Mak92]. elastoplasticity [Car99]. elastostatics [BCS96]. elastoviscoplasticity [BL86b, BLR89]. electrochemistry [HT95]. electromagnetic [PR93]. electromagnetics [HS83]. electromagnetism [de 95]. Electronic [Neu60, Neu61]. elektronischen [Neu60, Neu61]. Element [BO90, Bre75b, SB90, Yse86b, AKP95, AF88a, AFW93, AC92, AO93, AK99, AS97b, AP86, ADG84, AR93, ASW80, AHS84, ALS90, Bab71, BR72, Bab73, BO80, BD81, BV84, BS89, BS92a, Ban92, Ban78, BBFS90, BYI86, BS92b, BH92a, BG76b, BE86, BE87, BS91, BC91, BL93, BL94, BB95, BW95, BB97, BBG98, BS99a, Ben99, BP79, Ber89a, BP92, BB82a, BB82b, BR88, BL97, BHR77, BHR78, Bri83, Cai91b, CMQ82, CMQ84, CPT98, CKL98, Cas85, Cha98, Cha99, CL94, CC95, Che97, CZ98, CEL98, Cho89, CW71, CZ99, CC85, Con84, CD85, CS98, DD95, DG98, Dob81, DDW75, DG86, DPRW93, Dry84, DGP93, Dur91, Ern96, EJ81, EW92, Fei87, FŽ87, FŽ88, FN98, FN72, FGN81, For85, FM92a, FM92b]. element [FHLM88, GN87, Gas96, Gir76, GR79, GP79, Gol81, GH73, GLC89, GQ98, GB86a, GB86b, GB86c, GP83, Guo88, GS93b, GC96, GS98c, HC92, HN89, Han82, Han91a, Han92, HS83, HH76, HL98, Heb88, Heu98, HMS99, Heu01, HT95, Ish79, Ish81, Ish84a, Ish84b, Ish85, Jen91, JMT97, Joh73, Joh76, JM78, KJ90, KV85, Kik81, Kil75, KN84, LN99, Le 80b, LR89a, Les73, LZ80, Li90, Liu96, Lou79, LM98, Lub92a, MP96a, Mak92, Man81a, Man81b, Man83, Man84, MP95a, MP96b, Mel99, Mil85, Mis99, Mit80, Mit81, Miy76, Moc76, MG85, Mon92, NS95a, NS81, Nii90, Noc89, NW76, O’R84, OD95, OP96, Pav94a, Pav94b, PCV96, PB87, Per99, Pit79, Pit88, Pit92, Ran79, Ran84, RSSV97, Red88, RG88, Rei82, Ros75, Rua85, Rua88]. element [RW88, SK96, Sch94b, SW92, Sco73, Seg99, cS84, SJX97, Spa93, Ste86, SS96, Str72, SO86, SF86, Tor97, TS91, Tro90, Tsu99, VŽ95, VW92, Ver87a, Ver87b, Ver91a, Vog83a, Vog83b, Wan89, WhY88, Woh99, Won92a, Won92b, Yse86a, Yse90, Žen89, Žen90, Žen95, Zer94, Zha89, Zha92b, Zha95, Zlá68, Zlá69, Zlá70, de 95, FM92a, FM92b]. element-capacitance [Dry84]. Elementare [Kan73]. elementarsymmetrischen [Str73a]. elementary [Bau65a, RS86, Spe71, Str73a, Ves79b, Wil84, Kan73]. Elemente [Sch78b, Nat75b]. Elementen [Wei80]. Elements [MT90, Ste90a, AABM98, AM94a, AT79, AAG79, Arn81, Att77, BKP79, BD98, Ber73, BD82, BDT88, BLR86, BN84, Bra94, Bre96, ByS99, BDM85, BDDF87, BCS96, Cha97, CL75, Cha79, Dör90, Dur90, DMR91, FHLM88, HS97, HJR88b, HH99, Höh82, Koh87, Koh88, LLS95, LPDT70, LL85, Lin62, Man74, Man81b, MQ89, Mor89, Néd80, Néd82, Néd86, Osw92, PS96b, Prö91, QLV91, Sar97, Sch78b, Ste88, Won93, ŽV95, Néd82, AABM98, AAG79, Att77, Bre75b, CL75, Cha79, LL85, Tor97, Nat75b, Wei80]. eleven [Cur70, HP76a]. Elimination [Bau65b, Cry68, DK77, Sau78, Str69, vV77, Bus71, Huc98, Läu61, Sti60, Stu85a, Stu85b, Wri94, Läu61]. ellipsoidal [GK72]. ellipsoids [KP81]. Elliptic [BL90, Smi91, TW68, Woh63, Ain98, AK99, AW85, Auz87, Axe77, BS99a, BD96, BDM85, BDDF87, BS85, Bul65a, Bul65b, Bul69a, Bul69b, CHL85, Car79, Cha73b, CL75, Cha99, CZ98, CP94b, Cho89, CC85, CM92, Dou61, Dou62b, Dry84, DSW96, ELV91, FGM67, FŽ87, FŽ88, FG77, FGN81, Geo79, GP65, Gun64b, Guo88, Guo99, Hac79, Had69b, Han97, Hei89a, HvdR63, HMW86, Hwa99, JBC92, JN94, KM67, Kea68, KJ90, KS92, KW96, KP95, KKL97, Koc95, Kor94, Kor96, Kos85, LN99, LMW84, LA97, Lub92a, MP95a, MT78, Mat93a, Mat93b, McA68, Mid75, Mil85, MF64, Moc75, MH72, Mor78, Mor99, Noc88, NW76, Pao95, Par61, Par65b, PCV96, Pfl98, Reu88a, Ros75, SP72, Sar97, Sch95a, Sch91, Sch71b, Sch72b, SW98, SW99a, Sta77a]. elliptic [Sta97, Sti88, Tho71, VW92, Wan89, WhY88, Wit78b, YN93, Žen89, Žen90]. elliptic-hyperbolic [FG77]. elliptiques [Cha73b, CL75, SP72]. elliptischen [Hei89a, Wit78b]. elliptischer [Kos85, Woh63]. embedded [CH89]. embedding [GI99]. emboî [CR89, Rom85]. Empirical [Neu63]. empirische [Neu63]. enclosing [GKR86, Sch78b]. Enclosure [Leh66, Sch68b, Wet60, Wer69, Hof75]. end [CR80]. endlich [KS73]. endlich-dimensionalen [KS73]. endlichen [FN68, Kup71, WL79, Wag80]. endlicher [GL65, Neu60, Neu61]. Endomorphism [KS73]. Energy [LG76a, LG76b, Mar84, BB95, BB97, CE92, Lus96]. Enlargement [Tsu88]. Enlarging [Lor80]. ENO [AS97a, IS96]. ENO-schemes [IS96]. ensemble [Att68]. ensembles [Gas65]. entier [SA84]. entire [Wil87]. entrant [Ste93a]. Entropy [Son92, Yse99a, BLLN95, CL96a, DLD94, DM96]. entry [Cuy83a]. Entrywise [O’C93]. Entwicklung [Mar73]. Entwicklungen [Wol78a, Gri75a]. Entwurf [Lod70]. epsilon [Wyn64]. equal [Peh88]. Equality [Tay69a]. Equally [Dag90, Sal63, Cul86, DP88a, Pec74]. Equally-weighted [Sal63]. Equation [AU61, Alb61a, Bra65, Due59, Jou71a, ML90, MRMW90, SS66, Sch61a, SG62a, SG62b, SEG62, Uhl62, Wal67, vPS90, AGS98, ADK91, Akr92, AD93, BB95, BBG98, Ben64, Bra81a, BH62, CGH93, CM62, CM70, CE92, Cry67, DN89, DG98, Des70, DH80, Dob81, Dör96, EFM89, GA92, GLM89, GK94, GK85, GK66, Gre80, Grü97, yGL93, Gut81, Had71, Hai77, Hem98, Hof67, Hor75, HP81, HP83, Hüb82, JK91b, JY93, Jou71b, Kes84, KV85, KW99, KS71, Kot99, KP90, Kre68, KS83, KS93, Lin93, LB97, Lou79, Lu97, LM98, MQ81, MW99, Mel95, MS72, MSF91, Mey78, MC88a, MC88b, Mor62, Néd82, NS81, NV80a, O’R87, OP88, OSS90, PB87, Per66, RS70, Rat93, Raz79, Rei89a, RW85a, Sch64a, Seg99, Shi78, Spe75]. equation [Ste72, Sto67, Sun95, gS97, gS98b, Sym66, Tha76, TL91, TV98, Ver98b, WL79, Wag83b, Woo68, Yan94, ZG78, Zie85, dFSS97, de 95, Hai77, Néd82, RW85a, Wag76, Wag80]. Equations [AU61, Auc92, Bar69, Bau65b, Bit64, Boh69, Bol84, Bör90, BMPW66, Bra60, Bra61, Cai91a, CKP95, JP90, Kes79, Kin73, Kre63, Las92, Lig90, MPW66, MW67, Nab92, Nat70, OP64, Red67, Rei90, Rut60, SV93, Ste90a, Sto70, Tan71, Tör63, Wal68, Woź77a, AGS87, AKK74, AGP92, AM94a, ACM99, Alb60, Alb61b, AU62, Alb78, AH72, Ale86, AV86, AYB98, AM94b, AG88, And69, AS87, Ans63, Ant72, Arn79, AW85, ASW93, Aro63a, Aro63b, Atk67, Atk72, Atk73, Auz87, AF89, Axe77, BD83, BDS80, BJ85a, BJ86, Ban87, Ban78, BBFS90, BB92, Bel83, BZ85, BJZ88, BMM87, BS95, BB82a, BB82b, BSV68, BL86b, BD99, Bra82, Bri83, BF66b, Bru78, BN81, BM84, BS66b, BRS71, Bus77, BEHH75, CG88a, CGPM98, Cam70, Cam93]. equations [CMQ84, Car71, CKL98, Cas78, CP76, Cha74, CG88b, CS90, CS92, Cha92, CD73, CS76, CT91, CC95, CŠ79, CMR93, Cho89, Chr70a, Chr70b, CZ99, CC85, CFH68, CW79a, CG67, Coo71, CW78, CS88, CM92, CL96b, Cro80, Cry73, DJ69, Del73, DD98, DC74, Deu74, DE94, DE96, DLO^{+}77, Dix82, DMJ86, Dou61, Dou62b, DJ62, DD73a, DPRW93, Egg84, Egg86, Egg88, ES90, ENV85, EMN88, Ein71, ESS92, Elf80, EP95, Els84a, EG95, EM91, EK99, EGH99, Fae98, FGM67, FF94, FG73, FB89, FR84, FDD92, FL91, Fun88, Gar96, Gau61, Gau73, Gir76, GR79, GK85, GF73, GHT94, Gor70, GM68, GP78, GLC89, GP65, GHR65, GeN80a, GOS99, Gri85, Gue94, GQ98, Gug97, GH99, Gun64b, GP83, Guo88, GNV86]. equations [Gut89b, Hac77b, Hac77a, Hac79, Hac89, Had80, Had69b, HV64, Hai79, Hai80, Hai81, HL88, HR95, HS80, Han94, Han88, Han91b, HP75, He94, HL98, HJ90, Hel87, Her71, HL99, Hoc79, Hoc81, Höh82, Hop81, Hop86, HI71, HL60, Hwa99, iS91, Ise84, Ish84b, IK89, Jac79, Jac82, Jac84, JLR70, JM87, Jay93, JS95, Jeg68, Jer89, JS86, Jov89, Kac76, Kad67, KR84, KR79, KR97, Kau89, KM67, Kea68, Kem72, KJ90, KP95, KKL97, Kin72, Koc95, Kos86, KMS99, Koß81, Kot98, KS91, Kre59, Kre90, Ksc68, KM98, LG76a, LG76b, LS66, LSW85, Lan77, Lau98, LMW84, Le 80a, Le 80b, LP78, Lee61, Leh67, Lei71, Les73, LL80, Lin75, LS95, Liu95, LA97, Lub82, Lub91, Lub92b, LS92, Lub94, LO98, Lyc72, LRT64, LT68, Man85, Man78a]. equations [Man81b, MS96, Mar75, MPP99, Mar70b, Mat76, McA68, McC75, McL94, MS78, Mer67, Mey70, Mey83, MM74, Mil81, MF64, Mit80, Mit81, Miy76, MS95a, Mon92, Moo68, Müh85, Mül79a, Mül80, Nak85, NC93, OP81, OW69, Pal99, Pao85, Pao95, Pao98, Par59, Par61, Par62a, Par62b, Par65b, Per67, PS75, PS96a, Pfl98, Pir82, Pla96, Por78, PQS98, PR86, PS92b, QYK99, QLV91, Rak73, RS96, Rei82, RRS89, Reu95, Rhe88, Ric78, Ris74, RW65, RS68, Roa71, RF94, Roc88, Ros75, RS88, SSVH87, Sar89, Sau84, Saw82, Sch79a, Sch85b, Sch94a, SD67, Sch87a, Sch95a, Sch91, Sch79b, Sch61d, Sch61e, ST73, STR76, Sep72, Sey79b, Sey79a, Sey83, Sha67, She92, Shi78, SK73, Sik84, Slo81, SW98, SW99a, SS98, SG78, Spe78]. equations [Spe79, Squ69, Sre69, Sre71, Sta70, Sta80, Ste61, Ste63, Ste65, Ste93a, Ste73a, SWB91, Sül88, SSBB88, Tan74, Tan92, TT87, Tho90, Tho75, TW68, TW83, Tor91, TS91, Tro71, Tsu88, TK64, Tur78, Ura70, Van79, Ver87a, Ver89, Ver91a, Ver91b, Ver98a, Ver88, Ver92, Voi74, Voi79, VI86, Wag83a, Wah75b, Web94, WA72, WhY88, Wer68, Wer82, WH86, Wet60, Wit89, Wit90, Wol78b, Wri92, fX96, Yam75, Yam76, YN93, Zad76, Žen90, Zen86, Zen97, Zha89, Zha93, dHW73, dHW75, tMS86, vD73, vdHtR81, vPS96, vW77b, AG88, Cro80, Gue94, JLR70, Kes79, Le 80a, Ama72, Dre77, Ess77, För67, GS74, Gek72, Gri70a, Kil78, Mül79b, Nat69a, Nat75a, NW73, Web79, vW77a]. Equiconvergence [AJS90]. equidistant [MS92, Rei84a, Rei84b, RS86, Rei89b, Sch80a]. Equilibration [vdS69, vdS70a]. Equilibria [HISS90, Rap84]. Equilibrium [AT79, JP90, HH76, JM78, LV84]. Equivalence [Ans66, BY93, Eis67a, Hai86, SSS86]. equivalent [Lub88c]. Ergänzungen [Häm65]. Ergodic [BW72a, BW72b]. Ermakov [BMP81]. Ermittlung [Cro76, Wei80, Gai72]. Errata [BW72b, SSV76a]. Erratum [BV88, Bea86a, Ber89c, CS92, Chr70a, DKS90a, Heu01, JN82a, Koh88, Li87, Lub92b, MC88a, SV76a, SV77, Yse86a]. erreur [Att77, Che79, Gou77, LA88, LA89, Tor97]. Erreurs [LV74a, SA84]. Error [Alo96, AYB98, AL85, Auc92, Axe77, Bab71, BD81, Bar91a, BP79, Ber59b, Ber60, Bla72, Boh69, BS60, BS63, Bra61, BS81, BF99, Bre70, BHR77, BHR78, Bri83, CK76, Coo71, DP72, Due59, FMMS99, Fie78, Hae63, HR87, Häm66, HS79, Hel87, Hig87, JO74, Kam71, LW82, Lan66, Leh60, Leh61, Ley90, MvdS76, Nak85, NW76, Not93a, O’C93, OP64, Ort63, PK64, PCV96, PZ99, Por78, RS70, Rei84b, RJ72, Sha68, Slo76, Spa94, Spe78, Spe79, Ste66b, Ste95, Tsa79, VV88, Wah75b, Wil60, Yam80, Yam82, Yam86a, dBP77, vS77, AABM98, AF88a, AFW93, AC92, AO93, Ain96, Ain98, And88, Arn84, AM88, AF89, BKP79, Bak70, BJ85b, BJ89, Ban87, BS92b, BM96, BH92a, BG76b, BL93, BL94, BB95, Ber73, Ber59a]. error [BB82a, BB82b, BLR86, Bra94, Bra73, Bra85, BEN93, CD98, Cha61, Cha79, CL94, Che79, Cho89, CW78, CS98, DS91, Die96, Din83, DMR91, Dur91, Egg92, Eng77, Fae98, Fet87, FJ72, För67, For72, FM92b, Gao89, GN87, GS74, Gek78, GW95a, GF73, Gou77, Gra68, GK72, GB86a, GB86b, Gut72, HL84, HL97, Hai99, HW84, HR94, HP66, Hof67, HMSW99, JT76, Kac84, Kan73, Kin74, Koh87, Koh88, Kos85, Kra63, KT59, Lam70, LW99, LA88, LA89, LdD93, Mei67, MD77, Mer92, Mie79, Mir71, Moc75, MMS92, Nii90, NS95b, NS96, Noc88, Not90, Not93b, Nov86, O’C96, Oli67, Oli68a, PW93, PV72, PS96b, PP80, Pot82, Rei65, Rei84a, Rok72, SA84, Sch81a, Sch61d, Sch61e, Sch84, Sch96, Seg99, SW96, She92, She94, Sie96]. error [Ske86, SS80, Spi71, Spi76, Sre69, Sre71, Ste66a, Ste65, Str73b, Stu85a, Stu85b, SP94, SW80, TW83, Tom84, Tor97, Ver89, Ver91b, Ver98a, Ver98b, VH93, Vog83a, Vui90, Wal97, Wan89, Weg84, Wid77, Woh99, Woź78, Wyn59c, Yam86b, Žit89, Akr84, AB85, Alb61a, Alb62, Att77, Leh67, Mar75, Red77, Ric74b, Sau78, Sch82, Sch79a, SD67, Sch61b]. Error-bounds [Bab71]. Errors [Bra65, CI94, Ehr69, LV74a, Not93c, Cha71, FR84, Häm65, MSF91, Nev77, PB81, PR94, RW85b, SA84, Sha84, gS99, Yam75, Yam76, Zad76, Pot78]. Ersatzkernmethode [Sch79a]. erste [Spr76]. ersten [Lan87]. erster [Het75, Sch61d, Sch61e].

Erweiterung [NW74]. erzwungene [Wer69]. espace [DRW81, Lau67, LH66, Rob71]. espaces [SA84]. essentially [Kir86]. estimate [Auz88, Ber59a, Bra85, CGK92, Fet87, Moc75, RS70, Rei84a, SW96, Ske86, Vui90, SD67]. Estimates [Auc92, Gau75a, Mar87, Par88, SB90, Akr84, AB85, AP85, AP86, Axe77, BKP79, BD81, Ban87, BP79, Ber73, Ber97, BS81, BHR77, BHR78, BNvW89, CD98, Cas85, Cho89, CDR98, CS88, DA86, Dur91, ES90, FMMS99, FJ72, FM92b, GN87, Gek78, GF73, Hel87, HP66, Koh87, Koh88, Kos85, LW99, Liu96, LA88, LA89, MvdS76, Mer92, Nak85, Nii90, NS95b, NS96, Noc88, PCV96, Por78, Pot78, Rei84b, RJ72, Sch82, Sch79a, Sch81a, SD67, She92, She94, SS80, Spa94, Spe78, Spe79, Spi71, Stu85a, TW83, Tom84, VV88, Vog83a, Wah75b, Wan89, Weg84, vS77, SA84]. Estimateur [AABM98]. Estimating [Cha71, CW79b, HP76b, Hig92, Dum98, Man78b, Mei66, OW69]. Estimation [BO90, Gou77, Kat60, KP91, AF88a, AFW93, AO93, Ain98, BB82a, BB82b, Blu61, Bra94, Bra73, Gao89, HW84, Kim96, Kuh93, Mar75, MD77, MMS92, Oli68a, Red77, Sch94a, Sch84, Sch96, Seg99, Zad76, Žit89, Gou77]. Estimations [LA88, LA89, SA84, FM92b, GS74]. estimator [AABM98, DR92, Sie96, Woh99]. estimators [AC92, Ain96, Alo96, DMR91, Ver89, Ver91b, Ver98a, Ver98b]. Étude [Car72a, El 84, LV74a, Pau81a]. Études [Bre71]. Euclidean [BH60b]. Euler [Mar65, EV97, EM91, FF87, GL86, Hor75, LM70, Lyn85, Mar65, ML98, Nab92, Nie80, Pal99, PS96a, SS98]. Euler-Knopp [GL86]. Eulerian [GT78]. evaluated [ELR92]. Evaluating [Mat92, EP75, EP76, EV97, Hal60a, Hal60b]. Evaluation [LV75, Ziv83, Brä81b, CR74, CP70, DP85, Eve64, Hag87, Hen75, Hun72, Hun73, Kut75, MM67, Mon84, Mon87, PE72, Pag81, Pat76, Skr98, Ste66b, Wil59a, Wil59b, LV75, Rei75, Att77]. evaluations [Cur70]. Evans [Ale82b, Ale82b]. even [Sar88, dV93]. ever [Nie95]. evolution [BDS80, BLR89, Hel87, RW85a, vW77b, vW77a, MS95b]. Evolution-Galerkin [MS95b]. evolutionary [Cro80, Dzi91, SSVH87]. Evolutionsgleichungen [vW77b, vW77a]. Exact [All73, Dix82, Gün94, Spr80, AB74, DA71, Gue86, Hav84, MS94, Sar96, Wer78]. Exactness [Wil64, DMR91, DR92]. exakter [Wer78]. example [HS91]. examples [MP95b]. Exclusion [BF60, Ale86]. Existence [BS92b, BS97a, Kar68, Le 82, BLW74, HL60, Kuh93, Pau77, Utr85, Zhe89, För67, Pet70, Spr76]. Existenz [Zhe89, Pau77, Pet70, Spr76]. Existenzaussagen [För67]. existing [Wyn60]. Exit [För93]. exp [Luk61]. Expanding [HB59]. expansion [BLR86, CL94, EV97, Fra62b, LM70, Lyn85, ML98, SP94, VH93, Wes75]. expansions [AS97a, AF89, BD77, BG76a, BK98, CS98, Dix82, Fra62a, Fra63, HL84, Hoc79, Hof67, Lem98, LdD93, Ném65, RDM62, Ste65, Wal91, Wan89, Wol78a, Gri75a]. experimental [Pau81a]. expérimentation [Pau81a]. experiments [Gre62, Ish85]. Explicit [Jou71b, Jou71a, Moc76, vdH72, ACM99, ASW93, Cro80, Gun64a, Hai80, Ish84a, JN81, JN82a, Lee61, Len89, NV96, Rei97, VV88, Woo68]. explicit-implicit [Cro80]. explicite [Cro80]. Exponent [SS66, Sto67, Wag76, WL79, Wag80]. Exponenten [SS66, Sch61a, SEG62, Sto67, Wag76, WL79, Wag80]. Exponential [HMS99, KK97, CRZ99, FL70, GS95a, Mat92, Nii85, Nør75, Ren80, SU89, Sch82, Sug97]. exponentially [Cas81b, GS93b]. exponentials [Ise84]. Exponentialsummen [Sch82]. Exponents [Sch61a, SEG62, Lew71]. Extended [Wol78b, AS83, And71, BL74, Cas80b, CMO91, Shi82]. Extending [Dek71].

Extension [Rob72, BR91, Bul69a, Gas71b, NW74, Oli68a, PSS84, Tim65, Rut63a]. extensions [DDM92, Pat93, Sal61, SZ61, Ver88]. exterior [BD98, CHL85, Ern96, GH92, Gol81, HP83, KS83, Mas87, Sym67]. Extra [Had69a]. Extraneous [VG88]. Extrapolated [EL88, Had71, Mis84, SB83]. Extrapolation [BR88, BRZRS98, BEHH75, GM92, Hac77b, Hac77a, HL88, LL80, Sar89, Sch75b, Sid90, Tur78, BLR86, Bre72, Bre80, BS66a, BS66b, BS67, Cas82, Deu83, DHZ87, FJ75, Föß89, HO90a, Hoc81, JS95, Kin74, Lub89a, Lun85, MY79, Oli71, Qun91, Sal71, Sid82, Str73b, Wal89, Wer68, Wuy71]. Extrema [Wet70]. Extremal [Col61, Bau76, Bro65b, Har78, Men88, Opf79]. Extremaleigenschaft [Bau76]. Extremalprinzipien [Col61]. Extremalsignaturen [Bro65b]. extreme [Ste75b, Whi77]. extreme-discrepancy [Whi77].

F. [Rut69]. Faber [Eie89, ES88, SV93]. face [Per99]. face-edge [Per99]. Factor [Aum63, Din83, NF69, Wyn63, You71]. factoring [HS69, Ste69b, Ste70, Ste73b]. Factorization [Bea79, Bea86a, ER74, Fia90, SS79, BBdH86, Bra92, Cas96, Guo99, Heg91, LE81, Mag92, VC81, Wer72, Kar61]. factorizations [KM91, LA97, MS98, Mes95]. factors [Ben65, DV99, Gau72, Gut87, LTT92, Wyn59a, Wyn59b]. Fairly [Zur64]. Faktorenzerlegung [Kar61]. faktorisierung [Wer72]. Falk [Dur91]. Fall [Mül68]. falsi [PP81]. Familie [Eng77]. Families [Ste90a, VL73, BL70b, BDM85, Med73, MP99]. family [ADG84, Buc64, DML83, Eng77, FP93, FHLM88, GMF99, HP77, JK91b, Mor89, Néd86, Red77, Ste88]. Fan [Ste67]. far [Cro78]. far-reaching [Cro78]. Fast [BL80, CKL98, CGC95, GO94a, HJR88a, Mat96, Rat82, Rei90, Shi79, Sun98, Swe84, Swe91, BL92, Die87, Gir89, GK85, Gut81, Hac79, HN89, Hag87, Heg94, KS98, KC93, LSW85, PE97, Qia88, Web81]. faster [CW77]. Feedback [BV84, Tau76]. Fehler [Uhl62]. Fehlerabschätzung [Ber59a, Ber59b, Ber60, BS60, Bra61, Bra73, Bra85, Due59, Leh60, Rei65, Ric74b, Sch61e, Mar75, Red77, Sch61d]. Fehlerabschätzungen [Boh69, För67, For72, GS74, Kos85, Leh61, Akr84, AB85, Alb61a, Sch82, Sch79a, SD67]. Fehleranalyse [Sau78]. Fehlerdarstellung [Kan73]. Fehlerfortpflanzung [Sch61b]. Fehlerfunktion [Lam70]. fehlerhaften [Bra65]. Fehlerintegrals [PK64]. Fehlerschranken [Nic66b, Alb62, Eng77, Leh67, Mer92]. Fehlerschranken-Arithmetik [Nic66b]. FEM [CFS97, CD98, GH92, SSS99]. Fer [Cas96]. Field [Sta97, Bau62, DGL99, Eck76, KS99, KO98, LM75, Zen68]. Field-of-values [Sta97, KS99]. Fields [Bau68, Hen62, Sto70, Kem68, NS64, Vas88]. fifth [Str67]. Fike [wEC86]. filtered [LS95]. filtering [CRZ99, WW97, Wag97a, Wag97b]. filters [GP78]. final [JHW77]. find [YK88]. Finding [MN68, BWW73, CR93, Cro76, DL97, DD78, HP77, HM89, KvdD77, Lar81, PCT95, Sou69, Sre71, Yam86b]. finis [AABM98, AAG79, Att77, Bre75b, Con84, LL85, Néd82, Tor97, Tro90]. Finite [Akr92, BO90, BS92b, BE86, BS91, BC91, BL93, BB97, BBG98, BS99a, Bia90, Bre75b, BRR80, BRR81a, BRR81b, CK82, CZ98, Cho89, DGP93, FŽ87, FN98, FN68, GL65, GS93b, Han92, Höh82, HMW86, KM67, KKL97, KMS99, KS73, LR89a, Les73, Man81a, MT90, Mis99, Neu60, Ran84, Rau86, Sco73, Seo95, SB90, Ste90a, Ver87a, Ver91a, Yse86b, Žen89, AABM98, AF88a, AFW93, AC92, AK99, AM94a, Ale86, AT79, AP86, AAG79, Arn81, ADG84, Aro63a, Aro63b, AR93, Att77, ASW80, AHS84, ALS90, Bab71, BR72, Bab73, BKP79, BO80, BD81, BV84, BS89, BS92a, BV87, BV88, Ban92, Ban78, BBFS90, BD98, BYI86, BH92a, BG76b, BE87, BL94, BB95, BW95, Ben99, BP79, Ber73, Ber89a, BP92, BDT88, BB82a, BB82b]. finite [BLR86, BR88, BL97, BN84, BH62, Bra66, Bra94, BHR77, BHR78, BDM85, BDDF87, Bri83, BCS96, BEN93, Cai91b, Can71, CMQ82, CMQ84, CPT98, Cas85, Cha98, CGH93, Cha99, CT93, CL94, CC95, Che97, CEL98, CW71, CZ99, CC85, Con84, CD85, Cry73, DD95, Dob81, Dob89, DDW75, DG86, DPRW93, Dry84, Dur90, DMR91, Eis67a, EV97, Ern96, EJ81, EW92, EGH99, Fei87, FŽ88, FN72, FGN81, Föß89, FHLM88, FGM87, GN87, Gas96, Gir76, GR79, GP79, GW95a, Gol81, GLC89, GQ98, GB86a, GB86b, GB86c, GJ89, GP83, Guo88, GC96, GS98c, HC92, HS97, HH76, HL98, Hel68, HJR88b, HH99, HT95, Ish79, Ish81, Ish84a, Ish84b, Ish85, IJ87, JBC92, Jam70, Jen91, JMT97, Joh73, Joh76, JM78, Jov89, Kel65, KJ90]. finite [Kik81, Kil75, Koh87, Koh88, KN84, KNR95, Kup71, Kut75, Kut71, LLS95, LN99, LL85, Le 80b, Le 82, Li90, Liu96, Lou79, LM98, Lub92a, Mak92, Man74, Man81b, Man83, Man84, MQ89, MP95a, Mey79, Mil72, Mil85, Mit80, Mit81, Miy76, Moc76, MG85, ML98, Mon92, Mor89, NS95a, Nat75b, Néd80, Néd82, Néd86, NS81, Neu61, Nii90, Noc89, NW76, O’R84, OD95, Opf82, OSS90, OP96, Pag81, Pao85, Pao98, Par62b, Pat76, Pav94a, Pav94b, PCV96, PB87, PS96a, Per99, Pit79, Pit88, Pit92, PS96b, Prö91, QLV91, Ran79, RSSV97, Red88, RG88, Rei82, Roh98, Rua85, Rua88, Sah85, Seg99, Sep72, Ser99a, SJX97, Soc88, Söd86, SS98, Spi71, SV79, Ste86, Ste88, SS96, Ste61, Str72, SO86, SF86, Tau76, Tor97]. finite [TS91, Tro90, Tsu99, VŽ95, VW92, VV98, Ver87b, Vog83a, Vog83b, WL79, Wag80, Wag83b, Wan89, Wei80, Woh99, Won92a, Won92b, Won93, Yse86a, Yse90, Žen90, ŽV95, Žen95, Zer94, Zha89, Zha92b, Zha95, Zlá68, Zlá69, Zlá70, de 95, Koß81]. finite-difference [Jam70, Jov89, Kut71, OSS90, Par62b, Spi71, Ste61]. Finite-Dimensional [KS73, BRR80, BRR81a, Rau86, Seo95, Ale86, Söd86, SV79]. finite-element [Ern96, Joh73, Tro90]. Finite-part [Bia90, EV97, Kut75, ML98, Pag81]. finiten [Wei80]. Finiter [Nat75b]. First [vPS90, AU62, BL92, Bra66, Bru78, BEHH75, DMJ86, Egg86, EG95, FGN81, GP78, GHR65, HP75, Het75, HvdR63, KS93, Lan87, LP78, MGR87, Moc76, Mor99, Sch61d, Sch61e, Spr76, WA72, dHW73, vPS96, Ano99]. first-order [FGN81, MGR87, Sch61d, Sch61e]. fit [SW93]. fitted [Cas81b, GS93b]. Fitting [ST91, GA87, Spä86]. five [BP85, OO98]. Fixed [DD98, Ale71, BM87, GK83, HL84, Jer89, Kal74, Mil81, MO91, SD67, Sch78b, SV79, VG88]. fixed-point [BM87]. fixed-points [Sch78b]. fixed-stepsize [HL84]. Fixpunktsatz [SD67]. Fixpunktsatzes [Ale71]. Fixpunktverfahren [Sch78b]. flat [BDT88]. flexion [DN86]. Floating [Dek71, Pic72, Wil60]. Floating-Point [Dek71, Pic72, Wil60]. Flottante [Pic72]. flow [Agu89, Alt80, BN90, BS92b, BL93, BL94, Ber89a, BH71a, Cha98, DD95, Eng96, Fei87, GW94, GW95a, Man84, MP92, Mor62, NS95a, SS96, VV98, WKW97]. flow-theory [Agu89]. flows [ABF99, Dob89, Heb88, LR89a, NV96]. fluid [BS92b, Dör90, Gas96, NS95a]. fluides [AG88]. fluids [AG88, Yse97b, Yse99a]. Flux [SB90, PLCC92]. Fock [FDD92]. Fokker [DLD94, Lem98]. fold [BCD86]. following [GI99, MS85]. Fonctions [Att68, Aus71, Dia80]. Forced [Wer69, Tau76, vD73]. Form [Lin76, MW68b, MW68d, Axe77, Bre75a, Gri85, Ric74a, Gri85]. Formal [dJ77, Tho90, VB92a]. Forme [Bre75a]. formelles [Tho90]. Formeln [Wyn60].

Formes [HW87, Tao84]. Forms [Ric65, HW87, Lin76, MF64, Tao84]. formuale [AB85]. Formula [Bra60, Eck68, Mar65, BP98, DD75, GL87, Kra63, Lev82, Mal99, Mar70c, Mar73, PCT95, Sch77b, Sty79, Sug97, Xu94, Zla78, Zla81]. formulae [BR91, Ber72, Ber89b, Ber89c, BS81, Bro78, Cas80b, Cas81b, Cas81a, CH86, CH89, Dry94, DV97, Ghi68, GR70b, Gue86, Gue88, Gün75, Hun72, Nik92, Not90, Not93a, Not93b, Pal76, Pin75, Sch78a, Sha67, Ste66b, Str84, VC92a, Wal97, vdS72, vS77, Wal98]. Formulas [BH90, CD90, Loc73, Sal71, Sch72a, Akr84, BM96, BLW74, Bra85, CR80, CW71, Dic87, El 79, Eng77, For75, FF87, Fri70, GG86, HP76a, Hen79, JN86, Jet76, Kam71, Köh95, LS70, MS67, Mon84, Nov88, OPS76, Pau77, Peh88, Peh90, Sab78, Sal63, Sch87c, Str66, Str67, Str68, TM73, Won82, Wyn60, Yse80, vdH72, vdHtR81, CGG90, Loc80, Möl76, Pal77]. formulation [AGP92, AYB98, AF88b, Ber89a, BH62, FHLM88, Per99]. formulations [BL97, Mat93a, Mat93b, de 95]. formules [GG86, Sab78]. Fornberg [DHP99a]. Fornberg-like [DHP99a]. Fort [Jou71a]. Fortsetzung [Nie73b]. Forward [Stu85a, Stu85b, Lu97, LM98]. forward-backward [Lu97, LM98]. founding [Ano83b]. Four [Ste99, BP85, KR79]. Fourier [BS59, BS59, Bet70, CMQ82, Cle66, DA86, Gau72, GHT94, yGL93, Gut81, Gut87, Hag87, Hal92, Heg94, LSW85, Nat85, Pat76, Reu95, Ste66c, Web81]. Fourier-transforms [Ste66c]. Fouriertransformation [Hal92]. fourteen [Gün75]. Fourth [Qun91, BL86a, BBG98, FH90, Her90, LS70, Nat69a, Sch84, SS95, Tro76, tMS86]. fourth-order [BL86a, FH90, LS70, Nat69a]. Fraction [BW72b, Arn80, Fra62a, Fra62b, Hay65, RDM62, Sto83, Van81, Wyn63]. fractional [CM80, MM74, SA84]. fractional-order [SA84]. fractions [BJ85b, BJ89, BW72a, FJ72, Fie78, Gra68, Gut89a, HP66, Jac85, JW88, JTW89, Lor92, NW84, RW85b, Spe71, TW80, Wyn59a, Wyn59b]. Frage [Ans66]. framework [AM88, CD91, HA98]. Frankel [Jou71a]. Fréchet [Mat92]. Fredholm [Atk72, Cha74, GP78, HS80, HP75, Jeg68, Kau89, LP78, LL80, RS70, Rei90, Ric78]. Fredholmscher [Jeg68]. Free [Hae63, Mey84a, Alt80, BDG97, BB95, BB97, Bra71a, BRS92, Bro85, Bro78, BD72, CDV99, Cha97, CE92, Dry94, Enn95, Fag81b, Fag82, GK83, KjS95, Kio79, Kre71, Lub89b, MW99, MSF91, Mey70, Mey78, Noc88, PW93, Pau77, RBS82, SV76b, SV77, Wul73]. freien [Bro78, Pau77]. French [Tor97, AABM98, AG88, AAG79, Att68, Att77, Aus71, BN90, BK96, Bre71, Bre72, Bre75a, Bre75b, Car72a, Cha84, CR89, CL71, Cha73b, CL75, Cha79, Che79, Con84, Con80, Cro79, Cro80, Del80, DGB80, Del73, DD84b, DD84a, Des61, DN86, Dia80, DRW81, El 82, El 84, FM92a, FM92b, Gas63, Gas65, Gas71a, Gas71b, Gou77, GG86, Gue94, Hai77, Hil83, HW87, JLR70, Kes79, LV74a, LV74b, LV75, Laf75, Lau67, LL85, Le 80a, LV84, Leb70, Leg66, LH66, LA88, LA89, Mai67, Maî68, Mig76, MK80, Mor74, Néd82, Par73, Pau81a, Pau81b, Pic72, Rob66, Rob71, RM72, Rob76, RW85a, Rom85, Sab78, SP72, SA84, Tao84, Tho90, Tis72, Tro90]. frequencies [Par61]. frequency [Hac89, Hac92, Kim98a, Wag97a, Wag97b]. Fresnel [Bul67, Ném65]. Freud [DD99]. Friedrichs [SSS86, HMSW99]. Frobenius [Gre80, HL64a, HL65]. full [LdD93, Reu88a, Reu88b, ZL94]. Fully [Cho90, CZ99, Ain98, ADK91, LK78, MS67, Sac72]. Function [BGBMN91, Fil64, Lie90, Nat71, Neu63, Ziv83, AGP92, AYB98, BH61b, CGH76, CV70, CGT94, Cur70, DDW75, FS70, For72, GH73, GS95a, IS96, Jar94, KT59, Lam70, LTT92, Lev82, LW99, Luk61, MK60, Nit62, Nit69a, Nør75, Sch81b, Sch81c, Stu89, Tem83, YK88, Zen84b, Zet70, ZL94]. function-vorticity-pressure [ZL94]. functional [Bel83, Cry73, Eyd84, Eyd85, Hor75, Jac79, Jac82, Jac84, Kem72, Liu95, Mei97, Per66, Ste65, Sug97]. functional-differential [Bel83, Hor75, Jac82, Liu95]. functionals [BL76, BH71b, GS74, IS96, KK74, KK76, Nie73a, Tsa77]. Functions [Att68, Boe64, BD68, Dir69, Ehr69, Ess60, GK64, GG90, Uhl92, WB69, AS83, AV85, AB87, AL68, Ari98, Aus71, BJ89, BM96, BLW74, Bit70, BL81, Bra71a, BL67, BS68, Bul65a, Bul65b, Bul69b, Car72b, Cha71, CR74, CK74, CK76, Chu79, CMW63, Coc65, CW78, CW79b, CV84b, DGM88, DB77, Des63, Dia80, Dör66, Dry94, DM90, Elb91, EA88, Faß97, FP90, Goë68, vG72, Hay65, HSS96, Her70, HvdR63, HP81, HV85, Hsu61, Hsu62, HdH85, IK77, IS96, Jar94, JS91, Joh69, Kiw84, KWW85, Kre72b, KM91, Kup71, Lam80, LMS77, Lub88c, MRW76, MY79, Mil65a, Mor80, Nie73a, Nin66, Not90, Not93a, Not93b, NR96, NS83b, OW69, Pom71, Rei67, Rei71b, Riv62, Rob72, SSV76a, SSV76b, Sal75, Sal64, Sch83, SS65, Shi82, SW93]. functions [Spe71, Ste71, Sto61, Str73a, Str63, SP94, Tay69a, Tsu72, Tsu73, Tsu75, Utr83, VG88, Wah75a, Web91, Web81, Wer62, Wid77, Wil87, Wil97, Wyn60, Wyn62, vdGK86, vG82, Ant72, Bar84, Bro78, For75, For78, KK74, KK76, Kre71, Mar73, Pau77]. Fundamental [RR90, DGM88, DV99, Par61, RS95]. Funktion [For72, Nat71, Nit62, Zet70, Neu63]. Funktionale [KK74, KK76]. Funktionalen [GS74]. Funktionaleng [For78]. Funktionen [Dör66, Ehr69, Ess60, Kre72b, Kup71, Str73a, Wyn60, Bar84, Bro78, BL67, BS68, Dir69, For75, Häm66, KK74, KK76, Kre71, Mar73, Pau77, Sto61, Wer62, WB69]. Funktionenräumen [Nit69a]. Funktions [Häm66]. further [Bau61, BG71]. FVS [LLS95].

G [Dir69]. G. [Dir69]. Galerkin

[BS60, Kos85, AGS98, AM94a, ADK91, AS97b, AL96, Axe77, Bak82, Ber91, Ber97, Ber59b, Bla72, BS60, CGH76, CT93, Cho90, CS98, DP72, DD73a, DD73b, DD74, Els84a, Fae98, FJ75, Fis93, For72, FHLM88, Gek78, HHM76, HL98, Hel87, Kos85, Ley90, Lu97, MT90, Moo88, MMS92, MS95b, Nak85, PR94, PR86, Prö91, Ric78, RS88, SK96, SS97, SW96, Sül88, TW83, Wah75b, Web94, XS92]. Galerkin-characteristic [Ber91]. Galerkin-Näherungen [Kos85]. Galerkin-type [AS97b]. Galerkin-Verfahren [BS60, For72]. Galerkin-wavelet [XS92]. Galerkinschen [Ber59b]. Galois [Eck76]. GAMM [PS59]. gamma [AB87]. ganzer [Dir69]. ganzrationalen [Zur64]. Garabedian [De 68]. gas [Eng96]. Gauss [Bra78, Bra78, Sau78, Bor94, BD72, BC92, DH89, EL88, GK83, Häu86, Hun72, Kno89, KWW85, Not93a, Not93b, Par88, Peh88, Peh90, RM72, Sau78, Sch85a, Ste66a, Ste66b, Str84, Akr84, AB85]. Gauß-Elimination [Sau78]. Gauss-like [Bor94]. Gauß-Quadraturformeln [Akr84, AB85]. Gauss-Tchebycheff [Peh90]. Gauss-type [Hun72, Ste66a, Ste66b]. Gaussian [Akr84, AB85, BP98, Bus71, CM89, Cry68, DK77, Die96, För93, Huc98, Not90, SD72, Sch90, Str69, Stu85a, Stu85b, vV77, vS77]. Gautschi [HL99]. Gautschi-type [HL99]. geb [GB71, GB81]. Gebilde [Woh63]. gebrochen [MW60b, Wer62]. gegebener [Mit73]. Gelenkvierecks [WS59]. genauen [Hav84]. Genauigkeit [Kre59]. General [AM88, BD68, CD91, CV84b, FM92a, MW68d, Mül68, RS95, Sto93, WS59, Wat68, Wyn64, AGS98, AL68, BD82, Bre80, CM95, CG95, DNvW84, Döh82, FZ93, Fre94, GT84, Haa60, Kle81a, Kre72a, Mid75, Müh78, Müh79, Sch85a, Sch71a, Shi79, Shr91, Slo81, Ste97, Ste98, TT87, vW77a]. générales [Gas63]. Generalisation [ES88, LS66, Tay69b, Lau67]. generalisations [Bro64]. généralisé [MK80]. Generalised [NF69, Nin66]. généralisée [Mai67]. Generalization [Ans66, wEC86, Chu76, Lau67, Mar65, Mei63, Sid90, Bad89, Bau61, Luc70, Mil80, Pie72, PP81, Sid82, Ste67, Ste73b, Wil71]. Generalizations [BV68, Hou71, Hou73, NN87, FP66, GL86, Nie80]. Generalized [FNP82, Gas63, Jia98, KR79, LM69, Mai67, ML90, Zie97, Arn80, Axe87, AC94, Car92, CR93, CDR98, CW79b, DZ75, Fro88b, GR86, JT70, Joh77, Kea79, Kno89, Kot99, LT98a, LR93, LH88, Luk93, Mel99, Mey84b, MK80, Mor94, NSSS83, PW93, Pyl67, Sar96, Sch75c, Slo80, gS98a, Ver77a, Ver77b, VI86, Zha96, Van85]. generated [Eie89, Fis61, Nie85, Sto84]. generating [Med73]. Generation [Lin61, CHS96, Yse99a]. generator [LK78]. generators [DP88b, EHG89, Mar70a]. généraux [FM92a]. Geometric [SSV76a, SSV76b, SVN76, CCCD93, CKSS97, DM88, MH72, Par73]. Geometrical [Nic74]. géométrique [Par73]. Geometry [LL90, AR93, EJ81, MC88a, MC88b]. German [Sch68a, Ade69, Akr84, AB85, AU61, Alb61a, Alb61b, Alb62, Ale71, AH72, Ale73, Ale76, AV76, Ale82b, AV86, Ama72, Ans63, Ans64, Ans66, Ans67, Ans69, Ans73, Ant72, Arn79, Arn80, Aum63, Bar84, BS59, Bau76, Ber59a, Ber59b, Ber60, Bey78, Bit64, Bit68, Bit71, Bla77a, Bla77b, Bod68, Boh67, Boh69, BS60, BS70, Bra65, Bra70, Bra74, Bra77, Bra60, Bra61, Bra73, Bra78, Bra85, Bro78, Bro65b, Bro65a, BL67, BS68, BH74, Bul64, CS59, Col61, Com59, Cro76, Cro77, Dir69, Döh82, Dol81, Dör66, Dör70, Dre77, Due59, Eck68, EH70, Ehr69, Els69, Eng72, Eng77, Ess60, ES73, Ess77, EN80, Fel61, FN68, Fie87, Fil64, Foc65, For75, For78, För67, For68a, For72, För82, Fre60, Gai70, Gai72, GS74, Gek72, GL65]. German [GK80, Gor70, GB71, GB81, Gri70b, Gri70a, Gri75b, Gri75c, Gri75a, Gri85, GKR86, Gün75, Haa60, Had68, Had69a, Had72a, Had72b, Hae63, Hal92, Hal60a, Häm65, Häm66, Hav77, Hav84, Hei89a, Her74, Het74, Het75, Hof69, Hof70, Hof75, Hüb72, Hüb82, Jeg68, Jet76, Jet84, Kad67, Kan73, Kar61, Ker66, Kil78, Kle81a, Kle81b, Kna71, KK74, KK76, KR82, Kos85, Kos86, Koß81, Kra68, Kre59, Kre63, KS73, Kre71, Kre72b, Kul68, Kup71, Küp76a, Küp78, Lam70, Lan87, Läu61, Leh60, Leh61, Leh63, Leh66, Leh67, Lin76, Lin62, Loc73, Loc80, Lod70, Lor77, MW60a, MW60b, Mar75, Mar65, Mar70b, Mar70c, Mar73, Mei63, Men74, MS78, Mer91, Mer92, Mey84b, Mit73, Möl76, Mül68, Mül79a, Mül79b, Mül80, Mül82b, Nat69a]. German [Nat69b, Nat70, Nat71, NW74, Nat75a, Nat75b, Nat77, NR68, Neu63, Neu60, Neu61, NW73, Nic66a, Nic66b, NY83, Nie73b, Nit62, Nit68, Nit69a, Nit69b, Nit70, Nit76, NO68, Ort86, Pal76, Pal77, PK64, Pau77, PS66, Pet70, Pit73, Pit75, PS82, Pot78, Red67, Red77, Rei65, Rei71a, Rei75, Ric74b, RL66, Rut60, Rut63a, Rut63b, Rut66a, Sac73, Sau63, Sau73, Sau78, Sau84, Saw82, Sch73a, Sch82, Sch70a, SS66, Sch71a, Sch79a, Sch61a, SG62a, SG62b, SEG62, Sch70b, Sch77a, Sch72a, Sch64a, Sch66a, SD67, Sch68b, SV70, Sch78b, Sch64b, Sch75b, Sch80a, Sch61b, Sch61c, Sch63, Sch70c, Sch73c, Sch61d, Sch61e, Sch66b, ST73, STR76, Sch78c, Sch80b, Sch75d, See82, Sik59, Sik61, Smo90, Spr76, Sti59]. German [Sto61, Sto64b, Sto67, Sto70, Str73a, Tau76, Tör63, Tre62, Tro71, Tro76, Uhl62, Vol86, Vos75, Wag76, WL79, Wag80, Wal67, Wal68, WS59, Web79, Weg78, Wei64, Wei80, Wen68, Wer62, Wer69, WB69, Wer72, Wer78, Wer79, WH86, Wet60, Wet70, Wit77, Wit78a, Wit78b, Wit78c, Woh63, Wol78a, Wyn60, Yse80, Yse83, Zet70, Zhe89, Zur63, Zur64, vF70, vFvH73, vW77b, vW77a]. Gerschgorin [Car91, HVW70, JS70, MV68a, MV68b, Med71, Med73, Por66, Tod65, Var64, VL73]. Gershgorin [Bau68, DZ75, Els73, Ste67, Zen84a]. Gesamtschrittverfahren [Ker66]. gest [GB71, GB81]. gestellter [Nat77]. gestörter [NY83]. get [BCD86, Kac84]. gewisse [Ber59a, Wyn60]. gewissen [För67]. gewisser [Ans64]. gewöhnliche [Ber59a, Gri85, Sch61e]. gewöhnlichen [Arn79, För67, Web79, Wet60]. gewöhnlicher [Bey78, GS74, Gri70a, Mül79a, Mül79b, Mül80, Nat70, Nat71, Rut60]. gewönliche [Küp76a]. gewönlichen [Mar75, Sch61d, Wer69]. gewönlicher [Ess77, Gri70b, Nat69a, Nat75a, Tro76]. Gibbs [GS95a]. Ginzburg [Che97, DGP93, GHR97]. Gittern [Hal92]. Given [OP64, Mit73, Rap77, SD72]. Givens [MC84, Rat82]. Gleichmäßig [GKR86]. gleichmäßige [NY83]. gleichmäßigen [Mit73]. Gleichung [Gek72, Hüb82]. Gleichungen [MS78, Sau84, Bit64]. Gleichungssysteme [Alb61a, Alb61b, AV86]. Gleichungssystemen [Bra65, Due59, Wal67]. Gleichungsystemen [Sch64a]. Global [BR81, hDxY99, Han80, LPDT70, Pyt02, Arn84, Dan70, DD75, DH97, HL84, HW84, Har82, Moc75, Nev77, NS95b, NS96, RH93, Rus84, Sch81a, Ske86, Wer78, Vol86]. globale [Wer78, Vol86]. globalization [Kno89]. goes [HC92]. good [BP85, KZ74, OSS90]. governed [Arn82]. GPN [Tor91]. GPN-stability [Tor91]. Grad [Gün75, MSF91, Sik59]. Grad-Schlüter-Shafranov [MSF91]. grade [AG88]. Graded [SZ91a, Dah85, Mat96]. Gradient [Dan67, Fre90, AS83, AL86b, Axe87, AC94, BC93, BC78, BG78, DS91, Döh82, ELP90, For68b, FR86, GT84, Gin63, GT72, Gre79, Had72b, Jer89, KJ90, MK80, NBE97, NS81, PB87, Rod73, Shi78, Slo80]. Gradientenverfahren [Döh82, Had72b]. Gradients [Not93c, vdSvdV86, KN84, Ste75b]. grading [JY93]. Graeffe [Pól68]. Gram [Saa86]. grandes [LV84]. grands [Leg66]. graph [GK94, Har62, HL64a, SS98]. graph-norm [SS98]. graph-theoretic [HL64a]. graphs [Dij59, Hea66, KvdD77, vG82]. greatest [Akr88]. Green [CV70, For72, HS79, Nat71, Sus64]. Greenberg [She77]. Greensche [For72]. Greenschen [Nat71]. Greetings [Cia98]. Gregory [Mar70c, Mar73, dV93]. Gregoryschen [Mar70c, Mar73]. Grid [Hua90, Tör63, Yan94, AM94a, CHS96, GO94b, Grü97, HM83, Hac89, Hac92, Hop86, NR95, Ste93a, Tha76, Wit89, Wit90]. grids [CSZ96, ELV91, EW92, Föß89, GOS99, Gri83, Hal92, Koc95, KNR95, Pfl98, Rei84a, Rei84b, Sch95a, Sus64]. großen [Nat69a]. großer [Sch73c]. größten [Döh82]. groundwater [WKW97]. Group [Fel61, FN68, Dix67, Eck76, Lin62]. Groups [Com59, GL65, Neu60, BNZ72b, BNZ72a, Neu61]. growth [Rei71a]. Grundbereich [Sto70]. grundlegende [Kle81a]. Gruppe [Fel61, FN68]. Gruppen [GL65, Neu60, Neu61, Com59]. Gruppenelementen [Lin62].

H [BH87, Bra71b, Lin61]. H.

[Bra71b, FP90, GP85]. H.-O. [GP85]. Haar [Car72a, Van82]. Hadamard [Bia90, EV97, Pag81]. Hakopian [Wal97]. halbgeordneten [Boh67, Kra68]. halblinearer [AU61, Ans63, Ans67]. Half [AU61, ASW93, DRW81, Kra68, AS87, Ans63, Ans67, Boh67, Egg88]. Half-explicit [ASW93]. Half-Linear [AU61, Ans63, Ans67]. Half-Ordered [Kra68, Boh67]. Half-Space [DRW81]. Halley [DD75, Her91, Yao99]. Hamilton [FF94, Grü97, Hop86, KMS99, LS95]. Hamiltonian [BMX98, LV97, McL94, Rei97, jS99, Tou92]. Hammerstein [Gek72, Kil78]. Hammersteinschen [Gek72, Kil78]. Han [Sch81b, Sch81c]. Hand [OP64, CR94, Man78a, Ste90b]. handling [Ric74a]. Hanging [SSS99]. Hankel [Coc65, FS70, FZ93, HJR88a, HJ90, HSY91, Laf75, MK60, Ris74, Tyr94]. hardening [Car99]. hardly [Nie95]. harmonic [BHH88, Mel99, Mon92]. Hartree [FDD92]. Hausdorff [Gre82, PS66]. Hausdorffsche [PS66]. Having [Bra65, Eng77, Tay69a]. heat [Des70, Die87, GLM89, GK94, KV85, Lu97, LM98, Mor62, Raz69]. Helmholtz [DG98, EO99, Ern96, Gol81, Kim98b, KS83, KS93, Lin93, LB97, de 95]. help [Men74, WW84]. Henrici [GV88]. Hermann [OP96]. Hermann-Miyoshi [OP96]. Hermite [DD84b, Arn80, BSV68, BH71b, Cla78, DD84b, DD84a, ES73, GP71, GM82, Gou77, Her70, Kan73, Kre72a, LL90, Mel72, Müh81, OP81, Rei89b, RS91, Wei92b]. Hermite-Interpolation [Arn80, Kan73]. Hermite-spline-interpolation [Rei89b]. Hermitian [CD79, CL71, Fre90, FN91, GH95, Mue66, NN87, Ste69a, Ste76b, Yse80]. hermitienne [CL71]. Hermitischer [ES73]. Hessenberg [MW68c, MW68d, MPW70, Zur63]. hexahedral [Man84]. hierarchic [Ain98, Sch96]. Hierarchical [BDY88, AABM98, BB92, Hal92, Osw92, Ste97]. hiérarchique [AABM98]. hierarchischen [Hal92]. hierarchy [RS96]. High [Cas78, Cas81a, Dix67, GM68, Lau98, NR96, Pru75, Cas81b, CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, CC85, FGM67, FF94, vG72, GM94, Had69b, KW81, LR73, Oga89, Pat76, Sha67, Ste61, Ste90b, Sun95, Sun98, Ura70, Wag76]. High-dimensional [NR96, Oga89]. high-order [CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, FF94, Sun98, Ura70]. Higher [Ade69, Man74, PV72, vV78, Ans63, ADG84, CM95, CR93, GW86, Kre59, KNR95, PPV69, Rei97, She92, VD91, Vas88, Wal91]. Higher-Order [Ade69, CR93, PPV69, Rei97]. highest [Sri94]. highly [KW96]. Hilbert [BG78, Bro78, CK74, CK76, Ley90, MO91, Sac73, Yam68]. Hilbert-Raum [Sac73]. Hilbert-spaces [Bro78]. Hilbertian [Ben86]. Hilberträumen [Bro78]. Hilfe [Bit68, Bit71, Ess60, Mar73, Men74, Red77, Sau63, Tro76, Wen68, Zur64]. Hill [WL79, Wag80, Wag83b]. Hilliard [BB95, CE92, EFM89]. Hillschen [WL79, Wag80]. Hochschule [Sau67]. hoher [Kre59]. höherer [Ade69, Ans63, Wag76]. holomorphic [Pom71]. holonomic [Han92, RG88]. Holt [BV88, BV87]. homogène [Cha84]. Homogeneous [MRMW90, Cha84, Tro71, VH93, Wam94]. homogenization [ABFJ97]. homogenized [Mar87]. homotopies [LMPY85]. Homotopy [LR89b, BM84, Chu83, LSY87, LS87, Lor80, ZH93, Zul88]. Hopf [Bau65a, Ber82, CG88b, GS96a, LO98]. hopscotch [tMS86]. Horner [Rei71a, Fis61]. Hornerschema [Rei71a]. Householder [Ano83a, Ano86, H. 73, Wil78, Ano98a, Ano98b, BG65, CV94, Cup84, JHW77, MRW68, Mue66, Ort63, Wil62c]. Huang [Sta81]. Hybrid [BZ94, Bre75b, Qia88, SV93, FF97, MS96, Mil85]. hybrides [Bre75b]. Hyperbolic [Sau63, Tör63, Ans63, BS97b, Bri83, Bro85, CGH93, DH80, DJ62, EK99, Fer97, FG77, Gek74, GHT94, GM66a, GM66b, Hac77b, Hac77a, Hof75, Koß81, Les73, LT98b, Moc76, Roh98, Sta80, Ste61, Str86, Wag83a, Wer68]. hyperbolic-parabolic [EK99]. hyperbolische [Hof75, Sau63]. hyperbolischen [Ans63, Koß81]. hyperbolischer [Tör63]. hyperelastic [LV84]. hyperélastiques [LV84]. hypergeometric [Tem83]. Hypersingular [vPS90, Heu98, Heu01]. hysteresis [HM89].

IA [Kot98]. ice [CDV99]. ideal [Can71]. idempotent [HC63]. identification [GG87, GHMG85]. identities [Sus64]. Identity [Bre70, Cla78]. if [VN93]. ihre [Alb61b, Mar65, Sch66b]. II [Gri75c, Häm65, Kle81b, BS77, BB82b, Bit71, BLLN95, BHR78, BRR81a, BNZ72b, Bru78, Bul65b, Cha88, CK76, CSV68a, DD84a, DGP93, Dun68, Eyd85, Fag81a, FM92b, Fra62b, Fra77, FHL86b, Gau63, GG87, GV61b, GP65, Gri75c, GB86b, Hac92, Häm65, Hou68, Ish85, Kle81b, KO98, Kor96, LG76b, Lub88b, LdD93, MW60b, Mat93b, May85, MV68a, MM67, Mil65a, Mit81, NS96, Osb68, Pau81b, PV72, Rei71b, RS91, SEG62, Sch81c, Sch68b, Sch86, SSS99, STR76, Spe79, Ste73b, Str64, Str67, Stu85b, SP94, Tod88, Ver91a, Ver91b, Wil59b, Won92b, Wyn59b, Yam76, Yam82, Mül79b, STR76]. II. [LR89c]. III [Gri75a, BRR81b, Bul69b, CSV68b, ELV91, Fra63, FM68, Gau78, Gri75a, GB86c, MV68b, SV78, SS66, Mül80]. IIIC [Kot98]. Ill [Dem87, BRZRS98, CRZ99, Deu74, ELP90, Han91b, HNS95, Hào94, nJyH99, Kal98, Kin92, Nat77, PV90, Pla96, RvdS90, Rie97, Ruh70, Sch93, Sch98, Sch87d, Sch88b, Wil59a, Wil59b, Wil72]. ill-conditioned [BRZRS98, Deu74, Ruh70, Wil59a, Wil59b, Wil72]. Ill-Posed [Dem87, CRZ99, ELP90, Han91b, HNS95, Hào94, nJyH99, Kal98, Kin92, Nat77, PV90, Pla96, RvdS90, Rie97, Sch93, Sch98, Sch87d, Sch88b]. ILU [BW99, Ste94a]. im [Mül68, Sac73, Wen68]. Image [CL97, CCCD93, DRW81]. Imbedding [Bör90]. immune [BMR96]. Imperfect [AS83]. implementation [Akr80, BMP81, HS91, MP85, Sch81c, SS80]. implementations [CVV88]. implications [Küp76a, Küp78]. Implicit [ACM99, Kre63, MW68a, Ver69, dHW75, Ale82b, AF89, BD83, BZ85, Bia91, Cro80, Dub70, GM66b, Had71, Hun86, Lub89a, Spi86, Str86, Ura70]. Implicit-explicit [ACM99]. implicite [Cro80]. implicite-explicite [Cro80]. implicitly [PS82]. implizit [PS82]. implizite [Kre63]. improperly [HA98]. improve [Häm65]. Improved [DA86, MF64, PV72, ZG78, BL78, HL65, Mid75]. Improvement [IU82, BYI86, Sch61e]. Improving [Dum98, Egg86, Lev89, Cas82, Per66, Slo88]. impulses [DD98]. IMT [MI82]. IMT-type [MI82]. in-place [Heg94]. incertitude [LV75]. includes [Wal97]. including [Reu88b, AH72]. Inclusion [Bau66, Car91, DZ71, Leh63, MO64, Mor72, Ale86, BH73, Enn95, HVW70, JO72, JT76, MV80]. inclusions [Els73, SV79]. incompatible [Chi99]. Incomplete [BS99c, Fia90, Guo99, Lou86, LR89c, AB87, Bra92, Bul69a, Heg91, LA97, Mes95]. incomplete-factorization [Bra92]. incompressibility [Dör90]. incompressible [ABF99, Cha98, Dob89, FHLM88, HB97, Le 82, LR89a, Man84, SSBB88, Ver87a, Ver91a, Vog83a, Néd82]. incompressibles [Néd82]. inconsistent [CEG83, Elf80]. incorporation [IU82]. increase [Kac84]. increased [vdH72]. increasing [Cle71]. Incremental [CT91, CT93]. Indefinite [GG90, Guo99, Yse89]. independence [CP94a]. Independent [Due59, Sau63]. Index [Ano83c, Hek87, ST89, ASW93, CM95, CG95, HL64b, Jay93]. indicator [GW94, GW95a, SS98]. indicators [AYB98, Fae98]. indicatrix [Fag81b, Fag81a]. Inequalities [Row71, Arn82, BH60b, BM87, BHR77, BHR78, BS85, CC91, CD85, Dan73, Ess77, Fie66, Gre75, Gri85, HM83, Kar68, Kor94, Kor96, MO64, Mit78, Noc89, PSS90, RN98, Sah85, Sch66b, She77, Spa94, Küp78, Red77]. inequality [Bau61, Bau65a, Bea86b, BH74, Dun88, Fet87, HM86, Loi67, Mar62, Sch60, Vui90]. Inequation [Mei63]. inertia [CD79, Sch65]. Inexact [GO88, Bör89, MS94]. inexactQP [Kan98]. inexactQP-based [Kan98]. inf [Dur91]. inf-sup [Dur91]. infeasible [SW99b]. infeasible-interior-point [SW99b]. Infinite [BF70, Cle71, DG98, Fis93, Ghi68, Han82, IK89, Pat76, SVN76, Sim78, Tha76, Zhe89]. infinity [Nev77]. Inflections [Sak97]. influence [Ain96]. information [BLLN95]. infra [Lyn63]. infra-max [Lyn63]. inhomogene [Leh60, Leh61]. Inhomogeneous [Leh60, Leh61, Gek78, GP83]. Initial [AU61, Ans69, Sau63, Wat68, vFvH73, AU62, Ans64, Ans67, Arn84, BS96, BHK94, BEHH75, Cas81a, CEL98, DLO^{+}77, DH80, Dut99, Gek74, GK80, GeN80a, GN66, Hac77b, Hel87, Hof75, Hul71, JN81, JN82a, JN82b, Kra86, Kre59, Lub94, Mar75, Mül82b, Nev77, Spi66, Spi76, Spi83, Sta77b, Ste63, vdGK86]. Initial-Boundary [vFvH73, BS96, DH80, GK80, Lub94]. initial-value [Ans67, BHK94, Spi66]. instabilities [MO78]. Instability [Ste68, ELP90, He97, Wyn59d]. integer [Coh59, EFP98, Sch95a]. integers [BB62]. Integral [Ben64, Bra60, Bra61, BRS71, Rei90, Wal97, vPS90, Ama72, AS87, Ant72, Atk67, Atk72, Atk73, BJ85a, BJ86, BH65, BNZ72b, BNZ72a, Bru78, BN81, Cha74, CS90, CS92, Cha92, Che79, CGC95, Coh59, CS88, Cry67, DJ69, DMJ86, Egg84, Egg86, Egg88, ES90, EP95, Els84a, EG95, Fae98, GK85, GK66, GP78, GOS99, Gut81, HS80, HP75, Her71, Heu98, Heu01, Hoc79, Hoc81, HP81, HP83, Jeg68, JS86, Kac76, Kad67, Kau89, Kel65, Kil78, KP90, KS83, Kre90, KS93, Ksc68, LSW85, Lau98, LP78, LL80, LS92, Lub94, Lyn68, Lyn85, LT68, Man85, Mar70b, Mid75, PR86, PS92b, RS70, Rak73, Ric78, RW65, RS68, Roa71, RS88, Sch79a, Sch85b, Sch91, SK73, Slo81, SW98, Spe75, Spe78, Spe79, Sta70, Ste72, Sym66]. integral [TIK67, Ver92, Wal89, WhY88, Won82, dHW73, dHW75, de 95, vdHtR81, vPS96]. integral-interpolation [Kac76]. intégrale [Che79]. Integralgleichungen [Bra60, Jeg68, Kad67, Ama72, Ant72, Bra61, Kil78, Mar70b, Sch79a]. Integrals [Bia90, PK64, AS97b, Bul65a, Bul65b, Bul67, Bul69a, Bul69b, Car79, CR74, Cle66, CP70, CM89, DP85, DD99, Die96, ESS92, EP75, EP76, EV97, Hal60a, Hal60b, HvdR63, Hun72, Hun73, Jel73, Klo99, Kut75, LM69, Mon84, Mon87, ML98, MH72, Mor78, Mor99, Ném65, Nin66, PE72, Pag81, Pat76, RvdS90, SW92, Ste66b, VC92a, Wes75]. integrand [VH93]. Integrands [Dag90, DP88a, LR98, Pet93]. integrate [FP93]. Integration [AS97b, BO90, Bul64, Dag90, For78, GT78, Jel73, Kre72b, Mar65, Pau77, Rut60, Sch63, Sch70c, BLW74, Bet70, BMP81, BP98, Cas78, Cas80b, Cas81b, Cas81a, Cas83, CC60, DP88a, DJ69, Des70, ESS92, FF87, Fri70, Gao89, Gau61, GMF99, GG71, GHR65, Han64, Hsu61, Hsu62, Hun73, Imh63, Jet76, Jet84, JN94, Kut75, LG76a, LG76b, LV97, McL94, MS67, MY79, MP92, MC88a, MC88b, Nev77, NS95b, NS96, NR96, Oga89, Oli71, Pal76, SK96, Sch80a, Sim78, SS78, SS80, Smi65, Str66, Str67, Str68, Tsu73, Ura70, VŽ95, WA72, Wes75, Zad76, dFSS97, Bul64, Jet84, Mar65, Sch63, Sch70c, WL79]. Integrationsformeln [Pal76]. integrator [Cas96]. integrators [CM95, HL97, Sto97]. integro [Ein71, Han94, Lin75, Lub82, Mat76, RW85a, Sch87a, Tan92, RW85a]. integro-differential [Ein71, Han94, Lin75, Lub82, Mat76, RW85a, Sch87a, Tan92]. intégro-différentielle [RW85a]. integrodifferential [DJ62, NW73, Ver88]. Integrodifferentialgleichungen [NW73]. intense [LMGR94]. interest [LR89c]. interface [CZ98, Han82, HS83, KS98, Mey70]. Interior [Ber97, Liu96, BM96, Har78, HP83, Pet93, SW99b, TW68]. intermediate [Gue86, Wei63]. Internal [CC91]. International [Ano83b, Ano84]. interpolant [dV93]. interpolants [AL93, Ber89b, Ber89c, GM83, Gün94, Lub89b, Rei84a, Rei84b, RS86]. interpolated [Ber73]. interpolating [AL68, Bau76, vG72, Hen79, LA88, LA89]. Interpolation [BG76a, Boj93, BS70, Bra74, Dag90, Dör66, DM90, EH70, For78, Kre71, LL90, NR68, Nit69b, NSSS83, Sch61b, Sto61, Tis72, Web91, WB69, Wer72, Wyn60, Žen70, Zur64, Bak70, BSV68, BMP81, BH71b, CMQ82, Cha61, Che79, CD90, CD91, Cla76, Cla78, Cle92, CS66, CS68, CMO91, DP88a, DGM88, DP82, Din83, EM87, Egg86, El 79, ES73, GM82, GH73, Gou77, GMH81, Her70, Her71, JS91, Kac76, KC93, Kre72a, LW99, MS92, Müh79, Müh81, MS99, NR95, OP81, Pow94, RS91, RS95, RO70, Row71, SL86, SA84, SKL88, Sau97, SH95, Str73a, Tsa77, Tsu75, VB92a, Wal97, Wuy73, ŽV95, Arn80, BS70, EH70, Kan73, Nit69a, Sch61b, Tis72, Zur64, Rei89b, Sch73a, Wer72]. Interpolations [Wyn60, Mel72, Wal97]. Interpolations-Algorithmus [Wyn60]. Interpolationskoeffizienten [Str73a]. Interpolationspolynome [NR68]. Interpolationsprozesse [ES73]. interpolatorischer [Eng77, Fie87, For75]. Interpolatory [DD99, KRV74, KE82, AJS90, BS81, Fie87, For75, Kau81, Küt82]. Interpolierenden [Sch73a]. interpolierender [Bau76]. intersection [Pyl67]. Interval [Nic66b, Ale71, AH72, BV87, BV88, Cle71, FM89, Han80, KK97, May85, May86, MY79, MS99, Nit62, Rei88, Sim78]. Intervall [Nit62]. Intervallen [MW60a, MW60b, Bla77a, Bla77b]. Intervallmatrizen [AH72]. Intervallrechnung [Ale71]. Intervals [MW60a, MW60b, BL78, Bla77a, FGM87, Ghi68, Bla77b]. invariance [Ste84]. Invariant [KS73, NS95b, NS96, ACR94, Bey87, BK98, HR94, Lus96, Ste76b, Sto93]. invarianten [KS73]. Inverse [Faß97, GLM91, Ale82b, BKP79, BDG97, BK81, For68a, Geo79, Had68, Hal77, HV99, Joc80, Lue79, Mor74, Pyl67, RK84, Sch78b, Sch70d, SFR93, gS99, VN93, Vog83b, Wil62b, dO70, Dör66]. inverse-positivity [Sch70d]. Inversen [For68a]. inverser [Sch78b]. Inverses [Gau62, Gau63, Gau75a, Hen62, All73, Gau78, NC93, Not98]. Inversion [GO94a, Lew82, Ale82b, FM68, GLM89, GLR88, GMS95, Har62, HJR88a, Jer85, Nat78, Nat79]. Inversmonotonie [Lor77]. invertibility [CHS96]. Invertierung [AH72]. inverting [AH72, GL87]. Investigation [Neu63, BS98a, Neu60]. Investigations [Rei71a]. inviscid [QLV91]. involutory [HC63]. involving [BH60b, Hen75, KR84, Raz69]. Irreducible [Nod71, BNZ72b, Yam66]. irréductibles [HW87]. irregular [CM62, Ran84, Tho90]. irregularly [Sal64]. irrégulières [Tho90]. isolate [Por66]. Isolated [MV68a, MV68b, JS70, Med71, Med73, Ste75b, Tod65, Var64]. isoparametric [Man81b, ŽV95]. Iterated [Kat60, Per67, Per68, Böh81, Fra76, Fra77, Hai78, Ise84, Jel73, Sch93]. Iterates [Hen62, Yam86a]. iteratifs [Gas63]. Iteration [Alb62, BG94, Bit64, Bra65, BD68, RM72, Sch70b, Wal67, Wei64, Zaw91, RM72, Alb61b, AV85, BK81, BS85, DD75, DC74, Dou61, Dou62b, FR88, FDD92, Geo79, HNS95, HV85, JT70, Kio79, Man77, Man78b, Mei97, Mey83, NV80b, Nev90a, Nev90b, Ort86, Rut69, Rut70, Sán75, Sch77a, SV79, Ste69a, Ste70, Ste76b, Ves99, VG88, Wal91, Wil62b, Yao99, Ade69, Sac73]. iteration-variation [FDD92]. Iterationen [Sch64a]. Iterations [Ste94b, FS99, Gar76, Han91b, HS69, Ish84a, Kim98b, Kle65, PS84, Sah85, Sch64a, Ste69b, Ste73b, Wit97, Woź78]. Iterationsbeschleunigung [Bra65]. Iterationsfolge [Alb62]. Iterationsfolgen [Alb61b, Bra65, Ort86]. Iterationsmethode [Mit73]. Iterationsverfahren [Bit64, Bra70, For68a, Gek72, Sch77a, Wal67, Weg78, Ade69, Sac73, Sch70b, Wei64]. Iterative [Atk73, Bos71, BF66a, CRZ99, Cha87, Cha88, Gas63, GV61a, GV61b, GW66, GO88, GeN80a, HB60, JP90, Kin73, MPW66, NP78, SV93, Ale76, BWY90, Bar93, Bea79, Bea86a, BS97a, Bra70, BWW73, CR94, CGR94, Cas77, Cro78, Die87, DPRW93, EMN88, ENR90, Elf80, EGS94, EM91, For68a, FS92, FS94, Gek72, GP78, Gun64a, Gun64b, Gut81, GNV86, Gut89b, Had87, HV64, Heu98, Heu01, Hod67, Kac76, KR84, KW96, Kim98a, KZ94, LRS91, LP95, LR93, LV89, Man85, MS96, Mar69, MS94, Mat93a, Mat93b, McC75, Mis84, Mit73, Nic75, NV83, Pao85, Pao95, Pao98, Par59, Par61, PSS90, Pla96, Pom71, Sch98, Sch87d, Sch88b, Sch79b, Sta97, Tan74, TS98, VW92, Weg78, Weg84, Wol78b, Yam75, Yam76, Kil78, Sch75d, Ypm84]. iterativen [Ale82b]. IV [CSV68c].

J [Ale82b]. J. [BH87]. Jacobi [BW72b, FF94, Grü97, Hop86, Sch61c, AL76, AR85, Ask72, BD77, BW72a, CH60, Dol81, DH93b, EB68, Ebe70, Fer89, GH84, Har91, HvdR63, vK66a, vK66b, KMS99, LS95, MS78, MP85, RH93, Rut66b, Sch61c, Spe71, Ves76a, Ves76b, Ves79a, VW79, VH89, Ves93, Wil62d]. Jacobi-fractions [Spe71]. Jacobi-like [Dol81, Ves76b, Ves79a, VW79].

Jacobi-Matrizen [MS78]. jacobiähnliches [Dol81]. Jacobian [NR68]. Jacobians [MS78]. Jacobischen [NR68]. Jenkins [Hag87]. Jensen [Bea86b]. Johnson [BDT88]. Johnston [Med71]. Jordan [DH89, Läu61, Sti60]. Jordan-Elimination [Läu61]. Jordanization [Buc64]. Josef [FV94]. June [FV94].

Kaczmarz [CEG83]. Kahan [Mei86]. KAM [jS99]. Kantorovic [Bau61]. Kantorovich [And71, Had96, Häu86, Sch60, Tsu99, Yam86a, Yam86b]. Kantorovich-type [Häu86]. Kantorowitsch [Mei63]. Karman [Miy76, Kes79, Rei82]. Kergin [Wal97]. Kernel [ML90, BG76b, BH61b, CK74, CK76, HS80, PW86, Sch79a, Slo76]. Kernels [Sch72a, Atk72, BRS71, CR74, DP85, Die96, Fie87, HS79, LL80, Jet76]. Kernen [Jet76, Sch72a]. Ketten [Her74]. Kettenbruch [Arn80]. Kettenbruch-Algorithmus [Arn80]. Khachian [KP81]. Kikuchi [Kes79]. Kind [Rei90, vPS90, Atk72, BJ85a, BJ86, Bru78, BN81, Bul69a, DMJ86, Egg86, EG95, GP78, HS80, HP75, Hoc79, HvdR63, JS86, KS93, Lau98, LP78, Lor77, Man85, Mid75, Ric78, SW96, Slo81, WA72, Yse83, dHW73, dHW75, vdHtR81, vPS96]. kind-property [Lor77]. kinds [BF66b, Mor99]. kinetic [DLD94]. kinetics [ASS88]. Kirchgraber [NR92]. Kirchhoff [Kil75, Pit88]. Klasse [Cro76, Jeg68, Wei64]. Klaus [GB81]. kleinen [MW60a, MW60b]. kleinsten [Döh82, Läu61]. Kleiser [CS86]. Kleiser-Schumann [CS86]. Knopp [GL86]. knot [KK97, Pit73]. Knoten [Bro78, Pau77]. Knotenpolynomen [Pit73]. Knotenzahl [Möl76]. knots [Boj93, Bra71a, BM92, CH86, GK83, GG86, Gue88, MP96b, Sch78a, Wul73, Sch80a]. Kocan [Sch95a]. Koeffizienten [Bra65, Lam70, Sch82].

Koeffizientenfunktionen [Hav77]. Koerzitivität [Gri70b]. Kogbetliantz [CVV88, Fer89]. Kollokationsmethode [Kad67]. Kollokationsverfahren [Hei89a]. Komplementaritätsprobleme [Smo90]. Kondition [Sau84]. konforme [Men74]. konformen [Gai72, Weg78]. König [GH66, VG88]. Konsistenzordnung [Yse83]. Konstanten [Sik61]. Konstruktion [Sch73a]. Konstruktionsverfahren [Sch71a]. kontinuierlicher [NY83]. konvergente [Ale73]. konvergentes [Dol81, PS82]. Konvergenz [Ale76, AV76, Ale82b, AV86, Ans73, ES73, Gri70b, Gri70a, Kle81b, Mer91, Nat75b, Nit70, Sch61c, Wer78, Ans67, Ans69, EH70, Gri75b, Pal77, vFvH73]. Konvergenzaussagen [Hei89a, Wit77]. konvergenzbeschleunigenden [Sch66a, Sch68b]. Konvergenzbeschleunigung [Alb62, Bra70, Ort86]. Konvergenzkriterien [Ans64, Eng72]. Konvergenzordnung [Hav84, Mül82b, Wag76, Gri75c]. konvergien [Sch75d]. Korovkin [Ser99a]. Korovkin-type [Ser99a]. Korrektur [Fre60]. Korteweg [GA92, dFSS97]. Kreın [GK85]. Kreın-Sobolev [GK85]. Kreiss [GP85]. Kriterien [Het74, Het75]. Kriterium [Nit68]. Kronecker [CDL96, Lub88c]. Kronrod [BR91, Not93a]. Kryloff [Ber59b, Wet60, Wet60]. Kryloff-Bogoliubov [Wet60, Wet60]. Krylov [Ber59b, BS60, Mor94]. Krylow [BS60]. Kubaturformeln [Möl76]. Kufarev [HR79]. Kuramoto [Akr92]. kurze [Jet84]. Kutta [DKS90a, VHS90, ASW93, BZ85, BSW88, CGPM98, Cro79, Cur70, DKS90b, GW86, GMF99, Gug97, Hai82, Hai86, HISS90, HS81, Hun86, HS90, KR79, Kot99, KS91, Lub82, LS70, NW81, QYK99, Rei97, Ren85, SSVH87, Sch95b, Spi86, SW84, Ver88, Ver92, Ver69, Ver77b, Zen86, Zen88, Zen97, Zou97, dHW75, vdH72].

L [BGHS79, Mei63, Rut69, SV67]. L-splines [SV67]. L. [JBC92, Mei63]. Lagrange [AL93, BH92a, Ben99, BS70, CMO91, GM82, MT96, MS92, Pit79, SA84, Sül88]. Lagrange-Interpolation [BS70]. Lagrangescher [NR68]. Lagrangian [Bab73, BL97, KP91, NR68, Sch81b, Sch81c]. Laguerre [ELR92, Gar76, Web81, Zou99]. L’Algorithm [Leb70]. lambda [Lan64a]. lambda-matrices [Lan64a]. Lambert [SS84]. Lambert-McLeod [SS84]. laminated [Lus96]. Lanczos [BRS92, BZS99, Han97, Jia98, Sim98]. Lanczos-type [BZS99]. Landau [Che97, DGP93, GHR97, Lem98]. Landweber [Han91b, HNS95]. Länge [Sau78]. Language [vWMPK69, AKK^{+}68, BBG^{+}60, BBG^{+}63, PS59]. languages [PS59]. Laplace [Bri83, CM62, CM70, GL87, GLR88, GMS95, Gus85, Hof67, KS71, Li86, Li87, NW74, Par59, Tha76, Yse80]. Laplaceoperator [Yse80]. Laplaceschen [NW74]. Laplacian [SS88, Spa93]. Large [Leg66, Sch73c, Woź77b, BK98, CRZ99, Gir89, GT82b, Ho90b, JM87, Jia98, Kle65, LV84, Nat69a, Sad93b, Sad93a, Woź78, Zha96]. largest [Cry67, Döh82]. lattice [KZ74, Rei89b, RS95]. lattices [BP85, RS86]. law [BN90, BL93]. laws [CL96a, CM80, KNR95, LS95, LT98b, MO78, Roh98]. Lax [Ans66, Hed65, MO78, PSS84, Zwa73]. layer [AGS98, BYI86, IU82, Rat93, SSS99, Wen68]. Layered [Bit64]. LCNO [MP95b]. leading [Peñ98, Sad93b]. Least [ADdR89, Bau65b, BG65, Cam93, CP94a, Gan81, Gol65b, GW66, GR70a, GvM91, JO74, PCV96, Sau78, SS79, Sre71, vdS75, Axe87, Bar91b, Bjö88, BD72, BS99c, EL88, FH87, Gir89, Han88, HM86, HM89, Kuh93, Läu61, LH88, PW93, Pom68, Saa86, Sau84, Sch81c, SW93, Spä86, Sto64a, VB92b, VV89, Wei92a].

Least-Squares [ADdR89, PCV96, EL88, FH87, Han88, LH88]. Lebesgue [MS92]. leg [DML83]. Legendre [MQ81]. Lehmer [Lin61]. length [MS85]. less [BL81]. Level [Yse86b, Bre96, HV99, JK91b, Sch98, Yse90]. Levenberg [AO77b, BD72]. Levenberg-like [AO77b]. Levinson [GH95]. Lewy [SSS86]. Liapunov [CW78]. Lidskii [LM99]. Lidskii-Mirsky-Wielandt [LM99]. Lie [GeN80a]. life [HL97]. life-span [HL97]. Like [BD68, AG91, AO77b, BH92b, Bor94, Bus77, CT93, DHP99a, Den68, Dol81, Ley86, NC93, Por78, Ves76b, Ves79a, VW79, Yam87, Ypm84]. Limit [BRR81a, BJ85b, Hac77a, Jac85, JW88, MPP99, NW84, RW85b, Sep72, TW80, Wuy71]. limit-periodic [BJ85b]. Limitation [vF70]. Limited [Sac73, Lou86, Zet70]. limites [Pau81a, Pau81b]. limits [Ksc68, Pat76, Pau81a, Pau81b, Tod88]. Lindelöf [Nev90a, Nev90b]. Line [Ari98, GA87, Voi74, Wal68, AS87, Egg88, EGS94, Jar94, LP95, Par59, Par61, Sch68a, Sch81b, Sch81c, Wer78, vW77b, vW77a]. line-search [Jar94]. linéaire [Hil83, LV75, Maî68, Pau81a, Pau81b, Rob66, RW85a]. linéaires [Cha79, Des61, HW87, Le 80a, Leg66, Tho90, Tro90]. Linear [ABS84, AU61, Alb61a, AL85, Auc92, Aum63, Bar69, Bau65b, BL90, Bit64, Bol84, BMPW66, BG65, Des61, Due59, Fer89, Gol65b, GO88, HB60, Jer70, Kem72, Kin73, KvdD77, KS73, Leg66, Maî68, Mül68, NP78, Nev90a, NS96, OP64, Rob66, SS79, SV93, Sti59, Tan71, Wal67, WR71, Woź77b, dBP77, vdS70a, vdS70b, vdS75, AK99, Alb61b, ACP99, AT79, AO76, Ans63, Ans67, Atk67, BL70a, BL70b, BL76, Bar91b, BY66, Bey78, Bia91, BLR86, Bod68, BH71b, Bra82, BRS92, BZ94, BRZRS98, BZS99, Bro65b, BF66b, CR94, CGR94, CGPM98, Car71, CP94a, Cas80a, Cas83, CS76, Cha73b, Cha79, Cho89, CFH68, CW79a, CG67, DML83, DJ69, Dix82, Dol81, Dör96, DMR91, ENV85, EMN88, Elf80, EGS94, EM91, Fer97, Fis81]. linear [FR88, FQ99, FN91, Gar65, GT72, Gor70, Gov90, GLC89, Gre80, Gri70a, Gun64b, Gut89b, Had87, Han94, Han88, HP75, He94, HL64a, HL65, Hil83, HW87, Ise84, JK91a, JS95, KR84, Kiw94, Kle65, Kle66, Kna85, KK74, KK76, KZ94, Koh87, Koh88, Kub84, LV75, LRS91, LV97, Len89, LM75, LH88, Lub94, Mak92, Man77, MS96, MS94, Mat80, Mat76, McA68, Mei66, MSS67, Mey70, Mil81, Mis84, MC88a, MC88b, MSW91, Mor89, Müh85, NV80a, NN87, Nie85, Nie73a, NV83, Oli67, Oli68a, Oli68b, OP96, Pau81b, Pla96, Por78, Ran79, Red77, Rei65, Rei91, Reu91, Ris74, RL66, Sau73, Sau84, Sch71a, Sch81c, Sch64a, Sch87b, Sch70d, She77, Smo90, Spä86, SW87, Sre69, Sre71, Ste86, Ste61, Ste65, Ste73a]. linear [Sti60, SW99b, Sto93, Sto97, Str64, Tan74, TT87, Tho90, Tom84, Tro76, Tro90, Tsa77, Van79, Ver77a, Vol86, VI86, Wal89, Wer62, Wil62e, Wim74, Wit77, Woź78, Wyn59c, Yam75, Yam76, Zen84b, Zie85, vW77a, Kra68]. linear-quadratic [Kna85]. lineare [Gor70, KK76, Rei65, Smo90, Sti59, vW77a, Cha73b, Kra68]. linearen [Aum63, Bit64, Due59, KS73, Mül68, Sau84, Sch64a, Vol86, Wal67, Wit77]. linearer [Alb61a, Alb61b, Bey78, Bod68, Bro65b, Gri70a, KK74, Red77, Sau73, Sch71a, Tro76, Wer62]. lineares [RL66]. linearization [Kiw87]. Linearly [Lub89a, Hun86, Nie95]. Lines [RT92, AFW93, Mey78, Seg99, Voi79]. Linienmethode [Wal68]. Liouville [Hav77, Nat69b, Nit69b, AP86, Hal77, Hav77, MP95b, Mar90, Nat69b, Nit69b, Pru75, VD91, Wei63]. Liouville-Probleme [Hav77, Nat69b]. Liouville-Randwertproblemen [Nit69b]. Lipschitz [Sac73, BS91, BW95, DRW81, Sac73, Sik84, Tou92]. Lipschitz-beschränkte [Sac73]. Lipschitz-Limited [Sac73]. liquid [PR93]. lissage [Dia80]. LMMs [BMR96]. Lobatto [Kam71, Kot98, Not93b, Sch87c]. Local [ELV91, Fae98, GT82a, MS94, AGS87, Arn84, BS98b, CP94a, Cas82, CD91, Con80, DD73b, Egg86, LMR95, Nic75, Pav94b, Sch81a]. localization [MW99]. Locally [Joh77, EW92, Het74, Het75, Pie70, PS82, Ver84]. locaux [Con80]. Locher [Lev82]. Locking [BS92a, Zha95, CPT98, Cha97, CP94b, Pit92]. locking-free [Cha97]. logarithmic [BB95, CE92, Deu75, Jar94, KS93]. lognormal [Sim78]. loi [BN90]. lokal [Het74, Het75, PS82]. Long [Car71, DP88b, EHG89, PS92a, Wah75b]. Long-range [Car71, DP88b, EHG89]. Look [GH95, BZS99, FZ93, Fre94]. Look-ahead [GH95, BZS99, FZ93, Fre94]. Lossless [SV70]. Lösung [AU61, Alb61b, Ama72, Ans63, Ans64, Ans67, Ant72, Bit64, Gek72, Kle81a, Kle81b, KR82, Lan87, Leh66, MS78, Mül79a, Mül79b, Mül80, Mül82b, NW73, Sau78, Tör63, Vol86, Arn79, Tro76, WH86]. Lösungen [CS59, Dre77, Hüb82, Kil78, Kos86, Koß81, Mar70b, Pet70, Spr76, Tre62, Zhe89, Hof75]. low [Bea86b, Pit88]. low-order [Pit88]. Lower [Ban87, GI88, IK77, Sto68, BG95, Bjö88, Bod68, BS66a, CNV88, Küp76b, Sch75c, VC92a, tR91, Kna71]. lowest [GN87]. LU [MS98]. lumped [Ish81, Ish85, GS93b]. Lyapunov [MS72].

M [Sch95a, MS98]. M-matrices [MS98]. M. [JBC92]. MAC [Por78]. MAC-like [Por78]. Machine [Neu60, Neu61]. Machines [GL65]. Maclaurin [EV97, LM70, Lyn85, ML98, Sch77b]. magnetohydrodynamic [Rap77]. magnetostatic [Per99]. magnetostatics [Žen95]. Mandelbrot [ES92]. manifold [EP99]. manifolds [ACR94, BK98, DH97, FR84, NS95b, NS96, Rhe88, Spä86]. many [Tod88, Tsu72, Tsu73, Tsu75]. map [Jer89]. mapping [DHP99a, Ell79, HP81, HP83, Men74, Men88, PW86, Ste75a, Šve70, Sym66, Sym67, Sym69, Weg84, Weg89, Weg78]. mappings [BHH88, CHS96, KP90, Opf79, Opf80, PPSS97, Stu76, Wei80]. maps [Gut81]. March [Ano83b, GV88]. Marker [Gir76]. Markov [Bar93, Lev82, MS94, Mei97, O’C93, SZ91b, Ste93b]. Marquardt [BD72]. Marshall [Els69, Els69]. mass [Ish81, Ish85]. matching [BC91, CSZ96]. matchings [KvdD77]. materials [HB97, Vog83a]. Math [Aro63a, Bea86a, Ber60, CS92, Hal60a, Heu01, JN82a, Koh88, Lub92b, MC88a, SV77]. Mathematical [BP92, GK64, BS98a, Dör66, Zah77]. Mathematik [Ano83b]. mathematischen [Dör66]. Mathieu [SS66, SG62a, Wag76]. Mathieuschen [SS66, Sch61a, SG62a, SEG62, Wag76]. Mathiew [Sch61a, SEG62]. Matrice [LV74b, CL71, Mai67, Rob66, RM72]. Matrices [Ale73, AR85, Bar69, Bau63, Bau69, Bod68, BMRW68, Bus68, Fre90, FNP82, Gau62, Gau63, Gau75a, Gau75b, GI88, GO94a, Hen62, Kat60, Kau81, Lew82, MW65, MW67, MW68c, Mue66, Nab92, NP78, Nic74, PW70, RB68, Sch70a, Sch73c, Sha82, SZ91a, SW62, TT94, Tyr94, vdS69, Ale82a, ALM97, All70, All73, AL76, Bat90, Ben65, BBdH86, BdP72, Buc64, CGR90, CV84a, DML83, Dah85, DL97, DR75, Des63, DZ71, DV99, Dol81, Dub70, ENR90, EFP98, ER67, Els84b, Fre94, Gas65, Gas71a, Gas71b, Gau78, GW95b, GH84, Gre62, Har82, HR94, Hea66, HJR88a, Hei89b, Her74, Ho90b, HC63, Hou66, HVW70, Huc98, JO72, Kel72, vK66a, Kub84, Laf75, Lan64a, Lan64b, Lew71, LV89, LRZ95, Loi68, Lyn63, Mal99, Man70, MO68]. matrices [MRW70, MPW70, Mat96, May85, MS98, Mes95, MS72, MS73, MV80, MN85, NV80b, OD95, Paa71, PE97, PW84, Pyl67, Rei89a, Ris74, Ros64, RS63, Rut66b, Rut70, Saa86, Sac72, Sad93b, Sar86, Sar90, Sar96, SS76, Sch80b, Son93, Ste76b, Sto68, Sto84, gS91, Šve70, Tay69b, Tee64, VC81, Ves76a, Ves76b, Ves79a, Ves79b, VW79, Vit75, Wag97a, Wag97b, Wei92b, Wer72, Wil62c, Wil72, Wil84, Ye95, Yse89, Zen72, Zie97, dV73, vG80, Rut66a, AH72]. Matricial [Deu70]. Matrix [BD73, BF90, BL80, Bre70, BKP76, BGBMN91, EHM94, ER74, Fie66, Foc65, Gas65, Gas71b, Hig92, KM86, Leh66, Lig90, Mai67, MPW65, MW68d, Nod71, Rob66, Sto62, vF70, Ale82b, MBW67, BH60b, Ber84, BC78, Bra82, BP66, CD79, CL71, CMR86, Deu75, DC99, Dör90, Dry82, Ern96, ESR82, For68a, GK90, GP85, GW95b, HN89, HV64, Har62, HL64b, Hel68, HP92, HUY95, Hod67, Joh77, KvdD77, KM91, LV74b, LE81, Mal99, MRW68, Mat92, MS95a, Mor72, NS64, Osb69, PR69, Raz79, RM72, Rod73, Ruh70, Sch83, Sch68c, Ser99a, Shi78, Shr91, Sim98, Smi67, Ste69a, Ste82, Ste99, Sto64a, gS83, VN93, Ves93, Wil62a, Wil62b, Wil71, Wri92, Yam66, Zen68, Zha96, Zie85, vGS84, Ale82b, Bod68, For68a, FM68, vF70]. matrix-pair [gS83]. Matrizen [Kul68, Ale73, Bod68, Bra71b, Dol81, MS78, Rut66a, Sch70a, Sch73c, Sch80b, Wer72]. Matrizenaufgaben [Leh66]. Matrizennormen [Foc65]. Matrizielle [Bod68]. max [FS99, Lyn63]. maxima [Pie70, Pie72]. Maximal [Nod71, DV97, LMR95, Yse83, Mül82b]. maximale [Yse83, Mül82b]. Maximum [BLLN95, Ste63, TW83, Zen84b, Ein71, NW74, Tao84, Wid66, Žen95]. Maximum-norm [TW83]. Maximumsprinzips [NW74]. Maxwell [CZ99, Hop81, Mon92, RS96]. McLeod [SS84]. mean [CL96b, DD95, DGL99, MH72, Tod90, Wal97, vG80]. mean-field [DGL99]. means [FF87, JT76, Kna85, Run70, Tod88, Tro76, Wer68]. measure [BL70a]. measurements [HM89]. mechanical [FL91, Sch95b]. Mechanics [BB90]. media [GJ89, HN92, VV98]. medium [JK91a]. mehr [Sau63]. Mehrdimensionale [Sch70c]. mehrdimensionalen [Mit73]. mehrdimensionaler [EH70]. mehreren [Rei75]. mehrfacher [Kil78]. Mehrgitterverfahren [Smo90]. mehrparametriger [Had69a]. Mehrschrittverfahren [Ans67, Gri70b]. Mehrstellenformeln [Yse80]. mehrstufige [Gek72]. mehrstufiges [Bit64]. meilleure [Lau67, SA84]. meinen [Ber60]. Mellin [EP95]. membrane [Enn95, Pit92, RSSV97]. memoriam [BH87, CV94]. memory [NS91]. Mengen [Kup71]. Merhschichtverfahren [vW77b]. meromorphic [HV85]. Mersenne [Kar61]. Mersennescher [Kar61]. Merten [Neu63]. Mertensschen [Neu63]. Mesh [PS75, VV98, AGP92, AM78, AL96, BKP79, BO80, BS98b, CGH93, Cle71, DD98, Gol81, JY93, Sta77a, Sta80]. mesh-grading [JY93]. meshes [BX96, CSZ96, DS91, LZ80, Man84, Rum96, SSS99, Sta77b, Ste97, Ste98]. metals [PR93]. Method [Alb61a, AR85, BDY88, MBW67, BD73, Ber59b, Bit64, Bit68, Bit71, BS60, BB90, Bra60, CMG90, Cup81, Dan67, DH93b, Due59, EB68, ES88, GL65, GR90, Had68, Ker66, Kes79, Leh60, Ley90, LH66, Lin61, Mue66, Ort63, RR90, RT92, Sac73, Sau63, SS66, Sch61a, SEG62, SS79, SB90, ST90, SV93, SZ91b, Tan71, Tör63, Uhl62, VHS90, Wal67, Wal68, Woź77b, Ziv83, Zur63, vFvH73, vPS90, AGP92, AKP95, AC92, AK99, Akr88, AH72, Ale82a, ABFJ97, AM78, Alo96, AL76, And71, ABF99, AL68, AR68, Ans63, Ans67, Atk73, Att77, AL86b, Axe87, ASW80, AHS84, ALS90, AD64, Bab71, Bab73, BD81, BS89, Bad89, BK87, BB92, BC93, BYI86, BE86, Bar71, BZ85, Ben99, BMX98, BP79, BH61b, BH65, Ber91]. method [BP92, Ber84, Ber97, Bia91, BB82a, BB82b, Bjö88, Bla72, BC78, BG78, BR88, BMP81, Boh80, Bör89, BD96, BS98a, Bra81a, Bra73, Bro85, BHH88, BEHH75, Cai91b, CGK92, CGR94, CS86, Cas85, CEG83, Cha74, CL75, Cha99, CT93, CC95, CG59, CŠ79, CMR93, Chr70a, Chr70b, Chu83, CW71, CC85, Cle66, Cle71, CC60, CM80, CP70, CW79b, Cro80, hDxY99, DP72, Dan73, De 68, DD95, DK83, DN89, DC74, Deu74, DPR76, Die87, DH80, Dob81, Döh82, Dol81, Dör70, Dou60, DG86, DHP99b, Dre77, Dry82, Dry84, Dun88, Ebe70, EFM89, EG95, EJSS97, Ern96, EJ81, Eyd84, Eyd85, Fag81b, Fag81a, FMMS99, FŽ88, FK87, FR88, For68a, For68b, Fra76, Fra77, FR86, FN91, Fro88a, GT84, Geo79, Gin63, Gir76, GR79]. method [GLR88, Goë68, Gol62, Gol65a, Gol81, GG71, GHT94, GM66b, GeN80b, Gri80, GB86a, GB86b, GB86c, Gun64b, Guo88, yGL93, GC96, Gus73, HC92, HR89, Hac89, HN89, Hac92, Hal75, HHM76, Han82, Han91a, Han64, Han88, Han97, HP75, Hào94, Har82, HS83, Har78, HH76, Häu86, HL98, Hed65, Hei89a, HW78, HJR88b, Her91, Heu98, HMS99, Heu01, HL99, Hoc81, HM89, HR79, Hop81, Hor75, HP81, HP83, HS70, HHZ94, IM94, Imh63, JBC92, JLR70, Jen91, Jer89, JS86, JY93, Joh73, Kac76, Kad67, Kan98, KR97, Kea79, KS89, vK66a, vK66b, Kes84, KKL97, Kin74, Kir88, Kiw84, Kiw87, Kiw94, Kje61, Kno89, KW95, Kre90, Kri83, KN84, LV74b, Lam80, Lam91, Lan66, LT98a, LV84, LR93, Lev89, Ley86, Li86, Li87, LS87, Li90, Li95, LL80]. method [Lin75, LE81, Lor80, LH88, LM98, Lub92a, LR73, MT96, MS94, MGR87, MBL90, Mer91, MS72, MS73, Mey78, Mey84b, MY79, Mig76, Mil85, Mit73, Miy76, MP85, MK80, Mon92, Moo68, MO91, MSW92, Nak85, Nat75a, Neg90, NN87, Nie85, Nie73b, Nie89, Noc89, Oga89, OD95, Osw92, Pal99, PS92a, Pau81a, Pau81b, Pav94a, Pav94b, PB87, Pei91, PV72, Pie70, Pie72, Pit79, PSS90, Pól68, PS82, Pot82, Prö91, PS92b, Pyt02, QLV91, RS76, RK84, Rei89a, Ren79, Reu95, RH93, Rie97, RW65, Rob72, Rod73, Rok72, Rom85, Rua88, RW88, RS88, Rus84, RS72, RS63, Rut66b, Rut69, Rut70, Rya67, SV90, Sad93b, Sal61, SZ61, SS84, SKL88, Sar86, Sar90, Sar96, SK96, Sch71a, Sch77a, Sch81b]. method [Sch81c, Sch68b, SV70, SH95, Sch91, Sch86, Sch61c, ST73, Sch71b, Sch78c, Sch74b, Seg99, SW96, Shi78, Slo88, Smi65, Smo90, Sou69, Spa93, Spe75, Ste95, Ste73b, Ste75b, Ste90b, SB69, Str72, SO86, Sül88, Sun95, Sym66, Tau76, TL91, TS98, Tor97, TS91, TK64, Ura67, Ura70, VŽ95, Ves76a, VW79, Vog83a, Vog83b, Voi74, Voi79, VI86, Wag76, Wag80, WKW97, Wah75b, WFS95, Web94, Weg78, Weg84, Weg89, WA72, Wer68, Wer79, Wer82, Wil62a, Wil62c, Woo68, Yam66, Yam86b, You71, Žen90, Žen95, Zet70, Zha89, Zha95, ZL94, Zlá68, Zou99, d’H93a, dHH87, vW77a, VV98, Ale76, AV76, AV86, Nit76, Bra77]. Methode [Dre77, AG88, Att77, Bre75b, CL75, Con84, Cro80, JLR70, Kes79, LV84, LH66, Mig76, MK80, Pau81a, Pau81b, Rom85, Tor97, LV74b]. Methoden [Cro77]. méthodes [Cro79, Del73, Gue94, Par73, SP72, Tro90, Aus71, BK96, Hai77, Le 80a]. Methods [ABS84, Ade69, Ans69, BG95, BV68, Bol84, Bör90, BF66a, Bre75b, Col61, Cro79, Fil64, Fre90, GV61a, GV61b, Gol65b, GO88, HB59, HB60, Kin73, Kre63, Lie90, MT90, Mei66, NP78, Nit68, Rei90, Ver69, Wei64, Zha92a, AGS87, AS97a, AFW93, AO93, AM94a, ADK91, ACM99, Alb78, Alb85, ALM97, AM94b, And69, AL96, ADG84, ASW93, AR93, AM88, Aus71, Axe77, AL86a, AV89, AC94, ASS88, BO80, BMPS99, BJ85a, BJ86, Bak82, BR81, BWY90, BH92a, Bar93, Bar91b, BW93, Bea79, Bea86a, BM87, Bel83, BJZ88, Ben64, BS97a, BK96, BS95, Bey87, BD97, BS96, Böh81, BY93, BF70, Bou90, Bra70, BHR78, BD88, BSW88, BN81, BS66a, BS66b, BS88, Bus77, BC92, CG88a, CGPM98, CHL85, CPT98, CR93, Cas78]. methods [Cas83, CH60, CP76, Cha87, Cha88, CSZ96, CG88b, CS90, CS92, CS76, CK82, Che97, CZ98, CEL98, CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, CGT94, CW79a, CG67, Cro77, CS98, DML83, DL97, DHP99a, Dej66a, Dej66b, DKS90a, DKS90b, Del73, DG98, Den68, Deu79, Deu83, DHZ87, DMJ86, Dou62a, DJ62, DD73a, DD73b, DPRW93, DSW96, Dut99, Egg84, Egg92, ELP90, ENV85, EMN88, Eie89, ENR90, EN88, ESS92, Elf80, EGS94, Els84a, Els89, EM91, EL88, EW92, FJ75, FF97, FGS95, Fer97, Fer89, FP93, Fie87, FGN81, Föß89, FHL86a, FHL86b, Fro88b, Fro89, FS92, FS94, Fun88, Gab84, Gar96, GN87, Gas96, Gek72, Gek78, Gek82, GW86, GMS95, GMF99, GH73, GM66a, GM68, GM92, GR86, Gri83, GLL91, GKR86]. methods [Gue94, GQ98, Gug97, GP83, GS98c, Gut81, GNV86, Gut89b, Hac77b, HM83, Had71, Had87, HV64, HW76, Hai77, Hai79, Hai80, Hai81, Hai82, HL84, Hai86, HO90a, HISS90, Hai99, HP77, Har91, Heb88, Hei89b, Her71, HH99, Hoc79, Hod67, Hop86, HL60, Hul71, HS81, Hun86, HS90, iS91, IJ87, Jac79, Jac82, Jac84, Jay93, JS95, Jel73, JK78, Jel79, Jia98, JMT97, Joh76, JM78, Kal98, KR84, KR79, KW81, Kau89, Kem72, KS92, KV85, KW96, Kim98b, Kin72, Kio79, Kir86, Kle66, Kle81a, Kle81b, KL90, Kor94, Kor96, Koß81, Kot98, Kot99, LG76a, LG76b, LS66, LRS91, Le 80a, LP95, Lee61, Len89, Les73, LV89, Lin93, Liu95, Liu96, LBF98, LM91, Lu97, Lub89a, Lub91, Lub92b, Lyc72, LRT64, LM69, MP96a]. methods [MNS74, Mak92, Man85, Man78a, Man81a, Man83, MS96, MQ89, Mat93a, Mat93b, Mat96, Mat90, Mat76, MP96b, Mel99, Mey83, Mie79, Mis99, Mis84, MF64, MG85, Mor94, MMS92, MS95b, Müh85, NC93, NN97, NV83, NS95b, NS96, Nit69b, NW79, NW81, Not98, Nov86, O’R84, Oli71, OO88, OSS90, Pao85, Pao95, Pao98, PW86, Par73, Par59, Par61, Par88, Pat76, Pav94a, Pav94b, PCV96, PPV69, PCT95, PV90, Pla96, Pom71, PQS98, PR94, PR86, QYK99, Rai69, Ran79, Rat93, Red73, Rei97, Rei65, Ren85, RRS89, Reu91, Rit80, RF94, Roc88, RKL89, Rua85, Sad93a, SP72, Sar89, Sar97, Saw82, Sch85b, SHH85, Sch95b, Sch99, Sch87d, Sch88b, Sch79b, STR76, Sch97, Sch74a, SW92, Ser99b, Sha68, Sha84, She92, She94, Sim98, Slo76]. methods [Slo81, SW98, SW99a, Smi65, SG78, Spi66, Spi71, Spi86, Squ69, Sta81, Sta77a, Sta80, Sta97, Ste86, SS96, Ste61, Ste78b, Ste73a, Sto84, Sto93, Str64, SW84, SWB91, Tao84, Tay69b, TW83, Tro90, Tur78, VW92, Ver88, Ver92, VC92b, Ver77a, Ver77b, Ver84, Vos75, Wal89, Wan89, WhY88, Wer68, Wit77, Wit78b, Wit78c, Wit89, Wit90, Wol78b, Won92a, Won92b, XS92, Yam87, Ypm84, Yse97a, Zen86, Zen88, Zen97, ZH93, Zla78, Zla81, Zul88, Zwa73, dHW75, de 95, tMS86, vV78, vW77b, Gai72, Gri70b, Had72b, Nat77, Sch79a, Sch66a, Sch78b, Wit78a]. Metric [Ans73, BF90, PS66, Ale71, Brä81b, GT72, GT82b, Pie70, Rit80, Wer78]. Metrik [PS66, Wer78]. metrischen [Ans73]. MHD [Rap84]. micro [HS97]. micro-structures [HS97]. microstructure [Lus96]. microstructures [Chi99, Li97]. mid [BD83]. mid-point [BD83]. Midpoint [Cha92, AF89, SB83]. mild [DJ69, Rok72]. Mildly [GP65, Par65b, BMPS99, Car71, Dou61, Dou62b, Gun64b, Mey79, NW76, Reu88a]. Miller [MvdS76, Zah77]. minces [DN86, Tro90]. Mindlin [Pei91, Pit88]. minima [Con80]. Minimal [CKSS97, VL73, FN91, GG86, Gro89, Hin96, Kac84, Koc95, KP81, Sau97, Sch78a, Sch95a, WFS95, Wil62e]. minimal-error [Kac84]. minimalem [Lam70]. minimaler [Möl76, Sau78]. minimax [Jia94, Peh95, Sto63]. Minimieren [Zet70]. Minimierungsprobleme [Kle81a, Kle81b]. minimisation [Aus71, BS77].

Minimization [Shi82, Brä81b, CL97, CGT94, Dun88, FG89, GA87, Gol62, GP67, Kiw87, Kle81a, Kle81b, Kna85, Krä89, Sch97, Slo80, Sto77, Zet70]. minimize [Sre69, Sre71]. minimizers [BK87, Li95, Li97, Neg90]. Minimizing [EHM94, GK64, Ari98, DS91, Jar94, Kiw84]. Minimum [And88, Bar65, Bar68, Lam70, Pin75, Ste78a, vG80, Möl76]. Minimum-point [Möl76]. minmax [Law64]. minorantes [Rob66]. Minorants [Rob66]. Mirsky [LM99]. mittels [BS70, Kar61]. Mixed [AR93, BHR78, BDDF87, Che97, Gas96, Hua90, MP95a, Néd80, OP96, SSS99, SS96, Ver87b, AABM98, AGP92, AM94a, Alo96, ADG84, AF88b, AHS84, Ben62, BS96, BL97, Bra94, BDM85, BD88, BCS96, CD98, CEL98, Dob81, DG86, DPRW93, Dur90, El 84, EJSS97, EK99, EW92, FG77, FGN81, FM92a, FM92b, FHLM88, GN87, GR79, GP79, HH76, HH99, Ish79, JBC92, Joh73, KS71, LMW84, Le 80b, Le 82, Mak92, Mat93a, Mat93b, Miy76, Mor89, Néd82, Néd86, PCV96, Per99, Red88, Rei82, SS88, Ste86, Ste88, Tod88, VW92, Voi79, Wan89, AG88, Con84]. mixed-variables [Ben62]. mixte [AG88, FM92a, FM92b]. mixtes [AABM98, Con84, El 84, Néd82]. Miyoshi [OP96]. mode [Bou90]. Model [Com59, Arn81, BS91, BB97, CDV99, CCCD93, Che97, DGP93, DGL99, Edd93, GHR97, Had71, HH76, MBL90, OP96, Wah75b, Yse97b, Yse99a]. Modeling [Cia90, Eng96, Sch96]. modelling [BMR96]. models [Ain98, Cha87, Cha88, CGT94, RS96, Sch96, Zha92b]. Modification [BNS78, Her70, Yan94]. Modifications [Dan67, Jac85, Kub84]. Modified [GG90, Mag92, MW68c, Pat93, Ste94a, BJ89, Ben65, CR94, CGR94, Del73, Deu74, DPR76, DHP99b, ESR82, FS70, Hag87, HHM76, HI71, Mey84b, SD72, TL91, Pot78]. modifiées [Del73]. modifizierte [Mey84b, Pot78]. modules [Gai72, PS92a]. moduli [BNvW89]. Moduls [Gai72]. modulus [MH72, Mor78, Mor99, Rei84a]. Möglichkeit [Bra70]. Möglichkeiten [Häm65]. Molecular [FDD92]. Möller [VC92a]. mollification [Hào94, HA98]. Moment [FGM87, Gre82, LM91]. Moment-preserving [FGM87]. Moments [BH60a, GG90, Hou68, CGR94, Gau86, SD72]. momentum [FDD92, LG76a, LG76b]. Monde [LM66]. Monitoring [Bus71, ER74]. monosplines [BL74, Gro89, Str84, Bra74]. Monotone [Alb62, Bar72, Bel83, CSV69, Ish84a, KR84, Kor94, Kor96, Pao85, AV85, Cha74, Cho89, DDM92, EM87, Egg88, El 84, Fro88a, Kos86, LMR95, Lor77, Pao95, Pao98, Sac73, Sch64a, Sch72b, SW99b, Žen95, Alb61b, Bra65, Sto70]. monotonen [Alb62]. monotoner [Kos86]. Monotonic [Alb61b, BSW61, El 84, MP99, SL86]. Monotonically [Sch79b, Gal89]. Monotonicity [Spi86, Str84, Str73b, DE96, FP66, För93, Kra86, MVW79, Nik92, Ste68, Utr85, vdGK86, Hof75]. Monotonicty [Col61]. Monotonie [Loc80, Bra78, Col61, Sto70]. Monotonie-Eigenschaften [Sto70]. Monotoniesätze [Hof75]. Monotonity [Fia90]. Monotony [Küt82, Loc80, Bra78]. Monte [CdM70, CP70, Gir89, Han64, Ksc68, Oga89, TK64, TIK67]. Moore [Lue79]. Morley [PB87]. mortar [AKP95, Woh99, Ben99]. Moser [Hal75]. motion [CL96b, FL91, GK94, LG76a, LG76b]. mountain [CMR93]. moving [TS91]. MR [Bea86a, CS92, JN82a, Koh88, Lub92b, MC88a]. Muhlbach [AS97a]. Multi [Cia90, Hop86, Hua90, Kir86, Par61, Wit89, Yse86b, BB97, CH89, HM83, Hac89, Hac92, Hal60a, Hal60b, Han80, Hee68, JK91b, Rhe88, Sch98, Ste93a, Ura67, Wam94, Wit90, Yse90, d’H93a, vW77b]. multi-component [BB97]. multi-dimensional [CH89, Hal60a, Hal60b, Han80, Rhe88]. Multi-Grid [Hua90, Hop86, Wit89, HM83, Hac89, Ste93a, Wit90]. multi-homogeneous [Wam94]. Multi-Level [Yse86b, Yse90]. Multi-line [Par61, vW77b]. multi-point [Ura67]. multi-sector [Hee68]. Multi-step [Kir86]. Multi-Structures [Cia90, d’H93a]. multi-time-level [JK91b]. multibody [SFR93]. multicommodity [SD94]. multiderivative [JK78, Jel79]. Multidimensional [EH70, Ber72, BMP81, CC95, CP70, LLS98, Neg90, Sch91, SW93, ST89, Sch70c]. multidomain [SV90]. Multigrid [BDY88, BB90, Sch85b, Sch99, vPS90, BB92, BD96, Bra81a, BD99, Heb88, Kor94, Kor96, Man81b, Par88, PB87, Pei91, Reu88a, Reu91, Reu95, Smo90, WKW97, Web94, Zha89]. Multilevel [BW99, BD97, DK92, DSW96, HH99, Kin92, VW92, Zha92a, AV89, BY93, ELV91, HR89, LN99, Man85, Not98, Osw92, Pei85, Pfl98, Reu88b, Rie97, Yse86a]. multiparameter [Mül82a]. multipas [Cro80, Le 80a]. multipass [Cro80, Le 80a]. Multiple [BL76, SZ91a, SSBB88, CR94, CCK95, Cuy83a, DL97, Dan66, Deá80, Deu74, Deu79, DLO^{+}77, GK83, Kil78, Kio79, Lor80, Mal99, MBL90, MP96b, MS95a, Ral66, gS92, gS95b]. Multiplicability [Foc65]. Multiplication [BL80, Bre70, HN89]. multiplicative [BD97, EHG89, GO95, Lan72, LM99]. multiplicity [HP76b, Van82]. multipliers [Bab73, BH92a, Ben99, MT96, Pit79]. Multiplikativität [Foc65]. multiply [DHP99a, Ell79]. Multipoint [CW71, Ste94b, Wal98, BF70, Jer70]. Multipole [Lem98]. multisplitting [ALM97, Els89, Fro89]. multisplittings [FM89]. multistage [JN86]. Multistep [Lie90, ACM99, Alb78, Ans67, Cas83, CS76, DML83, Dej66a, Dej66b, EN88, ESS92, FP93, Gri70b, Gri83, Hai99, Hoc79, Jac79, Jel73, JK78, Jel79, JN86, Kem72, KL90, LS66, Len89, Lub91, Lub92b, Lub94, Lyc72, Mat76, NS96, Sch95b, Sch97, Sha67, Ver77a]. multistep-methods [Gri83]. multivariable [BL81]. Multivariate [CD90, MS95a, Wer86, Cuy82, Cuy83b, CV84b, CDL96, Gut87, Won84]. Munich [Ano83b]. MUSCL [CL96a, LS95]. my [Ber60, Dou62b].

nach [BS60, Bra60, Fel61, Kad67, Kre72b, Läu61, Sch79a]. Nachtrag [Fre60]. Näherung [Ale82b, Mit73]. Näherungen [Kos85, Men74]. Näherungslösungen [Leh67, Leh61]. Näherungsverfahren [Nit70]. näherungsweisen [AU61]. nährungswise [Ama72]. narrow [ŽV95]. Natural [Leh61, Utr83, Ver88, Zen88, BH92a]. natürliche [Leh61]. Navier [AGP92, AM94a, AYB98, BMM87, Cam70, CMQ84, CMG90, FB89, Gir76, GR79, Gue94, GQ98, GP83, HL98, JLR70, Le 80b, LO98, Pir82, Por78, She92, Sül88, SSBB88, TL91, Ver87a, Ver91a, Wit89, Wit90]. Near [ST90, Hed65, KO98, Rei88, Tho90, Tho71, Vas88]. near-field [KO98]. near-stability [Hed65]. nearby [Kri64]. Nearest [Dem87]. Nearly [SZ91b, Bar93, FHLM88, Ste93b, Vog83a]. nearness [Mit73]. Nebenbedingungen [Hof70, KK76, Cro77, Het75, Hof69, Sau84]. nebencharacteristics [Wer68]. Nebencharakteristiken [Sau63]. Necessarily [Sac73]. Necessary [Sha82]. Need [Nic66b]. négatifs [DGB80]. Negative [Nod71, HL64b, DGB80]. neighborhood [AD93]. neighbourhood [Mid75, ZG78]. neighbouring [Wil84]. Nennerpolynoms [Lam70]. Nested [MV69, CR89, GT87, LRS91, Rom85]. network [SD94]. networks [Eng96]. neuen [Lod70]. neues [Döh82]. Neumann [BE87, Bör89, Gek74, Han94, Kat60, KS71, MG85, Ple76, Shi78, VV84, Zou97]. neutral [BJZ88, Hor75, Jac82, Liu95, QYK99]. Nevanlinna [AY83, KC93]. Neville [Müh78]. Newman [CNV88, tR91]. Newton [AH72, Bra78, Kle81a, Kle81b, Mey84b, Wit78b, Wit78c, AM78, AR68, BR81, BK81, Boh80, Böh81, Bra77, BD68, Bus77, CG59, Col61, CV84b, Dan73, Den68, Deu74, Dun88, FN98, Fro88a, Gol65a, Gri80, GT82a, GR70b, Gut89a, Had68, Hal75, Häu86, Jer85, Jer89, Kal98, KS89, Kle81a, Kle81b, Kno89, LW82, Lan66, Lub89b, Mey84b, Mie79, MVW79, Moo68, MO91, Müh79, NC93, PP80, Ptá76, Ral66, RB68, Rob72, Rok72, Sch85a, SHH85, Sch71b, Sch74a, SG78, Tsa77, Yam86a, Yam86b, Yam87, Ypm84, Bra77]. Newton-interpolation-algorithm [Müh79]. Newton-Like [BD68, Bus77, Den68, NC93, Yam87]. Newton-like-iterative [Ypm84]. Newton-method [Bra77]. Newton-Moser [Hal75]. Newton-type [Mie79, SHH85]. Newton-Verfahren [AH72, Mey84b, Wit78b, Wit78c, Bra77, Had68]. Newtonian [AG88, BN90, BL93, BL94, Raz69]. newtoniens [AG88, BN90]. Newtonschen [Col61]. NGP [QYK99]. NGP-stability [QYK99]. nicht [Sac73, Sch80a, Wer78]. nicht-äquidistanten [Sch80a]. nicht-exakter [Wer78]. Nichtäquidistante [EN80]. nichtkompaktem [Sto70]. nichtkompakter [Gri85]. nichtlineare [AV86, Bra70, GKR86, SV70, Wit78c, Boh67, Hof70]. nichtlinearen [Arn79, Boh69, Cro77, För67, Hof69, Sto64b, Vos75, Wal68, Wer69]. nichtlinearer [Ama72, Bau76, Cro76, MS78, Pet70, Red77, Saw82]. nichtnegativer [Sch80b, Yse83]. Nichtnegativitäts [Gor70]. Nichtnegativitäts-und [Gor70]. nichtsymmetrische [AB85]. nichttrivialen [Bra71b, Sch70a]. no [Bea86a, CS92, JN82a, Koh88, Lub92b, MC88a, SV77]. nodal [HJR88b, dV93]. Nodes [Dag90, Bro78, CC85, DP88a, Dry94, EFP98, Fri70, MS92, Pau77, SSS99, TS91]. nœuds [GG86]. noise [GK66]. noisy [AY93, CW79b, Gir89, HdH85, Utr85, Wah75a]. nombre [GG86]. Non [Fre90, Hen62, Nod71, Str64, Ver91b, AK99, AG88, AB74, Atk67, Aus71, ASS88, Ban87, BL70a, BL70b, BS91, BB97, BM92, CGPM98, Car71, CSZ96, Cha73b, Cho89, DD95, Del80, DM96, EGS94, FGS95, Gun64b, HR94, HL64b, Hsu62, vK66a, Klo99, Le 80a, LL80, MSS67, Mis99, Mis84, MSW91, Pau81a, Pau81b, Ped96, PCV96, Rap77, SA84, Sar97, Sch80a, Ste65, Ste68, VI86, Wyn59c, BL94, BW95, CD79, FN91, GH95, Ste76b]. non- [Ste65]. non-compact [Rap77]. Non-conforming [Ver91b, Sar97]. non-convex [Ped96]. non-definite [Mis99]. non-differentiable [Aus71, ASS88]. non-distinct [vK66a]. non-equidistant [Sch80a]. non-exact [AB74]. Non-Hermitian [Fre90, CD79, FN91, GH95, Ste76b]. non-linéaire [Pau81a, Pau81b]. Non-linear [Str64, AK99, BL70a, BL70b, Car71, Cho89, Gun64b, MSS67, MSW91, Pau81b, VI86, Wyn59c]. non-Lipschitz [BS91, BW95]. non-matching [CSZ96]. non-monotonicity [Ste68]. Non-Negative [Nod71, HL64b]. non-Newtonian [AG88, BL94]. Non-Normal [Hen62]. non-parameterized [DM96]. non-parametric [DD95]. non-periodic [Hsu62]. non-quasiconvex [Klo99]. non-self-adjoint [HR94]. non-selfadjoint [PCV96]. non-smooth [Ban87, BB97, LL80]. non-symmetric [Mis84]. non-symmetrizable [EGS94]. non-uniform [BM92, FGS95]. non-unique [Atk67]. Noncompact [Sto70, Gri85]. nonconformal [Tro90]. nonconformes [Tro90]. Nonconforming [CPT98, JMT97, Ran79, Bre96, ByS99, CL94, EJ81, For85, GN87, LZ80, Li86, Li87, Liu96, Osw92, SJX97, Won92a, Won92b, Won93]. nonconvergent [Del80]. nonconvex [Chi91, YN93]. nondefective [gS92, gS95b]. nondifferentiable [Kiw87, Rap84]. nondominant [Lev89]. nonhomogeneous [BS89]. noninvertible [KR84]. Nonlinear [AM94a, Boh69, BK84, CT93, Hof69, MT90, MSW88, OSS90, Tsu75, VC92b, Wal68, Wer69, Woź77a, Žen89, AGS87, ADK91, AH72, AV86, AM78, Ama72, AO77a, AO77b, Arn79, Auz87, AC94, ALS90, BMPS99, BBG98, Bau76, Bit70, Bla72, BLR89, Bos71, Bra70, BRR80, BRR81a, BRR81b, BG71, BD72, Bus77, Cam93, CM95, CG95, CGH93, CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, CL96b, Cro76, Cro77, Dan73, DC74, Deu74, DPR76, DMJ86, Dou61, Dou62b, DD73a, Dun88, ES90, EKN92, Eyd84, Eyd85, FŽ87, FŽ88, FN98, För67, Fro89, GH92, Geo79, GP65, GeN80a, GeN80b, GKR86, Gut81, GNV86, Gut89b, Hac79, HR89, HNS95, HL98, Hoc79, Hoc81, Ish84a, ID86, JM87, nJyH99, JO80, Kal98, KR84, Kan98, Kar68, KR97]. nonlinear [KS92, Kin72, KM98, Le 80a, Le 82, Lub91, Lub92b, LR73, MO78, Man81a, Man83, McC75, Meh67, MS78, Mey79, Mey78, Mey84a, Mey83, Mit78, Mit80, Mit81, Moc74, Moo68, MSW92, Nak85, Neg90, Nii85, Noc89, NW76, OA75, Osb67, Osb68, Pao87, Pao95, Par65b, Pau81a, Pau81b, Per67, Per68, PPV69, Pet70, PR94, Rap84, Red77, Reu88a, Ric74a, Rob72, RW88, Sah85, Saw82, Sch93, Sch98, Şch73b, Şch75a, Sch81b, Sch81c, Sch94a, SV70, Sch79b, Sch71b, Sch72b, Seg99, Seo95, Ser99b, Sey79a, Sha68, Sik84, Söd86, SG78, Spr80, Sto63, Sto64b, Sül88, Tsu88, Tsu99, TK64, Voi74, Voi79, Vos75, Wer82, Wit78c, YM86, Žen90, Zen97, ZH93, Boh67, Hof70]. nonlinearities [BW95, CD73]. nonlinearity [BS91, BS99a]. nonlinearly [NMC89, Wol78a]. nonmonotone [GLL91].

Nonnegative [NP78, OO98, MN85, Sch80b, Son93, Yam66, Yse83]. nonnegativity [Mav77, Gor70]. nonoscillatory [LT98b]. nonoverlapping [NN97]. nonperiodic [BMM87]. nonselfadjoint [Cha73b]. nonsimple [RP85]. Nonsingular [ZH93, BRR80, DP72]. Nonsmooth [Cho90, BL86a, CEL98, Dut99, Hel87, RS88]. Nonstandard [Sar97]. nonstationary [Dob81, GL99, Had87]. nonstiff [BSW88]. Nonsymmetric [SV93, AB85, BDG97, CGR94, ENV85, Ho90b, KS99, Man77, Man78b, MS96, Sta97]. Nontrivial [Sch70a, Zen72]. nonuniform [Gol81, EN80]. Norm [EB68, Gas65, Gau75a, Lod70, Mai67, Nat71, Sch83, Ban87, Bar65, Bar68, BL94, Bau62, BY93, CD85, Dav62, Des63, Ebe70, Gro89, Hig92, Mal99, Mei86, Not90, Not93b, Owe77, Pin75, SS98, Sre69, Sre71, Ste78a, Tao84, TW83, Wid66]. norm-preserving [Mei86]. Normal [Hen62, Bar84, Bat90, CL71, Deá80, dV73]. normale [CL71]. normalen [Bar84]. Normalität [Bra74]. normality [Bra74, CDL96, Lub88c]. normalizable [Hou66]. normalized [VC94]. Norme [Mai67, Lau67, Tao84]. Normed [Lau67, SW62, GT72, LPDT70, Sch78c]. Normen [Nit70]. Normes [RM72, Gas63, Gas65, Gas71b, Maî68]. normierten [Sch78c]. Norms [BF60, BSW61, Foc65, Gas63, Gas71b, KV72, Kin73, Maî68, SS62, Sto62, VL73, AO76, AO77a, AO77b, BO80, Bac70, BH92a, Bar72, BH61a, Deu70, Deu75, FS99, Gri67, Joh77, Kri64, Lan72, Sto64a, RM72, Nit70]. Not-Necessarily [Sac73]. Note [GW66, Hsu61, Mig76, Nod71, Sti60, VH89, Wil62d, Wil72, Wyn63, dO70, vV77, AB84, Bak82, Ban92, Cas77, Cas80a, CGT94, Dij59, Dub70, Els84b, Her91, HVW70, HS81, Jet84, KL90, Lan72, LV89, Med71, Med73, Mey79, MN85, NBE97, Osb69, Osb65, Sch97, Sep72, Ves99, Wul73]. notice [JHW77]. notwendig [Sac73]. Notwendigkeit [Nic66b]. null [BHK^{+}85, Pot89, Ste84]. Nullstellen [BS70, Bra77, Dir69, Ker66]. Nullstellenbestimmung [Wer79]. Nullstelleneinschliessung [AH72]. Num [Aro63a, Ber60, Hal60a]. Number [EHM94, GI88, Bad89, Bra81a, Bre99, Coh60, DP88b, EHG89, Fri70, GG86, Hüb82, Kem68, Kle65, KS83, MP96a, Mar62, Mar70a, Saa86, Sah85, Sch78a, Šve70, YK88, Zhe89]. Numbers [AR85, Dem87, Kar61, Lin61, Lod70, SZ91a, vdS69, Ask72, Bit91, CdM70, DA71, Fis61, GT78, He97, Kim96, Nie85, OD95, RDM62, Rya67, Smi67, gS92, gS95b, Won93]. Numer [Bea86a, CS92, Heu01, JN82a, Koh88, Lub92b, MC88a, SV77]. numerator [Lam70]. Numerical [Agu89, Alb61a, AU62, Alb85, ACP99, AB87, AY83, Alt80, AB74, Ano83b, Ano84, AS87, Arn84, ASS88, BO90, Bao96, Bar69, BZ85, BH65, Bet70, BK98, BL86b, BLR89, BG78, BMR96, BP98, Bol84, Bra60, Bra61, Bro85, Bul65a, Bul65b, BS66b, Bul67, BS67, Bul69b, BGBMN91, CDV99, Car99, CR74, Chi91, CCK95, CM70, Chr70a, Chr70b, CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, Coh59, CE92, DL97, DJ69, DHP99a, DN89, DJ62, Ein71, FG73, FG89, FF87, FL91, Gau61, GK85, Gol65b, GP78, Gue88, HM86, HS91, Her71, HL60, Ish85, LV74a, Lan77, LM66, MPP99, Mar65, Mas87, Mel95, Mil65b, ML90, Nab92, Nat78, Nic66a, Nie80, OP81, Pao87, PR93, Rak73, RP85, Rou87, Rut60, RN98, Sch86, Sch84]. Numerical [Sey79b, Sey79a, SK73, Sim78, Sta70, Ste72, Ste66c, Sym67, Tör63, Tsu72, Tsu73, Ura67, VV98, Web81, Wer68, Woź77b, Woź77a, YN93, Zaw91, dJ77, vD73, AM78, AG91, And69, Ans64, AM88, Atk72, Atk73, BK87, Ban92, BS59, BR91, Ben64, Bit91, BZ82, Bus71, CPT98, Car71, Cas78, Cas80a, Cas81b, Cas81a, Cas83, Cha61, Cha71, Chi99, CŠ79, CMR93, Chu83, Cle66, Cle71, CC60, Coh60, Coo71, CS88, Cro76, Cry73, Des70, DPR76, DNvW84, DE94, DE96, DLO^{+}77, Dob89, Dou60, Eis67b, EM91, Eng72, Eng96, FGS95, FB89, GL87, GLR88, GS77, Goë68, GMF99, GG71, Gov90, GP65, GHR65, Gre62, Gun64a, Gus73, Hac77a, Had71, Hai81, HL97, Han82, HP75, Heb88, Hen75, Hin96, HvdR63, HP83, HMSW99, Hsu61]. numerical [Hsu62, Hun73, Imh63, iS91, Ish84b, IU82, JLR70, Jer85, JN94, Joc80, Kad67, KW95, Kra86, Kra63, KP90, Kre68, Kre72b, KS93, KM98, Kut75, LG76a, LG76b, LSW85, Lan87, LV84, LP78, Li95, Li97, Lue79, LT68, MNS74, Mar84, Mat93a, MS67, Mei97, MM74, MP92, Mik70, Mil72, Mon84, Mon87, Mor80, Mül79a, Mül79b, Mül80, Mül82b, MI82, NMC89, Neg90, NW73, Nev77, Nie73b, NS91, NS95b, NS96, Nov86, OP88, Oli68b, Oli71, Osb65, Osb67, Osb68, PE72, Pag81, Pao95, PW86, Par62a, Par65b, Pau81a, Pau81b, Ped96, PPV69, PB81, RS70, RL81, RK84, Rei82, Ren79, RKL89, SP72, SU89, Sal60, Sal63, SS76, SK96, Saw82, Sch64b, Sch66b, SW92, Soc88, Spi66, Spi76, Spi83, Ste95, Tan92, Tem83, Ura70]. numerical [Van79, VŽ95, Vei60, VV88, WL79, Wil64, Wit89, Woo68, Wyn62, Zad76, dFSS97, vdGK86, Aus71, BN90, Cro77, DN86, KO98, LV74b, Leg66, Nat69b, Pit73, WS59, Web79, Wei80, Wit78a]. numerically [BMX98, GH84, Gün98]. Numérique [LV74b, BN90, DN86, GG86, JLR70, LV84, Leg66, Pau81a, Pau81b, SP72]. Numériques [LV74a, Aus71]. Numerische [Ano83b, Bra60, Bra61, Eng72, Kad67, NW73, Nic66a, Sch66b, WL79, Cro77, Nat69b, Pit73, WS59, Web79, Wei80, Wit78a]. numerischen [Alb61a, Ans64, BS59, Kre72b, Lan87, Mar65, Mül79a, Mül79b, Mül80, Mül82b, Rut60, Saw82, Sch64b, Tör63]. numerisches [Nie73b]. Numerov [AP85]. Nyström [Atk73, CG88b, HW76, Hai77, Kre90, NV80a, Spe75].

O. [GP85]. O.D.E.s [Cas80b]. obéit [BN90]. Obere [Lin76]. obeying [BN90, BL93]. Obituary [GV88]. objective [Eyd84, Eyd85]. observables [MPP99]. observation [Par62a]. observations [JTW89]. obstacle [HHZ94, JN94, Kik81, Krä89, Sch86]. obstacles [ABF99, OO88]. obtain [Wyn66]. obtained [Sch81a, TM73, Zul88]. Occasion [Ano83b, FV94]. occurring [Kra86, Rei84a, Wer72, vdGK86]. odd [Mar70c, Rei84a]. ODE [Kac84, Kir88, NR92, EN80]. ODE-solver [Kir88, NR92]. ODEs [Alb85, Dah85, Ste90b, Wri94]. off [GF73, Woź78]. Ohne [AH72]. Olkin [Els69]. Olver [Oli68a, Van79]. One [BNS78, DHZ87, Jou71a, Man78a, Squ69, VH89, Ziv83, AFW93, BV84, BJZ88, Bey87, BSV68, Bou90, CP76, CC85, DML83, Fun88, FH90, GK94, GB86a, GB86b, GB86c, Jam70, Jou71b, Kir86, Kle81a, Kle81b, LSW85, Mey70, Nak85, PB81, RK84, Rei75, RF94, Seg99, Sha84, Sto93, Str82, Stu89, TW83, Ura70, Ver84, Wei92a, CSV67]. One-Dimensional [Jou71a, AFW93, BV84, Bou90, CC85, FH90, GB86a, GB86b, GB86c, Jam70, Jou71b, LSW85, RK84, Ver84]. one-leg [DML83]. one-phase [Mey70]. One-Sided [VH89, Str82]. One-step [DHZ87, Man78a, Squ69, BJZ88, Bey87, Kir86, PB81, Sha84, Ura70]. ones [vG82]. ONLINE [Ano99]. onto [MT78, Opf80]. open [HMS99]. opérateur [Cha73b, Rob66]. opérateurs [Cha73b, CL75]. operation [Dav62]. operational [Egg92, Lub88a, Lub88b]. operations [Lor92]. Operator [Ans73, Boh69, CSV69, Mel99, Rob66, Cha74, Cha73b, CDL96, DLD94, EJ81, FL70, FGM67, HM89, KR97, Kos86, Lei71, Lem98, MD77, Moo68, NC93, PL69, Per67, Rap77, Ste71, SF86, Tho75, Vog83b, Wol78b, NW74, Sch66b, Yse80]. operator-theoretical [MD77].

Operator-Ungleichungen [Sch66b].

Operatoren [Sac73, Had72a, Kna71, Wit77]. Operatorenfolgen [Ans73]. Operatorgleichungen [Kos86, Boh69]. Operators [KV72, Kna71, Nat71, Sac73, Bau65a, BG95, Bea79, Bea86a, BH65, Cha73b, CL75, Cha79, DP63, Gri70b, Han97, HSY91, Heu98, Heu01, Kel65, Kje61, KS83, Küp76b, Küp78, LMR95, LM75, MS90, MS94, May65, Mil67, Sch94b, Sch65, Sch70d, Ser99a, Söd86, Sti88, Wag83a, Yam68, Zen84b, Had72a, Küp76a, Wit77]. Optimal [BH90, BWW73, CK74, CK76, CH86, ENR90, GN87, HP94, Koh87, Koh88, Leh66, Lev82, Ley90, Mei67, MMS92, Rei88, Rou87, Sal60, Sik82b, Sik84, Ste78a, Str69, Tha90, Tro86, ZH93, vG82, AB84, Arn82, Bar68, BLW74, BL76, Bro78, DS91, EG95, Els84b, EG94, För82, GT67, GM77, GNP89, GM94, GHR97, HT95, IS96, KS89, KW99, KWW85, LS70, MM74, MRW76, OO88, Pal77, Pat93, Pau77, Rum96, Sch81a, Sik82a, Ste86, Tom84, Vog83a, Whi77, Wil97, Zer94, Zha89, For75, GS74, Jet76, KK74, KK76, Leh63, Pal76, Sch72a, Sch63]. optimale [Bro78, GS74, Jet76, KK74, KK76, Leh63, Pal76, Sch72a, Sch63]. optimalen [För82]. optimaler [Leh66, Pal77, Pau77]. Optimalität [For75, Nit68]. Optimality [Nit68, PHP90, ST90, Sug97, Luk93]. Optimally [Bau63, Bau69, Bus68, Gau75b, Sha82]. Optimierung [Kra68]. Optimierungs [Wet70]. Optimierungs- [Wet70]. Optimierungsaufgaben [Sau73]. Optimierungsprobleme [SV70]. Optimierungsproblemen [KR82]. Optimierungsverfahren [Sch75d]. Optimisation [Mav77, Sch75d]. Optimization [Kra68, Wet70, ABFJ97, GT82b, GLL91, Han80, McC75, Opf80, Opf82, OO88, Pot89, Reu91, SDK91, SV70]. Optimum [Had87, LT98a, DB77, For68b, NF69, You71]. orbit [AD93]. Orbits [CKP95, FP93, Had80, Had96]. ord [Sau67]. Order [Ade69, CGG90, Deu83, GV61a, GV61b, GH99, Hai81, NW79, Rut60, SWB91, Ver69, AF88a, AU62, Alb78, Ans63, ADG84, BO80, BDS80, BL86a, BBG98, Bel83, Bra66, BDM85, BDDF87, BS88, BEHH75, CR93, Cas78, Cas81b, Cas81a, CŠ79, Chr70a, Chr70b, CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, Coc65, CGT94, Cry73, Cur70, CV84b, Del73, DH80, DPRW93, EFM89, EG95, FF94, Fer97, FGN81, Fun88, FH90, GN87, GW86, GW95a, GHR65, GM94, Guo88, Hai78, Hai79, HHM76, Hav84, Her90, Het75, HL99, Jeg68, JO80, Kac84, KR79, KW81, Kio79, Kos85, KNR95, Küp78, LG76a, LG76b, Lau98, Lev89, LT98b, LR73, LS70, MK60, Man78a, Man74, Mar84, MGR87, Mil85, Moc76, Mül82b, Nat69a, PPV69, PS84, Pfl98, PV72]. order [Pit88, Pru75, Qun91, Rei97, SA84, SSVH87, Sch61d, Sch61e, Sch84, Sch74a, Ser99b, Sha67, She92, Son92, Spr76, SS95, Sun95, Sun98, Tho90, Tro76, Ura70, Van79, Vas88, Wag76, Wal91, Wan89, Wer79, Wer82, Yse83, Zha89, Zlá69, Zlá70, dV93, tMS86, vV78, AU61, ADK91, Gri75c, Het74]. Ordered [Kra68, Boh67, Sar96, Tay69b]. Ordering [MC84, Nic74, Ale82a, Bra92, Bro68, NF69, Rei89a]. Orderings [BV68]. orders [CL75]. ordinaires [Tho90]. Ordinary [CKP95, Nat70, Nat71, Rut60, Wer69, AKK74, Alb60, Alb78, Arn79, BD83, Ber59a, Bet70, Bey78, BS66b, Cas78, Chr70a, Chr70b, Coo71, CW78, DD98, DLO^{+}77, Ess77, FG73, För67, Gar96, Gau61, GS74, GF73, Gri70b, Gri85, Hai81, HL60, Ise84, KR79, Küp76a, Küp76b, LS66, Lan77, Lyc72, Man78a, Mar75, Mey70, Mül79a, Mül79b, Mül80, Nat69a, Nat75a, PS75, Sch61d, Sch61e, Sey79b, Sey83, Sha67, Squ69, Tho90, Tro76, Ura70, Web79, Wet60, Zad76, Zha93]. ordinate [Zla81]. ordinates [NV80a].

Ordinateurs [LV74a]. Ordnung

[Ade69, Ans63, Het74, Het75, Jeg68, Küp78, Tro76, Wer79, AU61, Kos85, Mar70c, Nat69a, Rut60, Sch61d, Sch61e, Spr76]. ordre [Del73, Tho90]. ordres [CL75]. oriented [Gün98]. origin [Lub88c]. Orlicz [Rob71]. Orr [Fis93]. Orthogonal [GR90, LBF98, Rat82, RF94, Rut66a, Spä86, Str63, Wei59, Bia91, BD97, Faß97, Fer97, Gau86, GZ95, HP76a, Hwa99, Sch95a, SW87, Sta77b, Ste69a, Wim74, Xu94, dHH87]. orthogonaler [Rut66a]. Orthogonality [Peh95, GLM91, Jia94]. orthogonalization [Qia88, Swe84, Swe91]. orthonormal [DV99, Ste82]. Orthosymplectic [LV97]. oscillating [AKS88]. Oscillation [DR75]. oscillations [Chi91, CCK95, vD73]. oscillators [GMF99]. oscillatory [Cla76, HL99, Won82]. other [Wyn60]. outer [NC93]. output [HO90a]. over-determined [Par62b]. Over-Relaxation [BV68, BK84, Bro64, Kje61]. overdetermined [Bra82, Sre69, Sre71]. overflow [Cha87, Cha88]. Overlapping [CSZ96, FS99, NN97]. Overrelaxation [GV61a, GV61b, ENR90, Kri83, NF69, Sar86, Sch74b, You71, Hüb72].

P [JBC92]. Padé [Cha84, Cuy82, Cuy83b, CV84b, CDL96, DD84b, DD84a, DMGVO96, FL70, Gut89a, Lub88c, Lub89b, SV75, SV76a, SV78, VC94, Wer86, Wyn66, vI84]. Padé-approximants [Cuy82, Cuy83b]. Padé-Hermite [DD84b]. Padé-type [DMGVO96, vI84]. pair [gS83]. pairs [Ves93, Zha96]. Pairwise [Bar91a]. panel [HN89]. pantograph [Kot99]. Paper [Aro63a, Ber60, Bra71b, Dou62b, Els69, Med71, RS65]. Parabolic [Cai91a, Cho90, Wal68, vFvH73, AF88a, AFW93, ACM99, And69, Aro63a, Aro63b, Axe77, BBG98, BW93, BB82a, BB82b, Car71, CP76, CZ98, CEL98, CL96b, Cro80, Dou60, DJ62, DD73a, EK99, FGM67, FJ75, Fet87, Gek78, GLM91, GK80, GM68, Hai80, Höh82, Ish84b, Jov89, KM67, KP91, Le 80a, Lee61, LS92, MF64, Osb65, Osb67, Osb68, Pal99, Seg99, SO89, Tee64, TW83, Tom84, Voi74, Voi79, Vui90, Wid66, Žen89, Zou97, tMS86, vW77b]. parabolic-elliptic [Žen89]. paraboliques [Cro80, Le 80a]. parabolische [GK80, vFvH73, vW77b]. parabolischen [Wal68]. paracontractions [EKN92]. Parallel [CMR86, FM89, Fro89, Gar76, HJ90, LMR95, BDG97, CKL98, DPRW93, Heg91, Heg94, LRZ95, LH88, Nie89, Shr91, VB92b, Wri94]. Parallelization [DP88b]. Parameter [BB90, GI99, MI82, Sch94a, Arn81, Blu61, Dia80, GG87, GR86, GHMG85, Hüb72, nJyH99, Kal98, Lan64b, Luk93, MS78, MVW79, Ric74a, Sch99, Seo95, Zet70, Wol78a]. parameterabhängiger [MS78]. parameterized [DM96, Rhe88]. parameternichtlinearer [Wol78a]. parameters [DL97, DA86, EG94, HR79, LT98a, Man78b, SKL88, Zet70]. paramètre [Dia80]. parametric [Cle92, DD95, Pla96, Sak97]. Parametrization [MS85]. Parametrized [Had69a, FR84]. parasitic [Ver77a]. Paraxial [LMGR94]. parmi [Rob66]. Parrinello [BS98a, BS99b]. part [BK96, Bia90, EV97, GMP95, KR84, Kut75, ML98, Pag81, Coh60, GB86a, GB86b, GB86c, Sto97, Wyn59a, Wyn59b]. Partial [AU61, Kre63, vV77, Alb60, AU62, AD93, BSV68, BEHH75, CP76, CT91, DPRW93, GP65, HI71, Kap96, Koc95, Kre59, Lee61, LRT64, May65, Mer67, Par65b, SSVH87, Sch95a, Ste63, Str64, Tur78]. partially [Hsu61]. Particle [CMG90, DN89, LG76a, MGR87, Yse97a, Yse97b, Yse99a]. Particles [Yse99b, LG76b]. Particular [Cor77, OO98]. partielle [Kre59, Kre63]. partieller [AU61]. partition [Ros64]. Partitioned [BdP72, BSW88, GT82b, Ren85, SW84, Die87, Dry84, GT82a, Hai81, JO72, MV80, Rei97, Ste82]. Partitionen [RL66]. Partitioning [Ben62, Har62, Heg91]. partitions [RL66]. pass [CMR93]. patched [FG77]. patching [FMMS99]. path [GI99, MS85, TIK67, vGS84]. path-following [GI99]. path-integral [TIK67]. paths [SW99b, ZH93]. patterns [BH71a, Gor70]. PCG [Ser99b]. PDE [Cha99]. PDEs [HS97, AK99]. Peaceman [Ale76, Ale76, MF64]. Peano [Die96, Fie87, Jet76]. Peano-kernels [Jet76]. Peano-Kernen [Jet76]. Peanokerne [Fie87]. penalization [ABF99]. Penalty [Hua90, SB90, BE86, Har78, Hop81, Kik81, Kin74, Sch86, She92]. penalty-finite [Kik81]. penalty-projection [She92]. pencils [BMX98, Kub84]. Penrose [Lue79]. Pereyra [Hig87]. Perfidious [Coh94]. performance [GL87]. Periodenverhältnisse [Woh63]. periodic [AD93, BJ85b, Bao96, Bel83, BMM87, BS95, Boj93, Bro78, Cas81a, CD73, CR74, Dry94, DGP93, For75, For78, vG72, Had80, Hsu61, Hsu62, Jac85, JW88, KK74, KK76, Koß81, Kre71, Kre72b, NW84, Osb65, Osb67, Osb68, Tee64, TW80, Zen84b, Zha93, Zhe89, CSV68c]. periodic-nonperiodic [BMM87]. periodischen [KK74, KK76, Zhe89]. periodischer [Bro78, For75, For78, Koß81, Kre71, Kre72b]. Periods [Woh63, Tod90]. Perturbation [Bar93, CL71, EP94, O’C93, gS96b, gS98a, Sun68, Dei89, Her90, Nii84, Nii85, NS95b, NS96, Osb69, SU89, Ste93b, SO87, gS83, gS93a, WH86, CL71, Sau73]. Perturbations [AD73, AKK74, Gau86, HP92]. Perturbed [NW81, Sto97, Asc85, Del73, GMF99, NY83, O’R87, O’R84, Ste84, SO86, SO89, SS95, Ver98b, vV78]. perturbées [Del73]. Peter [GV88]. Petrov [FHLM88, MMS92, PR94, Web94]. phase [BB97, Mey70, MP92, NS91, VV98]. phenomena [Ste68]. phenomenon [CP94b, GS95a, KN84]. physics [Dör66]. Physik [Dör66]. Picard [Nev90a, Nev90b]. Pick [AY83, KC93]. Piecewise [BSV68, DM88, FQ99, Hul71, Pre71, DP88a, DD74, GS95a, HHM76, Her70, JY93, Kiw94, Koh87, Koh88, Mel72, PR86, Ric78, Tou92, Vog83b, Wid77]. Pivot [Cry68]. pivoted [Ste99, Ves99]. Pivoting [DK77, Peñ98, vV77, vdS70a]. place [Heg94]. Planar [BS98b, CH86, Fri70, HP76a]. Planck [Lem98, DLD94]. plane [ADG84, DG86, GS96a, Rei91, Saa86, Sar88, SS88, Ste76a, Tho71]. plaques [DN86]. plaquette [Tor97]. plaquettes [Tor97]. plasma [DLD94, EJ81, Rap77]. plastic [Han92, RG88]. plasticity [Joh76]. plate [BDT88, Bre96, ByS99, Com85, Dur91, Joh73, Kut71, OP96, PS96b, Pow94, Ran79, Sch96, d’H93a]. plate-bending [PS96b]. plateau [Wil61]. plates [DN86, Ish79, NMC89, Pei91, Pit88]. plus [HJR88a]. plusieurs [Tis72]. Poincaré [Lev89, Sch94b]. Poincaré-type [Lev89]. Point [Dek71, Men88, MRMW90, Pic72, Wil60, Ale71, AM78, AKS88, BD83, Bak82, BWY90, BM87, Bla72, BF70, Bos71, CK82, DPR76, DMGVO96, DD73b, DD74, Fra76, Fra77, GS95a, GeN80b, GN66, IS96, IJ87, Jer89, KS99, Ley86, LR73, Men74, Mil81, Moc74, MO91, OA75, PPV69, Red73, RS76, Ren79, Şch73b, SD67, Sch75c, Son92, Ste98, SW99b, TS98, TIK67, Ura67, Whi77, XS92, Möl76]. Points [Sch63, BL81, BRR81a, BRR81b, DK83, GS96a, GR86, Hal60a, Hal60b, KZ74, KK97, Mor94, NR95, Pau81a, Pau81b, PS82, RP85, RO70, Row71, Sal60, Sch78b, Sey79b, Sey79a, SS78, SS80, SV79, Ste75b, SO89, VG88]. points-convergence [SS78]. Pointwise [Küp76b, Nii90, Noc89, Sch88b, SV79, KS89, Küp78, Nat75b]. Poisson [Bra81a, BH62, CR74, Lou79, Mey78, NS81, Rei89a, Shi78, Shi79, Sun95, Uhl62, ZG78]. Poissonsche [Uhl62]. polarization [Bao96]. poles [HSY91, Lub89b, SV75, SSV76a, SV76a, SSV76b, SV78, VC94]. pollution [BR88, RSSV97]. Polya [Dir69, Dir69]. polygon [Dry82]. polygonal [EJSS97, Hin96, KS98, YN93]. polygons [EG95, Lau98, Wil62e, vPS96]. polyhedra [Gre75]. polyhedral [AO76, AO77a, AO77b, GS98c, SD94, WW68]. polyhedron [MN68, Rat93]. polynômiale [SA84]. Polynom [BS70]. Polynom-Nullstellen [BS70]. Polynomapproximation [Pit73]. Polynome [Bro65b, Rei75, Ehr69]. Polynomen [Bra77, Ess60, Ker66, Sch82, Sik59]. Polynomen. [Sik61]. Polynomgleichungen [SD67]. Polynomgleichungssystemen [Dre77]. Polynomial [BS70, Coh94, HB59, Nic66a, Ric65, Sau97, Skr98, Uhl92, Ziv83, Akr88, Bak70, BH73, BM84, Car91, CK74, CK76, Din83, DD74, Dre77, DM88, ELR92, Els73, GH72, GMS95, Gut72, Hag87, HS69, Hul71, JT70, Jia94, Lam70, LSY87, MM67, MS95a, PL69, Peh95, Pit73, RS86, Ric78, SA84, SD67, ST91, SS65, Sha70, Sou69, Ste69b, Ste73b, VC92b, Wam94, Wid77, Wul73, Zul88, Lam70]. Polynomials [BS63, Ehr69, Ess60, Ker66, MRS95, NR68, Riv62, Sik61, TT94, AJS90, BD77, Bau66, Bea86b, Bet70, Boe64, Bra77, Bro65b, Car92, CR93, Dan66, Eie89, Eve64, Fro88b, Gau61, Gau86, GZ95, HP76a, HHM76, Hen79, IM94, Jia94, Kra86, LS87, Loc93, Mel99, Nør75, Peh95, Rei71a, Rei75, Row71, Rus84, SV76b, SV77, Sch82, Sik59, Skr98, Ste70, Str63, Tsa79, Vog83b, Wei59, Wil59a, Wil59b, Wim74, Xu94, Žen70, vdS70c, Pit73]. Polynominal [Fre90]. Polynoms [Nic66a]. polythermal [CDV99]. polytopes [HS92]. popularity [BB62]. porous [Alt80, JK91a, VV98]. portion [BK96].

Posed [Dem87, CRZ99, ELP90, Han91b, HNS95, Hào94, HA98, nJyH99, Kal98, Kin92, Kol78, Lam91, Nat77, PV90, Pla96, RvdS90, Rie97, Sch93, Sch98, Sch87d, Sch88b]. posedness [GP85]. positiv [Kna71]. positiv-definiter [Kna71]. Positive [AR85, DE94, EHM94, Kna71, MW65, MPW65, MPW66, Nic74, Sch65, dBP77, All70, Bau65a, Bra85, Coh59, FP66, GK90, Had72a, Huc98, Mat96, NN87, Pet93, Spr76]. positiver [Spr76]. Positivity [Ask72, Zen84a, Hel68, PS96a, Sch70d]. possessing [CK74, CK76]. possibilities [Häm65]. possibility [Bra70, Spi76]. Possible [Sto62, Mir71]. Possion [ML90]. postconditioner [Jer85]. Posteriori [Auc92, BS63, AABM98, AF88a, AFW93, AC92, AO93, Ain96, Ain98, BB82a, BB82b, Bra94, CD98, Cha79, Fae98, Gut72, HMSW99, nJyH99, Kal98, Maî68, Moc75, Nov86, PR94, Sch96, Seg99, SW96, Sie96, Ver89, Ver91b, Ver98a, Ver98b, Žit89]. Postscript [Hou73]. Potency [vF70]. potential [Kes84, LT68, Mar90, Pie70, Pie72, Rat93, Sar88, Ver87b]. potentials [Hal77, Wen68]. Potenzbeschränktheit [vF70]. Potenzreihenabschnitten [Rei71a]. pour [Aus71, Bre75b, CR89, Cha79, Cro80, Del80, Del73, FM92a, FM92b, Gue94, Hai77, Le 80a, LV84, Néd82, Pau81a, Pau81b, SP72, Tro90]. Powell [Sch81b, Sch81c]. Power [Lub88c, Nev90b, BN90, BL93, CD72, DML83, DVZ98, Kun74, Mil67, MSW92, SS65, Tho90, Wil87]. power-bounded [Mil67]. power-boundedness [DML83]. Powers [KV72, BP66, May85]. practical [BE87, Gau72]. preassigned [Gue88]. Precision [Dek71, CK74, CK76, DV97, Ish85, Pat76, Spe71, Sri94]. preconditioned [AL86b, EO99, Sta97]. preconditioner [Bre96, GC96, Hem98, Ste97]. Preconditioners [Fre90, AKP95, ELV91, GO94b, HV99, KS99, Yse90].

Preconditioning

[KP95, Won92a, Won92b, Won93, Yse89, AL86a, AV89, Bra92, DK92, DC99, Guo99, Jia94, KKL97, MP96a, Man90, Peh95]. Preconditionings [HN90]. predictor [FH87]. predictor-corrector [FH87]. Preisach [HM89]. prescribed [Opf80]. Presence [Not93c, DP63, GK66, Rau86]. Présentation [Par73, Par73]. Preserving [DE96, AL93, BMX98, CM95, Cle92, DM85, FGM87, KjS95, Kio79, Len89, Mei86, PS96a, SKL88]. pressure [BN84, Jen91, KP90, Rua85, She94, ZL94]. prewavelets [BM92, FQ99]. Primal [BHR77, Ros64, Car99, Kle66, Mil85, Sch99]. Primitive [Fil64, BP79, BB62, Lew71, Vit75]. primitivity [HL64b]. principal [CM89, DD99, Die96, EP75, EP76, Hun72, Hun73, Kut75, Lyn85, Mon84, Mon87, PE72, Ver77a, Wes75]. principe [RM72]. Principle [Bit68, Bit71, Rut63a, Dan70, GS77, NW74, Pla96, RM72, Ste78b, YK88]. Principles [Ans66, Col61, Wer69, FG77, PS96b, RG88, Zen84b]. Prinzips [Rut63a]. priori [Gek78, GF73, Kos86, McA68, PR94, FJ72]. priori-Schranken [Kos86]. probabilistic [Gao89]. probabilities [Deá80]. probability [MO64]. Problem [BB90, Bre75b, Dem87, DRW81, Leh61, Mor74, NO68, Ort63, Ple76, PC83, SB90, Wat68, AU62, AY83, AT79, And71, Arn81, ASW80, AHS84, AKS88, ALS90, BL92, BV87, BV88, BL86a, BE86, BE87, BS91, BC91, BW95, BS99a, BP79, Bit91, BZ82, BH62, Bra66, Bra69, BD88, Bri64, Bro85, BEHH75, Car99, Cas85, CSV67, Com85, CDR98, Dei89, DH80, Döh82, DD73b, DD74, Dry82, EP99, Fei87, FN98, Fis93, FG89, FP90, FM92a, FM92b, GH92, GP79, Gol81, Gre82, Gre80, Gri75b, Haa60, Hac77b, Had71, HHM76, Han92, Han94, HS83, HL64a, HL65, Her90, HT95, HS70, HHZ94, HMW86, Ish84a, Jen91, Joc80, KM67, Kea68, KS98, Kik81, Kim98b, KP90, KS83, Lan87, Law64, Li90, LRZ95, LM91, Mar84, MP95a, Mar90, Mei86]. problem [MS73, Mey70, Moc74, MG85, Nag63, NS95a, Nii84, Nii85, Osb69, Osb65, Osb67, Osb68, OP96, PW93, Par62b, Pau81a, Pau81b, Pei85, Pit92, RvdS90, Rap84, Red88, RG88, RL81, RK84, Rei91, Ric74b, Rod73, Rou87, Rua85, RL66, Sac88, SV90, SDK91, Sau84, Şch73b, SS88, Sch99, Sch86, Shi78, Shi79, Smi67, Spa93, Spe75, Ste86, Ste88, SO86, SO87, SS95, gS99, Utr83, VB92a, VV84, Voi74, WFS95, WZ95, ZG78, Zer94, Zha95, d’H93a, dO70, vV78, Mar75, RL66, Sau78, Hav77, SV70]. Probleme [Hav77, Mar65, Nat77, DRW81, Bre75b, FM92a, FM92b, Mor74, Pau81a, Pau81b, Bra70, Lor77, Nat69b]. Problemen [Het75]. problèmes [Con84, LV84, SP72, Tro90]. Problemklassen [Cro77]. Problems [AU61, Ans69, ADdR89, Aum63, Bau65b, Ber59b, BL90, Cho90, Cia90, CS59, Gol65b, GvM91, Had68, Had69a, Hua90, Leh60, Leh66, Mar65, MW68b, RR90, Sau63, SS79, Smi91, Tre62, Wer69, Wet70, vdS75, vFvH73, Ain98, ACP99, AM78, Alt80, AG91, AO76, AO77a, AO77b, AS97b, AR68, AD73, Ans64, Ans67, Arn82, Arn84, Asc85, ASS88, BO80, BV84, BS92a, Bak82, BWY90, BD98, BYI86, BW93, Ben62, Ben64, BH61b, BH65, BD82, Ber59a, BD97, BK81, Bjö88, Bla72, BLR89, BC78, BG78, BS96, Boh67, Boh80, BD96, BF70, Bos71, Bra70, BRR80, BRR81a, BRR81b, BDM85, BDDF87, BHK94, CRZ99, CHL85, CH71, Cas81a, Cha98, CL97, CK82, CZ98, CEL98, CP94b, Chi91, CCK95, Chi99, CŠ79, CSV67, CSV68a, CSV68b]. problems [CSV68c, CSV69, CNV70, CHS96, Con84, Cry73, DP72, DPR76, DLO^{+}77, Die87, Dij59, Dry84, DSW96, Dun88, Dut99, ES90, ELP90, EO99, EJSS97, Ern96, Ess77, EN80, EL88, ELV91, Eyd84, Eyd85, Fag81b, FJ75, FF97, FŽ87, FŽ88, Fer97, FN72, FG77, För67, Fra76, Fra77, FR86, FH90, GS74, Gek74, Gek78, Geo79, GLM91, Gir89, GeN80b, GT82b, Gri70a, Gri75c, Gri75a, GKR86, GN66, GB86a, GB86b, GB86c, GHR97, GS93b, GS98c, Guo99, Gus73, Hal77, Han82, Han91a, HNS95, HV99, Hào94, Har78, Hav77, HS92, Hee68, HA98, Hei89a, Hof67, Hof75, Hul71, Ish81, Ish85, IU82, IJ87, JBC92, Jam70, JN81, JN82a, JN82b, Jer70, JN94, nJyH99, JMT97, Joh73, Joh76, JM78, Kal98, Kan98, KS89, KW96, Kin92, KS71, KS99, Kle65]. problems [Kle81a, Kle81b, Kna85, KZ94, Kol78, KR82, Kos85, Kra86, Krä89, Kre59, Kub84, Lan77, LN99, Le 82, LV84, Ley86, Li86, Li87, LR89b, Lin76, LH88, LBF98, Lou86, LR89c, Lub92a, Lub94, LR73, Man85, Man84, Mar75, MP95b, Mas87, Mat93a, Mat93b, MBL90, McC75, Mer91, Mer92, Mey79, Mey84a, Mil85, Mis99, Mit81, Moc75, MT89, MO91, Mül82a, Mül82b, Nat77, Neg90, Nev77, NY83, Noc88, NW76, OA75, Opf82, Pao87, Ped96, PCV96, Per68, PS75, PPV69, Per99, Pet70, PV72, PB81, PV90, Pon66, PR94, Ran79, Ran84, Rau86, Raz69, Red73, RS76, Red77, Ren79, Reu88a, Ric74a, Rie97, Rua88, Run70, RW88, RS72, Rus74, Sad93a, SP72, SU89, Sán75, SSVH87, Sar88, Sar97, Sav93, Sch93, Sch98, Şch75a]. problems [Sch87b, Sch87d, Sch88b, Sch86, ST73, Sch75c, STR76, Sch84, Sch97, Sch71b, Sch72b, Sch78c, Seo95, Ser99b, Sey83, Sha68, SW96, Smo90, Soc88, Spi66, Spi76, Spi83, Spr80, Sta77a, Sta77b, Sta97, SS96, Ste63, Ste68, SW99b, Str86, Str64, SO89, Tee64, Tom84, TS98, Tou92, Tro76, Tro90, TIK67, Ura67, VV89, VŽ95, VW92, VV88, VV98, Vol86, Vos75, WKW97, Wan89, Web79, Wei92a, Wei63, Wen68, Wet60, Wil61, Wol78a, XS92, Žen89, Žen95, Zlá69, Zou97, vdGK86, Bey78, GK80, Het75, Lor77, Nat69b, Nit69b, Sau73, Sch66b, Spr76, Wit78b, Wit78c]. procédés [Bre72, Gas63]. Procedure [Ghi68, Blu61, DD78, DPRW93, FG89, GMP95, Gir89, LMS77, Lue79, Man78b, MQ89, McC75, Mil65b, MH72, OA75, Şch73b, Şch75a, Sto67, Tim65, Tsu88, Zlá70]. Procedures [Gas63, HS80, AB74, Ans63, Ben62, BZ94, Cro76, Cro78, DP63, Kim98a, Mil72, Nic75, Pea62, Ves76b, Ves79a, Wyn64, Zlá69, Eng72]. Process [Sid90, Ale82b, BK81, Cur70, JM87, LW82, PP80, Ptá76, Ral66, Sid82, Wil62d, vV76]. processes [BWW73, Bre72, GT67, Mar69, Nit70, Sep72, Tan74, Wal91, ES73]. processing [CCCD93, Heg94]. Procrustes [EP99]. Product [DP88a, Dag90, MC88a, MC88b, SS78, SS80, BJ86, Ban78, BM96, DD99, Dav62, Gas71a, LSY87, LRT64, Mon84, WA72, Wes75, Sch80a]. Product-integration [SS78]. production [Son92]. Products [Cup84, Bet70, SSS99, ST91]. produits [Gas71a]. Produktintegration [Sch80a]. Professor [Sau67, FV94]. Program [FN68, Neu61]. program-controlled [Neu61]. Programm [FN68]. Programmable [Kar61, Neu60]. Programmation [DD84a, Con80]. programmgesteuerten [Neu60, Neu61]. programmgesteuerter [Kar61]. Programmierung [Fel61, Sto64b]. Programming [GK64, Ben62, Bit70, CG59, Con80, Dij60, GK66, Kiw94, Kle65, Kle66, KR82, Meh67, MT89, PS59, Ros64, Sau73, Sch81b, Sch81c, Sti60, Sto63, Sto64b, DD84a, Fel61]. programs [Kal74]. progressively [LR89a]. projected [Dun88]. Projection [BF90, LM75, Red73, Tan71, DDW75, GR86, GN66, Gue94, GQ98, HS92, He94, LH88, LM91, LR73, Nat77, NN87, PV90, Pyl67, Sch77a, She92, She94, Wit77, Wit78a, Zen88, Vos75]. projection-iteration [Sch77a]. projections [DD73b, Hav84, LR89c, Weg89]. Projectionsverfahren [Wit77, Nat77, Vos75]. Projective [Wit78b, Wit78c]. Projektionen [Hav84].

Projektions [Sch77a].

Projektions-Iterationsverfahren

[Sch77a]. Projektionsverfahren [Wit78a]. Projektive [Wit78b, Wit78c]. prolongations [Nev90b]. proof [Bau65a, Hai78, See82]. proofs [Weg84]. propagated [Zad76]. propagation [DG86, GJ89, Oli67, Wyn59c]. proper [CL75, Cha79]. Properties [Ruh70, VHS90, BS77, Bar68, Bre89, CGR90, EKL91, FHL86a, FHL86b, Gug97, JK78, Kle81a, Law64, Men88, Mik70, Nov88, Opf79, PW86, PL69, Rie91, Sch73a, SS78, Sto77, SW99b, Ver77b, Zen68, Zen86, Zla78, Zla81, dFSS97, Gas65]. property [Bau76, Ein71, Lor77, Nag63, Prö91, SJX97]. proposed [FP90]. proposition [Pit75, Wer62]. propres [CL75, Cha79, Mor74, Par73]. Propriétés [Gas65]. Proving [Sch70d]. proximal [Kiw94]. Pseudo [CdM70, Lin61, Ale71, DA71, EHG89, Fis61, Fun88, Nie85, Par62b]. pseudo-metric [Ale71]. Pseudo-random [CdM70, DA71, Fis61, Nie85]. pseudo-stability [Par62b]. pseudodifferential [BS95, Sar89, Wag83a]. pseudoinverses [Wei95]. Pseudometrische [Ale71]. pseudorandom [Bit91]. Pseudospectra [TT94]. pseudospectral [FH90, yGL93]. Pseudozeros [TT94]. Publications [Ano99]. puissance [BN90]. Punksystemen [Men74]. Punkte [Sch63]. Punkten [Gün75]. punktweise [Küp78, Nat75b]. purpose [Wyn64].

q [Wyn59d]. q-d [Wyn59d]. QD [Cuy83b]. QD-algorithm [Cuy83b]. QMR [FN91]. QR [Par67]. Quadraten [Läu61]. Quadratic [DH93b, GvM91, KRV74, Tha90, AS83, Ale86, BS77, BCD86, Coh59, Con80, For85, Fra62a, Fra62b, Fra63, Gan81, GK90, GK66, Har91, vK66b, Kiw84, Kle81b, Kna85, KR82, Lei71, Lin76, MPP99, MT89, Rua88, Shi82, Sto77, Tao84, Wil62d, Wil84]. Quadratically [GvM91, VW79]. quadratique [Con80]. quadratiques [Tao84]. quadratische [Kle81b]. quadratischen [KR82]. Quadratur [Wit78a, För82, Gai70]. Quadrature [Bia90, BH90, Bra60, Dry94, Eck68, Hae63, Häm66, Loc73, LdD93, PS92b, SS97, Sch72a, Smi65, TM73, Won82, Akr84, AB85, AB84, BJ85a, Ban92, BR91, Ber72, BP98, Bor94, Bra73, Bra78, BS81, Bra85, BF99, Bro78, BS67, Cle71, CM89, Dic87, Die96, DV97, Egg92, EP75, EP76, Eng77, Fie87, For75, FP90, För93, Ghi68, GG86, Gün75, Häm65, Kam71, Köh95, KWW85, Lev82, Loc80, Lub88a, Lub88b, LK78, LG82, Mar70c, Mar73, MY79, MI82, Nik92, NV96, Not90, Not93a, Not93b, Nov88, OPS76, Pat93, Peh88, Peh90, Pin75, Pot78, Rat93, RJ72, Sab78, SD72, Sal75, Sch77b, Sch88a, Sik82b, Slo81, Slo88, Sri94, Str84, Wil64, Wit78a, vdS72, vS77, Gai70]. quadrature-based [Slo88]. Quadratures [KE82, Bar65, Bar68, DD99, För82, GM77, GK83, Kau81, Küt82, Mik70, Mon87, Pec74, Pet93, Ste66a]. Quadratureverfahren [Bra73, Bra78]. Quadraturfehler [Häm65, Hae63]. Quadraturfehlern [Häm66]. Quadraturformel [Mar70c, Mar73]. Quadraturformelmethode [Bra60]. Quadraturformeln [Bro78, Eng77, For75, Gün75, Jet76, Pau77, Sch72a, Akr84, AB85, Bra85, Eck68, Loc73, Loc80, Pot78]. Quadraturverfahren [Fie87]. Quadrilateral [LLS95, LZ80, Man84, cS84, ŽV95]. Quadrilaterals [WS59, Gai72, PS92a]. Qualitative [Gri75b, Sto97]. Qualocation [SW98, CS90, CS92, SW99a]. quantitative [Auz88]. Quartic [SBH84]. Quasi [BL94, CD72, CD90, Nit68, BN90, BM87, CC91, FN91, Hal60a, Hal60b, RH93, Ste61, Tro71, Zou99, Kle81a, Kle81b, BN90, BL93, GT82a, KS89, MO91, Sch74a, SG78]. quasi-cyclic [RH93]. quasi-homogeneous [Tro71]. Quasi-interpolation [CD90]. quasi-Laguerre [Zou99]. quasi-linear [Ste61]. quasi-minimal [FN91]. Quasi-Newton [Kle81a, Kle81b, GT82a, KS89, MO91, Sch74a, SG78]. Quasi-Newton-Verfahren [Kle81a, Kle81b]. quasi-Newtonian [BN90, BL93]. quasi-newtoniens [BN90]. Quasi-norm [BL94]. Quasi-Optimalität [Nit68]. Quasi-Optimality [Nit68]. Quasi-random [CD72, Hal60a, Hal60b]. quasi-variational [BM87, CC91]. quasiconformal [Wei80]. quasiconvex [Klo99]. quasidegree [LG82]. quasidifferentiable [Kiw84]. quasihomogenen [Tro71]. quasikonformer [Wei80]. Quasilinear [Ans69, Tör63, vFvH73, ACM99, AU62, Axe77, Kos85, Lub92a, Mil85, Moc75, Wer68]. quasilineare [Ans69, vFvH73]. quasilinearer [Kos85, Tör63]. Quasisolenoidal [Rua85]. quasistatic [BL86b]. quaternion [BGBM89]. quelconque [LL85]. quelques [Con84]. Question [Ans66]. Questions [BL90]. queueing [Cha87]. queuing [Cha88]. Quotient [Rut63b, Eck76, gS91, See82]. Quotient-Difference [Rut63b, See82]. Quotienten [Rut63b, See82]. Quotienten-Differenzen-Algorithmus [Rut63b, See82]. quotients [McA68, Wyn66].

Rachford [Ale76, MF64].

Rachford-Verfahrens [Ale76]. Radau [Kam71, Kot98]. radial [DM90, IS96, JS91, SW93]. radii [Lyn63, NV80b]. radius [KT59, Luk61, MV69, Sch80b]. Radon [Nat78, Nat79, Ros95]. Ralston [Sch88a]. Rand [Gri70a, Nat75a]. Rand- [Gri70a]. Rand-Eigenwertaufgaben [Nat75a]. Randbedingungen [Vos75]. Random [Hea66, Lin61, BL81, CdM70, CD72, DP88b, DA71, EHG89, Fis61, Hal60a, Hal60b, HI71, Lam80, LSY87, LMS77, Mar70a, Nie85, Web91]. randomized [Ste95]. Randsingularitäten [Nit76]. Randwertaufgaben [Ess77, För67, GS74, ST73, STR76, Spr76, Tro76, Vos75, Wen68, Wer69, Ber59a, Ber59b, Bey78, CS59, EN80, GKR86, Hei89a, Leh60, Sch66b, Tre62, Vol86, Wit78b, Wit78c]. Randwertproblem [Leh61]. Randwertprobleme [Kos85, vFvH73, NY83, Pet70]. Randwertproblemen [Nit69b]. Randwertproblems [Lan87]. Rang [Kle81a, Kle81b]. Rang-Eins-Typ [Kle81a, Kle81b]. range [Car71, DP88b, EHG89, SS76, Küp76a, Küp78]. Range-domain [Küp76a, Küp78]. Ranges [vF70, Tay69a]. Rank [BNS78, Bra82, FH95, Kle81a, Kle81b, Sto84]. rank-2 [Sto84]. rank-one [Kle81a]. rank-one-type [Kle81b]. rank-revealing [FH95]. Raphson [Lan66]. rapid [VI86]. Rapidly [Lei71, AKS88]. Rate [CDR98, Not93c, vdSvdV86, AL86b, CGK92, GT67, LP95, LE81, NN97, Ptá76, Rit80, Ser99b, WFS95]. rates [BD77, Han97, Las92, Sch93, Utr85]. Rational [CS66, CFH68, Hwa99, MW60b, Ral65, RB68, Ric65, Tre81, WSB67, WB69, Ziv83, Zur64, AAG79, Arn80, Bar84, BL78, Bla77b, Boe64, Bro65a, BL67, BS68, CL61, CL62, Cla76, Cla78, Cle92, Dun67, Dun68, Gol63, GMH81, GM83, Kim98b, Lam70, Law64, LL85, Lub88c, Lub89b, Lun85, Nør75, Pom68, Riv62, Sac72, SSV76a, SSV76b, SVN76, SL86, Sak97, Sch83, Sch88b, Sto61, VB92a, VB92b, VG88, Wei92b, Wer72, Wuy71, Wuy73, Wyn60, Wyn62, vdGK86, Bla77a, Bra74, CS68]. rationale [Lam70, MW60b, Wyn60, Bla77a, Bra74]. rationalen [Arn80, Bar84, Bro65a, BL67, BS68, Sto61, WB69, Wer72]. rationaler [Bla77b]. rationnels [AAG79, LL85]. Ratios [Sto62]. Raum [Sac73]. Räume [Ale71]. Räumen [Ade69, Ans73, Boh67, Dör70, Sch78c, Mey84b]. Raumkurven [PS82]. Raviart [BCS96, Cha99]. ray [Lou86, LR89c]. Rayleigh [CL75, Mer91, Mer92, CL75, JT70, KL83, Mer91, Mer92, OA75, PV72, gS91, Žit89]. re [Ste93a]. re-entrant [Ste93a]. reaching [Cro78]. reaction [ASS88, BS91, BW95, LO98, Pao85, Pao98, SS95, Ver98b, YM86]. reaction-diffusion [ASS88, BS91, BW95, LO98, Pao85, Pao98, SS95, Ver98b, YM86]. Real [AR85, BD73, BMPW66, Dir69, PW70, BMX98, BM84, Car72b, Coh60, DD78, Dol81, Har82, HP92, vK66a, Kem68, Kup71, LMPY85, Loc93, MPW70, MS78, Rut66b, SSV76a, SSV76b, Ves76a, Ves79a, VW79]. real-valued [Car72b]. Realistic [SW80]. receiving [Gor70]. Rechenanlagen [Sch61d]. Rechenautomaten [Lin62, Nic66b]. Rechengeräte [Kar61]. Recherche [Rob66]. Rechteckgittern [Tör63]. Rechteckregel [Kre72b]. rechten [Tro71]. reciprocals [Kun74]. reconstruction [BLLN95, GH84, Nat85]. recourse [Kal74]. Recovering [GS95a]. recovery [AS97a, CL97, IS96, MRW76]. rectangles [Die87, Fer97]. Rectangular [HN90, Tör63, CMR86, Dur90, Hof67, MT78]. Recurrence [Met65b, Cas77, Cas80a, Lev89, Müh79, Oli67, Oli68a, Oli68b, Met65a]. recursions [Mat80, Sch75b, Wyn66]. Recursive [Dij60, Lei71, Oli67]. red [Ale82a, Bra92, Par88]. red-black [Ale82a, Bra92, Par88]. Reduced [Mel72, EGS94, GT84, HV64, MK80, PS96b]. reduced-shear [PS96b]. Reducibility [Hil83]. reducible [BJ85a, BNZ72a]. Reducing [EB68, BR88, Ebe70]. Réductibilité [Hil83]. Reduction [GN66, MW68b, MW68d, BV87, BV88, KW99, LB97, Rei89a, Run70, SSVH87, Sch70d, STR76, VC92b, ST73]. réduit [MK80]. Reduktionsverfahren [ST73, STR76]. redundant [Gre75]. reelle [Dol81]. reeller [Kup71]. reentrant [BR88, Qun91]. REENTRY [Mar75]. REENTRY-Problem [Mar75]. refinable [HSS96]. refined [EW92]. Refinement [GW66, MPW66, AL96, BS98b, ELV91, Mat93a, Mat93b, Pav94b, Rei88, Sie96, SS98, vV76]. refinements [AM78, BKP79, Sty81, Tha76]. Reflections [Cup84, RW85b]. reflector [KO98]. reflexive [Pie72]. reformulation [Van79]. Regeln [See82]. Regelungssysteme [Tau76]. region [CM70, Dry82, LR89c, xY95]. region-of-interest [LR89c]. Regions [SSS86, Sar96, Bra77, CM62, CH86, DHP99a, Die87, Dry84, Fri70, HP76a, JO72, JT76, MT78, Opf80, PPSS97, Rei91, RW85b, SV76b, SV77, Shi79, Str67]. regression [CP70]. regula [PP81]. Regular [AV85, Kul68, KM98, Ple76, BS98b, CV84a, Els89, Föß89, Lor92, Sch80b]. reguläre [Kul68, Sch80b]. Regularisierung [Nat77]. Regularities [Mar70a]. Regularity [LMPY85, LS87, Tom84, Auz88, TW68]. regularization [DA86, EG94, FH87, HHZ94, Jer85, nJyH99, Kal98, Luk93, PV90, Rie97, Nat77]. regularized [HV99, Sch93]. Rehabilitation [DH89]. Reibung [Tau76]. Reissner [Pei91, Pit88]. Rekursionen [Sch75b]. Related [MW68b, CL97, Die96, Faß97, Jer70, Kau81, LTT92, Man85, Rap84, SS78]. relatifs [Rom85]. Relation [PC83, BS77, DKS90a, DKS90b, He97, Hod67, JT70, Lam70, Müh79, Nit62, Sto77]. Relations [AR85, Cas77, Cas80a, Lev89, Met65a, Oli67, Oli68a, Oli68b, Wei92a, Met65b]. Relationship [Woh63]. relationships [PHP90]. Relative [O’C96, Oli67, And88, Dei89, EJ81, Wet70, Rom85]. Relative-error [O’C96]. Relaxation [Alb61a, BV68, BS97b, BK84, Bro64, Hod67, Kje61, Klo99, MQ89, Tay69b, You71]. Relaxationsverfahren [Alb61a, AV76]. relaxed [Li97]. Reliable [GMH81, YK88]. remainder [Akr88, BG76a, FP90, LM70, Mar70c, Mar73, Mei66, Pet93, Sch77b, vdS72]. remainders [Bar65]. remark [Bra71b, EHG89, Els73, SF86, BS60, Bul64, Els69, Red67, Rut60, Sch64b, Wal67, Zur63]. Remarks [Bau69, ER67, GG87, GHMG85, Han64, She94, NR92, For68a, För82, Gue94, Had72a, Had72b, Smo90]. Remarques [Gue94]. Remes [Car72a, Lau67, Ral65, Car72a]. Remes-algorithm [Car72a]. Remez [BL70b, NS83b]. Repeated [Jac85, Bra92, MI82, Ste66b]. Report [BBG^{+}60, PS59, vWMPK69, BBG^{+}63]. representation [Bit70, Che79, Coh59, Kan73, Lin62, Mar70c, RS86, Sch77b, Wal89, Che79, Mar73]. representations [Hom95, Mor80]. representing [HSY91]. reproducing [CK74, CK76]. requiring [Kle65]. Residual [gS97, gS98b, AO93, FN91, Jia94, Oli80, Peh95, Stu85a, WFS95, Woh99]. residuals [hDxY99, Pyt02]. Residue [Fel61, Kre71, BS70]. Residuenabschätzung [BS70]. residues [CD72]. Resolution [ML90, Tsu88, Yse99a, Del73, JLR70, Leg66]. resolvent [Mil67]. Resommation [Tho90]. Respect [KS73, BL81, Gar96, Kri64, Nit70]. response [BMR96]. restart [Kle81b, Kle81b]. Restart-Varianten [Kle81b]. Restglied [Kre71]. Restgliedabschätzungen [Pot78]. Restglieddartstellung [Mar70c]. Restgliedes [Mar73]. Restklassenabzählung [Fel61]. restricted [Kuh93, Sto84, Tay69a]. résultats [El 82, DGB80, FM92a, LV74a]. Results [LV74a, Bar93, Ben64, BLLN95, BS66a, DGB80, El 82, FM92a, FH90, ID86, Kri83, Le 82, LMPY85, LS87, Mat93a, Mei97, PV72, Sch95b, SWB91, Wer82, Wit89]. Resummation [Tho90]. retarded [Arn84, FG73, GF73]. retournement [Pau81a, Pau81b]. revealing [FH95]. Revised [BBG^{+}63]. revisited [OP96]. Riccati [DE94, DE96, Las92, O’R87, Raz79, gS97, gS98b, fX96]. Richard [Cia98]. Richardson [Sch75b, BLR86, FR88, Föß89, GV61a, GV61b, GO88, Sch75b, Sid82, Sid90, Zaw91]. Richardson-Extrapolation [Sch75b]. Richtmyer [PSS84]. Riemann [BS97b, DM96]. Right [OP64, CR94, Man78a, Ste90b, Tro71, Vog83b]. Right-Hand [OP64, Man78a, Ste90b]. right-inverse [Vog83b]. Rigorous [CKP95]. Ritz [Ber59b, BS60, Mer91, Mer92, Nit69b, Ber59b, BS60, CH71, CGH76, CL75, Goë68, Leh60, Mer91, Mer92, Nit69b, Nit76, OA75, PV72, Şch73b, Şch75a, SF86, Žit89]. Ritz-Approximationen [Mer92]. Ritz-Verfahrens [Mer91]. Ritzschen [Leh60, Nit76]. RK [NS95b]. RK-methods [NS95b]. Robert [GB71]. Robust [KW96, Ver98b, Reu95, Ste97, WKW97]. Robustness [Ste93a, Egg84, Zha95]. rods [AR93, Cha97]. Roe [BS97b, DM96]. Roe-type [BS97b]. Romberg [BF99, Bul64, Hun73, Jet84, LM69, Mei66, Rut63a, Sch70c, Str73b]. Romberg-Integration [Bul64, Jet84, Sch70c]. Rombergschen [Rut63a]. Root [Lar81, BWW73, CR93, HP76b, HP77, PS84, Ver77a, Yam66]. root-finding [BWW73]. Roots [BH60a, Nic66a, Uhl92, BH61a, Bau66, Bra77, BB62, Car91, Dan66, Hou68, Sou69, Ver77a, vdS70c]. Rosenberg [May86, Mil80, Rob76]. Rosenbrock [KW81, NW79, RRS89, Roc88, SWB91]. Rosenbrock-type [KW81]. Rosenbrock-Wanner [RRS89]. rotationally [Lus96]. Rotations [Rat82, Ste76a]. Round [GF73, Woź78]. Round-off [GF73, Woź78]. rounded [Kul71]. Rounding [Lod70, Not93c]. Roundoff [Mil77]. routines [DLO^{+}77]. Row [Bau65b, Fer89, Mar69, MO68, Gün98]. rows [Peñ98]. Rückkehrpunkten [PS82]. Ruhelage [Tro71]. Rule [Bia90, RS90, AB87, AD64, BD83, EP75, EP76, Fis61, SB83, Sri94, dV93, Kre72b, See82]. rules [AF89, Bor94, Bra78, BNvW89, CK74, CK76, Cor77, CM89, Die96, GI99, LK78, LG82, Mei66]. Runden [Lod70]. Runge [DKS90a, ASW93, BZ85, BSW88, CGPM98, Cro79, Cur70, DKS90b, GW86, GMF99, Gug97, Hai82, Hai86, HISS90, HS81, Hun86, HS90, KR79, Kot99, KS91, Lub82, LS70, NW81, QYK99, Rei97, Ren85, SSVH87, Sch95b, Spi86, SW84, Ver88, Ver92, Ver69, Ver77b, VHS90, Zen86, Zen88, Zen97, Zou97, dHW75, vdH72]. Rutishauser [See82, See82].

S [Ano83a, Ano86, Cia98, JHW77, H. 73, Wil78]. S.O.R [NV80b]. saddle [BWY90, KS99, TS98]. Samelson [GB81, GB81, Ste70]. Sample [Bar91a]. sampling [Bit91, FGS95, GLM89]. sämtlicher [Dre77]. Sard [BG76b]. satisfying [CL96a, Mil67, Sik84]. Saturation [Gai70]. Satz [Wer62]. Sauer [GB71]. scalar [Bel83, Kac76, Kim98a, KC93, KNR95, LS95, MO78, Tro86]. scale [CGC95, Ho90b, Wal97]. Scaled [Bau63, Bau69, Bus68, Sha82, Sch81a, Šve70, Wei95]. Scaling [MO68, Ari98, Els84b, Ste84, vG80]. scalings [vGS84]. scattered [ST91, SW93]. scattering [KO98]. Schaefer [Bra71b]. Schauder [Sab78]. Scheme [Sch70b, Auz88, BR72, Cam70, CGH93, CC85, CL96a, DLD94, Dob89, GA92, GLM89, Grü97, Her90, Ish81, Ish85, Kil75, Kno89, LT98b, MO78, Nii84, Nii85, Nii90, O’R87, Reu88a, Reu88b, Sán75, SO87, SS95, VV88, YM86, Zer94, Rei71a]. schemes [Asc85, BG94, BL97, BEN93, CL96b, Eis67a, EGH99, FF94, FB89, GK80, GS93b, Had69b, JK91a, JK91b, Jer70, Jov89, KMS99, KNR95, LMW84, LS95, Mir71, Moc76, NS95b, NS96, PS96a, Pit88, Rei75, Roh98, SSVH87, Sei63, Son92, SV79, Str73b, SO89, Sun68, Tho71, VV98, Wid66, ZG78, Žen89, IS96]. Schichtenverfahren [vW77a]. schlecht [Nat77]. Schlüter [MSF91]. Schranken [Bod68, Hae63, Häm65, Kna71, Kos86, Lin76, Nat70, Nat71, Tre62]. Schritt [Kle81b]. Schrittweitenbestimmung [Wer78]. Schröder [VG88]. Schrödinger [ADK91, Dör96, KL83, MPP99, RF94]. Schrödinger-type [RF94, MPP99]. Schulz [For68a, For68a]. Schumann [CS86]. Schur [Bre99, Dav62, GH95, KW99, Man90, WKW97]. Schur-complement [WKW97]. schwach [GKR86, Hav77, Kos85]. schwachen [Sch75d]. Schwankungsbereiche [Sau73]. Schwarz [Bad89, Bre96, Cai91a, CSZ96, DSW96, FS99, GO94b, GO95, Hac92, HR79, Mat93a, Mat93b, Mil65b, Pav94a, Pav94b, RKL89, Sar97, Zha92a]. Schwingungen [Tau76, Wer69]. Scott [CV94]. search [Ari98, DD78, GA87, Jar94, Sch81b, Sch81c, Rob66]. searches [Wer78]. Sebastiao [Hou73, Chu76, Hou71]. Secant [PQS98, MS85, Pot82, Ste73b]. Second [AF88a, Fun88, GV61a, GV61b, RR90, Rei90, Atk72, BO80, BDS80, BJ85a, BJ86, Bel83, Bra66, BDM85, BDDF87, BN81, CŠ79, CGT94, Cry73, Del73, DH80, DPRW93, EFM89, Fer97, GW95a, Hai79, HHM76, HS80, Het75, HL99, Hoc79, HvdR63, JO80, JS86, Küp78, Lau98, Lev89, Man85, Mid75, Mil85, Mor99, Nat69a, Pfl98, Ric78, Ser99b, SW96, Slo81, Son92, Spr76, Tho90, Van79, Wan89, Zlá69, Zlá70, dHW75, vdHtR81, Het74, ADK91]. second- [Nat69a]. second-kind [SW96]. Second-order [AF88a, BDS80, CGT94, HHM76, HL99, JO80, Lev89, Son92, Spr76, Tho90, Van79, Het74, ADK91]. Secondary [Sau63]. section [Mor72]. sector [Hee68]. sectors [GS96a, SVN76]. secular [Mel95]. seepage [RBS82]. segment [NSS86]. segmentation [CKSS97]. segmented [Law64]. segments [Rei71a]. Seidel [BdP72, EL88, Par88, RM72]. Seiten [Tro71]. Selbstabbildung [KS73]. selbstadjungierte [Ber59a]. selbstadjungierter [Kna71, Mer92]. Selected [Mar65]. Selection [BF90, NO68, Ain96, GI99, PS75]. selective [Sch77a]. selektives [Sch77a]. self [BG95, Fri70, HR94, Heg94, Jer70, Meh67]. self-adjoint [BG95, Jer70]. self-contained [Fri70]. self-duality [Meh67]. self-sorting [Heg94]. Selfadjoint [Kna71, Mer92, PCV96]. Semi [AGS98, GV61a, GV61b, Sch87d, ASW93, BD83, BE87, EMN88, Fis93, Gun64a, Hem98, Hod67, Lee61, LE81, MPP99, Rei97, Sim78, VV88]. semi-circulant [Hem98]. semi-classical [MPP99]. semi-definite [BE87]. semi-direct [LE81]. Semi-discrete [AGS98]. semi-explicit [ASW93, Gun64a, Lee61, Rei97, VV88]. semi-implicit [BD83]. semi-infinite [Fis93, Sim78]. Semi-Iterative [GV61a, GV61b, Sch87d, EMN88, Hod67]. semicoercive [Spa94]. semiconductor [CC95, DN89]. Semidefinite [Sch72a, Jet76, NN87]. semidefiniten [Jet76, Sch72a]. semidiscrete [Hel87]. semidiscretizations [Str86]. semigroups [Can71, Las92]. semiinfinite [Opf82]. semiinfinite-finite [Opf82]. semiiterative [ENV85, Eie89]. semilinear [BS99a, CMR93, EGH99, GS74, Hel87, LN99, Nii84, Noc88, SO87, SS95]. semilinearen [GS74]. seminormed [Mei67]. seminorms [LM75]. semismooth [PQS98]. sense [Vei60]. Sensitivity [BH92b, BHK94, HUY95, fX96, Gau86]. separable [DG98]. Separation [Hou66, BB97]. September [GV88]. Sequence [Alb62, Bre85, Bre89, MVW79, Wim74]. Sequences [Ans73, Bra65, Akr88, Alb61b, Bre78, BS88, CD72, Del80, Gal89, GM92, Hal60a, Hal60b, MP99, Nik92, Ort86, SV76b, SV77, Tod88, Tro86, Van94, Her74]. sequential [EKN92]. Serendip [LL85]. serial [DA71, Har91, Nie85]. Series [ES88, CS94, DVZ98, Gab84, GeN80a, GeN80b, KL83, Kun74, Lub88c, MSW92, Nie80, Ren79, SS65, SK73, Sto83, Tho90, Ura67, Wil87, Wyn62, Mar70c]. Set [Att68, Har80, BY66, Car72b, CGC95, ES92, LPDT70, Mal99]. Sets [Red67, VL73, Alb61b, Ale86, Bau66, Bit70, CW79a, KL90, Whi77, Kup71]. seven [Gün75]. seventh [Str68]. Seventieth [Cia98, FV94]. several [Bou90, Gek72, Hsu62, KNR95, Nie91, Tis72, Wei59, Zwa73]. shading [LRT93, Tou92]. Shadowing [CKP95, Had96]. Shafranov [MSF91]. shallow [Red88]. Shape [ABFJ97, LRT93, AL93, CM62, DM85, GM94, NS91, SDK91, Tou92]. Shape-from-shading [LRT93, Tou92]. shape-optimization [SDK91]. shaped [Rei91]. shaping [PR93]. Sharp [BNvW89, Noc88, PP80, Har91, Yam86b]. sharpness [NV80b, SV76b, SV77]. shear [DN86, PS96b]. sheets [CDV99]. shell [BDT88, Man81a, Tro90]. shells [Man83, Pit92, RSSV97]. Shift [RB68, JT70, Leb70]. Shifted [Wit97, Bat90, DT71]. shifting [Lub88c]. shifting-origin [Lub88c]. shock [GW94, Yse99a]. shooting [BHK94, Deu74, DPR76, Deu79, DLO^{+}77, Lam91, Lin93, Lor80]. short [Dub70, Jet84]. shortest [hDxY99, Pyt02, vGS84]. Shortley [Auz88]. side [CR94, JW88, KK76, Sau84, Ste90b]. Sided [VH89, Spi76, Str82, Zha93]. Sides [OP64, Man78a, Tro71]. sieve [Rya67]. Sigmoidal [EV97]. sign [Kem68, KvdD77, Loi68, GK80]. Sign-stability [GK80]. signature [Bro65b]. signed [Str63]. Signorini [Han91a, Spa93]. Silva [Hou73, Chu76, Hou71]. similar [Wuy73]. Similarity [MW68d, Rat82, Por66]. Simple [BRR81b, BCD86, Dan70, SL86, SU89, SW80, Tsu88, Wen68, Zhe89]. simplex [AD64, Bar71, GT78, Hee68, Nag63, Str66]. simplification [Sty79]. simplified [Edd93, GHR97, Sch75b]. Simulating [Gün98]. simulation [Eng96, PR93, Kup71]. simulations [BS99b]. simultanen [Döh82]. Simultaneous [BH73, Car72b, Li97, Rut70, Ste76b, AV86, CR93, Chr70a, Chr70b, Döh82, Els73, Fro88b, Rut69, Sch84]. SINC [Ber89c, Ber89b, Bia90, MBL90]. SINC- [Ber89c, Ber89b]. sine [Bul67]. Single [CG67, AGS98, BS88, LG76a]. single-step [BS88]. Singular [BN78, BGBMN91, Cha73a, CI94, Dag90, GR70a, Lou86, Rei91, Tan71, AKK74, AS97b, Asc85, Atk72, BK87, BS99a, Cas85, Cha92, CVV88, CK82, CNV70, DP85, DP72, DK83, DP63, EMN88, Eis67a, Els84a, FP94, GI99, Gus73, Had87, Hav77, Her90, Her71, HP81, HC63, IJ87, Jam70, Jeg68, LRZ95, Li95, LT68, MS78, MSW91, MSW92, NC93, Nat69b, Nat75a, Neg90, Nii84, Nii85, NS95b, NS96, PR86, PS92b, RvdS90, Red73, RS76, RW85a, SU89, Sch87a, Sch91, SW92, Sco73, Shi78, SO87, gS96b, gS98a, Tan74, Tan92, Tsu88, Van85, Vas88, WA72, Won93, Wri92, Zha96, NY83]. singuläre [Nat75a]. singulãrem [Jeg68]. singulären [Hav77, MS78]. singulärer [Nat69b]. Singularité [LV74b]. singularités [Tho90]. singularities [Gri80, Li86, Li87, MY79, Rei88, Tha76, Tho90, Tsu88, VH93]. singularity [BM96, LV74b, LdD93, Pet93, Sak97]. Singularly [O’R84, Asc85, NY83, O’R87, SO86, SO89, SS95, Ver98b, vV78]. singularly-perturbed [Asc85]. singulière [RW85a]. Sivashinsky [Akr92]. Size [Cry68, Cle71, Hoc81, Yse99b]. sizes [Gol81]. skew [Paa71]. skew-symmetric [Paa71]. slab [MC88a, MC88b]. slip [Ver87a, Ver91a]. slopes [EM87]. slowly [Wyn62]. Small [MW60a, MW60b, Mar90, BL86b, BLR89, BZ82, BB62]. Smallest [MV68a, MV68b, Döh82, GK90, Tod65, Var64]. Smith [Med71]. Smooth [KM91, Ban87, BB97, LL80, MRW76, NR96, Owe77, Sar88, SP94]. Smoothed [LM91]. smoother [Ste94a]. smoothers [Wit89, Wit90]. Smoothing [CW79b, HdH85, Rei67, Rei71b, SB83, Utr85, Wah75a, AL68, AY93, Cul86, DA86, DR75, Dia80, EA88, HS86, Jou71b, LA88, LA89, SS90, dHH87]. smoothings [Par88]. smoothness [AS90, BNvW89]. Sobolev [GZ95, GK85, HM86, MM74, Nak85, Nit69a, Rie91, SA84]. Sobolevschen [Nit69a]. solenoidal [Dob89]. Solid [BB90]. Solution [AU61, Alb61a, Bar69, Bit64, BMPW66, BF66a, CM62, EB68, Ebe70, FR84, GW66, GK66, Har80, HS92, HI71, JK91a, Leg66, Leh61, Leh66, Lig90, MPW66, MW67, Mey83, Moc74, Nab92, OP64, Rei90, Ris74, SS79, Sch72b, SZ91b, Tör63, Tre62, Woź77b, Alb60, AM78, Alt80, Ama72, AM94b, And69, AR68, Ans63, Ans67, Ant72, Arn84, Atk67, Atk72, Atk73, AKS88, BV84, BL86a, BZ85, BP79, BH65, Bla72, BG78, BMR96, BF70, Bos71, Bra82, BHR77, BHR78, Bri83, Bro85, BF66b, BSW88, CR94, Car71, Cas77, Cas80a, Cha98, Cha74, CMR93, CM70, Chr70a, Chr70b, CS88, Cry73, Del73, Deu74, DPR76, DE94, DE96, DV99, DLO^{+}77, Dix82, Dou60, DPRW93, EMN88, Eis67b, EP95, EM91, FŽ87, FG73, FZ93, FL91, Gek72]. solution [GK85, GP65, Gun64a, Had71, Had87, HS80, Han91b, HL64a, Heg91, HW78, Hof75, Hop81, HS70, IU82, JLR70, Joc80, Kac76, KM67, KS71, Kle81a, Kle81b, KO98, KR82, Kra86, Kre68, KS93, LV75, LSW85, Lan87, Lan77, LP78, Lei71, LRT64, LT68, Man81b, Mas87, McC75, Mei86, Mei97, Mel95, Mer67, MM74, Mit78, Mit80, Mit81, Miy76, Mül79a, Mül79b, Mül80, Mül82b, NS81, NW73, OP88, Oli67, Oli68b, Osb65, Osb67, Osb68, Par62a, Par65b, Per66, PS75, Pfl98, RS70, Rak73, Ran84, Raz69, Raz79, RL81, RK84, Ren79, Rhe88, Sah85, SU89, Sar88, Saw82, Sch87d, Sch86, SK73, Sik84, Spi66, Spi76, Spi83, Spr80, Sta70, Ste72, Tha76, Tho75, Ura67, Van79, Vol86, Vui90, Wei92a, Wer68, Wer82]. solution [Wol78b, Yam75, Yam76, Zie85, dHW73, vdGK86, Arn79, Nit76, WH86]. Solutions [BG65, CS59, GR70a, Sre69, Tho90, vdS70b, vdS75, Ale86, ACP99, AS87, Ban87, BYI86, BS92b, Bel83, BP92, Bra69, BRR80, BG71, CD73, Coo71, CW78, DC74, Dre77, Egg86, FF94, FH95, GP79, Gre75, GeN80a, Hac77a, Hac79, Han82, Hüb82, Hwa99, Ish84b, Kar68, Kil78, Koc95, Kos86, Koß81, KM98, Las92, LMW84, Leh67, Lev89, LRT93, Lou79, Mar84, Mar70b, Mat80, McA68, MSW91, MSW92, NW76, Owe77, Pao87, Pao95, Pet70, Ral66, Rei91, RS95, RS88, RN98, Sar89, Sch93, Sch95a, Sch75c, Spr76, Sre71, Ste63, gS97, gS98b, SSBB88, Tan92, Tsu88, WZ95, YN93, Zha93, Zhe89, Zul88]. solvability [KS91]. Solve [CMG90, Pla96]. solved [MS95a]. solver [BL92, BS97b, BG94, CR89, DM96, KS98, SS84, Kir88, NR92]. solvers [Bör89, CKL98, GLC89, Shi79]. solveur [CR89]. Solving [AL85, Bau65b, BL90, Bol84, BM84, GP71, Gol65b, GO88, Gut81, He94, Hig87, HB60, Ise84, JM87, Kin73, Opf82, Tan71, Woź77a, Alb61b, Ans64, Ban78, BWY90, Ben62, BH61b, Bra81a, BRS92, BZ94, Bus77, CT91, Die87, DH80, Dob81, Eyd84, Eyd85, Fer97, FG89, Gir76, GeN80b, Gus73, GNV86, Hac77b, Had69b, HHM76, HP75, Ish81, Ish85, JS95, JN81, JN82a, Kim98a, Krä89, LV84, Li86, Li87, LS87, LH88, MS78, MS72, MS73, MF64, MO91, Müh85, NS95a, NV80a, PV90, Rua88, Sch98, Sep72, She77, Sta77b, Ste73a, Ste90b, Sun95, Tan74, Tee64, TL91, VC92b, Ves76a, VI86, Wil62e, Woo68, ZH93, Zlá69, STR76, Tro76]. Some [BBFS90, BH60b, Ben64, BL90, Bit70, Bro64, Bro68, El 82, FGM67, FP66, Fis61, FH90, GL86, Gar65, Gre62, HS69, Hun72, JTW89, JM78, Kri83, Kul68, Lyn63, May65, NR92, RO70, Sal61, SZ61, Sha67, Sik61, Son93, Ste90a, Ste69b, Ste73b, Str66, Str67, Str68, Ver69, Ves76b, Wer82, de 95, AJS90, All70, Ans64, Asc85, Ask72, Bre72, Bre78, ES90, ELP90, EO99, Fie87, Had69b, HW84, He97, JS95, Kal98, KV85, MNS74, Mar69, McA68, MV80, Mor80, Müh78, Nak85, NN97, NV80b, PW93, Pao87, Pit88, RKL89, RW88, Saa86, Saw82, She92, SG78, Sto83, Tho71, Zen72, Zlá69, Eck68, Gri85, Had69a, Tao84]. Somme [Pic72]. Sommerfeld [Fis93]. Sonderfälle [Rut63b]. SOR [AV76, Ale82a, EKL91, EL88, HP94, NN87, Nic74, Nie89, PHP90]. sorting [Heg94]. source [KjS95, Mey84a]. source-free [KjS95]. sowie [Mül68]. Space [DRW81, KS73, Lau67, LH66, MRMW90, Sac73, SW62, Ale71, BHK^{+}85, BL81, CP76, Dör96, Dou62a, FDD92, GM66b, GN66, KNR95, LPDT70, MBL90, Nak85, PS82, RF94, Rou87, Seg99, Sto64a, TW83, Yam68]. space-adaptive [Dör96]. Space-Homogeneous [MRMW90]. Spaced [Dag90, Cul86, DP88a]. Spaces [Ade69, Ans73, BF90, Kra68, Ley90, LW99, Yse86b, Ale86, AS90, BG78, Boh67, Bro78, CGH76, CK74, CK76, DD74, DDW75, DM88, Gal89, GZ95, GT72, Gue86, IS96, LW82, Mei67, Mey84b, MM74, Mil80, MO91, Nit69a, Pie70, Pie72, Rie91, Rob71, Sal64, SA84, Sar97, Sik82b, Söd86, Vog83b, Yam87, Yse86a, Yse90, Zen84a, Zen84b, Dör70, Sch78c]. span [HL97]. spanning [Wil62e]. Sparse [ADdR89, ER74, GOS99, Pot89, BHK^{+}85, Bjö88, GO94b, Hal92, Hei89b, Mil77, Pfl98, Sad93b, Ste99, Zha96]. spatially [KP91]. Special [Hee68, Rut63b, BJ89, CMW63, Deu79, Dör66, vK66b, Mil65a, Mor80, Sch75b, Sha67, Sid82]. Specific [BD68]. specified [MO68]. SPECT [BS97c]. Spectral [BMM87, CHL85, Dut99, Hei89b, Hen62, PL69, Sac88, BDG97, CMQ84, CDR98, CGR90, Des63, Fis93, Fun88, GH84, HP92, Kub84, Lyn63, MQ81, MP96a, MV69, Mel99, NV80b, OD95, RSSV97, SV90, Sch80b, Wei92b, BK96]. spectrales [BK96]. spectre [Cha73b]. spectrum [Cha73b, EJ81, Rap77, Ste75b]. Speed [RS90, Dix67, Soc88]. Spektralradius [Sch80b]. Spezialfällen [Sch75b]. speziellen [Dör66]. Sphäroiddifferentialgleichung [SG62b]. sphere [Str68, Tis72, Tis72]. spherelike [ST91]. spheres [FP93]. spherically [Gau84, GM86, Kre68]. Spheroidal [SG62b]. Spline [Ben86, BM92, CS90, CS92, CM92, Dag90, Gai70, GM86, IJ87, Nat70, Nit69b, Sch73a, Sch87a, SW99a, AS90, AL68, Ant72, AW85, Att68, Bia91, BD97, Bra71a, CW79b, CS98, DP88a, DGM88, DR75, Dia80, DP82, DM85, El 79, EA88, Fer97, FGM87, vG72, Gün94, HS86, HdH85, KP95, KKL97, LBF98, Mar73, Nie73a, NS83a, NS83b, NR95, Pow94, PR86, Rei84a, Rei84b, RS86, Rei89b, Rei67, Rei71b, Ren80, RF94, SL86, SKL88, Sar88, Sar89, Sch91, Sun98, Utr83, Wah75a, dV93, Att68, Mar73, Mül79a, Mül79b, Mül80, Nat70]. Spline-Approximation [Gai70, Nat70]. spline-functions [AL68]. Spline-Funktionen [Mar73]. spline-in-tension [SKL88]. Spline-Interpolation [Nit69b]. Splineapproximationen [Mül79a, Mül79b, Mül80]. Splinefunktionen [WH86, Ant72]. Splines [KRV74, Tha90, AGS98, AY93, AD74, BS79, Boj93, Cha73a, Cul86, Els84a, Gme87, HS79, HP81, JS91, KK97, LA88, LA89, Luc70, Man72, MP96b, Mül82b, NSSS83, OO98, OO88, Pre71, RS91, SU89, SS90, SV67, ST91, Sha70, SBH84, Tor97, WH86, Won84, Wul73, dHH87, Arn79, Bau76, Mül82b]. Split [Cas83]. Splitting [Pon66, Yse86b, EFM89, EK99, FGM67, KR97, Man70, RM72, Tsa79, Ver84, Yse86a, Yse90]. Splittings [Ale73, HN90, Ple76, AV85, BS97a, CV84a, Els89, FS92, FS99, Gre79, Lyn63, MS90, Sch80b, Son93, Tro76]. spurious [BP92]. square [Axe87, Whi77]. Squares [ADdR89, Bau65b, BG65, Gol65b, GW66, GR70a, GvM91, JO74, Sau78, SS79, vdS75, Bar91b, Bjö88, BD72, BS99c, Cam93, Coh59, EL88, FH87, Gan81, Gir89, Han88, HM89, Kuh93, Läu61, LH88, PW93, PCV96, Pom68, Saa86, Sau84, Sch81c, SW93, Spä86, Sre71, VB92b, VV89, Wei92a]. SSOR [AV86, CW77, LV89]. SSOR-method [AV86]. SSOR-Verfahrens [AV86]. st [Chr70a, Chr70b]. Stabile [Rut63b, vW77b]. stabiles [vW77a]. Stabilität [Bey78, Hav84, NY83, Tro71]. Stabilitätsungleichung [Ess77]. Stabilitätsungleichungen [Gri85]. stabilité [Cro79]. Stability [Ans73, BJ85a, BJ86, BJZ88, BL80, Bol84, CGG90, Cam70, Egg84, ER74, Gri83, Hun86, iS91, JN81, JN82a, JN82b, JK91b, Kot98, Kot99, Lie90, Liu95, MW60b, Mor62, OA75, PW86, PE97, Par62b, RT92, Şch73b, Sch91, Str86, Stu85b, Wat68, Wid66, Woź77b, Woź77a, Zaw91, Zla78, dJ77, vdS70b, vdS75, And71, Aro63a, Aro63b, AM88, Bey78, BN84, BS97c, Bus71, CG88a, Cas82, Cro79, DKS90a, DKS90b, DDW75, Eis67a, EJ81, Gar96, Gek82, Gol63, GHT94, GP85, Gri85, Gug97, GH99, Hai86, Hav84, HL98, Hed65, HS70, HS81, Jac84, JK78, Jel79, JN86, JKR88, KJ90, KvdD77, Loc93, MNS74, Mes95, MS73, Mil72, NY83, OSS90, PR94, QYK99, Rap77, SSS86, Şch75a, Sei63, Spi66, SW84, Stu76, Sül88, TW83]. stability [Tro71, Uhl97, Ver92, Ver77b, Wei95, Zen86, Zen97, Zla81, vV83, vdH72, Ess77, GK80, Tor91]. stability-problem [MS73]. Stabilization [ACR94, SB69, Bar71, Cha98, GLL91]. Stabilized [BD88, AL96, Rie97]. Stable [Gus85, MSS67, Ste98, Wri94, Ziv83, BMX98, Cas81a, FR88, FHLM88, GM66b, GH84, Hai79, Hai80, Hai82, Har78, Mat76, Nev77, Rai69, Sac72, Sav93, Sto83, vW77a, Rut63b, vW77b]. stage [BMPS99, FS92, FS94, Gek72, Gun64b, JT70]. stagewise [Blu61]. staircase [MS99]. Stammfunktion [Fil64]. Standard [Lod70, MW68b]. Starlike [Gri80, BHH88]. stars [GH99]. start [Wam94]. Starting [Sch64a]. state [Alt80, Rou87, SSBB88, Tro71]. state-space [Rou87]. statement [För67, Hav84]. Statements [WH86]. states [MSS67, NMC89, Ste93b]. stationären [Men74]. Stationary [Gut89b, BS97a, Die87, GP83, Had87, Hop81, Men74, Mey83, Tan74, TL91]. statistical [DA86, LV74a]. Statistique [LV74a]. steady [Alt80, SSBB88, Tro71]. steady-state [Alt80, SSBB88, Tro71]. Stefan [AG91, Joc80, Kna85, Mey70, RL81, RK84, Rou87, VV88]. Stefan-like [AG91]. Stefan-type [VV88]. Steffensen [Nie91]. steifer [Saw82]. Stein [May86, Mil80, Rob76]. Steklov [Sch94b]. Stencils [Kul68]. Stenger [Sty79]. Step [Due59, Ker66, AS97a, BC93, BJZ88, Bey87, BS88, Chr70a, Chr70b, CG67, Cur70, DHZ87, GNV86, Gut89b, Hoc81, Jar94, Kir86, Kle81b, KZ94, Man78a, Nic75, NV83, PB81, Sha84, Spi66, Squ69, Sto93, Ura70, Zwa73]. step-by-step [Spi66]. Steplength [Reu91]. stepping [NV96]. steps [CM80, KL83]. Stepsize [CGG90, Arn84, Deu79, Deu83, GHR65, HL84, KR79, Ren85, RRS89, Zla78, Zla81]. stetig [BL67]. stetiger [Ess60]. Stetigkeit [Ans73, Bla77b]. Stetigkeitssätze [MW60b]. Stiefel [BGHS79, Del73, EP99].

Stieltjes’ [Lyn68, HP66, VN93]. stiff [AM88, AF89, BD83, BMR96, BSW88, CGPM98, Cas80b, Cas81b, Cas83, Dah85, HL88, Jel79, KR79, Lub91, Lub92b, NS95b, NS96, Saw82, Sch95b, SWB91]. stiffened [d’H93a]. stiffness [Dör90, Ren85]. Stochastic [Buc62, Nov88, Sch70a, DZ71, Kal74, Zen72]. stochastischer [Bra71b, Sch70a]. Stoer [FV94]. Stokes [AGP92, AM94a, AYB98, BP79, BMM87, Bör90, BD99, BD88, Cam70, CMQ84, Con84, CMG90, Dob81, FF97, FB89, FM92a, FM92b, Gir76, GR79, GP79, Gue94, GQ98, GP83, HL98, JLR70, Jen91, KW99, KP90, Le 80b, LO98, Néd82, Pir82, Por78, QLV91, Rua85, Sac88, SV90, She92, Ste90a, SS96, Sül88, SSBB88, TL91, Ver87a, Ver89, Ver91a, Ver91b, WFS95, Web94, Wit89, Wit90]. Stokes-type [Con84]. storage [Ste76a]. Störungstheorie [WH86]. straight [BD82, Cha97]. strains [BL86b, BLR89]. Strassen [Dum98]. strategies [Kal98, Mag92, Peñ98]. strategy [nJyH99]. stratified [GJ89]. Stream [EP86, AGP92, AYB98, ZL94]. Streamline [Lub92a, JMT97, Nii90, TL91]. streamlinediffusion-finite-element-methods [JMT97]. stress [DN86, Vas88]. Strict [TT87, Bra82]. strictly [Owe77, VN93]. Strong [CEG83, Cro78, JO80]. Stronger [Dej66b, Dej66a]. strongly [AW85, AF89, CM92, Sch91, WhY88]. structural [Gar96, May65, Pot89]. Structure [Loc73, Ans64, Bar84, BMX98, Can71, ESS92, Gas96, IS96, Spi71]. Structured [BH92b, GT82b, Zha96]. Structures [Cia90, Woh63, Bao96, HC92, HS97, d’H93a]. Struktur [Ans64, Bar84, Loc73]. Studies [Bre71]. Study [El 84, LV74a, CR89, Coh59, Coh60, ENV85, Jer89, KW81, Opf82, Pau81a, VV98, Car72a]. studying [DC74]. Sturm [MP95b, Nat69b, Nit69b, AP86, Hal77, Hav77, Her74, Mar90, Pru75, VD91, Wei63]. Sturmsche [Her74]. Stützstellen [Sch80a]. sub [Lan72]. sub-multiplicative [Lan72]. subdivision [DH97]. subdomains [Bad89, Bör89, NN97]. subgradient [Stu89]. Subgroups [Fel61, GL65, Neu61]. subject [Osb67, Osb68]. Sublinear [DK83]. submanifolds [FR84]. Suboptimal [BR91]. subordinate [Bau62]. subproblems [CGT94, Sch81c]. subregions [Dry84]. Subroup [Neu60]. Subsequent [SS79]. subsonic [BH71a]. subspace [HR94, Mor94]. subspaces [Ain96, Ben86, Man84, Pit79, Ste76b, gS93a, gS96b, gS98a]. substance [Gor70]. substance-receiving [Gor70]. substanzerhaltende [Gor70]. substructures [KW96]. Substructuring [AKP95, JP90, Heu98, Heu01, MT96]. successful [SH95]. Successive [BV68, GV61a, GV61b, BK84, Bro64, ENR90, Gal89, Kje61, Kri83, NF69, Sar86, You71]. Sufficient [Sha82, Sty81, Tor91, Wyn59d]. suitable [Heg94, Mor80]. suites [Del80]. Sum [Pic72, Coh59, Hen75, Mar69, Sch68a]. summability [NV83]. Summation [Bar91a, Mar65, DVZ98, FF87]. Summenformel [Mar65]. sums [Bac70, DM90, Kap96, MO68, Sch82]. sup [Dur91]. super [CC85, HW87, Prö91, Gas63]. super-approximation [Prö91]. super-irréductibles [HW87]. superconductivity [Che97, DGL99, GHR97]. superconductors [DGP93]. Superconvergence [Bra94, BN81, CL94, CEL98, DD73b, DG86, Dur90, KN84, PLCC92, Ric78, Tan92, GR86, SJX97, Vol86, Zha92b]. superfast [HJ90]. Superirreducible [HW87]. Superkonvergenz [Vol86]. Superlinear [Han97, Sta81, Rit80, Spr76]. superlinearen [Spr76]. superlinearly [PS82]. Supplement [Fre60]. supplementary [Wer82]. supported [CP94a, DGM88, RS95]. supraconvergence [MS95b]. supraconvergent [GA92]. Supremum [SS62]. Sur-decomposition [Gas63]. surds [Fra62a, Fra62b, Fra63]. sure [AY93]. surface [MW99, RBS82, SDK91, SW92, Str68]. surfaces [CKSS97, Dzi91, HMS99, Hin96, KO98, ST91, Wil62e]. survivable [SD94]. Swiss [HS91]. switching [Sey83]. Symm [JY93]. Symmetric [MBW67, BMRW68, BCS96, BKP76, BNS78, Com59, Cup81, DK77, DH93b, Huc98, KM86, Leh66, MW65, MPW65, MW67, MW68b, Ort63, RB68, AV76, Ale82a, AL76, Ber84, CR94, CKL98, Dub70, ESR82, Gau84, GM86, HP76a, Hal77, HUY95, vK66a, KK97, Kre68, Les73, LR89b, LK78, MRW68, MRW70, MGR87, MS67, Mis84, Moc76, Paa71, PW84, RS63, Rut66b, Rut70, SS76, Sch68c, Sta81, Str73a, Str67, Ves76b, Ves99, Wag97a, Wil62a, Wil62b, Wil62c]. symmetries [Rau86]. symmetrischen [AV76, Com59]. symmetrischer [Leh66]. symmetrizable [EGS94]. symmetrizers [BC92]. Symmetry [Meh67, vFvH73, VC92a]. Symplectic [McL94, MP92, BMX98, Cas96, ESS92, Rei97, jS99, Sto97]. synthesis [Bou90, Mey75]. System [AL85, ADdR89, CMG90, Maî68, MPW66, Tan71, BG71, BEHH75, Dou60, GLC89, Hil83, KP91, LG76b, LV75, Pao85, SFR93, Tan74, Wam94, Wil62e, Yam75, Yam76, Men74]. systematic [MO78, Mey75, Tay69b]. systematically [Cle71]. Systeme [AU61, Bey78, Mar75, LV75, Hil83, Maî68]. Systemen [Ans63]. systèmes [HW87, Leg66]. Systems [ABS84, Alb61a, BH92b, BE73, BMPW66, Bra65, Due59, GP71, GO88, HN90, Hig87, HB60, Kin73, Leg66, LM66, NP78, SV93, Wal67, Woź77b, dBP77, vdS70a, vdS70b, AKK74, AF88a, AFW93, AU62, AV86, Ans63, Aro63a, Aro63b, BD83, BMPS99, Bar91b, BS97b, Bey78, BK98, Bia91, BMR96, BS96, BS97c, Bra82, BRS92, BZ94, BRZRS98, BZS99, BSW88, BM84, Bus77, CR94, CGR94, Cas80b, Cas81b, Cas83, CEG83, CL96a, Deu74, DHZ87, Dre77, ENV85, EMN88, EGS94, EM91, FR88, FGN81, FN91, FZ93, Gas96, GM66a, GM66b, Gov90, GNV86, Gut89b, GH95, Had87, Heg91, HJ90, HW87, HMSW99, HMW86, JK91a, JM87, JS95, JT76, Kac84, KjS95, Kar68, KS92, KL90, LRS91, LV97, LSY87, LS87, Lub89a, Man77, MS96, Mar75]. systems [MS94, MGR87, MT78, Men88, MP92, Mis84, Moc76, MSW91, MSW92, Müh85, NN87, NV83, NS95b, NS96, Pao87, QYK99, Rei97, Roh98, Rus74, Saw82, Sch94a, Sch64a, Sch95b, Sch79b, jS99, SG78, Ste73a, Sto97, Sun98, TT87, TK64, VC92b, VI86, Wer68, Woź78, Wyn66, Zen97, ZH93, Zul88, Tau76]. Szego [Not90, Not93b, Par62a].

T [Bla77b]. T-Approximationen [Bla77b]. table [Cla78, Cuy83a, Gut89a, KZ74, Wyn66]. tableaux [Nag63]. Tables [BP85, Lub88c]. take [ABF99]. Tangent [Kin72]. Tangential [Wag97a, Wag97b]. tau [Sac88]. Taylor [BG76a, BK98, CW71, Hul71, Ren79, Wal98]. Tchebycheff [CG59, Peh90]. Tchebychev [Ho90b, Man77, Man78b]. Tchebyscheff [Sti60]. Technique [Dek71, FH87, Gun64a, Ho90b, Jou71b, KP91, Oli68a, PZ99, RL81, SDK91, Wuy71, Wyn62]. Techniques [DK77, BBFS90, BR88, Boh80, BRZRS98, BG71, BHK94, DPR76, Deu79, ELV91, HW84, HN92, Mei97, Opf80, Opf82, Rei82]. Technische [Ano83b]. Technischen [Sau67]. tée [Rom85]. tées [CR89]. Teil [Hav77]. temperature [CDV99, Zhe89]. Temperaturregelung [Zhe89]. temporal [ADK91]. temporally [KP91]. tension [SKL88]. Tensor [RR90, Ban78, LRT64, SSS99, ST91]. Tensorial [Laf75]. tensorielle [Laf75]. ter [Kos85, Rut60]. term [Cha98, FP90, LM70, Mar73, MS99, vdS72]. termination [Ari98, BS77, Sto77]. terms [BNvW89, Mey84a, Vas88]. test [Loc93, Nie85]. testing [KvdD77]. tests [Fis61, Mat90]. th [ELR92, Rut60]. th-degree [ELR92]. their [Alb61b, Bit70, CK82, CZ98, Chi99, Coc65, FR84, GP65, He97, Hen79, KM91, KM98, Lub94, Mon87, MS95b, Par65b, Rob71, SU89, Sch66b, SW84, Utr83, Won84]. Theodorsen [Gut81, Hüb82]. Theodorsen-Gleichung [Hüb82]. Theorem [GH66, Akr80, Ale71, BDS80, Car91, Dav70, Dei89, Els73, GT84, GK72, Had96, HVW70, LM99, Lub88c, May86, Mil80, Osa96, Osb69, Par62a, Rob76, Sch65, jS99, Ste67, Sto63, Sto64b, Tou92, Tsu99, Wet60, Yam86a, Yam87, ŽV95, wEC86]. théorème [Rob76]. Théorèmes [Lau67]. Theorems [BF60, BD68, SS62, BG76b, Bit70, CDL96, Dic87, GP85, Gro89, Hof75, Hou66, HL60, MS90, MO64, May65, MV80, MN85, Mor72, SV76b, SV77, SSS86, Son93, Str84, Stu85b, Tom84, Lau67]. theoretic [Har62, HL64a]. Theoretical [Pau81a, GS77, MD77, MK80, Per99, VV98]. Theorie [Com59, Hof69, NW73, Sik61, Wyn60, DD84b]. théorique [MK80, Pau81a]. Theory [BV68, CSV69, Com59, DD84b, Geu82, Hof69, Nic74, O’C93, Sik61, Sto77, Ura67, Agu89, AK99, Alb85, BG76a, BY93, BHR77, Bro64, CL71, DLD94, FGS95, FŽ88, Gao89, GO95, HW76, HW84, Ksc68, Lub82, LT68, Mat93b, Mis84, NN87, NW73, NV83, PSS84, Ran79, Sep72, Ser99a, Sun68, Uhl97, Wei63, WH86, Wit90, Wyn60, HV99]. thermomechanical [NS91]. thickness [HC92]. thin [Ain98, KW96, Pow94, RSSV97, Tro90]. thin-shell [Tro90]. Third [LT98b, Bul69a, KM67, Kea68]. Thirteen [Ske86]. Thomas [BCS96]. Thomas-type [BCS96]. Three [Deá80, Smi91, AG88, BDDF87, CKSS97, Dou62a, DSW96, EP86, For85, Han91a, Han94, Heu98, Heu01, HH99, JT70, KZ74, MT78, MS99, Rua85, Son92, Ste98, Zla81]. three-dimensional [EP86, For85, Han94, MT78, Rua85]. three-ordinate [Zla81]. three-point [Son92, Ste98]. three-stage [JT70]. three-term [MS99]. Tightening [Ber59b, Ber60]. Tikhonov [nJyH99, Rie97, Sch93]. Time [LS92, Sch86, Bar91b, BS97b, CGC95, CZ99, Dör96, Fag82, FF94, FN72, Gek78, GS93b, Höh82, JN81, JN82a, JN82b, JK91b, Kau89, Lub94, Mon92, NV96, Pal99, PB81, gS98b]. time- [Dör96]. Time-dependent [Sch86, BS97b, CZ99, Dör96, Fag82, FN72, GS93b, Pal99]. time-harmonic [Mon92]. time-inhomogeneous [Gek78]. time-scale [CGC95]. time-stepping [NV96]. Timoshenko [Li90]. Titzche [Nit68]. Titzschen [Nit68]. TM [Bao96]. Toeplitz [Bar69, BBdH86, Fre94, GH95, HJR88a, Huc98, Laf75, PE97, Qia88, Ris74, Ser99a, Ser99b, Swe84, Swe91, Wim74]. Toeplitz-plus-Hankel [HJR88a]. tomography [Lou86, LR89c, Nat85]. Topological [LH66, Bre75a, Ste75a, Sty79, Sty81]. topologique [Bre75a, LH66]. toroidal [EJ81]. torsion [And71]. torus [CM92]. Total [SB90, BS99c, CL97, Kuh93, PW93, VV89, Wei92a]. Totally [dBP77, Coh59, MP99]. tracing [Chu83]. Tracking [HSY91]. trajectory [SS84]. transcendental [Chu79]. transferable [Lam91]. Transfinite [GH73]. transform [BS59, Bri83, BRS71, CGC95, GL87, GLR88, GMS95, Gut81, Hag87, Hal92, Heg94, LSW85, Lou86, Nat78, Nat79, Rie91, Ros95, Web81]. Transformation [MT78, RB68, Sch73c, Bul69a, Eyd84, Eyd85, GL86, Hom95, Kio79, MSS67, MI82, Nie80, RS63, TM73, Wyn62, BS59]. Transformationen [Sch73c]. Transformations [BG65, Rat82, SW62, Bod68, Bre85, Bre89, Dah85, EV97, LM98, Por66, Tro86, Wim74, dHH87]. Transformed [Dun67, Dun68]. Transforming [Wit89, Wit90]. transforms [GMS95, Gus85, Hek87, Lun85, ST89, Ste66c, Won82]. transient [Ste93b]. transitions [NS91]. translates [HSS96, LTT92, LW99]. Transmission [Sch61b, GS98c, Rei91]. transonic [Ber89a, GW94, GW95a]. Transport [MW99, MRMW90, Ben64, MC88a, MC88b, NV80a, Pir82, WKW97]. transport-diffusion [Pir82]. transportation [Nag63]. transversality [Run70]. transverse [DN86]. Trapezoidal [RS90, AB87, AF89, Kre72b]. Traub [Hag87]. treating [Won92a, Won92b]. Treatment [Bra60, WS59, AU62, Alb85, BS89, BZ82, BS66b, Chu83, Ein71, Heb88, Her71, Kad67, Nat69b, OP81, Web79, Wen68]. triangle [Žen70]. triangles [BG76a]. Triangular [Ben65, Bre70, Cup84, ER74, BG76b, Bra94, Buc64, CGH93, DS91, DMR91, KS99, Man74]. Triangularization [PW70, PE97]. triangularizing [Gre62]. triangulations [FQ99, Gme87]. Tricomi [ASW80, ALS90, GL87, VV84]. Tridiagonal [MBW67, Cup81, KM86, Lew82, RB68, Ber84, Dub70, ESR82, Heg91, Her74, Rei89a, Sac72, Sch87b, Wil62a, Wil62b, Ye95]. Tridiagonalematrizen [Her74]. tridiagonalisation [Zie97]. Tridiagonalization [Sch68c, MRW68, Rai69]. tridiagonals [PW84]. tridimensional [PR93]. Trigonometric [Zur64, DV97, Gau61, Hen79, IM94, Kre71, Kre72a, Müh79, ST91]. trigonometrische [Kre71]. trigonometrischer [Zur64]. Triplex [AKK^{+}68]. Triplex-ALGOL [AKK^{+}68]. trockener [Tau76]. troisième [AG88]. Truncated [LR89c, GT78, Ste99]. Truncation [BJ85b, BJ89, Cha61, Gra68, HP66, JT76, FJ72, Mir71, RW85b]. trust [xY95]. Tschebyscheff [Bla77a, BL67, BS68, Cro77, Dör70, Fil64, För82, Hof69, Hof70, Wer62, Zur64, Bro65a, Cro76, Des61, MW60a, MW60b, Sti59]. Tschebyscheff-Approximation [Bla77a, BL67, BS68, Cro77, Fil64, Hof69, Hof70, Wer62, Zur64]. Tschebyscheff-Approximationen [Bro65a, Cro76, MW60a, MW60b, Sti59]. Tschebyscheff-Typ [För82]. Tschebyscheff-Verfahren [Dör70]. Tschebyscheffschen [Mül68, Nit62]. TSVD [FH95]. Tuning [AGP92, MI82]. turning [GR86, Mor94, Pau81a, Pau81b, PS82, RP85, SO89]. Two [AL93, BDM85, CL61, HV99, Müh85, Sau63, Sha70, Yse90, Zha93, AF88a, AM94a, AM78, AKS88, BMPS99, Bak82, BSV68, Bla72, BF70, Bos71, Bre96, BEHH75, CH71, Cha84, Cha99, CK82, DP85, DPR76, DMGVO96, Dij59, DD73b, DD74, FB89, Fra76, Fra77, Fri70, FS92, FS94, Gab84, GM66b, GeN80b, GN66, Gun64b, yGL93, GC96, HS83, IJ87, JM78, KK97, KZ94, Ley86, LR73, Moc74, Mon84, MMS92, NV80a, Nic75, OA75, Par59, PPV69, Pow94, Red73, RS76, Ren79, Sac88, Sal64, Şch73b, Sch75c, Seo95, Shi79, Spi76, Sto61, Stu89, TIK67, VV98, XS92, Zwa73, Gün75, Hüb72]. Two-dimensional [Sha70, AF88a, CH71, DP85, yGL93, HS83, JM78, Mon84, NV80a, Sac88, Shi79]. two-grid [AM94a]. Two-level [HV99, Bre96]. two-line [Par59]. two-parameter [Seo95]. two-point [AM78, Bak82, Bla72, Bos71, CK82, DPR76, DMGVO96, Fra76, Fra77, GeN80b, GN66, Ley86, OA75, PPV69, Ren79, Şch73b, Sch75c, TIK67, XS92]. Two-sided [Zha93, Spi76]. two-stage [BMPS99, FS92, FS94, Gun64b]. two-step [KZ94]. Typ [För82, Kle81a, Kle81b, Ric74b]. Type [EB68, Fre90, GP71, Jou71a, AL76, AHS84, BRS92, BZS99, BCS96, BS88, CR89, CR74, Cho89, Con84, CDL96, DMGVO96, Ebe70, Egg84, EP95, FGM67, For75, För93, FR86, GMF99, GK66, GLC89, Hal75, Häu86, Hel87, HJR88b, HL99, Huc98, IM94, IK89, JK91a, Kal98, KW81, Kle81a, Kle81b, LLS95, LW82, LMR95, Lev82, Lub88c, Lub89b, MPP99, May86, Mei97, Nii85, NS95b, NS96, NW79, Not90, Not93b, NS83b, OPS76, RRS89, Sab78, Sch81b, Sch81c, Ser99a, Ste66a, Ste66b, Str84, SWB91, Str73b, vI84, vdHtR81, AS97b, BS97b, DGP93, För82, GM77, Hun72, Lev89, Mie79, MI82, RF94, SHH85, VV88, vW77b]. type-II [DGP93]. types [DJ62].

überlinear [PS82]. Überrelaxation [Hüb72]. ultrametric [VN93]. ultrarelativistic [LMGR94]. unabhängigen [Sau63]. unbeschränkten [Bla77a, Bla77b]. unbounded [Bla77a, Bla77b, HB97, Las92]. uncentered [CR80]. uncertainty [LV75]. Unconditionally [Hai79, Hai80]. unconstrained [Brä81b, GLL91, KS89, Kle81a, Kle81b]. uncontrollability [He97]. Uncoupled [SZ91b, Bar93]. Underdetermined [AL85, Mey83]. underrelaxation [CEG83]. unendlich [Zhe89]. ungerader [Mar70c]. Ungleichung [Mei63]. Ungleichungen [Küp78, Sch66b]. Unicity [Str82]. Unified [Mie79, AO93, FF97, Fro88b]. Uniform [CMO91, Her90, Joh69, BZ85, BP66, Buc64, BM92, CM89, DML83, Dej66a, Dej66b, FGS95, Gar65, Gün94, Hel68, Lub89b, Mit73, NY83, O’R87, Pow94, Riv62, SO87]. Uniformly [SO89, Vog83a, McA68, Nii84, SS95, GKR86]. unifying [DC99]. unilateral [Com85]. unique [Atk67, KS91]. Uniqueness [Hal77, Mül68, BS97a, Bro65a, BL67, Cro78, JO80, Kuh93, Pet70]. unit [Tao84, Tre81]. Unitäre [Sch73c]. Unitary [GR90, Faß97, Sch73c]. unité [Tao84]. Units [Neu60]. univalence [KT59, Luk61]. univariate [SH95, vG82]. Universität [Ano83b]. unknowns [CT91, CT93]. unnormalized [FM68, MM67, Met65a, Met65b]. Unrestricted [MM67]. unrestringierter [Kle81a, Kle81b]. unsmoothed [GHT94]. unstable [DH97, Hin96]. unsteady [GQ98]. unstructured [BX96, CSZ96, KNR95]. Unsymmetric [MW67, Jia98, Kri83, Sad93b, Sad93a, Sar86, Wag97b]. untere [Bod68, Kna71]. Untergruppen [Fel61]. Untergruppenverbände [Neu61]. Untergruppenverbandes [GL65, Neu60]. Untersuchung [Neu63]. Untersuchungen [Neu60, Rei71a]. updates [GT82a, GT82b]. Updating [BN78, Cup84]. Upon [Wyn66, Jer89]. Upper [Sch75c, Wei95, Bjö88, BS66a, Küp76b, NV80b, Sto64a, Lin76]. Upperbounds [vdS70c]. upstream [CGH93]. Upwind [Roh98, KNR95]. upwinding [BBFS90]. Use [Hof70, Leh66, Mar65, AS97a, Alb61b, Com59, HP81, Kim98b, Red77, Sch66b, Tha76, Wyn62]. used [Gre79, Peh95]. useful [Cla78, Wil71]. Uses [Had69a]. Using [Bre70, EG94, Not98, SB90, AO93, And69, BO80, BE86, BLLN95, Bra92, Cas80b, CSZ96, Chr70a, Chr70b, CR80, Cle71, CS66, CS68, CFH68, CGT94, Dix82, DD74, GeN80a, Gus85, Had71, HP76a, Han80, JS91, JY93, Jou71b, Küp78, Läu61, MQ89, Mar73, MP96b, MS99, NC93, Sar97, ST91, Sim78, Spe71, Tor97, VV98, Web81, dHH87]. USSOR [LV89, Sar90, Sar96]. utilisant [Tor97]. utilisation [Goë68].

V [Ano83a, Mei63, CSV69, GS95a]. v. [Fis61]. valeurs [Mor74, Par73]. valid [Vog83a]. validation [ACP99, CW79b, Gir89, Luk93]. Value [AU61, Ans69, Ber59b, BN78, BGBMN91, CI94, Cho90, CS59, GR70a, Leh60, Leh61, Sau63, Sik61, Tre62, Wat68, Wer69, vFvH73, Ain98, AU62, AM78, AM94b, AR68, Ans64, Ans67, Arn84, Asc85, AM88, AHS84, AKS88, BO80, BV84, BL92, BL86a, BH61b, BH65, Ber59a, Bey78, Bla72, BS96, BF70, Bos71, Bra66, BHK94, BEHH75, CH71, Cas81a, CVV88, CK82, CŠ79, CSV67, CSV68a, CSV68b, CSV68c, CSV69, CNV70, CM89, Cry73, DP72, DD99, DPR76, DLO^{+}77, Die96, DH80, Dut99, ES90, EP75, EP76, EJSS97, Ess77, EN80, För67, Fra76, Fra77, GS74, Gek74, GK80, GeN80a, GeN80b, GKR86, GN66, Gus73, Hac77b, HHM76, Hei89a, Hof67, Hof75, Hul71, Ish84a, IJ87, Jam70, JN81, JN82a, JN82b, Jer70, Kea68]. value [KS71, Kos85, Kra86, Kre59, Kut75, Lan87, LN99, Ley86, Li86, Li87, LRZ95, Lou86, Lub94, LR73, Lyn85, Mar75, Moc74, Mon84, Mon87, Mül82b, Nev77, NY83, Nit69b, NW76, OA75, PE72, Pao87, Per68, PPV69, Pet70, Red73, RS76, Ren79, Reu88a, RW88, RS72, Rus74, RW85b, Sán75, Şch73b, SS88, Sch66b, ST73, Sch75c, STR76, Sch84, Sch71b, Sch72b, Ser99b, Sha68, Spi66, Spi76, Spi83, Spr76, Sta77a, Sta77b, Ste63, Ste68, SO86, Tro76, TIK67, Ura67, Van85, Vol86, Vos75, Wal97, Wen68, Wes75, Wit78b, Wit78c, Won93, Wri92, Wri94, XS92, Zlá69, dO70, vdGK86]. value-vector [AR68]. value/alternative [Sán75]. Valued [BGBMN91, Car72b, GM83, Joh69, KM91]. Values [Hen62, Bau62, Bau68, FP94, GS95a, HC63, Hun72, Hun73, JTW89, Jeg68, KS99, LM75, NS64, Rob72, Sta97, Zen68, Zha96]. values/vectors [Zha96]. Vandermonde [BH92b, BE73, CGR90, EFP98, GP71, Gau62, Gau63, Gau75a, Gau75b, Gau78, GI88, GO94a, Hig87]. Vandermonde-like [BH92b]. vanishing [KK97]. Varga [Cia98]. Variable [CGG90, Str63, Wer78, Ziv83, Brä81b, GM68, GT82b, Gri83, JT70, Jov89, LM98, Pal99, Rit80, RF94, SW99a, TM73, TW83, Yse99b, Zla78, Zla81, Rei75]. variable-coefficient [SW99a]. Variable-Metrik-Verfahren [Wer78]. variable-shift [JT70]. Variable-signed [Str63]. Variablen [Rei75]. Variables [Sau63, Ben62, BP79, BSV68, BDDF87, Cha84, Dou62a, GM66b, Hsu62, MH72, Nie91, Sal64, Sch99, Stu89, Tis72, Tsu72, Tsu73, Tsu75, Wei59, Bro65b]. Variance [Bar91a]. Varianten [Kle81b]. variants [Atk73, Kle81b]. Variation [Hen62, CL97, FDD92, Zet70]. Variational [Bit68, Bit71, Cia90, RG88, And69, Arn82, BM87, BHR77, BHR78, BH74, BS85, CC91, CV70, CD85, Dan70, DB77, Eyd84, Eyd85, Fet87, FG77, For72, HM83, Klo99, KW95, Kor94, Kor96, Man72, Mit78, Noc89, Ped96, PSS90, Rum96, Run70, RN98, Sha68, Spa94, Vui90, Žen89]. variationally [Kol78, SG78]. variations [HP92]. Variationsbereich [Zet70]. Variationsmethoden [Bit68, Bit71]. Variationsungleichung [BH74]. various [Gut81, Nit70]. varying [Bar91b, GHR65, KW96, KP91]. Vector [Bar69, GM83, Kra68, KS73, AR68, Cor77, GM92, GN66, Heg94, JS95, Joh69, Ver87b, Wyn64]. vector-valued [Joh69]. vectorial [Van94]. vectors [CR94, EP86, Sch64a, Ste84, Zha96]. Vektorräumen [Kra68]. Vektorraums [KS73]. velocity [BN84, Rua85]. velocity-pressure [BN84, Rua85]. verallgemeinerten [Mey84b]. verallgemeinerter [Arn80]. Verallgemeinerung [Ans66, Mar65, Mei63]. Veränderlichen [Sau63, Bro65b]. Verbesserung [Ale82b, Häm65, Sch61e]. Vereinfachte [Sch75b]. Verfahren [AH72, Ans63, Ber59b, BS60, Cro76, Dol81, Dör70, GL65, Kle81a, Leh60, Mey84b, Nie73b, PS82, Saw82, SS66, Sch61a, SEG62, Sch68b, SV70, Sto67, Wag76, Wer78, Wer79, Wit78b, Wit78c, Zet70, Zur63, Bra77, Col61, For72, Had68, Kle81a, Kle81b, Nit69b, Nit76, Sch66a]. Verfahrens [Ale82b, Fil64, Mer91, Ale76, AV86, Nit68, Sch61c]. Vergleichsfaktor [Aum63]. Verhalten [Fie87, Ort86]. verification [RN98]. verifications [YN93]. Verschärfung [Ber60, Ber59b]. verschiedener [Nit70]. version [AK99, BS89, GB86a, GB86c, Guo88, GC96, GS98c, Heu98, HMS99, Heu01, Jen91, MP96a, OD95, Pav94a, Pav94b, Sac72, Vog83a, Vog83b]. versions [BD81, GB86a, GB86b, GB86c, Li90, LM99]. Verteilung [NW73]. vertices [MN68]. Verwendung [Alb61b, Leh66]. very [Ruh70, Tsu72, Tsu73, Tsu75, Wil72]. Verzweigungsprobleme [Sch77a]. Verzweigungsproblemen [Web79]. VI [Ano86]. via [CL97, DD73b, Fel61, KP90, MS85, O’C96, Opf80, QLV91, Sch90, Ser99a]. Vibrations [Wer69, Tau76]. vicinity [MY79]. vielen [Zhe89]. Vierecken [Gai72]. vierter [Nat69a]. viewpoint [Moo88]. Vincent [Akr80]. Virgule [Pic72]. viscoelastic [BS92b, LR89a, NS95a]. viscosité [BN90]. viscosity [FF94, Koc95, LRT93, Sch95a]. Viscous [EK99, ABF99, Heb88, Man84, QLV91, Tau76]. voisinage [Tho90]. vollständigen [GL65]. Volterra [BJ85a, Bru78, BN81, DMJ86, Egg86, Egg88, Hoc79, Hoc81, IK89, Jac79, Kau89, Lub82, Mat76, Rak73, SW96, Ste72, Tan92, Ver88, WA72, dHW73, dHW75, vdHtR81]. Volume [KjS95, BC91, BW93, Cai91b, CGH93, Cha99, EGH99, KMS99, KNR95, Mis99, PS96a, Roh98, SS98, VV98]. volume-matching [BC91]. Volume-preserving [KjS95]. Volumes [Ano83c]. vom [Kle81a, Kle81b, Nit62, Ric74b]. Voraussetzungen [Sch75d]. Vorticity [AGP92, AYB98, yGL93, ZL94]. Vorzeichenstabile [GK80]. Vries [GA92, dFSS97].

W [Had72a]. Wachspress [AAG79]. Wachstrum [Rei71a]. walk [HI71]. Wanner [RRS89]. Watther [Sau67]. wave [BS97b, CMR93, DG86, McL94, Mor62, TV98, Yan94]. waveguides [BP92]. Wavelet [WZ95, vPS96, Ber97, CT93, CGC95, Rie91, Rie97, Ste98, SP94, XS92]. wavelet-like [CT93]. waves [GJ89, Kim98a, Wah75b]. ways [Ske86, Sto83]. Weak [Ros75, Stu76, Bra69, Els89, MS90, Sch75d, Hav77]. Weak-element [Ros75]. weakly [GKR86, Kos85, Roh98, Tan92]. Weber [Coh60]. wedge [Rei91]. wedge-shaped [Rei91]. Weeks [GLR88]. Weierstrass [Jar94, PCT95, Tod90]. Weight [GG90, DB77, Lev82, Not90, Not93b, Sal75, Str63]. Weighted [Bau65b, FS99, DN89, GK94, Lu97, Pec74, Sal63, vG82]. Weights [KE82, DD99, Mon87, Peh88]. Weinberger [Mei86]. Weiser [DR92]. Well [Mül68, Uhl92, CCK95, GP85, Lam91, Lus96]. Well-Conditioned [Uhl92]. well-posed [Lam91]. well-posedness [GP85]. Weller [Auz88]. wells [Chi99]. Wendroff [Hed65, MO78, Zwa73]. Wert [Jeg68, Sik61]. Wertebereich [vF70]. Wetterling [Had72a]. Weyl [Bit91]. where [Har80]. Which [Bus68, Nab92, FR86, Gue86, Por66, Sre69, Sre71, Wal97, Wyn66]. whose [Boe64, vI84]. Wielandt [LM99, Loi67]. Wiener [CG88b]. Wilf [Eck68]. Wilfscher [Eck68]. Wilkinson [BH87, PC83]. Wilson [CL94, LZ80, Sch81b, Sch81c, cS84, SJX97]. Wing [RS65]. Winograd [Bre70]. without [AH72, BP92, Brä81b, CPT98, Car72a, Eng77, Häm65, SO89, Wer78]. work [Dir69, Had72a]. worst [gS95b]. worst-case [gS95b]. wrap [Heg91]. wrap-around [Heg91]. Wurzeln [Nic66a].

X [Ano98a, Ano98b, Lou86, LR89c]. X-ray [Lou86, LR89c].

Yamamoto [Dav70]. Yuan [Pyt02].

zäher  [Tau76].  Zahlen

[Kar61]. Zählerpolynom [Lam70]. Zeilensummenkriterium [Sch68a]. Zerfallende [Her74]. Zerlegung [KS73]. Zerlegungen [Ale73, Kul68, Sch80b]. Zero [CGG90, Zla81, HC92, PCT95, SV76b, SV77, Ver77a]. zero-finding [PCT95]. zero-free [SV76b, SV77]. zero-parasitic [Ver77a]. Zerostability [Zla81]. Zeros [BS63, BS70, Dir69, FS70, Ker66, AH72, BL76, BH73, Car92, Chu79, Coc65, DD78, Elb91, Els73, Fro88b, GH72, Gut72, IK77, JT70, Kap96, Kio79, MK60, MS95a, SV75, SV76a, SV78, Van82, VC94, Wer79, Wil59a, Wil59b, YK88, Gün75]. Zolotukhin [BMP81]. zum [Ber59b, Bra73, Due59, For68a, For72, Had72b, Leh60, Zet70, Zur63, Lod70, NO68, Sch68a, Sch70b, Wet60]. zur [AU61, Alb61a, Alb61b, Ale82b, Ans63, Ans67, Arn80, Bit64, BS60, Bra70, Bul64, Cro76, Cro77, Dir69, Döh82, Dre77, ES73, For68a, För82, GL65, Häm65, Hav84, Jet84, Ker66, Kle81a, Kle81b, Koß81, KS73, Mit73, Mül79a, Mül79b, Mül80, Neu63, Nie73b, PS82, Rut60, Sau78, SS66, Sch61a, SEG62, Sch64b, Sch75b, Sto61, Sto67, Tör63, Wag76, Wag80, Weg78, Wer79, Ale76, AV76, Ale82b, AV86, Ans64, Ans66, BS59, Bey78, Bla77b, Bra61, BL67, BS68, Ehr69, For75, For78, Häm66, Hof69, Jeg68, KR82, Kre72b, Lan87, Leh66, Loc73, Lor77, Mar70c, MS78, Mer91, Mül68, Nit70, PK64, Pau77, Ric74b, Sau84, SG62a, SG62b, Sch61c, Zur64]. Zusammenhang [WB69]. zwei [Sau63, Sto61]. Zweidimensionale [Gün75]. Zweiparametrige [Hüb72]. zweiter [Het74, Het75, Küp78, Nat69a, Spr76]. zyklisch [KS73]. zyklisches [Dol81]. Zylindersymmetrie [vFvH73].

References

Achchab:1998:EEP

[AABM98] B. Achchab, A. Agouzal, J. Baranger, and J. F. Maitre. Estimateur d’erreur a posteriori hiérarchique. application aux éléments finis mixtes (French) [A posteriori hierarchical error estimator. application to mixed finite elements]. Numerische Mathematik, 80(2): 159–179, August 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                                   URL http://link.springer.
                                                   de/link/service/journals/
                                                   00211/bibs/8080002/80800159.
                                                   htm;   http://link.springer.
                                                   de/link/service/journals/
                                                   00211/papers/8080002/80800159.
                                                   pdf;   http://link.springer.
                                                   de/link/service/journals/
                                                   00211/papers/8080002/80800159.
                                                   ps.gz.

Apprato:1979:EFR

[AAG79] D. Apprato, R. Arcangéli, and J. L. Gout. Sur les éléments finis rationnels de Wachspress. (French) [On the finite rational elements of Wachspress]. Numerische Mathematik, 32(3): 247–270, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anderssen:1974:NDP

[AB74] R. S. Anderssen and P. Bloomfield. Numerical differentiation procedures for non-exact data. Numerische Mathematik, 22(3):

157–182, June 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Andersson:1984:NOQ

[AB84] Jan-Erik Andersson and Borislav D.

Bojanov. A note on the optimal quadrature in H^{p}. Numerische Mathematik, 44(2):301–308, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Akrivis:1985:FNG

[AB85] G. Akrivis and A. Burgstaller.

Fehlerabschätzungen für nichtsymmetrische Gauß-Quadraturformeln. (German) [Error estimates for nonsymmetric Gaussian quadrature formuale]. Numerische Mathematik, 47(4):535–543, December 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Allasia:1987:NCI

[AB87] Giampietro Allasia and Renata Besenghi. Numerical calculation of incomplete gamma functions by the trapezoidal rule. Numerische Mathematik, 50(4): 419–428, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Angot:1999:PMT

[ABF99] Philippe Angot, Charles-Henri Bruneau, and Pierre Fabrie. A penalization method to take into account obstacles in incompressible viscous flows. Numerische Mathematik, 81(4):497–520, February 1999.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9081004/90810497.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9081004/90810497.pdf.

Allaire:1997:SOH

[ABFJ97] Grégoire Allaire, Eric Bonnetier, Gilles Francfort, and François Jouve. Shape optimization by the homogenization method. Numerische Mathematik, 76(1): 27–68, March 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076001/70760027.
htm;   http://science.springer.
de/nmee/bibs/7076001/70760027.
htm.

Abaffy:1984:CDM

[ABS84] Joszef Abaffy, Charles G. Broyden, and Emilio Spedicato. A class of direct methods for linear systems. Numerische Mathematik, 45(3):361–376, December 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ainsworth:1992:PEE

[AC92] Mark Ainsworth and Alan Craig. A posteriori error estimators in the finite element method. Numerische Mathematik, 60(4):429–463, January 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Axelsson:1994:NGC

[AC94] O. Axelsson and A. T. Chronopoulos. On nonlinear generalized conjugate gradient methods. Numerische Mathematik, 69(1): 1–15, November 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069001/40690001.
htm;   http://science.springer.
de/nmee/bibs/4069001/40690001.
htm.

Akrivis:1999:IEM

[ACM99] Georgios Akrivis, Michel Crouzeix, and Charalambos Makridakis. Implicit-explicit multistep methods for quasilinear parabolic equations. Numerische Mathematik, 82(4):521–541, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082004/90820521.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082004/90820521.pdf.

Alefeld:1999:NVS

[ACP99] G. E. Alefeld, X. Chen, and F. A. Potra. Numerical validation of solutions of linear complementarity problems. Numerische Mathematik, 83(1):1–23, July 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083001/90830001.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083001/90830001.pdf.

Ascher:1994:SDI

[ACR94] Uri M. Ascher, Hongsheng Chin, and Sebastian Reich. Stabilization of DAEs and invariant manifolds. Numerische Mathematik, 67(2): 131–149, March 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/4067002/40670131.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4067002/40670131.
                                        htm.

Azpeitia:1964:DRS

[AD64] A. G. Azpeitia and D. J. Dickinson. A decision rule in the simplex method that avoids cycling. Numerische Mathematik, 6(1): 329–331, December 1964. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anselone:1973:PBA

[AD73] Philip M. Anselone and Joel Davis. Perturbations of best approximation problems. Numerische Mathematik, 21(1):63– 69, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Astor:1974:DS

[AD74] P. H. Astor and C. S. Duris.

Discrete L-splines. Numerische Mathematik, 22(5):393–402, October 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alouges:1993:DPD

[AD93] François Alouges and Arnaud Debussche. On the discretization of a partial differential equation in the neighborhood of a periodic orbit. Numerische Mathematik, 65(2): 143–175, June 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arioli:1989:ASA

[ADdR89] Mario Arioli, Iain S. Duff, and Peter P. M. de Rijk. On the augmented system approach to sparse least-squares problems. Numerische Mathematik, 55(6): 667–684, August 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ade:1969:IHO

[Ade69] Hans Ade. Iterationsverfahren höherer Ordnung in Banach-Räumen. (German) [Iteration methods of higher-order in Banach spaces]. Numerische Mathematik, 13(1):39–50, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arnold:1984:FHO

[ADG84] Douglas N. Arnold, Jim Douglas, Jr., and Chaitan P. Gupta. A family of higher order mixed finite element methods for plane elasticity. Numerische Mathematik, 45(1):1–22, October 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Akrivis:1991:FDG

[ADK91] Georgios D. Akrivis, Vassilios A. Dougalis, and Ohannes A. Karakashian. On fully discrete Galerkin methods of secondorder temporal accuracy for the nonlinear schrödinger equation. Numerische Mathematik, 59(1): 31–53, April 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Adjerid:1988:SOF

[AF88a] Slimane Adjerid and Joseph E. Flaherty. Second-order finite element approximations and a posteriori error estimation for two-dimensional parabolic systems. Numerische Mathematik, 53(1/2):183–198, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arnold:1988:NMF

[AF88b] Douglas N. Arnold and Richard S. Falk. A new mixed formulation for elasticity. Numerische Mathematik, 53(1/2):13–30, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Auzinger:1989:AEE

[AF89] W. Auzinger and R. Frank.

Asymptotic error expansions for stiff equations: an analysis for the implicit midpoint and trapezoidal rules in the strongly stiff case. Numerische Mathematik, 56(5):469–499, November 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Adjerid:1993:PEE

[AFW93] Slimane Adjerid, Joseph E. Flaherty, and Yun J. Wang. A posteriori error estimation with finite element methods of lines for one-dimensional parabolic systems. Numerische Mathematik, 65(1):1–21, May 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Amrouche:1988:MAM

[AG88] Chérif Amrouche and Vivette Girault. Une méthode d’approximation mixte des équations des fluides non newtoniens de troisième grade. (French) [Mixed approximations of the equations of non-Newtonian fluids of grade three]. Numerische Mathematik, 53(3): 315–349, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Amiez:1991:NAS

[AG91] Geneviève Amiez and Pierre-Alain Gremaud. On a numerical approach to Stefan-like problems. Numerische Mathematik, 59(1):71–89, April 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Achdou:1992:TMM

[AGP92] Yves Achdou, Roland Glowinski, and Olivier Pironneau. Tuning the mesh of a mixed method for the stream function. vorticity formulation of the Navier–Stokes equations. Numerische Mathematik, 63(2): 145–163, November 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Abaffy:1987:LCA

[AGS87] József Abaffy, Aurél Galántai, and Emilio Spedicato. The local convergence of ABS methods for nonlinear algebraic equations. Numerische Mathematik, 51(4): 429–439, August 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ainsworth:1998:SGA

[AGS98] Mark Ainsworth, Rolf Grigorieff, and Ian Sloan. Semidiscrete Galerkin approximation of the single layer equation by general splines. Numerische Mathematik, 79(2): 157–174, April 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079002/80790157.
htm;   http://science.springer.

de/nmee/bibs/8079002/80790157.
htm.

Aguilera:1989:NAC

[Agu89] Néstor E. Aguilera. Numerical analysis of cavitational flowtheory. Numerische Mathematik, 55(1):1–32, April 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1972:IZN

[AH72] G. Alefeld and J. Herzberger.

Nullstelleneinschliessung Mit Dem Newton-Verfahren Ohne Invertierung Von Intervallmatrizen. (German) [Including zeros of nonlinear equations by the Newton method without inverting interval matrices]. Numerische Mathematik, 19(1):56– 64, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aziz:1984:FEM

[AHS84] A. K. Aziz, Houde Han, and M. Schneider. A finite element method for a boundary value problem of mixed type. Numerische Mathematik, 44(2): 191–200, August 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ainsworth:1996:ISS

[Ain96] Mark Ainsworth. The influence and selection of subspaces for a posteriori error estimators. Numerische Mathematik, 73(4): 399–418, June 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073004/60730399.
htm;   http://science.springer.
de/nmee/bibs/6073004/60730399.
htm.

Ainsworth:1998:PEE

[Ain98] Mark Ainsworth. A posteriori error estimation for fully discrete hierarchic models of elliptic boundary value problems on thin domains. Numerische Mathematik, 80(3):325–362, September 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080003/80800325.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800325.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800325.
ps.gz.

Akhlaghi:1990:ESC

[AJS90] M. R. Akhlaghi, A. Jakimovski, and A. Sharma. Equiconvergence of some complex interpolatory polynomials. Numerische Mathematik, 57(6/7): 635–649, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Ainsworth:1999:ATP

[AK99] Mark Ainsworth and David Kay. The approximation theory for the p-version finite element method and application to non-linear elliptic PDEs. Numerische Mathematik, 82(3):351–388, May 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082003/90820351.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082003/90820351.pdf.

Apostolatos:1968:ALT

[AKK^{+}68] N. Apostolatos, U. Kulisch, R. Krawczyk, B. Lortz, K. Nickel, and H. W. Wippermann. The algorithmic language Triplex-ALGOL 60. Numerische Mathematik, 11(2):175–180, February 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). From the first page: “triplex numbers, previously called Fehlerschrankenzahlen by K. Nickel and interval numbers by R. E. Moore. … Triplex numbers are intervals with a main value enclosed, which corresponds to conventional numbers of real type. The first number following [ is the lower bound of the interval associated with a triplex number, the next number is the previous mentioned main value, and the last number followed by ] is the upper bound of the interval. The following relation has to be fulfilled: lower bound ≤ main value ≤ upper bound.”.

Abrahamsson:1974:DAS

[AKK74] L. R. Abrahamsson, H. B. Keller, and H. O. Kreiss. Difference approximations for singular perturbations of systems of ordinary differential equations. Numerische Mathematik, 22(5): 367–391, October 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Achdou:1995:SPM

[AKP95] Y. Achdou, Yu. A. Kuznetsov, and O. Pironneau. Substructuring preconditioners for the Q_{1} mortar element method. Numerische Mathematik, 71(4): 419–449, October 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/5071004/
                                        50710419.htm;   http://science.
                                        springer.de/nmee/bibs/5071004/
                                        50710419.htm.

Akritas:1980:IVT

[Akr80] Alkiviadis G. Akritas. An implementation of Vincent’s theorem. Numerische Mathematik, 36(1): 53–62, December 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Akrivis:1984:FFG

[Akr84] G. Akrivis. Fehlerabschätzungen für Gauß-Quadraturformeln.

(German) [Error estimates for Gaussian quadrature formulas]. Numerische Mathematik, 44(2): 261–278, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Akritas:1988:NMC

[Akr88] Alkiviadis G. Akritas. A new method for computing polynomial greatest common divisors and polynomial remainder sequences. Numerische Mathematik, 52(2):119–127, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Akrivis:1992:FDD

[Akr92] Georgios D. Akrivis. Finite difference discretization of the Kuramoto–Sivashinsky equation. Numerische Mathematik, 63(1):1–11, October 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aziz:1988:TPB

[AKS88] A. K. Aziz, R. B. Kellogg, and A. B. Stephens. A two point boundary value problem with a rapidly oscillating solution. Numerische Mathematik, 53(1/2): 107–121, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anselone:1968:GMC

[AL68] P. M. Anselone and P. J. Laurent. A general method for the construction of interpolating or smoothing spline-functions.

Numerische Mathematik, 12(1): 66–82, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anderson:1976:JTM

[AL76] P. J. Anderson and G. Loizou. A Jacobi type method for complex symmetric matrices. Numerische Mathematik, 25(4): 347–363, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arioli:1985:EAA

[AL85] M. Arioli and A. Laratta. Error analysis of an algorithm for solving an underdetermined linear system. Numerische Mathematik, 46(2):255–268, May 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Axelsson:1986:EDC

[AL86a] Owe Axelsson and Gunhild Lindskog. On the eigenvalue distribution of a class of preconditioning methods. Numerische Mathematik, 48(5): 479–498, May 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Axelsson:1986:RCP

[AL86b] Owe Axelsson and Gunhild Lindskog. On the rate of convergence of the preconditioned conjugate gradient method. Numerische Mathematik, 48(5): 499–523, May 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Archer:1993:TSP

[AL93] J. C. Archer and E. Le Gruyer.

Two shape preserving Lagrange C^{∈}-interpolants. Numerische Mathematik, 64(1):1–11, January 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Apel:1996:AMR

[AL96] Thomas Apel and Gert Lube.

Anisotropic mesh refinement in stabilized Galerkin methods. Numerische Mathematik, 74(3): 261–282, September 1996. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/6074003/
60740261.htm;   http://science.
springer.de/nmee/bibs/6074003/
60740261.htm.

Albrecht:1960:ASP

[Alb60] Rudolf F. Albrecht. Approximation to the solution of partial differential-equations by the solution of ordinary differentialequations. Numerische Mathematik, 2(1):245–262, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Albrecht:1961:FRN

[Alb61a] J. Albrecht. Fehlerabschätzungen bei Relaxationsverfahren zur numerischen Auflösung linearer Gleichungssysteme. (German)

[Error bounds for the relaxation method for numerical solution of linear equation systems]. Numerische Mathematik, 3:188– 201, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Albrecht:1961:MII

[Alb61b]  J. Albrecht.  Monotone It-

erationsfolgen und ihre Verwendung zur Lösung linearer Gleichungssysteme. (German) [monotonic iteration sequences and their use in solving sets of linear equations]. Numerische Mathematik, 3(1):345– 358, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
345.

Albrecht:1962:FKM

[Alb62]  Julius Albrecht.  Fehler-

schranken und Konvergenzbeschleunigung bei einer monotonen oder alternierenden Iterationsfolge. (German) [Error bounds and convergence acceleration with a monotone or alternating iteration sequence]. Numerische Mathematik, 4:196–208, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Albrecht:1978:OCM

[Alb78] P. Albrecht. On the order of composite multistep methods for ordinary differential equations. Numerische Mathematik, 29(4):381–396, April 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Albrecht:1985:NTO

[Alb85]  Peter Albrecht.  Numerical

treatment of ODEs: the theory of A-methods. Numerische Mathematik, 47(1):59–87, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1971:AFP

[Ale71]  G. Alefeld.  Anwendungen

Des Fixpunktsatzes Für Pseudometrische Räume In Der Intervallrechnung. (German) [Application of the fixed point theorem for pseudo-metric space to interval computation]. Numerische Mathematik, 17(1):33– 39, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1973:UKZ

[Ale73]  G. Alefeld.  Über konver-

gente Zerlegungen von Matrizen. (German) [On convergent splittings of matrices]. Numerische Mathematik, 20(4): 312–316, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1976:CPI

[Ale76]  G. Alefeld.  Zur Konver-

genz des Peaceman–Rachford-Verfahrens. (German) [On the convergence of the Peaceman– Rachford iterative method]. Numerische Mathematik, 26(4): 409–419, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1982:CSS

[Ale82a] G. Alefeld. On the convergence of the symmetric SOR method for matrices with red-black ordering. Numerische Mathematik, 39(1):113–117, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1982:KVD

[Ale82b] G. Alefeld. Zur Konvergenz eines Verfahrens von D. J. Evans zur iterativen Verbesserung einer Näherung für die inverse Matrix. (German) [On the convergence of D. J. Evans’ implicit matrix inversion process]. Numerische Mathematik, 39(2): 163–173, August 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1986:CIE

[Ale86] G. Alefeld. Componentwise inclusion and exclusion sets for solutions of quadratic equations in finite-dimensional spaces. Numerische Mathematik, 48(4): 391–416, April 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Allgower:1970:CPD

[All70] E. L. Allgower. Criteria for positive definiteness of some band matrices. Numerische Mathematik, 16(2):157–162, November 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Allgower:1973:EIC

[All73] E. L. Allgower. Exact inverses of certain band matrices. Numerische Mathematik, 21 (4):279–284, August 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1997:MMB

[ALM97] Götz Alefeld, Ingrid Lenhardt, and Günter Mayer. On multisplitting methods for band matrices. Numerische Mathematik, 75(3):267–292, January 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7075003/
70750267.htm;   http://science.
springer.de/nmee/bibs/7075003/
70750267.htm.

Alonso:1996:EEM

[Alo96] A. Alonso. Error estimators for a mixed method. Numerische Mathematik, 74(4):385–395, October 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074004/60740385.

                                        htm;   http://science.springer.
                                        de/nmee/bibs/6074004/60740385.
                                        htm.

Aziz:1990:FEM

[ALS90] A. K. Aziz, R. Lemmert, and M. Schneider. A finite element method for the nonlinear Tricomi problem. Numerische Mathematik, 58(1):95– 108, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alt:1980:NSS

[Alt80] Hans Wilhelm Alt. Numerical solution of steady-state porous flow free boundary problems. Numerische Mathematik, 36(1): 73–98, December 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Allgower:1978:NMM

[AM78] E. L. Allgower and S. F. Mc-Cormick. Newton’s method with mesh refinements for numerical solution of nonlinear twopoint boundary value problems. Numerische Mathematik, 29(3): 237–260, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ascher:1988:GFS

[AM88] U. M. Ascher and R. M. M. Mattheij. General framework, stability and error analysis for numerical stiff boundary value methods. Numerische Mathematik, 54(3):355–372, December 1988. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

AitOuAmmi:1994:NGM

[AM94a] Ali Ait Ou Ammi and Martine Marion. Nonlinear Galerkin methods and mixed finite elements: two-grid algorithms for the Navier–Stokes equations. Numerische Mathematik, 68(2): 189–213, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068002/40680189.
htm;   http://science.springer.
de/nmee/bibs/4068002/40680189.
htm.

Amodio:1994:BVM

[AM94b] P. Amodio and F. Mazzia.

Boundary value methods for the solution of differential-algebraic equations. Numerische Mathematik, 66(4):411–421, January 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Amann:1972:NLN

[Ama72] Herbert Amann. Über die nährungswise Lösung nichtlinearer Integralgleichungen. (German) [On the approximate solution of nonlinear integral equations]. Numerische Mathematik, 19(1):29–45, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anderssen:1969:NSP

[And69] R. S. Anderssen. The numerical solution of parabolic differential equations using variational methods. Numerische Mathematik, 13(2):129–145, May 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anderssen:1971:SAE

[And71] R. S. Anderssen. A stability analysis for the extended Kantorovich method applied to the torsion problem. Numerische Mathematik, 17(3): 239–246, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anderson:1988:MRE

[And88] Ned Anderson. Minimum relative error approximations for 1/t. Numerische Mathematik, 54(2):117–124, November 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anonymous:1983:AAH

[Ano83a]  Anonymous.  Announcemen:

Alston S. Householder Award V (1984). Numerische Mathematik, 42(2):258, November 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anonymous:1983:AIC

[Ano83b] Anonymous. Announcing an International Conference on Numerical Analysis on the Occasion of the 25th Anniversary of the founding of Numerische Mathematik, March 19–21, 1984 at the Technische Universität Munich. Numerische Mathematik, 42(1):124, 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anonymous:1983:AIV

[Ano83c] Anonymous. Author index: Volumes 1–40 (1959–1982). Numerische Mathematik, 41(3): I–LII, July 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). This appears as a separatelynumbered section at the end of the issue.

Anonymous:1984:AIC

[Ano84] Anonymous. Announcing an International Conference on Numerical Analysis. Numerische Mathematik, 43(2):328, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anonymous:1986:AAH

[Ano86] Anonymous. Announcement:

Alston S. Householder Award VI (1987). Numerische Mathematik, 49(4):459, July 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anonymous:1998:AHAa

[Ano98a] Anonymous. Announcement:

Householder Award X. Numerische Mathematik, 81(1):

161–162, November 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anonymous:1998:AHAb

[Ano98b] Anonymous. Announcement: Householder Award X. Numerische Mathematik, 81(2): 321–322, December 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anonymous:1999:PDO

[Ano99] Anonymous. Publications in the directory ONLINE First. Numerische Mathematik, 83(2): 177, August 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ansorge:1963:AVC

[Ans63]  R. Ansorge.  Die Adams–

Verfahren als Charakteristikenverfahren höherer Ordnung zur Lösung von hyperbolischen Systemen halblinearer Differentialgleichungen. (German) [the Adams method for characteristic procedures of higher order for the solution of hyperbolic systems of half-linear differential equations]. Numerische Mathematik, 5(1):443–460, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=5&issue=1&spage=
443.

Ansorge:1964:SGK

[Ans64] Rainer Ansorge. Zur Struktur gewisser Konvergenzkriterien bei der numerischen Lösung von Anfangswertaufgaben. (German) [On the structure of some convergence criteria for numerical solving of certain initial value problems]. Numerische Mathematik, 6(1):224– 234, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ansorge:1966:FVA

[Ans66] Rainer Ansorge. Zur Frage der Verallgemeinerung des Äquivalenzsatzes von Lax. (German) [On the question of generalization of the Lax equivalence principles]. Numerische Mathematik, 8(2):178–185, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ansorge:1967:KML

[Ans67]  Rainer Ansorge.  Konver-

genz von Mehrschrittverfahren zur Lösung halblinearer Anfangswertaufgaben. (German) [Convergence of a multistep method for solution of halflinear initial-value problems]. Numerische Mathematik, 10(3): 209–219, October 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ansorge:1969:KDQ

[Ans69] Rainer Ansorge. Konvergenz von Differenzenverfahren für quasilineare Anfangswertaufgaben. (German) [Convergence of difference methods for quasilinear initial value problems]. Numerische Mathematik, 13(3): 217–225, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ansorge:1973:FKF

[Ans73] Rainer Ansorge. f-Konvergenz und f-Stetigkeit von Operatorenfolgen auf metrischen Räumen. (German) [f-convergence and f-stability of operator sequences in metric spaces]. Numerische Mathematik, 20(4): 280–287, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Antes:1972:SLI

[Ant72] Heinz Antes. Splinefunktionen bei der Lösung von Integralgleichungen. (German) [On the solution of integral equations with spline functions]. Numerische Mathematik, 19(2): 116–126, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anderson:1976:DLA

[AO76] D. H. Anderson and M. R. Osborne. Discrete, linear approximation problems in polyhedral norms. Numerische Mathematik, 26(2):179–189, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anderson:1977:DNAa

[AO77a] D. H. Anderson and M. R. Osborne. Discrete, nonlinear approximation problems in polyhedral norms. Numerische Mathematik, 28(2):143–156, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anderson:1977:DNAb

[AO77b] D. H. Anderson and M. R. Osborne. Discrete, nonlinear approximation problems in polyhedral norms: a Levenberg-like algorithm. Numerische Mathematik, 28(2):157–170, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ainsworth:1993:UAP

[AO93] Mark Ainsworth and J. Tinsley Oden. A unified approach to a posteriori error estimation using element residual methods. Numerische Mathematik, 65(1): 23–50, May 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Andrew:1985:CNE

[AP85] Alan L. Andrew and John W.

Paine. Correction of Numerov’s eigenvalue estimates. Numerische Mathematik, 47(2): 289–300, September 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Andrew:1986:CFE

[AP86] Alan L. Andrew and John W. Paine. Correction of finite element estimates for Sturm– Liouville eigenvalues. Numerische Mathematik, 50(2): 205–215, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anselone:1968:SCV

[AR68] P. M. Anselone and L. B. Rall. The solution of characteristic value-vector problems by Newton’s method. Numerische Mathematik, 11(1):38–45, January 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arioli:1985:RBC

[AR85] Mario Arioli and Francesco Romani. Relations between condition numbers and the convergence of the Jacobi method for real positive definite matrices. Numerische Mathematik, 46(1): 31–42, March 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arunakirinathar:1993:MFE

[AR93] K. Arunakirinathar and B. D. Reddy. Mixed finite element methods for elastic rods of arbitrary geometry. Numerische Mathematik, 64(1):13–43, January 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ariyawansa:1998:LST

[Ari98]  K. A. Ariyawansa.  Line

search termination criteria for collinear scaling algorithms for minimizing a class of convex functions. Numerische Mathematik, 80(3):363–376, September 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080003/80800363.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800363.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800363.
ps.gz.

Arndt:1979:LGD

[Arn79] Herbert Arndt. Lösung von gewöhnlichen Differentialgleichungen mit nichtlinearen Splines. (German) [Solution of ordinary differential equations with nonlinear splines]. Numerische Mathematik, 33(3): 323–338, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arndt:1980:VKA

[Arn80]  Herbert Arndt.  Ein ve-

rallgemeinerter Kettenbruch-Algorithmus zur rationalen Hermite-Interpolation. (German) [A generalized continued fraction algorithm for rational Hermite interpolation]. Numerische Mathematik, 36(1):99– 107, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arnold:1981:DFE

[Arn81] Douglas N. Arnold. Discretization by finite elements of a model parameter dependent problem. Numerische Mathematik, 37(3):405–421, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arnautu:1982:AOD

[Arn82]  Viorel Arnăutu.  Approxi-

mation of optimal distributed control problems governed by variational inequalities. Numerische Mathematik, 38(3): 393–416, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arndt:1984:NSR

[Arn84] Herbert Arndt. Numerical solution of retarded initial value problems: local and global error and stepsize control. Numerische Mathematik, 43(3): 343–360, April 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aronson:1963:CPS

[Aro63a] D. G. Aronson. A correction to the paper “On the stability of certain finite difference approximations to parabolic systems of differential equations” [Num. Math. 5, 118–137 (1963)]. Numerische Mathematik, 5(1): 290, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Aro63b].

Aronson:1963:SCF

[Aro63b] D. G. Aronson. On the stability of certain finite difference approximations to parabolic systems of differential equations. Numerische Mathematik, 5(1): 118–137, December 1963. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See correction [Aro63a].

Abaffy:1983:ICG

[AS83] József Abaffy and Fridrich Sloboda. Imperfect conjugate gradient algorithms for extended quadratic functions. Numerische Mathematik, 42(1):97– 105, September 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Anselone:1987:NSI

[AS87] P. M. Anselone and I. H.

Sloan. Numerical solutions of integral equations on the half line. I. the compact case. Numerische Mathematik, 51(6): 599–614, November 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alfeld:1990:DBS

[AS90] Peter Alfeld and Larry L. Schumaker. On the dimension of bivariate spline spaces of smoothness r and degree d = 3r+1. Numerische Mathematik, 57(6/7): 651–661, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Abgrall:1997:UME

[AS97a] Rémi Abgrall and Thomas Sonar. On the use of Muhlbach expansions in the recovery step of ENO methods. Numerische Mathematik, 76(1): 1–25, March 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076001/70760001.
htm;   http://science.springer.
de/nmee/bibs/7076001/70760001.
htm.

Andra:1997:ISG

[AS97b] H. Andrä and E. Schnack. Integration of singular Galerkintype boundary element integrals for 3D elasticity problems. Numerische Mathematik, 76(2): 143–165, April 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076002/70760143.
htm;   http://science.springer.
de/nmee/bibs/7076002/70760143.
htm.

Ascher:1985:SDS

[Asc85] Uri Ascher. On some difference schemes for singular singularlyperturbed boundary value problems. Numerische Mathematik, 46(1):1–30, March 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Askey:1972:PCN

[Ask72] Richard Askey. Positivity of the Cotes numbers for some Jacobi abscissas. Numerische Mathematik, 19(1):46–48, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aziz:1988:NMR

[ASS88] A. K. Aziz, A. B. Stephens, and Manil Suri. Numerical methods for reaction-diffusion problems with non-differentiable kinetics. Numerische Mathematik, 53(1/2):1–11, July 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aziz:1980:FEM

[ASW80] A. K. Aziz, M. Schneider, and A. Werschulz. A finite element method for the Tricomi problem. Numerische Mathematik, 35(1):13–20, March 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arnold:1993:HER

[ASW93] M. Arnold, K. Strehmel, and R. Weiner. Half-explicit Runge– Kutta methods for semi-explicit differential-algebraic equations of index 1. Numerische Mathematik, 64(4):409–431, April 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Amara:1979:EFE

[AT79] M. Amara and J. M. Thomas. Equilibrium finite elements for the linear elastic problem. Numerische Mathematik, 33(4): 367–383, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Atkinson:1967:SNU

[Atk67] Kendall E. Atkinson. The solution of non-unique linear integral equations. Numerische Mathematik, 10(2): 117–124, July 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Atkinson:1972:NSF

[Atk72] Kendall Atkinson. The numerical solution of Fredholm integral equations of the second kind with singular kernels. Numerische Mathematik, 19(3): 248–259, June 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Atkinson:1973:IVN

[Atk73]  Kendall Atkinson.  Itera-

tive variants of the Nyström method for the numerical solution of integral equations. Numerische Mathematik, 22(1):17– 31, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Atteia:1968:FSD

[Att68] Marc Atteia. Fonctions ‘spline’ définies sur un ensemble convexe. (French) [Spline functions defined on a convex set]. Numerische Mathematik, 12(3): 192–210, October 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Atteia:1977:EEM

[Att77]  M. Atteia.  Évaluation de

l’erreur dans la méthode des éléments finis. (French) [Error in the finite elements method]. Numerische Mathematik, 28(3): 295–306, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Albrecht:1961:DNL

[AU61] Rudolf Albrecht and Wolfram Ulrich. Ein Differenzenverfahren zur näherungsweisen Lösung des Anfangswertproblems für Systeme halblinearer partieller Differentialgleichungen. 1. Ordnung. (German) [A difference equation for approximate solution of initial value problems for half-linear partial differential equations. 1. Order]. Numerische Mathematik, 3:131– 146, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Albrecht:1962:NTI

[AU62] R. F. Albrecht and W. Urich.

Numerical treatment of the initial value problem for systems of quasilinear partial differential equations of first order. Numerische Mathematik, 4:253– 261, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Auchmuty:1992:PEE

[Auc92] Giles Auchmuty. A posteriori error estimates for linear equations. Numerische Mathematik, 61(1):1–6, February 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Aumann:1963:UVL

[Aum63] Georg Aumann. Über den Vergleichsfaktor bei linearen Approximationsproblemen. (German) [On the comparison factor for linear approximation problems]. Numerische Mathematik, 5(1):68–72, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Auslender:1971:MNP

[Aus71]  A. Auslender.  Méthodes

numériques pour la décomposition et la minimisation de fonctions non différentiables. (French) [Numerical methods for the decomposition and minimisation of non-differentiable functions]. Numerische Mathematik, 18(3): 213–223, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Auzinger:1987:DCN

[Auz87]  Winfried Auzinger.  Defect

correction for nonlinear elliptic difference equations. Numerische Mathematik, 51(2): 199–208, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Auzinger:1988:QDR

[Auz88] Winfried Auzinger. A quantitative discrete H^{2}-regularity estimate for the Shortley–Weller scheme in convex domains. Numerische Mathematik, 52(5): 523–537, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1976:KSR

[AV76] G. Alefeld and R. S. Varga. Zur Konvergenz des symmetrischen Relaxationsverfahren. (German) [On the convergence of the symmetric SOR method]. Numerische Mathematik, 25(3): 291–295, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1985:RSM

[AV85] Götz Alefeld and Peter Volkmann. Regular splittings and monotone iteration functions. Numerische Mathematik, 46(2): 213–228, May 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Alefeld:1986:KSV

[AV86] Götz Alefeld and Peter Volkmann. Zur Konvergenz des SSOR-Verfahrens für nichtlineare Gleichungssysteme. (German) [Convergence of the SSOR-method for nonlinear systems of simultaneous equations]. Numerische Mathematik, 50(1): 111–121, January 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Axelsson:1989:AMP

[AV89] O. Axelsson and P. S. Vassilevski. Algebraic multilevel preconditioning methods. I. Numerische Mathematik, 56(2/3): 157–177, October 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arnold:1985:CSC

[AW85] Douglas N. Arnold and Wolfgang L. Wendland. The convergence of spline collocation for strongly elliptic equations on curves. Numerische Mathematik, 47(3):317–341, November 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Axelsson:1977:EEG

[Axe77] Owe Axelsson. Error estimates for Galerkin methods for quasilinear parabolic and elliptic differential equations in divergence form. Numerische Mathematik, 28(1):1–14, March 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Axelsson:1987:GCG

[Axe87] O. Axelsson. A generalized conjugate gradient, least square method. Numerische Mathematik, 51(2):209–227, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Allison:1983:NAN

[AY83] A. C. Allison and N. J. Young.

Numerical algorithms for the Nevanlinna–Pick problem. Numerische Mathematik, 42(2): 125–145, November 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Arcangeli:1993:ASC

[AY93] Rémi Arcangéli and Bernard Ycart. Almost sure convergence of smoothing D^{m}-splines for noisy data. Numerische Mathematik, 66(3):281–294, December 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Amara:1998:EIN

[AYB98] M. Amara, M. Ben Younes, and C. Bernardi. Error indicators for the Navier–Stokes equations in stream function and vorticity formulation. Numerische Mathematik, 80(2): 181–206, August 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080002/80800181.

                                        htm;   http://link.springer.
                                        de/link/service/journals/
                                        00211/papers/8080002/80800181.
                                        pdf;   http://link.springer.
                                        de/link/service/journals/
                                        00211/papers/8080002/80800181.
                                        ps.gz.

Babuska:1971:EBF

[Bab71] Ivo Babuška. Error-bounds for finite element method. Numerische Mathematik, 16(4): 322–333, January 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:1973:FEM

[Bab73] Ivo Babuška. The finite element method with Lagrangian multipliers. Numerische Mathematik, 20(3):179–192, June 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bacopoulos:1970:ASC

[Bac70] Alex Bacopoulos. On approximation by sums of Chebyshev norms. Numerische Mathematik, 16(3):243–247, December 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Badea:1989:GSA

[Bad89] Lori Badea. A generalization of the Schwarz alternating method to an arbitrary number of subdomains. Numerische Mathematik, 55(1):61–81, April 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baker:1970:EPI

[Bak70] Christopher T. H. Baker. The error in polynomial interpolation. Numerische Mathematik, 15(4):315–319, August 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bakker:1982:NGM

[Bak82] Miente Bakker. A note on C^{∘} Galerkin methods for twopoint boundary problems. Numerische Mathematik, 38(3): 447–453, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bank:1978:EAS

[Ban78] Randolph E. Bank. Efficient algorithms for solving tensor product finite element equations. Numerische Mathematik, 31(1):49–61, March 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Banerjee:1987:LNE

[Ban87] U. Banerjee. Lower norm error estimates for approximate solutions of differential equations with non-smooth coefficients. Numerische Mathematik, 51(3):303–321, July 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Banerjee:1992:NEN

[Ban92] Uday Banerjee. A note on the effect of numerical quadrature in finite element eigenvalue approximation. Numerische Mathematik, 61(2):145–152, March 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bao:1996:NAD

[Bao96] Gang Bao. Numerical analysis of diffraction by periodic structures: TM polarization. Numerische Mathematik, 75(1):1– 16, November 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6075001/60750001.
htm;   http://science.springer.
de/nmee/bibs/6075001/60750001.
htm.

Barnhill:1965:CQR

[Bar65] Robert E. Barnhill. Complex quadratures with remainders of minimum norm. Numerische Mathematik, 7(5):384–390, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barnhill:1968:APM

[Bar68] Robert E. Barnhill. Asymptotic properties of minimum norm and optimal quadratures. Numerische Mathematik, 12(5): 384–393, December 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bareiss:1969:NSL

[Bar69] Erwin H. Bareiss. Numerical solution of linear equations with Toeplitz and vector Toeplitz matrices. Numerische Mathematik, 13(5):404–424, October 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bartels:1971:SSM

[Bar71] Richard H. Bartels. A stabilization of the simplex method. Numerische Mathematik, 16(5): 414–434, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barker:1972:MN

[Bar72] George Phillip Barker. Monotone norms. Numerische Mathematik, 18(4):321–326, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bartke:1984:SNR

[Bar84] Klaus Bartke. Die Struktur der normalen rationalen Funktionen. (German) [The structure of normal rational functions]. Numerische Mathematik, 43(3):379–388, April 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barlow:1991:EAP

[Bar91a] Jesse L. Barlow. Error analysis of a pairwise summation algorithm to compute the sample variance. Numerische Mathematik, 58(6):583–590, January 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrlund:1991:CLS

[Bar91b] Anders Barrlund. Constrained least squares methods for linear time varying DAE systems. Numerische Mathematik, 60(2): 145–161, November 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barlow:1993:PRN

[Bar93] Jesse L. Barlow. Perturbation results for nearly uncoupled Markov chains with applications to iterative methods. Numerische Mathematik, 65(1): 51–62, May 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Batterson:1990:CSA

[Bat90] Steve Batterson. Convergence of the shifted QR algorithm on 3 × 3 normal matrices. Numerische Mathematik, 58(4): 341–352, November 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1961:FGK

[Bau61] F. L. Bauer. A further generalization of the Kantorovic inequality. Numerische Mathematik, 3:117–119, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1962:FVS

[Bau62] F. L. Bauer. On the field of values subordinate to a norm. Numerische Mathematik, 4:103– 113, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1963:OSM

[Bau63] F. L. Bauer. Optimally scaled matrices. Numerische Mathematik, 5(1):73–87, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1965:EPH

[Bau65a] F. L. Bauer. An elementary proof of the Hopf inequality for positive operators. Numerische Mathematik, 7(4):331–337, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1965:EWR

[Bau65b] F. L. Bauer. Elimination with weighted row combinations for solving linear equations and least squares problems. Numerische Mathematik, 7(4):338– 352, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 119–133].

Bauer:1966:ISR

[Bau66] F. L. Bauer. Inclusion sets for roots of polynomials. Numerische Mathematik, 9(2):173– 176, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1968:FVG

[Bau68] F. L. Bauer. Fields of values and Gershgorin disks. Numerische Mathematik, 12(2):91–95, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1969:ROS

[Bau69] F. L. Bauer. Remarks on optimally scaled matrices. Numerische Mathematik, 13(1):1– 3, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baumeister:1976:ENI

[Bau76] J. Baumeister. Über die Extremaleigenschaft nichtlinearer interpolierender Splines. (German) [On the extremal property of nonlinear interpolating splines]. Numerische Mathematik, 25(4):433–445, December 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brenner:1962:PSI

[BB62] C. Brenner and J. L. Brenner. The popularity of small integers as primitive roots. Numerische Mathematik, 4:336–342, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bieterman:1982:FEMa

[BB82a] M. Bieterman and I. Babuška.

The finite element method for parabolic equations. I. A posteriori error estimation. Numerische Mathematik, 40(3): 339–371, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bieterman:1982:FEMb

[BB82b] M. Bieterman and I. Babuška.

The finite element method for parabolic equations. II. A posteriori error estimation and adaptive approach. Numerische Mathematik, 40(3): 373–406, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:1990:MMP

[BB90] D. Braess and C. Blömer. A multigrid method for a parameter dependent problem in solid mechanics. Numerische Mathematik, 57(8):747–761, August 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bank:1992:HBM

[BB92] Randolph E. Bank and Mohamed Benbourenane. The hierarchical basis multigrid method for convection-diffusion equations. Numerische Mathematik, 61(1):7–37, February 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1995:EBF

[BB95] John W. Barrett and James F. Blowey. An error bound for the finite element approximation of the Cahn–Hilliard equation with logarithmic free energy. Numerische Mathematik, 72(1):1– 20, November 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5072001/50720001.
htm;   http://science.springer.
de/nmee/bibs/5072001/50720001.
htm.

Barrett:1997:FEA

[BB97] John W. Barrett and James F. Blowey. Finite element approximation of a model for phase separation of a multi-component alloy with non-smooth free energy. Numerische Mathematik, 77(1):1–34, July 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077001/70770001.
htm;   http://science.springer.
de/nmee/bibs/7077001/70770001.
htm.

Bojanczyk:1986:FTM

[BBdH86] A. W. Bojánczyk, R. P. Brent, and F. R. de Hoog. QR factorization of Toeplitz matrices. Numerische Mathematik, 49(1): 81–94, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bank:1990:SUT

[BBFS90] Randolph E. Bank, Josef F.

Bürgler, Wolfgang Fichtner, and R. Kent Smith. Some upwinding techniques for finite element approximations of convection-diffusion equations. Numerische Mathematik, 58(2): 185–202, September 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Backus:1960:RAL

[BBG^{+}60] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language Algol 60. Numerische Mathematik, 2(1):106–136, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/
10.1007/BF01386216.

Backus:1963:RRA

[BBG^{+}63] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. Mc-Carthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language AL-GOL 60. Numerische Mathematik, 4:420–453, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Edited by Peter Naur. Dedicated to the memory of William Turanski.

Barrett:1998:FEA

[BBG98] John W. Barrett, James F. Blowey, and Harald Garcke. Finite element approximation of a fourth order nonlinear degenerate parabolic equation. Numerische Mathematik, 80(4): 525–556, October 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/8080004/
                                        80800525.htm;   http://link.
                                        springer.de/link/service/
                                        journals/00211/papers/8080004/
                                        80800525.pdf.

Blum:1978:CGM

[BC78] E. K. Blum and A. R. Curtis. A convergent gradient method for matrix eigenvector-eigentuple problems. Numerische Mathematik, 31(3):247–263, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1991:FEAb

[BC91] John W. Barrett and Roma Chakrabarti. Finite element approximation of the volume-matching problem. Numerische Mathematik, 60(3): 291–313, December 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Butcher:1992:SGM

[BC92] J. C. Butcher and R. P. K. Chan.

On symmetrizers for Gauss methods. Numerische Mathematik, 60(4):465–476, January 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bank:1993:ACS

[BC93] Randolph E. Bank and Tony F.

Chan. An analysis of the composite step biconjugate gradient method. Numerische Mathematik, 66(3):295–319, December 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezzi:1986:HGA

[BCD86] F. Brezzi, M. Cornalba, and A. Di Carlo. How to get around a simple quadratic fold. Numerische Mathematik, 48(4): 417–427, April 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brink:1996:SCB

[BCS96] Ulrich Brink, Carsten Carstensen, and Erwin Stein. Symmetric coupling of boundary elements and Raviart–Thomastype mixed finite elements in elastostatics. Numerische Mathematik, 75(2):153–174, December 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6075002/60750153.

                                        htm;   http://science.springer.
                                        de/nmee/bibs/6075002/60750153.
                                        htm.

Brown:1968:NLI

[BD68] Kenneth M. Brown and J. E. Dennis, Jr. On Newton-like iteration functions: General convergence theorems and a specific algorithm. Numerische Mathematik, 12(3):186–191, October 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brown:1972:DFA

[BD72] Kenneth M. Brown and J. E. Dennis, Jr. Derivative free analogues of the Levenberg– Marquardt and Gauss algorithms for nonlinear least squares approximation. Numerische Mathematik, 18(4): 289–298, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beavers:1973:CME

[BD73] A. N. Beavers, Jr. and E. D. Denman. A computational method for eigenvalues and eigenvectors of a matrix with real eigenvalues. Numerische Mathematik, 21(5):389–396, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bain:1977:CRE

[BD77] M. Bain and L. M. Delves. The convergence rates of expansions in Jacobi polynomials. Numerische Mathematik, 27(2):

219–225, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:1981:EEC

[BD81] Ivo Babuška and Milo R. Dorr.

Error estimates for the combined h and p versions of the finite element method. Numerische Mathematik, 37(2): 257–277, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bernadou:1982:AGA

[BD82] M. Bernadou and Y. Ducatel. Approximation of general arch problems by straight beam elements (elasticity). Numerische Mathematik, 40(1):1– 29, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bader:1983:SIM

[BD83] G. Bader and P. Deuflhard.

A semi-implicit mid-point rule for stiff systems of ordinary differential equations. Numerische Mathematik, 41(3): 373–398, July 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezzi:1988:SMM

[BD88] Franco Brezzi and Jim Douglas, Jr. Stabilized mixed methods for the Stokes problem. Numerische Mathematik, 53(1/2): 225–235, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bornemann:1996:CMM

[BD96] Folkmar A. Bornemann and Peter Deuflhard. The cascadic multigrid method for elliptic problems. Numerische Mathematik, 75(2):135–152, December 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6075002/60750135.
htm;   http://science.springer.
de/nmee/bibs/6075002/60750135.
htm.

Bialecki:1997:MAM

[BD97] Bernard Bialecki and Maksymilian Dryja. Multilevel additive and multiplicative methods for orthogonal spline collocation problems. Numerische Mathematik, 77(1): 35–58, July 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077001/70770035.
htm;   http://science.springer.
de/nmee/bibs/7077001/70770035.
htm.

Bansch:1998:AFE

[BD98] E. Bänsch and W. Dörfler. Adaptive finite elements for exterior domain problems. Numerische Mathematik, 80(4): 497–523, October 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/

journals/00211/bibs/8080004/
80800497.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8080004/
80800497.pdf.

Braess:1999:CMA

[BD99] Dietrich Braess and Wolfgang Dahmen. A cascadic multigrid algorithm for the Stokes equations. Numerische Mathematik, 82(2):179–191, April 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082002/90820179.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082002/90820179.pdf.

Brezzi:1987:MFE

[BDDF87] Franco Brezzi, Jim Douglas, Jr., Ricardo Durán, and Michel Fortin. Mixed finite elements for second order elliptic problems in three variables. Numerische Mathematik, 51(2): 237–250, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bai:1997:IFP

[BDG97] Zhaojun Bai, James Demmel, and Ming Gu. An inverse free parallel spectral divide and conquer algorithm for nonsymmetric eigenproblems. Numerische Mathematik, 76(3): 279–308, May 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/7076003/70760279.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/7076003/70760279.
                                        htm.

Brezzi:1985:TFM

[BDM85] Franco Brezzi, Jim Douglas, Jr., and L. D. Marini. Two families of mixed finite elements for second order elliptic problems. Numerische Mathematik, 47(2): 217–235, September 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brenner:1972:PMS

[BdP72] J. L. Brenner and J. de Pillis. Partitioned matrices and Seidel convergence. Numerische Mathematik, 19(1):76–80, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baker:1980:ATS

[BDS80] Garth A. Baker, Vassilios A. Dougalis, and Steven M. Serbin. An approximation theorem for second-order evolution equations. Numerische Mathematik, 35(2):127–142, June 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bernadou:1988:ACC

[BDT88] M. Bernadou, Y. Ducatel, and P. Trouvé. Approximation of a circular cylindrical shell by Clough-Johnson flat plate finite elements. Numerische Mathematik, 52(2):187–217, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bank:1988:HBM

[BDY88] Randolph E. Bank, Todd F.

Dupont, and Harry Yserentant. The hierarchical basis multigrid method. Numerische Mathematik, 52(4):427–458, April 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bjorck:1973:ACV

[BE73] Åke Björck and Tommy Elfving. Algorithms for confluent Vandermonde systems. Numerische Mathematik, 21(2): 130–137, April 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1986:FEA

[BE86] John W. Barrett and Charles M.

Elliott. Finite element approximation of the Dirichlet problem using the boundary penalty method. Numerische Mathematik, 49(4):343–366, July 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1987:PFE

[BE87] John W. Barrett and Charles M.

Elliott. A practical finite element approximation of a semidefinite Neumann problem on a curved domain. Numerische Mathematik, 51(1):23–36, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beauwens:1979:FIM

[Bea79]  Robert Beauwens.  Factor-

ization iterative methods, Moperators and H-operators. Numerische Mathematik, 31(4): 335–357, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Bea86a].

Beauwens:1986:EIM

[Bea86a] R. Beauwens. Erratum: “Factorization iterative methods, M-operators and H-operators” [Numer. Math. 31 1978/1979, no. 4, 335–357 and MR 81e:65014]. Numerische Mathematik, 49(4):457, July 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Bea79].

Beauzamy:1986:JIP

[Bea86b] Bernard Beauzamy. Jensen’s inequality for polynomials with concentration at low degrees. Numerische Mathematik, 49 (2/3):221–225, March 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Busch:1975:EAM

[BEHH75] W. Busch, R. Esser, W. Hackbusch, and U. Herrmann. Extrapolation applied to the method of characteristics for a first order system of two partial differential equations. I. the initial value problem. Numerische Mathematik, 24(4):331–353, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bellen:1983:MMP

[Bel83] A. Bellen. Monotone methods for periodic solutions of second order scalar functionaldifferential equations. Numerische Mathematik, 42(1):15– 30, September 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Benders:1962:PPS

[Ben62] J. F. Benders. Partitioning procedures for solving mixedvariables programming problems. Numerische Mathematik, 4:238–252, December 1962. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bennett:1964:IEM

[Ben64] J. H. Bennett. Integral equation methods for transport problems. some numerical results. Numerische Mathematik, 6(1):49– 54, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bennett:1965:TFM

[Ben65] John M. Bennett. Triangular factors of modified matrices. Numerische Mathematik, 7 (3):217–221, June 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Benbourhim:1986:SAH

[Ben86] Mohammed-Najib Benbourhim. Spline approximant in Hilbertian subspaces of ⌋(Ω). Numerische Mathematik, 49(2/3): 291–303, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buttgenbach:1993:CEB

[BEN93] B. Büttgenbach, H. Esser, and R. J. Nessel. On the comparison of error bounds for finite difference schemes. Numerische Mathematik, 64(4): 477–486, April 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

BenBelgacem:1999:MFE

[Ben99]  Faker Ben Belgacem.  The

Mortar finite element method with Lagrange multipliers. Numerische Mathematik, 84 (2):173–197, December 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9084002/90840173.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9084002/90840173.pdf.

Bertram:1959:FGS

[Ber59a] G. Bertram. Eine Fehlerabschätzung für gewisse selbstadjungierte, gewöhnliche Randwertaufgaben. (German) [An error estimate for certain adjoint ordinary boundary value problems]. Numerische Mathematik, 1(1):181–185, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
181.

Bertram:1959:VFR

[Ber59b] G. Bertram. Verschärfung einer Fehlerabschätzung zum Ritz– Galerkinschen Verfahren von Kryloff für Randwertaufgaben. (German) [Tightening of an error bound for the Ritz–Galerkin method for Krylov boundary value problems]. Numerische Mathematik, 1:135–141, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See correction [Ber60].

Bertram:1960:BMA

[Ber60] G. Bertram. Berichtigung zu meinen Arbeit ,,Verschärfung einer Fehlerabschätzung … [Num. Math. 1, 135–141 (1959)]. (German) [Correction to my paper “Tightening of an error bound … ”]. Numerische Mathematik, 2(1):54, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Ber59b].

Bereanu:1972:CCM

[Ber72] Bernard Bereanu. On the convergence of Cartesian multidimensional quadrature formulae.

Numerische Mathematik, 19(4): 348–350, August 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berger:1973:EEF

[Ber73] Alan E. Berger. L^{2} error estimates for finite elements with interpolated boundary conditions. Numerische Mathematik, 21(4): 345–349, August 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bernardi:1982:AHB

[Ber82] C. Bernardi. Approximation of Hopf bifurcation. Numerische Mathematik, 39(1):15–37, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bernstein:1984:ABM

[Ber84] Herbert J. Bernstein. An accelerated bisection method for the calculation of eigenvalues of a symmetric tridiagonal matrix. Numerische Mathematik, 43(1): 153–160, January 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berger:1989:CFE

[Ber89a]  Harald Berger.  A conver-

gent finite element formulation for transonic flow. Numerische Mathematik, 56(5): 425–447, November 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berrut:1989:BFC

[Ber89b] Jean-Paul Berrut. Barycentric formulae for cardinal (SINC-) interpolants. Numerische Mathematik, 54(6):703–718, March 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Ber89c].

Berrut:1989:EFC

[Ber89c] Jean-Paul Berrut. Erratum:

“Barycentric formulae for cardinal (SINC-)interpolants”. Numerische Mathematik, 55(6): 747, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Ber89b].

Bermejo:1991:AAG

[Ber91] Rodolfo Bermejo. Analysis of an algorithm for the Galerkincharacteristic method. Numerische Mathematik, 60(2): 163–194, November 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bertoluzza:1997:IEW

[Ber97] Silvia Bertoluzza. Interior estimates for the wavelet Galerkin method. Numerische Mathematik, 78(1):1–20, November 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.de/link/service/
journals/00211/bibs/7078001/

70780001.htm;   http://science.
springer.de/nmee/bibs/7078001/
70780001.htm.

Bettis:1970:NIP

[Bet70] D. G. Bettis. Numerical integration of products of Fourier and ordinary polynomials. Numerische Mathematik, 14(5): 421–434, April 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=14&issue=5&spage=
421.

Beyn:1978:SDS

[Bey78]  Wulf-Jürgen Beyn.  Zur

Stabilität von Differenzenverfahren für Systeme linearer gewöhnlicher Randwertaufgaben. (German) [On the stability of difference approximations to systems of linear ordinary boundary value problems]. Numerische Mathematik, 29(2):209–226, January 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beyn:1987:ICC

[Bey87] Wolf-Jürgen Beyn. On invariant closed curves for one-step methods. Numerische Mathematik, 51(1):103–122, May 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1960:NET

[BF60] F. L. Bauer and C. T. Fike.

Norms and exclusion theorems. Numerische Mathematik, 2(1): 137–141, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bradbury:1966:NIM

[BF66a] W. W. Bradbury and R. Fletcher.

New iterative methods for solution of the eigenproblem. Numerische Mathematik, 9(3):259– 267, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Broyden:1966:ASC

[BF66b] C. G. Broyden and F. Ford.

An algorithm for the solution of certain kinds of linear equations. Numerische Mathematik, 8(4):307–323, June 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bosarge:1970:IDM

[BF70] W. E. Bosarge, Jr. and Peter L.

Falb. Infinite dimensional multipoint methods and the solution of two point boundary value problems. Numerische Mathematik, 14(3):264–286, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berens:1990:CSM

[BF90] H. Berens and M. Finzel. A continuous selection of the metric projection in matrix spaces, addendum. Numerische Mathematik, 57(6/7):663–667, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Brass:1999:EBR

[BF99] H. Brass and J.-W. Fischer. Error bounds for Romberg quadrature. Numerische Mathematik, 82(3):389–408, May 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082003/90820389.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082003/90820389.pdf.

Businger:1965:LLS

[BG65] Peter A. Businger and Gene H. Golub. Linear least squares solutions by Householder transformations. Numerische Mathematik, 7(3):269–276, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 111– 118].

Brown:1971:DTC

[BG71] Kenneth M. Brown and William B. Gearhart. Deflation techniques for the calculation of further solutions of a nonlinear system. Numerische Mathematik, 16(4): 334–342, January 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barnhill:1976:IRT

[BG76a] Robert E. Barnhill and John A.

Gregory. Interpolation remainder theory from Taylor expansions on triangles. Numerische Mathematik, 25(4): 401–408, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barnhill:1976:SKT

[BG76b] Robert E. Barnhill and John A.

Gregory. Sard kernel theorems on triangular domains with application to finite element error bounds. Numerische Mathematik, 25(3):215–229, September 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Blum:1978:NSE

[BG78] E. K. Blum and P. B. Geltner. Numerical solution of eigentuple-eigenvector problems in Hilbert spaces by a gradient method. Numerische Mathematik, 31(3):231–246, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bini:1994:ISD

[BG94] Dario Bini and Luca Gemignani.

Iteration schemes for the divideand-conquer eigenvalue solver. Numerische Mathematik, 67(4): 403–425, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067004/40670403.
htm;   http://science.springer.
de/nmee/bibs/4067004/40670403.
htm.

Beattie:1995:MCL

[BG95] Christopher Beattie and Friedrich Goerisch. Methods for computing lower bounds to eigenvalues of self-adjoint operators. Numerische Mathematik, 72(2): 143–172, December 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5072002/
50720143.htm;   http://science.
springer.de/nmee/bibs/5072002/
50720143.htm.

Bunse-Gerstner:1989:QA

[BGBM89] Angelika Bunse-Gerstner, Ralph Byers, and Volker Mehrmann. A quaternion QR algorithm. Numerische Mathematik, 55(1): 83–95, April 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bunse-Gerstner:1991:NCA

[BGBMN91] Angelika Bunse-Gerstner, Ralph Byers, Volker Mehrmann, and Nancy K. Nichols. Numerical computation of an analytic singular value decomposition of a matrix valued function. Numerische Mathematik, 60(1):1– 39, November 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1979:ES

[BGHS79] F. L. Bauer, G. H. Golub, A. S.

Householder, and K. Samelson. Eduard L. Stiefel: 4/21/1909– 11/27/1978. Numerische Mathematik, 32(4):480–481, December 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). With a German translation.

Bauer:1960:MCR

[BH60a] F. L. Bauer and A. S. Householder. Moments and characteristic roots. Numerische Mathematik, 2(1):42–53, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1960:SII

[BH60b] F. L. Bauer and A. S. Householder. Some inequalities involving the Euclidean condition of a matrix. Numerische Mathematik, 2(1):308–311, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1961:ANC

[BH61a] F. L. Bauer and A. S. Householder. Absolute norms and characteristic roots. Numerische Mathematik, 3:241–246, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bergman:1961:AMK

[BH61b] Stefan Bergman and John G. Herriot. Application of the method of the kernel function for solving boundary- value problems. Numerische Mathematik, 3:209–225, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bramble:1962:FFD

[BH62] J. H. Bramble and B. E. Hubbard. On the formulation of finite difference analogues of the Dirichlet problem for Poisson’s equation. Numerische Mathematik, 4:313–327, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bergmann:1965:NSB

[BH65] Stefan Bergmann and John G. Herriot. Numerical solution of boundary-value problems by the method of integral operators. Numerische Mathematik, 7(1): 42–65, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bergman:1971:CCS

[BH71a] S. Bergman and J. G. Herriot. Computation of a class of subsonic flow patterns. Numerische Mathematik, 17(3): 247–252, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bramble:1971:BCL

[BH71b] J. H. Bramble and S. R. Hilbert.

Bounds for a class of linear functionals with applications to Hermite interpolation. Numerische Mathematik, 16(4): 362–369, January 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:1973:SIZ

[BH73] D. Braess and K. P. Hadeler.

Simultaneous inclusion of the zeros of a polynomial. Numerische Mathematik, 21(2): 161–165, April 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brosowski:1974:VAG

[BH74] Bruno Brosowski and K.-H.

Hoffmann. Eine Variationsungleichung und Anwendungen. (German) [A variational inequality and its application]. Numerische Mathematik, 22(2): 137–147, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1987:MJH

[BH87] F. L. Bauer and A. S. Householder. In memoriam: J. H. Wilkinson (1919–1986). Numerische Mathematik, 51(1): 1–2, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bojanov:1990:COQ

[BH90] Borislav Bojanov and Daren Huang. Comparison of optimal quadrature formulas. Numerische Mathematik, 56(8): 817–825, February 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barbosa:1992:BLM

[BH92a] Helio J. C. Barbosa and Thomas J. R. Hughes. Boundary Lagrange multipliers in finite element methods: error analysis in natural norms. Numerische Mathematik, 62(1):1–15, June 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bartels:1992:SSV

[BH92b] Sven G. Bartels and Desmond J. Higham. The structured sensitivity of Vandermonde-like systems. Numerische Mathematik, 62(1):17–33, June 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bshouty:1988:CMS

[BHH88] D. Bshouty, N. Hengartner, and W. Hengartner. A constructive method for starlike harmonic mappings. Numerische Mathematik, 54(2):167–178, November 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berry:1985:ACS

[BHK^{+}85] M. W. Berry, M. T. Heath, I. Kaneko, M. Lawo, R. J. Plemmons, and R. C. Ward. An algorithm to compute a sparse basis of the null space. Numerische Mathematik, 47(4): 483–504, December 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buchauer:1994:SAI

[BHK94] O. Buchauer, P. Hiltmann, and M. Kiehl. Sensitivity analysis of initial-value problems with application to shooting techniques. Numerische Mathematik, 67(2): 151–159, March 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067002/40670151.
htm;   http://science.springer.
de/nmee/bibs/4067002/40670151.
htm.

Brezzi:1977:EEF

[BHR77] Franco Brezzi, William W.

Hager, and P. A. Raviart. Error estimates for the finite element solution of variational inequalities. I. primal theory. Numerische Mathematik, 28(4): 431–443, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezzi:1978:EEF

[BHR78] Franco Brezzi, William W.

Hager, and P.-A. Raviart. Error estimates for the finite element solution of variational inequalities. II. mixed methods. Numerische Mathematik, 31(1): 1–16, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bialecki:1990:SQR

[Bia90]  Bernard Bialecki.  A Sinc

quadrature rule for Hadamard finite-part integrals. Numerische Mathematik, 57(3): 263–269, May 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bialecki:1991:ADI

[Bia91] Bernard Bialecki. An alternating direction implicit method for orthogonal spline collocation linear systems. Numerische Mathematik, 59(5):413–429, August 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bittner:1964:UMI

[Bit64] L. Bittner. Über ein mehrstufiges Iterationsverfahren zur Lösung von linearen Gleichungen. (German) [On a more layered iteration method for solution of linear equations]. Numerische Mathematik, 6(1):161– 180, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bittner:1968:AVH

[Bit68] L. Bittner. Abschätzungen bei Variationsmethoden mit Hilfe von Dualitätssätzen. I. (German) [Bounds of the variational method with the aid of the duality principle. I]. Numerische Mathematik, 11(2):

129–143, February 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bittner:1970:SRT

[Bit70] L. Bittner. Some representation theorems for functions and sets and their application to nonlinear programming. Numerische Mathematik, 16(1):32– 51, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bittner:1971:AVH

[Bit71] L. Bittner. Abschätzungen bei Variationsmethoden mit Hilfe von Dualitätssätzen. II. (German) [Bounds of the variational method with the aid of the duality principle. II]. Numerische Mathematik, 16(4): 285–303, January 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bittner:1991:NSP

[Bit91] Leonhard Bittner. A numerical sampling problem and the Weyl pseudorandom numbers. Numerische Mathematik, 59(7): 637–645, October 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bakke:1985:SRQ

[BJ85a] V. L. Bakke and Z. Jackiewicz.

Stability of reducible quadrature methods for Volterra integral equations of the second kind. Numerische Mathematik, 47(2):159–173, September 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baltus:1985:TEB

[BJ85b] Christopher Baltus and William B. Jones. Truncation error bounds for limit-periodic continued fractions K(a_{n}/1) with lim a_{n} = 0. Numerische Mathematik, 46(4): 541–569, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bakke:1986:SAP

[BJ86] V. L. Bakke and Z. Jackiewicz. Stability analysis of product θ-methods for Abel integral equations of the second kind. Numerische Mathematik, 48(2):127–136, February 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baltus:1989:TEB

[BJ89] Christopher Baltus and William B. Jones. Truncation error bounds for modified continued fractions with applications to special functions. Numerische Mathematik, 55(3):281–307, June 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bjorck:1988:DMS

[Bjö88] Åke Björck. A direct method for sparse least squares problems with lower and upper bounds. Numerische Mathematik, 54(1):

19–32, October 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bellen:1988:SAO

[BJZ88] A. Bellen, Z. Jackiewicz, and M. Zennaro. Stability analysis of one-step methods for neutral delay-differential equations. Numerische Mathematik, 52(6): 605–619, June 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Biegler-Konig:1981:NIP

[BK81] Friedrich W. Biegler-König. A Newton iteration process for inverse eigenvalue problems. Numerische Mathematik, 37(3): 349–354, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brewster:1984:NSR

[BK84] M. E. Brewster and R. Kannan. Nonlinear successive overrelaxation. Numerische Mathematik, 44(2):309–315, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ball:1987:NMD

[BK87] J. M. Ball and G. Knowles. A numerical method for detecting singular minimizers. Numerische Mathematik, 51(2): 181–197, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bernardi:1996:MSP

[BK96] Christine Bernardi and Andreas Karageorghis. Méthodes spectrales dans une portion de disque. (French) [Spectral methods in a part of the disk]. Numerische Mathematik, 73(3): 265–289, May 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073003/60730265.
htm;   http://science.springer.
de/nmee/bibs/6073003/60730265.
htm.

Beyn:1998:NTE

[BK98] Wolf-Jürgen Beyn and Winfried Kleß. Numerical Taylor expansions of invariant manifolds in large dynamical systems. Numerische Mathematik, 80(1):1–38, July 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080001/80800001.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800001.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800001.
ps.gz.

Bunch:1976:DSM

[BKP76] James R. Bunch, Linda Kaufman, and Beresford N. Parlett. Decomposition of a symmetric matrix. Numerische Mathematik, 27(1):95–109, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:1979:DIE

[BKP79] I. Babuška, R. B. Kellogg, and J. Pitkäranta. Direct and inverse error estimates for finite elements with mesh refinements. Numerische Mathematik, 33(4): 447–471, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brosowski:1967:ERT

[BL67] Bruno Brosowski and Henry L.

Loeb. Zur Eindeutigkeit der rationalen Tschebyscheff-Approximation an stetig differenzierbare Funktionen. (German) [On the uniqueness of rational Chebyshev approximation of continuous differentiable functions]. Numerische Mathematik, 10(1):51–55, May 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrar:1970:CMN

[BL70a] Richard B. Barrar and Henry L.

Loeb. On the convergence in measure of non-linear Chebyshev approximations. Numerische Mathematik, 14(3): 305–312, February 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrar:1970:RAN

[BL70b] Richard B. Barrar and Henry L. Loeb. On the Remez algorithm for non-linear families. Numerische Mathematik, 15(5): 382–391, September 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrar:1974:AEM

[BL74] R. B. Barrar and Henry L. Loeb. Analytic extended monosplines. Numerische Mathematik, 22(2): 119–125, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrar:1976:MZA

[BL76] R. B. Barrar and Henry L. Loeb. Multiple zeros and applications to optimal linear functionals. Numerische Mathematik, 25(3): 251–262, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Belogus:1978:DIA

[BL78] David Belogus and Nadav Liron. DCR2: an improved algorithm for ℓ_{∞} rational approximation on intervals. Numerische Mathematik, 31(1):17–29, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bini:1980:SFA

[BL80] Dario Bini and Grazia Lotti. Stability of fast algorithms for matrix multiplication. Numerische Mathematik, 36(1):63– 72, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bozzini:1981:AMF

[BL81] M. Bozzini and L. Lenarduzzi.

Approximation of multivariable functions with respect to random points less than 2^{k}, k dimension of space. Numerische Mathematik, 37(2): 193–203, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Banerjee:1986:ASF

[BL86a] U. Banerjee and A. Lutoborski.

Approximation of the solution of a fourth-order boundary value problem with nonsmooth coefficient. Numerische Mathematik, 50(2):125–145, March 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Blanchard:1986:NAE

[BL86b] D. Blanchard and P. Le Tallec. Numerical analysis of the equations of small strains quasistatic elastoviscoplasticity. Numerische Mathematik, 50(2): 147–169, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Birkhoff:1990:SCQ

[BL90] Garrett Birkhoff and Robert E.

Lynch. Some current questions on solving linear elliptic problems. Numerische Mathematik, 57(6/7):527–545, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Bahlmann:1992:FSF

[BL92] Dieter Bahlmann and Ulrich Langer. A fast solver for the first biharmonic boundary value problem. Numerische Mathematik, 63(3):297–313, December 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1993:FEE

[BL93] John W. Barrett and W. B. Liu. Finite element error analysis of a quasi-Newtonian flow obeying the Carreau or power law. Numerische Mathematik, 64(4):433–453, April 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1994:QNE

[BL94] John W. Barrett and W. B. Liu. Quasi-norm error bounds for the finite element approximation of a non-Newtonian flow. Numerische Mathematik, 68(4): 437–456, October 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4068004/

40680437.htm;   http://science.
springer.de/nmee/bibs/4068004/
40680437.htm.

Boffi:1997:ANA

[BL97] Daniele Boffi and Carlo Lovadina. Analysis of new augmented Lagrangian formulations for mixed finite element schemes. Numerische Mathematik, 75(4):405–419, February 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7075004/70750405.
htm;   http://science.springer.
de/nmee/bibs/7075004/70750405.
htm.

Blair:1972:EBS

[Bla72] J. J. Blair. Error bounds for the solution of nonlinear twopoint boundary value problems by Galerkin’s method. Numerische Mathematik, 19(2):99– 109, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Blatt:1977:RTA

[Bla77a] Hans-Peter Blatt. Rationale Tschebyscheff-Approximation über unbeschränkten Intervallen. (German) [Rational Chebyshev approximation on unbounded intervals]. Numerische Mathematik, 27(2): 179–190, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Blatt:1977:SRA

[Bla77b] Hans-Peter Blatt. Zur Stetigkeit rationaler T-Approximationen über unbeschränkten Intervallen. (German) [Continuity of best rational approximations on unbounded intervals]. Numerische Mathematik, 27(2): 191–202, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Borwein:1995:MER

[BLLN95] J. M. Borwein, A. S. Lewis, M. N. Limber, and D. Noll. Maximum entropy reconstruction using derivative information. II. computational results. Numerische Mathematik, 69(3): 243–256, January 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/5069003/
                                        50690243.htm;   http://science.
                                        springer.de/nmee/bibs/5069003/
                                        50690243.htm.

Blum:1986:AEE

[BLR86] H. Blum, Q. Lin, and R. Rannacher. Asymptotic error expansion and Richardson extrapolation for linear finite elements. Numerische Mathematik, 49(1): 11–37, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Blanchard:1989:NAE

[BLR89] Dominique Blanchard, Patrick Le Tallec, and Michel Ravachol. Numerical analysis of evolution problems in nonlinear small strains elastoviscoplasticity. Numerische Mathematik, 55 (2):177–195, May 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Blum:1961:SPE

[Blu61] M. Blum. A stagewise parameter estimation procedure for correlated data. Numerische Mathematik, 3:202–208, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrar:1974:EOI

[BLW74] R. B. Barrar, Henry L. Loeb, and Helmut Werner. On the existence of optimal integration formulas for analytic functions. Numerische Mathematik, 23(2): 105–117, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brunovsky:1984:SSP

[BM84] Pavol Brunovský and Pavol Meravý. Solving systems of polynomial equations by bounded and real homotopy. Numerische Mathematik, 43(3): 397–418, April 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Belbas:1987:AFP

[BM87] S. A. Belbas and I. D. Mayergoyz. Applications of fixedpoint methods to discrete variational and quasi-variational inequalities. Numerische Mathematik, 51(6):631–654, November 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buhmann:1992:SPN

[BM92] Martin D. Buhmann and Charles A. Micchelli. Spline prewavelets for non-uniform knots. Numerische Mathematik, 61(4): 455–474, May 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baratella:1996:AEB

[BM96] Paola Baratella and Paola Moroni. Asymptotic error bounds of product formulas for functions with interior singularity. Numerische Mathematik, 74(2): 137–151, August 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074002/60740137.
htm;   http://science.springer.
de/nmee/bibs/6074002/60740137.
htm.

Bernardi:1987:SAP

[BMM87] Christine Bernardi, Yvon Maday, and Brigitte Métivet. Spectral approximation of the periodic-nonperiodic Navier– Stokes equations. Numerische Mathematik, 51(6):655–700, November 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bogues:1981:IME

[BMP81] Katrina Bogues, Corbett R. Morrow, and Thomas N. L.

Patterson. An implementation of the method of Ermakov and Zolotukhin for multidimensional integration and interpolation. Numerische Mathematik, 37(1):49–60, May 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bai:1999:BAT

[BMPS99] Zhong-Zhi Bai, Violeta Migallón, José Penadés, and Daniel B. Szyld. Block and asynchronous two-stage methods for mildly nonlinear systems. Numerische Mathematik, 82(1):1–20, March 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082001/90820001.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082001/90820001.pdf.

Bowdler:1966:SRC

[BMPW66] H. J. Bowdler, R. S. Martin, G. Peters, and J. H. Wilkinson. Solution of real and complex systems of linear equations. Numerische Mathematik, 8(3): 217–234, May 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 93–110].

Bocharov:1996:NSL

[BMR96] G. A. Bocharov, G. I. Marchuk, and A. A. Romanyukha. Numerical solution by LMMs of stiff delay differential systems modelling an immune response. Numerische Mathematik, 73(2): 131–148, April 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073002/60730131.
htm;   http://science.springer.
de/nmee/bibs/6073002/60730131.
htm.

Bowdler:1968:ASM

[BMRW68] Hilary J. Bowdler, R. S. Martin, C. Reinsch, and J. H. Wilkinson. The QR and QL algorithms for symmetric matrices. Numerische Mathematik, 11(4): 293–306, May 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 227–240].

Benner:1998:NSS

[BMX98] Peter Benner, Volker Mehrmann, and Hongguo Xu. A numerically stable, structure preserving method for computing the eigenvalues of real Hamiltonian or symplectic pencils. Numerische Mathematik, 78(3):329–358, January 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078003/70780329.
htm;   http://science.springer.
de/nmee/bibs/7078003/70780329.
htm.

Bunch:1978:USV

[BN78] James R. Bunch and Christopher P. Nielsen. Updating the singular value decomposition. Numerische Mathematik, 31(2):111–129, June 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brunner:1981:SCM

[BN81] Hermann Brunner and Syvert P.

Nørsett. Superconvergence of collocation methods for Volterra and Abel integral equations of the second kind. Numerische Mathematik, 36(4): 347–358, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Boland:1984:SBC

[BN84] J. M. Boland and R. A.

Nicolaides. On the stability of bilinear-constant velocitypressure finite elements. Numerische Mathematik, 44(2): 219–222, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baranger:1990:ANE

[BN90] Jacques Baranger and Khalid Najib. Analyse numérique des écoulements quasi-newtoniens dont la viscosité obéit à la loi puissance ou la loi de carreau. (French) [Numerical analysis of quasi-Newtonian flow obeying the power law or the Carreau law]. Numerische Mathematik, 58(1):35–49, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bunch:1978:ROM

[BNS78] James R. Bunch, Christopher P. Nielsen, and Danny C. Sorensen. Rank one modification of the symmetric eigenproblem. Numerische Mathematik, 31(1):31– 48, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buttgenbach:1989:SET

[BNvW89] B. Büttgenbach, R. J. Nessel, and E. van Wickeren. Sharp estimates in terms of moduli of smoothness for compound cubature rules. Numerische Mathematik, 56(4):359–369, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brown:1972:IGR

[BNZ72a] H. Brown, J. Neubüser, and H. Zassenhaus. On integral groups. I: The reducible case. Numerische Mathematik, 19(5): 386–399, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brown:1972:IGI

[BNZ72b] H. Brown, J. Neubüser, and H. Zassenhaus. On integral groups. II: The irreducible case. Numerische Mathematik, 20(1): 22–31, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:1980:AFE

[BO80] I. Babuška and J. Osborn. Analysis of finite element methods for second order boundary value problems using mesh dependent norms. Numerische Mathematik, 34(1):41–62, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Banerjee:1990:EEN

[BO90] Uday Banerjee and John E. Osborn. Estimation of the effect of numerical integration in finite element eigenvalue approximation. Numerische Mathematik, 56(8):735–762, February 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bode:1968:MUS

[Bod68] Albert Bode. Matrizielle untere Schranken linearer Abbildungen und M-Matrizen. (German) [Matrix lower bounds of linear transformations and M matrices]. Numerische Mathematik, 11(5), June 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Boehm:1964:FWB

[Boe64] B. Boehm. Functions whose best rational Chebyshev approximation are polynomials. Numerische Mathematik, 6(1):235– 242, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bohl:1967:NAH

[Boh67]  Erich Bohl.  Nichtlineare

Aufgaben in halbgeordneten Räumen. (German) [Nonlinear problems in half-ordered spaces]. Numerische Mathematik, 10(3):220–231, October 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bohl:1969:UFN

[Boh69] Erich Bohl. Über Fehlerabschätzungen bei nichtlinearen Operatorgleichungen. (German) [On error bounds for nonlinear operator equations]. Numerische Mathematik, 13(3): 226–237, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bohl:1980:CTN

[Boh80] Erich Bohl. Chord techniques and Newton’s method for discrete bifurcation problems. Numerische Mathematik, 34(2): 111–124, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bohmer:1981:DNM

[Böh81] Klaus Böhmer. Discrete Newton methods and iterated defect corrections. Numerische Mathematik, 37(2):167–192, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bojanov:1993:IPS

[Boj93] Borislav Bojanov. Interpolation by periodic splines with Birkhoff knots. Numerische Mathematik, 65(1):63–75, May 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bollen:1984:NSD

[Bol84] Jo A. M. Bollen. Numerical stability of descent methods for solving linear equations. Numerische Mathematik, 43(3): 361–377, April 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Borgers:1989:NDD

[Bör89]  Christoph Börgers.  The

Neumann–Dirichlet domain decomposition method with inexact solvers on the subdomains. Numerische Mathematik, 55(2): 123–136, May 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Borgers:1990:DIM

[Bör90] Christoph Börgers. Domain imbedding methods for the Stokes equations. Numerische Mathematik, 57(5):435–451, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Borges:1994:CGL

[Bor94] Carlos F. Borges. On a class of Gauss-like quadrature rules. Numerische Mathematik, 67(3): 271–288, April 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067003/40670271.
htm;   http://science.springer.

                                        de/nmee/bibs/4067003/40670271.
                                        htm.

Bosarge:1971:ICS

[Bos71] W. E. Bosarge, Jr. Iterative continuation and the solution of nonlinear two-point boundary value problems. Numerische Mathematik, 17(4):268–283, August 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bourquin:1990:ACS

[Bou90] Frédéric Bourquin. Analysis and comparison of several component mode synthesis methods on one-dimensional domains. Numerische Mathematik, 58(1):11– 33, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buchanan:1966:UCM

[BP66] M. L. Buchanan and B. N. Parlett. The uniform convergence of matrix powers. Numerische Mathematik, 9(1):51– 54, November 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bercovier:1979:EEF

[BP79] M. Bercovier and O. Pironneau. Error estimates for finite element method solution of the Stokes problem in the primitive variables. Numerische Mathematik, 33(2):211–224, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bourdeau:1985:TGL

[BP85] Marc Bourdeau and Alain Pitre.

Tables of good lattices in four and five dimensions. Numerische Mathematik, 47(1):39– 43, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bermudez:1992:MAF

[BP92] Alfredo Bermúdez and Dolores G. Pedreira. Mathematical analysis of a finite element method without spurious solutions for computation of dielectric waveguides. Numerische Mathematik, 61(1):39– 57, February 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bojanov:1998:NID

[BP98] Borislav Bojanov and Guergana Petrova. Numerical integration over a disc. A new Gaussian quadrature formula. Numerische Mathematik, 80(1): 39–59, July 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080001/80800039.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800039.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800039.
ps.gz.

Babuska:1972:FES

[BR72] Ivo Babuška and Michael B. Rosenzweig. A finite element scheme for domains with corners. Numerische Mathematik, 20(1):1–21, February 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bank:1981:GAN

[BR81] R. E. Bank and D. J. Rose. Global approximate Newton methods. Numerische Mathematik, 37(2):279–295, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Blum:1988:ETR

[BR88] H. Blum and R. Rannacher. Extrapolation techniques for reducing the pollution effect of reentrant corners in the finite element method. Numerische Mathematik, 52(5): 539–564, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Begumisa:1991:SKE

[BR91] Annuntiato Begumisa and Ian Robinson. Suboptimal Kronrod extension formulae for numerical quadrature. Numerische Mathematik, 58(8): 807–818, March 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brakhage:1960:UNB

[Bra60]  H. Brakhage.  Über die

numerische Behandlung von Integralgleichungen nach der Quadraturformelmethode. (German) [On the numerical treatment of integral equations for the quadrature formula method]. Numerische Mathematik, 2(1):183–196, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brakhage:1961:FNE

[Bra61] H. Brakhage. Zur Fehlerabschätzung für die numerische Eigenwertbestimmung bei Integralgleichungen. (German) [Toward error bounds for numerical eigenvalue determination in integral equations]. Numerische Mathematik, 3:174–179, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:1965:MIG

[Bra65] Dietrich Braess. Monotone Iterationsfolgen bei Gleichungssystemen mit fehlerhaften Koeffizienten und Iterationsbeschleunigung. (German) [Monotone iteration sequences for equation systems with coefficients having errors, and iteration acceleration]. Numerische Mathematik, 7(1):32–41, February 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bramble:1966:SOF

[Bra66] James H. Bramble. A second order finite difference analog of the first biharmonic boundary value problem. Numerische Mathematik, 9(3):236–249, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bramble:1969:CDA

[Bra69] J. H. Bramble. On the convergence of difference approximations to weak solutions of Dirichlet’s problem. Numerische Mathematik, 13(2): 101–111, May 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:1970:MKI

[Bra70] Dietrich Braess. Eine Möglichkeit zur Konvergenzbeschleunigung bei Iterationsverfahren für bestimmte nichtlineare Probleme. (German) [A possibility for convergence acceleration of iterative methods for certain nonlinear problems]. Numerische Mathematik, 14(5): 468–475, April 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0029-
                                        599X&volume=14&issue=5&spage=
                                        468.

Braess:1971:CAS

[Bra71a] Dietrich Braess. Chebyshev approximation by spline functions with free knots. Numerische Mathematik, 17(5):357–366, October 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brauer:1971:RPH

[Bra71b] Alfred Brauer. A remark on the paper of H. H. Schaefer, “Abschätzung der nichttrivialen Eigenwerte stochastischer Matrizen”. Numerische Mathematik, 17(2):163–165, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brass:1973:FQC

[Bra73] Helmut Braß. Eine Fehlerabschätzung zum Quadratureverfahren von Clenshaw und Curtis. (German) [An error estimation of the quadrature method of Clenshaw and Curtis]. Numerische Mathematik, 21(5): 397–403, October 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:1974:RIN

[Bra74]  Dietrich Braess.  Rationale

Interpolation, Normalität und Monosplines. (German) [Rational interpolation, normality and monosplines]. Numerische Mathematik, 22(3): 219–232, June 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:1977:ENP

[Bra77]  Dietrich Braess.  Über die

Einzugsbereiche der Nullstellen von Polynomen beim Newton-Verfahren. (German) [On the regions of attraction when computing the roots of polynomials with the Newton-method]. Numerische Mathematik, 29(1): 123–132, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brass:1978:MQG

[Bra78] Helmut Braß. Monotonie bei den Quadratureverfahren von Gauß und Newton–Cotes. (German) [Monotony in the quadrature rules of Gauss and Newton– Cotes]. Numerische Mathematik, 30(3):349–354, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Braess:1981:CNM

[Bra81a] Dietrich Braess. The contraction number of a multigrid method for solving the Poisson equation. Numerische Mathematik, 37(3):387–404, July 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brauninger:1981:VMA

[Brä81b]  J. Bräuninger.  A variable

metric algorithm for unconstrained minimization without evaluation of derivatives. Numerische Mathematik, 36(4): 359–373, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brannigan:1982:SCS

[Bra82]  M. Brannigan.  The strict

Chebyshev solution of overdetermined systems of linear equations with rank deficient matrix. Numerische Mathematik, 40(3): 307–318, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brass:1985:FPQ

[Bra85] Helmut Braß. Eine Fehlerabschätzung für positive Quadraturformeln. (German) [An error estimate for positive quadrature formulas]. Numerische Mathematik, 47(3):395–399, November 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brand:1992:IFP

[Bra92]  Clemens W. Brand.  An

incomplete-factorization preconditioning using repeated redblack ordering. Numerische Mathematik, 61(4):433–454, May 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brandts:1994:SPE

[Bra94] Jan H. Brandts. Superconvergence and a posteriori error estimation for triangular mixed finite elements. Numerische Mathematik, 68(3):311–324, September 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.

                                        de/link/service/journals/
                                        00211/bibs/4068003/40680311.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4068003/40680311.
                                        htm.

Brent:1970:EAA

[Bre70] R. P. Brent. Error analysis of algorithms for matrix multiplication and triangular decomposition using Winograd’s identity. Numerische Mathematik, 16(2): 145–156, November 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1971:EAF

[Bre71] C. Brezinski. Études sur les ϵ- et ρ-algorithmes. (French) [Studies on ϵ and ρ-algorithms]. Numerische Mathematik, 17(2): 153–162, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1972:CAC

[Bre72] Claude Brezinski. Conditions d’application et de convergence de procédés d’extrapolation. (French) [Application and convergence conditions for some extrapolation processes]. Numerische Mathematik, 20(1):64– 79, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1975:FCA

[Bre75a] C. Brezinski. Forme confluente de l’ε-algorithme topologique. (French) [Confluent form of the topological ε-algorithm]. Numerische Mathematik, 23(4): 363–370, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezzi:1975:MEF

[Bre75b] F. Brezzi. Sur la méthode des éléments finis hybrides pour le problème biharmonique. (French) [On hybrid finite element methods for the biharmonic problem]. Numerische Mathematik, 24(2): 103–131, April 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1978:CAS

[Bre78] C. Brezinski. Convergence acceleration of some sequences by the ϵ-algorithm. Numerische Mathematik, 29(2):173–177, January 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1980:GEA

[Bre80] C. Brezinski. A general extrapolation algorithm. Numerische Mathematik, 35(2): 175–187, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1985:CST

[Bre85] Claude Brezinski. Composite sequence transformations. Numerische Mathematik, 46(3): 311–321, June 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1989:CPS

[Bre89] Claude Brezinski. Contraction properties of sequence transformations. Numerische Mathematik, 54(5):565–574, February 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brenner:1996:TLA

[Bre96] Susanne C. Brenner. A twolevel additive Schwarz preconditioner for nonconforming plate elements. Numerische Mathematik, 72(4):419–447, February 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6072004/60720419.
htm;   http://science.springer.
de/nmee/bibs/6072004/60720419.
htm.

Brenner:1999:CNS

[Bre99] Susanne C. Brenner. The condition number of the Schur complement in domain decomposition. Numerische Mathematik, 83(2):187–203, August 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083002/90830187.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083002/90830187.pdf.

Briones:1964:AAC

[Bri64] F. Briones. On the alternants appearing in Chebyshev’s best approximation problem. Numerische Mathematik, 6(1):211– 223, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brilla:1983:EAL

[Bri83] Jozef Brilla. Error analysis for Laplace transform—finite element solution of hyperbolic equations. Numerische Mathematik, 41(1):55–62, April 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Broyden:1964:SGT

[Bro64] C. G. Broyden. Some generalisations of the theory of successive over-relaxation. Numerische Mathematik, 6(1):269– 284, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brosowski:1965:ERT

[Bro65a] Bruno Brosowski. Über die Eindeutigkeit der rationalen Tschebyscheff-Approximationen. (German) [On the uniqueness of rational Chebyshev approximations]. Numerische Mathematik, 7(2):176–186, April 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brosowski:1965:ELP

[Bro65b] Bruno Brosowski. Über Extremalsignaturen linearer Polynome in n Veränderlichen. (German) [On the extremal signature of linear polynomials in n variables]. Numerische Mathematik, 7(5):396–405, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Broyden:1968:SAC

[Bro68] C. G. Broyden. Some aspects of consistent ordering. Numerische Mathematik, 12(1):47–56, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brombeer:1978:OQF

[Bro78] Rüdiger Brombeer. Über optimale Quadraturformeln mit freien Knoten zu Hilberträumen periodischer Funktionen. (German) [On optimal quadrature formulae with free nodes for Hilbert-spaces of periodic functions]. Numerische Mathematik, 30(2):149–164, June 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brokate:1985:NSH

[Bro85] M. Brokate. Numerical solution of a hyperbolic free boundary problem with a method of characteristics. Numerische Mathematik, 46(1):85–99, March 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezzi:1980:FDA

[BRR80] F. Brezzi, J. Rappaz, and P. A. Raviart. Finite-dimensional approximation of nonlinear problems. I. branches of nonsingular solutions. Numerische Mathematik, 36(1):1–25, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezzi:1981:FDAa

[BRR81a] F. Brezzi, J. Rappaz, and P. A.

Raviart. Finite-dimensional approximation of nonlinear problems. II. limit points. Numerische Mathematik, 37(1): 1–28, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezzi:1981:FDAb

[BRR81b] F. Brezzi, J. Rappaz, and P. A. Raviart. Finite dimensional approximation of nonlinear problems. III. simple bifurcation points. Numerische Mathematik, 38(1):1–30, October 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Burgmeier:1971:IET

[BRS71] J. W. Burgmeier, R. D. Russell, and L. F. Shampine. Integral equations with transform kernels and the eigenproblem KCx = λx. Numerische Mathematik, 17(4):253–267, August 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1992:BFL

[BRS92] C. Brezinski, M. Redivo Zaglia, and H. Sadok. A breakdown-free Lanczos type algorithm for solving linear systems. Numerische Mathematik, 63(1):29–38, October 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brunner:1978:DVI

[Bru78] Hermann Brunner. Discretization of Volterra integral equations of the first kind. II. Numerische Mathematik, 30(2): 117–136, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1998:ETI

[BRZRS98] C. Brezinski, M. Redivo-Zaglia, G. Rodriguez, and S. Seatzu. Extrapolation techniques for illconditioned linear systems. Numerische Mathematik, 81(1):1– 29, November 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8081001/80810001.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081001/80810001.
pdf.

Bauer:1959:NFT

[BS59] F. L. Bauer and H. J. Stetter. Zur numerischen Fourier– Transformation. (German) [The numerical Fourier transform]. Numerische Mathematik, 1(1): 208–220, December 1959. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
208.

Borsch-Supan:1960:BFB

[BS60] W. Börsch-Supan. Bemerkungen zur Fehlerabschätzung beim Ritz–Galerkin-Verfahren nach Krylow. (German) [Remark on the error bounds of the Ritz– Galerkin method of Krylov]. Numerische Mathematik, 2(1):79– 83, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Borsch-Supan:1963:PEB

[BS63] W. Börsch-Supan. A posteriori error bounds for the zeros of polynomials. Numerische Mathematik, 5(1):380–398, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1966:AUL

[BS66a] Roland Bulirsch and Josef Stoer.

Asymptotic upper and lower bounds for results of extrapolation methods. Numerische Mathematik, 8(2):93–104, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1966:NTO

[BS66b] Roland Bulirsch and Josef Stoer.

Numerical treatment of ordinary differential equations by extrapolation methods. Numerische Mathematik, 8(1):1–13, March 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1967:NQE

[BS67] Roland Bulirsch and Josef Stoer. Numerical quadrature by extrapolation. Numerische Mathematik, 9(4):271–278, February 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brosowski:1968:RTA

[BS68] Bruno Brosowski and Josef Stoer. Zur rationalen Tschebyscheff-Approximation differenzierbarer und analytischer Funktionen. (German) [On a rational Chebyshev approximation of differentiable and analytic functions]. Numerische Mathematik, 12(1): 57–65, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Borsch-Supan:1970:RPN

[BS70] Wolfgang Börsch-Supan. Residuenabschätzung für Polynom-Nullstellen mittels Lagrange-Interpolation. (German) [Residue bounds for polynomial zeros through Lagrange interpolation]. Numerische Mathematik, 14(3): 287–296, February 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baptist:1977:RBQ

[BS77] P. Baptist and J. Stoer. On the relation between quadratic termination and convergence properties of minimisation algorithms. II. applications. Numerische Mathematik, 28(4): 367–391, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrow:1979:EAS

[BS79] D. L. Barrow and P. W. Smith.

Efficient L_{2} approximation by splines. Numerische Mathematik, 33(1):101–114, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brass:1981:EEI

[BS81] H. Brass and G. Schmeisser.

Error estimates for interpolatory quadrature formulae. Numerische Mathematik, 37(3): 371–386, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bruch:1985:AIE

[BS85] John C. Bruch, Jr. and James M. Sloss. Alternating iteration and elliptic variational inequalities. Numerische Mathematik, 47(3):459–481, November 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Burmeister:1988:OCS

[BS88] W. Burmeister and Jochen W.

Schmidt. On the R-order of coupled sequences arising in single-step type methods. Numerische Mathematik, 53(6):

653–661, September 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:1989:TND

[BS89] I. Babuška and Manil Suri. The treatment of nonhomogeneous Dirichlet boundary conditions by the p-version of the finite element method. Numerische Mathematik, 55(1):97– 121, April 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1991:FEAa

[BS91] John W. Barrett and Robert M. Shanahan. Finite element approximation of a model reactiondiffusion problem with a non-Lipschitz nonlinearity. Numerische Mathematik, 59(3): 217–242, June 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:1992:LEF

[BS92a] Ivo Babuška and Manil Suri. Locking effects in the finite element approximation of elasticity problems. Numerische Mathematik, 62(4):439–463, September 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Baranger:1992:FEA

[BS92b] J. Baranger and D. Sandri. Finite element approximation of viscoelastic fluid flow: Existence of approximate solutions and error bounds. I. discontinuous constraints. Numerische Mathematik, 63(1):13–27, October 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Berthold:1995:CCM

[BS95] D. Berthold and B. Silbermann. Corrected collocation methods for periodic pseudodifferential equations. Numerische Mathematik, 70(4): 397–425, June 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070004/50700397.
htm;   http://science.springer.
de/nmee/bibs/5070004/50700397.
htm.

Bodenmann:1996:CDM

[BS96] Rudolf Bodenmann and H. Joachim Schroll. Compact difference methods applied to initialboundary value problems for mixed systems. Numerische Mathematik, 73(3):291–309, May 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.de/link/service/
journals/00211/bibs/6073003/
60730291.htm;   http://science.
springer.de/nmee/bibs/6073003/
60730291.htm.

Benzi:1997:EUS

[BS97a] Michele Benzi and Daniel B.

Szyld. Existence and uniqueness of splittings for stationary iterative methods with applications to alternating methods. Numerische Mathematik, 76(3): 309–321, May 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076003/70760309.
htm;   http://science.springer.
de/nmee/bibs/7076003/70760309.
htm.

Bereux:1997:RTR

[BS97b] François Bereux and Lionel Sainsaulieu. A Roe-type Riemann solver for hyperbolic systems with relaxation based on time-dependent wave decomposition. Numerische Mathematik, 77(2): 143–185, August 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077002/70770143.
htm;   http://science.springer.
de/nmee/bibs/7077002/70770143.
htm.

Borwein:1997:SAD

[BS97c] J. M. Borwein and W. Sun. The stability analysis of dynamic SPECT systems. Numerische Mathematik, 77(3):283–298, September 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077003/70770283.
htm;   http://science.springer.

de/nmee/bibs/7077003/70770283.
htm.

Bornemann:1998:MIC

[BS98a] Folkmar A. Bornemann and Christof Schütte. A mathematical investigation of the Car– Parrinello method. Numerische Mathematik, 78(3):359–376, January 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078003/70780359.
htm;   http://science.springer.
de/nmee/bibs/7078003/70780359.
htm.

Buss:1998:PMR

[BS98b] J. F. Buss and R. B. Simpson. Planar mesh refinement cannot be both local and regular. Numerische Mathematik, 79 (1):1–10, March 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079001/80790001.
htm;   http://science.springer.
de/nmee/bibs/8079001/80790001.
htm.

Barrett:1999:FEA

[BS99a] John W. Barrett and Mathias Schneider. Finite element approximation of a semilinear elliptic problem with a singular nonlinearity. Numerische Mathematik, 82(1):21–56, March 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9082001/90820021.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9082001/90820021.pdf.

Bornemann:1999:AAC

[BS99b] Folkmar A. Bornemann and Christof Schütte. Adaptive accuracy control for Car– Parrinello simulations. Numerische Mathematik, 83(2): 179–186, August 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/journals/
                                        00211/bibs/9083002/90830179.
                                        htm;   http://link.springer-
                                        ny.com/link/service/journals/
                                        00211/papers/9083002/90830179.
                                        pdf.

Bruntjen:1999:ITL

[BS99c] K. Brüntjen and H. Späth. Incomplete total least squares. Numerische Mathematik, 81 (4):521–538, February 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9081004/90810521.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9081004/90810521.pdf.

Birkhoff:1968:PHI

[BSV68] G. Birkhoff, M. H. Schultz, and R. S. Varga. Piecewise Hermite interpolation in one and two variables with applications to partial differential equations. Numerische Mathematik, 11(3): 232–256, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bauer:1961:AMN

[BSW61] F. L. Bauer, J. Stoer, and C. Witzgall. Absolute and monotonic norms. Numerische Mathematik, 3:257–264, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bruder:1988:PAR

[BSW88] Jürgen Bruder, Karl Strehmel, and Rüdiger Weiner. Partitioned adaptive Runge–Kutta methods for the solution of nonstiff and stiff systems. Numerische Mathematik, 52(6): 621–638, June 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buck:1962:SA

[Buc62] R. C. Buck. Stochastic ascent.

Numerische Mathematik, 4:177– 186, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Buchmann:1964:CUJ

[Buc64] M. Buchmann. A condition for uniform Jordanization of a family of triangular matrices. Numerische Mathematik, 6(1):335– 345, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1964:BRI

[Bul64] R. Bulirsch. Bemerkungen zur Romberg-Integration. (German) [Remark on Romberg integration]. Numerische Mathematik, 6(1):6–16, December 1964. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1965:NCEa

[Bul65a] Roland Bulirsch. Numerical calculation of elliptic integrals and elliptic functions. Numerische Mathematik, 7(1):78–90, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1965:NCEb

[Bul65b] Roland Bulirsch. Numerical calculation of elliptic integrals and elliptic functions. II. Numerische Mathematik, 7(4):353– 354, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1967:NCS

[Bul67] Roland Bulirsch. Numerical calculation of the sine, cosine and Fresnel integrals. Numerische Mathematik, 9(5):380– 385, April 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1969:EBT

[Bul69a] R. Bulirsch. An extension of the Bartky-transformation to incomplete elliptic integrals of the third kind. Numerische Mathematik, 13(3):266–284, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1969:NCE

[Bul69b] R. Bulirsch. Numerical calculation of elliptic integrals and elliptic functions. III. Numerische Mathematik, 13(4):305–315, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Businger:1968:MWC

[Bus68]  P. A. Businger.  Matrices

which can be optimally scaled. Numerische Mathematik, 12(4): 346–348, November 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Businger:1971:MNS

[Bus71] P. A. Businger. Monitoring the numerical stability of Gaussian elimination. Numerische Mathematik, 16(4):360–361, January 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bus:1977:CNM

[Bus77] J. C. P. Bus. Convergence of Newton-like methods for solving systems of nonlinear equations. Numerische Mathematik, 27(3): 271–281, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bernal:1968:GTC

[BV68] M. J. M. Bernal and J. H. Verner. On generalizations of the theory of consistent orderings for successive overrelaxation methods. Numerische Mathematik, 12(3):215–222, October 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Babuska:1984:FAF

[BV84] I. Babuška and M. Vogelius. Feedback and adaptive finite element solution of onedimensional boundary value problems. Numerische Mathematik, 44(1):75–102, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Balla:1987:RHP

[BV87] Katalin Balla and Mária Vicsek. On the reduction of Holt’s problem to a finite interval. Numerische Mathematik, 51(3): 291–302, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [BV88].

Balla:1988:ERH

[BV88] Katalin Balla and Mária Vicsek. Erratum: “On the reduction of Holt’s problem to a finite interval”. Numerische Mathematik, 52(4):479, April 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [BV87].

Beyer:1972:ECC

[BW72a] W. A. Beyer and M. S. Waterman. Ergodic computations with continued fractions and Jacobi’s algorithm. Numerische Mathematik, 19(3): 195–205, June 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See errata [BW72b].

Beyer:1972:EEC

[BW72b] W. A. Beyer and M. S. Waterman. Errata: “Ergodic Computations with Continued Fraction and Jacobi’s Algorithm”. Numerische Mathematik, 20 (5):430, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=20&issue=5&spage=

430. See [BW72a].

Baughman:1993:CVM

[BW93] Lisa A. Baughman and Noel J.

Walkington. Co-volume methods for degenerate parabolic problems. Numerische Mathematik, 64(1):45–67, January 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barrett:1995:FEA

[BW95] John W. Barrett and Patrick A.

Wood. The finite element approximation of a coupled reaction-diffusion problem with non-Lipschitz nonlinearities.

Numerische Mathematik, 71(2): 135–157, August 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071002/50710135.
htm;   http://science.springer.
de/nmee/bibs/5071002/50710135.
htm.

Bank:1999:MID

[BW99] Randolph E. Bank and Christian Wagner. Multilevel ILU decomposition. Numerische Mathematik, 82(4):543–576, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082004/90820543.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082004/90820543.pdf.

Brent:1973:OIP

[BWW73] Richard Brent, Shmuel Winograd, and Philip Wolfe. Optimal iterative processes for rootfinding. Numerische Mathematik, 20(5):327–341, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bank:1990:CIM

[BWY90] Randolph E. Bank, Bruno D. Welfert, and Harry Yserentant. A class of iterative methods for solving saddle point problems. Numerische Mathematik, 56(7): 645–666, January 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bank:1996:ACU

[BX96] Randolph E. Bank and Jinchao Xu. An algorithm for coarsening unstructured meshes. Numerische Mathematik, 73(1): 1–36, March 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073001/60730001.
htm;   http://science.springer.
de/nmee/bibs/6073001/60730001.
htm.

Barrodale:1966:ABL

[BY66] Ian Barrodale and Andrew Young. Algorithms for best L_{1} and L_{∞} linear approximations on a discrete set. Numerische Mathematik, 8(3):295–306, May 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bornemann:1993:BNE

[BY93] Folkmar Bornemann and Harry Yserentant. A basic norm equivalence for the theory of multilevel methods. Numerische Mathematik, 64(4): 455–476, April 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bar-Yoseph:1986:AFE

[BYI86] Pinchas Bar-Yoseph and Moshe Israeli. An asymptotic finite element method for improvement of solutions of boundary layer problems. Numerische Mathematik, 49(4):425–438, July 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brenner:1999:BDD

[ByS99] Susanne C. Brenner and Li yeng Sung. Balancing domain decomposition for nonconforming plate elements. Numerische Mathematik, 83(1):25–52, July 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083001/90830025.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083001/90830025.pdf.

Braess:1982:NTS

[BZ82] D. Braess and E. Zehnder. On the numerical treatment of a small divisor problem. Numerische Mathematik, 39(2): 269–292, August 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bellen:1985:NSD

[BZ85] Alfredo Bellen and Marino Zennaro. Numerical solution of delay differential equations by uniform corrections to an implicit Runge–Kutta method. Numerische Mathematik, 47(2): 301–316, September 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1994:HPS

[BZ94] C. Brezinski and M. Redivo Zaglia. Hybrid procedures for solving linear systems. Numerische Mathematik, 67(1):1– 19, February 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067001/40670001.
htm;   http://science.springer.
de/nmee/bibs/4067001/40670001.
htm.

Brezinski:1999:NLA

[BZS99] C. Brezinski, M. Redivo Zaglia, and H. Sadok. New lookahead Lanczos-type algorithms for linear systems. Numerische Mathematik, 83(1):53–85, July 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083001/90830053.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083001/90830053.pdf.

Cai:1991:ASA

[Cai91a]  Xiao-Chuan Cai.  Addi-

tive Schwarz algorithms for parabolic convection-diffusion equations. Numerische Mathematik, 60(1):41–61, November 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cai:1991:FVE

[Cai91b]  Zhiqiang Cai.  On the fi-

nite volume element method. Numerische Mathematik, 58(7): 713–735, February 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Campbell:1970:SAD

[Cam70] N. G. Campbell. Stability analysis of a difference scheme for the Navier–Stokes equations. Numerische Mathematik, 14(5): 435–447, April 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0029-
                                        599X&volume=14&issue=5&spage=
                                        435.

Campbell:1993:LSC

[Cam93] Stephen L. Campbell. Least squares completions for nonlinear differential algebraic equations. Numerische Mathematik, 65(1):77–94, May 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cannon:1971:CIS

[Can71] John J. Cannon. Computing the ideal structure of finite semigroups. Numerische Mathematik, 18(3):254–266, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carasso:1971:LRN

[Car71] Alfred Carasso. Long-range numerical solution of mildly non-linear parabolic equations. Numerische Mathematik, 16(4): 304–321, January 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carasso:1972:EAR

[Car72a] Claude Carasso. Étude de l’algorithme de Rémès en l’absence de condition de Haar. (French) [Study of the Remesalgorithm without the Haar condition]. Numerische Mathematik, 20(2):165–178, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carroll:1972:SAC

[Car72b] M. P. Carroll. Simultaneous L_{1} approximation of a compact set of real-valued functions. Numerische Mathematik, 19(2): 110–115, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carlson:1979:CEI

[Car79] B. C. Carlson. Computing elliptic integrals by duplication. Numerische Mathematik, 33(1): 1–16, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:1991:IRP

[Car91] C. Carstensen. Inclusion of the roots of a polynomial based on Gerschgorin’s theorem. Numerische Mathematik, 59(4): 349–360, July 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carpenter:1992:AZG

[Car92] Amos J. Carpenter. Asymptotics for the zeros of the generalized Bessel polynomials. Numerische Mathematik, 62(4): 465–482, September 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:1999:NAP

[Car99] Carsten Carstensen. Numerical analysis of the primal problem of elastoplasticity with hardening. Numerische Mathematik, 82(4):577–597, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082004/90820577.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082004/90820577.pdf.

Cash:1977:NIS

[Cas77] J. R. Cash. A note on the iterative solution of recurrence relations. Numerische Mathematik, 27(2):165–170, June 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cash:1978:HOM

[Cas78] J. R. Cash. High order methods for the numerical integration of ordinary differential equations. Numerische Mathematik, 30(4): 385–409, December 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cash:1980:NNS

[Cas80a]  J. R. Cash.  A note on

the numerical solution of linear recurrence relations. Numerische Mathematik, 34(4): 371–386, December 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cash:1980:ISS

[Cas80b] J. R. Cash. On the integration of stiff systems of O.D.E.s using extended backward differentiation formulae. Numerische Mathematik, 34(3):235–246, September 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cash:1981:HOS

[Cas81a] J. R. Cash. High order P -stable formulae for the numerical integration of periodic initial value problems. Numerische Mathematik, 37(3):355–370, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cash:1981:DHO

[Cas81b] J. R. Cash. On the design of high order exponentially fitted formulae for the numerical integration of stiff systems. Numerische Mathematik, 36(3): 253–266, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cash:1982:IAS

[Cas82] J. R. Cash. On improving the absolute stability of local extrapolation. Numerische Mathematik, 40(3):329–337, December 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cash:1983:SLM

[Cas83] J. R. Cash. Split linear multistep methods for the numerical integration of stiff differential systems. Numerische Mathematik, 42(3):299–310, December 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Casas:1985:EFE

[Cas85] Eduardo Casas. L^{2} estimates for the finite element method for the Dirichlet problem with singular data. Numerische Mathematik, 47(4):627–632, December 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Casas:1996:FFS

[Cas96] F. Casas. Fer’s factorization as a symplectic integrator. Numerische Mathematik, 74(3): 283–303, September 1996. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/6074003/
60740283.htm;   http://science.
springer.de/nmee/bibs/6074003/
60740283.htm.

Clenshaw:1960:MNI

[CC60] C. W. Clenshaw and A. R. Curtis. A method for numerical integration on an automatic computer. Numerische Mathematik, 2(1):197–205, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ciavaldini:1985:FEM

[CC85] J. F. Ciavaldini and M. Crouzeix.

A finite element method scheme for one-dimensional elliptic equations with high super convergence at the nodes. Numerische Mathematik, 46(3): 417–427, June 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Capatina:1991:IAQ

[CC91] Anca Radoslovescu Căpăţînă and M. Cocu. Internal approximation of quasi-variational inequalities. Numerische Mathematik, 59(4):385–398, July 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1995:AFE

[CC95] Zhangxin Chen and Bernardo Cockburn. Analysis of a finite element method for the drift-diffusion semiconductor device equations: The multidimensional case. Numerische Mathematik, 71(1): 1–28, July 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071001/50710001.
htm;   http://science.springer.
de/nmee/bibs/5071001/50710001.
htm.

Caselles:1993:GMA

[CCCD93] Vicent Caselles, Francine Catté, Tomeu Coll, and Françoise Dibos. A geometric model for active contours in image processing. Numerische Mathematik, 66(1):1–31, October 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chipot:1995:NAO

[CCK95] M. Chipot, C. Collins, and D. Kinderlehrer. Numerical analysis of oscillations in multiple well problems. Numerische Mathematik, 70(3): 259–282, May 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070003/50700259.
htm;   http://science.springer.
de/nmee/bibs/5070003/50700259.
htm.

Cenacchi:1972:QRS

[CD72] G. Cenacchi and A. De Matteis. Quasi-random sequences by power residues. Numerische Mathematik, 20(1):54– 63, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandra:1973:APS

[CD73] Jagdish Chandra and Paul William Davis. On the approximation of periodic solutions of equations with bounded nonlinearities. Numerische Mathematik, 21(2):96–108, April 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carlson:1979:ECI

[CD79] David Carlson and Biswa Nath Datta. On the effective computation of the inertia of a non-Hermitian matrix. Numerische Mathematik, 33(3): 315–322, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cortey-Dumont:1985:FEA

[CD85] Ph. Cortey-Dumont. On finite element approximation in the L^{∞}-norm of variational inequalities. Numerische Mathematik, 47(1):45–57, August 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chui:1990:CMQ

[CD90] Charles K. Chui and Harvey Diamond. A characterization of multivariate quasi-interpolation formulas and its applications. Numerische Mathematik, 57(2): 105–121, April 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chui:1991:GFL

[CD91] Charles K. Chui and Harvey Diamond. A general framework for local interpolation. Numerische Mathematik, 58(6): 569–581, January 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:1998:PEE

[CD98] Carsten Carstensen and Georg Dolzmann. A posteriori error estimates for mixed FEM in elasticity. Numerische Mathematik, 81(2):187–209, December 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8081002/80810187.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081002/80810187.
pdf.

Cuyt:1996:KTT

[CDL96] Annie Cuyt, Kathy Driver, and Doron Lubinsky. Kronecker type theorems, normality and continuity of the multivariate Padé operator. Numerische Mathematik, 73(3): 311–327, May 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073003/60730311.

htm;   http://science.springer.
de/nmee/bibs/6073003/60730311.
htm.

Cenacchi:1970:PNC

[CdM70] G. Cenacchi and A. de Matteis. Pseudo-random numbers for comparative Monte Carlo calculations. Numerische Mathematik, 16(1):11–15, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Conca:1998:RCE

[CDR98] Carlos Conca, Mario Durán, and Jacques Rappaz. Rate of convergence estimates for the spectral approximation of a generalized eigenvalue problem. Numerische Mathematik, 79(3): 349–369, May 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079003/80790349.
htm;   http://science.springer.
de/nmee/bibs/8079003/80790349.
htm.

Calvo:1999:NAT

[CDV99] N. Calvo, J. Durany, and C. Vázquez. Numerical approach of temperature distribution in a free boundary model for polythermal ice sheets. Numerische Mathematik, 83(4):557–580, October 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/

                                        9083004/90830557.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083004/90830557.pdf.

Copetti:1992:NAC

[CE92] M. I. M. Copetti and C. M. Elliott. Numerical analysis of the Cahn–Hilliard equation with a logarithmic free energy. Numerische Mathematik, 63(1):39– 65, October 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Censor:1983:SUK

[CEG83] Yair Censor, Paul P. B. Eggermont, and Dan Gordon. Strong underrelaxation in kaczmarz’s method for inconsistent systems. Numerische Mathematik, 41(1): 83–92, April 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1998:SMF

[CEL98] Hongsen Chen, Richard Ewing, and Raytcho Lazarov. Superconvergence of mixed finite element methods for parabolic problems with nonsmooth initial data. Numerische Mathematik, 78(4):495–521, February 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/7078004/70780495.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/7078004/70780495.
                                        htm.

Cody:1968:RCA

[CFH68] W. J. Cody, W. Fraser, and J. F.

Hart. Rational Chebyshev approximation using linear equations. Numerische Mathematik, 12(4):242–251, November 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:1997:ACF

[CFS97] C. Carstensen, S. A. Funken, and E. P. Stephan. On the adaptive coupling of FEM and BEM in 2-d-elasticity. Numerische Mathematik, 77(2): 187–221, August 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077002/70770187.
htm;   http://science.springer.
de/nmee/bibs/7077002/70770187.
htm.

Cheney:1959:NMC

[CG59] E. W. Cheney and A. A.

Goldstein. Newton’s method for convex programming and Tchebycheff approximation. Numerische Mathematik, 1(1):253– 268, December 1959. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
253.

Cooper:1967:SSM

[CG67] G. J. Cooper and E. Gal.

Single step methods for linear differential equations. Numerische Mathematik, 10(4): 307–315, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Calvo:1988:ASM

[CG88a] M. Calvo and T. Grande. On the asymptotic stability of θmethods for delay differential equations. Numerische Mathematik, 54(3):257–269, December 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandler:1988:CNM

[CG88b] G. A. Chandler and I. G. Graham. The convergence of Nyström methods for Wiener– Hopf equations. Numerische Mathematik, 52(3):345–364, March 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Campbell:1995:IGN

[CG95] Stephen L. Campbell and C. William Gear. The index of general nonlinear DAEs. Numerische Mathematik, 72(2): 173–196, December 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5072002/
50720173.htm;   http://science.
springer.de/nmee/bibs/5072002/
50720173.htm.

Chui:1995:FIW

[CGC95] C. K. Chui, J. C. Goswami, and A. K. Chan. Fast integral wavelet transform on a dense set of the time-scale domain. Numerische Mathematik, 70(3): 283–302, May 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070003/50700283.
htm;   http://science.springer.
de/nmee/bibs/5070003/50700283.
htm.

Calvo:1990:ZSV

[CGG90] M. Calvo, T. Grande, and Rolf Dieter Grigorieff. On the zero stability of the variable order variable stepsize BDF-Formulas. Numerische Mathematik, 57(1):39–50, March 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cavendish:1976:RAB

[CGH76] James C. Cavendish, William J.

Gordon, and Charles A. Hall. Ritz–Galerkin approximations in blending function spaces. Numerische Mathematik, 26(2): 155–178, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Champier:1993:CUF

[CGH93] S. Champier, T. Gallouët, and R. Herbin. Convergence of an upstream finite volume scheme for a nonlinear hyperbolic equation on a triangular mesh.

Numerische Mathematik, 66(2): 139–157, November 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cai:1992:CRE

[CGK92] Xiao-Chuan Cai, William D. Gropp, and David E. Keyes. Convergence rate estimate for a domain decomposition method. Numerische Mathematik, 61(2): 153–169, March 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Calvo:1998:CRM

[CGPM98] M. Calvo, S. Gonzalez-Pinto, and J. I. Montijano. On the convergence of Runge–Kutta methods for stiff non linear differential equations. Numerische Mathematik, 81(1):31– 51, November 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8081001/80810031.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081001/80810031.
pdf.

Cordova:1990:VMC

[CGR90] Antonio Córdova, Walter Gautschi, and Stephan Ruscheweyh. Vandermonde matrices on the circle: spectral properties and conditioning. Numerische Mathematik, 57(6/7):577–591, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Calvetti:1994:ACI

[CGR94] D. Calvetti, G. H. Golub, and L. Reichel. An adaptive Chebyshev iterative method for nonsymmetric linear systems based on modified moments. Numerische Mathematik, 67(1):21– 40, February 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067001/40670021.
htm;   http://science.springer.
de/nmee/bibs/4067001/40670021.
htm.

Conn:1994:NUA

[CGT94] A. R. Conn, Nick Gould, and Ph. L. Toint. A note on using alternative second-order models for the subproblems arising in barrier function methods for minimization. Numerische Mathematik, 68(1): 17–33, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680017.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680017.
htm.

Causey:1960:CAE

[CH60] Robert L. Causey and Peter Henrici. Convergence of approximate eigenvectors in Jacobi methods. Numerische Mathematik, 2(1):67–78, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carlson:1971:RAT

[CH71] R. E. Carlson and C. A. Hall. Ritz approximations to twodimensional boundary value problems. Numerische Mathematik, 18(2):171–181, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cools:1986:OAK

[CH86] Ronald Cools and Ann Haegemans. Optimal addition of knots to cubature formulae for planar regions. Numerische Mathematik, 49(2/3):269–274, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cools:1989:CMD

[CH89] Ronald Cools and Ann Haegemans. On the construction of multi-dimensional embedded cubature formulae. Numerische Mathematik, 55(6):735–745, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chakravarti:1961:TEI

[Cha61] P. C. Chakravarti. Truncation error in interpolation and numerical differentiation. Numerische Mathematik, 3(1):279– 284, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
279.

Chawla:1971:EEN

[Cha71] M. M. Chawla. Estimating errors of numerical approximation for analytic functions. Numerische Mathematik, 16(4): 370–374, January 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:1973:SS

[Cha73a] Paula P. Chan. Singular splines.

Numerische Mathematik, 20(5): 342–349, October 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chatelin:1973:ADS

[Cha73b] F. Chatelin. Approximation du spectre d’un opérateur linéare: application aux opérateurs differentiels elliptiques non autoadjoints. (French) [Approximation of the spectrum of a linear operator: application to nonselfadjoint differential elliptic operators]. Numerische Mathematik, 20(3):193–204, June 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:1974:MOM

[Cha74] Paula P. Chan. A monotone operator method for the solution of Fredholm integral equations. Numerische Mathematik, 22(5): 403–408, October 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chatelin:1979:BEP

[Cha79] Françoise Chatelin. Sur les bornes d’erreur a posteriori pour les éléments propres d’opérateurs linéaires. (French) [On the a posteriori error bounds for the proper elements of linear operators]. Numerische Mathematik, 32(3): 233–246, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chaffy:1984:CHA

[Cha84] Claudine Chaffy. Une construction “homogène” d’approximants de Padé à deux variables complexes. (French) [A “homogeneous” construction of Padé approximants in two complex variables]. Numerische Mathematik, 45(1):149–164, October 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:1987:IMO

[Cha87] R. H. Chan. Iterative methods for overflow queueing models. I. Numerische Mathematik, 51(2):143–180, May 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:1988:IMO

[Cha88] R. H. Chan. Iterative methods for overflow queuing models. II. Numerische Mathematik, 54 (1):57–78, October 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandler:1992:MCC

[Cha92] G. A. Chandler. Midpoint collocation for Cauchy singular integral equations. Numerische Mathematik, 62(4): 483–509, September 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chapelle:1997:LFA

[Cha97] Dominique Chapelle. A lockingfree approximation of curved rods by straight beam elements. Numerische Mathematik, 77(3): 299–322, September 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7077003/
70770299.htm;   http://science.
springer.de/nmee/bibs/7077003/
70770299.htm.

ChaconRebollo:1998:TTS

[Cha98] Tomás Chacón Rebollo. A term by term stabilization algorithm for finite element solution of incompressible flow problems. Numerische Mathematik, 79(2): 283–319, April 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079002/80790283.
htm;   http://science.springer.
de/nmee/bibs/8079002/80790283.
htm.

Chatzipantelidis:1999:FVM

[Cha99] Panagiotis Chatzipantelidis. A finite volume method based on the Crouzeix–Raviart element for elliptic PDE’s in two dimensions. Numerische Mathematik, 82(3):409–432, May 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082003/90820409.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082003/90820409.pdf.

Chenin:1979:RIE

[Che79] Patrick Chenin. Une représentation intégrale de l’erreur d’interpolation. (French) [An integral representation of interpolation error]. Numerische Mathematik, 32(4): 461–479, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1997:MFE

[Che97] Zhiming Chen. Mixed finite element methods for a dynamical Ginzburg–Landau model in superconductivity. Numerische Mathematik, 76(3): 323–353, May 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076003/70760323.
htm;   http://science.springer.
de/nmee/bibs/7076003/70760323.
htm.

Chipot:1991:NAO

[Chi91] M. Chipot. Numerical analysis of oscillations in nonconvex problems. Numerische Mathematik, 59(8):747–767, October 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chipot:1999:AMP

[Chi99] Michel Chipot. The appearance of microstructures in problems with incompatible wells and their numerical approach. Numerische Mathematik, 83 (3):325–352, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083003/90830325.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083003/90830325.pdf.

Canuto:1985:SME

[CHL85] C. Canuto, S. I. Hariharan, and L. Lustman. Spectral methods for exterior elliptic problems. Numerische Mathematik, 46(4): 505–520, August 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chow:1989:FEE

[Cho89] S.-S. Chow. Finite element error estimates for non-linear elliptic equations of monotone type. Numerische Mathematik, 54(4): 373–393, January 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Choudury:1990:FDG

[Cho90] G. Choudury. Fully discrete Galerkin approximations of parabolic boundary-value problems with nonsmooth boundary data. Numerische Mathematik, 57(2):179–203, April 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Christiansen:1970:ENS

[Chr70a] Jan Christiansen. Erratum: Numerical solution of ordinary simultaneous differential equations of the 1st order using a method for automatic step change. Numerische Mathematik, 15(1):v, May 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Chr70b].

Christiansen:1970:NSO

[Chr70b]  Jan Christiansen.  Numeri-

cal solution of ordinary simultaneous differential equations of the 1st order using a method for automatic step change. Numerische Mathematik, 14(4):317–324, March 1970. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Chr70a].

Clement:1996:IMA

[CHS96] Ph. Clément, R. Hagmeijer, and G. Sweers. On the invertibility of mappings arising in 2D grid generation problems. Numerische Mathematik, 73(1): 37–51, March 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073001/60730037.
htm;   http://science.springer.
de/nmee/bibs/6073001/60730037.
htm.

Chung:1976:GAA

[Chu76] Soon Park Chung. Generalization and acceleration of an algorithm of Sebastiao e Silva and its duals. Numerische Mathematik, 25(4):365–377, December 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chung:1979:AZT

[Chu79] Soon Park Chung. An algorithm for the zeros of transcendental functions. Numerische Mathematik, 32(4):359–371, December 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chu:1983:NTC

[Chu83] M. T. Chu. On a numerical treatment for the curve-tracing of the homotopy method. Numerische Mathematik, 42(3):

323–329, December 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandrasekaran:1994:BEE

[CI94] Shivkumar Chandrasekaran and Ilse C. F. Ipsen. Backward errors for eigenvalue and singular value decompositions. Numerische Mathematik, 68(2): 215–223, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068002/40680215.
htm;   http://science.springer.
de/nmee/bibs/4068002/40680215.
htm.

Ciarlet:1990:NCV

[Cia90] P. G. Ciarlet. A new class of variational problems arising in the modeling of elastic multistructures. Numerische Mathematik, 57(6/7):547–560, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Ciarlet:1998:ERV

[Cia98] Philippe Ciarlet. Editorial: [to Richard S. Varga, seventieth birthday greetings]. Numerische Mathematik, 80(4):495, October 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See their earliest joint paper [CSV67].

Chawla:1974:ORP

[CK74] M. M. Chawla and Veena Kaul.

Optimal rules with polynomial precision for Hilbert spaces possessing reproducing kernel functions. Numerische Mathematik, 22(3):207–218, June 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chawla:1976:ORP

[CK76] M. M. Chawla and V. Kaul.

Optimal rules with polynomial precision for Hilbert spaces possessing reproducing kernel functions. II. error bounds. Numerische Mathematik, 25(4): 379–382, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chawla:1982:FDM

[CK82] M. M. Chawla and C. P. Katti.

Finite difference methods and their convergence for a class of singular two-point boundary value problems. Numerische Mathematik, 39(3):341–350, October 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:1998:FPS

[CKL98] C. Carstensen, M. Kuhn, and U. Langer. Fast parallel solvers for symmetric boundary element domain decomposition equations. Numerische Mathematik, 79(3): 321–347, May 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079003/80790321.
htm;   http://science.springer.
de/nmee/bibs/8079003/80790321.
htm.

Coomes:1995:RCS

[CKP95] Brian A. Coomes, Hüseyin Koçak, and Kenneth J. Palmer. Rigorous computational shadowing of orbits of ordinary differential equations. Numerische Mathematik, 69(4):401–421, February 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069004/50690401.
htm;   http://science.springer.
de/nmee/bibs/5069004/50690401.
htm.

Caselles:1997:MSG

[CKSS97] Vicent Caselles, Ron Kimmel, Guillermo Sapiro, and Catalina Sbert. Minimal surfaces: a geometric three dimensional segmentation approach. Numerische Mathematik, 77(4): 423–451, October 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7077004/
70770423.htm;   http://science.
springer.de/nmee/bibs/7077004/
70770423.htm.

Cheney:1961:TNA

[CL61] E. W. Cheney and H. L. Loeb.

Two new algorithms for rational approximation. Numerische Mathematik, 3(1):72– 75, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cheney:1962:RCA

[CL62] E. W. Cheney and H. L. Loeb.

On rational Chebyshev approximation. Numerische Mathematik, 4:124–127, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chatelin-Laborde:1971:PMH

[CL71] Françoise Chatelin-Laborde.

Perturbation d’une matrice hermitienne ou normale. (French) [Perturbation theory for Hermitian (normal) matrix]. Numerische Mathematik, 17(4): 318–337, August 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chatelin:1975:MRR

[CL75] F. Chatelin and M. J. Lemordant. La méthode de Rayleigh– Ritz appliquée à des opérateurs différentielles elliptiques — ordres de convergence des éléments propres. (French) [The Rayleigh–Ritz method applied to elliptic differential operators — orders of convergence of proper elements]. Numerische Mathematik, 23(3): 215–222, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1994:SAE

[CL94] Hongsen Chen and Bo Li. Superconvergence analysis and error expansion for the Wilson nonconforming finite element. Numerische Mathematik, 69(2): 125–140, December 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4069002/
40690125.htm;   http://science.
springer.de/nmee/bibs/4069002/
40690125.htm.

Coquel:1996:ESM

[CL96a] Frédéric Coquel and Philippe G. LeFloch. An entropy satisfying MUSCL scheme for systems of conservation laws. Numerische Mathematik, 74(1): 1–33, July 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074001/60740001.
htm;   http://science.springer.
de/nmee/bibs/6074001/60740001.
htm.

Crandall:1996:CDS

[CL96b] Michael G. Crandall and Pierre-Louis Lions. Convergent difference schemes for nonlinear parabolic equations and mean curvature motion. Numerische Mathematik, 75(1):17– 41, November 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6075001/60750017.
htm;   http://science.springer.
de/nmee/bibs/6075001/60750017.
htm.

Chambolle:1997:IRT

[CL97] Antonin Chambolle and Pierre-Louis Lions. Image recovery via total variation minimization and related problems. Numerische Mathematik, 76(2): 167–188, April 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076002/70760167.
htm;   http://science.springer.
de/nmee/bibs/7076002/70760167.
htm.

Claessens:1976:NAO

[Cla76] G. Claessens. A new algorithm for oscillatory rational interpolation. Numerische Mathematik, 27(1):77–83, March 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Claessens:1978:UIR

[Cla78] G. Claessens. A useful identity for the rational Hermite interpolation table. Numerische Mathematik, 29(2):227–231, January 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Clendenin:1966:MNC

[Cle66] W. W. Clendenin. A method for numerical calculation of Fourier integrals. Numerische Mathematik, 8(5):422–436, August 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Clendenin:1971:MNQ

[Cle71] W. W. Clendenin. A method for numerical quadrature over an infinite interval using systematically increasing mesh size. Numerische Mathematik, 16(5): 397–407, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Clements:1992:CPP

[Cle92] John C. Clements. A convexitypreserving C^{2} parametric rational cubic interpolation. Numerische Mathematik, 63(2): 165–171, November 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chow:1962:SLE

[CM62] Tse-Sun Chow and Harold W. Milnes. Solution of Laplace’s equation by boundary contraction over regions of irregular shape. Numerische Mathematik, 4:209–225, December 1962. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chow:1970:NSL

[CM70] T. S. Chow and H. W. Milnes.

Numerical solution of Laplace equation for annular region. Numerische Mathematik, 16(3): 231–242, December 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Crandall:1980:MFS

[CM80] Michael Crandall and Andrew Majda. The method of fractional steps for conservation laws. Numerische Mathematik, 34(3):285–314, September 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Criscuolo:1989:UCG

[CM89] G. Criscuolo and G. Mastroianni. On the uniform convergence of Gaussian quadrature rules for Cauchy principal value integrals. Numerische Mathematik, 54(4):445–461, January 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Costabel:1992:SCS

[CM92] Martin Costabel and William McLean. Spline collocation for strongly elliptic equations on the torus. Numerische Mathematik, 62(4):511–538, September 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Campbell:1995:CPI

[CM95] Stephen L. Campbell and Edward Moore. Constraint preserving integrators for general nonlinear higher index DAEs. Numerische Mathematik, 69(4): 383–399, February 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5069004/
50690383.htm;   http://science.
springer.de/nmee/bibs/5069004/
50690383.htm.

Cottet:1990:PMS

[CMG90] G. H. Cottet and S. Mas-Gallic. A particle method to solve the Navier–Stokes system. Numerische Mathematik, 57(8): 805–827, August 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Criscuolo:1991:UCD

[CMO91] Giuliana Criscuolo, Giuseppe Mastroianni, and Donatella Occorsio. Uniform convergence of derivatives of extended Lagrange interpolation. Numerische Mathematik, 60(2): 195–218, November 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Canuto:1982:ACF

[CMQ82] C. Canuto, Y. Maday, and A. Quarteroni. Analysis of the combined finite element and Fourier interpolation. Numerische Mathematik, 39(2): 205–220, August 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Canuto:1984:CFE

[CMQ84] Claudio Canuto, Yvon Maday, and Alfio Quarteroni. Combined finite element and spectral approximation of the Navier– Stokes equations. Numerische Mathematik, 44(2):201–217, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cosnard:1986:PDR

[CMR86] M. Cosnard, J.-M. Muller, and Y. Robert. Parallel QR decomposition of a rectangular matrix. Numerische Mathematik, 48(2): 239–249, February 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Choi:1993:MPM

[CMR93] Y. S. Choi, P. J. McKenna, and M. Romano. A mountain pass method for the numerical solution of semilinear wave equations. Numerische Mathematik, 64(4):487–509, April 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Clenshaw:1963:ASF

[CMW63] C. W. Clenshaw, G. F. Miller, and M. Woodger. Algorithms for special functions. I. Numerische Mathematik, 4:403–419, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ciarlet:1970:NMH

[CNV70] P. G. Ciarlet, F. Natterer, and R. S. Varga. Numerical methods of high-order accuracy for singular nonlinear boundary value problems. Numerische Mathematik, 15(2):87–99, June 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Csordas:1988:LBB

[CNV88] G. Csordas, T. S. Norfolk, and R. S. Varga. A lower bound for the de Bruijn-Newman constant Λ. Numerische Mathematik, 52 (5):483–497, May 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cochran:1965:ZHF

[Coc65] James Alan Cochran. The zeros of Hankel functions as functions of their order. Numerische Mathematik, 7(3):238–250, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cohn:1959:NSR

[Coh59] Harvey Cohn. Numerical study of the representation of a totally positive quadratic integer as the sum of quadratic integral squares. Numerische Mathematik, 1:121–134, December 1959. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Cohn:1960:NSW

[Coh60] Harvey Cohn. A numerical study of Weber’s real class number calculation. Part I. Numerische Mathematik, 2(1):347– 362, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=2&issue=1&spage=
347.

Cohen:1994:PP

[Coh94] A. M. Cohen. Is the polynomial so perfidious? Numerische Mathematik, 68(2): 225–238, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068002/40680225.
htm;   http://science.springer.
de/nmee/bibs/4068002/40680225.
htm.

Collatz:1961:MEB

[Col61]  Lothar Collatz.  Monotonie

und Extremalprinzipien beim Newtonschen Verfahren. (German) [Monotonicty and extremal principles of Newton methods]. Numerische Mathematik, 3:99–106, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Comet:1959:UAB

[Com59] Stig Comét. Über die Anwendung von Binärmodellen in der Theorie der Charaktere der symmetrischen Gruppen. (German) [On the use of the binary model in the theory of characters of symmetric groups]. Numerische Mathematik, 1:90–109, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Comodi:1985:ABP

[Com85] M. I. Comodi. Approximation of a bending plate problem with a boundary unilateral constraint. Numerische Mathematik, 47(3): 435–458, November 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

ContesseB:1980:CCM

[Con80] Luis Contesse B. Une caractérisation complète des minima locaux en programmation quadratique. (French) [A complete characterization of local minima in quadratic programming]. Numerische Mathematik, 34(3):315–332, September 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Conca:1984:AQP

[Con84]  Carlos Conca.  Approxima-

tion de quelques problèmes de type Stokes par une méthode d’éléments finis mixtes. (French) [Mixed finite element approximation of Stokes-type problems]. Numerische Mathematik, 45(1):75–91, October 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cooper:1971:EBN

[Coo71] G. J. Cooper. Error bounds for numerical solutions of ordinary differential equations. Numerische Mathematik, 18(2): 162–170, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cordellier:1977:PRV

[Cor77] F. Cordellier. Particular rules for the vector ϵ-algorithm. Numerische Mathematik, 27(2): 203–207, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cranley:1970:RMM

[CP70] R. Cranley and T. N. L. Patterson. A regression method for the Monte Carlo evaluation of multidimensional integrals. Numerische Mathematik, 16(1):58– 72, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cerutti:1976:CMP

[CP76] John H. Cerutti and Seymour V.

Parter. Collocation methods for parabolic partial differential equations in one space dimension. Numerische Mathematik, 26(3):227–254, September 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carnicer:1994:LSB

[CP94a] J. M. Carnicer and J. M. Peña. Least supported bases and local linear independence. Numerische Mathematik, 67(3): 289–301, April 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067003/40670289.
htm;   http://science.springer.
de/nmee/bibs/4067003/40670289.
htm.

Chenais:1994:LPC

[CP94b] D. Chenais and J. C. Paumier. On the locking phenomenon for a class of elliptic problems. Numerische Mathematik, 67(4): 427–440, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067004/40670427.
htm;   http://science.springer.
de/nmee/bibs/4067004/40670427.
htm.

Capatina-Papaghiuc:1998:NFE

[CPT98] Daniela Capatina-Papaghiuc and Jean-Marie Thomas. Nonconforming finite element methods without numerical locking. Numerische Mathematik, 81(2): 163–186, December 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/

journals/00211/bibs/8081002/
80810163.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8081002/
80810163.pdf.

Chawla:1974:NEI

[CR74] M. M. Chawla and T. R. Ramakrishnan. Numerical evaluation of integrals of periodic functions with Cauchy and Poisson type kernels. Numerische Mathematik, 22(4):317–323, August 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Christiansen:1980:DCU

[CR80] J. Christiansen and R. D. Russell. Deferred corrections using uncentered end formulas. Numerische Mathematik, 35(1): 21–33, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Charrier:1989:ACC

[CR89] P. Charrier and J. Roman. Algorithmique et calculs de complexité pour un solveur de type dissections emboî tées. (French) [Algorithmic study and complexity bounds for a nested dissection solver]. Numerische Mathematik, 55(4): 463–476, July 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carstensen:1993:CHO

[CR93] Carsten Carstensen and Martin Reinders. On a class of higher-order methods for simultaneous root finding of generalized polynomials. Numerische Mathematik, 64(1):69–84, January 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Calvetti:1994:ABM

[CR94] D. Calvetti and L. Reichel. Application of a block modified Chebyshev algorithm to the iterative solution of symmetric linear systems with multiple right hand side vectors. Numerische Mathematik, 68(1): 3–16, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680003.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680003.
htm.

Cromme:1976:KVE

[Cro76] Ludwig Cromme. Eine Klasse von Verfahren zur Ermittlung bester nichtlinearer Tschebyscheff-Approximationen. (German) [A class of numerical procedures for finding best nonlinear Chebyshev approximations]. Numerische Mathematik, 25(4): 447–459, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cromme:1977:NMB

[Cro77] Ludwig Cromme. Numerische Methoden zur Behandlung einiger Problemklassen der nichtlinearen Tschebyscheff-Approximation mit Nebenbedingungen. (German) [Numerical methods for constrained nonlinear Chebyshev approximation]. Numerische Mathematik, 28(1):101–117, March 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cromme:1978:SUF

[Cro78]  Ludwig Cromme.  Strong

uniqueness: a far-reaching criterion for the convergence analysis of iterative procedures. Numerische Mathematik, 29(2): 179–193, January 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Crouzeix:1979:SMR

[Cro79]  Michel Crouzeix.  Sur la

B-stabilité des méthodes de Runge–Kutta. (French) [On the B-stability of Runge–Kutta methods]. Numerische Mathematik, 32(1):75–82, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Crouzeix:1980:MMI

[Cro80] Michel Crouzeix. Une méthode multipas implicite-explicite pour l’approximation des équations d’évolution paraboliques. (French) [A multipass explicit-implicit method for the approximation of parabolic evolutionary equations]. Numerische Mathematik, 35(3):257–276, October 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cryer:1967:CLE

[Cry67] C. W. Cryer. On the calculation of the largest eigenvalue of an integral equation. Numerische Mathematik, 10(2): 165–176, July 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cryer:1968:PSG

[Cry68] C. W. Cryer. Pivot size in Gaussian elimination. Numerische Mathematik, 12(4): 335–345, November 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cryer:1973:NSB

[Cry73] Colin W. Cryer. The numerical solution of boundary value problems for second order functional differential equations by finite differences. Numerische Mathematik, 20(4):288–299, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Calvetti:1999:IEF

[CRZ99] D. Calvetti, L. Reichel, and Q. Zhang. Iterative exponential filtering for large discrete ill-posed problems. Numerische Mathematik, 83(4):535–556, October 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083004/90830535.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083004/90830535.pdf.

Collatz:1959:ELR

[CS59] Lothar Collatz and J. Schroeder.

Einschließen der Lösungen von Randwertaufgaben. (German) [Bracketing solutions to boundary value problems]. Numerische Mathematik, 1:61– 72, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cody:1966:RCA

[CS66] W. J. Cody and J. Stoer.

Rational Chebyshev approximations using interpolation. Numerische Mathematik, 9(3):177– 188, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [CS68].

Cody:1968:CRC

[CS68] W. J. Cody and J. Stoer. Corrigendum: “Rational Chebyshev approximations using interpolation”. Numerische Mathematik, 12(3):230, October 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [CS66].

Chartres:1976:CLM

[CS76] Bruce A. Chartres and Robert S.

Stepleman. Convergence of linear multistep methods for differential equations with discontinuities. Numerische Mathematik, 27(1):1–10, March 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chocholaty:1979:NMB

[CŠ79] Pavel Chocholaty and L’udomír Šlahor. A numerical method to boundary value problems for second order delay differential equations. Numerische Mathematik, 33(1):69–75, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shi:1984:CCQ

[cS84]  Zhong ci Shi.  A conver-

gence condition for the quadrilateral Wilson element. Numerische Mathematik, 44(3): 349–361, September 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Canuto:1986:AKS

[CS86] Claudio Canuto and Giovanni Sacchi Landriani. Analysis of the Kleiser-Schumann method. Numerische Mathematik, 50(2): 217–243, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Costabel:1988:DEN

[CS88] Martin Costabel and Ernst P. Stephan. Duality estimates for the numerical solution of integral equations. Numerische Mathematik, 54(3):

339–353, December 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chandler:1990:SQM

[CS90] G. A. Chandler and I. H.

Sloan. Spline qualocation methods for boundary integral equations. Numerische Mathematik, 58(5):537–567, December 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [CS92].

Chandler:1992:EQM

[CS92] G. A. Chandler and I. H. Sloan.

Erratum: “Spline qualocation methods for boundary integral equations” [Numer. Math. 58 (1990), no. 5, 537–567 and MR 91m:65323]. Numerische Mathematik, 62(2):295, July 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [CS90].

Calvo:1994:CS

[CS94] M. P. Calvo and J. M. Sanz-Serna. Canonical B-series. Numerische Mathematik, 67(2): 161–175, March 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067002/40670161.
htm;   http://science.springer.
de/nmee/bibs/4067002/40670161.
htm.

Crouzeix:1998:AEE

[CS98] Michel Crouzeix and Francisco-Javier Sayas. Asymptotic expansions of the error of spline Galerkin boundary element methods. Numerische Mathematik, 78(4):523–547, February 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078004/70780523.
htm;   http://science.springer.
de/nmee/bibs/7078004/70780523.
htm.

Ciarlet:1967:NMH

[CSV67] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. I: One dimensional problem. Numerische Mathematik, 9(5):394–430, April 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See birthday greetings [Cia98].

Ciarlet:1968:NMHa

[CSV68a] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. II: Nonlinear boundary conditions. Numerische Mathematik, 11(4):331–345, May 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ciarlet:1968:NMHb

[CSV68b] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. III: Eigenvalue problems. Numerische Mathematik, 12(2): 120–133, August 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ciarlet:1968:NMHc

[CSV68c] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. IV: Periodic boundary conditions. Numerische Mathematik, 12(4):266–279, November 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ciarlet:1969:NMH

[CSV69] P. G. Ciarlet, M. H. Schultz, and R. S. Varga. Numerical methods of high-order accuracy for nonlinear boundary value problems. V: Monotone operator theory. Numerische Mathematik, 13(1):51–77, March 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chan:1996:OSM

[CSZ96] Tony F. Chan, Barry F.

Smith, and Jun Zou. Overlapping Schwarz methods on unstructured meshes using nonmatching coarse grids. Numerische Mathematik, 73(2): 149–167, April 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.

                                        de/link/service/journals/
                                        00211/bibs/6073002/60730149.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/6073002/60730149.
                                        htm.

Chen:1991:IUS

[CT91] Min Chen and Roger Temam. Incremental unknowns for solving partial differential equations. Numerische Mathematik, 59(3):255–271, June 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1993:NGM

[CT93] Min Chen and Roger Temam. Nonlinear Galerkin method in the finite difference case and wavelet-like incremental unknowns. Numerische Mathematik, 64(3):271–294, March 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Culpin:1986:CCS

[Cul86] David Culpin. Calculation of cubic smoothing splines for equally spaced data. Numerische Mathematik, 48(6):627–638, June 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cuppen:1981:DCM

[Cup81] J. J. M. Cuppen. A divide and conquer method for the symmetric tridiagonal eigenproblem. Numerische Mathematik, 36(2):177–195, April 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cuppen:1984:UTP

[Cup84] J. J. M. Cuppen. On updating triangular products of Householder reflections. Numerische Mathematik, 45(3): 403–409, December 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Curtis:1970:EOR

[Cur70] A. R. Curtis. An eighth order Runge–Kutta process with eleven function evaluations per step. Numerische Mathematik, 16(3):268–277, December 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cuyt:1982:AMP

[Cuy82]  Annie A. M. Cuyt.  The

ε-algorithm and multivariate Padé-approximants. Numerische Mathematik, 40(1):39– 46, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cuyt:1983:ACT

[Cuy83a]  Annie A. M. Cuyt.  Ac-

celerating the convergence of a table with multiple entry. Numerische Mathematik, 41(2): 281–286, June 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cuyt:1983:QAM

[Cuy83b]  Annie A. M. Cuyt.  The

QD-algorithm and multivariate Padé-approximants. Numerische Mathematik, 42(3): 259–269, December 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ciarlet:1970:DVG

[CV70] P. G. Ciarlet and R. S. Varga. Discrete variational Green’s function. Numerische Mathematik, 16(2):115–128, November 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Csordas:1984:CRS

[CV84a] George Csordas and Richard S. Varga. Comparisons of regular splittings of matrices. Numerische Mathematik, 44(1): 23–35, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cuyt:1984:GON

[CV84b] Annie A. M. Cuyt and Brigitte M. Verdonk. General order Newton–Padé approximants for multivariate functions. Numerische Mathematik, 43(2): 293–307, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ciarlet:1994:MAS

[CV94] Philippe G. Ciarlet and Richard S. Varga. In memoriam: Alston Scott Householder, 1904– 1993. Numerische Mathematik, 68(2):187, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068002/40680187.
htm;   http://science.springer.
de/nmee/bibs/4068002/40680187.
htm.

Charlier:1988:EIK

[CVV88] J. P. Charlier, M. Vanbegin, and P. Van Dooren. On efficient implementations of Kogbetliantz’s algorithm for computing the singular value decomposition. Numerische Mathematik, 52(3):279–300, March 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ciarlet:1971:MTF

[CW71] P. G. Ciarlet and C. Wagschal.

Multipoint Taylor formulas and applications to the finite element method. Numerische Mathematik, 17(1):84–100, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Conrad:1977:FSA

[CW77] V. Conrad and Y. Wallach. A faster SSOR algorithm. Numerische Mathematik, 27(4): 371–372, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cooper:1978:LFE

[CW78] G. J. Cooper and F. C. P. Whitworth. Liapunov functions and error bounds for approximate solutions of ordinary differential equations. Numerische Mathematik, 30(4):411–414, December 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Conrad:1979:AMS

[CW79a] V. Conrad and Y. Wallach. Alternating methods for sets of linear equations. Numerische Mathematik, 32(1): 105–108, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Craven:1979:SND

[CW79b] Peter Craven and Grace Wahba. Smoothing noisy data with spline functions. estimating the correct degree of smoothing by the method of generalized crossvalidation. Numerische Mathematik, 31(4):377–403, December 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1998:FEM

[CZ98] Zhiming Chen and Jun Zou. Finite element methods and their convergence for elliptic and parabolic interface problems. Numerische Mathematik, 79(2): 175–202, April 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079002/80790175.
htm;   http://science.springer.

de/nmee/bibs/8079002/80790175.
htm.

Ciarlet:1999:FDF

[CZ99] P. Ciarlet, Jr. and Jun Zou.

Fully discrete finite element approaches for time-dependent Maxwell’s equations. Numerische Mathematik, 82(2): 193–219, April 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer-
ny.com/link/service/journals/
00211/bibs/9082002/90820193.
htm;   http://link.springer-
ny.com/link/service/journals/
00211/papers/9082002/90820193.
pdf.

Dieter:1971:EDS

[DA71] U. Dieter and J. Ahrens. An exact determination of serial correlations of pseudo-random numbers. Numerische Mathematik, 17(2):101–123, April 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Davies:1986:IES

[DA86] A. R. Davies and R. S. Anderssen. Improved estimates of statistical regularization parameters in Fourier differentiation and smoothing. Numerische Mathematik, 48(6): 671–697, June 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dagnino:1990:PIS

[Dag90] Catterina Dagnino. Product integration of singular integrands based on cubic spline interpolation at equally spaced nodes. Numerische Mathematik, 57(1): 97–104, March 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dahlquist:1985:TGM

[Dah85] Germund Dahlquist. On transformations of graded matrices, with applications to stiff ODEs. Numerische Mathematik, 47(3): 363–385, November 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Daniel:1966:CAM

[Dan66] James W. Daniel. Correcting approximations to multiple roots of polynomials. Numerische Mathematik, 9(1):99– 102, November 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Daniel:1967:CCG

[Dan67] James W. Daniel. Convergence of the conjugate gradient method with computationally convenient modifications. Numerische Mathematik, 10(2): 125–131, July 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Daniel:1970:SGC

[Dan70] James W. Daniel. A simple, global, complementary variational principle. Numerische Mathematik, 16(1):52– 57, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Daniel:1973:NMN

[Dan73] James W. Daniel. Newton’s method for nonlinear inequalities. Numerische Mathematik, 21(5):381–387, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Davis:1962:NSP

[Dav62] Chandler Davis. The norm of the Schur product operation. Numerische Mathematik, 4:343– 344, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Davis:1970:TY

[Dav70] Chandler Davis. On a theorem of Yamamoto. Numerische Mathematik, 14(3): 297–298, February 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Delves:1977:OCW

[DB77] L. M. Delves and M. Bain. On the optimum choice of weight functions in a class of variational calculations. Numerische Mathematik, 27(2):209–218, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

deBoor:1977:BEA

[dBP77] Carl de Boor and Allan Pinkus.

A backward error analysis for totally positive linear systems.

Numerische Mathematik, 27(4): 485–490, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demoulin:1974:IMS

[DC74] Yves-Marie J. Demoulin and Y. M. Chen. An iteration method for studying the bifurcation of solutions of the nonlinear equations, L(λ)u + ϵR(λ, u) = 0. Numerische Mathematik, 23 (1):47–61, February 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DiBenedetto:1999:UAA

[DC99] Fabio Di Benedetto and Stefano Serra Capizzano. A unifying approach to abstract matrix algebra preconditioning. Numerische Mathematik, 82(1):57–90, March 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082001/90820057.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082001/90820057.pdf.

Douglas:1973:GMP

[DD73a] Jim Douglas, Jr. and Todd Dupont. Galerkin methods for parabolic equations with nonlinear boundary conditions. Numerische Mathematik, 20(3): 213–237, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Douglas:1973:SGM

[DD73b] Jim Douglas, Jr. and Todd Dupont. Superconvergence for Galerkin methods for the two point boundary problem via local projections. Numerische Mathematik, 21(3): 270–278, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Douglas:1974:GAT

[DD74] Jim Douglas, Jr. and Todd Dupont. Galerkin approximations for the two point boundary problem using continuous, piecewise polynomial spaces. Numerische Mathematik, 22(2):99–109, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Davies:1975:GCH

[DD75] M. Davies and B. Dawson. On the global convergence of Halley’s iteration formula. Numerische Mathematik, 24(2): 133–135, April 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Davies:1978:ASP

[DD78] M. Davies and B. Dawson.

An automatic search procedure for finding real zeros. Numerische Mathematik, 31(3): 299–312, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DellaDora:1984:APHb

[DD84a] J. Della Dora and C. Di Crescenzo. Approximants de Padé–Hermite. II. programmation. (French) [Padé–Hermite approximants. II. Programming]. Numerische Mathematik, 43(1):41–57, January 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DellaDora:1984:APHa

[DD84b] J. Della Dora and C. Di Crescenzo. Approximants de Padé-hermite. I. théorie. (French) [Padé–Hermite approximants. I. theory]. Numerische Mathematik, 43(1):23–39, January 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deckelnick:1995:CFE

[DD95] Klaus Deckelnick and Gerhard Dziuk. Convergence of a finite element method for non-parametric mean curvature flow. Numerische Mathematik, 72(2):197–222, December 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5072002/50720197.
htm;   http://science.springer.
de/nmee/bibs/5072002/50720197.
htm.

Delfour:1998:FMA

[DD98] Michel Delfour and François Dubeau. Fixed mesh approximation of ordinary differential equations with impulses. Numerische Mathematik, 78(3): 377–401, January 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7078003/
70780377.htm;   http://science.
springer.de/nmee/bibs/7078003/
70780377.htm.

Damelin:1999:IPQ

[DD99] S. B. Damelin and Kai Diethelm. Interpolatory product quadratures for Cauchy principal value integrals with Freud weights. Numerische Mathematik, 83(1):87–105, July 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083001/90830087.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083001/90830087.pdf.

Dahmen:1992:MEB

[DDM92] Wolfgang Dahmen, Ronald A.

DeVore, and Charles A. Micchelli. On monotone extensions of boundary data. Numerische Mathematik, 60(4): 477–492, January 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Douglas:1975:SPF

[DDW75] Jim Douglas, Jr., Todd Dupont, and Lars Wahlbin. The stability in L^{q} of the L^{2}-projection into finite element function spaces. Numerische Mathematik, 23(3): 193–197, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DeMeersman:1968:MG

[De 68] R. De Meersman. The method of Garabedian. Numerische Mathematik, 11(3):257–263, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

deLaBourdonnaye:1995:SFC

[de 95] Armel de La Bourdonnaye. Some formulations coupling finite element and integral equation methods for Helmholtz equation and electromagnetism. Numerische Mathematik, 69(3): 257–268, January 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/5069003/
                                        50690257.htm;   http://science.
                                        springer.de/nmee/bibs/5069003/
                                        50690257.htm.

Dieci:1994:PDN

[DE94] Luca Dieci and Timo Eirola. Positive definiteness in the numerical solution of Riccati differential equations. Numerische Mathematik, 67(3): 303–313, April 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067003/40670303.
htm;   http://science.springer.
de/nmee/bibs/4067003/40670303.
htm.

Dieci:1996:PMN

[DE96] Luca Dieci and Timo Eirola.

Preserving monotonicity in the numerical solution of Riccati differential equations. Numerische Mathematik, 74(1): 35–47, July 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074001/60740035.
htm;   http://science.springer.
de/nmee/bibs/6074001/60740035.
htm.

Deak:1980:TDA

[Deá80] I. Deák. Three digit accurate multiple normal probabilities. Numerische Mathematik, 35(4):369–380, November 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deif:1989:RBP

[Dei89]  A. Deif.  A relative back-

ward perturbation theorem for the eigenvalue problem. Numerische Mathematik, 56(6): 625–626, December 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dejon:1966:AST

[Dej66a]  B. Dejon.  Addendum to

“stronger than uniform convergence of multistep difference methods”. Numerische Mathematik, 9(3):268–270, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Dej66b].

Dejon:1966:STU

[Dej66b] B. Dejon. Stronger than uniform convergence of multistep difference methods. Numerische Mathematik, 8(1):29–41, March 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See addendum [Dej66a].

Dekker:1971:FPT

[Dek71]  Theodorus J. Dekker.  A

floating-point technique for extending the available precision. Numerische Mathematik, 18(3): 224–242, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://www-gdz.sub.
                                        uni-goettingen.de/cgi-bin/
                                        digbib.cgi?PPN362160546_0018.

Delchambre:1973:AMD

[Del73] M. Delchambre. De l’avantage des méthodes aux différences modifiées de Stiefel–Bettis pour la résolution d’équations différentielles du second ordre perturbées. (French) [The advantage of modified Stiefel–Bettis difference methods for the solution of perturbed second order differential equations]. Numerische Mathematik, 21(1):33– 36, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Delahaye:1980:APS

[Del80]  Jean-Paul Delahaye.  Algo-

rithmes pour suites non convergentes. (French) [Algorithms for nonconvergent sequences]. Numerische Mathematik, 34(3): 333–347, September 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demmel:1987:CND

[Dem87] James Weldon Demmel. On condition numbers and the distance to the nearest ill-posed problem. Numerische Mathematik, 51(3):251–289, July 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dennis:1968:NLM

[Den68] J. E. Dennis, Jr. On Newton-like methods. Numerische Mathematik, 11(4):324–330, May 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Descloux:1961:DAT

[Des61] J. Descloux. Dégénérescence dans les approximations de Tschebyscheff linéaires et discrètes. (French) [Degeneration in linear and discrete Chebyshev approximations]. Numerische Mathematik, 3:180–187, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Descloux:1963:BSN

[Des63] Jean Descloux. Bounds for the spectral norm of functions of matrices. Numerische Mathematik, 5(1):185–190, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Descloux:1970:NIH

[Des70] J. Descloux. On the numerical integration of the heat equation. Numerische Mathematik, 15(5): 371–381, September 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deutsch:1970:MN

[Deu70]  Emeric Deutsch.  Matricial

norms. Numerische Mathematik, 16(1):73–84, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deuflhard:1974:MNM

[Deu74]  P. Deuflhard.  A modified

Newton method for the solution of ill-conditioned systems of nonlinear equations with application to multiple shooting. Numerische Mathematik, 22(4): 289–315, August 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deutsch:1975:MNL

[Deu75] Emeric Deutsch. On matrix norms and logarithmic norms. Numerische Mathematik, 24(1): 49–51, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deuflhard:1979:SCC

[Deu79] P. Deuflhard. A stepsize control for continuation methods and its special application to multiple shooting techniques. Numerische Mathematik, 33(2): 115–146, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deuflhard:1983:OSC

[Deu83] P. Deuflhard. Order and stepsize control in extrapolation methods. Numerische Mathematik, 41(3):399–422, July 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

deFrutos:1997:ACP

[dFSS97] J. de Frutos and J. M. Sanz-Serna. Accuracy and conservation properties in numerical integration: The case of the Korteweg–de Vries equation. Numerische Mathematik, 75(4): 421–445, February 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7075004/
70750421.htm;   http://science.
springer.de/nmee/bibs/7075004/
70750421.htm.

Douglas:1986:SMF

[DG86] Jim Douglas, Jr. and Chaitan P. Gupta. Superconvergence for a mixed finite element method for elastic wave propagation in a plane domain. Numerische Mathematik, 49(2/3): 189–202, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demkowicz:1998:CIE

[DG98] L. Demkowicz and K. Gerdes. Convergence of the infinite element methods for the Helmholtz equation in separable domains. Numerische Mathematik, 79(1): 11–42, March 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/8079001/80790011.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/8079001/80790011.
                                        htm.

Delahaye:1980:RNA

[DGB80] J. P. Delahaye and B. Germain-Bonne. Résultats négatifs en accélération de la convergence. (French) [Negative results in convergence acceleration]. Numerische Mathematik, 35(4): 443–457, November 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Du:1999:ACM

[DGL99] Qiang Du, Max D. Gunzburger, and Hyesuk Kwon Lee. Analysis and computation of a meanfield model for superconductivity. Numerische Mathematik, 81(4):539–560, February 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9081004/90810539.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9081004/90810539.pdf.

Dahmen:1988:CSF

[DGM88] W. Dahmen, T. N. T. Goodman, and Charles A. Micchelli. Compactly supported fundamental functions for spline interpolation. Numerische Mathematik, 52(6):639–664, June 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Du:1993:FEA

[DGP93] Qiang Du, Max Gunzburger, and Janet Peterson. Finite element approximation of a periodic Ginzburg–Landau model for type-II superconductors. Numerische Mathematik, 64(1):85– 114, January 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dittrich:1980:MCS

[DH80] Susanne Dittrich and Wolfgang Hackbusch. A method of characteristics solving the initialboundary value problem of a hyperbolic differential equation of second order. Numerische Mathematik, 34(3):217–234, September 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dekker:1989:RGJ

[DH89] T. J. Dekker and W. Hoffmann. Rehabilitation of the Gauss–Jordan algorithm. Numerische Mathematik, 54(5): 591–599, February 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

dHennezel:1993:DDM

[d’H93a] Frédéric d’Hennezel. Domain decomposition method and elastic multi-structures: the stiffened plate problem. Numerische Mathematik, 66(2): 181–197, November 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Drmac:1993:QCB

[DH93b] Zlatko Drmač and Vjeran Hari. On quadratic convergence bounds for the J-symmetric Jacobi method. Numerische Mathematik, 64(2):147–180, February 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dellnitz:1997:SAC

[DH97] Michael Dellnitz and Andreas Hohmann. A subdivision algorithm for the computation of unstable manifolds and global attractors. Numerische Mathematik, 75(3):293–317, January 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7075003/70750293.
htm;   http://science.springer.
de/nmee/bibs/7075003/70750293.
htm.

deHoog:1987:EMC

[dHH87] F. R. de Hoog and M. F.

Hutchinson. An efficient method for calculating smoothing splines using orthogonal transformations. Numerische Mathematik, 50(3):311–319, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DeLillo:1999:NCM

[DHP99a] Thomas K. DeLillo, Mark A.

Horn, and John A. Pfaltzgraff. Numerical conformal mapping of multiply connected regions by Fornberg-like methods. Numerische Mathematik, 83(2):205–230, August 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083002/90830205.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083002/90830205.pdf.

Douglas:1999:MMC

[DHP99b] Jim Douglas, Jr., Chieh-Sen Huang, and Felipe Pereira. The modified method of characteristics with adjusted advection. Numerische Mathematik, 83(3):353–369, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083003/90830353.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083003/90830353.pdf.

deHoog:1973:SVI

[dHW73] Frank de Hoog and Richard Weiss. On the solution of Volterra integral equations of the first kind. Numerische Mathematik, 21(1):22– 32, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

deHoog:1975:IRM

[dHW75] F. de Hoog and R. Weiss. Implicit Runge–Kutta methods for second kind Volterra integral equations. Numerische Mathematik, 23(3):199–213, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deuflhard:1987:OSE

[DHZ87] P. Deuflhard, E. Hairer, and J. Zugck. One-step and extrapolation methods for differentialalgebraic systems. Numerische Mathematik, 51(5):501–516, October 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Diaz:1980:CDP

[Dia80] F. Utreras Diaz. Sur le choix du paramètre d’ajustement dans le lissage par fonctions spline. (French) [On the choice of the adjustment parameter in smoothing by spline functions]. Numerische Mathematik, 34(1): 15–28, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dickmeis:1987:CTC

[Dic87] W. Dickmeis. Comparison theorems for compound quadrature formulas. Numerische Mathematik, 50(5):547–556, March 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dietel:1987:FIM

[Die87] Carmen Dietel. A fast iterative method for solving stationary heat conduction problems on regions partitioned into rectangles. Numerische Mathematik, 51(4):415–428, August 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Diethelm:1996:PKB

[Die96] Kai Diethelm. Peano kernels and bounds for the error constants of Gaussian and related quadrature rules for Cauchy principal value integrals. Numerische Mathematik, 73(1): 53–63, March 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073001/60730053.
htm;   http://science.springer.
de/nmee/bibs/6073001/60730053.
htm.

Dijkstra:1959:NTP

[Dij59] E. W. Dijkstra. A note on two problems in connexion with graphs. Numerische Mathematik, 1(1):269–271, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
269.

Dijkstra:1960:RP

[Dij60] E. W. Dijkstra. Recursive programming. Numerische Mathematik, 2(1):312–318, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dingyi:1983:ABP

[Din83] Mu Dingyi. Asymptotic behavior of the polynomial factor in the interpolation error. Numerische Mathematik, 42(2): 247–257, November 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dirschmid:1969:BAG

[Dir69] H. J. Dirschmid. Bemerkungen zu einer Arbeit von G. Polya zur Bestimmung der Nullstellen ganzer Funktionen. (German) [Comment on a work by G. Polya for determining zeros of real functions]. Numerische Mathematik, 13(4):344–348, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dixon:1967:HSC

[Dix67] John D. Dixon. High speed computation of group characters. Numerische Mathematik, 10(5):446–450, November 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dixon:1982:ESL

[Dix82] John D. Dixon. Exact solution of linear equations using P -adic expansions. Numerische Mathematik, 40(1):137–141, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Douglas:1962:NMI

[DJ62] J. Douglas, Jr. and B. F. Jones, Jr. Numerical methods for integrodifferential equations of parabolic and hyperbolic types. Numerische Mathematik, 4:96– 102, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Davis:1969:NIL

[DJ69] Joel Davis and Jerome J. Jacoby. Numerical integration of linear integral equations with mild discontinuity. Numerische Mathematik, 13(4):357–361, August 1969. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

deJong:1977:TFD

[dJ77] Lieuwe Sytse de Jong. Towards a formal definition of numerical stability. Numerische Mathematik, 28(2):211–219, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dax:1977:PTS

[DK77] Achiya Dax and S. Kaniel. Pivoting techniques for symmetric Gaussian elimination. Numerische Mathematik, 28(2): 221–241, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Decker:1983:SCC

[DK83] D. W. Decker and C. T. Kelley. Sublinear convergence of the chord method at singular points. Numerische Mathematik, 42(2): 147–154, November 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dahmen:1992:MP

[DK92] Wolfgang Dahmen and Angela Kunoth. Multilevel preconditioning. Numerische Mathematik, 63(3):315–344, December 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dekker:1990:ERB

[DKS90a] K. Dekker, J. F. B. M. Kraaijevanger, and J. Schneid. Erratum: “On the relation between algebraic stability and B-convergence for Runge–Kutta methods”. Numerische Mathematik, 58(2):230, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [DKS90b].

Dekker:1990:RBA

[DKS90b] K. Dekker, J. F. B. M. Kraaijevanger, and J. Schneid. On the relation between algebraic stability and B-convergence for Runge–Kutta methods. Numerische Mathematik, 57(3): 249–262, May 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [DKS90a].

Dai:1997:NMF

[DL97] Hua Dai and Peter Lancaster.

Numerical methods for finding multiple eigenvalues of matrices depending on parameters. Numerische Mathematik, 76(2): 189–208, April 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076002/70760189.
htm;   http://science.springer.
de/nmee/bibs/7076002/70760189.
htm.

Degond:1994:ESF

[DLD94] Pierre Degond and Brigitte Lucquin-Desreux. An entropy scheme for the Fokker– Planck collision operator of plasma kinetic theory. Numerische Mathematik, 68(2):

239–262, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/4068002/40680239.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4068002/40680239.
                                        htm.

Diekhoff:1977:CRN

[DLO^{+}77] H.-J. Diekhoff, P. Lory, H. J. Oberle, H.-J. Pesch, P. Rentrop, and R. Seydel. Comparing routines for the numerical solution of initial value problems of ordinary differential equations in multiple shooting. Numerische Mathematik, 27(4): 449–469, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dodd:1985:ACS

[DM85] S. L. Dodd and D. F. McAllister. Algorithms for computing shape preserving spline approximations to data. Numerische Mathematik, 46(2): 159–174, May 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dyn:1988:PPS

[DM88] Nira Dyn and Charles A. Micchelli. Piecewise polynomial spaces and geometric continuity of curves. Numerische Mathematik, 54(3):319–337, December 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dyn:1990:ISR

[DM90] Nira Dyn and Charles A. Micchelli. Interpolation by sums of radial functions. Numerische Mathematik, 58(1):1–9, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dubois:1996:NPE

[DM96] François Dubois and Guillaume Mehlman. A non-parameterized entropy correction for Roe’s approximate Riemann solver. Numerische Mathematik, 73(2): 169–208, April 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073002/60730169.
htm;   http://science.springer.
de/nmee/bibs/6073002/60730169.
htm.

Diaz-Mendoza:1996:CTP

[DMGVO96] Carlos Díaz-Mendoza, Pablo González-Vera, and Ramón Orive. On the convergence of two-point Padé-type approximants. Numerische Mathematik, 72(3):295–312, January 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6072003/60720295.
htm;   http://science.springer.
de/nmee/bibs/6072003/60720295.
htm.

Dixon:1986:CAD

[DMJ86] Jennifer Dixon, Sean McKee, and Rolf Jeltsch. Convergence analysis of discretization methods for nonlinear first kind Volterra integral equations. Numerische Mathematik, 49(1):67– 80, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dahlquist:1983:UPB

[DML83] Germund Dahlquist, Huang Mingyou, and Randall LeVeque. On the uniform powerboundedness of a family of matrices and the applications to one-leg and linear multistep methods. Numerische Mathematik, 42(1):1–13, September 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Duran:1991:AEE

[DMR91] Ricardo Durán, María Amelia Muschietti, and Rodolfo Rodríguez. On the asymptotic exactness of error estimators for linear triangular finite elements. Numerische Mathematik, 59(2): 107–127, May 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Destuynder:1986:AND

[DN86] P. Destuynder and J.-C. Nédélec. Approximation numérique du cisaillement transverse dans les plaques minces en flexion. (French) [Numerical approximation of transverse shear stress in bent plates]. Numerische Mathematik, 48(3):281–302, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Degond:1989:NAW

[DN89] P. Degond and B. Niclot. Numerical analysis of the weighted particle method applied to the semiconductor Boltzmann equation. Numerische Mathematik, 55(5):599–618, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dickmeis:1984:GAC

[DNvW84] W. Dickmeis, R. J. Nessel, and E. van Wickeren. A general approach to counterexamples in numerical analysis. Numerische Mathematik, 43(2): 249–263, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

deOliveira:1970:NIC

[dO70] G. N. de Oliveira. Note on an inverse characteristic value problem. Numerische Mathematik, 15(4):345–347, August 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dobrowolski:1981:MFE

[Dob81] Manfred Dobrowolski. A mixed finite element method for solving the nonstationary Stokes equation. Numerische Mathematik, 36(3):225–236, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dobrowolski:1989:DSF

[Dob89] Manfred Dobrowolski. A discrete solenoidal finite difference scheme for the numerical approximation of incompressible flows. Numerische Mathematik, 54(5):533–542, February 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dohler:1982:NGS

[Döh82] B. Döhler. Ein neues Gradientenverfahren zur simultanen Berechnung der kleinsten oder größten Eigenwerte des allgemeinen Eigenwertproblems. (German) [A new gradient method for the simultaneous calculation of the smallest or largest eigenvalues of the general eigenvalue problem]. Numerische Mathematik, 40(1):79– 91, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dollinger:1981:LKZ

[Dol81] E. Dollinger. Ein linear konvergentes zyklisches jacobiähnliches Verfahren für beliebige reelle Matrizen. (German) [A linear convergent cyclic Jacobi-like method for arbitrary real matrices]. Numerische Mathematik, 38(2):245–253, December 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Doring:1966:IIS

[Dör66]  Boro Döring.  Inverse In-

terpolation bei den speziellen Funktionen der mathematischen Physik. (German) [Inverse interpolation in the special functions of mathematical physics]. Numerische Mathematik, 8(2): 123–136, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Doring:1970:TVB

[Dör70] Boro Döring. Das Tschebyscheff-Verfahren in Banach-Räumen. (German) [Chebyshev’s method in Banach spaces]. Numerische Mathematik, 15(3): 175–195, July 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dorfler:1990:CSM

[Dör90] W. Dörfler. The conditioning of the stiffness matrix for certain elements approximating the incompressibility condition in fluid dynamics. Numerische Mathematik, 58(2): 203–214, September 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dorfler:1996:TSA

[Dör96]  Willy Dörfler.  A time-

and space-adaptive algorithm for the linear time-dependent Schrödinger equation. Numerische Mathematik, 73(4): 419–448, June 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/6073004/60730419.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/6073004/60730419.
                                        htm.

Douglas:1960:NMS

[Dou60]  Jim Douglas.  A numeri-

cal method for the solution of a parabolic system. Numerische Mathematik, 2(1):91– 98, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Douglas:1961:ADI

[Dou61] J. Douglas, Jr. Alternating direction iteration for mildly nonlinear elliptic difference equations. Numerische Mathematik, 3:92–98, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See correction [Dou62b].

Douglas:1962:ADM

[Dou62a] Jim Douglas, Jr. Alternating direction methods for three space variables. Numerische Mathematik, 4:41–63, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Douglas:1962:CMP

[Dou62b] Jim Douglas, Jr. A correction to my paper “Alternating direction iteration for mildly nonlinear elliptic difference equations”. Numerische Mathematik, 4:301–302, December 1962. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Dou61].

Douglas:1963:CAD

[DP63] J. Douglas, Jr. and C. M.

Pearcy. On convergence of alternating direction procedures in the presence of singular operators. Numerische Mathematik, 5(1):175–184, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dailey:1972:EBG

[DP72] J. W. Dailey and J. G. Pierce.

Error bounds for the Galerkin method applied to singular and nonsingular boundary value problems. Numerische Mathematik, 19(3):266–282, June 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dikshit:1982:DCS

[DP82] H. P. Dikshit and P. Powar. Discrete cubic spline interpolation. Numerische Mathematik, 40(1): 71–78, November 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dagnino:1985:ECT

[DP85] C. Dagnino and A. Palamara Orsi. On the evaluation of certain two-dimensional singular integrals with Cauchy kernels. Numerische Mathematik, 46(1):121–130, March 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dagnino:1988:PIP

[DP88a] C. Dagnino and A. Palamara Orsi. Product integration of piecewise continuous integrands based on cubic spline interpolation at equally spaced nodes. Numerische Mathematik, 52(4): 459–466, April 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

DeMatteis:1988:PRN

[DP88b] A. De Matteis and S. Pagnutti. Parallelization of random number generators and longrange correlations. Numerische Mathematik, 53(5):595–608, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deuflhard:1976:MCM

[DPR76] P. Deuflhard, H.-J. Pesch, and P. Rentrop. A modified continuation method for the numerical solution of nonlinear two-point boundary value problems by shooting techniques. Numerische Mathematik, 26(3): 327–343, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Douglas:1993:PIP

[DPRW93] J. Douglas, Jr., P. J. Paes Leme, J. E. Roberts, and Junping Wang. A parallel iterative procedure applicable to the approximate solution of second order partial differential equations by mixed finite element methods. Numerische Mathematik, 65(1):95–108, May 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Demmler:1975:OMS

[DR75] A. Demmler and C. Reinsch.

Oscillation matrices with spline smoothing. Numerische Mathematik, 24(5):375–382, October 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Duran:1992:AEB

[DR92] Ricardo Durán and Rodolfo Rodríguez. On the asymptotic exactness of Bank-Weiser’s estimator. Numerische Mathematik, 62(3):297–303, August 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Drexler:1977:MBS

[Dre77] Franz-Josef Drexler. Eine Methode zur Berechnung sämtlicher Lösungen von Polynomgleichungssystemen. (German) [A method for computing all solutions of systems of polynomial equations]. Numerische Mathematik, 29(1):45–58, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Duchon:1981:PDI

[DRW81] Jean Duchon, Raoul Robert, and Patrick Witomski. Problème de Dirichlet dans l’image bilipschitzienne d’un demi-espace. (French) [The Dirichlet problem in the bi-Lipschitz image of a half-space]. Numerische Mathematik, 36(2):129–149, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dryja:1982:CMM

[Dry82] Maksymilian Dryja. A capacitance matrix method for Dirichlet problem on polygon region. Numerische Mathematik, 39(1): 51–64, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dryja:1984:FEC

[Dry84] Maksymilian Dryja. A finite element-capacitance method for elliptic problems on regions partitioned into subregions. Numerische Mathematik, 44(2): 153–168, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dryanov:1994:QFF

[Dry94] Dimiter P. Dryanov. Quadrature formulae with free nodes for periodic functions. Numerische Mathematik, 67(4): 441–464, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/

00211/bibs/4067004/40670441.
htm;   http://science.springer.
de/nmee/bibs/4067004/40670441.
htm.

DAzevedo:1991:OTM

[DS91] E. F. D’Azevedo and R. B.

Simpson. On optimal triangular meshes for minimizing the gradient error. Numerische Mathematik, 59(4):321–348, July 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dryja:1996:MSM

[DSW96] Maksymilian Dryja, Marcus V.

Sarkis, and Olof B. Widlund. Multilevel Schwarz methods for elliptic problems with discontinuous coefficients in three dimensions. Numerische Mathematik, 72(3):313–348, January 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6072003/60720313.
htm;   http://science.springer.
de/nmee/bibs/6072003/60720313.
htm.

Dekker:1971:ASA

[DT71] T. J. Dekker and J. F. Traub.

An analysis of the shifted LR algorithm. Numerische Mathematik, 17(3):179–188, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dubrulle:1970:SNI

[Dub70] A. Dubrulle. A short note on the implicit QL algorithm for symmetric tridiagonal matrices. Numerische Mathematik, 15(5): 450, September 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dueck:1959:FEL

[Due59] W. Dueck. Eine Fehlerabschätzung zum Einzelschrittverfahren bei linearen Gleichungssystemen. (German) [An error bound for the independent step method in linear equation systems]. Numerische Mathematik, 1:73–77, December 1959. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dumitrescu:1998:IEA

[Dum98] Bogdan Dumitrescu. Improving and estimating the accuracy of Strassen’s algorithm. Numerische Mathematik, 79(4): 485–499, June 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/8079004/80790485.
                                        htm;   http://link.springer.
                                        de/link/service/journals/
                                        00211/papers/8079004/80790485.
                                        pdf;   http://link.springer.
                                        de/link/service/journals/
                                        00211/papers/8079004/80790485.
                                        ps.gz.

Dunham:1967:TRC

[Dun67] Charles B. Dunham. Transformed rational Chebyshev approximation. Numerische Mathematik, 10(2):147–152, July 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dunham:1968:TRC

[Dun68] Charles B. Dunham. Transformed rational Chebyshev approximation. II. Numerische Mathematik, 12(1):8–10, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dunn:1988:PNM

[Dun88] J. C. Dunn. A projected Newton method for minimization problems with nonlinear inequality constraints. Numerische Mathematik, 53(4):377–409, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Duran:1990:SRM

[Dur90] Ricardo Durán. Superconvergence for rectangular mixed finite elements. Numerische Mathematik, 58(3):287–298, October 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Duran:1991:ISC

[Dur91] Ricardo Durán. The inf-sup condition and error estimates for the Arnold-Falk plate bending element. Numerische Mathematik, 59(8):769–778, October 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dutt:1999:SMI

[Dut99] P. Dutt. Spectral methods for initial boundary value problems with nonsmooth data. Numerische Mathematik, 81(3): 323–344, January 1999. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/9081003/
90810323.htm;   http://link.
springer.de/link/service/
journals/00211/papers/9081003/
90810323.pdf.

deVries:1973:ENM

[dV73] H. J. de Vries. On the eigenproblem for normal matrices. Numerische Mathematik, 21(1):37– 42, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

deVilliers:1993:NSI

[dV93] J. M. de Villiers. A nodal spline interpolant for the Gregory rule of even order. Numerische Mathematik, 66(1):123–137, October 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dryanov:1997:QFM

[DV97] D. Dryanov and V. Vatchev. On quadrature formulae with maximal trigonometric degree of precision. Numerische Mathematik, 77(1):59–82, July 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077001/70770059.
htm;   http://science.springer.
de/nmee/bibs/7077001/70770059.
htm.

Dieci:1999:COF

[DV99] Luca Dieci and Erik S. Van Vleck. Computation of orthonormal factors for fundamental solution matrices. Numerische Mathematik, 83 (4):599–620, October 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083004/90830599.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083004/90830599.pdf.

Dassie:1998:ASP

[DVZ98] Silvia Dassiè, Marco Vianello, and Renato Zanovello. Asymptotic summation of power series. Numerische Mathematik, 80(1):61–73, July 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080001/80800061.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800061.
pdf;   http://link.springer.
de/link/service/journals/

                                        00211/papers/8080001/80800061.
                                        ps.gz.

Deutsch:1971:IDE

[DZ71] Eckart Deutsch and Christoph Zenger. Inclusion domains for the eigenvalues of stochastic matrices. Numerische Mathematik, 18(2):182–192, April 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Deutsch:1975:BGG

[DZ75] E. Deutsch and C. Zenger. On Bauer’s generalized Gershgorin discs. Numerische Mathematik, 24(1):63–70, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dziuk:1991:AES

[Dzi91] G. Dziuk. An algorithm for evolutionary surfaces. Numerische Mathematik, 58(6): 603–611, January 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elfving:1988:ACC

[EA88] Tommy Elfving and Lars-Erik Andersson. An algorithm for computing constrained smoothing spline functions. Numerische Mathematik, 52(5): 583–595, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eberlein:1968:SEN

[EB68] P. J. Eberlein and J. Boothroyd.

Solution to the eigenproblem by a norm reducing Jacobi type method. Numerische Mathematik, 11(1):1–12, January 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 327–338].

Eberlein:1970:SCE

[Ebe70] P. J. Eberlein. Solution to the complex eigenproblem by a norm reducing Jacobi type method. Numerische Mathematik, 14(3):232–245, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 404–417].

Eckhardt:1968:EEW

[Eck68] Ulrich Eckhardt. Einige Eigenschaften Wilfscher Quadraturformeln. (German) [Some eigenvalues of the Wilf quadrature formula]. Numerische Mathematik, 12(1):1–7, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eckmann:1976:CGG

[Eck76] J.-P. Eckmann. On the computation of the Galois group over the quotient field of C(λ). Numerische Mathematik, 26(1): 27–37, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Edder:1993:ASM

[Edd93] Olivier Edder. Analysis of a simplified model of drop combustion. Numerische Mathematik, 65(3):273–284, July 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elliott:1989:SOS

[EFM89] C. M. Elliott, D. A. French, and F. A. Milner. A second order splitting method for the Cahn–Hilliard equation. Numerische Mathematik, 54(5): 575–590, February 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eisinberg:1998:VMI

[EFP98] Alfredo Eisinberg, Giuseppe Franzé, and Paolo Pugliese. Vandermonde matrices on integer nodes. Numerische Mathematik, 80(1):75–85, July 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8080001/
80800075.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8080001/
80800075.pdf;   http://link.
springer.de/link/service/
journals/00211/papers/8080001/
80800075.ps.gz.

Engl:1994:UCD

[EG94] Heinz W. Engl and Wilhelm Grever. Using the l-curve for determining optimal regularization parameters. Numerische Mathematik, 69(1):25– 31, November 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069001/40690025.
htm;   http://science.springer.
de/nmee/bibs/4069001/40690025.
htm.

Elschner:1995:OOC

[EG95] J. Elschner and I. G. Graham.

An optimal order collocation method for first kind boundary integral equations on polygons. Numerische Mathematik, 70(1): 1–31, March 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070001/50700001.
htm;   http://science.springer.
de/nmee/bibs/5070001/50700001.
htm.

Eggermont:1984:SRC

[Egg84] P. P. B. Eggermont. Stability and robustness of collocation methods for Abel-type integral equations. Numerische Mathematik, 45(3):431–445, December 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eggermont:1986:IAC

[Egg86] P. P. B. Eggermont. Improving the accuracy of collocation solutions of Volterra integral equations of the first kind by local interpolation. Numerische Mathematik, 48(3):263–279, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eggermont:1988:MAV

[Egg88]  P. P. B. Eggermont.  On

monotone Abel–Volterra integral equations on the half line. Numerische Mathematik, 52(1): 65–79, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eggermont:1992:QEO

[Egg92] P. P. B. Eggermont. On the quadrature error in operational quadrature methods for convolutions. Numerische Mathematik, 62(1):35–48, June 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eymard:1999:CFV

[EGH99] Robert Eymard, Thierry Gallouët, and Raphaèle Herbin. Convergence of finite volume schemes for semilinear convection diffusion equations. Numerische Mathematik, 82(1):91–116, March 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9082001/90820091.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9082001/90820091.pdf.

Elman:1994:CLI

[EGS94] Howard C. Elman, Gene H.

Golub, and Gerhard Starke. On the convergence of line iterative methods for cyclically reduced non-symmetrizable linear systems. Numerische Mathematik, 67(2):177–190, March 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4067002/
40670177.htm;   http://science.
springer.de/nmee/bibs/4067002/
40670177.htm.

Ehlich:1970:KMI

[EH70] H. Ehlich and W. Haussmann.

Konvergenz mehrdimensionaler Interpolation. (German) [Convergence of multidimensional interpolation]. Numerische Mathematik, 15(2):165–174, June 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eichenauer-Herrmann:1989:RLR

[EHG89] Jürgen Eichenauer-Herrmann and Holger Grothe. A remark on long-range correlations in multiplicative congruential pseudo random number generators. Numerische Mathematik, 56(6): 609–611, December 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elsner:1994:MCN

[EHM94] L. Elsner, C. He, and V. Mehrmann.

Minimizing the condition number of a positive definite matrix by completion. Numerische Mathematik, 69(1):17– 23, November 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069001/40690017.
htm;   http://science.springer.
de/nmee/bibs/4069001/40690017.
htm.

Ehresmann:1969:ABA

[Ehr69] Dietbert Ehresmann. Zur Abschätzung des Bestapproximationsfehlers bei der Approximation differenzierbarer Funktionen durch Polynome. (German) [Toward approximation of best approximation errors for the approximation of differentiable functions by polynomials]. Numerische Mathematik, 13(1): 94–100, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eiermann:1989:SMG

[Eie89] Michael Eiermann. On semiiterative methods generated by Faber polynomials. Numerische Mathematik, 56(2/3): 139–156, October 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Einarsson:1971:NTI

[Ein71]  Bo Einarsson.  Numerical

treatment of integro-differential equations with a certain maximum property. Numerische Mathematik, 18(3):267–288, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eisen:1967:ESC

[Eis67a]  Dennis Eisen.  The equiv-

alence of stability and convergence for finite difference schemes with singular coefficients. Numerische Mathematik, 10(1):20–29, May 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eisen:1967:NS

[Eis67b] Dennis Eisen. On the numerical solution of u_{t} = u_{rr} + \frac{2}{r}u_{r}. Numerische Mathematik, 10(5): 397–409, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Evequoz:1981:NFE

[EJ81] H. Evéquoz and Y. Jaccard.

A nonconforming finite element method to compute the spectrum of an operator relative to the stability of a plasma in toroidal geometry. Numerische Mathematik, 36(4): 455–465, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elschner:1997:CMM

[EJSS97] J. Elschner, Y. Jeon, I. H. Sloan, and E. P. Stephan. The collocation method for mixed boundary value problems on domains with curved polygonal boundaries.

Numerische Mathematik, 76(3): 355–381, May 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076003/70760355.
htm;   http://science.springer.
de/nmee/bibs/7076003/70760355.
htm.

Evje:1999:VSA

[EK99] Steinar Evje and Kenneth Hvistendahl Karlsen. Viscous splitting approximation of mixed hyperbolic-parabolic convectiondiffusion equations. Numerische Mathematik, 83(1):107–137, July 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083001/90830107.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083001/90830107.pdf.

Elsner:1991:CPA

[EKL91] L. Elsner, I. Koltracht, and P. Lancaster. Convergence properties of ART and SOR algorithms. Numerische Mathematik, 59(1):91–106, April 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elsner:1992:CSA

[EKN92] L. Elsner, I. Koltracht, and M. Neumann. Convergence of sequential and asynchronous nonlinear paracontractions. Numerische Mathematik, 62(3):

305–319, August 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

ElTom:1979:BCF

[El 79] M. E. A. El Tom. On best cubature formulas and spline interpolation. Numerische Mathematik, 32(3):291–306, September 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

ElTarazi:1982:SCR

[El 82] Mouhamed Nabih El Tarazi.

Some convergence results for asynchronous algorithms. (French) [Des résultats de convergence des algorithmes asynchrones]. Numerische Mathematik, 39(3): 325–340, October 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Curiously, the title of this paper is in English, while the body is in French!

ElTarazi:1984:AMA

[El 84] Mouhamed Nabih El Tarazi. Algorithmes mixtes asynchrones. étude de convergence monotone. (French) [Asynchronous mixed algorithms. study of monotonic convergence]. Numerische Mathematik, 44(3): 363–369, September 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Evans:1988:EGS

[EL88] D. J. Evans and C. Li. Extrapolated Gauss–Seidel 1 and SOR methods for least-squares problems. Numerische Mathematik, 53(4):485–498, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elbert:1991:AZB

[Elb91] Á. Elbert. An approximation for the zeros of Bessel functions. Numerische Mathematik, 59(7): 647–657, October 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elfving:1980:BIM

[Elf80] Tommy Elfving. Block-iterative methods for consistent and inconsistent linear equations. Numerische Mathematik, 35(1):1– 12, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ellacott:1979:ACM

[Ell79] S. W. Ellacott. On the approximate conformal mapping of multiply connected domains. Numerische Mathematik, 33(4): 437–446, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eicke:1990:ISG

[ELP90] Bertolt Eicke, Alfred K. Louis, and Robert Plato. The instability of some gradient methods for ill-posed problems. Numerische Mathematik, 58(1): 129–134, September 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elbert:1992:ADL

[ELR92] Árpád Elbert, Andrea Laforgia, and Lucia G. Rodonò. Asymptotics for the nth-degree Laguerre polynomial evaluated at n. Numerische Mathematik, 63(2):173–182, November 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elsner:1969:BAM

[Els69] L. Elsner. Bemerkung zu einer Arbeit von Marshall und Olkin. (German) [Remark on a paper of Marshall and Olkin]. Numerische Mathematik, 13(3): 245–246, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elsner:1973:RSI

[Els73] Ludwig Elsner. A remark on simultaneous inclusions of the zeros of a polynomial by Gershgorin’s theorem. Numerische Mathematik, 21(5):425–427, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elschner:1984:GMS

[Els84a] J. Elschner. Galerkin methods with splines for singular integral equations over (0, 1). Numerische Mathematik, 43(2): 265–281, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elsner:1984:NOB

[Els84b] L. Elsner. A note on optimal block-scaling of matrices. Numerische Mathematik, 44(1): 127–128, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elsner:1989:CWR

[Els89] Ludwig Elsner. Comparisons of weak regular splittings and multisplitting methods. Numerische Mathematik, 56(2/3): 283–289, October 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ewing:1991:LRT

[ELV91] R. E. Ewing, R. D. Lazarov, and P. S. Vassilevski. Local refinement techniques for elliptic problems on cell-centered grids. III. algebraic multilevel BEPS preconditioners. Numerische Mathematik, 59(5):431–452, August 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Edelman:1987:ASM

[EM87] Alan Edelman and Charles A. Micchelli. Admissible slopes for monotone and convex interpolation. Numerische Mathematik, 51(4):441–458, August 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elsner:1991:CBI

[EM91] Ludwig Elsner and Volker Mehrmann. Convergence of block iterative methods for linear systems arising in the numerical solution of Euler equations. Numerische Mathematik, 59(6):541–559, September 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eiermann:1988:SSL

[EMN88] Michael Eiermann, Ivo Marek, and Wilhelm Niethammer. On the solution of singular linear systems of algebraic equations by semi-iterative methods. Numerische Mathematik, 53(3): 265–283, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Esser:1980:NDR

[EN80] Henning Esser and Klaus Niederdrenk. Nichtäquidistante Diskretisierungen von Randwertaufgaben. (German) [Nonuniform discretizations of boundary value problems in ODE]. Numerische Mathematik, 35(4): 465–478, November 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eirola:1988:WDM

[EN88] Timo Eirola and Olavi Nevanlinna. What do multistep methods approximate? Numerische Mathematik, 53(5):559–569, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Engels:1972:KND

[Eng72] Hermann Engels. Über Konvergenzkriterien für numerische Differentiationsverfahren. (German) [On convergence criteria for numerical differentiation procedures]. Numerische Mathematik, 19(3):230–237, June 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Engels:1977:FIQ

[Eng77] Hermann Engels. Eine Familie interpolatorischer Quadraturformeln mit ableitungsfreien Fehlerschranken. (German) [A family of quadrature formulas having error bounds without derivatives]. Numerische Mathematik, 28(1):49–58, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Engl:1996:MNS

[Eng96] Gabriele Engl. The modeling and numerical simulation of gas flow networks. Numerische Mathematik, 72(3):349–366, January 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/6072003/60720349.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/6072003/60720349.
                                        htm.

Ennenbach:1995:IFM

[Enn95] R. Ennenbach. The inclusion of free membrane eigenvalues.

Numerische Mathematik, 70(1): 33–44, March 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070001/50700033.
htm;   http://science.springer.
de/nmee/bibs/5070001/50700033.
htm.

Eiermann:1990:OSO

[ENR90] Michael Eiermann, Wilhelm Niethammer, and Arden Ruttan. Optimal successive overrelaxation iterative methods for P - cyclic matrices. Numerische Mathematik, 57(6/7):593–606, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Eiermann:1985:SSM

[ENV85] M. Eiermann, W. Niethammer, and R. S. Varga. A study of semiiterative methods for nonsymmetric systems of linear equations. Numerische Mathematik, 47(4):505–533, December 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elman:1999:ESP

[EO99] Howard C. Elman and Dianne P.

O’Leary. Eigenanalysis of some preconditioned Helmholtz problems. Numerische Mathematik, 83(2):231–257, August 1999.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083002/90830231.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083002/90830231.pdf.

Elliott:1975:CQR

[EP75] David Elliott and D. F. Paget. On the convergence of a quadrature rule for evaluating certain Cauchy principal value integrals. Numerische Mathematik, 23(4):311–319, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elliott:1976:CQR

[EP76] David Elliott and D. F. Paget. On the convergence of a quadrature rule for evaluating certain Cauchy principal value integrals: an addendum. Numerische Mathematik, 25(3): 287–289, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

ElDabaghi:1986:SVT

[EP86] Fadi El Dabaghi and Olivier Pironneau. Stream vectors in three-dimensional aerodynamics. Numerische Mathematik, 48 (5):561–589, May 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elden:1994:PAB

[EP94] Lars Eldén and Haesun Park.

Perturbation analysis for block downdating of a Cholesky decomposition. Numerische Mathematik, 68(4):457–467, October 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068004/40680457.
htm;   http://science.springer.
de/nmee/bibs/4068004/40680457.
htm.

Elliott:1995:AAS

[EP95] David Elliott and Siegfried Prössdorf. An algorithm for the approximate solution of integral equations of Mellin type. Numerische Mathematik, 70(4): 427–452, June 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070004/50700427.
htm;   http://science.springer.
de/nmee/bibs/5070004/50700427.
htm.

Elden:1999:PPS

[EP99] Lars Eldén and Haesun Park.

A Procrustes problem on the Stiefel manifold. Numerische Mathematik, 82(4):599–619, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082004/90820599.htm;   http:

                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9082004/90820599.pdf.

Elsner:1967:RBM

[ER67] L. Elsner and R. M. Redheffer. Remarks on band matrices. Numerische Mathematik, 10(2):153–161, July 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Erisman:1974:MST

[ER74] A. M. Erisman and J. K. Reid. Monitoring the stability of the triangular factorization of a sparse matrix. Numerische Mathematik, 22(3): 183–186, June 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ernst:1996:FEC

[Ern96] Oliver G. Ernst. A finiteelement capacitance matrix method for exterior Helmholtz problems. Numerische Mathematik, 75(2):175–204, December 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer.de/link/service/
                                        journals/00211/bibs/6075002/
                                        60750175.htm;   http://science.
                                        springer.de/nmee/bibs/6075002/
                                        60750175.htm.

Esser:1973:BKH

[ES73] H. Esser and K. Scherer. Eine Bemerkungen zur Konvergenz Hermitischer Interpolationsprozesse. (German)

[Convergence of Hermite interpolation processes]. Numerische Mathematik, 21(3): 220–222, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ellacott:1988:CMG

[ES88] S. W. Ellacott and E. B. Saff.

On Clenshaw’s method and a generalisation to Faber series. Numerische Mathematik, 52(5): 499–509, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eggermont:1990:EBI

[ES90] P. P. B. Eggermont and J. Saranen. L^{p} estimates of boundary integral equations for some nonlinear boundary value problems. Numerische Mathematik, 58(5): 465–478, December 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ewing:1992:AMS

[ES92] John H. Ewing and Glenn Schober. The area of the Mandelbrot set. Numerische Mathematik, 61(1):59–72, February 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Evans:1982:MBA

[ESR82] D. J. Evans, J. Shanehchi, and C. C. Rick. A modified bisection algorithm for the determination of the eigenvalues of a symmetric tridiagonal matrix. Numerische Mathematik, 38(3):

417–419, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Esseen:1960:UAB

[Ess60]  C. G. Esseen.  Über die

asymptotisch beste Approximation stetiger Funktionen mit Hilfe von Bernstein-Polynomen. (German) [On the asymptotically best approximation of continuous functions with the aid of Berstein polynomials]. Numerische Mathematik, 2(1):206– 213, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Esser:1977:SDR

[Ess77] Henning Esser. Stabilitätsungleichung für Diskretisierungen von Randwertaufgaben gewönlicher Differentialgleichungen. (German) [Stability inequalities for discretizations of boundary value problems in ordinary differential equations]. Numerische Mathematik, 28(1):69–100, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eirola:1992:CIS

[ESS92] T. Eirola and J. M. Sanz-Serna. Conservation of integrals and symplectic structure in the integration of differential equations by multistep methods. Numerische Mathematik, 61(3): 281–290, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Elliott:1997:STE

[EV97] David Elliott and Ezio Venturino. Sigmoidal transformations and the Euler– Maclaurin expansion for evaluating certain Hadamard finitepart integrals. Numerische Mathematik, 77(4):453–465, October 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077004/70770453.
htm;   http://science.springer.
de/nmee/bibs/7077004/70770453.
htm.

Eve:1964:EP

[Eve64] J. Eve. The evaluation of polynomials. Numerische Mathematik, 6(1):17–21, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ewing:1992:AMF

[EW92] Richard E. Ewing and JunPing Wang. Analysis of mixed finite element methods on locally refined grids. Numerische Mathematik, 63(2):183–194, November 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eydeland:1984:MSN

[Eyd84] Alexander Eydeland. A method of solving nonlinear variational problems by nonlinear transformation of the objective functional. I. Numerische Mathematik, 43(1):59–82, January 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Eydeland:1985:MSN

[Eyd85] Alexander Eydeland. A method of solving nonlinear variational problems by nonlinear transformation of the objective functional. II. Numerische Mathematik, 46(4):599–609, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Faermann:1998:LPE

[Fae98] Birgit Faermann. Local aposteriori error indicators for the Galerkin discretization of boundary integral equations. Numerische Mathematik, 79(1): 43–76, March 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/8079001/80790043.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/8079001/80790043.
                                        htm.

Fage:1981:IMI

[Fag81a] Dimitri M. Fagé. The indicatrix method. II. Numerische Mathematik, 38(2):255–261, December 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fage:1981:IMF

[Fag81b] Dimitri M. Fagé. The indicatrix method in free boundary problems. Numerische Mathematik, 38(1):39–52, October 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fage:1982:ATD

[Fag82] D. Fagé. Approximation of time-dependent free boundaries. Numerische Mathematik, 40(2): 179–199, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fassbender:1997:IUE

[Faß97] Heike Faßbender. Inverse unitary eigenproblems and related orthogonal functions. Numerische Mathematik, 77(3): 323–345, September 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7077003/
70770323.htm;   http://science.
springer.de/nmee/bibs/7077003/
70770323.htm.

FernandezCara:1989:CTN

[FB89] Enrique Fernández Cara and Mercedes Marín Beltrán. The convergence of two numerical schemes for the Navier–Stokes equations. Numerische Mathematik, 55(1):33–60, April 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fischer:1992:MHF

[FDD92] Patrick Fischer, Mireille Defranceschi, and Joseph Delhalle.

Molecular Hartree–Fock equations for iteration-variation calculations in momentum space. Numerische Mathematik, 63(1): 67–82, October 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Feistauer:1987:FEA

[Fei87] Miloslav Feistauer. On the finite element approximation of a cascade flow problem. Numerische Mathematik, 50(6): 655–684, April 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Felsch:1961:PRG

[Fel61] H. Felsch. Programmierung der Restklassenabzählung einer Gruppe nach Untergruppen. (German) [Programming of residue classes of a group via its subgroups]. Numerische Mathematik, 3:250–256, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fernando:1989:LCR

[Fer89] K. Vince Fernando. Linear convergence of the row cyclic Jacobi and Kogbetliantz methods. Numerische Mathematik, 56(1): 73–91, September 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fernandes:1997:EOS

[Fer97]  Ryan I. Fernandes.  Effi-

cient orthogonal spline collocation methods for solving linear second order hyperbolic problems on rectangles. Numerische Mathematik, 77(2): 223–241, August 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077002/70770223.
htm;   http://science.springer.
de/nmee/bibs/7077002/70770223.
htm.

Fetter:1987:EEA

[Fet87] Alfredo Fetter. L^{∞}-error estimate for an approximation of a parabolic variational inequality. Numerische Mathematik, 50 (5):557–565, March 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Freeden:1987:NIM

[FF87] W. Freeden and J. Fleck. Numerical integration by means of adapted Euler summation formulas. Numerische Mathematik, 51(1):37–64, May 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Falcone:1994:DTH

[FF94] Marizio Falcone and Roberto Ferretti. Discrete time highorder schemes for viscosity solutions of Hamilton–Jacobi– Bellman equations. Numerische Mathematik, 67(3): 315–344, April 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/4067003/40670315.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4067003/40670315.
                                        htm.

Farhloul:1997:DHM

[FF97] Mohamed Farhloul and Michel Fortin. Dual hybrid methods for the elasticity and the Stokes problems: a unified approach. Numerische Mathematik, 76(4): 419–440, June 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/7076004/70760419.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/7076004/70760419.
                                        htm.

Feldstein:1973:NSO

[FG73] Alan Feldstein and Richard Goodman. Numerical solution of ordinary and retarded differential equations with discontinuous derivatives. Numerische Mathematik, 21(1):1–13, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fix:1977:PVP

[FG77] George J. Fix and Morton E. Gurtin. On patched variational principles with application to elliptic and mixed elliptichyperbolic problems. Numerische Mathematik, 28(3): 259–271, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Floria:1989:NPS

[FG89] Luis M. Floría and Robert B.

Griffiths. Numerical procedure for solving a minimization eigenvalue problem. Numerische Mathematik, 55(5):565–574, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fairweather:1967:SHA

[FGM67] G. Fairweather, A. R. Gourlay, and A. R. Mitchell. Some high accuracy difference with a splitting operator for equations of parabolic and elliptic type. Numerische Mathematik, 10(1):56–66, May 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frontini:1987:MPS

[FGM87] Marco Frontini, Walter Gautschi, and Gradimir V. Milovanović. Moment-preserving spline approximation on finite intervals. Numerische Mathematik, 50(5): 503–518, March 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fix:1981:MFE

[FGN81] G. J. Fix, M. D. Gunzburger, and R. A. Nicolaides. On mixed finite element methods for first-order elliptic systems. Numerische Mathematik, 37(1): 29–48, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Feichtinger:1995:ENM

[FGS95] Hans G. Feichtinger, Karlheinz Gröchenig, and Thomas Strohmer. Efficient numerical methods in non-uniform sampling theory. Numerische Mathematik, 69(4):423–440, February 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069004/50690423.
htm;   http://science.springer.
de/nmee/bibs/5069004/50690423.
htm.

Friedrich:1987:PCT

[FH87] Volkmar Friedrich and Bernd Hofmann. A predictor-corrector technique for constrained leastsquares regularization. Numerische Mathematik, 51(3): 353–367, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Funaro:1990:SRA

[FH90] Daniele Funaro and Wilhelm Heinrichs. Some results about the pseudospectral approximation of one-dimensional fourth-order problems. Numerische Mathematik, 58(4): 399–418, November 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fierro:1995:ATS

[FH95] Ricardo D. Fierro and Per Christian Hansen. Accuracy of TSVD solutions computed from rankrevealing decompositions. Numerische Mathematik, 70(4): 453–471, June 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070004/50700453.
htm;   http://science.springer.
de/nmee/bibs/5070004/50700453.
htm.

Frank:1986:CPDa

[FHL86a] R. Frank, J. Hertling, and H. Lehner. B-convergence properties of defect correction methods. I. Numerische Mathematik, 49(2/3):139–162, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frank:1986:CPDb

[FHL86b] R. Frank, J. Hertling, and H. Lehner. B-convergence properties of defect correction methods. II. Numerische Mathematik, 49(2/3):163–188, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Franca:1988:NFS

[FHLM88] Leopoldo P. Franca, Thomas J. R. Hughes, Abimael F. D. Loula, and Isidoro Miranda. A new family of stable elements for nearly incompressible elasticity based on a mixed Petrov–Galerkin finite element formulation. Numerische Mathematik, 53(1/2):123–141, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fiala:1990:MIF

[Fia90] Tibor Fiala. On the monotonity of incomplete factorization. Numerische Mathematik, 57(5):473–479, June 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fiedler:1966:MI

[Fie66] Miroslav Fiedler. Matrix inequalities. Numerische Mathematik, 9(2):109–119, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Field:1978:EBC

[Fie78] David A. Field. Error bounds for continued fractions K(1/b_{n}). Numerische Mathematik, 29(3): 261–267, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fiedler:1987:AVP

[Fie87]  H. Fiedler.  Das asympto-

tische Verhalten der Peanokerne einiger interpolatorischer Quadraturverfahren. (German) [The asymptotic behavior of the Peano kernels of some interpolatory quadrature methods]. Numerische Mathematik, 51(5):

571–581, October 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Filippi:1964:ATA

[Fil64] Siegfried Filippi. Angenäherte Tschebyscheff-Approximation einer Stammfunktion des Verfahrens von Clenshaw und Curtis. (German) [Chebyshev approximation of a primitive function of the Clenshaw–Curtis methods]. Numerische Mathematik, 6(1):320–328, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fisser:1961:STA

[Fis61] H. Fisser. Some tests applied to pseudo-random numbers generated by v. Hörner’s rule. Numerische Mathematik, 3:247–249, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fischer:1981:ADL

[Fis81] Jürgen Fischer. An algorithm for discrete linear L_{p} approximation. Numerische Mathematik, 38(1):129–139, October 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fischer:1993:SGA

[Fis93] Thomas M. Fischer. A spectral Galerkin approximation of the Orr–Sommerfeld eigenvalue problem in a semi-infinite domain. Numerische Mathematik, 66(2):159–179, November 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Field:1972:PET

[FJ72] David A. Field and Willson B. Jones. A priori estimates for truncation error of continued fractions K(1/b_{n}). Numerische Mathematik, 19(4):283–302, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fairweather:1975:EGM

[FJ75] G. Fairweather and J. P. Johnson. On the extrapolation of Galerkin methods for parabolic problems. Numerische Mathematik, 23(3):269–287, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fiala:1987:CDB

[FK87] Tibor Fiala and Anna Krebsz. On the convergence and divergence of Bairstow’s method. Numerische Mathematik, 50(4): 477–482, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fair:1970:PAO

[FL70] Wyman Fair and Yudell L. Luke. Padé approximations to the operator exponential. Numerische Mathematik, 14(4): 379–382, March 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fuhrer:1991:NSD

[FL91] C. Führer and B. J. Leimkuhler. Numerical solution of differential-algebraic equations for constrained mechanical motion. Numerische Mathematik, 59(1):55–69, April 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fraser:1968:AUA

[FM68] M. Fraser and N. Metropolis. Algorithms in unnormalized arithmetic. III: Matrix inversion. Numerische Mathematik, 12(5):416–428, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frommer:1989:PIM

[FM89] Andreas Frommer and Günter Mayer. Parallel interval multisplittings. Numerische Mathematik, 56(2/3):255–267, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fortin:1992:AEMa

[FM92a] M. Fortin and Z. Mghazli.

Analyse d’un élément mixte pour le problème de Stokes. I. Résultats généraux. (French) [Analysis of a mixed element for the Stokes problem. I. general results]. Numerische Mathematik, 62(2):149–160, July 1992.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fortin:1992:AEMb

[FM92b] M. Fortin and Z. Mghazli. Analyse d’un élément mixte pour le problème de Stokes. II. construction et estimations d’erreur. (French) [Analysis of a mixed element for the Stokes problem. II. construction and error estimates]. Numerische Mathematik, 62(2): 161–188, July 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fdez-Manin:1999:EEP

[FMMS99] G. Fdez-Manín and R. Muñoz-Sola. Error estimates for the patching method in the bidimensional case. Numerische Mathematik, 82(4):621–634, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082004/90820621.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082004/90820621.pdf.

Felsch:1968:UPB

[FN68] V. Felsch and J. Neubüser. Über ein Programm zu Berechnung der Automorphismengruppe einer endlichen Gruppe. (German) [On a program for computation of the automorphism group of a finite group]. Numerische Mathematik, 11(3):

277–292, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fix:1972:FEA

[FN72] G. Fix and N. Nassif. On finite element approximations to time-dependent problems. Numerische Mathematik, 19(2): 127–135, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Freund:1991:QQM

[FN91] Roland W. Freund and Noël M.

Nachtigal. QMR: a quasiminimal residual method for non-Hermitian linear systems. Numerische Mathematik, 60(3): 315–339, December 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Feistauer:1998:FEA

[FN98] Miloslav Feistauer and Karel Najzar. Finite element approximation of a problem with a nonlinear Newton boundary condition. Numerische Mathematik, 78(3):403–425, January 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078003/70780403.
htm;   http://science.springer.
de/nmee/bibs/7078003/70780403.
htm.

Funderlic:1982:DGD

[FNP82] R. E. Funderlic, M. Neumann, and R. J. Plemmons. LU decompositions of generalized diagonally dominant matrices. Numerische Mathematik, 40(1):57– 69, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Focke:1965:UMM

[Foc65]  J. Focke.  Über die Mul-

tiplikativität von Matrizennormen. (German) [On the multiplicability of matrix norms]. Numerische Mathematik, 7(3): 251–254, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forster:1967:EFG

[För67] Peter Förster. Existenzaussagen und Fehlerabschätzungen bei gewissen nichtlinearen Randwertaufgaben mit gewöhnlichen Differentialgleichungen. (German) [Existence statement and error bounds for certain nonlinear boundary value problems with ordinary differential equations]. Numerische Mathematik, 10(5):410–422, November 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forster:1968:BIS

[For68a] Peter Forster. Bemerkungen zum Iterationsverfahren von Schulz zur Bestimmung der Inversen einer Matrix. (German) [Remarks on Schulz’s iterative method for computing the inverse of a matrix]. Numerische Mathematik, 12(3):211–214, October 1968. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Forsythe:1968:ADD

[For68b] George E. Forsythe. On the asymptotic directions of the sdimensional optimum gradient method. Numerische Mathematik, 11(1):57–76, January 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forster:1972:DGF

[For72] Peter Forster. Die diskrete Greensche Funktion und Fehlerabschätzungen zum Galerkin-Verfahren. (German) [Discrete variational Green’s function and error bounds for Galerkin approximations]. Numerische Mathematik, 19(5):407–418, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forst:1975:OIQ

[For75] Wilhelm Forst. Zur Optimalität interpolatorischer Quadraturformeln periodischer Funktionen. (German) [Optimal formulas of interpolatory type for the quadrature of periodic functions]. Numerische Mathematik, 25(1):15–21, March 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forst:1978:IID

[For78] Wilhelm Forst. Zur Interpolation und Integration differenzierbarer periodischer Funktionaleng. (German) [On interpolation and integration of differentiable periodic functions]. Numerische Mathematik, 30(2): 137–147, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forster:1982:BOT

[För82]  Klaus-Jürgen Förster.  Be-

merkungen zur optimalen Tschebyscheff-Typ Quadratur. (German) [Remarks on optimal Chebyshevtype quadratures]. Numerische Mathematik, 38(3):421–425, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fortin:1985:TDQ

[For85] M. Fortin. A three-dimensional quadratic nonconforming element. Numerische Mathematik, 46(2):269–279, May 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forster:1993:ECM

[För93] Klaus-Jürgen Förster. Exit criteria and monotonicity in compound quadrature of Gaussian type. Numerische Mathematik, 66(3):321–327, December 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fossmeier:1989:REF

[Föß89]  Reinhard Fößmeier.  On

Richardson extrapolation for finite difference methods on regular grids. Numerische Mathematik, 55(4):451–462, July 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fiedler:1966:SGP

[FP66] Miroslav Fiedler and Vlastimil Pták. Some generalizations of positive definiteness and monotonicity. Numerische Mathematik, 9(2):163–172, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Forster:1990:PPH

[FP90] Klaus-Jürgen Förster and Knut Petras. On a problem proposed by H. Braß concerning the remainder term in quadrature for convex functions. Numerische Mathematik, 57(8):763–777, August 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ferrandiz:1993:FMM

[FP93] José M. Ferrándiz and M. Teresa Pérez. A family of multistep methods to integrate orbits on spheres. Numerische Mathematik, 65(3):285–300, July 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fernando:1994:ASV

[FP94] K. Vince Fernando and Beresford N. Parlett. Accurate singular values and differential qd algorithms. Numerische Mathematik, 67(2): 191–229, March 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067002/40670191.
htm;   http://science.springer.
de/nmee/bibs/4067002/40670191.
htm.

Floater:1999:PLP

[FQ99] Michael S. Floater and Ewald G. Quak. Piecewise linear prewavelets on arbitrary triangulations. Numerische Mathematik, 82(2):221–252, April 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082002/90820221.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082002/90820221.pdf.

Fink:1984:SMS

[FR84] James P. Fink and Werner C. Rheinboldt. Solution manifolds and submanifolds of parametrized equations and their discretization errors. Numerische Mathematik, 45(3): 323–343, December 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Freund:1986:CCA

[FR86] Roland Freund and Stephan Ruscheweyh. On a class of Chebyshev approximation problems which arise in connection with a conjugate gradient type method. Numerische Mathematik, 48(5):525–542, May 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fischer:1988:SRI

[FR88] Bernd Fischer and Lothar Reichel. A stable Richardson iteration method for complex linear systems. Numerische Mathematik, 54(2):225–242, November 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frank:1962:CFEa

[Fra62a] E. Frank. On continued fraction expansions for binomial quadratic surds. Numerische Mathematik, 4:85–95, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frank:1962:CFEb

[Fra62b] Evelyn Frank. On continued fraction expansion for binomial quadratic surds. II. Numerische Mathematik, 4:303–309, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frank:1963:CEB

[Fra63] Evelyn Frank. On continued expansions for binomial quadratic surds. III. Numerische Mathematik, 5(1):113–117, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frank:1976:MID

[Fra76] Reinhard Frank. The method of iterated defect-correction and its application to two-point boundary value problems. I. Numerische Mathematik, 25(4): 409–419, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frank:1977:MID

[Fra77] R. Frank. The method of iterated defect-correction and its application to two-point boundary value problems. II. Numerische Mathematik, 27(4): 407–420, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Freud:1960:NK

[Fre60] G. Freud. Nachtrag bei der Korrektur. (German) [Supplement with correction]. Numerische Mathematik, 2(1): 150, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=2&issue=1&spage=
150.

Freund:1990:CGT

[Fre90] Roland Freund. On conjugate gradient type methods and polynominal preconditioners for a class of complex non-Hermitian matrices. Numerische Mathematik, 57(3):285–312, May 1990.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Freund:1994:LAB

[Fre94] Roland W. Freund. A lookahead Bareiss algorithm for general Toeplitz matrices. Numerische Mathematik, 68(1): 35–69, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680035.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680035.
htm.

Fritsch:1970:NNS

[Fri70] F. N. Fritsch. On the number of nodes in self-contained integration formulas of degree two for compact planar regions. Numerische Mathematik, 16(3): 224–230, December 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frommer:1988:CBN

[Fro88a] Andreas Frommer. Comparison of Brown’s and Newton’s method in the monotone case. Numerische Mathematik, 52(5): 511–521, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frommer:1988:UAM

[Fro88b] Andreas Frommer. A unified approach to methods for the simultaneous computation of all zeros of generalized polynomials.

Numerische Mathematik, 54(1): 105–116, October 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frommer:1989:PNM

[Fro89] Andreas Frommer. Parallel nonlinear multisplitting methods. Numerische Mathematik, 56(2/3):269–282, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ferreira:1970:ZMH

[FS70] Erasmo M. Ferreira and Javier Sesma. Zeros of the modified Hankel function. Numerische Mathematik, 16(3): 278–284, December 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frommer:1992:STS

[FS92] Andreas Frommer and Daniel B. Szyld. H-splittings and twostage iterative methods. Numerische Mathematik, 63(3): 345–356, December 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Frommer:1994:ATS

[FS94] Andreas Frommer and Daniel B. Szyld. Asynchronous two-stage iterative methods. Numerische Mathematik, 69(2):141–153, December 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069002/40690141.
htm;   http://science.springer.
de/nmee/bibs/4069002/40690141.
htm.

Frommer:1999:WMN

[FS99] Andreas Frommer and Daniel B.

Szyld. Weighted max norms, splittings, and overlapping additive Schwarz iterations. Numerische Mathematik, 83 (2):259–278, August 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083002/90830259.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083002/90830259.pdf.

Funaro:1988:DDM

[Fun88] Daniele Funaro. Domain decomposition methods for pseudo spectral approximations. I. second order equations in one dimension. Numerische Mathematik, 52(3):329–344, March 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Freund:1994:CEP

[FV94] Roland W. Freund and Richard S.

Varga. Congratulations [to editor Professor Josef Stoer on the occasion of his seventieth birthday, June 21, 1994]. Numerische Mathematik, 68(1):1, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Xu:1996:SAA

[fX96]  Shu fang Xu.  Sensitivity

analysis of the algebraic Riccati equations. Numerische Mathematik, 75(1):121–134, November 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/6075001/60750121.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/6075001/60750121.
                                        htm.

Feistauer:1987:FES

[FŽ87] Miloslav Feistauer and Alexander Ženíšek. Finite element solution of nonlinear elliptic problems. Numerische Mathematik, 50(4):451–475, July 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Feistauer:1988:CMF

[FŽ88] Miloslav Feistauer and Alexander Ženíšek. Compactness method in the finite element theory of nonlinear elliptic problems. Numerische Mathematik, 52(2):147–163, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Freund:1993:LAA

[FZ93] Roland W. Freund and Hong Yuan Zha. A look-ahead algorithm for the solution of general Hankel systems. Numerische Mathematik, 64(3):295–321, March 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gaviano:1987:LSC

[GA87] M. Gaviano and J. Abaffy.

Line search by curve fitting in minimization algorithms. Numerische Mathematik, 50(3): 321–335, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Garcia-Archilla:1992:SSK

[GA92]  Bosco García-Archilla.  A

supraconvergent scheme for the Korteweg–de Vries equation. Numerische Mathematik, 61(3): 291–310, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gabutti:1984:TMA

[Gab84] Bruno Gabutti. On two methods for accelerating convergence of series. Numerische Mathematik, 43(3):439–461, April 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gaier:1970:SSA

[Gai70]  Dieter Gaier.  Saturation

bei Spline-Approximation und Quadratur. (German) [Saturation in spline approximation and quadrature]. Numerische Mathematik, 16(2):129–140, November 1970. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Gaier:1972:EKM

[Gai72]  Dieter Gaier.  Ermittlung

des konformen Moduls von Vierecken mit Differenzenmethoden. (German) [Determination of the conformal modules of quadrilaterals by difference methods]. Numerische Mathematik, 19(2):179–194, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gal:1989:CMC

[Gal89] S. G. Gal. A construction of monotonically convergent sequences from successive approximations in certain Banach spaces. Numerische Mathematik, 56(1):67–71, September 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gander:1981:LSQ

[Gan81] Walter Gander. Least squares with a quadratic constraint. Numerische Mathematik, 36(3): 291–307, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gao:1989:PTE

[Gao89] Feng Gao. A probabilistic theory for error estimation in automatic integration. Numerische Mathematik, 56(4):309–329, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Garsia:1965:SBB

[Gar65]  Adriano M. Garsia.  Some

bounds on best uniform linear approximation. Numerische Mathematik, 7(3):197–205, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gargantini:1976:PLI

[Gar76] Irene Gargantini. Parallel Laguerre iterations: the complex case. Numerische Mathematik, 26(3):317–323, September 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Garay:1996:SSO

[Gar96]  Barnabas M. Garay.  On

structural stability of ordinary differential equations with respect to discretization methods. Numerische Mathematik, 72(4):449–479, February 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6072004/60720449.
htm;   http://science.springer.
de/nmee/bibs/6072004/60720449.
htm.

Gastinel:1963:DNG

[Gas63] N. Gastinel. Sur-decomposition de normes générales et procédés iteratifs. (French) [Super decomposition of generalized norms and iterative procedures]. Numerische Mathematik, 5(1):142– 151, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gastinel:1965:PCE

[Gas65] N. Gastinel. Propriétés de certains ensembles normes de matrices. (French) [Properties of certain matrix norm ensembles]. Numerische Mathematik, 7(3): 255–260, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gastinel:1971:CPM

[Gas71a] N. Gastinel. Sur le calcul des produits de matrices. (French) [Calculation of the product of matrices]. Numerische Mathematik, 17(3):222–229, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gastinel:1971:ENA

[Gas71b] N. Gastinel. Sur l’extension de normes sur des algèbres de matrices. (French) [On the extension of norms on matrix algebras]. Numerische Mathematik, 17(1):71–83, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gastaldi:1996:MFE

[Gas96] Lucia Gastaldi. Mixed finite element methods in fluid structure systems. Numerische Mathematik, 74(2):153–176, August 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer.de/link/service/

journals/00211/bibs/6074002/
60740153.htm;   http://science.
springer.de/nmee/bibs/6074002/
60740153.htm.

Gautschi:1961:NIO

[Gau61]  Walter Gautschi.  Numeri-

cal integration of ordinary differential equations based on trigonometric polynomials. Numerische Mathematik, 3(1):381– 397, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
381.

Gautschi:1962:IVC

[Gau62] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. Numerische Mathematik, 4:117– 123, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1963:IVC

[Gau63] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. II. Numerische Mathematik, 5(1):425– 430, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1972:AFP

[Gau72] Walter Gautschi. Attenuation factors in practical Fourier analysis. Numerische Mathematik, 18(5):373–400, October 1972.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1973:CAE

[Gau73] Walter Gautschi. On the condition of algebraic equations. Numerische Mathematik, 21(5): 405–424, October 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1975:NEI

[Gau75a] Walter Gautschi. Norm estimates for inverses of Vandermonde matrices. Numerische Mathematik, 23(4):337–347, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1975:OCV

[Gau75b] Walter Gautschi. Optimally conditioned Vandermonde matrices. Numerische Mathematik, 24(1):1–12, February 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1978:IVC

[Gau78] Walter Gautschi. On inverses of Vandermonde and confluent Vandermonde matrices. III. Numerische Mathematik, 29(4): 445–450, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1984:DAS

[Gau84]  Walter Gautschi.  Discrete

approximations to spherically symmetric distributions. Numerische Mathematik, 44(1): 53–60, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1986:SOP

[Gau86] Walter Gautschi. On the sensitivity of orthogonal polynomials to perturbations in the moments. Numerische Mathematik, 48(4):369–382, April 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gotze:1971:RSG

[GB71] H. Götze and F. L. Bauer.

Robert Sauer, geb. 16.9.1898 – gest. 22.8.1970. (German) Robert Sauer, born 16.9.1898 – died 22.8.1970. Numerische Mathematik, 16(4):i–ii, January 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gotze:1981:KSG

[GB81] H. Götze and F. L. Bauer. Klaus Samelson: geb. 21.12.1918, gest. 25.5.1980 (German) [Klaus Samelson: born 21.12.1981, died 25.5.1980]. Numerische Mathematik, 36(2):109, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gui:1986:VFEa

[GB86a] W. Gui and I. Babuška. The h, p and h − p versions of the finite element method for onedimensional problems. Part I:

The error analysis of the pversion. Numerische Mathematik, 49(6):577–612, November 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gui:1986:VFEb

[GB86b] W. Gui and I. Babuška. The h, p and h − p versions of the finite element method for onedimensional problems. Part II: The error analysis of the h and p versions. Numerische Mathematik, 49(6):613–657, November 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gui:1986:VFEc

[GB86c] W. Gui and I. Babuška. The h, p and h − p versions of the finite element method for one-dimensional problems. Part III: The adaptive h − p version. Numerische Mathematik, 49(6): 659–683, November 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guo:1996:PVF

[GC96] Benqi Guo and Weiming Cao. A preconditioner for the h-p version of the finite element method in two dimensions. Numerische Mathematik, 75(1):59– 77, November 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6075001/60750059.

htm;   http://science.springer.
de/nmee/bibs/6075001/60750059.
htm.

Gekeler:1972:MIL

[Gek72] Eckart Gekeler. Über mehrstufige Iterationsverfahren und die Lösung der Hammersteinschen Gleichung. (German) [On several stage iterative methods and the solution of the Hammerstein equations]. Numerische Mathematik, 19(4):351–360, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gekeler:1974:CND

[Gek74] E. Gekeler. On the convergence of the von Neumann difference approximation to hyperbolic initial boundary value problems. Numerische Mathematik, 23(2): 167–174, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gekeler:1978:AEE

[Gek78] E. Gekeler. A-priori error estimates of Galerkin backward differentiation methods in timeinhomogeneous parabolic problems. Numerische Mathematik, 30(4):369–383, December 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gekeler:1982:SBD

[Gek82] Eckart Gekeler. On the stability of backward differentiation methods. Numerische Mathematik, 38(3):467–471, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greig:1980:ISN

[GeN80a] D. M. Greig and M. A. Abd el Naby. Iterative solutions of nonlinear initial value differential equations in Chebyshev series using Lie series. Numerische Mathematik, 34(1):1–13, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greig:1980:SMS

[GeN80b] D. M. Greig and M. A. Abd el Naby. A series method for solving nonlinear two-point boundary value problems. Numerische Mathematik, 34(2): 87–98, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Georg:1979:CII

[Geo79] Kurt Georg. On the convergence of an inverse iteration method for nonlinear elliptic eigenvalue problems. Numerische Mathematik, 32(1):69–74, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Geurts:1982:CTC

[Geu82] A. J. Geurts. A contribution to the theory of condition. Numerische Mathematik, 39(1):85–96, June 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goodman:1973:RER

[GF73] Richard Goodman and Alan Feldstein. Round-off error for retarded ordinary differential equations: a priori bounds and estimates. Numerische Mathematik, 21(5):355–372, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Good:1971:CMN

[GG71] I. J. Good and R. A. Gaskins. The centroid method of numerical integration. Numerische Mathematik, 16(4): 343–359, January 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gout:1986:FQN

[GG86] J. L. Gout and A. Guessab.

Sur les formules de quadrature numérique à nombre minimal de nœuds d’intégration. (French) [Cubature formulas with a minimal number of knots]. Numerische Mathematik, 49(4): 439–455, July 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gerlach:1987:RPI

[GG87] Juergen Gerlach and Ronald Guenther. Remarks on parameter identification. II. convergence of the algorithm. Numerische Mathematik, 51(1): 3–9, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1990:MMI

[GG90] Gene H. Golub and Martin H. Gutknecht. Modified moments for indefinite weight functions. Numerische Mathematik, 57 (6/7):607–624, July 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Gragg:1966:TK

[GH66] W. B. Gragg and A. S. Householder. On a theorem of König. Numerische Mathematik, 8(5): 465–468, August 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gargantini:1972:CAD

[GH72] Irene Gargantini and Peter Henrici. Circular arithmetic and the determination of polynomial zeros. Numerische Mathematik, 18(4):305–320, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gordon:1973:TEM

[GH73] William J. Gordon and Charles A. Hall. Transfinite element methods: blending-function interpolation over arbitrary curved element domains. Numerische Mathematik, 21(2):109–129, April 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gragg:1984:NSR

[GH84] William B. Gragg and William J.

Harrod. The numerically stable reconstruction of Jacobi matrices from spectral data. Numerische Mathematik, 44(3): 317–335, September 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gatica:1992:CBF

[GH92] Gabriel N. Gatica and George C.

Hsiao. On the coupled BEM and FEM for a nonlinear exterior Dirichlet problem in r^{2}. Numerische Mathematik, 61(2): 171–214, March 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gutknecht:1995:LAL

[GH95] Martin H. Gutknecht and Marlis Hochbruck. Look-ahead Levinson and Schur algorithms for non-Hermitian Toeplitz systems. Numerische Mathematik, 70(2):181–227, April 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5070002/
50700181.htm;   http://science.
springer.de/nmee/bibs/5070002/
50700181.htm.

Guglielmi:1999:OSS

[GH99] Nicola Guglielmi and Ernst Hairer. Order stars and stability for delay differential equations. Numerische Mathematik, 83(3):371–383, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083003/90830371.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083003/90830371.pdf.

Ghizzetti:1968:PCQ

[Ghi68] Aldo Ghizzetti. Procedure for constructing quadrature formulae on infinite intervals. Numerische Mathematik, 12(2): 111–119, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guenther:1985:RPI

[GHMG85] Ronald Guenther, Robert Hudspeth, William McDougal, and Jürgen Gerlach. Remarks on parameter identification. I. Numerische Mathematik, 47(3): 355–361, November 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greenspan:1965:VSN

[GHR65] H. Greenspan, W. Hafner, and M. Ribaric. On varying stepsize in numerical integration of first order differential equations. Numerische Mathematik, 7(4): 286–291, August 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gunzburger:1997:AAO

[GHR97] Max D. Gunzburger, L. Steven Hou, and S. S. Ravindran. Analysis and approximation of optimal control problems for a simplified Ginzburg–Landau model of superconductivity. Numerische Mathematik, 77(2): 243–268, August 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077002/70770243.
htm;   http://science.springer.
de/nmee/bibs/7077002/70770243.
htm.

Goodman:1994:SUF

[GHT94] Jonathan Goodman, Thomas Hou, and Eitan Tadmor. On the stability of the unsmoothed Fourier method for hyperbolic equations. Numerische Mathematik, 67(1):93– 129, February 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067001/40670093.
htm;   http://science.springer.
de/nmee/bibs/4067001/40670093.
htm.

Gautschi:1988:LBC

[GI88] Walter Gautschi and Gabriele Inglese. Lower bounds for the condition number of Vandermonde matrices. Numerische Mathematik, 52(3): 241–250, March 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grossmann:1999:PSR

[GI99] C. Grossmann and A. Ismail. Parameter selection rules for path-following with singular embedding. Numerische Mathematik, 82(4):647–675, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9082004/90820647.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9082004/90820647.pdf.

Ginsburg:1963:CGM

[Gin63] Theo Ginsburg. The conjugate gradient method. Numerische Mathematik, 5(1):191– 200, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Girault:1976:CFE

[Gir76]  V. Girault.  A combined

finite element and Marker and Cell method for solving Navier–Stokes equations. Numerische Mathematik, 26(1): 39–59, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Girard:1989:FMC

[Gir89] Didier A. Girard. A fast “Monte Carlo cross-validation” procedure for large least squares problems with noisy data. Numerische Mathematik, 56(1):1– 23, September 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guillot:1989:AFD

[GJ89] J. C. Guillot and P. Joly. Approximation by finite differences of the propagation of acoustic waves in stratified media. Numerische Mathematik, 54(6): 655–702, March 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldstein:1964:MPM

[GK64] A. A. Goldstein and R. B.

Kripke. Mathematical programming by minimizing differentiable functions. Numerische Mathematik, 6(1):47– 48, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gorenflo:1966:SAT

[GK66] Rudolf Gorenflo and Yehudith Kovetz. Solution of an Abeltype integral equation in the presence of noise by quadratic programming. Numerische Mathematik, 8(4):392–406, June 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guderley:1972:BTC

[GK72] K. G. Guderley and C. L. Keller.

A basic theorem in the computation of ellipsoidal error bounds. Numerische Mathematik, 19(3): 218–229, June 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Glashoff:1980:VDP

[GK80] K. Glashoff and H. Kreth. Vorzeichenstabile Differenzenverfahren für parabolische Anfangsrandwertaufgaben. (German) [Sign-stability in difference schemes for parabolic initial-boundary value problems]. Numerische Mathematik, 35(3):343–354, October 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1983:CGQ

[GK83] G. H. Golub and J. Kautský. Calculation of Gauss quadratures with multiple free and fixed knots. Numerische Mathematik, 41(2):147–163, June 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gohberg:1985:NSI

[GK85] I. Gohberg and I. Koltracht. Numerical solution of integral equations, fast algorithms and Kreĭn-Sobolev equation. Numerische Mathematik, 47(2): 237–288, September 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Garus:1990:CSP

[GK90] Rainer Garus and Tassilo Küpper. Computation of the smallest positive eigenvalue of a quadratic λ-matrix. Numerische Mathematik, 57(4): 393–403, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Girao:1994:CCA

[GK94] Pedro Martins Girão and Robert V. Kohn. Convergence of a crystalline algorithm for the heat equation in one dimension and for the motion of a graph by weighted curvature. Numerische Mathematik, 67(1): 41–70, February 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067001/40670041.
htm;   http://science.springer.
de/nmee/bibs/4067001/40670041.
htm.

Grossmann:1986:GED

[GKR86] C. Großmann, M. Krätzschmar, and H.-G. Roos. Gleichmäßig einschließende Diskretisierungsverfahren für schwach nichtlineare Randwertaufgaben. (German) [Uniformly enclosing discretization methods for weakly nonlinear boundary value problems]. Numerische Mathematik, 49(1): 95–110, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gerhards:1965:VBV

[GL65] L. Gerhards and W. Lindenberg.

Ein Verfahren zur Berechnung des vollständigen Untergruppenverbandes endlicher Gruppen auf Dualmaschinen. (German) [A method for calculation of the complete collection of subgroups of finite groups from dual machines]. Numerische Mathematik, 7(1):1–10, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gabutti:1986:SGE

[GL86] Bruno Gabutti and J. N. Lyness. Some generalizations of the Euler-Knopp transformation. Numerische Mathematik, 48(2):199–220, February 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gabutti:1987:NPT

[GL87] Bruno Gabutti and Paolo Lepora. The numerical performance of Tricomi’s formula for inverting the Laplace transform. Numerische Mathematik, 51(3): 369–380, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goodman:1999:CNC

[GL99] T. N. T. Goodman and S. L. Lee. Convergence of nonstationary cascade algorithms. Numerische Mathematik, 84(1):1–33, November 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9084001/90840001.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9084001/90840001.pdf.

Greenbaum:1989:CLS

[GLC89] Anne Greenbaum, Congming Li, and Han Zheng Chao. Comparison of linear system solvers applied to diffusion-type finite element equations. Numerische Mathematik, 56(6): 529–546, December 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grippo:1991:CNS

[GLL91] L. Grippo, F. Lampariello, and S. Lucidi. A class of nonmonotone stabilization methods in unconstrained optimization. Numerische Mathematik, 59(8): 779–805, October 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gilliam:1989:DSI

[GLM89] David S. Gilliam, John R.

Lund, and Clyde F. Martin. A discrete sampling inversion scheme for the heat equation. Numerische Mathematik, 54(5): 493–506, February 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gilliam:1991:IPP

[GLM91] D. S. Gilliam, J. R. Lund, and C. F. Martin. Inverse parabolic problems and discrete orthogonality. Numerische Mathematik, 59(4):361–383, July 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Giunta:1988:MWM

[GLR88] G. Giunta, G. Laccetti, and M. R. Rizzardi. More on the Weeks method for the numerical inversion of the Laplace transform. Numerische Mathematik, 54(2):193–200, November 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gourlay:1966:ADM

[GM66a] A. R. Gourlay and A. R. Mitchell. Alternating direction methods for hyperbolic systems. Numerische Mathematik, 8(2): 137–149, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gourlay:1966:SID

[GM66b] A. R. Gourlay and A. R. Mitchell. A stable implicit difference method for hyperbolic systems in two space variables. Numerische Mathematik, 8(4): 367–375, June 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gourlay:1968:HAD

[GM68] A. R. Gourlay and A. R. Mitchell. High accuracy A.D.I. methods for parabolic equations with variable coefficients. Numerische Mathematik, 12(3): 180–185, October 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1977:OCQ

[GM77] Walter Gautschi and Giovanni Monegato. On optimal Chebyshev-type quadratures. Numerische Mathematik, 28(1):59–67, March 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gasca:1982:LHI

[GM82] M. Gasca and J. I. Maeztu. On Lagrange and Hermite interpolation in r^{k}. Numerische Mathematik, 39(1):1–14, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Graves-Morris:1983:VVR

[GM83] P. R. Graves-Morris. Vector valued rational interpolants. I. Numerische Mathematik, 42(3): 331–348, December 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gautschi:1986:SAS

[GM86] Walter Gautschi and Gradimir V.

Milovanović. Spline approximations to spherically symmetric distributions. Numerische Mathematik, 49(2/3): 111–121, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Graves-Morris:1992:EMV

[GM92] P. R. Graves-Morris. Extrapolation methods for vector sequences. Numerische Mathematik, 61(4):475–487, May 1992.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guillaume:1994:CHO

[GM94] Ph. Guillaume and M. Masmoudi. Computation of high order derivatives in optimal shape design. Numerische Mathematik, 67(2): 231–250, March 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067002/40670231.
htm;   http://science.springer.
de/nmee/bibs/4067002/40670231.
htm.

GmeligMeyling:1987:ACS

[Gme87] R. H. J. Gmelig Meyling. Approximation by cubic C^{1}-splines on arbitrary triangulations. Numerische Mathematik, 51(1): 65–85, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gonzalez:1999:NFR

[GMF99] Ana B. González, Pablo Martín, and José M. Farto. A new family of Runge–Kutta type methods for the numerical integration of perturbed oscillators. Numerische Mathematik, 82(4):635–646, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082004/90820635.htm;   http:

//link.springer-ny.com/link/
service/journals/00211/papers/
9082004/90820635.pdf.

Graves-Morris:1981:RRI

[GMH81] P. R. Graves-Morris and T. R.

Hopkins. Reliable rational interpolation. Numerische Mathematik, 36(2):111–128, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ghaddar:1995:APD

[GMP95] Chahid Ghaddar, Yvon Maday, and Anthony T. Patera. Analysis of a part design procedure. Numerische Mathematik, 71(4):465–510, October 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071004/50710465.
htm;   http://science.springer.
de/nmee/bibs/5071004/50710465.
htm.

Giunta:1995:ABT

[GMS95] G. Giunta, A. Murli, and G. Schmid. An analysis of bilinear transform polynomial methods of inversion of Laplace transforms. Numerische Mathematik, 69(3):269–282, January 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069003/50690269.
htm;   http://science.springer.

                                        de/nmee/bibs/5069003/50690269.
                                        htm.

Guderley:1966:RTP

[GN66] Karl G. Guderley and Paul J. Nikolai. Reduction of twopoint boundary value problems in a vector space to initial value problems by projection. Numerische Mathematik, 8(3): 270–289, May 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gastaldi:1987:OEE

[GN87] Lucia Gastaldi and Ricardo Nochetto. Optimal L^{∞}-error estimates for nonconforming and mixed finite element methods of lowest order. Numerische Mathematik, 50(5):587–611, March 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Graf:1989:BOA

[GNP89] Siegfried Graf, Erich Novak, and Anargyros Papageorgiou. Bisection is not optimal on the average. Numerische Mathematik, 55(4):481–491, July 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gutknecht:1986:SIM

[GNV86] M. H. Gutknecht, W. Niethammer, and R. S. Varga. kstep iterative methods for solving nonlinear systems of equations. Numerische Mathematik, 48(6):699–712, June 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1988:CIC

[GO88] Gene H. Golub and Michael L.

Overton. The convergence of inexact Chebyshev and Richardson iterative methods for solving linear systems. Numerische Mathematik, 53(5):571–593, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gohberg:1994:FIC

[GO94a] I. Gohberg and V. Olshevsky.

Fast inversion of Chebyshev– Vandermonde matrices. Numerische Mathematik, 67(1):71– 92, February 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067001/40670071.
htm;   http://science.springer.
de/nmee/bibs/4067001/40670071.
htm.

Griebel:1994:ASP

[GO94b] M. Griebel and P. Oswald. On additive Schwarz preconditioners for sparse grid discretizations. Numerische Mathematik, 66(4):449–463, January 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Griebel:1995:ATA

[GO95] M. Griebel and P. Oswald.

On the abstract theory of additive and multiplicative Schwarz algorithms. Numerische Mathematik, 70(2): 163–180, April 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070002/50700163.
htm;   http://science.springer.
de/nmee/bibs/5070002/50700163.
htm.

Goel:1968:CBF

[Goë68] J.-J. Goël. Construction of basic functions for numerical utilisation of Ritz’s method. Numerische Mathematik, 12(5): 435–447, December 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldstein:1962:CMM

[Gol62] A. A. Goldstein. Cauchy’s method of minimization. Numerische Mathematik, 4:146– 150, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldstein:1963:SRA

[Gol63] A. A. Goldstein. On the stability of rational approximation. Numerische Mathematik, 5(1): 431–438, December 1963. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=5&issue=1&spage=
431.

Goldstein:1965:NM

[Gol65a] A. A. Goldstein. On Newton’s method. Numerische Mathematik, 7(5):391–393, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1965:NMS

[Gol65b] G. H. Golub. Numerical methods for solving linear least squares problems. Numerische Mathematik, 7(3):206–216, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldstein:1981:FEM

[Gol81] C. I. Goldstein. The finite element method with nonuniform mesh sizes applied to the exterior Helmholtz problem. Numerische Mathematik, 38(1):61– 82, October 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gorenflo:1970:NSD

[Gor70] Rudolf Gorenflo. Nichtnegativitäts-und substanzerhaltende Differenzenschemata für lineare Diffusionsgleichungen. (German) [Nonnegativity and substance-receiving difference patterns for linear diffusion equations]. Numerische Mathematik, 14(5):448–467, April 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-

599X&volume=14&issue=5&spage=
448.

Griebel:1999:SGB

[GOS99] M. Griebel, P. Oswald, and T. Schiekofer. Sparse grids for boundary integral equations. Numerische Mathematik, 83(2):279–312, August 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083002/90830279.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083002/90830279.pdf.

Gout:1977:EEI

[Gou77] J. L. Gout. Estimation de l’erreur d’interpolation d’Hermite dans R^{n}. (French) [Estimation of the Hermite interpolation error in R^{n}]. Numerische Mathematik, 28(4):407–429, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Govaerts:1990:BAL

[Gov90] W. Govaerts. Bordered augmented linear systems in numerical continuation and bifurcation. Numerische Mathematik, 58(4):353–368, November 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greenspan:1965:MNE

[GP65] Donald Greenspan and Seymour V. Parter. Mildly nonlinear elliptic partial differential equations and their numerical solution. II. Numerische Mathematik, 7(2):129–146, April 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Goldstein:1967:EAM

[GP67] A. A. Goldstein and J. F. Price.

An effective algorithm for minimization. Numerische Mathematik, 10(3):184–189, October 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Galimberti:1971:SCV

[GP71] G. Galimberti and V. Pereyra.

Solving confluent Vandermonde systems of Hermite type. Numerische Mathematik, 18(1):44– 60, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Graves:1978:NIF

[GP78] Jack Graves and P. M. Prenter.

Numerical iterative filters applied to first kind Fredholm integral equations. Numerische Mathematik, 30(3):281–299, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Glowinski:1979:MFE

[GP79] R. Glowinski and O. Pironneau. On a mixed finite element approximation of the Stokes problem. I. convergence of the approximate solutions. Numerische Mathematik, 33(4):

397–424, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gunzburger:1983:CFE

[GP83] Max D. Gunzburger and Janet S. Peterson. On conforming finite element methods for the inhomogeneous stationary Navier–Stokes equations. Numerische Mathematik, 42(2): 173–194, November 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gorlich:1985:AWP

[GP85] E. Görlich and D. Pontzen. Alpha well-posedness, alpha stability, and the matrix theorems of H.-O. Kreiss. Numerische Mathematik, 46(1): 131–146, March 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guermond:1998:AUN

[GQ98] J.-L. Guermond and L. Quartapelle. On the approximation of the unsteady Navier– Stokes equations by finite element projection methods. Numerische Mathematik, 80(2): 207–238, August 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080002/80800207.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080002/80800207.

pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080002/80800207.
ps.gz.

Golub:1970:SVD

[GR70a] G. H. Golub and C. Reinsch. Singular value decomposition and least squares solutions. Numerische Mathematik, 14(5):403–420, April 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 134– 151].

Guenther:1970:NCF

[GR70b] R. B. Guenther and E. L.

Roetman. Newton–Cotes formulae in n-dimensions. Numerische Mathematik, 14(4): 330–345, March 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Girault:1979:AMF

[GR79] V. Girault and P. A. Raviart.

An analysis of a mixed finite element method for the Navier–Stokes equations. Numerische Mathematik, 33(3): 235–271, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Griewank:1986:AGT

[GR86] A. Griewank and G. W. Reddien. The approximation of generalized turning points by projection methods with superconvergence to the critical parameter. Numerische Mathematik, 48(5):591–606, May 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gragg:1990:DCM

[GR90] W. B. Gragg and L. Reichel. A divide and conquer method for unitary and orthogonal eigenproblems. Numerische Mathematik, 57(8):695–718, August 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gragg:1968:TEB

[Gra68] William B. Gragg. Truncation error bounds for g-fractions. Numerische Mathematik, 11(5): 370–379, June 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greenstadt:1962:SNE

[Gre62] John Greenstadt. Some numerical experiments in triangularizing matrices. Numerische Mathematik, 4:187–195, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greenberg:1975:ADR

[Gre75] Harold Greenberg. An algorithm for determining redundant inequalities and all solutions to convex polyhedra. Numerische Mathematik, 24(1):19– 26, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greenbaum:1979:CSU

[Gre79] A. Greenbaum. Comparison of splittings used with the conjugate gradient algorithm. Numerische Mathematik, 33(2): 181–193, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greenberg:1980:ALD

[Gre80] Harold Greenberg. An algorithm for a linear Diophantine equation and a problem of Frobenius. Numerische Mathematik, 34(4):349–352, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greaves:1982:AHM

[Gre82] G. Greaves. An algorithm for the Hausdorff moment problem. Numerische Mathematik, 39(2): 231–238, August 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gries:1967:CCC

[Gri67]  David Gries.  Characteriza-

tions of certain classes of norms. Numerische Mathematik, 10(1): 30–41, May 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grigorieff:1970:KRE

[Gri70a] Rolf Dieter Grigorieff. Die Konvergenz des Rand- und Eigenwertproblems linearer gewöhnlicher Differenzengleichungen. (German) [The convergence of boundary- and eigenvalue problems of linear difference equations]. Numerische Mathematik, 15(1):15–48, May 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grigorieff:1970:KGD

[Gri70b] Rolf Dieter Grigorieff. Über die Koerzitivität gewönlicher Differenzenoperatoren und die Konvergenz von Mehrschrittverfahren. (German) [On the coerciveness of ordinary difference operators and the convergence of multistep methods]. Numerische Mathematik, 15(3): 196–218, July 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grigorieff:1975:DAEc

[Gri75a] Rolf Dieter Grigorieff. Diskrete Approximation von Eigenwerteproblemen. III. Asymptotische Entwicklungen. (German) [Discrete approximation of eigenvalue-problems. III. Asymptotic expansions]. Numerische Mathematik, 25(1): 79–97, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grigorieff:1975:DAEa

[Gri75b] Rolf Dieter Grigorieff. Diskrete Approximation von Eigenwertproblemen. I. Qualitative Konvergenz. (German) [Discrete approximation of the eigenvalue problem. I. Qualitative convergence]. Numerische Mathematik, 24(4):355–374, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grigorieff:1975:DAEb

[Gri75c] Rolf Dieter Grigorieff. Diskrete Approximation von Eigenwertproblemen. II. Konvergenzordnung. (German) [Discrete approximation of eigenvalue problems. II. Convergence order]. Numerische Mathematik, 24(5): 415–433, October 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Griewank:1980:SDC

[Gri80] A. O. Griewank. Starlike domains of convergence for Newton’s method at singularities. Numerische Mathematik, 35(1): 95–111, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grigorieff:1983:SMM

[Gri83] Rolf Dieter Grigorieff. Stability of multistep-methods on variable grids. Numerische Mathematik, 42(3):359–377, December 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grigorieff:1985:ESG

[Gri85] Rolf Dieter Grigorieff. Einige Stabilitätsungleichungen für gewöhnliche Differenzengleichungen in nichtkompakter Form. (German) [Some stability inequalities for ordinary difference equations in noncompact form]. Numerische Mathematik, 47(4): 565–576, December 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grozev:1989:CTM

[Gro89] Georgi R. Grozev. Comparison theorems for L-monosplines of minimal norm. Numerische Mathematik, 56(4):331–343, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Grune:1997:AGS

[Grü97]  Lars Grüne.  An adap-

tive grid scheme for the discrete Hamilton–Jacobi–Bellman equation. Numerische Mathematik, 75(3):319–337, January 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.de/link/service/
journals/00211/bibs/7075003/
70750319.htm;   http://science.
springer.de/nmee/bibs/7075003/
70750319.htm.

Geistbeck:1974:OFD

[GS74] F. Geistbeck and A. Sachs. Optimale Fehlerabschätzungen mit dualen Funktionalen bei semilinearen Randwertaufgaben gewöhnlicher Differentialgleichungen. (German) [Optimal error estimations with dual functionals for boundary value problems of ordinary semilinear differential equations]. Numerische Mathematik, 22(2):

127–136, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Glashoff:1977:TNA

[GS77] K. Glashoff and E. Sachs.

On theoretical and numerical aspects of the bang-bang-principle. Numerische Mathematik, 29(1):93–113, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sun:1983:PBE

[gS83] Ji guang Sun. The perturbation bounds for eigenspaces of a definite matrix-pair. Numerische Mathematik, 41(3): 321–343, July 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sun:1991:ERQ

[gS91] Ji guang Sun. Eigenvalues of Rayleigh quotient matrices. Numerische Mathematik, 59(6): 603–614, September 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sun:1992:CNN

[gS92]  Ji guang Sun.  On condi-

tion numbers of a nondefective multiple eigenvalue. Numerische Mathematik, 61(2): 265–275, March 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sun:1993:BPA

[gS93a]  Ji guang Sun.  Backward

perturbation analysis of certain characteristic subspaces. Numerische Mathematik, 65(3): 357–382, July 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guo:1993:FEA

[GS93b] Wen Guo and Martin Stynes. Finite element analysis of exponentially fitted Lumped schemes for time-dependent convection-diffusion problems. Numerische Mathematik, 66(3): 347–371, December 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gottlieb:1995:GPV

[GS95a] David Gottlieb and Chi-Wang Shu. On the Gibbs phenomenon. V. recovering exponential accuracy from collocation point values of a piecewise analytic function. Numerische Mathematik, 71(4):511–526, October 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071004/50710511.
htm;   http://science.springer.
de/nmee/bibs/5071004/50710511.
htm.

Sun:1995:WCC

[gS95b] Ji guang Sun. On worst-case condition numbers of a nondefective multiple eigenvalue.

Numerische Mathematik, 69(3): 373–382, January 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5069003/
50690373.htm;   http://science.
springer.de/nmee/bibs/5069003/
50690373.htm.

Govaerts:1996:DHP

[GS96a] W. Govaerts and A. Spence. Detection of Hopf points by counting sectors in the complex plane. Numerische Mathematik, 75(1): 43–58, November 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6075001/60750043.
htm;   http://science.springer.
de/nmee/bibs/6075001/60750043.
htm.

Sun:1996:PAS

[gS96b] Ji guang Sun. Perturbation analysis of singular subspaces and deflating subspaces. Numerische Mathematik, 73(2): 235–263, April 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073002/60730235.
htm;   http://science.springer.
de/nmee/bibs/6073002/60730235.
htm.

Sun:1997:RBA

[gS97] Ji guang Sun. Residual bounds of approximate solutions of the algebraic Riccati equation. Numerische Mathematik, 76(2): 249–263, April 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076002/70760249.
htm;   http://science.springer.
de/nmee/bibs/7076002/70760249.
htm.

Sun:1998:PAG

[gS98a]  Ji guang Sun.  Perturba-

tion analysis of generalized singular subspaces. Numerische Mathematik, 79(4): 615–641, June 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079004/80790615.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790615.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790615.
ps.gz.

Sun:1998:RBA

[gS98b] Ji guang Sun. Residual bounds of approximate solutions of the discrete-time algebraic Riccati equation. Numerische Mathematik, 78(3):463–478, January 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078003/70780463.
htm;   http://science.springer.

de/nmee/bibs/7078003/70780463.
htm.

Guo:1998:VCF

[GS98c] Benqi Guo and Ernst P.

Stephan. The h-p version of the coupling of finite element and boundary element methods for transmission problems in polyhedral domains. Numerische Mathematik, 80(1): 87–107, July 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080001/80800087.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800087.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800087.
ps.gz.

Sun:1999:BEI

[gS99] Ji guang Sun. Backward errors for the inverse eigenvalue problem. Numerische Mathematik, 82(2):339–349, April 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082002/90820339.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082002/90820339.pdf.

Gaier:1967:RCO

[GT67] Dieter Gaier and John Todd.

On the rate of convergence of optimal ADI processes. Numerische Mathematik, 9(5):452– 459, April 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golomb:1972:MGN

[GT72] M. Golomb and R. A. Tapia. The metric gradient in normed linear spaces. Numerische Mathematik, 20(2):115–124, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Good:1978:IST

[GT78] I. J. Good and T. N. Tideman. Integration over a simplex, truncated cubes, and Eulerian numbers. Numerische Mathematik, 30(4):355–367, December 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Griewank:1982:LCA

[GT82a] A. Griewank and Ph. L. Toint. Local convergence analysis for partitioned quasi-Newton updates. Numerische Mathematik, 39(3):429–448, October 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Griewank:1982:PVM

[GT82b] A. Griewank and Ph. L. Toint. Partitioned variable metric updates for large structured optimization problems. Numerische Mathematik, 39(1): 119–137, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gaviano:1984:GCT

[GT84] M. Gaviano and F. Testa. A general convergence theorem for the reduced gradient method. Numerische Mathematik, 43(2): 241–247, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gilbert:1987:AND

[GT87] John R. Gilbert and Robert Endre Tarjan. The analysis of a nested dissection algorithm. Numerische Mathematik, 50(4): 377–404, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guessab:1986:CFW

[Gue86] A. Guessab. Cubature formulae which are exact on spaces P, intermediate betweenP_{k} and Q_{k′}. Numerische Mathematik, 49(5): 561–576, September 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guessab:1988:NCF

[Gue88] A. Guessab. Numerical cubature formulae with preassigned knots. Numerische Mathematik, 52(4):467–478, April 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guermond:1994:RMP

[Gue94] Jean-Luc Guermond. Remarques sur les méthodes de projection pour l’approximation des équations de Navier–Stokes. (French) [Remarks on projection methods for approximating Navier–Stokes equations]. Numerische Mathematik, 67(4): 465–473, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067004/40670465.
htm;   http://science.springer.
de/nmee/bibs/4067004/40670465.
htm.

Guglielmi:1997:ASP

[Gug97]  Nicola Guglielmi.  On the

asymptotic stability properties of Runge–Kutta methods for delay differential equations. Numerische Mathematik, 77(4): 467–485, October 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7077004/
70770467.htm;   http://science.
springer.de/nmee/bibs/7077004/
70770467.htm.

Gunn:1964:NSS

[Gun64a] J. E. Gunn. The numerical solution of ∇ · a∇u = f by a semi-explicit alternatingdirection iterative technique. Numerische Mathematik, 6(1): 181–184, December 1964. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gunn:1964:TSI

[Gun64b] J. E. Gunn. On the twostage iterative method of Douglas for mildly non-linear elliptic difference equations. Numerische Mathematik, 6(1):243– 249, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gunther:1975:ZQG

[Gün75] Claus Günther. Zweidimensionale Quadraturformeln von Grad 7 mit 14 Punkten. (German) [Two dimensional quadrature formulae of degree seven with fourteen common zeros]. Numerische Mathematik, 24(4): 309–316, August 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gunttner:1994:EBU

[Gün94] R. Günttner. Exact bounds for the uniform approximation by cardinal spline interpolants. Numerische Mathematik, 68(2): 263–267, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068002/40680263.
htm;   http://science.springer.
de/nmee/bibs/4068002/40680263.
htm.

Gunther:1998:SDC

[Gün98] Michael Günther. Simulating digital circuits numerically — a charge-oriented ROW approach. Numerische Mathematik, 79(2):

203–212, April 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079002/80790203.
htm;   http://science.springer.
de/nmee/bibs/8079002/80790203.
htm.

Guo:1988:VFE

[Guo88] Ben Qi Guo. The h-p version of the finite element method for elliptic equations of order 2m. Numerische Mathematik, 53(1/2): 199–224, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Guo:1999:IBF

[Guo99] Chun-Hua Guo. Incomplete block factorization preconditioning for indefinite elliptic problems. Numerische Mathematik, 83(4):621–639, October 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083004/90830621.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083004/90830621.pdf.

Gustafsson:1973:NMS

[Gus73] Bertil Gustafsson. A numerical method for solving singular boundary value problems. Numerische Mathematik, 21(4): 328–344, August 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gustafson:1985:SCA

[Gus85] Sven-Åke Gustafson. Stable convergence acceleration using Laplace transforms. Numerische Mathematik, 47(3): 387–394, November 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gutknecht:1972:PEB

[Gut72] Martin Gutknecht. A posteriori error bounds for the zeros of a polynomial. Numerische Mathematik, 20(2):139–148, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gutknecht:1981:STI

[Gut81] Martin H. Gutknecht. Solving Theodorsen’s integral equation for conformal maps with the fast Fourier transform and various nonlinear iterative methods. Numerische Mathematik, 36(4):405–429, May 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gutknecht:1987:AFM

[Gut87] Martin H. Gutknecht. Attenuation factors in multivariate Fourier analysis. Numerische Mathematik, 51(6): 615–629, November 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gutknecht:1989:CFA

[Gut89a] Martin H. Gutknecht. Continued fractions associated with the Newton–Padé table. Numerische Mathematik, 56(6): 547–589, December 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gutknecht:1989:SAS

[Gut89b] Martin H. Gutknecht. Stationary and almost stationary iterative (k, l)-step methods for linear and nonlinear systems of equations. Numerische Mathematik, 56(2/3):179–213, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1961:CSIa

[GV61a] Gene H. Golub and Richard S. Varga. Chebyshev semi-iterative methods, successive overrelaxation iterative methods, and second order Richardson iterative methods. I. Numerische Mathematik, 3:147–156, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1961:CSIb

[GV61b] Gene H. Golub and Richard S. Varga. Chebyshev semi-iterative methods, successive overrelaxation iterative methods, and second order Richardson iterative methods. II. Numerische Mathematik, 3:157–168, December 1961. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1988:OPH

[GV88] Gene H. Golub and Richard S.

Varga. Obituary to Peter Henrici (13. September 1923–13. March 1987). Numerische Mathematik, 52(5):481–482, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1991:QCL

[GvM91] Gene H. Golub and Urs von Matt. Quadratically constrained least squares and quadratic problems. Numerische Mathematik, 59(6):561–580, September 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golub:1966:NIR

[GW66] G. H. Golub and J. H. Wilkinson. Note on the iterative refinement of least squares solution. Numerische Mathematik, 9(2): 139–148, December 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gekeler:1986:OCR

[GW86] E. Gekeler and R. Widmann. On the order conditions of Runge– Kutta methods with higher derivatives. Numerische Mathematik, 50(2):183–203, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gohner:1994:SIA

[GW94] U. Göhner and G. Warnecke. A shock indicator for adaptive transonic flow computations. Numerische Mathematik, 66(4):423–448, January 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gohner:1995:SOF

[GW95a] U. Göhner and G. Warnecke. A second order finite difference error indicator for adaptive transonic flow computations. Numerische Mathematik, 70(2): 129–161, April 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070002/50700129.
htm;   http://science.springer.
de/nmee/bibs/5070002/50700129.
htm.

Govaerts:1995:CBM

[GW95b] Willy Govaerts and Bodo Werner. Continuous bordering of matrices and continuous matrix decompositions. Numerische Mathematik, 70(3): 303–310, May 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070003/50700303.
htm;   http://science.springer.
de/nmee/bibs/5070003/50700303.
htm.

Gautschi:1995:COP

[GZ95] Walter Gautschi and Minda Zhang. Computing orthogonal polynomials in Sobolev spaces. Numerische Mathematik, 71(2): 159–183, August 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071002/50710159.
htm;   http://science.springer.
de/nmee/bibs/5071002/50710159.
htm.

S:1973:AHA

[H. 73] H. S. Alston S. Householder Award. Numerische Mathematik, 21(1):i, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hegland:1998:MFI

[HA98] M. Hegland and R. S. Anderssen. A mollification framework for improperly posed problems. Numerische Mathematik, 78(4): 549–575, February 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7078004/
70780549.htm;   http://science.
springer.de/nmee/bibs/7078004/
70780549.htm.

Haas:1960:BEA

[Haa60] Hans-Lothar Haas. Bestimmung der Eigenwerte des allgemeinen Eigenwertproblems. (German)

[Determination of the eigenvalues of the general eigenvalue problem]. Numerische Mathematik, 2(1):341–343, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=2&issue=1&spage=
341.

Hackbusch:1977:ELN

[Hac77a] W. Hackbusch. Extrapolation to the limit for numerical solutions of hyperbolic equations. Numerische Mathematik, 28(4): 455–474, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:1977:EAC

[Hac77b] Wolfgang Hackbusch. Extrapolation applied to certain discretization methods solving the initial value problem for hyperbolic differential equations. Numerische Mathematik, 28(2): 121–142, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:1979:FSN

[Hac79]  Wolfgang Hackbusch.  On

the fast solutions of nonlinear elliptic equations. Numerische Mathematik, 32(1): 83–95, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:1989:FDM

[Hac89]  Wolfgang Hackbusch.  The

frequency decomposition multigrid method. I. application to anisotropic equations. Numerische Mathematik, 56(2/3): 229–245, October 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:1992:FDM

[Hac92]  Wolfgang Hackbusch.  The

frequency decomposition multi grid method. II. Convergence analysis based on the additive Schwarz method. Numerische Mathematik, 63(4): 433–453, December 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadeler:1968:NVI

[Had68]  K. P. Hadeler.  Newton-

Verfahren für inverse Eigenwertaufgaben. (German) [Newton method for inverse eigenvalue problems]. Numerische Mathematik, 12(1):35–39, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadeler:1969:EAM

[Had69a] K. P. Hadeler. Einige Anwendungen mehrparametriger Eigenwertaufgaben. (German) [Some uses of extra parametrized eigenvalue problems]. Numerische Mathematik, 13(3): 285–292, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadjidimos:1969:SHA

[Had69b]  Apostolos Hadjidimos.  On

some high accuracy difference schemes for solving elliptic equations. Numerische Mathematik, 13(5):396–403, October 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadjidimos:1971:NSM

[Had71] A. Hadjidimos. The numerical solution of a model problem biharmonic equation by using extrapolated alternating direction implicit methods. Numerische Mathematik, 17(4):301–317, August 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadeler:1972:BAW

[Had72a] K. P. Hadeler. Bemerkungen zu einer Arbeit von W. Wetterling über positive Operatoren. (German) [Remarks on a work by W. Wetterling on positive operators]. Numerische Mathematik, 19(3):260–265, June 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadeler:1972:BGG

[Had72b] K. P. Hadeler. Bemerkungen zum Gradientenverfahren. (German) [Remarks on gradient methods]. Numerische Mathematik, 18(5):413–422, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadeler:1980:ECP

[Had80] K. P. Hadeler. Effective computation of periodic orbits and bifurcation diagrams in delay equations. Numerische Mathematik, 34(4):457–467, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadjidimos:1987:OSN

[Had87] A. Hadjidimos. Optimum stationary and nonstationary iterative methods for the solution of singular linear systems. Numerische Mathematik, 51(5): 517–530, October 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadeler:1996:SOK

[Had96] K. P. Hadeler. Shadowing orbits and Kantorovich’s theorem. Numerische Mathematik, 73(1): 65–73, March 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073001/60730065.
htm;   http://science.springer.
de/nmee/bibs/6073001/60730065.
htm.

Haemmerlin:1963:UAS

[Hae63] Günther Haemmerlin. Über ableitungsfreie Schranken für Quadraturfehler. (German) [On derivative-free bounds for quadrature error]. Numerische Mathematik, 5(1):226–233, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hager:1987:MFF

[Hag87] William W. Hager. A modified fast Fourier transform for polynomial evaluation and the Jenkins–Traub algorithm. Numerische Mathematik, 50(3): 253–261, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1977:MNP

[Hai77] E. Hairer. Méthodes de Nyström pour l’équation différentielle y′′ = f(x, y). (French) [Nyström methods for the differential equation y′′ = f(x, y)]. Numerische Mathematik, 27(3): 283–300, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1978:OID

[Hai78]  E. Hairer.  On the order

of iterated defect correction — an algebraic proof. Numerische Mathematik, 29(4): 409–424, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1979:USM

[Hai79]  E. Hairer.  Unconditionally

stable methods for second order differential equations. Numerische Mathematik, 32(4):

373–379, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1980:USE

[Hai80]  E. Hairer.  Unconditionally

stable explicit methods for parabolic equations. Numerische Mathematik, 35(1): 57–68, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1981:OCN

[Hai81] E. Hairer. Order conditions for numerical methods for partitioned ordinary differential equations. Numerische Mathematik, 36(4):431–445, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1982:CCS

[Hai82] E. Hairer. Constructive characterization of A-stable approximations to exp(z) and its connection with algebraically stable Runge–Kutta methods. Numerische Mathematik, 39(2): 247–258, August 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1986:SRK

[Hai86] E. Hairer. A- and B-stability for Runge–Kutta methods— characterizations and equivalence. Numerische Mathematik, 48(4):383–389, April 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1999:BEA

[Hai99] Ernst Hairer. Backward error analysis for multistep methods. Numerische Mathematik, 84(2):199–232, December 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9084002/90840199.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9084002/90840199.pdf.

Halton:1960:BEC

[Hal60a] J. H. Halton. Berichtigung: “On the efficiency of certain quasirandom sequences of points in evaluating multi-dimensional integrals” [Num. Math. 2, 84– 90 (1960)]. (German) [Correction]. Numerische Mathematik, 2(1):196, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Hal60b].

Halton:1960:ECQ

[Hal60b] J. H. Halton. On the efficiency of certain quasi-random sequences of points in evaluating multi-dimensional integrals. Numerische Mathematik, 2(1): 84–90, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See correction [Hal60a].

Hald:1975:NTM

[Hal75] Ole H. Hald. On a Newton-Moser type method. Numerische Mathematik, 23(5): 411–426, October 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hald:1977:DIS

[Hal77] Ole H. Hald. Discrete inverse Sturm–Liouville problems. I. uniqueness for symmetric potentials. Numerische Mathematik, 27(2):249–256, June 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hallatschek:1992:FDG

[Hal92] Klaus Hallatschek. Fouriertransformation auf dünnen Gittern mit hierarchischen Basen. (German) [Fourier transform on sparse grids with hierarchical bases]. Numerische Mathematik, 63(1):83–97, October 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hammerlin:1965:ASQ

[Häm65]  Günther Hämmerlin.  Über

ableitungsfreie Schranken für Quadraturfehler. II: Ergänzungen und Möglichkeiten zur Verbesserung. (German) [On constraints without derivatives for quadrature errors. II: Completions and possibilities to improve]. Numerische Mathematik, 7(3):232– 237, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hammerlin:1966:AQA

[Häm66]  Günther Hämmerlin.  Zur

Abschätzung von Quadraturfehlern für analytische Funktionen. (German) [On bounds for quadrature error for analytic funktions]. Numerische Mathematik, 8(4):334–344, June 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Handscomb:1964:RMC

[Han64] D. C. Handscomb. Remarks on a Monte Carlo integration method. Numerische Mathematik, 6(1):261–268, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hansen:1980:GOU

[Han80] Eldon Hansen. Global optimization using interval analysis — the multi-dimensional case. Numerische Mathematik, 34(3): 247–270, September 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Han:1982:NSI

[Han82]  Houde Han.  The numeri-

cal solutions of interface problems by infinite element method. Numerische Mathematik, 39(1): 39–50, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hanke:1988:LSC

[Han88] Michael Hanke. On a leastsquares collocation method for linear differential-algebraic equations. Numerische Mathematik, 54(1):79–90, October 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Han:1991:BEM

[Han91a] Houde Han. A boundary element method for Signorini problems in three dimensions. Numerische Mathematik, 60(1):63– 75, November 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hanke:1991:ALI

[Han91b] Martin Hanke. Accelerated Landweber iterations for the solution of ill-posed equations. Numerische Mathematik, 60(3): 341–373, December 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Han:1992:FEA

[Han92]  Weimin Han.  Finite ele-

ment analysis of a holonomic elastic-plastic problem. Numerische Mathematik, 60(4): 493–508, January 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Han:1994:BID

[Han94]  Houde Han.  The bound-

ary integro-differential equations of three-dimensional Neumann problem in linear elasticity. Numerische Mathematik, 68 (2):269–281, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068002/40680269.
htm;   http://science.springer.
de/nmee/bibs/4068002/40680269.
htm.

Hanke:1997:SCR

[Han97] Martin Hanke. Superlinear convergence rates for the Lanczos method applied to elliptic operators. Numerische Mathematik, 77(4):487–499, October 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077004/70770487.
htm;   http://science.springer.
de/nmee/bibs/7077004/70770487.
htm.

Hao:1994:MMI

[Hào94] Dinh Nho Hào. A mollification method for ill-posed problems. Numerische Mathematik, 68(4): 469–506, October 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4068004/
40680469.htm;   http://science.
springer.de/nmee/bibs/4068004/
40680469.htm.

Harary:1962:GTA

[Har62] Frank Harary. A graph theoretic approach to matrix inversion by partitioning. Numerische Mathematik, 4:128–135, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hartung:1978:SIP

[Har78]  Joachim Hartung.  A sta-

ble interior penalty method for convex extremal problems. Numerische Mathematik, 29(2): 149–158, January 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hartfiel:1980:CSS

[Har80] D. J. Hartfiel. Concerning the solution set of Ax = b where P ≤ A ≤ Q and p ≤ b ≤ q. Numerische Mathematik, 35(3): 355–359, October 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hari:1982:GCE

[Har82] V. Hari. On the global convergence of the Eberlein method for real matrices. Numerische Mathematik, 39(3):361–369, October 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hari:1991:SQC

[Har91] Vjeran Hari. On sharp quadratic convergence bounds for the serial Jacobi methods. Numerische Mathematik, 60(3): 375–406, December 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Haussler:1986:KTC

[Häu86] W. M. Häussler. A Kantorovichtype convergence analysis for the Gauss–Newton method. Numerische Mathematik, 48(1): 119–125, January 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Haverkamp:1977:SLP

[Hav77] Richard Haverkamp. Sturm– Liouville-Probleme mit schwach singulären Koeffizientenfunktionen. Teil I. Eigenwertproblem. (German) [Weak singular Sturm–Liouville problems. I. Eigenvalue problem]. Numerische Mathematik, 27(4): 471–484, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Haverkamp:1984:ASG

[Hav84] Richard Haverkamp. Eine Aussage zur L_{∞}-Stabilität und zur genauen Konvergenzordnung der H_{0}^{1}-Projektionen. (German) [A statement concerning L_{∞}- stability and the exact order of convergence of H_{0}^{1}-projections]. Numerische Mathematik, 44(3): 393–405, September 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hayden:1965:CFA

[Hay65] T. L. Hayden. Continued fraction approximation to functions. Numerische Mathematik, 7(4): 292–309, August 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1959:CME

[HB59] Alston S. Householder and Friedrich L. Bauer. On certain methods for expanding the characteristic polynomial. Numerische Mathematik, 1:29– 37, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1960:CIM

[HB60] A. S. Householder and F. L.

Bauer. On certain iterative methods for solving linear systems. Numerische Mathematik, 2(1):55–59, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Han:1997:ABC

[HB97] H. Han and W. Bao. The artificial boundary conditions for incompressible materials on an unbounded domain. Numerische Mathematik, 77(3):347–363, September 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/

                                        00211/bibs/7077003/70770347.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/7077003/70770347.
                                        htm.

Householder:1963:SVI

[HC63] A. S. Householder and John A. Carpenter. The singular values of involutory and of idempotent matrices. Numerische Mathematik, 5(1):234–237, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Habbal:1992:DFE

[HC92] A. Habbal and D. Chenais. Deterioration of a finite element method for arch structures when the thickness goes to zero. Numerische Mathematik, 62(3): 321–341, August 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hutchinson:1985:SND

[HdH85] M. F. Hutchinson and F. R. de Hoog. Smoothing noisy data with spline functions. Numerische Mathematik, 47(1):99– 106, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Dai:1999:GCM

[hDxY99] Yu hong Dai and Ya xiang Yuan. Global convergence of the method of shortest residuals. Numerische Mathematik, 83(4):581–598, October 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083004/90830581.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083004/90830581.pdf. See

comments [Pyt02].

He:1994:SCL

[He94] Bingsheng He. Solving a class of linear projection equations. Numerische Mathematik, 68(1): 71–80, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680071.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680071.
htm.

He:1997:DUD

[He97] C. He. On the distance to uncontrollability and the distance to instability and their relation to some condition numbers in control. Numerische Mathematik, 76(4): 463–477, June 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076004/70760463.
htm;   http://science.springer.
de/nmee/bibs/7076004/70760463.
htm.

Heap:1966:RMG

[Hea66] B. R. Heap. Random matrices and graphs. Numerische Mathematik, 8(2):114–122, April 1966.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hebeker:1988:NTV

[Heb88] F. K. Hebeker. On the numerical treatment of viscous flows against bodies with corners and edges by boundary element and multigrid methods. Numerische Mathematik, 52(1):81–99, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hedstrom:1965:NSL

[Hed65] G. W. Hedstrom. The nearstability of the Lax–Wendroff method. Numerische Mathematik, 7(1):73–77, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heesterman:1968:SSA

[Hee68] A. R. G. Heesterman. Special simplex algorithm for multi-sector problems. Numerische Mathematik, 12(4): 288–306, November 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hegland:1991:PST

[Heg91] Markus Hegland. On the parallel solution of tridiagonal systems by wrap-around partitioning and incomplete LU factorization. Numerische Mathematik, 59(5):453–472, August 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hegland:1994:SSP

[Heg94] Markus Hegland. A self-sorting in-place fast Fourier transform algorithm suitable for vector and parallel processing. Numerische Mathematik, 68(4): 507–547, October 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4068004/
40680507.htm;   http://science.
springer.de/nmee/bibs/4068004/
40680507.htm.

Heinrichs:1989:KKE

[Hei89a] Wilhelm Heinrichs. Konvergenzaussagen für Kollokationsverfahren bei elliptischen Randwertaufgaben. (German) [Convergence of the collocation method for elliptic boundary value problems]. Numerische Mathematik, 54(6):619–637, March 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heinrichs:1989:SMS

[Hei89b] Wilhelm Heinrichs. Spectral methods with sparse matrices. Numerische Mathematik, 56(1): 25–41, September 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hekrdla:1987:ITD

[Hek87] Josef Hekrdla. Index transforms for N-dimensional DFT’s.

Numerische Mathematik, 51(4): 469–480, August 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heller:1968:UPC

[Hel68] Sidney Heller. On the uniform positivity of a certain finite difference matrix. Numerische Mathematik, 11(1):46–52, January 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Helfrich:1987:EES

[Hel87] Hans-Peter Helfrich. Error estimates for semidiscrete Galerkin type approximations to semilinear evolution equations with nonsmooth initial data. Numerische Mathematik, 51(5): 559–569, October 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hemmingsson:1998:SPC

[Hem98] Lina Hemmingsson. A semicirculant preconditioner for the convection-diffusion equation. Numerische Mathematik, 81(2): 211–248, December 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8081002/
80810211.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8081002/
80810211.pdf.

Henrici:1962:BII

[Hen62] Peter Henrici. Bounds for iterates, inverses, spectral variation and fields of values of nonnormal matrices. Numerische Mathematik, 4:24–39, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Henrich:1975:NES

[Hen75] C. J. Henrich. On the numerical evaluation of a sum involving binomial coefficients. Numerische Mathematik, 24(1):81– 83, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Henrici:1979:BFI

[Hen79] Peter Henrici. Barycentric formulas for interpolating trigonometric polynomials and their conjugates. Numerische Mathematik, 33(2):225–234, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hertling:1970:APC

[Her70]  Jörg Hertling.  Approxima-

tion of piecewise continuous functions by a modification of piecewise Hermite interpolation. Numerische Mathematik, 15(5): 404–414, September 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hertling:1971:NTS

[Her71] Jörg Hertling. Numerical treatment of singular integral equations by interpolation methods. Numerische Mathematik, 18(2): 101–112, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Herz:1974:ZTA

[Her74] B. Herz. Zerfallende Tridiagonalematrizen und aneinandergesetzte Sturmsche Ketten. (German) [Decomposable tridiagonal matrices and contiguous Sturm sequences]. Numerische Mathematik, 23(1):95–98, February 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Herceg:1990:UFO

[Her90] Dragoslav Herceg. Uniform fourth order difference scheme for a singular perturbation problem. Numerische Mathematik, 56(7):675–693, January 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

HernandezVeron:1991:NHM

[Her91] Miguel A. Hernández Verón. A note on Halley’s method. Numerische Mathematik, 59(3): 273–276, June 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hettich:1974:KZO

[Het74] R. Hettich. Kriterien zweiter Ordnung für lokal beste Approximationen. (German) [Secondorder conditions for locally best approximations]. Numerische Mathematik, 22:409–417, October 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hettich:1975:KEZ

[Het75] R. Hettich. Kriterien erster und zweiter Ordnung für lokal beste Approximationen bei Problemen mit Nebenbedingungen. (German) [On first and second order conditions for locally best approximations for constraint problems]. Numerische Mathematik, 25(1):109–122, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heuer:1998:ISM

[Heu98]  Norbert Heuer.  An itera-

tive substructuring method for the p-version of the boundary element method for hypersingular integral operators in three dimensions. Numerische Mathematik, 79(3): 371–396, May 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079003/80790371.
htm;   http://science.springer.
de/nmee/bibs/8079003/80790371.

htm. See erratum [Heu01].

Heuer:2001:EIS

[Heu01] Norbert Heuer. Erratum: “An iterative substructuring method for the p-version of the boundary element method for hypersingular integral operators in three dimensions” Numer. Math. (1998) 79: 371–396. Numerische Mathematik, 87(4):793–794, February 2001. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        1087004/10870793.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        1087004/10870793.pdf. See

[Heu98].

Haslinger:1976:MFE

[HH76] J. Haslinger and I. Hlaváček. A mixed finite element method close to the equilibrium model. Numerische Mathematik, 26(1): 85–97, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hiptmair:1999:MMM

[HH99] Ralf Hiptmair and Ronald H. W. Hoppe. Multilevel methods for mixed finite elements in three dimensions. Numerische Mathematik, 82(2):253–279, April 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9082002/90820253.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9082002/90820253.pdf.

Hallet:1976:GMM

[HHM76] P. Hallet, J. P. Hennart, and E. H. Mund. A Galerkin method with modified piecewise polynomials for solving a secondorder boundary value problem. Numerische Mathematik, 27(1): 11–20, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Huang:1994:RMO

[HHZ94] Hongci Huang, Weimin Han, and Jinshi Zhou. The regularization method for an obstacle problem. Numerische Mathematik, 69(2):155–166, December 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069002/40690155.
htm;   http://science.springer.
de/nmee/bibs/4069002/40690155.
htm.

Hoshino:1971:SPD

[HI71] S. Hoshino and K. Ichida. Solution of partial differential equations by a modified random walk. Numerische Mathematik, 18(1):61–72, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Higham:1987:EAB

[Hig87] Nicholas J. Higham. Error analysis of the Björck-Pereyra algorithms for solving Vandermonde systems. Numerische Mathematik, 50(5):613–632, March 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Higham:1992:EMN

[Hig92] Nicholas J. Higham. Estimating the matrix p-norm. Numerische Mathematik, 62(4): 539–555, September 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hilali:1983:RSD

[Hil83] Abdelaziz Hilali. Réductibilité d’un système différentiel linéaire. (French) [Reducibility of a linear differential system]. Numerische Mathematik, 41(1):1–17, April 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hinze:1996:NAU

[Hin96] Michael Hinze. On the numerical approximation of unstable minimal surfaces with polygonal boundaries. Numerische Mathematik, 73(1):95– 118, March 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/6073001/60730095.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/6073001/60730095.
                                        htm.

Hairer:1990:ERK

[HISS90] E. Hairer, A. Iserles, and J. M. Sanz-Serna. Equilibria of Runge–Kutta methods. Numerische Mathematik, 58(3): 243–254, October 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heinig:1990:PSA

[HJ90] G. Heinig and P. Jankowski.

Parallel and superfast algorithms for Hankel systems of equations. Numerische Mathematik, 58(1):109–127, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heinig:1988:FIA

[HJR88a] G. Heinig, P. Jankowski, and K. Rost. Fast inversion algorithms of Toeplitz-plus-Hankel matrices. Numerische Mathematik, 52(6):665–682, June 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hennart:1988:CMD

[HJR88b] Jean-Pierre Hennart, Jérôme Jaffre, and Jean Elizabeth Roberts. A constructive method for deriving finite elements of nodal type. Numerische Mathematik, 53(6):701–738, September 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hull:1960:NME

[HL60] T. E. Hull and W. A. J. Luxemburg. Numerical methods and existence theorems for ordinary differential equations. Numerische Mathematik, 2(1):30– 41, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heap:1964:GTA

[HL64a] B. R. Heap and M. S. Lynn. A graph-theoretic algorithm for the solution of a linear Diophantine problem of Frobenius. Numerische Mathematik, 6(1):346– 354, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heap:1964:IPN

[HL64b] B. R. Heap and M. S. Lynn. The index of primitivity of a non-negative matrix. Numerische Mathematik, 6(1):120– 141, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heap:1965:LDP

[HL65] B. R. Heap and M. S. Lynn. On a linear Diophantine problem of Frobenius: an improved algorithm. Numerische Mathematik, 7(3):226–231, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1984:AEG

[HL84] E. Hairer and C. Lubich. Asymptotic expansions of the global error of fixed-stepsize methods. Numerische Mathematik, 45(3):345–360, December 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1988:EAS

[HL88] E. Hairer and C. Lubich. Extrapolation at stiff differential equations. Numerische Mathematik, 52(4):377–400, April 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1997:LSB

[HL97] E. Hairer and Ch. Lubich. The life-span of backward error analysis for numerical integrators. Numerische Mathematik, 76(4): 441–462, June 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076004/70760441.
htm;   http://science.springer.
de/nmee/bibs/7076004/70760441.
htm.

He:1998:CSF

[HL98] Yinnian He and Kaitai Li. Convergence and stability of finite element nonlinear Galerkin method for the Navier–Stokes equations. Numerische Mathematik, 79(1):77–106, March 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.de/link/service/
journals/00211/bibs/8079001/
80790077.htm;   http://science.
springer.de/nmee/bibs/8079001/
80790077.htm.

Hochbruck:1999:GTM

[HL99] Marlis Hochbruck and Christian Lubich. A Gautschitype method for oscillatory second-order differential equations. Numerische Mathematik, 83(3):403–426, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083003/90830403.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083003/90830403.pdf.

Hackbusch:1983:MGM

[HM83] W. Hackbusch and H. D. Mittelmann. On multi-grid methods for variational inequalities. Numerische Mathematik, 42(1): 65–76, September 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hegland:1986:NCL

[HM86] M. Hegland and J. T. Marti. Numerical computation of least constants for the Sobolev inequality. Numerische Mathematik, 48(6):607–616, June 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hoffmann:1989:LSM

[HM89] K.-H. Hoffmann and G. H. Meyer. A least squares method for finding the Preisach hysteresis operator from measurements. Numerische Mathematik, 55(6): 695–710, August 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Heuer:1999:ECH

[HMS99] Norbert Heuer, Matthias Maischak, and Ernst P. Stephan. Exponential convergence of the hp-version for the boundary element method on open surfaces. Numerische Mathematik, 83(4):641–666, October 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083004/90830641.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083004/90830641.pdf.

Houston:1999:PEA

[HMSW99] P. Houston, J. A. Mackenzie, E. Süli, and G. Warnecke. A posteriori error analysis for numerical approximations of Friedrichs systems. Numerische Mathematik, 82(3):433–470, May 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082003/90820433.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082003/90820433.pdf.

Huy:1986:FDA

[HMW86] C. U. Huy, P. J. McKenna, and W. Walter. Finite difference approximations to the Dirichlet problem for elliptic systems. Numerische Mathematik, 49(2/3): 227–237, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:1989:FMM

[HN89] W. Hackbusch and Z. P. Nowak. On the fast matrix multiplication in the boundary element method by panel clustering. Numerische Mathematik, 54 (4):463–491, January 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hanke:1990:PSR

[HN90] Martin Hanke and Michael Neumann. Preconditionings and splittings for rectangular systems. Numerische Mathematik, 57(1):85–95, March 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hu:1992:CTA

[HN92] Xiaohua Hu and R. A. Nicolaides. Covolume techniques for anisotropic media. Numerische Mathematik, 61(2): 215–234, March 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hanke:1995:CAL

[HNS95] Martin Hanke, Andreas Neubauer, and Otmar Scherzer. A convergence analysis of the Landweber iteration for nonlinear ill-posed problems. Numerische Mathematik, 72(1):21– 37, November 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5072001/50720021.
htm;   http://science.springer.
de/nmee/bibs/5072001/50720021.
htm.

Hairer:1990:DOE

[HO90a] E. Hairer and A. Ostermann.

Dense output for extrapolation methods. Numerische Mathematik, 58(4):419–439, November 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ho:1990:TAT

[Ho90b]  Diem Ho.  Tchebychev ac-

celeration technique for large scale nonsymmetric matrices. Numerische Mathematik, 56(7): 721–734, January 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hock:1979:AEM

[Hoc79] Wilhelm Hock. Asymptotic expansions for multistep methods applied to nonlinear Volterra integral equations of the second kind. Numerische Mathematik, 33(1):77–100, March 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hock:1981:EMS

[Hoc81] Wilhelm Hock. An extrapolation method with step size control for nonlinear Volterra integral equations. Numerische Mathematik, 38(2):155–178, December 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hodgkins:1967:RBD

[Hod67] W. R. Hodgkins. On the relation between dynamic relaxation and semi-iterative matrix methods. Numerische Mathematik, 9 (5):446–451, April 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hofmann:1967:AED

[Hof67] Peter Hofmann. Asymptotic expansions of the discretization error of boundary value problems of the Laplace equation in rectangular domains. Numerische Mathematik, 9(4):302– 322, February 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hoffman:1969:TNT

[Hof69] Karl-Heinz Hoffman. Zur Theorie der nichtlinearen Tschebyscheff-Approximation mit Nebenbedingungen. (German) [Toward a theory of nonlinear Chebyshev approximation with constraints]. Numerische Mathematik, 14(1):24–41, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hoffmann:1970:NTA

[Hof70] Karl-Heinz Hoffmann. Nichtlineare Tschebyscheff-Approximation mit Nebenbedingungen: Anwendungen. (German) [Nonlinear Chebyshev approximation with constraints: Use]. Numerische Mathematik, 14(4): 383–393, March 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hofmann:1975:MHA

[Hof75] Wolf Hofmann. Monotoniesätze für hyperbolische Anfangswertaufgaben und Einschließung von Lösungen. (German) [Monotonicity theorems for hyperbolic initial value problems and solution enclosure]. Numerische Mathematik, 24(2): 137–149, April 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hohn:1982:FEP

[Höh82] W. Höhn. Finite elements for parabolic equations backwards in time. Numerische Mathematik, 40(2):207–227, December 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Homeier:1995:DRJ

[Hom95] Herbert H. H. Homeier. Determinantal representations for the J transformation. Numerische Mathematik, 71(3):275–288, September 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071003/50710275.
htm;   http://science.springer.
de/nmee/bibs/5071003/50710275.
htm.

Hoppe:1981:PMA

[Hop81] Ronald H. W. Hoppe. A penalty method for the approximate solution of stationary Maxwell equations. Numerische Mathematik, 36(4):389–403, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hoppe:1986:MMH

[Hop86] Ronald H. W. Hoppe. Multigrid methods for Hamilton– Jacobi–Bellman equations. Numerische Mathematik, 49(2/3): 239–254, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hornung:1975:EMN

[Hor75] U. Hornung. Euler method for neutral functional-differential equation. Numerische Mathematik, 24(3):233–240, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1966:STN

[Hou66] A. S. Householder. Separation theorems for normalizable matrices. Numerische Mathematik, 9(1):46–50, November 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1968:MCR

[Hou68] A. S. Householder. Moments and characteristic roots. II. Numerische Mathematik, 11(2):

126–128, February 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1971:GAS

[Hou71]  A. S. Householder.  Gen-

eralizations of an algorithm of Sebastião e Silva. Numerische Mathematik, 16(4): 375–382, January 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1973:PGA

[Hou73] A. S. Householder. Postscript to: “Generalizations of an algorithm of Sebastiao e Silva”. Numerische Mathematik, 20(3): 205–207, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Henrici:1966:TEE

[HP66] Peter Henrici and Pia Pflüger.

Truncation error estimates for Stieltjes fractions. Numerische Mathematik, 9(2):120– 138, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hanson:1975:ANM

[HP75] Richard J. Hanson and James L.

Phillips. An adaptive numerical method for solving linear Fredholm integral equations of the first kind. Numerische Mathematik, 24(4):291–307, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Haegemans:1976:CCF

[HP76a] Ann Haegemans and Robert Piessens. Construction of cubature formulas of degree eleven for symmetric planar regions, using orthogonal polynomials. Numerische Mathematik, 25(2): 139–148, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hansen:1976:EMR

[HP76b] Eldon Hansen and Merrell Patrick. Estimating the multiplicity of a root. Numerische Mathematik, 27(1): 121–131, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hansen:1977:FRF

[HP77] Eldon Hansen and Merrell Patrick. A family of root finding methods. Numerische Mathematik, 27(3):257–269, September 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hough:1981:USS

[HP81] D. M. Hough and N. Papamichael. The use of splines and singular functions in an integral equation method for conformal mapping. Numerische Mathematik, 37(1): 133–147, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hough:1983:IEM

[HP83] D. M. Hough and N. Papamichael. An integral equation method for the numerical conformal mapping of interior, exterior and doubly-connected domains. Numerische Mathematik, 41(3):287–307, July 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hinrichsen:1992:SVU

[HP92] D. Hinrichsen and A. J.

Pritchard. On spectral variations under bounded real matrix perturbations. Numerische Mathematik, 60(4): 509–524, January 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hadjidimos:1994:OCS

[HP94] Apostolos Hadjidimos and Robert J. Plemmons. Optimal p-cyclic SOR. Numerische Mathematik, 67(4): 475–490, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067004/40670475.
htm;   http://science.springer.
de/nmee/bibs/4067004/40670475.
htm.

Hopkins:1979:KMD

[HR79] T. R. Hopkins and D. E.

Roberts. Kufarev’s method for determining the Schwarz– Christoffel parameters. Numerische Mathematik, 33(4): 353–365, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Halpern:1987:EAA

[HR87] Laurence Halpern and Jeffrey Rauch. Error analysis for absorbing boundary conditions. Numerische Mathematik, 51(4): 459–467, August 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:1989:ADN

[HR89] W. Hackbusch and A. Reusken. Analysis of a damped nonlinear multilevel method. Numerische Mathematik, 55(2): 225–246, May 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Haviv:1994:BEA

[HR94] Moshe Haviv and Ya’acov Ritov. Bounds on the error of an approximate invariant subspace for non-self-adjoint matrices. Numerische Mathematik, 67 (4):491–500, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/4067004/40670491.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4067004/40670491.
                                        htm.

Halpern:1995:ABC

[HR95] L. Halpern and J. Rauch. Absorbing boundary conditions for diffusion equations. Numerische Mathematik, 71(2):

185–224, August 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071002/50710185.
htm;   http://science.springer.
de/nmee/bibs/5071002/50710185.
htm.

Householder:1969:CSI

[HS69] A. S. Householder and G. W.

Stewart. Comments on “Some iterations for factoring a polynomial”. Numerische Mathematik, 13(5):470–471, October 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Howland:1970:CMS

[HS70] James Lucien Howland and John Albert Senez. A constructive method for the solution of the stability problem. Numerische Mathematik, 16(1): 1–7, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hammerlin:1979:EBA

[HS79] Günther Hämmerlin and Larry L.

Schumaker. Error bounds for the approximation of Green’s kernels by splines. Numerische Mathematik, 33(1): 17–22, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hammerlin:1980:PKA

[HS80] Günther Hämmerlin and Larry L.

Schumaker. Procedures for kernel approximation and solution of Fredholm integral equations of the second kind. Numerische Mathematik, 34(2): 125–141, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hundsdorfer:1981:NSR

[HS81] W. H. Hundsdorfer and M. N. Spijker. A note on B-stability of Runge–Kutta methods. Numerische Mathematik, 36(3): 319–331, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hariharan:1983:BEM

[HS83] S. I. Hariharan and E. Stephan. A boundary element method for a two-dimensional interface problem in electromagnetics. Numerische Mathematik, 42(3):311–322, December 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hu:1986:CSS

[HS86] Chiu Li Hu and Larry L. Schumaker. Complete spline smoothing. Numerische Mathematik, 49(1):1–10, January 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hundsdorfer:1990:ACC

[HS90] Willem H. Hundsdorfer and Josef Schneid. An algebraic characterization of Bconvergent Runge–Kutta methods. Numerische Mathematik, 56(7):695–705, January 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hersch:1991:NIC

[HS91] Joseph Hersch and William Sawyer. Numerical implementation of coherence for the example of the “Swiss cross”. Numerische Mathematik, 59(7): 659–665, October 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

He:1992:SPP

[HS92] Bingsheng He and Josef Stoer.

Solution of projection problems over polytopes. Numerische Mathematik, 61(1):73– 90, February 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hackbusch:1997:CFE

[HS97] W. Hackbusch and S. A. Sauter.

Composite finite elements for the approximation of PDEs on domains with complicated micro-structures. Numerische Mathematik, 75(4):447–472, February 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7075004/70750447.
htm;   http://science.springer.
de/nmee/bibs/7075004/70750447.
htm.

Heil:1996:ATR

[HSS96] Christopher Heil, Gilbert Strang, and Vasily Strela. Approximation by translates of refinable functions. Numerische Mathematik, 73(1): 75–94, March 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/6073001/60730075.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/6073001/60730075.
                                        htm.

Hsu:1961:NNI

[Hsu61] L. C. Hsu. Note on the numerical integration of periodic functions and of partially periodic functions. Numerische Mathematik, 3:169–173, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hsu:1962:NIN

[Hsu62] L. C. Hsu. On the numerical integration of non-periodic functions of several variables. Numerische Mathematik, 4:330– 335, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Helton:1991:TPR

[HSY91] J. W. Helton, P. G. Spain, and N. J. Young. Tracking poles and representing Hankel operators directly from data. Numerische Mathematik, 58(6): 641–660, January 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hou:1995:AFE

[HT95] L. Steven Hou and James C.

Turner. Analysis and finite element approximation of an optimal control problem in electrochemistry with current density controls. Numerische Mathematik, 71(3):289–315, September 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071003/50710289.
htm;   http://science.springer.
de/nmee/bibs/5071003/50710289.
htm.

Huang:1990:MGA

[Hua90] Ziping Huang. A multi-grid algorithm for mixed problems with penalty. Numerische Mathematik, 57(3):227–247, May 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hubner:1972:ZUG

[Hüb72] Otto Hübner. Zweiparametrige Überrelaxation. (German) [Two parameter overrelaxation]. Numerische Mathematik, 18(4): 354–366, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hubner:1982:ALD

[Hüb82] Otto Hübner. Über die Anzahl der Lösungen der diskreten Theodorsen-Gleichung. (German) [On the number of solutions of the discrete Theodorsen equation]. Numerische Mathematik, 39(2):195–204, August 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Huckle:1998:SGE

[Huc98]  Thomas Huckle.  Symmet-

ric Gaussian elimination for Cauchy-type matrices with application to positive definite Toeplitz matrices. Numerische Mathematik, 79(2): 213–229, April 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079002/80790213.
htm;   http://science.springer.
de/nmee/bibs/8079002/80790213.
htm.

Hulme:1971:PPT

[Hul71] Bernie L. Hulme. Piecewise polynomial Taylor methods for initial value problems. Numerische Mathematik, 17(5): 367–381, October 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hunter:1972:SGF

[Hun72] D. B. Hunter. Some Gausstype formulae for the evaluation of Cauchy principal values of integrals. Numerische Mathematik, 19(5):419–424, October 1972. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Hunter:1973:NEC

[Hun73] D. B. Hunter. The numerical evaluation of Cauchy principal values of integrals by Romberg integration. Numerische Mathematik, 21(3):185–192, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hundsdorfer:1986:SCL

[Hun86] W. H. Hundsdorfer. Stability and B-convergence of linearly implicit Runge–Kutta methods. Numerische Mathematik, 50(1): 83–95, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hiriart-Urruty:1995:SAA

[HUY95] J.-B. Hiriart-Urruty and D. Ye.

Sensitivity analysis of all eigenvalues of a symmetric matrix. Numerische Mathematik, 70(1): 45–72, March 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070001/50700045.
htm;   http://science.springer.
de/nmee/bibs/5070001/50700045.
htm.

Hageman:1964:BIM

[HV64] L. A. Hageman and R. S.

Varga. Block iterative methods for cyclically reduced matrix equations. Numerische Mathematik, 6(1):106–119, December 1964. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Howland:1985:ACI

[HV85] James Lucien Howland and Rémi Vaillancourt. Attractive cycles in the iteration of meromorphic functions. Numerische Mathematik, 46(3): 323–337, June 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hanke:1999:TLP

[HV99] Martin Hanke and Curtis R. Vogel. Two-level preconditioners for regularized inverse problems I: Theory. Numerische Mathematik, 83(3):385–402, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083003/90830385.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083003/90830385.pdf.

Hofsommer:1963:NCE

[HvdR63] D. J. Hofsommer and R. van de Riet. On the numerical calculation of elliptic integrals of the first and second kind and the elliptic functions of Jacobi. Numerische Mathematik, 5(1): 291–302, December 1963. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Householder:1970:NGI

[HVW70] Alston S. Householder, Richard S.

Varga, and James H. Wilkinson. A note on Gerschgorin’s inclusion theorem for eigenvalues of matrices. Numerische Mathematik, 16(2):141–144, November 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hairer:1976:TNM

[HW76] E. Hairer and G. Wanner. A theory for Nyström methods. Numerische Mathematik, 25(4): 383–400, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Henderson:1978:NMS

[HW78] D. S. Henderson and A. Wassyng.

A new method for the solution of Ax=b. Numerische Mathematik, 29(3):287–289, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hanson:1984:ATG

[HW84] P. M. Hanson and J. E. Walsh.

Asymptotic theory of the global error and some techniques of error estimation. Numerische Mathematik, 45(1):51–74, October 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hilali:1987:FSI

[HW87] A. Hilali and A. Wazner. Formes super-irréductibles des systèmes différentiels linéaires. (French) [Superirreducible forms of linear differential systems]. Numerische Mathematik, 50(4): 429–449, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hwang:1999:RAO

[Hwa99] J. S. Hwang. Rational approximation to orthogonal bases and of the solutions of elliptic equations. Numerische Mathematik, 81(4):561–575, February 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9081004/90810561.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9081004/90810561.pdf.

Iusem:1986:CRA

[ID86] Alfredo N. Iusem and Alvaro Rodolfo De Pierro. Convergence results for an accelerated nonlinear Cimmino algorithm. Numerische Mathematik, 49(4):367–378, July 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Iyengar:1987:SFD

[IJ87] S. R. K. Iyengar and Pragya Jain. Spline finite difference methods for singular two point boundary value problems. Numerische Mathematik, 50(3): 363–376, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ifantis:1977:LBZ

[IK77] Evangelos K. Ifantis and Christos B. Kouris. Lower bounds for the zeros of analytic functions. Numerische Mathematik, 27(2):239–247, June 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ito:1989:AID

[IK89] K. Ito and F. Kappel. Approximation of infinite delay and Volterra type equations. Numerische Mathematik, 54(4): 405–444, January 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ichim:1994:BTM

[IM94] Ion Ichim and Ionel Molnar.

A Bairstow’s type method for trigonometric polynomials. Numerische Mathematik, 67(2): 251–259, March 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067002/40670251.
htm;   http://science.springer.
de/nmee/bibs/4067002/40670251.
htm.

Imhof:1963:MNI

[Imh63] J. P. Imhof. On the method for numerical integration of Clenshaw and Curtis. Numerische Mathematik, 5(1):138– 141, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

IntHout:1991:SAN

[iS91] K. J. in ’t Hout and M. N. Spijker. Stability analysis of numerical methods for delay differential equations. Numerische Mathematik, 59(8):807–814, October 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Iske:1996:SFS

[IS96] Armin Iske and Thomas Sonar. On the structure of function spaces in optimal recovery of point functionals for ENOschemes by radial basis functions. Numerische Mathematik, 74(2):177–201, August 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/6074002/
                                        60740177.htm;   http://science.
                                        springer.de/nmee/bibs/6074002/
                                        60740177.htm.

Iserles:1984:SLO

[Ise84] A. Iserles. Solving linear ordinary differential equations by exponentials of iterated commutators. Numerische Mathematik, 45(2):183–199, November 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ishihara:1979:MFE

[Ish79] Kazuo Ishihara. On the mixed finite element approximation for the buckling of plates. Numerische Mathematik, 33(2): 195–210, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ishihara:1981:FEL

[Ish81] Kazuo Ishihara. A finite element lumped mass scheme for solving eigenvalue problems of circular arches. Numerische Mathematik, 36(3):267–290, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ishihara:1984:MEI

[Ish84a] Kazuo Ishihara. Monotone explicit iterations of the finite element approximations for the nonlinear boundary value problem. Numerische Mathematik, 43(3):419–437, April 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ishihara:1984:NAB

[Ish84b] Kazuo Ishihara. On numerical asymptotic behavior of finite element solutions for parabolic equations. Numerische Mathematik, 44(2):285–300, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ishihara:1985:FEL

[Ish85] Kazuo Ishihara. A finite element lumped mass scheme for solving eigenvalue problems of circular arches. II. numerical experiments with Double precision arithmetic. Numerische Mathematik, 46(4):499–504, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Israeli:1982:INS

[IU82] Moshe Israeli and Marius Ungarish. Improvement of numerical solution of boundary layer problems by incorporation of asymptotic approximations. Numerische Mathematik, 39(3): 309–324, October 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jackiewicz:1979:CMM

[Jac79] Zdzisław Jackiewicz. Convergence of multistep methods for Volterra functional differential equations. Numerische Mathematik, 32(3):307–332, September 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jackiewicz:1982:AMN

[Jac82] Zdzisław Jackiewicz. Adams methods for neutral functionaldifferential equations. Numerische Mathematik, 39(2): 221–230, August 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jackiewicz:1984:ASA

[Jac84] Zdzisław Jackiewicz. Asymptotic stability analysis of θmethods for functional differential equations. Numerische Mathematik, 43(3):389–396, April 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jacobsen:1985:RML

[Jac85] Lisa Jacobsen. Repeated modifications of limit k-periodic continued fractions. Numerische Mathematik, 47(4): 577–595, December 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jamet:1970:CFD

[Jam70] Pierre Jamet. On the convergence of finite-difference approximations to one-dimensional singular boundary-value problems. Numerische Mathematik, 14(4): 355–378, March 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jarre:1994:NLS

[Jar94] Florian Jarre. A new line-search step based on the Weierstrass ℘- function for minimizing a class of logarithmic barrier functions. Numerische Mathematik, 68(1): 81–94, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680081.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680081.
htm.

Jay:1993:CMD

[Jay93] Laurent Jay. Collocation methods for differential-algebraic equations of index 3. Numerische Mathematik, 65(4): 407–421, August 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jai:1992:LPD

[JBC92] M. Jai, G. Bayada, and M. Chambat. The L. P. D. E. M.: a mixed finite difference method for elliptic problems. Numerische Mathematik, 63(2):195–211, November 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jeggle:1968:DKF

[Jeg68] Hansgeorg Jeggle. Zur Diskretisierung einer Klasse Fredholmscher Integralgleichungen mit singulãrem Wert der Ordnung 1. (German) [on the discretisation of a class of Fredholm integral equations with singular values of order 1]. Numerische Mathematik, 11(5): 427–443, June 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jeltsch:1973:III

[Jel73] Rolf Jeltsch. Integration of iterated integrals by multistep methods. Numerische Mathematik, 21(4):303–316, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jeltsch:1979:SSS

[Jel79] Rolf Jeltsch. A_{0}-stability and stiff stability of brown’s multistep multiderivative methods.

Numerische Mathematik, 32(2): 167–181, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jensen:1991:CPV

[Jen91] Soren Jensen. On computing the pressure by the p version of the finite element method for Stokes problem. Numerische Mathematik, 59(6):581–601, September 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jerome:1970:LSM

[Jer70] Joseph W. Jerome. Linear self-adjoint multipoint boundary value problems and related approximation schemes. Numerische Mathematik, 15(5): 433–449, September 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jerome:1985:ANA

[Jer85] Joseph W. Jerome. An adaptive Newton algorithm based on numerical inversion: regularization as postconditioner. Numerische Mathematik, 47(1):123–138, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jerome:1989:NMG

[Jer89] Joseph W. Jerome. Newton’s method for gradient equations based upon the fixed point map: convergence and complexity study. Numerische Mathematik, 55(6):619–632, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jetter:1976:OQS

[Jet76] Kurt Jetter. Optimale Quadraturformeln mit semidefiniten Peano-Kernen. (German) [Optimal integration formulas with semidefinite Peano-kernels]. Numerische Mathematik, 25(3): 239–249, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jetter:1984:KAR

[Jet84] Kurt Jetter. Eine kurze Anmerkung zur Romberg-Integration. (German) [A short note on Romberg integration]. Numerische Mathematik, 45(2): 275–281, November 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

JHW:1977:AHA

[JHW77] JHW. Alston S. Householder award (1977)—final notice. Numerische Mathematik, 28(1): 120, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jiang:1994:ORP

[Jia94]  Hong Jiang.  On the or-

thogonality of residual polynomials of minimax polynomial preconditioning. Numerische Mathematik, 67(3): 345–364, April 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067003/40670345.
htm;   http://science.springer.
de/nmee/bibs/4067003/40670345.
htm.

Jia:1998:GBL

[Jia98] Zhongxiao Jia. Generalized block Lanczos methods for large unsymmetric eigenproblems. Numerische Mathematik, 80(2): 239–266, August 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080002/80800239.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080002/80800239.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080002/80800239.
ps.gz.

Jeltsch:1978:SPB

[JK78] Rolf Jeltsch and L. Kratz.

On the stability properties of Brown’s multistep multiderivative methods. Numerische Mathematik, 30(1): 25–38, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jager:1991:SPM

[JK91a] Willi Jäger and Jozef Kačur. Solution of porous medium type systems by linear approximation schemes. Numerische Mathematik, 60(3):407–427, December 1991. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Jeltsch:1991:SFM

[JK91b] Rolf Jeltsch and Peywand Kiani. Stability of a family of multitime-level difference schemes for the advection equation. Numerische Mathematik, 60(1):77– 95, November 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jeltsch:1988:CSB

[JKR88] Rolf Jeltsch, Peywand Kiani, and Klaus Raczek. Counterexamples to a stability barrier. Numerische Mathematik, 52(3): 301–316, March 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jamet:1970:MRN

[JLR70] P. Jamet, P. Lascaux, and P.-A. Raviart. Une méthode de résolution numérique des équations de Navier–Stokes. (French) [A method of numerical solution of the Navier–Stokes equations]. Numerische Mathematik, 16(2):93–114, November 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Johnson:1978:SEF

[JM78] C. Johnson and B. Mercier. Some equilibrium finite element methods for two-dimensional elasticity problems. Numerische Mathematik, 30(1): 103–116, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jarausch:1987:SLN

[JM87] Helmut Jarausch and Wolfgang Mackens. Solving large nonlinear systems of equations by an adaptive condensation process. Numerische Mathematik, 50(6):633–653, April 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

John:1997:NSD

[JMT97] V. John, J. M. Maubach, and L. Tobiska. Nonconforming streamline-diffusion-finite-element-methods for convectiondiffusion problems. Numerische Mathematik, 78(2):165–188, December 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078002/70780165.
htm;   http://science.springer.
de/nmee/bibs/7078002/70780165.
htm.

Jeltsch:1981:SET

[JN81] Rolf Jeltsch and Olavi Nevanlinna. Stability of explicit time discretizations for solving initial value problems. Numerische Mathematik, 37(1):61–91, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [JN82a].

Jeltsch:1982:EET

[JN82a] R. Jeltsch and O. Nevanlinna. Erratum: “Stability of explicit time discretizations for solving initial value problems” [Numer. Math. 37 (1981), no. 1, 61–91 and MR 82g:65042]. Numerische Mathematik, 39(1):155, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [JN81].

Jeltsch:1982:SAT

[JN82b] Rolf Jeltsch and Olavi Nevanlinna. Stability and accuracy of time discretizations for initial value problems. Numerische Mathematik, 40(2): 245–296, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jeltsch:1986:ASM

[JN86] Rolf Jeltsch and Olavi Nevanlinna. Accuracy and stability of multistage multistep formulas. Numerische Mathematik, 48 (1):33–83, January 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jiang:1994:ENI

[JN94] Xun Jiang and Ricardo H. Nochetto. Effect of numerical integration for elliptic obstacle problems. Numerische Mathematik, 67(4): 501–512, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067004/40670501.
htm;   http://science.springer.
de/nmee/bibs/4067004/40670501.
htm.

Johnston:1972:CIR

[JO72] R. L. Johnston and D. D.

Olesky. On the computation of inclusion regions for partitioned matrices. Numerische Mathematik, 19(3):238–247, June 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jennings:1974:DEA

[JO74] L. S. Jennings and M. R. Osborne. A direct error analysis for least squares. Numerische Mathematik, 22(4):325–332, August 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jittorntrum:1980:SUS

[JO80] Krisorn Jittorntrum and M. R.

Osborne. Strong uniqueness and second-order convergence in nonlinear discrete approximation. Numerische Mathematik, 34(4):439–455, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jochum:1980:NSI

[Joc80] Peter Jochum. The numerical solution of the inverse Stefan problem. Numerische Mathematik, 34(4):411–429, December 1980. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Johnson:1969:UAV

[Joh69] Lee W. Johnson. Uniform approximation of vector-valued functions. Numerische Mathematik, 13(3):238–244, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Johnson:1973:CMF

[Joh73] Claes Johnson. On the convergence of a mixed finite-element method for plate bending problems. Numerische Mathematik, 21(1):43–62, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Johnson:1976:FEM

[Joh76] Claes Johnson. On finite element methods for plasticity problems. Numerische Mathematik, 26(1):79–84, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Johnson:1977:LCG

[Joh77] Charles R. Johnson. Locally compatible generalized matrix norms. Numerische Mathematik, 27(4):391–394, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Joubert:1971:EDAb

[Jou71a] G. R. Joubert. Explicit difference approximations of Du Fort–Frankel type for the onedimensional diffusion equation. Numerische Mathematik, 18(1): 18–25, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Joubert:1971:EDAa

[Jou71b] G. R. Joubert. Explicit difference approximations of the onedimensional diffusion equation, using a smoothing technique. Numerische Mathematik, 17(5): 409–430, October 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jovanovic:1989:CFD

[Jov89] Boško S. Jovanović. On the convergence of finite-difference schemes for parabolic equations with variable coefficients. Numerische Mathematik, 54(4): 395–404, January 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

James:1990:ISA

[JP90] Douglas James and Robert J.

Plemmons. An iterative substructuring algorithm for equilibrium equations. Numerische Mathematik, 57(6/7):625–633, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Johnston:1970:CBI

[JS70] R. L. Johnston and B. T. Smith. Calculation of best isolated Gerschgorin discs. Numerische Mathematik, 16(1):22– 31, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Joe:1986:BMS

[JS86] Stephen Joe and Ian H. Sloan. On Bateman’s method for second kind integral equations. Numerische Mathematik, 49(5): 499–510, September 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jetter:1991:ACI

[JS91] Kurt Jetter and Joachim Stöckler. Algorithms for cardinal interpolation using box splines and radial basis functions. Numerische Mathematik, 60(1):97–114, November 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jbilou:1995:ASV

[JS95] K. Jbilou and H. Sadok. Analysis of some vector extrapolation methods for solving systems of linear equations. Numerische Mathematik, 70(1): 73–89, March 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/5070001/50700073.

htm;   http://science.springer.
de/nmee/bibs/5070001/50700073.
htm.

Shang:1999:KTS

[jS99] Zai jiu Shang. KAM theorem of symplectic algorithms for Hamiltonian systems. Numerische Mathematik, 83(3): 477–496, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083003/90830477.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083003/90830477.pdf.

Jenkins:1970:TSV

[JT70] M. A. Jenkins and J. F. Traub.

A three-stage variable-shift iteration for polynomial zeros and its relation to generalized Rayleigh iteration. Numerische Mathematik, 14(3): 252–263, February 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jones:1976:TEA

[JT76] William B. Jones and W. J.

Thron. Truncation error analysis by means of approximant systems and inclusion regions. Numerische Mathematik, 26(2): 117–154, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jacobsen:1989:SOD

[JTW89] L. Jacobsen, W. J. Thron, and H. Waadeland. Some observations on the distribution of values of continued fractions. Numerische Mathematik, 55(6): 711–733, August 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jacobsen:1988:CAL

[JW88] Lisa Jacobsen and Haakon Waadeland. Convergence acceleration of limit periodic continued fractions under asymptotic side conditions. Numerische Mathematik, 53(3): 285–298, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Joe:1993:PCC

[JY93] S. Joe and Y. Yan. A piecewise constant collocation method using cosine mesh-grading for Symm’s equation. Numerische Mathematik, 65(4):423–433, August 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kacewicz:1976:IIM

[Kac76] Bolesław Kacewicz. An integralinterpolation iterative method for the solution of scalar equations. Numerische Mathematik, 26(4):355–365, December 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kacewicz:1984:HIO

[Kac84] B. Z. Kacewicz. How to increase the order to get minimal-error algorithms for systems of ODE. Numerische Mathematik, 45(1): 93–104, October 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kadner:1967:NBI

[Kad67] Horst Kadner. Die numerische Behandlung von Integralgleichungen nach der Kollokationsmethode. (German) [The numerical treatment of integral equations in the collocation method]. Numerische Mathematik, 10(3):241–260, October 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kall:1974:ASP

[Kal74] Peter Kall. Approximations to stochastic programs with complete fixed recourse. Numerische Mathematik, 22(4):333–339, August 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kaltenbacher:1998:PPC

[Kal98] Barbara Kaltenbacher. A posteriori parameter choice strategies for some Newton type methods for the regularization of nonlinear ill-posed problems. Numerische Mathematik, 79(4): 501–528, June 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.

de/link/service/journals/
00211/bibs/8079004/80790501.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790501.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790501.
ps.gz.

Kambo:1971:EBL

[Kam71] N. S. Kambo. Error bounds for the Lobatto and Radau quadrature formulas. Numerische Mathematik, 16(5): 383–388, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kansy:1973:EFA

[Kan73]  Klaus Kansy.  Elementare

Fehlerdarstellung für Ableitungen bei der Hermite-Interpolation. (German) [Elementary error representation for derivatives by Hermite interpolation]. Numerische Mathematik, 21(4): 350–354, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kanzow:1998:IBM

[Kan98] Christian Kanzow. An inexactQPbased method for nonlinear complementarity problems. Numerische Mathematik, 80(4): 557–577, October 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8080004/

80800557.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8080004/
80800557.pdf.

Kappert:1996:ZPS

[Kap96]  M. Kappert.  On the ze-

ros of the partial sums of cos(z) and sin(z). Numerische Mathematik, 74(4):397–417, October 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074004/60740397.
htm;   http://science.springer.
de/nmee/bibs/6074004/60740397.
htm.

Karst:1961:FMZ

[Kar61] E. Karst. Faktorenzerlegung Mersennescher Zahlen mittels programmgesteuerter Rechengeräte. (German) [Factorization of Mersenne numbers with programmable computers]. Numerische Mathematik, 3:79– 86, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Karamardian:1968:ESC

[Kar68]  S. Karamardian.  Existence

of solutions of certain systems of nonlinear inequalities. Numerische Mathematik, 12(4): 327–334, November 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kato:1960:EIM

[Kat60] Tosio Kato. Estimation of iterated matrices, with application to the von Neumann condition. Numerische Mathematik, 2 (1):22–29, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kautsky:1981:MRI

[Kau81] Jaroslav Kautský. Matrices related to interpolatory quadratures. Numerische Mathematik, 36(3):309–318, April 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kauthen:1989:CTC

[Kau89]  J.-P. Kauthen.  Continu-

ous time collocation methods for Volterra–Fredholm integral equations. Numerische Mathematik, 56(5):409–424, November 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Koc:1993:FAS

[KC93]  Çetin Kaya Koç and Guanrong

Chen. A fast algorithm for scalar Nevanlinna–Pick interpolation. Numerische Mathematik, 64(1): 115–126, January 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kautsky:1982:CWI

[KE82] J. Kautský and S. Elhay. Calculation of the weights of interpolatory quadratures. Numerische Mathematik, 40(3): 407–422, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Keast:1968:TBV

[Kea68] P. Keast. The third boundary value problem for elliptic equations. Numerische Mathematik, 12(4):322–326, November 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kearfott:1979:EDC

[Kea79] Baker Kearfott. An efficient degree-computation method for a generalized method of bisection. Numerische Mathematik, 32(2):109–127, June 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Keller:1965:AFD

[Kel65] Herbert B. Keller. On the accuracy of finite difference approximations to the eigenvalues of differential and integral operators. Numerische Mathematik, 7 (5):412–419, October 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kellogg:1972:CEM

[Kel72] R. B. Kellogg. On compiles eigenvalues of M and P matrices. Numerische Mathematik, 19 (2):170–175, April 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kempfert:1968:SDR

[Kem68] H. Kempfert. On sign determinations in real algebraic number fields. Numerische Mathematik, 11(2):170–174, February 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kemper:1972:LMM

[Kem72] Gene A. Kemper. Linear multistep methods for a class of functional differential equations. Numerische Mathematik, 19(5): 361–372, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kerner:1966:GBN

[Ker66] Immo O. Kerner. Ein Gesamtschrittverfahren zur Berechnung der Nullstellen von Polynomen. (German) [A complete step method for the computation of zeros of polynomials]. Numerische Mathematik, 8(3):290–294, May 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kesavan:1979:MKA

[Kes79] S. Kesavan. La méthode de Kikuchi appliquée aux équations de von Kármán. (French) [Kikuchi’s method applied to von Kármán’s equations]. Numerische Mathematik, 32(2): 209–232, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kesavan:1984:PMB

[Kes84] S. Kesavan. A potential method for the biharmonic equation. Numerische Mathematik, 43(1): 105–119, January 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kikuchi:1981:CPF

[Kik81] Noboru Kikuchi. Convergence of a penalty-finite element approximation for an obstacle problem. Numerische Mathematik, 37(1):105–120, May 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kiluchi:1975:FES

[Kil75] Fumio Kiluchi. On a finite element scheme based on the discrete Kirchhoff assumption. Numerische Mathematik, 24(3): 211–231, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kilberth:1978:IBM

[Kil78] K. Kilberth. Iterative Berechnung mehrfacher Lösungen von Hammersteinschen Integralgleichungen. (German) [Iterative calculation of multiple solutions of Hammerstein integral equations]. Numerische Mathematik, 29(2):195–207, January 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kimura:1996:AEC

[Kim96]  Masato Kimura.  Asymp-

totic estimation for the condition numbers in BEM. Numerische Mathematik, 73(2): 209–233, April 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073002/60730209.
htm;   http://science.springer.
de/nmee/bibs/6073002/60730209.
htm.

Kim:1998:DDI

[Kim98a] Seongjai Kim. Domain decomposition iterative procedures for solving scalar waves in the frequency domain. Numerische Mathematik, 79(2): 231–259, April 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079002/80790231.
htm;   http://science.springer.
de/nmee/bibs/8079002/80790231.
htm.

Kim:1998:URI

[Kim98b] Seongjai Kim. On the use of rational iterations and domain decomposition methods for the Helmholtz problem. Numerische Mathematik, 79(4): 529–552, June 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079004/80790529.

htm;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790529.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790529.
ps.gz.

King:1972:TMN

[Kin72] Richard F. King. Tangent methods for nonlinear equations. Numerische Mathematik, 18(4): 298–304, August 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kincaid:1973:CNI

[Kin73] David R. Kincaid. A class of norms of iterative methods for solving systems of linear equations. Numerische Mathematik, 20(5):392–408, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

King:1974:NEB

[Kin74] J. T. King. New error bounds for the penalty method and extrapolation. Numerische Mathematik, 23(2):153–165, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

King:1992:MAI

[Kin92] J. Thomas King. Multilevel algorithms for ill-posed problems. Numerische Mathematik, 61(3): 311–334, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kioustelidis:1979:DFT

[Kio79]  John B. Kioustelidis.  A

derivative-free transformation preserving the order of convergence of iteration methods in case of multiple zeros. Numerische Mathematik, 33(4): 385–389, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kirchgraber:1986:MME

[Kir86] Urs Kirchgraber. Multi-step methods are essentially one-step methods. Numerische Mathematik, 48(1):85–90, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kirchgraber:1988:OSB

[Kir88] U. Kirchgraber. An ODE-solver based on the method of averaging. Numerische Mathematik, 53(6):621–652, September 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kiwiel:1984:QAM

[Kiw84]  Krzysztof C. Kiwiel.  A

quadratic approximation method for minimizing a class of quasidifferentiable functions. Numerische Mathematik, 45(3): 411–430, December 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kiwiel:1987:CLM

[Kiw87] Krzysztof C. Kiwiel. A constraint linearization method for nondifferentiable convex minimization. Numerische Mathematik, 51(4):395–414, August 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kiwiel:1994:CDM

[Kiw94] Krzysztof C. Kiwiel. A Cholesky dual method for proximal piecewise linear programming. Numerische Mathematik, 68(3): 325–340, September 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4068003/
40680325.htm;   http://science.
springer.de/nmee/bibs/4068003/
40680325.htm.

Kerkhoven:1990:SFE

[KJ90] Thomas Kerkhoven and Joseph W.

Jerome. L_{∞} stability of finite element approximations to elliptic gradient equations. Numerische Mathematik, 57(6/7): 561–575, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Kjellberg:1961:SRM

[Kje61] Göran Kjellberg. On the successive over-relaxation method for cyclic operators. Numerische Mathematik, 3:87–91, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kang:1995:VPA

[KjS95] Feng Kang and Zai jiu Shang. Volume-preserving algorithms for source-free dynamical systems. Numerische Mathematik, 71(4):451–463, October 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5071004/
50710451.htm;   http://science.
springer.de/nmee/bibs/5071004/
50710451.htm.

Knauff:1974:OAL

[KK74] W. Knauff and R. Kreß. Optimale Approximation linearer Funktionale auf periodischen Funktionen. (German) [Optimal approximation of linear functionals on periodic functions]. Numerische Mathematik, 22(3): 187–205, June 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Knauff:1976:OAN

[KK76] W. Knauff and R. Kreß. Optimale Approximation mit Nebenbedingungen an lineare Funktionale auf periodischen Funktionen. (German) [Optimal approximation with side conditions to linear functionals on periodic functions]. Numerische Mathematik, 25(2):

149–152, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kim:1997:EDC

[KK97] Sang Dong Kim and Seokchan Kim. Exponential decay of C^{1}-cubic splines vanishing at two symmetric points in each knot interval. Numerische Mathematik, 76(4): 479–488, June 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076004/70760479.
htm;   http://science.springer.
de/nmee/bibs/7076004/70760479.
htm.

Kim:1997:FDP

[KKL97] Hong Oh Kim, Sang Dong Kim, and Yong Hun Lee. Finite difference preconditioning cubic spline collocation method of elliptic equations. Numerische Mathematik, 77(1): 83–103, July 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077001/70770083.
htm;   http://science.springer.
de/nmee/bibs/7077001/70770083.
htm.

Kulkarni:1983:SCA

[KL83] R. P. Kulkarni and B. V.

Limaye. On the steps of convergence of approximate eigenvectors in the Rayleigh– Schrödinger series. Numerische Mathematik, 42(1):31–50, September 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kloeden:1990:NMM

[KL90] P. E. Kloeden and J. Lorenz. A note on multistep methods and attracting sets of dynamical systems. Numerische Mathematik, 56(7):667–673, January 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Klee:1965:CLP

[Kle65] Victor Klee. A class of linear programming problems requiring a large number of iterations. Numerische Mathematik, 7(4):313–321, August 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Klee:1966:CPD

[Kle66] Victor Klee. A comparison of primal and dual methods of linear programming. Numerische Mathematik, 9(3):227– 235, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kleinmichel:1981:QNVa

[Kle81a] Helmut Kleinmichel. Quasi-Newton-Verfahren vom Rang-Eins-Typ zur Lösung unrestringierter Minimierungsprobleme. I. Verfahren und grundlegende Eigenschaften. (German) [Quasi-Newton methods of rankone type for the solution of unconstrained minimization problems. I. methods and general properties]. Numerische Mathematik, 38(2):219–228, December 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kleinmichel:1981:QNVb

[Kle81b] Helmut Kleinmichel. Quasi-Newton-Verfahren vom Rang-Eins-Typ zur Lösung unrestringierter Minimierungsprobleme. II. n-Schritt-quadratische Konvergenz für Restart-Varianten. (German) [Quasi-Newton methods of rank-one-type for the solution of unconstrained minimization problems. II. nstep quadratic convergence for restart variants]. Numerische Mathematik, 38(2): 229–244, December 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Klouvcek:1999:RNQ

[Klo99] Petr Klouček. The relaxation of non-quasiconvex variational integrals. Numerische Mathematik, 82(2):281–311, April 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082002/90820281.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082002/90820281.pdf.

Keast:1967:FDS

[KM67] P. Keast and A. R. Mitchell. Finite difference solution of the third boundary problem in elliptic and parabolic equations. Numerische Mathematik, 10(1): 67–75, May 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Krishnakumar:1986:EST

[KM86] A. S. Krishnakumar and Martin Morf. Eigenvalues of a symmetric tridiagonal matrix: a divide-and-conquer approach. Numerische Mathematik, 48(3): 349–368, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kunkel:1991:SFM

[KM91] Peter Kunkel and Volker Mehrmann. Smooth factorizations of matrix valued functions and their derivatives. Numerische Mathematik, 60(1):115–131, November 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kunkel:1998:RSN

[KM98] Peter Kunkel and Volker Mehrmann. Regular solutions of nonlinear differential-algebraic equations and their numerical determination. Numerische Mathematik, 79(4):581–600, June 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/8079004/

80790581.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8079004/
80790581.pdf;   http://link.
springer.de/link/service/
journals/00211/papers/8079004/
80790581.ps.gz.

Kossioris:1999:FVS

[KMS99] G. Kossioris, Ch. Makridakis, and P. E. Souganidis. Finite volume schemes for Hamilton–Jacobi equations. Numerische Mathematik, 83 (3):427–442, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083003/90830427.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083003/90830427.pdf.

Krizek:1984:SPF

[KN84] Michal Křížek and Pekka Neittaanmäki. Superconvergence phenomenon in the finite element method arising from averaging gradients. Numerische Mathematik, 45(1):105–116, October 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Knauer:1971:USE

[Kna71] B. Knauer. Untere Schranken für die Eigenwerte positivdefiniter, selbstadjungierter Operatoren. (German) [Lower bounds for the eigenvalues of positive definite selfadjoint operators]. Numerische Mathematik, 17(2):166–171, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Knabner:1985:CSP

[Kna85] Peter Knabner. Control of Stefan problems by means of linearquadratic defect minimization. Numerische Mathematik, 46(3): 429–442, June 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Knoth:1989:GSG

[Kno89] O. Knoth. A globalization scheme for the generalized Gauss–Newton method. Numerische Mathematik, 56(6): 591–607, December 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kroner:1995:CHO

[KNR95] Dietmar Kröner, Sebastian Noelle, and Mirko Rokyta. Convergence of higher order upwind finite volume schemes on unstructured grids for scalar conservation laws in several space dimensions. Numerische Mathematik, 71(4):527–560, October 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/5071004/50710527.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/5071004/50710527.
                                        htm.

Kochengin:1998:DRS

[KO98] Sergey A. Kochengin and Vladimir I. Oliker. Determination of reflector surfaces from near-field scattering data II. Numerical solution. Numerische Mathematik, 79(4): 553–568, June 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079004/80790553.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790553.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790553.
ps.gz.

Kocan:1995:AVS

[Koc95] Maciej Kocan. Approximation of viscosity solutions of elliptic partial differential equations on minimal grids. Numerische Mathematik, 72(1):73– 92, November 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5072001/50720073.
htm;   http://science.springer.
de/nmee/bibs/5072001/50720073.
htm.

Kohaupt:1987:OEE

[Koh87] Ludwig Kohaupt. Optimal L^{∞}- error estimates for piecewise linear finite elements in r^{n} for 1 ≤ n ≤ 3. Numerische Mathematik, 50(3):353–362, May 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Koh88].

Kohaupt:1988:EEE

[Koh88] Ludwig Kohaupt. Erratum: “Optimal L^{∞}-error estimates for piecewise linear finite elements in R^{n} for 1 ≤ n ≤ 3” [Numer. Math. 50 (1987), no. 3, 353–362 and MR 88i:65131]. Numerische Mathematik, 52(5): 604, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Koh87].

Kohler:1995:CAB

[Köh95] P. Köhler. Construction of asymptotically best quadrature formulas. Numerische Mathematik, 72(1):93–116, November 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer.de/link/service/
                                        journals/00211/bibs/5072001/
                                        50720093.htm;   http://science.
                                        springer.de/nmee/bibs/5072001/
                                        50720093.htm.

Kolata:1978:AVP

[Kol78] William G. Kolata. Approximation in variationally posed eigenvalue problems. Numerische Mathematik, 29(2): 159–171, January 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kornhuber:1994:MMM

[Kor94] Ralf Kornhuber. Monotone multigrid methods for elliptic variational inequalities. I. Numerische Mathematik, 69(2): 167–184, December 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4069002/
40690167.htm;   http://science.
springer.de/nmee/bibs/4069002/
40690167.htm.

Kornhuber:1996:MMM

[Kor96] Ralf Kornhuber. Monotone multigrid methods for elliptic variational inequalities. II. Numerische Mathematik, 72(4): 481–499, February 1996. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/6072004/
60720481.htm;   http://science.
springer.de/nmee/bibs/6072004/
60720481.htm.

Kosler:1985:FGN

[Kos85] R. Kosler. W^{m,2}-Fehlerabschätzungen für Galerkin-Näherungen schwach elliptischer quasilinearer Randwertprobleme 2m-ter Ordnung. (German) [W^{m,2}-error estimates for Galerkin approximations of weakly elliptic quasilinear boundary value problems of order 2m]. Numerische Mathematik, 47(4):597–609, December 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kosler:1986:PSD

[Kos86]  Robert Kosler.  A priori-

Schranken für diskrete Lösungen monotoner Operatorgleichungen und Anwendungen. (German) [A priori bounds for discrete solutions of monotone operator equations and applications]. Numerische Mathematik, 50(1): 45–55, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kossmann:1981:DBP

[Koß81] H. Koßmann. Differenzenverfahren zur Berechnung periodischer Lösungen von hyperbolischen Differentialgleichungen. (German) [Finite difference methods for the computation of periodic solutions of hyperbolic differential equations]. Numerische Mathematik, 37(2): 219–233, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Koto:1998:SRI

[Kot98]  Toshiyuki Koto.  Stability

of the Radau IA and Lobatto IIIC methods for delay differential equations. Numerische Mathematik, 79(4): 569–580, June 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079004/80790569.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790569.
pdf;   http://link.springer.
de/link/service/journals/

00211/papers/8079004/80790569.
ps.gz.

Koto:1999:SRK

[Kot99] Toshiyuki Koto. Stability of Runge–Kutta methods for the generalized pantograph equation. Numerische Mathematik, 84(2):233–247, December 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9084002/90840233.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9084002/90840233.pdf.

Konig:1981:KAM

[KP81] Hermann König and Diethard Pallaschke. On Khachian’s algorithm and minimal ellipsoids. Numerische Mathematik, 36(2): 211–223, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kratz:1990:NAS

[KP90] W. Kratz and A. Peyerimhoff.

A numerical algorithm for the Stokes problem based on an integral equation for the pressure via conformal mappings. Numerische Mathematik, 58(3): 255–272, October 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kunisch:1991:ETS

[KP91] K. Kunisch and G. Peichl. Estimation of a temporally and spatially varying diffusion coefficient in a parabolic system by an augmented Lagrangian technique. Numerische Mathematik, 59(5):473–509, August 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kim:1995:PCS

[KP95] Sang Dong Kim and Seymour V. Parter. Preconditioning cubic spline collocation — discretizations of elliptic equations. Numerische Mathematik, 72(1):39– 72, November 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5072001/50720039.
htm;   http://science.springer.
de/nmee/bibs/5072001/50720039.
htm.

Kaps:1979:GRK

[KR79] Peter Kaps and Peter Rentrop. Generalized Runge–Kutta methods of order four with stepsize control for stiff ordinary differential equations. Numerische Mathematik, 33(1):55– 68, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Korner:1982:ELB

[KR82] Frank Körner and Claus Richter. Zur effektiven Lösung von booleschen, quadratischen Optimierungsproblemen. (German) [On the efficient solution of Boolean quadratic programming problems]. Numerische Mathematik, 40(1):99–109, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kannan:1984:MIM

[KR84] R. Kannan and M. B. Ray.

Monotone iterative methods for nonlinear equations involving a noninvertible linear part. Numerische Mathematik, 45(2): 219–225, November 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Karlsen:1997:OSM

[KR97] Kenneth Hvistendahl Karlsen and Nils Henrik Risebro. An operator splitting method for nonlinear convection-diffusion equations. Numerische Mathematik, 77(3):365–382, September 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7077003/
70770365.htm;   http://science.
springer.de/nmee/bibs/7077003/
70770365.htm.

Kranzer:1963:EFN

[Kra63] H. C. Kranzer. An error formula for numerical differentiation. Numerische Mathematik, 5(1):439–442, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-

                                        599X&volume=5&issue=1&spage=
                                        439.

Krabs:1968:LOH

[Kra68] Werner Krabs. Lineare Optimierung in halbgeordneten Vektorräumen. (German) [Linear optimization in half-ordered vector spaces]. Numerische Mathematik, 11(3):220–231, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kraaijevanger:1986:AMP

[Kra86] J. F. B. M. Kraaijevanger. Absolute monotonicity of polynomials occurring in the numerical solution of initial value problems. Numerische Mathematik, 48(3):303–322, March 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kratzschmar:1989:DDA

[Krä89]  Michael Krätzschmar.  A

decomposition-dualization approach for solving constrained convex minimization problems with applications to discretized obstacle problems. Numerische Mathematik, 54(5): 507–531, February 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kreiss:1959:DHG

[Kre59] Heinz-Otto Kreiss. Über die Differenzapproximation hoher Genauigkeit bei Anfangswertproblemen für partielle Differentialgleichungen. (German) [on the difference approximation of higher accuracy with initial value problems for partial differential equations]. Numerische Mathematik, 1(1):186– 202, December 1959. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
186.

Kreiss:1963:UID

[Kre63] Heinz-Otto Kreiss. Über implizite Differenzmethoden für partielle Differentialgleichungen. (German) [On implicit difference methods for partial differential equations]. Numerische Mathematik, 5(1):24– 47, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kreiss:1968:NSS

[Kre68] Heinz-Otto Kreiss. On the numerical solution of the spherically symmetric diffusion equation. Numerische Mathematik, 12(3):223–225, October 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kress:1971:ART

[Kre71] Rainer Kreß. Ein ableitungsfreies Restglied für die trigonometrische Interpolation periodischer analytischer Funktionen. (German) [A derivative-free residue for the trigonometric interpolation of periodic analytic functions]. Numerische Mathematik, 16(5): 389–396, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kress:1972:GHT

[Kre72a] Rainer Kreß. On general Hermite trigonometric interpolation. Numerische Mathematik, 20(2):125–138, April 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kress:1972:NIP

[Kre72b] Rainer Kreß. Zur numerischen Integration periodischer Funktionen nach der Rechteckregel. (German) [On the numerical integration of periodic functions by the trapezoidal rule]. Numerische Mathematik, 20(1):87– 92, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kress:1990:NMB

[Kre90]  Rainer Kress.  A Nyström

method for boundary integral equations in domains with corners. Numerische Mathematik, 58(2):145–161, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kripke:1964:BAR

[Kri64] Bernard R. Kripke. Best approximation with respect to nearby norms. Numerische Mathematik, 6(1):103–105, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Krishna:1983:SNR

[Kri83] Lala B. Krishna. Some new results on unsymmetric successive overrelaxation method. Numerische Mathematik, 42(2): 155–160, November 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kammerer:1974:QIS

[KRV74] W. J. Kammerer, G. W. Reddien, and R. S. Varga. Quadratic interpolatory splines. Numerische Mathematik, 22(4): 241–259, August 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kiyono:1971:BCS

[KS71] Takeshi Kiyono and Masaaki Shimasaki. Boundary contraction solution of the Neumann and mixed boundary value problems of the Laplace equation. Numerische Mathematik, 16(5): 442–450, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kreisselmeier:1973:AZI

[KS73] Gerhard Kreisselmeier and Dietrich Suschowk. Ein Algorithmus zur zyklisch invarianten Zerlegung eines endlichdimensionalen Vektorraums bezüglich einer linearen Selbstabbildung.

(German) [An algorithm for the cyclic invariant decompositions of a finite-dimensional vector space with respect to linear endomorphism]. Numerische Mathematik, 22(1):69– 74, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kress:1983:CNB

[KS83] R. Kreß and W. T. Spassov. On the condition number of boundary integral operators for the exterior Dirichlet problem for the Helmholtz equation. Numerische Mathematik, 42(1):77– 95, September 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kelley:1989:PQN

[KS89] C. T. Kelley and E. W. Sachs. A pointwise quasi-Newton method for unconstrained optimal control problems. Numerische Mathematik, 55(2):159–176, May 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kraaijevanger:1991:USR

[KS91] J. F. B. M. Kraaijevanger and J. Schneid. On the unique solvability of the Runge–Kutta equations. Numerische Mathematik, 59(2):129–157, May 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kerkhoven:1992:AMC

[KS92] T. Kerkhoven and Y. Saad. On acceleration methods for coupled nonlinear elliptic systems. Numerische Mathematik, 60(4): 525–548, January 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kress:1993:NSL

[KS93] Rainer Kreß and Ian H. Sloan.

On the numerical solution of a logarithmic integral equation of the first kind for the Helmholtz equation. Numerische Mathematik, 66(2):199–214, November 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Khoromskij:1998:FIS

[KS98] B. N. Khoromskij and G. Schmidt.

A fast interface solver for the biharmonic Dirichlet problem on polygonal domains. Numerische Mathematik, 78(4):577–596, February 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078004/70780577.
htm;   http://science.springer.
de/nmee/bibs/7078004/70780577.
htm.

Klawonn:1999:BTP

[KS99] Axel Klawonn and Gerhard Starke. Block triangular preconditioners for nonsymmetric saddle point problems: field-of-values analysis. Numerische Mathematik, 81(4):577–594, February 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9081004/90810577.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9081004/90810577.pdf.

Kschwendt:1968:CLM

[Ksc68] H. Kschwendt. Convergence limits in the Monte Carlo theory of integral equations. Numerische Mathematik, 11(4): 307–314, May 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kreyszig:1959:RUE

[KT59] Erwin Kreyszig and John Todd. The radius of univalence of the error function. Numerische Mathematik, 1:78–89, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kuenzi:1965:DA

[KT65] H. P. Kuenzi and H. Tzschach. The duoplex-algorithm. Numerische Mathematik, 7(3):222– 225, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kublanovskaya:1984:AMS

[Kub84] V. N. Kublanovskaya. ABalgorithm and its modifications for the spectral problems of linear pencils of matrices. Numerische Mathematik, 43(3): 329–342, April 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kuhn:1993:AEU

[Kuh93] Dieter Kuhn. About existence and uniqueness of the restricted total least squares estimation. Numerische Mathematik, 65(4): 435–444, August 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kulisch:1968:RZM

[Kul68]  U. Kulisch.  Über reguläre

Zerlegungen von Matrizen und einige Anwendungen. (German) [On regular dismantlings of stencils and some applications]. Numerische Mathematik, 11(5): 444–449, June 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kulisch:1971:AAR

[Kul71] U. Kulisch. An axiomatic approach to rounded computations. Numerische Mathematik, 18(1):1–17, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kung:1974:CRP

[Kun74] H. T. Kung. On computing reciprocals of power series. Numerische Mathematik, 22(5): 341–348, October 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kupka:1971:SRA

[Kup71] I. Kupka. Simulation reeller Arithmetik und reeller Funktionen in endlichen Mengen. (German) [Simulation of real arithmetic and real functions in finite sets]. Numerische Mathematik, 17(2):143–152, April 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kupper:1976:EGD

[Küp76a] Tassilo Küpper. Einschließungsenaussagen für gewönliche Differentialoperatoren. (German) [Range-domain implications for ordinary differential operators]. Numerische Mathematik, 25(2): 201–214, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kupper:1976:PLU

[Küp76b] Tassilo Küpper. Pointwise lower and upper bounds for eigenfunctions of ordinary differential operators. Numerische Mathematik, 27(1):111–119, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kupper:1978:EDZ

[Küp78] Tassilo Küpper. Einschließungsaussagen bei Differentialoperatoren zweiter Ordnung durch punktweise Ungleichungen. (German) [Range-domain implications for second order differential operators using pointwise inequalities]. Numerische Mathematik, 30(1):93–101, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kuttler:1971:FAE

[Kut71] James R. Kuttler. A finitedifference approximation for the eigenvalues of the clamped plate. Numerische Mathematik, 17(3):230–238, June 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kutt:1975:NEP

[Kut75] H. R. Kutt. The numerical evaluation of principal value integrals by finite-part integration. Numerische Mathematik, 24(3): 205–210, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kutz:1982:MIQ

[Küt82] Martin Kütz. Monotony in interpolatory quadratures. Numerische Mathematik, 39(3): 421–428, October 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kammerer:1972:ABN

[KV72] W. J. Kammerer and R. S.

Varga. On asymptotically best norms for powers of operators. Numerische Mathematik, 20(2): 93–98, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kesavan:1985:SBE

[KV85] S. Kesavan and A. S. Vasudevamurthy. On some boundary element methods for the heat equation. Numerische Mathematik, 46(1):101–120, March 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Klee:1977:LAT

[KvdD77] Victor Klee and Pauline van den Driessche. Linear algorithms for testing the sign stability of a matrix and for finding Z-matrix matchings an acyclic graphs. Numerische Mathematik, 28(3): 273–285, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kaps:1981:SRT

[KW81] P. Kaps and G. Wanner. A study of Rosenbrock-type methods of high order. Numerische Mathematik, 38(2): 279–298, December 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Knowles:1995:VMN

[KW95] Ian Knowles and Robert Wallace. A variational method for numerical differentiation. Numerische Mathematik, 70(1):91– 110, March 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/

00211/bibs/5070001/50700091.
htm;   http://science.springer.
de/nmee/bibs/5070001/50700091.
htm.

Khoromskij:1996:RIM

[KW96] B. N. Khoromskij and G. Wittum. Robust iterative methods for elliptic problems with highly varying coefficients in thin substructures. Numerische Mathematik, 73(4): 449–472, June 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073004/60730449.
htm;   http://science.springer.
de/nmee/bibs/6073004/60730449.
htm.

Khoromskij:1999:AOS

[KW99] Boris N. Khoromskij and Gabriel Wittum. An asymptotically optimal Schur complement reduction for the Stokes equation. Numerische Mathematik, 81(3):345–375, January 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/9081003/90810345.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/9081003/90810345.
pdf.

Kowalski:1985:GQO

[KWW85] M. A. Kowalski, A. G. Werschulz, and H. Woźniakowski. Is Gauss quadrature optimal for analytic functions? Numerische Mathematik, 47(1):89–98, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kedem:1974:TGL

[KZ74] Gershon Kedem and S. K. Zaremba. A table of good lattice points in three dimensions. Numerische Mathematik, 23(2): 175–180, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kocvara:1994:ITS

[KZ94] Michal Kočvara and Jochem Zowe. An iterative twostep algorithm for linear complementarity problems. Numerische Mathematik, 68(1): 95–106, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/4068001/40680095.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4068001/40680095.
                                        htm.

LopezdeSilanes:1988:EEA

[LA88] M. C. López de Silanes and D. Apprato. Estimations de l’erreur d’approximation sur un domaine borné de r^{n} par D^{m}-splines d’interpolation et d’ajustement discrètes. (French) [Approximation error estimates on a bounded domain in r^{n} for interpolating and smoothing discrete D^{m}-splines ]. Numerische Mathematik, 53(3):

367–376, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LopezdeSilanes:1989:EEA

[LA89] M. C. López de Silanes and R. Arcangéli. Estimations de l’erreur d’approximation par splines d’interpolation et d’ajustement d’ordre (m, s). (French) [Approximation error estimates for interpolating and smoothing (m, s)-splines]. Numerische Mathematik, 56(5): 449–467, November 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lu:1997:CAB

[LA97] Hao Lu and Owe Axelsson. Conditioning analysis of block incomplete factorizations and its application to elliptic equations. Numerische Mathematik, 78(2):189–209, December 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078002/70780189.
htm;   http://science.springer.
de/nmee/bibs/7078002/70780189.
htm.

Lafon:1975:BTM

[Laf75] Jean-Claude Lafon. Base tensorielle de matrices de Hankel (ou de Toeplitz): Applications. (French) [Tensorial bases of Hankel (or Toeplitz) matrices applications]. Numerische Mathematik, 23(4):349–361, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lamprecht:1970:AFR

[Lam70]  Gúnther Lamprecht.  Über

die Abhängigkeit der Fehlerfunktion bei der rationale L_{p}- Approximation von den Koeffizienten des Nennerpolynoms bei minimalem Zählerpolynom. (German) [On the relation between the error function with rational L_{p} approximation and the coefficients of the denominator polynomial with minimum numerator polynomial]. Numerische Mathematik, 15(5): 392–403, September 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lambert:1980:ACM

[Lam80] Joseph M. Lambert. An alternative computational method for the approximation of random functions. Numerische Mathematik, 35(4):361–368, November 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lamour:1991:WPS

[Lam91] René Lamour. A well-posed shooting method for transferable DAE’s. Numerische Mathematik, 59(8):815–829, October 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lancaster:1964:ALM

[Lan64a]  P. Lancaster.  Algorithms

for lambda-matrices. Numerische Mathematik, 6(1):388– 394, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lancaster:1964:EMD

[Lan64b] P. Lancaster. On eigenvalues of matrices dependent on a parameter. Numerische Mathematik, 6(1):377–387, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lancaster:1966:EAN

[Lan66] P. Lancaster. Error analysis for the Newton–Raphson method. Numerische Mathematik, 9(1): 55–68, November 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lancaster:1972:NSM

[Lan72]  P. Lancaster.  A note on

sub-multiplicative norms. Numerische Mathematik, 19(3): 206–208, June 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Langford:1977:NSB

[Lan77] W. F. Langford. Numerical solution of bifurcation problems for ordinary differential equations. Numerische Mathematik, 28(2):171–190, June 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Langer:1987:NLE

[Lan87] Ulrich Langer. Zur numerischen Lösung des ersten biharmonischen Randwertproblems. (German) [On the numerical solution of the first biharmonic boundary value problem]. Numerische Mathematik, 50(3): 291–310, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Larkin:1981:RFD

[Lar81] F. M. Larkin. Root finding by divided differences. Numerische Mathematik, 37(1):93–104, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lasiecka:1992:CRA

[Las92] Irena Lasiecka. Convergence rates for the approximations of the solutions to algebraic Riccati equations with unbounded coefficients: case of analytic semigroups. Numerische Mathematik, 63(3):357–390, December 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lauchli:1961:JEA

[Läu61] Peter Läuchli. Jordan-Elimination und Ausgleichung nach kleinsten Quadraten. (German) [Jordan elimination and balancing using least squares]. Numerische Mathematik, 3:226–240, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Laurent:1967:TCM

[Lau67] P. J. Laurent. Théorèmes de caractérisation d’une meilleure approximation dans un espace norme et généralisation de l’algorithme de Remes. (French) [Theorems for characterizing the best approximation in a normed space and generalization of the Remes algorithm]. Numerische Mathematik, 10(3):190–208, October 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Laubin:1998:HOC

[Lau98] Pascal Laubin. High order convergence for collocation of second kind boundary integral equations on polygons. Numerische Mathematik, 79(1): 107–140, March 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079001/80790107.
htm;   http://science.springer.
de/nmee/bibs/8079001/80790107.
htm.

Lawson:1964:CPS

[Law64] Ch.L. Lawson. Characteristic properties of the segmented rational minmax approximation problem. Numerische Mathematik, 6(1):293–301, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Liu:1997:DRH

[LB97] Kang-Man Liu and Ivo Babuška.

Dimensional reduction for Helmholtz’s equation on a bounded domain. Numerische Mathematik, 77(4): 501–533, October 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7077004/
70770501.htm;   http://science.
springer.de/nmee/bibs/7077004/
70770501.htm.

Lou:1998:OSC

[LBF98] Zhou-Ming Lou, Bernard Bialecki, and Graeme Fairweather. Orthogonal spline collocation methods for biharmonic problems. Numerische Mathematik, 80(2):267–303, August 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8080002/
80800267.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8080002/
80800267.pdf;   http://link.
springer.de/link/service/
journals/00211/papers/8080002/
80800267.ps.gz.

Lyness:1993:QEE

[LdD93] J. N. Lyness and E. de Doncker. Quadrature error expansions. II. the full corner singularity. Numerische Mathematik, 64 (3):355–370, March 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LeRoux:1980:MMP

[Le 80a] Marie-Noëlle Le Roux. Méthodes multipas pour des équations paraboliques non linéaires. (French) [Multipass methods for nonlinear parabolic equations]. Numerische Mathematik, 35(2): 143–162, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LeTallec:1980:MFE

[Le 80b] P. Le Tallec. A mixed finite element approximation of the Navier–Stokes equations. Numerische Mathematik, 35(4): 381–404, November 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LeTallec:1982:EAR

[Le 82] P. Le Tallec. Existence and approximation results for nonlinear mixed problems: application to incompressible finite elasticity. Numerische Mathematik, 38 (3):365–382, March 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lipitakis:1981:RCA

[LE81] Elias A. Lipitakis and David J.

Evans. The rate of convergence of an approximate matrix factorization semi-direct method. Numerische Mathematik, 36(3): 237–251, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lebaud:1970:LDA

[Leb70] Colette Lebaud. L’algorithm double QR avec <<shift>>. (French) [The double QR algorithm with ‘shift’]. Numerische Mathematik, 16(2): 163–180, November 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lees:1961:ADS

[Lee61] Milton Lees. Alternating direction and semi-explicit difference methods for parabolic partial differential equations. Numerische Mathematik, 3(1): 398–412, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
398.

Legras:1966:RNG

[Leg66] J. Legras. Résolution numérique des grands systèmes différentiels linéaires. (French) [Numerical solution of large systems of linear differentials]. Numerische Mathematik, 8(1):14–28, March 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lehmann:1960:FRV

[Leh60] N. J. Lehmann. Eine Fehlerabschätzung zum Ritzschen Verfahren für inhomogene Randwertaufgaben. (German) [An error bound for the Ritz method for inhomogeneous boundary value problems]. Numerische Mathematik, 2(1):60– 66, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lehmann:1961:INR

[Leh61] N. J. Lehmann. Das inhomogene natürliche Randwertproblem und Fehlerabschätzungen für Näherungslösungen. (German) [The inhomogeneous natural boundary value problem and error bounds for approximate solution]. Numerische Mathematik, 3:1–29, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lehmann:1963:OE

[Leh63] N. J. Lehmann. Optimale Eigenwerteinschliessungen. (German) [Optimal eigenvalue inclusion]. Numerische Mathematik, 5(1): 246–272, December 1963. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lehmann:1966:VOE

[Leh66] N. J. Lehmann. Zur Verwendung optimaler Eigenwerteingrenzungen bei der Lösung symmetrischer Matrizenaufgaben. (German) [On the use of optimal eigenvalue enclosure for the solution of symmetric matrix problems]. Numerische Mathematik, 8(1):42–55, March 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lehmann:1967:FND

[Leh67] N. J. Lehmann. Fehlerschranken für Näherungslösungen bei Differentialgleichungen. (German) [Error bounds for approximate solutions of differential equations]. Numerische Mathematik, 10(3):261–288, October 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Leipnik:1971:RCR

[Lei71] R. B. Leipnik. Rapidly convergent recursive solution of quadratic operator equations. Numerische Mathematik, 17(1): 1–16, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lemou:1998:MEF

[Lem98] M. Lemou. Multipole expansions for the Fokker–Planck– Landau operator. Numerische Mathematik, 78(4):597–618, February 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/7078004/70780597.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/7078004/70780597.
                                        htm.

Lenferink:1989:CPE

[Len89] H. W. J. Lenferink. Contractivity preserving explicit linear multistep methods. Numerische Mathematik, 55(2):

213–223, May 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lesaint:1973:FEM

[Les73] P. Lesaint. Finite element methods for symmetric hyperbolic equations. Numerische Mathematik, 21(3):244–255, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Levin:1982:OQF

[Lev82] Moshe Levin. Optimal quadrature formula of Markov’s and Locher’s type with weight function. Numerische Mathematik, 40(1):31–37, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Levrie:1989:IMC

[Lev89]  Paul Levrie.  Improving a

method for computing nondominant solutions of certain second-order recurrence relations of Poincaré-type. Numerische Mathematik, 56(5): 501–512, November 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lewin:1971:EPM

[Lew71]  Mordechai Lewin.  On ex-

ponents of primitive matrices. Numerische Mathematik, 18(2): 154–161, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lewis:1982:ITM

[Lew82] Jerry W. Lewis. Inversion of tridiagonal matrices. Numerische Mathematik, 38(3): 333–345, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Leyk:1986:CLM

[Ley86]  Zbigniew Leyk.  A C^{0}-

collocation-like method for twopoint boundary value problems. Numerische Mathematik, 49(1): 39–53, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Leyk:1990:GMO

[Ley90]  Zbigniew Leyk.  Galerkin

method and optimal error algorithms in Hilbert spaces. Numerische Mathematik, 57(1): 1–14, March 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LaBudde:1976:EMCa

[LG76a] Robert A. LaBudde and Donald Greenspan. Energy and momentum conserving methods of arbitrary order for the numerical integration of equations of motion. I. motion of a single particle. Numerische Mathematik, 25(4):323–346, December 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LaBudde:1976:EMCb

[LG76b] Robert A. LaBudde and Donald Greenspan. Energy and momentum conserving methods of arbitrary order for the numerical integration of equations of motion. II. motion of a system of particles. Numerische Mathematik, 26(1):1–16, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lyness:1982:QQR

[LG82] J. N. Lyness and Luigi Gatteschi. On quasidegree quadrature rules. Numerische Mathematik, 39(2):259–267, August 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lieu:1966:MCE

[LH66] Bui Trong Liêü and P. Huard.

La méthode des centres dans un espace topologique. (French) [The method of centers in a topological space]. Numerische Mathematik, 8(1):56–67, March 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lou:1988:PPM

[LH88] Gang Lou and Shih-Ping Han.

A parallel projection method for solving generalized linear least-squares problems. Numerische Mathematik, 53(3): 255–264, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:1986:NCM

[Li86] Zi-Cai Li. A nonconforming combined method for solving Laplace’s boundary value problems with singularities. Numerische Mathematik, 49(5): 475–497, September 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Li87].

Li:1987:ENC

[Li87] Zi-Cai Li. Erratum: “A nonconforming combined method for solving Laplace’s boundary value problems with singularities”. Numerische Mathematik, 51(5):597, October 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Li86].

Li:1990:DTB

[Li90] Likang Li. Discretization of the Timoshenko beam problem by the p and the h-p versions of the finite element method. Numerische Mathematik, 57(4): 413–420, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:1995:NMC

[Li95]  Zhiping Li.  A numerical

method for computing singular minimizers. Numerische Mathematik, 71(3):317–330, September 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071003/50710317.
htm;   http://science.springer.
de/nmee/bibs/5071003/50710317.
htm.

Li:1997:SNA

[Li97] Zhiping Li. Simultaneous numerical approximation of microstructures and relaxed minimizers. Numerische Mathematik, 78(1):21–38, November 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.de/link/service/
journals/00211/bibs/7078001/
70780021.htm;   http://science.
springer.de/nmee/bibs/7078001/
70780021.htm.

Lie:1990:SFM

[Lie90] Ivar Lie. The stability function for multistep collocation methods. Numerische Mathematik, 57(8):779–787, August 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Light:1990:CSC

[Lig90] W. A. Light. The Chebyshev solution of certain matrix equations. Numerische Mathematik, 57(8):789–803, August 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Liniger:1961:MDH

[Lin61] Werner Liniger. On a method by D. H. Lehmer for the generation of pseudo random numbers. Numerische Mathematik, 3:265–270, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lindenberg:1962:UDG

[Lin62] Wolfgang Lindenberg. Über eine Darstellung von Gruppenelementen in digitalen Rechenautomaten. (German) [On a representation of group elements in computers]. Numerische Mathematik, 4:151–153, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Linz:1975:CDM

[Lin75] Peter Linz. Convergence of a discretization method for integro-differential equations. Numerische Mathematik, 25(1): 103–107, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Linden:1976:OSE

[Lin76]  Hanjörg Linden.  Obere

Schranken für Eigenwerte von Eigenwertaufgaben der Form (λ^{2} − λA − B)x = 0. (German) [Upper bounds for eigenvalues of quadratic eigenvalue problems of the forms (λ^{2}λA − B)x = 0]. Numerische Mathematik, 26(1): 17–26, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lindae:1993:SMD

[Lin93] Andreas Lindae. On shooting methods for the discrete Helmholtz equation with constant coefficients. Numerische Mathematik, 64(1): 127–145, January 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Liu:1995:SAM

[Liu95] Yunkang Liu. Stability analysis of θ-methods for neutral functional-differential equations. Numerische Mathematik, 70(4):473–485, June 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5070004/
50700473.htm;   http://science.
springer.de/nmee/bibs/5070004/
50700473.htm.

Liu:1996:IEN

[Liu96]  Xiaobo Liu.  Interior esti-

mates for nonconforming finite element methods. Numerische Mathematik, 74(1): 49–67, July 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074001/60740049.
htm;   http://science.springer.
de/nmee/bibs/6074001/60740049.
htm.

Lyness:1978:GFS

[LK78] J. N. Lyness and P. Keast.

On p-generator fully symmetric quadrature rules. Numerische Mathematik, 30(1):51– 63, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lin:1980:EMF

[LL80] Qun Lin and Jia Quan Liu. Extrapolation method for Fredholm integral equations with non-smooth kernels. Numerische Mathematik, 35(4): 459–464, November 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Laydi:1985:EFR

[LL85] M. Laydi and P. Lesaint. Sur les éléments finis rationnels de Serendip de degré quelconque. (French) [On rational Serendip finite elements of arbitrary degree]. Numerische Mathematik, 46(2):175–187, May 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lorentz:1990:BHI

[LL90] G. G. Lorentz and R. A.

Lorentz. Bivariate Hermite interpolation and applications to algebraic geometry. Numerische Mathematik, 57(6/7): 669–680, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Laghchim-Lahlou:1995:QFE

[LLS95] M. Laghchim-Lahlou and P. Sablonnière. Quadrilateral finite elements of FVS type and class C^{ρ}. Numerische Mathematik, 70(2):

229–243, April 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070002/50700229.
htm;   http://science.springer.
de/nmee/bibs/5070002/50700229.
htm.

Lawton:1998:CMC

[LLS98] W. Lawton, S. L. Lee, and Zuowei Shen. Convergence of multidimensional cascade algorithm. Numerische Mathematik, 78(3):427–438, January 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7078003/
70780427.htm;   http://science.
springer.de/nmee/bibs/7078003/
70780427.htm.

Lyness:1966:VMS

[LM66] J. N. Lyness and C. B. Moler.

Van der Monde systems and numerical differentiation. Numerische Mathematik, 8(5):458– 464, August 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lyness:1969:GRM

[LM69] J. N. Lyness and C. B. Moler.

Generalized Romberg methods for integrals of derivatives. Numerische Mathematik, 14(1):1– 13, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lyness:1970:RTD

[LM70] J. N. Lyness and B. J. J. McHugh. On the remainder term in the N-dimensional Euler Maclaurin expansion. Numerische Mathematik, 15(4): 333–344, August 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lightbourne:1975:PSF

[LM75] J. H. Lightbourne, III and R. H. Martin, Jr. Projection seminorms and the field of values of linear operators. Numerische Mathematik, 24(2): 151–161, April 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Louis:1991:SPM

[LM91] A. K. Louis and P. Maaß. Smoothed projection methods for the moment problem. Numerische Mathematik, 59(3): 277–294, June 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lu:1998:FEM

[LM98] Hao Lu and Joseph Maubach. A finite element method and variable transformations for a forward-backward heat equation. Numerische Mathematik, 81(2):249–272, December 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/8081002/
                                        80810249.htm;   http://link.

springer.de/link/service/
journals/00211/papers/8081002/
80810249.pdf.

Li:1999:LTA

[LM99] Chi-Kwong Li and Roy Mathias.

The Lidskii-Mirsky-Wielandt theorem — additive and multiplicative versions. Numerische Mathematik, 81(3):377–413, January 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/9081003/90810377.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/9081003/90810377.
pdf.

Laval:1994:PAU

[LMGR94] G. Laval, S. Mas-Gallic, and P.-A. Raviart. Paraxial approximation of ultrarelativistic intense beams. Numerische Mathematik, 69(1):33– 60, November 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069001/40690033.
htm;   http://science.springer.
de/nmee/bibs/4069001/40690033.
htm.

Li:1985:RRR

[LMPY85] Tien-Yien Li, John Mallet-Paret, and James A. Yorke. Regularity results for real analytic homotopies. Numerische Mathematik, 46(1):43–50, March 1985.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Layton:1995:PAM

[LMR95] William J. Layton, Joseph M. Maubach, and Patrick J. Rabier. Parallel algorithms for maximal monotone operators of local type. Numerische Mathematik, 71(1):29–58, July 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071001/50710029.
htm;   http://science.springer.
de/nmee/bibs/5071001/50710029.
htm.

Ling:1977:CPA

[LMS77] William H. Ling, Harry W. McLaughlin, and Mary Lynn Smith. A computational procedure for the approximation of random functions. Numerische Mathematik, 27(3): 301–305, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lazarov:1984:CDS

[LMW84] Rajčo D. Lazarov, Vladimir L. Makarov, and Wilfried Weinelt. On the convergence of difference schemes for the approximation of solutions u ∈ W_{2}^{m} (m > 0.5) of elliptic equations with mixed derivatives. Numerische Mathematik, 44(2):223–232, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Larson:1999:AMF

[LN99] Mats G. Larson and A. Jonas Niklasson. Adaptive multilevel finite element approximations of semilinear elliptic boundary value problems. Numerische Mathematik, 84 (2):249–274, December 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9084002/90840249.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9084002/90840249.pdf.

Lubich:1998:HBR

[LO98] Christian Lubich and Alexander Ostermann. Hopf bifurcation of reaction-diffusion and Navier–Stokes equations under discretization. Numerische Mathematik, 81(1):53– 84, November 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8081001/80810053.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081001/80810053.
pdf.

Locher:1973:SQG

[Loc73] Franz Locher. Zur Struktur von Quadraturformeln. (German) [On the structure of quadrature formulas]. Numerische Mathematik, 20(4):317–326, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Locher:1980:DDM

[Loc80]  F. Locher.  Dividierte Dif-

ferenzen und Monotonie von Quadraturformeln. (German) [Divided differences and monotony of quadrature formulas]. Numerische Mathematik, 34(2): 99–109, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Locher:1993:STR

[Loc93] F. Locher. A stability test for real polynomials. Numerische Mathematik, 66(1):33–40, October 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lode:1970:NNE

[Lod70] W. Lode. Zum neuen Norm Entwurf DIN 1333 “Abbrechen und Runden von Dezimalzahlen”. (German) [Toward a new standard design DIN 1333 “calculation and rounding in decimal numbers”]. Numerische Mathematik, 14(3):313–316, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Loizou:1967:BIW

[Loi67] G. Loizou. Bounds for the inequality of Wielandt. Numerische Mathematik, 10(2): 142–146, July 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Loizou:1968:MCS

[Loi68] G. Loizou. On matrices with a checkerboard sign distribution. Numerische Mathematik, 12(5): 429–430, December 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lorenz:1977:IDP

[Lor77] Jens Lorenz. Zur Inversmonotonie diskreter Probleme. (German) [On the monotone kindproperty of discrete problems]. Numerische Mathematik, 27(2): 227–238, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lory:1980:EDC

[Lor80] Peter Lory. Enlarging the domain of convergence for multiple shooting by the homotopy method. Numerische Mathematik, 35(2):231–240, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lorentzen:1992:AOR

[Lor92] Lisa Lorentzen. Arithmetic operations with regular Cfractions. Numerische Mathematik, 61(4):489–499, May 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Louis:1979:ACF

[Lou79] Alfred Louis. Acceleration of convergence for finite element solutions of the Poisson equation. Numerische Mathematik, 33(1):43–53, March 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Louis:1986:IDP

[Lou86] Alfred K. Louis. Incomplete data problems in X-ray computerized tomography I. singular value decomposition of the limited angle transform. Numerische Mathematik, 48(3): 251–262, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lee:1978:ANS

[LP78] J. W. Lee and P. M. Prenter. An analysis of the numerical solution of Fredholm integral equations of the first kind. Numerische Mathematik, 30(1):1– 23, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lee:1995:RCB

[LP95] Chang-Ock Lee and Seymour V. Parter. On the rate of convergence of the k × k block, kline iterative methods: k → ∞. Numerische Mathematik, 71 (1):59–90, July 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071001/50710059.
htm;   http://science.springer.
de/nmee/bibs/5071001/50710059.
htm.

Laurent:1970:GAC

[LPDT70] P. J. Laurent and Pham-Dinh-Tuan. Global approximation of a compact set by elements of a convex set in a normed space. Numerische Mathematik, 15(2): 137–150, June 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lucas:1973:HOP

[LR73] Thomas R. Lucas and George W.

Reddien. A high order projection method for nonlinear two point boundary value problems. Numerische Mathematik, 20(4): 257–270, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LeTallec:1989:FEA

[LR89a] Patrick Le Tallec and Michel Ravachol. Finite element approximation of viscoelastic progressively incompressible flows. Numerische Mathematik, 56(1): 43–66, September 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:1989:HAS

[LR89b] T. Y. Li and N. H. Rhee. Homotopy algorithm for symmetric eigenvalue problems. Numerische Mathematik, 55(3): 265–280, June 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Louis:1989:IDP

[LR89c] Alfred K. Louis and Andreas Rieder. Incomplete data problems in X-ray computerized tomography. II. truncated projections and region-of-interest tomography. Numerische Mathematik, 56(4):371–383, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Levenberg:1993:GAI

[LR93] N. Levenberg and L. Reichel. A generalized ADI iterative method. Numerische Mathematik, 66(2):215–233, November 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lyness:1998:CIC

[LR98] J. N. Lyness and U. Rüde. Cubature of integrands containing derivatives. Numerische Mathematik, 78(3):439–461, January 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078003/70780439.
htm;   http://science.springer.
de/nmee/bibs/7078003/70780439.
htm.

Lanzkron:1991:CNC

[LRS91] Paul J. Lanzkron, Donald J. Rose, and Daniel B. Szyld. Convergence of nested classical iterative methods for linear systems. Numerische Mathematik, 58(7):685–702, February 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lynch:1964:DSP

[LRT64] R. Lynch, J. Rice, and D. Thomas. Direct solution of partial difference equations by tensor product methods. Numerische Mathematik, 6(1):185– 199, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lions:1993:SSV

[LRT93] P. L. Lions, E. Rouy, and A. Tourin. Shape-from-shading, viscosity solutions and edges. Numerische Mathematik, 64(3): 323–353, March 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:1995:EAP

[LRZ95] T. Y. Li, Noah H. Rhee, and Zhong Gang Zeng. An efficient and accurate parallel algorithm for the singular value problem of bidiagonal matrices. Numerische Mathematik, 69(3):283–301, January 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069003/50690283.
htm;   http://science.springer.
de/nmee/bibs/5069003/50690283.
htm.

Lambert:1966:GMM

[LS66] John D. Lambert and Brian Shaw. A generalisation of multistep methods for ordinary differential equations. Numerische Mathematik, 8(3):250–263, May 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Luther:1970:OCF

[LS70] H. A. Luther and H. G. Sierra. On the optimal choice of fourthorder Runge–Kutta formulas. Numerische Mathematik, 15(4): 354–358, August 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:1987:RRS

[LS87] T.-Y. Li and Tim Sauer. Regularity results for solving systems of polynomials by homotopy method. Numerische Mathematik, 50(3):283–289, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lubich:1992:TDP

[LS92] Ch. Lubich and R. Schneider. Time discretization of parabolic boundary integral equations. Numerische Mathematik, 63(4): 455–481, December 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lions:1995:CMF

[LS95] P.-L. Lions and P. E. Souganidis. Convergence of MUSCL and filtered schemes for scalar conservation laws and Hamilton– Jacobi equations. Numerische Mathematik, 69(4):441–470, February 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069004/50690441.
htm;   http://science.springer.
de/nmee/bibs/5069004/50690441.
htm.

Lamp:1985:FFT

[LSW85] U. Lamp, K. T. Schleicher, and W. L. Wendland. The fast Fourier transform and the numerical solution of onedimensional boundary integral equations. Numerische Mathematik, 47(1):15–38, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:1987:RPH

[LSY87] T.-Y. Li, Tim Sauer, and James A. Yorke. The random product homotopy and deficient polynomial systems. Numerische Mathematik, 51(5): 481–500, October 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lynn:1968:NSS

[LT68] M. Stuart Lynn and William P.

Timlake. The numerical solution of singular integral equations of potential theory. Numerische Mathematik, 11(1):77–98, January 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LeBailly:1998:OPG

[LT98a] B. Le Bailly and J. P. Thiran. Optimum parameters for the generalized ADI method. Numerische Mathematik, 80(3): 377–395, September 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8080003/
80800377.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8080003/
80800377.pdf;   http://link.
springer.de/link/service/
journals/00211/papers/8080003/
80800377.ps.gz.

Liu:1998:TON

[LT98b] Xu-Dong Liu and Eitan Tadmor. Third order nonoscillatory central scheme for hyperbolic conservation laws. Numerische Mathematik, 79(3): 397–425, May 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079003/80790397.
htm;   http://science.springer.
de/nmee/bibs/8079003/80790397.
htm.

Lee:1992:ATF

[LTT92] S. L. Lee, Roger C. E. Tan, and W. S. Tang. L_{2}-approximation by the translates of a function and related attenuation factors. Numerische Mathematik, 60(4): 549–568, January 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lu:1997:GWG

[Lu97] Hao Lu. Galerkin and weighted Galerkin methods for a forwardbackward heat equation. Numerische Mathematik, 75(3): 339–356, January 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7075003/
70750339.htm;   http://science.
springer.de/nmee/bibs/7075003/
70750339.htm.

Lubich:1982:RKT

[Lub82]  Ch. Lubich.  Runge–Kutta

theory for Volterra integrodifferential equations. Numerische Mathematik, 40(1): 119–135, November 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lubich:1988:CQDa

[Lub88a] C. Lubich. Convolution quadrature and discretized operational calculus. I. Numerische Mathematik, 52(2):129–145, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lubich:1988:CQDb

[Lub88b] C. Lubich. Convolution quadrature and discretized operational calculus. II. Numerische Mathematik, 52(4):413–425, April 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lubinsky:1988:PSE

[Lub88c] D. S. Lubinsky. Power series equivalent to rational functions: a shifting-origin Kronecker type theorem, and normality of Padé tables. Numerische Mathematik, 54(1):33–39, October 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lubich:1989:LIE

[Lub89a]  C. Lubich.  Linearly im-

plicit extrapolation methods for differential-algebraic systems. Numerische Mathematik, 55(2): 197–211, May 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lubinsky:1989:UCR

[Lub89b] D. S. Lubinsky. On uniform convergence of rational, Newton– Padé interpolants of type (n, n) with free poles as n → ∞. Numerische Mathematik, 55(3): 247–264, June 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lubich:1991:CMM

[Lub91] C. Lubich. On the convergence of multistep methods for nonlinear stiff differential equations. Numerische Mathematik, 58(8): 839–853, March 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Lub92b].

Lube:1992:SDF

[Lub92a]  G. Lube.  Streamline dif-

fusion finite element method for quasilinear elliptic problems. Numerische Mathematik, 61(3): 335–357, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lubich:1992:ECM

[Lub92b] C. Lubich. Erratum: “On the convergence of multistep methods for nonlinear stiff differential equations” [Numer. Math. 58 (1991), no. 8, 839–853 and MR 92d:65127]. Numerische Mathematik, 61(2):277–279, March 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Lub91].

Lubich:1994:MTD

[Lub94] Ch. Lubich. On the multistep time discretization of linear initial-boundary value problems and their boundary integral equations. Numerische Mathematik, 67(3): 365–389, April 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067003/40670365.
htm;   http://science.springer.
de/nmee/bibs/4067003/40670365.
htm.

Lucas:1970:GS

[Luc70] Thomas R. Lucas. A generalization of L-splines. Numerische Mathematik, 15(5): 359–370, September 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Luecke:1979:NPC

[Lue79]  Glenn R. Luecke.  A nu-

merical procedure for computing the Moore–Penrose inverse. Numerische Mathematik, 32(2): 129–137, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Luke:1961:RUF

[Luk61] Yudell L. Luke. The radius of univalence of the function exp z^{2} ∫_{0}^{z} exp (−t^{2})dt. Numerische Mathematik, 3:76–78, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lukas:1993:AOG

[Luk93] Mark A. Lukas. Asymptotic optimality of generalized crossvalidation for choosing the regularization parameter. Numerische Mathematik, 66(1):41– 66, October 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lund:1985:BTR

[Lun85] John Lund. Bessel transforms and rational extrapolation. Numerische Mathematik, 47(1):1– 14, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Luskin:1996:ALM

[Lus96]  Mitchell Luskin.  Approxi-

mation of a laminated microstructure for a rotationally invariant, double well energy density. Numerische Mathematik, 75(2):205–221, December 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6075002/60750205.
htm;   http://science.springer.
de/nmee/bibs/6075002/60750205.
htm.

LaPorte:1974:ESE

[LV74a] M. La Porte and J. Vignes.

Étude statistique des erreurs dans l’arithmétique des ordinateurs; application au contrôle des Résultats d’algorithmes numériques. (French) [Statistical study of errors in computer arithmetic: Application to the control of the results of numerical algorithms]. Numerische Mathematik, 23(1):63– 72, February 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LaPorte:1974:MND

[LV74b] M. La Porte and J. Vignes. Méthode numérique de Détection de la singularité d’une matrice. (French) [Numerical method of detection of matrix singularity]. Numerische Mathematik, 23(1):73–81, February 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LaPorte:1975:EIS

[LV75] M. La Porte and J. Vignes. Evaluation de l’incertitude sur la solution d’un système linéaire.

(French) [Evaluation of the uncertainty in the solution of a linear system]. Numerische Mathematik, 24(1):39–47, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

LeTallec:1984:MNP

[LV84] Patrick Le Tallec and Marina Vidrascu. Une méthode numérique pour les problèmes d’équilibre de corps hyperélastiques compressibles en grandes déformations. (French) [A numerical method for solving equilibrium problems of compressible hyperelastic bodies in large deformations]. Numerische Mathematik, 43(2): 199–224, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Li:1989:NSU

[LV89] Xiezhang Li and Richard S. Varga. A note on the SSOR and USSOR iterative methods applied to p-cyclic matrices. Numerische Mathematik, 56(2/3): 109–121, October 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Leimkuhler:1997:OIL

[LV97] Benedict J. Leimkuhler and Erik S. Van Vleck. Orthosymplectic integration of linear Hamiltonian systems. Numerische Mathematik, 77(2): 269–282, August 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077002/70770269.
htm;   http://science.springer.
de/nmee/bibs/7077002/70770269.
htm.

Lai:1982:EBN

[LW82] Hang Chin Lai and Pou Yah Wu. Error bounds of Newton type process on Banach spaces. Numerische Mathematik, 39(2): 175–193, August 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Light:1999:SDI

[LW99] Will Light and Henry Wayne.

Spaces of distributions, interpolation by translates of a basis function and error estimates. Numerische Mathematik, 81(3):415–450, January 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/9081003/90810415.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/9081003/90810415.
pdf.

Lyche:1972:CMM

[Lyc72]  Tom Lyche.  Chebyshevian

multistep methods for ordinary differential equations. Numerische Mathematik, 19(1):65– 75, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lynn:1963:SIM

[Lyn63] M. Stuart Lynn. Some inframax bounds for the spectral radii of splittings of H-matrices. Numerische Mathematik, 5(1): 152–174, December 1963. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lyness:1968:CSI

[Lyn68] J. N. Lyness. The calculation of Stieltjes’ integral. Numerische Mathematik, 12(4): 252–265, November 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lyness:1985:EME

[Lyn85]  J. N. Lyness.  The Euler

Maclaurin expansion for the Cauchy principal value integral. Numerische Mathematik, 46(4): 611–622, August 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lesaint:1980:CNW

[LZ80] P. Lesaint and M. Zlámal. Convergence of the nonconforming Wilson element for arbitrary quadrilateral meshes. Numerische Mathematik, 36(1):33– 52, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Magolu:1992:MBA

[Mag92] Monga-Made Magolu. Modified block-approximate factorization strategies. Numerische Mathematik, 61(1):91–110, February 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Maitre:1967:NCN

[Mai67] Jean-François Maitre. Norme composée et norme associée généralisée d’une matrice. (French) [Composed and associative generalized matrix norm]. Numerische Mathematik, 10(2): 132–141, July 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Maitre:1968:CNA

[Maî68] Jean-François Maître. Sur une classe de normes et l’analyse a posteriori d’un système linéaire. (French) [On a class of norms and the A posteriori analysis of a linear system ]. Numerische Mathematik, 12(2):106–110, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Makridakis:1992:MFE

[Mak92] Ch. G. Makridakis. On mixed finite element methods for linear elastodynamics. Numerische Mathematik, 61(2): 235–260, March 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Malyshev:1999:FND

[Mal99] Alexander N. Malyshev. A formula for the 2-norm distance from a matrix to the set of matrices with multiple eigenvalues. Numerische Mathematik, 83(3):443–454, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083003/90830443.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083003/90830443.pdf.

Mangasarian:1970:CSM

[Man70] O. L. Mangasarian. A convergent splitting of matrices. Numerische Mathematik, 15(4): 351–353, August 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mansfield:1972:VCC

[Man72] Lois Mansfield. On the variational characterization and convergence of bivariate splines. Numerische Mathematik, 20(2): 99–114, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mansfield:1974:HOC

[Man74] Lois Mansfield. Higher order compatible triangular finite elements. Numerische Mathematik, 22(2):89–97, April 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Manteuffel:1977:TIN

[Man77] Thomas A. Manteuffel. The Tchebychev iteration for nonsymmetric linear systems. Numerische Mathematik, 28(3): 307–327, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mannshardt:1978:OSM

[Man78a] Reinhold Mannshardt. One-step methods of any order for ordinary differential equations with discontinuous right-hand sides. Numerische Mathematik, 31(2): 131–152, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Manteuffel:1978:APE

[Man78b] Thomas A. Manteuffel. Adaptive procedure for estimating parameters for the nonsymmetric Tchebychev iteration. Numerische Mathematik, 31(2): 183–208, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mansfield:1981:FEM

[Man81a] Lois Mansfield. Finite element methods for nonlinear shell analysis. Numerische Mathematik, 37(1):121–131, May 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mansfield:1981:MSF

[Man81b] Lois Mansfield. On the multigrid solution of finite element equations with isoparametric elements. Numerische Mathematik, 37(3):423–432, July 1981.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mansfield:1983:AFE

[Man83] Lois Mansfield. Analysis of finite element methods for the nonlinear dynamic analysis of shells. Numerische Mathematik, 42(2): 213–235, November 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mansfield:1984:FES

[Man84] Lois Mansfield. On finite element subspaces over quadrilateral and hexahedral meshes for incompressible viscous flow problems. Numerische Mathematik, 45(2):165–172, November 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mandel:1985:MIM

[Man85] Jan Mandel. On multilevel iterative methods for integral equations of the second kind and related problems. Numerische Mathematik, 46(1): 147–157, March 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mandel:1990:BDS

[Man90] Jan Mandel. On block diagonal and Schur complement preconditioning. Numerische Mathematik, 58(1):79–93, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marcus:1962:ICC

[Mar62] Marvin Marcus. An inequality connecting the P -condition number and the determinant. Numerische Mathematik, 4:350– 353, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marsal:1965:VSE

[Mar65]  Dietrich Marsal.  Eine Ve-

rallgemeinerung der Summenformel von Euler und ihre Anwendung auf ausgewählte Probleme der numerischen Integration. (German) [A generalization of Euler’s summation formula and its use for selected problems of numerical integration]. Numerische Mathematik, 7(2):147–154, April 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marek:1969:RSC

[Mar69]  Ivo Marek.  On the row

sum criterion and the convergence of some iterative processes. Numerische Mathematik, 13(3):207–216, July 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marsaglia:1970:RCR

[Mar70a] George Marsaglia. Regularities in congruential random number generators. Numerische Mathematik, 16(1):8–10, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marsal:1970:ELI

[Mar70b] D. Marsal. Eingabelung der Lösungen von Integralgleichungen. (German) [Bifurcation of solutions of integral equations]. Numerische Mathematik, 15(4): 348–350, August 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Martensen:1970:RGQ

[Mar70c] Erich Martensen. Zur Restglieddartstellung der Gregoryschen Quadraturformel ungerader Ordnung. (German) [On the representation of the remainder of the Gregory quadrature formula of an odd series]. Numerische Mathematik, 15(3): 229–233, July 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Martensen:1973:DER

[Mar73]  Erich Martensen.  Darstel-

lung und Entwicklung des Restgliedes der Gregoryschen Quadraturformel mit Hilfe von Spline-Funktionen. (German) [Representation and development of the remainder term of Gregory’s quadrature formula using spline functions]. Numerische Mathematik, 21(1):70– 80, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marcowitz:1975:FAS

[Mar75]  U. Marcowitz.  Fehlerab-

schätzung bei Anfangswertaufgaben für Systeme von gewönlichen Differentialgleichungen mit Anwendungen auf das REENTRY-Problem. (German) [Error estimation in initial value problems for systems of ordinary differential equations with an application to the REENTRY problem]. Numerische Mathematik, 24(3):249–275, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marciniak:1984:ECA

[Mar84] Andrzej Marciniak. Energy conserving, arbitrary order numerical solutions of the N-body problem. Numerische Mathematik, 45(2):207–218, November 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marini:1987:EAB

[Mar87] Luisa Donatella Marini. Estimates from above and from below for the homogenized coefficients. Numerische Mathematik, 50(5):533–545, March 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marti:1990:SPC

[Mar90] Jürg T. Marti. Small potential corrections for the discrete Eigenvalues of the Sturm–Liouville problem. Numerische Mathematik, 57(1): 51–62, March 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Masmoudi:1987:NSE

[Mas87] M. Masmoudi. Numerical solution for exterior problems. Numerische Mathematik, 51(1): 87–101, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Matthys:1976:ALM

[Mat76] J. Matthys. A-stable linear multistep methods for Volterra integro-differential equations. Numerische Mathematik, 27(1): 85–94, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mattheij:1980:CDD

[Mat80] R. M. M. Mattheij. Characterizations of dominant and dominated solutions of linear recursions. Numerische Mathematik, 35(4):421–442, November 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Matos:1990:AMB

[Mat90] Ana C. Matos. Acceleration methods based on convergence tests. Numerische Mathematik, 58(3):329–340, October 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mathias:1992:EFD

[Mat92] Roy Mathias. Evaluating the Fréchet derivative of the matrix exponential. Numerische Mathematik, 63(2):213–226, November 1992. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Mathew:1993:SAIa

[Mat93a] Tarek P. Mathew. Schwarz alternating and iterative refinement methods for mixed formulations of elliptic problems. I. algorithms and numerical results. Numerische Mathematik, 65(4): 445–468, August 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mathew:1993:SAIb

[Mat93b] Tarek P. Mathew. Schwarz alternating and iterative refinement methods for mixed formulations of elliptic problems. II. convergence theory. Numerische Mathematik, 65(4):469–492, August 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mathias:1996:FAE

[Mat96] Roy Mathias. Fast accurate eigenvalue methods for graded positive definite matrices. Numerische Mathematik, 74(1): 85–103, July 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074001/60740085.
htm;   http://science.springer.
de/nmee/bibs/6074001/60740085.
htm.

Mavrides:1977:OUN

[Mav77] Lazaros P. Mavrides. Optimisation under nonnegativity constraints. Numerische Mathematik, 28(3):287–293, September 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Maybee:1965:SST

[May65] J. S. Maybee. Some structural theorems for partial difference operators. Numerische Mathematik, 7(1):66–72, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mayer:1985:CPI

[May85] Günter Mayer. On the convergence of powers of interval matrices. II. Numerische Mathematik, 46(1):69–83, March 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mayer:1986:TSR

[May86] Günter Mayer. On a theorem of Stein–Rosenberg type in interval analysis. Numerische Mathematik, 50(1):17–26, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

McArthur:1990:SMM

[MBL90] Kelly M. McArthur, Kenneth L. Bowers, and John Lund. The Sinc method in multiple space dimensions: model problems. Numerische Mathematik, 56(8): 789–816, February 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Barth:1967:CES

[MBW67] R. S. Martin, W. Barth, and J. H. Wilkinson. Calculation of the eigenvalues of a symmetric tridiagonal matrix by the method of bisection. Numerische Mathematik, 9(5):386– 393, April 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 249–256].

Modi:1984:AGO

[MC84] J. J. Modi and M. R. B. Clarke.

An alternative Givens ordering. Numerische Mathematik, 43(1): 83–90, January 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Monegato:1988:EIL

[MC88a] G. Monegato and V. Colombo.

Erratum: “Product integration for the linear transport equation in slab geometry” [Numer. Math. 52 (1988), no. 2, 219–240 and MR 88k:65133]. Numerische Mathematik, 53(6):739, September 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [MC88b].

Monegato:1988:PIL

[MC88b] G. Monegato and V. Colombo.

Product integration for the linear transport equation in slab geometry. Numerische Mathematik, 52(2):219–240, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [MC88a].

McAllister:1968:PBD

[McA68] G. T. McAllister. A priori bounds on difference quotients of solutions to some linear uniformly elliptic difference equations. Numerische Mathematik, 11(1):13–37, January 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

McCormick:1975:IPS

[McC75] Steve F. McCormick. An iterative procedure for the solution of constrained nonlinear equations with application to optimization problems. Numerische Mathematik, 23(5):371–385, October 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

McLachlan:1994:SIH

[McL94] Robert McLachlan. Symplectic integration of Hamiltonian wave equations. Numerische Mathematik, 66(4):465–492, January 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meinguet:1977:OAE

[MD77] Jean Meinguet and Jean Descloux. An operator-theoretical approach to error estimation. Numerische Mathematik, 27(3): 307–326, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Medley:1971:NPJ

[Med71]  Helen I. Medley.  A note

on a paper of Johnston and Smith’s concerning best isolated Gerschgorin disks. Numerische Mathematik, 16(5): 435–441, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Medley:1973:NGF

[Med73] H. I. Medley. A note on Ggenerating families and isolated Gerschgorin disks. Numerische Mathematik, 21(2):93–95, April 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mehndiratta:1967:SSD

[Meh67] S. L. Mehndiratta. Symmetry and self-duality in nonlinear programming. Numerische Mathematik, 10(2): 103–109, July 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meinardus:1963:UVU

[Mei63] Günter Meinardus. Über eine Verallgemeinerung einer Ungleichung von L. V. Kantorowitsch. (German) [On a generalization of an inequation of L. V. Kantorowitsch]. Numerische Mathematik, 5(1):14–23, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meinguet:1966:MER

[Mei66] J. Meinguet. Methods for estimating the remainder in linear rules of approximation. Application to the Romberg algorithm. Numerische Mathematik, 8(4):345–366, June 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meinguet:1967:OAE

[Mei67]  J. Meinguet.  Optimal ap-

proximation and error bounds in seminormed spaces. Numerische Mathematik, 10(4): 370–388, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meinguet:1986:DKW

[Mei86] Jean Meinguet. On the Davis-Kahan-Weinberger solution of the norm-preserving dilation problem. Numerische Mathematik, 49(2/3):331–341, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meini:1997:NCR

[Mei97] Beatrice Meini. New convergence results on functional iteration techniques for the numerical solution of M/G/1 type Markov chains. Numerische Mathematik, 78(1):39– 58, November 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.

de/link/service/journals/
00211/bibs/7078001/70780039.
htm;   http://science.springer.
de/nmee/bibs/7078001/70780039.
htm.

Melkes:1972:RPB

[Mel72]  František Melkes.  Reduced

piecewise bivariate Hermite interpolations. Numerische Mathematik, 19(4):326–340, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Melman:1995:NSS

[Mel95] A. Melman. Numerical solution of a secular equation. Numerische Mathematik, 69(4): 483–493, February 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5069004/
50690483.htm;   http://science.
springer.de/nmee/bibs/5069004/
50690483.htm.

Melenk:1999:OAS

[Mel99] J. M. Melenk. Operator adapted spectral element methods I: harmonic and generalized harmonic polynomials. Numerische Mathematik, 84(1):35–69, November 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9084001/90840035.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9084001/90840035.pdf.

Menke:1974:BNK

[Men74] Klaus Menke. Bestimmung von Näherungen für die konforme Abbildung mit Hilfe von stationären Punksystemen. (German) [Determination of the best approximation for conformal mapping with help of stationary point system]. Numerische Mathematik, 22(2): 111–117, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Menke:1988:PSE

[Men88]  Klaus Menke.  Point sys-

tems with extremal properties and conformal mapping. Numerische Mathematik, 54(2): 125–143, November 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Merzrath:1967:DSP

[Mer67] Ernst Merzrath. Direct solution of partial difference equations. Numerische Mathematik, 9(5):431–436, April 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mertins:1991:KRR

[Mer91] Ulrich Mertins. Zur Konvergenz des Rayleigh–Ritz-Verfahrens bei Eigenwertaufgaben. (German) [On the convergence of the Rayleigh–Ritz method for eigenvalue problems]. Numerische Mathematik, 59(7):667–682, October 1991. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Mertins:1992:AFF

[Mer92]  U. Mertins.  Asymptotische

Fehlerschranken für Rayleigh– Ritz-Approximationen selbstadjungierter Eigenwertaufgaben. (German) [Asymptotic error estimates for Rayleigh–Ritz approximations of selfadjoint eigenvalue problems]. Numerische Mathematik, 63(2): 227–241, November 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Messaoudi:1995:SIF

[Mes95] A. Messaoudi. On the stability of the incomplete LUfactorizations and characterizations of H-matrices. Numerische Mathematik, 69(3):321–331, January 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069003/50690321.
htm;   http://science.springer.
de/nmee/bibs/5069003/50690321.
htm.

Metropolis:1965:AUA

[Met65a]  N. Metropolis.  Algorithms

in unnormalized arithmetic. I: Recurrence relations. Numerische Mathematik, 7(2):104– 112, April 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Met65b].

Metropolis:1965:BAU

[Met65b] N. Metropolis. Berichtigung: Algorithms in unnormalized arithmetic. I: Recurrence relations. Numerische Mathematik, 7(4):354, October 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Met65a].

Meyer:1970:FIP

[Mey70] G. H. Meyer. On a free interface problem for linear ordinary differential equations and the one-phase Stefan problem. Numerische Mathematik, 16(3): 248–267, December 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyer:1975:SAS

[Mey75] G. G. L. Meyer. A systematic approach to the synthesis of algorithms. Numerische Mathematik, 24(4):277–289, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyer:1978:MLP

[Mey78] Gunter H. Meyer. The method of lines for Poisson’s equation with nonlinear or free boundary conditions. Numerische Mathematik, 29(3):329–344, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyer-Spasche:1979:NAM

[Mey79] Rita Meyer-Spasche. A note on the approximation of mildly nonlinear Dirichlet problems by finite differences. Numerische Mathematik, 33(3): 303–313, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyn:1983:SUN

[Mey83] Klaus-Hermann Meyn. Solution of underdetermined nonlinear equations by stationary iteration methods. Numerische Mathematik, 42(2): 161–172, November 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyer:1984:FBP

[Mey84a] Gunter H. Meyer. Free boundary problems with nonlinear source terms. Numerische Mathematik, 43(3):463–483, April 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyer:1984:MNV

[Mey84b] Peter Wilhelm Meyer. Das modifizierte Newton-Verfahren in verallgemeinerten Banach-Räumen. (German) [The modified Newton method in generalized Banach spaces]. Numerische Mathematik, 43(1):91– 104, January 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mitchell:1964:IFA

[MF64] A. Mitchell and G. Fairweather.

Improved forms of the alternating direction methods of Douglas, Peaceman, and Rachford for solving parabolic and elliptic equations. Numerische Mathematik, 6(1):285–292, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Molchanov:1985:FEM

[MG85] I. N. Molchanov and Ē. F. Galba. On finite element methods for the Neumann problem. Numerische Mathematik, 46(4): 587–598, August 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mas-Gallic:1987:PMF

[MGR87] S. Mas-Gallic and P. A. Raviart. A particle method for firstorder symmetric systems. Numerische Mathematik, 51(3): 323–352, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morita:1972:CAG

[MH72] Tohru Morita and Tsuyoshi Horiguchi. Convergence of the arithmetic-geometric mean procedure for the complex variables and the calculation of the complete elliptic integrals with complex modulus. Numerische Mathematik, 20(5): 425–430, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-

599X&volume=20&issue=5&spage=
425.

Murota:1982:PTR

[MI82] Kazuo Murota and Masao Iri.

Parameter tuning and repeated application of the IMT-type transformation in numerical quadrature. Numerische Mathematik, 38(3):347–363, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Midy:1975:ICG

[Mid75] P. Midy. An improved calculation of the general elliptic integral of the second kind in the neighbourhood of x = 0. Numerische Mathematik, 25(1): 99–101, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miel:1979:UEA

[Mie79] George J. Miel. Unified error analysis for Newton-type methods. Numerische Mathematik, 33(4):391–396, December 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mignotte:1976:NMB

[Mig76] Maurice Mignotte. Note sur la méthode Bernoulli. (French) [A note on Bernoulli’s method]. Numerische Mathematik, 26(3): 325–326, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miklosko:1970:APC

[Mik70] Jozef Mikloško. Asymptotic properties and the convergence of numerical quadratures. Numerische Mathematik, 15(3): 234–249, July 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miller:1965:ASF

[Mil65a] G. F. Miller. Algorithms for special functions. II. Numerische Mathematik, 7(2):194– 196, April 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miller:1965:NAS

[Mil65b] Keith Miller. Numerical analogs to the Schwarz alternating procedure. Numerische Mathematik, 7(2):91–103, April 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miller:1967:PBO

[Mil67] John J. H. Miller. On powerbounded operators and operators satisfying a resolvent condition. Numerische Mathematik, 10(5):389–396, November 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miller:1972:SFN

[Mil72] Webb Miller. On the stability of finite numerical procedures. Numerische Mathematik, 19(5): 425–432, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miller:1977:RAS

[Mil77] Webb Miller. Roundoff analysis and sparse data. Numerische Mathematik, 29(1):37– 43, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Milaszewicz:1980:GSR

[Mil80] J. P. Milaszewicz. A generalization of the Stein–Rosenberg theorem to Banach spaces. Numerische Mathematik, 34(4): 403–409, December 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Milaszewicz:1981:FPL

[Mil81] J. P. Milaszewicz. On fixed point linear equations. Numerische Mathematik, 38(1):53–59, October 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Milner:1985:PHF

[Mil85] Fabio A. Milner. A primal hybrid finite element method for quasilinear second order elliptic problems. Numerische Mathematik, 47(1):107–122, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miranker:1971:DSB

[Mir71] W. L. Miranker. Difference schemes with best possible truncation error. Numerische Mathematik, 17(2):124–142, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Missirlis:1984:CTE

[Mis84] Nikolaos M. Missirlis. Convergence theory of extrapolated iterative methods for a certain class of non-symmetric linear systems. Numerische Mathematik, 45(3):447–458, December 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mishev:1999:FVE

[Mis99] Ilya D. Mishev. Finite volume element methods for nondefinite problems. Numerische Mathematik, 83(1):161–175, July 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083001/90830161.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083001/90830161.pdf.

Mittermayer:1973:IMG

[Mit73] E. Mittermayer. Eine Iterationsmethode zur mehrdimensionalen gleichmäßigen Approximation bei gegebener Näherung. (German) [An iterative method for the more dimensional uniform approximation by given nearness]. Numerische Mathematik, 21(4):298–302, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mittelmann:1978:ASN

[Mit78] Hans-Detlef Mittelmann. On the approximate solution of nonlinear variational inequalities. Numerische Mathematik, 29(4): 451–462, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mittelmann:1980:ESN

[Mit80] Hans-Detlef D. Mittelmann. On the efficient solution of nonlinear finite element equations. I. Numerische Mathematik, 35(3): 277–291, October 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mittelmann:1981:ESN

[Mit81] Hans Detlef Mittelmann. On the efficient solution of nonlinear finite element equations. II. bound-constrained problems. Numerische Mathematik, 36(4): 375–387, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miyoshi:1976:MFE

[Miy76] Tetsuhiko Miyoshi. A mixed finite element method for the solution of the von Karman equations. Numerische Mathematik, 26(3):255–269, September 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Magnus:1960:ZHF

[MK60] Wilhelm Magnus and Leon Kotin. The zeros of the Hankel function as a function of its order. Numerische Mathematik, 2(1):228–244, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mokhtar-Kharroubi:1980:CTM

[MK80] H. Mokhtar-Kharroubi. Sur la convergence théorique de la méthode du gradient réduit généralisé. (French) [On the theoretical convergence of the reduced generalized gradient method]. Numerische Mathematik, 34(1):73–85, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Monegato:1990:NRG

[ML90] G. Monegato and P. Lepora. On the numerical resolution of the generalized airfoil equation with Possion kernel. Numerische Mathematik, 56(8): 775–787, February 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Monegato:1998:EEF

[ML98] G. Monegato and J. N. Lyness. The Euler–Maclaurin expansion and finite-part integrals. Numerische Mathematik, 81(2):273–291, December 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8081002/
80810273.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8081002/
80810273.pdf.

Menzel:1967:AUA

[MM67] M. Menzel and N. Metropolis. Algorithms in unnormalized arithmetic II. unrestricted polynomial evaluation. Numerische Mathematik, 10(5): 451–462, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Micchelli:1974:AOA

[MM74] C. A. Micchelli and W. L. Miranker. Asymptotically optimal approximation in fractional Sobolev spaces and the numerical solution of differential equations. Numerische Mathematik, 22(2):75–87, April 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morton:1992:OEE

[MMS92] K. W. Morton, T. Murdoch, and E. Süli. Optimal error estimation for Petrov–Galerkin methods in two dimensions. Numerische Mathematik, 61(3): 359–372, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Manas:1968:FAV

[MN68] Miroslav Maňas and Josef Nedoma. Finding all vertices of a convex polyhedron. Numerische Mathematik, 12(3):226–229, October 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miller:1985:NCT

[MN85] Valerie A. Miller and Michael Neumann. A note on comparison theorems for nonnegative matrices. Numerische Mathematik, 47(3):427–434, November 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Makela:1974:CCC

[MNS74] Matti Mäkelä, Olavi Nevanlinna, and Aarne H. Sipilä. On the concepts of convergence, consistency, and stability in connection with some numerical methods. Numerische Mathematik, 22(4):261–274, August 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marshall:1964:ITE

[MO64] A. W. Marshall and I. Olkin. Inclusion theorems for eigenvalues from probability inequalities. Numerische Mathematik, 6(1):98–102, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marshall:1968:SMA

[MO68] Albert W. Marshall and Ingram Olkin. Scaling of matrices to achieve specified row and column sums. Numerische Mathematik, 12(1):83–90, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Majda:1978:SAC

[MO78] Andrew Majda and Stanley Osher. A systematic approach for correcting nonlinear instabilities — the Lax–Wendroff scheme for scalar conservation laws. Numerische Mathematik, 30(4):429–452, December 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Moret:1991:QNM

[MO91] Igor Moret and Pierpaolo Omari. A quasi-Newton method for solving fixed point problems in Hilbert spaces. Numerische Mathematik, 59(2): 159–177, May 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mock:1974:SNT

[Moc74] M. S. Mock. Solution of a nonlinear two point boundary value problem. Numerische Mathematik, 23(1):37–46, February 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mock:1975:GPE

[Moc75] M. S. Mock. A global a posteriori error estimate for quasilinear elliptic problems. Numerische Mathematik, 24(1):53– 61, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mock:1976:EFE

[Moc76] M. S. Mock. Explicit finite element schemes for first order symmetric hyperbolic systems. Numerische Mathematik, 26(4): 367–378, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Moller:1976:KMK

[Möl76]  H. M. Möller.  Kubatur-

formeln mit minimaler Knotenzahl. (German) [Minimumpoint cubature formulas]. Numerische Mathematik, 25(2): 185–200, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Monegato:1984:CPF

[Mon84] Giovanni Monegato. Convergence of product formulas for the numerical evaluation of certain two-dimensional Cauchy principal value integrals. Numerische Mathematik, 43(2): 161–173, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Monegato:1987:WCQ

[Mon87] Giovanni Monegato. On the weights of certain quadratures for the numerical evaluation of Cauchy principal value integrals and their derivatives. Numerische Mathematik, 50(3): 273–281, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Monk:1992:FEM

[Mon92] Peter Monk. A finite element method for approximating the time-harmonic Maxwell equations. Numerische Mathematik, 63(2):243–261, November 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Moore:1968:ANO

[Moo68] Robert H. Moore. Approximations to nonlinear operator equations and Newton’s method. Numerische Mathematik, 12(1):23–34, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Moore:1988:DCG

[Moo88] Gerald Moore. Defect correction from a Galerkin viewpoint. Numerische Mathematik, 52(5): 565–582, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morimoto:1962:SWE

[Mor62] Hirofumi Morimoto. Stability in the wave equation coupled with heat flow. Numerische Mathematik, 4:136–145, December 1962. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Morris:1972:ITS

[Mor72] Wm. L. Morris. Inclusion theorems for a section of a matrix. Numerische Mathematik, 18(5): 457–464, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morel:1974:PIV

[Mor74] Paul Morel. Sur le problème inverse des valeurs propres. (French) [On the inverse eigenvalue problem]. Numerische Mathematik, 23(1):83–94, February 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morita:1978:CCE

[Mor78]  Tohru Morita.  Calculation

of the complete elliptic integrals with complex modulus. Numerische Mathematik, 29(2): 233–236, January 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mori:1980:ARS

[Mor80] Masatake Mori. Analytic representations suitable for numerical computation of some special functions. Numerische Mathematik, 35(2):163–174, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morley:1989:FMF

[Mor89] Mary E. Morley. A family of mixed finite elements for linear elasticity. Numerische Mathematik, 55(6):633–666, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Moret:1994:AGT

[Mor94]  Igor Moret.  The approxi-

mation of generalized turning points by Krylov subspace methods. Numerische Mathematik, 68(3):341–353, September 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068003/40680341.
htm;   http://science.springer.
de/nmee/bibs/4068003/40680341.
htm.

Morita:1999:CEI

[Mor99] T. Morita. Calculation of the elliptic integrals of the first and second kinds with complex modulus. Numerische Mathematik, 82(4):677–688, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082004/90820677.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082004/90820677.pdf.

Modi:1985:EIJ

[MP85] J. J. Modi and J. D. Pryce. Efficient implementation of Jacobi’s diagonalization method on the DAP. Numerische Mathematik, 46(3):443–454, June 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miesbach:1992:SPF

[MP92] S. Miesbach and H. J. Pesch. Symplectic phase flow approximation for the numerical integration of canonical systems. Numerische Mathematik, 61(4): 501–521, May 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marini:1995:MFE

[MP95a] Donatella Marini and Paola Pietra. Mixed finite element approximation of a degenerate elliptic problem. Numerische Mathematik, 71(2): 225–236, August 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071002/50710225.
htm;   http://science.springer.
de/nmee/bibs/5071002/50710225.
htm.

Marletta:1995:LSL

[MP95b] Marco Marletta and John D. Pryce. LCNO Sturm–Liouville problems: computational difficulties and examples. Numerische Mathematik, 69(3): 303–320, January 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5069003/
50690303.htm;   http://science.
springer.de/nmee/bibs/5069003/
50690303.htm.

Maitre:1996:CND

[MP96a] Jean-François Maitre and Olivier Pourquier. Condition number and diagonal preconditioning: comparison of the p-version and the spectral element methods. Numerische Mathematik, 74(1): 69–84, July 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074001/60740069.
htm;   http://science.springer.
de/nmee/bibs/6074001/60740069.
htm.

McLean:1996:BEC

[MP96b] W. McLean and S. B. Prößdorf.

Boundary element collocation methods using splines with multiple knots. Numerische Mathematik, 74(4):419–451, October 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074004/60740419.
htm;   http://science.springer.
de/nmee/bibs/6074004/60740419.
htm.

Mortreux:1999:TMF

[MP99] P. Mortreux and M. Prévost.

On totally monotonic families of sequences. Numerische Mathematik, 84(1):71–96, November 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9084001/90840071.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9084001/90840071.pdf.

Markowich:1999:NAQ

[MPP99] Peter A. Markowich, Paola Pietra, and Carsten Pohl. Numerical approximation of quadratic observables of Schrödingertype equations in the semiclassical limit. Numerische Mathematik, 81(4):595–630, February 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer-
                                        ny.com/link/service/journals/
                                        00211/bibs/9081004/90810595.
                                        htm;   http://link.springer-
                                        ny.com/link/service/journals/
                                        00211/papers/9081004/90810595.
                                        pdf.

Martin:1965:SDPb

[MPW65] R. S. Martin, G. Peters, and J. H. Wilkinson. Symmetric decomposition of a positive definite matrix. Numerische Mathematik, 7(5):362–383, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 9–30].

Martin:1966:IRS

[MPW66] R. S. Martin, G. Peters, and J. H. Wilkinson. Iterative refinement of the solution of a positive definite system of equations. Numerische Mathematik, 8(3):203–216, May 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 31– 44].

Martin:1970:ARH

[MPW70] R. S. Martin, G. Peters, and J. H. Wilkinson. The QR algorithm for real Hessenberg matrices. Numerische Mathematik, 14(3):219–231, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 359– 371].

Maday:1981:LCS

[MQ81] Yvon Maday and Alfio Quarteroni. Legendre and Chebyshev spectral approximations of Burgers’ equation. Numerische Mathematik, 37(3): 321–332, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marini:1989:RPD

[MQ89] L. D. Marini and A. Quarteroni. A relaxation procedure for domain decomposition methods using finite elements. Numerische Mathematik, 55(5): 575–598, August 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Motta:1990:PAS

[MRMW90] Santo Motta, Giovanni Russo, Hardy Moock, and Joachim Wick. Point approximation of a space-homogeneous transport equation. Numerische Mathematik, 56(8):763–774, February 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mejia:1995:PCC

[MRS95] Diego Mejía, Stephan Ruscheweyh, and Luis C. Salinas. Polynomials on the Cauchy circle. Numerische Mathematik, 69(4): 471–481, February 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/5069004/
                                        50690471.htm;   http://science.
                                        springer.de/nmee/bibs/5069004/
                                        50690471.htm.

Martin:1968:HTS

[MRW68] R. S. Martin, C. Reinsch, and J. H. Wilkinson. Householder’s tridiagonalization of a symmetric matrix. Numerische Mathematik, 11(3):181–195, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 212–226].

Martin:1970:ABS

[MRW70] R. S. Martin, C. Reinsch, and J. H. Wilkinson. The QR algorithm for band symmetric matrices. Numerische Mathematik, 16(2):85–92, November 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See also [WR71, pp. 266–272].

Micchelli:1976:ORS

[MRW76] C. A. Micchelli, T. J. Rivlin, and S. Winograd. The optimal recovery of smooth functions. Numerische Mathematik, 26(2):191–200, June 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

McNamee:1967:CFS

[MS67] J. McNamee and F. Stenger.

Construction of fully symmetric numerical integration formulas. Numerische Mathematik, 10(4): 327–344, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyer-Spasche:1972:CMS

[MS72] Rita Meyer-Spasche. A constructive method of solving the Lyapunov equation for complex matrices. Numerische Mathematik, 19(5):433–438, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyer-Spasche:1973:MSS

[MS73] R. Meyer-Spasche. A method of solving the stability-problem for complex matrices. Numerische Mathematik, 20(5):364–371, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Menzel:1978:LPN

[MS78] Reinhard Menzel and Hubert Schwetlick. Zur Lösung parameterabhängiger nichtlinearer Gleichungen mit singulären Jacobi-Matrizen. (German) [On solving nonlinear equations depending on a real parameter in case of singular Jacobians]. Numerische Mathematik, 30(1):65– 79, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Menzel:1985:PSL

[MS85] Reinhard Menzel and Hubert Schwetlick. Parametrization via secant length and application to path following. Numerische Mathematik, 47(3): 401–412, November 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marek:1990:CTW

[MS90] Ivo Marek and Daniel B. Szyld. Comparison theorems for weak splittings of bounded operators. Numerische Mathematik, 58(4): 387–397, November 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mills:1992:LCL

[MS92] T. M. Mills and S. J. Smith. The Lebesgue constant for Lagrange interpolation on equidistant nodes. Numerische Mathematik, 61(1):111–115, February 1992. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Marek:1994:LCE

[MS94] Ivo Marek and Daniel B. Szyld.

Local convergence of the (exact and inexact) iterative aggregation method for linear systems and Markov operators. Numerische Mathematik, 69(1):61– 82, November 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069001/40690061.
htm;   http://science.springer.
de/nmee/bibs/4069001/40690061.
htm.

Moller:1995:MPE

[MS95a] H. Michael Möller and Hans J.

Stetter. Multivariate polynomial equations with multiple zeros solved by matrix eigenproblems. Numerische Mathematik, 70(3): 311–329, May 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070003/50700311.
htm;   http://science.springer.
de/nmee/bibs/5070003/50700311.
htm.

Morton:1995:EGM

[MS95b] K. W. Morton and E. Süli.

Evolution-Galerkin methods and their supraconvergence. Numerische Mathematik, 71(3): 331–355, September 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5071003/
50710331.htm;   http://science.
springer.de/nmee/bibs/5071003/
50710331.htm.

Manteuffel:1996:HIM

[MS96] Thomas A. Manteuffel and Gerhard Starke. On hybrid iterative methods for nonsymmetric systems of linear equations. Numerische Mathematik, 73(4): 489–506, June 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073004/60730489.
htm;   http://science.springer.
de/nmee/bibs/6073004/60730489.
htm.

McDonald:1998:BLF

[MS98] J. J. McDonald and H. Schneider. Block LU factorizations of M-matrices. Numerische Mathematik, 80(1): 109–130, July 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080001/80800109.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800109.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080001/80800109.
ps.gz.

Mulansky:1999:CII

[MS99] Bernd Mulansky and Jochen W.

Schmidt. Convex interval interpolation using a three-term staircase algorithm. Numerische Mathematik, 82(2):313–337, April 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082002/90820313.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082002/90820313.pdf.

Meyer-Spasche:1991:DEF

[MSF91] Rita Meyer-Spasche and Bengt Fornberg. Discretization errors at free boundaries of the Grad-Schlüter-Shafranov equation. Numerische Mathematik, 59(7):683–710, October 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Metropolis:1967:SSN

[MSS67] N. Metropolis, M. L. Stein, and P. R. Stein. Stable states of a non-linear transformation. Numerische Mathematik, 10(1): 1–19, May 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Moss:1988:NEA

[MSW88] William F. Moss, Philip W.

Smith, and Joseph D. Ward. Nonlinear eigenvalue approximation. Numerische Mathematik, 52(4):365–375, April 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morgan:1991:CSS

[MSW91] Alexander P. Morgan, Andrew J. Sommese, and Charles W. Wampler. Computing singular solutions to non-linear analytic systems. Numerische Mathematik, 58(7):669–684, February 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morgan:1992:PSM

[MSW92] Alexander P. Morgan, Andrew J. Sommese, and Charles W. Wampler. A power series method for computing singular solutions to nonlinear analytic systems. Numerische Mathematik, 63(3):391–409, December 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mastin:1978:TTR

[MT78] C. Wayne Mastin and Joe F. Thompson. Transformation of three-dimensional regions onto rectangular regions by elliptic systems. Numerische Mathematik, 29(4):397–407, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

More:1989:ABC

[MT89] Jorge J. Moré and Gerardo Toraldo. Algorithms for bound constrained quadratic programming problems. Numerische Mathematik, 55(4): 377–400, July 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marion:1990:NGM

[MT90] M. Marion and R. Temam.

Nonlinear Galerkin methods: The finite elements case. Numerische Mathematik, 57(3): 205–226, May 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mandel:1996:CSM

[MT96] Jan Mandel and Radek Tezaur.

Convergence of a substructuring method with Lagrange multipliers. Numerische Mathematik, 73 (4):473–487, June 1996. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/6073004/
60730473.htm;   http://science.
springer.de/nmee/bibs/6073004/
60730473.htm.

Mueller:1966:HMC

[Mue66] Dennis J. Mueller. Householder’s method for complex matrices and eigensystems of Hermitian matrices. Numerische Mathematik, 8(1):72– 92, March 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Muhlbach:1978:GNA

[Müh78]  G. Mühlbach.  The gen-

eral Neville–Aitken-algorithm and some applications. Numerische Mathematik, 31(1):97– 110, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Muhlbach:1979:GRR

[Müh79] G. Mühlbach. The general recurrence relation for divided differences and the general Newton-interpolation-algorithm with applications to trigonometric interpolation. Numerische Mathematik, 32(4): 393–408, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Muhlbach:1981:AAH

[Müh81] G. Mühlbach. An algorithmic approach to Hermite– Birkhoff interpolation. Numerische Mathematik, 37(3): 339–347, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Muhlbach:1985:TCM

[Müh85] G. Mühlbach. Two composition methods for solving certain systems of linear equations. Numerische Mathematik, 46(3): 339–349, June 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mueller:1968:ELA

[Mül68] Heinrich N. Müller. Zur Eindeutigkeit bei der linearen Approximation im allgemeinen sowie Tschebyscheffschen Fall. (German) [On the uniqueness of the linear approximation in general as well as Chebyshev cases]. Numerische Mathematik, 11(2): 144–150, February 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Multhei:1979:SBDa

[Mül79a] H. N. Mülthei. Splineapproximationen von beliebigem Defekt zur numerischen Lösung gewöhnlicher Differentialgleichungen. I. (German) [Spline approximations of arbitrary defect for the numerical solution of ordinary differential equations. I]. Numerische Mathematik, 32 (2):147–157, June 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Multhei:1979:SBDb

[Mül79b] Heinrich N. Mülthei. Splineapproximationen von beliebigem Defekt zur numerischen Lösung gewöhnlicher Differentialgleichungen. II. (German) [Spline approximations of arbitrary defect for the numerical solution of ordinary differential equations]. Numerische Mathematik, 32(3): 343–358, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Multhei:1980:SBD

[Mül80] Heinrich N. Mülthei. Splineapproximationen von beliebigem Defekt zur numerischen Lösung gewöhnlicher Differentialgleichungen. III. (German) [Spline approximations of arbitrary defect for the numerical solution of ordinary differential equations. III]. Numerische Mathematik, 34(2):143–154, June 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Muller:1982:DME

[Mül82a] Rolf E. Müller. Discretization of multiparameter eigenvalue problems. Numerische Mathematik, 40(3):319–328, December 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Multhei:1982:MKN

[Mül82b] H. N. Mülthei. Maximale Konvergenzordnung bei der numerischen Lösung von Anfangswertproblemen mit Splines. (German) [Maximal order of convergence in the numerical solution of initial value problems with splines]. Numerische Mathematik, 39(3):449–463, October 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Medley:1968:SIGa

[MV68a] Helen I. Medley and Richard S. Varga. On smallest isolated Gerschgorin disks for eigenvalues. II. Numerische Mathematik, 11 (4):320–323, May 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Medley:1968:SIGb

[MV68b] Helen I. Medley and Richard S.

Varga. On smallest isolated Gerschgorin disks for eigenvalues. III. Numerische Mathematik, 11 (5):361–369, June 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Marek:1969:NBS

[MV69] Ivo Marek and Richard S. Varga.

Nested bounds for the spectral radius. Numerische Mathematik, 14(1):49–70, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Meyer:1980:SNI

[MV80] D. Meyer and K. Veselić. On some new inclusion theorems for the eigenvalues of partitioned matrices. Numerische Mathematik, 34(4):431–437, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mattheij:1976:EEM

[MvdS76] R. M. M. Mattheij and A. van der Sluis. Error estimates for Miller’s algorithm. Numerische Mathematik, 26(1):61–78, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Mooney:1979:DCP

[MVW79] J. W. Mooney, H. Voss, and B. Werner. The dependence of critical parameter bounds on the monotonicity of a Newton sequence. Numerische Mathematik, 33(3):291–301, September 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Maehly:1960:TAKa

[MW60a] H. Maehly and Ch. Witzgall. Tschebyscheff-Approximationen in kleinen Intervallen. I. (German) [Chebyshev approximations in small intervals. I]. Numerische Mathematik, 2(1):142– 150, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Maehly:1960:TAKb

[MW60b] H. Maehly and Ch. Witzgall. Tschebyscheff-Approximationen in kleinen Intervallen. II: Stetigkeitssätze für gebrochen rationale Approximationen. (German) [Chebyshev approximations in small intervals. II. stability conditions for broken rational approximations]. Numerische Mathematik, 2(1):293– 307, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Martin:1965:SDPa

[MW65] R. S. Martin and J. H. Wilkinson. Symmetric decomposition of positive definite band matrices. Numerische Mathematik, 7 (5):355–361, October 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 50– 56].

Martin:1967:SSU

[MW67] R. S. Martin and J. H. Wilkinson. Solution of symmetric and unsymmetric band equations and the calculation of eigenvectors of band matrices. Numerische Mathematik, 9(4): 279–301, February 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 70– 92].

Martin:1968:IA

[MW68a] R. S. Martin and J. H. Wilkinson. The implicit QL algorithm. Numerische Mathematik, 12(5): 377–383, December 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 241– 248], where A. A. Dubrulle is listed as the lead author.

Martin:1968:RSE

[MW68b] R. S. Martin and J. H. Wilkinson. Reduction of the symmetric eigenproblem Ax = λBx and related problems to standard form. Numerische Mathematik, 11(2):99–110, February 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 303–314].

Martin:1968:MAC

[MW68c] Roger S. Martin and J. H.

Wilkinson. The modified LR algorithm for complex Hessenberg matrices. Numerische Mathematik, 12(5):369–376, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 396–403].

Martin:1968:SRG

[MW68d] Roger S. Martin and J. H. Wilkinson. Similarity reduction of a general matrix to Hessenberg form. Numerische Mathematik, 12(5):349–368, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 339–358].

Maitre:1999:TEB

[MW99] E. Maitre and P. Witomski. Transport equation with boundary conditions for free surface localization. Numerische Mathematik, 84(2):275–303, December 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9084002/90840275.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9084002/90840275.pdf.

Midy:1979:EMQ

[MY79] P. Midy and Y. Yakovlev. An extrapolation method for the quadrature of functions with singularities in the vicinity of the interval of integration. Numerische Mathematik, 32(2): 183–196, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nabben:1992:CMW

[Nab92] Reinhard Nabben. On a class of matrices which arise in the numerical solution of Euler equations. Numerische Mathematik, 63(3):411–431, December 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nagler:1963:PST

[Nag63] H. Nagler. On a property of simplex tableaux in the transportation problem. Numerische Mathematik, 4:398–402, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nakao:1985:EEG

[Nak85] Mitsuhiro T. Nakao. Error estimates of a Galerkin method for some nonlinear Sobolev equations in one space dimension. Numerische Mathematik, 47(1): 139–157, August 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1969:EGE

[Nat69a] Frank Natterer. Einschließungen für die großen Eigenwerte gewönlicher Differentialgleichungen zweiter und vierter Ordnung. (German) [Bounds for the large eigenvalues of second- and fourth-order ordinary differential equations]. Numerische Mathematik, 13(1): 78–93, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1969:NBS

[Nat69b] Frank Natterer. Numerische Behandlung singulärer Sturm– Liouville-Probleme. (German) [Numerical treatment of singular Sturm–Liouville problems]. Numerische Mathematik, 13(5): 434–447, October 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1970:SEG

[Nat70]  Frank Natterer.  Schranken

für die Eigenwerte gewöhnlicher Differentialgleichungen durch Spline-Approximation. (German) [Bounds for the eigenvalue of ordinary differential equations by spline approximation]. Numerische Mathematik, 14(4): 346–354, March 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1971:SNG

[Nat71] Frank Natterer. Schranken für die Norm der Greenschen Funktion gewöhnlicher Differentialoperatoren. (German) [Bounds for the norm of the Green’s function of ordinary differential operators ]. Numerische Mathematik, 17(5):402–408, October 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1975:DSR

[Nat75a]  Frank Natterer.  Das Dif-

ferenzenverfahren für singuläre Rand-Eigenwertaufgaben gewönlicher Differentialgleichungen. (German) [The difference method for singular boundary/eigenvalues of ordinary differential equations]. Numerische Mathematik, 23(5):387–409, October 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1975:PKF

[Nat75b] Frank Natterer. Uber die punktweise Konvergenz Finiter Elemente. (German) [On the pointwise convergence of finite elements]. Numerische Mathematik, 25(1):67–77, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1977:RSG

[Nat77]  Frank Natterer.  Regular-

isierung schlecht gestellter Probleme durch Projectionsverfahren. (German) [Regularization of ill-posed problems by projection methods]. Numerische Mathematik, 28(3): 329–341, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1978:NIR

[Nat78] Frank Natterer. Numerical inversion of the Radon transform. Numerische Mathematik, 30(1): 81–91, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1979:IAR

[Nat79] Frank Natterer. On the inversion of the attenuated Radon transform. Numerische Mathematik, 32(4):431–438, December 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1985:FRT

[Nat85] Frank Natterer. Fourier reconstruction in tomography. Numerische Mathematik, 47(3): 343–353, November 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Naiman:1997:NCG

[NBE97] Aaron E. Naiman, Ivo M. Babuška, and Howard C. Elman. A note on conjugate gradient convergence. Numerische Mathematik, 76(2): 209–230, April 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/7076002/70760209.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/7076002/70760209.
                                        htm.

Nashed:1993:CNL

[NC93] M. Z. Nashed and X. Chen. Convergence of Newton-like methods for singular operator equations using outer inverses. Numerische Mathematik, 66(2): 235–257, November 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nedelec:1980:MFE

[Néd80] J.-C. Nédélec. Mixed finite elements in r^{3}. Numerische Mathematik, 35(3):315–341, October 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nedelec:1982:EFM

[Néd82] J.-C. Nédélec. Éléments finis mixtes incompressibles pour l’équation de Stokes dans r^{3}. (French) [Incompressible mixed finite elements for the Stokes equation in r^{3}]. Numerische Mathematik, 39(1):97–112, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nedelec:1986:NFM

[Néd86] J.-C. Nédélec. A new family of mixed finite elements in r^{3}. Numerische Mathematik, 50(1): 57–81, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

NegronMarrero:1990:NMD

[Neg90] Pablo V. Negrón Marrero. A numerical method for detecting singular minimizers of multidimensional problems in nonlinear elasticity. Numerische Mathematik, 58(2):135–144, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nemeth:1965:CEF

[Ném65] G. Németh. Chebyshev expansions for Fresnel integrals. Numerische Mathematik, 7(4):310– 312, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neubuser:1960:UUE

[Neu60]  J. Neubüser.  Untersuchun-

gen des Untergruppenverbandes endlicher Gruppen auf einer programmgesteuerten elektronischen Dualmaschine. (German) [Investigation of subroup units of finite groups with a programmable electronic dual machine]. Numerische Mathematik, 2(1):280–292, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neubuser:1961:BUE

[Neu61] J. Neubüser. Bestimmung der Untergruppenverbände endlicher p-Gruppen auf einer programmgesteuerten elektronischen Dualmaschine. (German) [Determination of subgroups of finite p-groups on a programcontrolled electronic Dual machine]. Numerische Mathematik, 3(1):271–278, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
271.

Neubauer:1963:EUM

[Neu63]  Gerhard Neubauer.  Eine

empirische Untersuchung zur Mertensschen Funktion. (German) [An empirical investigation of the Merten function]. Numerische Mathematik, 5(1): 1–13, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nevanlinna:1977:BGE

[Nev77] Olavi Nevanlinna. On the behaviour of global errors at infinity in the numerical integration of stable initial value problems. Numerische Mathematik, 28(4):445–454, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nevanlinna:1990:LAP

[Nev90a] Olavi Nevanlinna. Linear acceleration of Picard–Lindelöf iteration. Numerische Mathematik, 57(2):147–156, April 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nevanlinna:1990:PBP

[Nev90b]  Olavi Nevanlinna.  Power

bounded prolongations and Picard–Lindelöf iteration. Numerische Mathematik, 58(5): 479–501, December 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nichols:1969:GCO

[NF69] Nancy K. Nichols and L. Fox. Generalised consistent ordering and the optimum successive overrelaxation factor. Numerische Mathematik, 13(5): 425–433, October 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nickel:1966:NBW

[Nic66a] Karl Nickel. Die numerische Berechnung der Wurzeln eines Polynoms. (German) [The numerical computation of polynomial roots]. Numerische Mathematik, 9(1):80–98, November 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nickel:1966:UNF

[Nic66b] Karl Nickel. Über die Notwendigkeit einer Fehlerschranken-Arithmetik für Rechenautomaten. (German) [On the need for interval arithmetic on computers]. Numerische Mathematik, 9(1):69– 79, November 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nicolaides:1974:GAS

[Nic74] R. A. Nicolaides. On a geometrical aspect of SOR and the theory of consistent ordering for positive definite matrices. Numerische Mathematik, 23(2):99–104, April 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nicolaides:1975:LCC

[Nic75] R. A. Nicolaides. On the local convergence of certain two step iterative procedures. Numerische Mathematik, 24(2):95– 101, April 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nielson:1973:BSF

[Nie73a] Gregory M. Nielson. Bivariate spline functions and the approximation of linear functionals. Numerische Mathematik, 21 (2):138–160, April 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niethammer:1973:NVA

[Nie73b]  Wilhelm Niethammer.  Ein

numerisches Verfahren zur analytischen Fortsetzung. (German) [A numerical method for analytical continuation]. Numerische Mathematik, 21(1):81– 92, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niethammer:1980:NAE

[Nie80] Wilhelm Niethammer. Numerical application of Euler’s series transformation and its generalizations. Numerische Mathematik, 34(3):271–283, September 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niederreiter:1985:STP

[Nie85] Harald Niederreiter. The serial test for pseudo-random numbers generated by the linear congruential method. Numerische Mathematik, 46(1): 51–68, March 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niethammer:1989:SMP

[Nie89]  Wilhelm Niethammer.  The

SOR method on parallel computers. Numerische Mathematik, 56(2/3):247–254, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nievergelt:1991:ASA

[Nie91]  Yves Nievergelt.  Aitken’s

and Steffensen’s accelerations in several variables. Numerische Mathematik, 59(3): 295–310, June 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nievergelt:1995:BHE

[Nie95]  Yves Nievergelt.  Bisection

hardly ever converges linearly. Numerische Mathematik, 70(1): 111–118, March 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/5070001/50700111.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/5070001/50700111.
                                        htm.

Niijima:1984:UCD

[Nii84]  Koichi Niijima.  A uni-

formly convergent difference scheme for a semilinear singular perturbation problem. Numerische Mathematik, 43(2): 175–198, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niijima:1985:DSE

[Nii85] Koichi Niijima. On a difference scheme of exponential type for a nonlinear singular perturbation problem. Numerische Mathematik, 46(4):521–539, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niijima:1990:PEE

[Nii90] Koichi Niijima. Pointwise error estimates for a streamline diffusion finite element scheme. Numerische Mathematik, 56(7): 707–719, January 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nikolov:1992:MSQ

[Nik92] Geno Nikolov. On the monotonicity of sequences of quadrature formulae. Numerische Mathematik, 62(4):557–565, September 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ninham:1966:GFD

[Nin66] B. W. Ninham. Generalised functions and divergent integrals. Numerische Mathematik, 8(5):444–457, August 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nitsche:1962:ATA

[Nit62] Johannes C. C. Nitsche. Über die Abhängigkeit der Tschebyscheffschen Approximierenden einer differenzierbaren Funktion vom Intervall. (German) [On the relation of Chebyshev approximations for a differentiable interval function]. Numerische Mathematik, 4:262–276, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nitsche:1968:KQO

[Nit68] Joachim A. Nitsche. Ein Kriterium für die Quasi-Optimalität des Titzschen Verfahrens. (German) [A criterion for the quasioptimality of Titzche methods]. Numerische Mathematik, 11(4): 346–348, May 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nitsche:1969:ISF

[Nit69a] J. Nitsche. Interpolation in Sobolevschen Funktionenräumen. (German) [Interpolation in Sobolev function spaces]. Numerische Mathematik, 13(4): 334–343, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nitsche:1969:VRS

[Nit69b]  Joachim A. Nitsche.  Ver-

fahren von Ritz und Spline-Interpolation bei Sturm–Liouville-Randwertproblemen. (German) [Ritz methods and spline interpolation of Sturm–Liouville boundary value problems]. Numerische Mathematik, 13(3): 260–265, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nitsche:1970:KNB

[Nit70] J. Nitsche. Zur Konvergenz von Näherungsverfahren bezüglich verschiedener Normen. (German) [On the convergence of approximation processes with respect to various norms]. Numerische Mathematik, 15(3): 224–228, July 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nitsche:1976:ERB

[Nit76] J. A. Nitsche. Der Einfluß von Randsingularitäten beim Ritzschen Verfahren. (German) [Solution effects of the Ritz method]. Numerische Mathematik, 25(3):263–278, September 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Jin:1999:PPC

[nJyH99] Qi nian Jin and Zong yi Hou. On an a posteriori parameter choice strategy for Tikhonov regularization of nonlinear ill-posed problems. Numerische Mathematik, 83(1):139–159, July 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083001/90830139.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083001/90830139.pdf.

Negron-Marrero:1989:NCC

[NMC89] Pablo V. Negrón-Marrero and Carlos Carbonera. The numerical computation of compressed states of nonlinearly elastic anisotropic plates. Numerische Mathematik, 56(1):93– 107, September 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nelson:1987:GPM

[NN87] Stuart Nelson and Michael Neumann. Generalizations of the projection method with applications to SOR theory for Hermitian positive semidefinite linear systems. Numerische Mathematik, 51(2):123–141, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nataf:1997:CRS

[NN97] F. Nataf and F. Nier. Convergence rate of some domain decomposition methods for overlapping and nonoverlapping subdomains. Numerische Mathematik, 75(3):357–377, January 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/7075003/70750357.

htm;   http://science.springer.
de/nmee/bibs/7075003/70750357.
htm.

Noltemeier:1968:A

[NO68] H. Noltemeier and M. Ost. Zum Auswahlproblem. (German) [On the selection problem]. Numerische Mathematik, 12(5): 431–434, December 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nochetto:1988:SEE

[Noc88] Ricardo H. Nochetto. Sharp L^{∞}-error estimates for semilinear elliptic problems with free boundaries. Numerische Mathematik, 54(3):243–255, December 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nochetto:1989:PAF

[Noc89] Ricardo H. Nochetto. Pointwise accuracy of a finite element method for nonlinear variational inequalities. Numerische Mathematik, 54(6):601–618, March 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Noda:1971:NCM

[Nod71] Takashi Noda. Note on the computation of the maximal eigenvalue of a non-negative irreducible matrix. Numerische Mathematik, 17(5):382–386, October 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Norsett:1975:PRA

[Nør75] Syvert P. Nørsett. C-polynomials for rational approximation to the exponential function. Numerische Mathematik, 25(1): 39–56, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Notaris:1990:ENG

[Not90] Sotirios E. Notaris. The error norm of Gaussian quadrature formulae for weight functions of Bernstein–Szegő type. Numerische Mathematik, 57(3): 271–283, May 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Notaris:1993:EBG

[Not93a]  Sotirios E. Notaris.  Er-

ror bounds for Gauss–Kronrod quadrature formulae of analytic functions. Numerische Mathematik, 64(3):371–380, March 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Notaris:1993:ENG

[Not93b] Sotirios E. Notaris. The error norm of Gauss–Lobatto quadrature formulae for weight functions of Bernstein–Szegő type. Numerische Mathematik, 64(3): 381–393, March 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Notay:1993:CRC

[Not93c] Yvan Notay. On the convergence rate of the conjugate gradients in presence of rounding errors. Numerische Mathematik, 65(3):301–317, July 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Notay:1998:UAI

[Not98]  Y. Notay.  Using approxi-

mate inverses in algebraic multilevel methods. Numerische Mathematik, 80(3):397–417, September 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080003/80800397.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800397.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800397.
ps.gz.

Novak:1986:APE

[Nov86] Erich Novak. The average a posteriori error of numerical methods. Numerische Mathematik, 50(2):245–252, March 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Novak:1988:SPQ

[Nov88] Erich Novak. Stochastic properties of quadrature formulas. Numerische Mathematik, 53(5): 609–620, August 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neumann:1978:CNM

[NP78] M. Neumann and R. J. Plemmons. Convergent nonnegative matrices and iterative methods for consistent linear systems. Numerische Mathematik, 31(3): 265–279, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neckermann:1968:UAD

[NR68] L. Neckermann and P. O. Runck. Über Approximationseigenschaften differenzierter Lagrangescher Interpolationspolynome mit Jacobischen Abszissen. (German) [On approximate eigenvalues of differentiated Lagrangian interpolation polynomials with Jacobian abscissas]. Numerische Mathematik, 12(3):159–169, October 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Novo:1992:SRO

[NR92] Sylvia Novo and Jesús Rojo. Some remarks on an ODEsolver of Kirchgraber. Numerische Mathematik, 61(2): 261–264, March 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nurnberger:1995:BSI

[NR95] G. Nürnberger and Th. Riessinger. Bivariate spline interpolation at grid points. Numerische Mathematik, 71(1):91–119, July 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5071001/
50710091.htm;   http://science.
springer.de/nmee/bibs/5071001/
50710091.htm.

Novak:1996:HDI

[NR96] Erich Novak and Klaus Ritter.

High-dimensional integration of smooth functions over cubes. Numerische Mathematik, 75(1): 79–97, November 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6075001/60750079.
htm;   http://science.springer.
de/nmee/bibs/6075001/60750079.
htm.

Nirschl:1964:BFV

[NS64] N. Nirschl and H. Schneider.

The Bauer fields of values of a matrix. Numerische Mathematik, 6(1):355–365, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neittaanmaki:1981:FEA

[NS81] P. Neittaanmäki and J. Saranen. On finite element approximation of the gradient for solution of Poisson equation. Numerische Mathematik, 37(3): 333–337, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nurnberger:1983:ABS

[NS83a] Günther Nürnberger and Manfred Sommer. Alternation for best spline approximations. Numerische Mathematik, 41(2): 207–221, June 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nurnberger:1983:RTA

[NS83b] Günther Nürnberger and Manfred Sommer. A Remez type algorithm for spline functions. Numerische Mathematik, 41(1): 117–146, April 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niezgodka:1991:CNA

[NS91] Marek Niezgódka and Jürgen Sprekels. Convergent numerical approximations of the thermomechanical phase transitions in shape memory alloys. Numerische Mathematik, 58(7): 759–778, February 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Najib:1995:DAS

[NS95a] K. Najib and D. Sandri. On a decoupled algorithm for solving a finite element problem for the approximation of viscoelastic fluid flow. Numerische Mathematik, 72(2):223–238, December 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/

00211/bibs/5072002/50720223.
htm;   http://science.springer.
de/nmee/bibs/5072002/50720223.
htm.

Nipp:1995:IMG

[NS95b] K. Nipp and D. Stoffer. Invariant manifolds and global error estimates of numerical integration schemes applied to stiff systems of singular perturbation type. I. RK-methods. Numerische Mathematik, 70(2): 245–257, April 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070002/50700245.
htm;   http://science.springer.
de/nmee/bibs/5070002/50700245.
htm.

Nipp:1996:IMG

[NS96] K. Nipp and D. Stoffer. Invariant manifolds and global error estimates of numerical integration schemes applied to stiff systems of singular perturbation type. II. linear multistep methods. Numerische Mathematik, 74(3):305–323, September 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074003/60740305.
htm;   http://science.springer.
de/nmee/bibs/6074003/60740305.
htm.

Nurnberger:1986:ASA

[NSS86] G. Nürnberger, M. Sommer, and H. Strauß. An algorithm for segment approximation. Numerische Mathematik, 48(4):463–477, April 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nurnberger:1983:IGS

[NSSS83] G. Nürnberger, L. L. Schumaker, M. Sommer, and H. Strauss. Interpolation by generalized splines. Numerische Mathematik, 42(2): 195–212, November 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nelson:1980:CTD

[NV80a] P. Nelson, Jr. and H. D. Victory, Jr. Convergence of twodimensional Nyström discreteordinates in solving the linear transport equation. Numerische Mathematik, 34(4): 353–370, December 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Neumann:1980:SSU

[NV80b] Michael Neumann and Richard S. Varga. On the sharpness of some upper bounds for the spectral radii of S.O.R. iteration matrices. Numerische Mathematik, 35(1):69–79, March 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niethammer:1983:ASI

[NV83] W. Niethammer and R. S.

Varga. The analysis of k-step iterative methods for linear systems from summability theory. Numerische Mathematik, 41(2): 177–206, June 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nochetto:1996:CEE

[NV96] Ricardo H. Nochetto and Claudio Verdi. Combined effect of explicit time-stepping and quadrature for curvature driven flows. Numerische Mathematik, 74(1): 105–136, July 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074001/60740105.
htm;   http://science.springer.
de/nmee/bibs/6074001/60740105.
htm.

Neunzert:1973:TAV

[NW73] H. Neunzert and J. Wick.

Die Theorie der asymptotischen Verteilung und die numerische Lösung von Integrodifferentialgleichungen. (German) [The theory of asymptotic distribution and the numerical solution of integrodifferential equations]. Numerische Mathematik, 21(3):234–243, June 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Natterer:1974:EML

[NW74] Frank Natterer and Bodo Werner. Eine Erweiterung des Maximumsprinzips für den Laplaceschen Operator. (German) [An extension of the maximum principle for the Laplace operator]. Numerische Mathematik, 22(2):149–155, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Noor:1976:EBF

[NW76] M. A. Noor and J. R. Whiteman. Error bounds for finite element solutions of mildly nonlinear elliptic boundary value problems. Numerische Mathematik, 26(1):107–116, March 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Norsett:1979:OCR

[NW79] Syvert P. Nørsett and Arne Wolfbrandt. Order conditions for Rosenbrock type methods. Numerische Mathematik, 32(1): 1–15, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Norsett:1981:PCR

[NW81] S. P. Nørsett and G. Wanner. Perturbed collocation and Runge–Kutta methods. Numerische Mathematik, 38(2): 193–208, December 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niethammer:1984:ALP

[NW84] W. Niethammer and H. Wietschorke. On the acceleration of limit periodic continued fractions. Numerische Mathematik, 44(1):129–137, June 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Niederdrenk:1983:GSS

[NY83] Klaus Niederdrenk and Harry Yserentant. Die gleichmäßige Stabilität singulär gestörter diskreter und kontinuierlicher Randwertprobleme. (German) [The uniform stability of singularly perturbed discrete and continuous boundary value problems]. Numerische Mathematik, 41(2):223–253, June 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Omodei:1975:SRP

[OA75] B. J. Omodei and R. S. Anderssen. Stability of the Rayleigh– Ritz procedure for nonlinear two-point boundary value problems. Numerische Mathematik, 24(1):27–38, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

OCinneide:1993:EPT

[O’C93] Colm Art O’Cinneide. Entrywise perturbation theory and error analysis for Markov chains. Numerische Mathematik, 65(1): 109–120, May 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

OCinneide:1996:REB

[O’C96] Colm Art O’Cinneide. Relativeerror bounds for the LU decomposition via the GT H algorithm. Numerische Mathematik, 73(4):507–519, June 1996. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/6073004/
60730507.htm;   http://science.
springer.de/nmee/bibs/6073004/
60730507.htm.

Olsen:1995:BSC

[OD95] Elwood T. Olsen and Jim Douglas, Jr. Bounds on spectral condition numbers of matrices arising in the p-version of the finite element method. Numerische Mathematik, 69(3):333–352, January 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069003/50690333.
htm;   http://science.springer.
de/nmee/bibs/5069003/50690333.
htm.

Ogata:1989:MCM

[Oga89] Yosihiko Ogata. A Monte Carlo method for high-dimensional integration. Numerische Mathematik, 55(2):137–157, May 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oliver:1967:REP

[Oli67] J. Oliver. Relative error propagation in the recursive solution of linear recurrence relations. Numerische Mathematik, 9(4): 323–340, February 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oliver:1968:EOE

[Oli68a] J. Oliver. An extension of Olver’s error estimation technique for linear recurrence relations. Numerische Mathematik, 12(5):459–467, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oliver:1968:NSL

[Oli68b] J. Oliver. The numerical solution of linear recurrence relations. Numerische Mathematik, 11(4):349–360, May 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oliver:1971:EEM

[Oli71]  J. Oliver.  The efficiency

of extrapolation methods for numerical integration. Numerische Mathematik, 17(1):17– 32, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oliveira:1980:CRC

[Oli80] F. Aleixo Oliveira. Collocation and residual correction. Numerische Mathematik, 36(1):27– 31, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Opfer:1988:DCS

[OO88] Gerhard Opfer and Hans Joachim Oberle. The derivation of cubic splines with obstacles by methods of optimization and optimal control. Numerische Mathematik, 52(1):17–31, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oberle:1998:NSP

[OO98] H. J. Oberle and G. Opfer. Nonnegative splines, in particular of degree five. Numerische Mathematik, 79(3): 427–450, May 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079003/80790427.
htm;   http://science.springer.
de/nmee/bibs/8079003/80790427.
htm.

Oettli:1964:CAS

[OP64] W. Oettli and W. Prager. Compatibility of approximate solution of linear equations with given error bounds for coefficients and right-hand sides. Numerische Mathematik, 6(1):405– 409, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oberle:1981:NTD

[OP81] H. J. Oberle and H. J. Pesch. Numerical treatment of delay differential equations by Hermite interpolation. Numerische Mathematik, 37(2): 235–255, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oliker:1988:NSE

[OP88] V. I. Oliker and L. D. Prussner.

On the numerical solution of the equation (∂^{2}z/∂x^{2})(∂^{2}z/∂y^{2}) − ((∂^{2}z/∂x∂y))^{2} = f and its discretizations. I. Numerische Mathematik, 54(3): 271–293, December 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oukit:1996:MFE

[OP96] A. Oukit and R. Pierre. Mixed finite element for the linear plate problem: The Hermann-Miyoshi model revisited. Numerische Mathematik, 74(4):453–477, October 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074004/60740453.
htm;   http://science.springer.
de/nmee/bibs/6074004/60740453.
htm.

Opfer:1979:NEP

[Opf79]  Gerhard Opfer.  New ex-

tremal properties for constructing conformal mappings. Numerische Mathematik, 32(4): 423–429, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Opfer:1980:CMP

[Opf80] Gerhard Opfer. Conformal mappings onto prescribed regions via optimization techniques. Numerische Mathematik, 35(2): 189–200, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Opfer:1982:SCA

[Opf82] Gerhard Opfer. Solving complex approximation problems by semiinfinite-finite optimization techniques: a study of convergence. Numerische Mathematik, 39(3):411–420, October 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Omladic:1976:NTQ

[OPS76] M. Omladič, S. Pahor, and A. Suhadolc. On a new type of quadrature formulas. Numerische Mathematik, 25(4): 421–426, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

ORiordan:1984:SPF

[O’R84] Eugene O’Riordan. Singularly perturbed finite element methods. Numerische Mathematik, 44(3):425–434, September 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

OReilly:1987:USS

[O’R87] M. J. O’Reilly. A uniform scheme for the singularly perturbed Riccati equation. Numerische Mathematik, 50(4): 483–501, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ortega:1963:EAH

[Ort63] James M. Ortega. An error analysis of Householder’s method for the symmetric eigenvalue problem. Numerische Mathematik, 5(1):211–225, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ortloff:1986:AVK

[Ort86] H. Ortloff. Asymptotisches Verhalten und Konvergenzbeschleunigung von Iterationsfolgen. (German) [Asymptotic behavior and acceleration of convergence of iteration sequences]. Numerische Mathematik, 49(5): 545–559, September 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Osada:1996:ATA

[Osa96] Naoki Osada. An acceleration theorem for the ρ-algorithm. Numerische Mathematik, 73(4): 521–531, June 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073004/60730521.
htm;   http://science.springer.
de/nmee/bibs/6073004/60730521.
htm.

Osborne:1965:NNS

[Osb65] M. R. Osborne. A note on the numerical solution of a periodic parabolic problem. Numerische Mathematik, 7(2):155– 158, April 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Osborne:1967:NSP

[Osb67]  M. R. Osborne.  The nu-

merical solution of a periodic parabolic problem subject to a nonlinear boundary condition. Numerische Mathematik, 10(1): 76–79, May 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Osborne:1968:NSP

[Osb68]  M. R. Osborne.  The nu-

merical solution of a periodic parabolic problem subject to a nonlinear boundary condition. II. Numerische Mathematik, 12(4):280–287, November 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Osborn:1969:NPT

[Osb69]  John E. Osborn.  A note

on a perturbation theorem for the matrix eigenvalue problem. Numerische Mathematik, 13(2): 152–153, May 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ortega:1990:NSC

[OSS90] T. Ortega and J. M. Sanz-Serna. Nonlinear stability and convergence of finite-difference methods for the “good” Boussinesq equation. Numerische Mathematik, 58(2):215–229, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Oswald:1992:HBM

[Osw92] P. Oswald. On a hierarchical basis multilevel method with nonconforming p1 elements. Numerische Mathematik, 62(2): 189–212, July 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Osborne:1969:EBA

[OW69] M. R. Osborne and G. A. Watson. On estimating best approximations of functions defined by differential equations. Numerische Mathematik, 13(3): 199–206, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Owens:1977:ABA

[Owe77]  R. W. Owens.  An algo-

rithm for best approximate solutions of Ax=b with a smooth strictly convex norm. Numerische Mathematik, 29(1):83– 91, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paardekooper:1971:EAS

[Paa71] M. H. C. Paardekooper. An eigenvalue algorithm for skewsymmetric matrices. Numerische Mathematik, 17(3): 189–202, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paget:1981:NEH

[Pag81] D. F. Paget. The numerical evaluation of Hadamard finite-part integrals. Numerische Mathematik, 36(4):447–453, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pallaschke:1976:ODI

[Pal76]  Diethard Pallaschke.  Opti-

male Differentiations- und Integrationsformeln in C_{0}[a, b]. (German) [Optimal differentiation and integration formulae in C_{0}(a, b)]. Numerische Mathematik, 26(2):201–210, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pallaschke:1977:KOD

[Pal77] Diethard Pallaschke. Konvergenz optimaler Differentiationsformeln. (German) [Convergence of optimal differentiation formulas]. Numerische Mathematik, 27(4):421–426, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Palencia:1999:BEM

[Pal99] C. Palencia. Backward Euler method for abstract timedependent parabolic equations with variable domains. Numerische Mathematik, 82(3): 471–490, May 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer-
ny.com/link/service/journals/
00211/bibs/9082003/90820471.
htm;   http://link.springer-
ny.com/link/service/journals/
00211/papers/9082003/90820471.
pdf.

Pao:1985:MIM

[Pao85] C. V. Pao. Monotone iterative methods for finite difference system of reaction-diffusion equations. Numerische Mathematik, 46(4):571–586, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pao:1987:NSS

[Pao87] C. V. Pao. Numerical solutions for some coupled systems of nonlinear boundary value problems. Numerische Mathematik, 51(4): 381–394, August 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pao:1995:BMI

[Pao95] C. V. Pao. Block monotone iterative methods for numerical solutions of nonlinear elliptic equations. Numerische Mathematik, 72(2):239–262, December 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5072002/50720239.
htm;   http://science.springer.

de/nmee/bibs/5072002/50720239.
htm.

Pao:1998:AMI

[Pao98]  C. V. Pao.  Accelerated

monotone iterative methods for finite difference equations of reaction-diffusion. Numerische Mathematik, 79(2): 261–281, April 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079002/80790261.
htm;   http://science.springer.
de/nmee/bibs/8079002/80790261.
htm.

Parter:1959:SLI

[Par59]  Seymour V. Parter.  On

“two-line” iterative methods for the Laplace and biharmonic difference equations. Numerische Mathematik, 1(1):240– 252, December 1959. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
240.

Parter:1961:SLI

[Par61] Seymour V. Parter. “multiline” iterative methods for elliptic difference equations and fundamental frequencies. Numerische Mathematik, 3(1):305– 319, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
305.

Parter:1962:ONS

[Par62a] Seymour V. Parter. An observation on the numerical solution of difference equations and a theorem of Szegő. Numerische Mathematik, 4:293–295, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Parter:1962:SCP

[Par62b] Seymour V. Parter. Stability, convergence, and pseudostability of finite-difference equations for an over-determined problem. Numerische Mathematik, 4:277–292, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Parlett:1965:CA

[Par65a] Beresford Parlett. Convergence of the QR algorithm. Numerische Mathematik, 7(2):187– 193, April 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Parter:1965:MNE

[Par65b] Seymour V. Parter. Mildly nonlinear elliptic partial differential equations and their numerical solution. I. Numerische Mathematik, 7(2):113–128, April 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Parlett:1967:CCQ

[Par67] Beresford Parlett. Correction to convergence of the QR algorithm. Numerische Mathematik, 10(2):163–164, July 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Parlett:1973:PGM

[Par73] B. N. Parlett. Présentation géométrique des méthodes de calcul des valeurs propres. (French) [Geometric presentation of methods for calculation of eigenvalues]. Numerische Mathematik, 21(3): 223–233, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Parter:1988:EMM

[Par88] Seymour V. Parter. Estimates for multigrid methods based on red-black Gauss–Seidel smoothings. Numerische Mathematik, 52(6):701–723, June 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Patterson:1976:HPM

[Pat76] T. N. L. Patterson. On high precision methods for the evaluation of Fourier integrals with finite and infinite limits. Numerische Mathematik, 27(1):41– 52, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Patterson:1993:MOQ

[Pat93] T. N. L. Patterson. Modified optimal quadrature extensions. Numerische Mathematik, 64(4): 511–520, April 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paulik:1977:EOQ

[Pau77] Augustin Paulik. Zur existenz optimaler Quadraturformeln mit freien Knoten bei Integration analytischer Funktionen. (German) [On the existence of optimal integration formulas with free nodes for analytic functions]. Numerische Mathematik, 27(4):395–405, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paumier:1981:MNPa

[Pau81a]  Jean-Claude Paumier.  Une

méthode numérique pour le calcul des points de retournement. application à un problème aux limites non-linéaire. I. étude théorique et expérimentation de la méthode. (French) [A numerical method for the calculation of turning points. application to a problem at nonlinear limits. I. theoretical and experimental study of the method]. Numerische Mathematik, 37(3): 433–444, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paumier:1981:MNPb

[Pau81b]  Jean-Claude Paumier.  Une

méthode numérique pour le calcul des points de retournement. application à un problème aux limites non-linéaire. II. analyse numérique d’un problème aux limites non-linéaire. (French) [A numerical method for the calculation of turning points. application to a problem at non-linear limits. II. numerical analysis of a problem at nonlinear limits.]. Numerische Mathematik, 37(3): 445–452, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pavarino:1994:ASM

[Pav94a] Luca F. Pavarino. Additive Schwarz methods for the pversion finite element method. Numerische Mathematik, 66(4): 493–515, January 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pavarino:1994:SML

[Pav94b] Luca F. Pavarino. Schwarz methods with local refinement for the p-version finite element method. Numerische Mathematik, 69(2):185–211, December 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069002/40690185.
htm;   http://science.springer.
de/nmee/bibs/4069002/40690185.
htm.

Pissanetzky:1981:NEO

[PB81] S. Pissanetzky and F. G. Basombrío. On the numerical errors in one-step algorithms for time dependent problems. Numerische Mathematik, 38(1):31– 37, October 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peisker:1987:CGM

[PB87] P. Peisker and D. Braess. A conjugate gradient method and a multigrid algorithm for Morley’s finite element approximation of the biharmonic equation. Numerische Mathematik, 50(5):567–586, March 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Puschmann:1983:CPR

[PC83] Heinrich Puschmann and Joaquín Cortés. The coordinex problem and its relation to the conjecture of Wilkinson. Numerische Mathematik, 42(3):291–297, December 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Petkovic:1995:WFZ

[PCT95] Miodrag S. Petković, Carsten Carstensen, and Miroslav Trajković. Weierstrass formula and zero-finding methods. Numerische Mathematik, 69(3): 353–372, January 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5069003/
50690353.htm;   http://science.

springer.de/nmee/bibs/5069003/
50690353.htm.

Pehlivanov:1996:LSM

[PCV96] A. I. Pehlivanov, G. F. Carey, and P. S. Vassilevski. Leastsquares mixed finite element methods for non-selfadjoint elliptic problems: I. error estimates. Numerische Mathematik, 72(4):501–522, February 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/6072004/
60720501.htm;   http://science.
springer.de/nmee/bibs/6072004/
60720501.htm.

Paget:1972:ANE

[PE72] D. F. Paget and David Elliott. An algorithm for the numerical evaluation of certain Cauchy principal value integrals. Numerische Mathematik, 19(5):373–385, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Park:1997:SAF

[PE97] Haesun Park and Lars Eldén. Stability analysis and fast algorithms for triangularization of Toeplitz matrices. Numerische Mathematik, 76(3): 383–402, May 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076003/70760383.
htm;   http://science.springer.

de/nmee/bibs/7076003/70760383.
htm.

Pearcy:1962:CAD

[Pea62] C. Pearcy. On convergence of alternating direction procedures. Numerische Mathematik, 4:172– 176, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pecka:1974:EWQ

[Pec74] F. Pecka. On equally weighted quadratures. Numerische Mathematik, 23(1):19–22, February 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pedregal:1996:NAN

[Ped96] Pablo Pedregal. On the numerical analysis of non-convex variational problems. Numerische Mathematik, 74(3):325–336, September 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074003/60740325.
htm;   http://science.springer.
de/nmee/bibs/6074003/60740325.
htm.

Peherstorfer:1988:GQF

[Peh88] Franz Peherstorfer. On Gauss quadrature formulas with equal weights. Numerische Mathematik, 52(3):317–327, March 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peherstorfer:1990:GTQ

[Peh90]  Franz Peherstorfer.  Gauss-

Tchebycheff quadrature formulas. Numerische Mathematik, 58 (3):273–286, October 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peherstorfer:1995:ORP

[Peh95] Franz Peherstorfer. Orthogonality of residual polynomials used in minimax polynomial preconditioning. Numerische Mathematik, 71(3):357–363, September 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071003/50710357.
htm;   http://science.springer.
de/nmee/bibs/5071003/50710357.
htm.

Peisker:1985:MAB

[Pei85] Petra Peisker. A multilevel algorithm for the biharmonic problem. Numerische Mathematik, 46(4):623–634, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peisker:1991:MMR

[Pei91] Petra Peisker. A multigrid method for Reissner–Mindlin plates. Numerische Mathematik, 59(5):511–528, August 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pena:1998:PSL

[Peñ98] J. M. Peña. Pivoting strategies leading to diagonal dominance by rows. Numerische Mathematik, 81(2):293–304, December 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8081002/80810293.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081002/80810293.
pdf.

Pereyra:1966:IAS

[Per66] Victor Pereyra. On improving an approximate solution of a functional equation by deferred corrections. Numerische Mathematik, 8(4):376–391, June 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pereyra:1967:IDC

[Per67] Victor Pereyra. Iterated deferred corrections for nonlinear operator equations. Numerische Mathematik, 10(4): 316–323, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pereyra:1968:IDC

[Per68] Victor Pereyra. Iterated deferred corrections for nonlinear boundary value problems. Numerische Mathematik, 11(2): 111–125, February 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Perugia:1999:MFM

[Per99] I. Perugia. A mixed formulation for 3D magnetostatic problems: theoretical analysis and face-edge finite element approximation. Numerische Mathematik, 84(2):305–326, December 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9084002/90840305.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9084002/90840305.pdf.

Petry:1970:EEF

[Pet70]  W. Petry.  Existenz und

Eindeutigkeit für Lösungen nichtlinearer Randwertprobleme. (German) [Existence and uniqueness for solutions of nonlinear boundary value problems]. Numerische Mathematik, 14(5):476–494, April 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=14&issue=5&spage=
476.

Petras:1993:ARC

[Pet93] Knut Petras. Asymptotics for the remainder of a class of positive quadratures for integrands with an interior singularity. Numerische Mathematik, 65(1):121–133, May 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pflaum:1998:MAS

[Pfl98] Christoph Pflaum. A multilevel algorithm for the solution of second order elliptic differential equations on sparse grids. Numerische Mathematik, 79(1): 141–155, March 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079001/80790141.
htm;   http://science.springer.
de/nmee/bibs/8079001/80790141.
htm.

Pierce:1990:ORC

[PHP90] Daniel J. Pierce, Apostolos Hadjidimos, and Robert J. Plemmons. Optimality relationships for p-cyclic SOR. Numerische Mathematik, 56(7): 635–643, January 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pichat:1972:CSA

[Pic72] M. Pichat. Correction d’une somme en arithmétique à virgule flottante. (French) [Correction of a sum in floatingpoint arithmetic]. Numerische Mathematik, 19(5):400–406, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pietrzykowski:1970:PMC

[Pie70] Tomasz Pietrzykowski. The potential method for conditional maxima in the locally compact metric spaces. Numerische Mathematik, 14(4): 325–329, March 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pietrzykowski:1972:GPM

[Pie72] Tomasz Pietrzykowski. A generalization of the potential method for conditional maxima on the Banach, reflexive spaces. Numerische Mathematik, 18(4): 367–372, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pinkus:1975:AMN

[Pin75] Allan Pinkus. Asymptotic minimum norm quadrature formulae. Numerische Mathematik, 24 (2):163–175, April 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pironneau:1982:TDA

[Pir82] O. Pironneau. On the transportdiffusion algorithm and its applications to the Navier–Stokes equations. Numerische Mathematik, 38(3):309–332, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pittnauer:1973:NPK

[Pit73] E. Pittnauer. Numerische Polynomapproximation mit Knotenpolynomen. (German) [Numerical polynomial approximation by knot polynomials]. Numerische Mathematik, 21(3): 256–263, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pittnauer:1975:AAG

[Pit75] E. Pittnauer. Eine Aussage über Alternanten. (German) [A proposition on alternants]. Numerische Mathematik, 23(5): 427–432, October 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pitkaranta:1979:BSF

[Pit79] Juhani Pitkäranta. Boundary subspaces for the finite element method with Lagrange multipliers. Numerische Mathematik, 33(3):273–289, September 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pitkaranta:1988:ASL

[Pit88] Juhani Pitkäranta. Analysis of some low-order finite element schemes for Mindlin–Reissner and Kirchhoff plates. Numerische Mathematik, 53(1/2): 237–254, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pitkaranta:1992:PML

[Pit92] Juhani Pitkäranta. The problem of membrane locking in finite element analysis of cylindrical shells. Numerische Mathematik, 61(4):523–542, May 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Patry:1964:BF

[PK64] J. Patry and J. Keller. Zur Berechnung des Fehlerintegrals. (German) [Toward the computation of error integrals]. Numerische Mathematik, 6(1):89– 97, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pattabhiraman:1969:SPP

[PL69] M. V. Pattabhiraman and P. Lancaster. Spectral properties of a polynomial operator. Numerische Mathematik, 13(3): 247–259, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Plato:1996:DPI

[Pla96] Robert Plato. On the discrepancy principle for iterative and parametric methods to solve linear ill-posed equations. Numerische Mathematik, 75(1):99– 120, November 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/6075001/60750099.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/6075001/60750099.
                                        htm.

Pehlivanov:1992:SAA

[PLCC92] A. I. Pehlivanov, R. D. Lazarov, G. F. Carey, and S. S. Chow. Superconvergence analysis of approximate boundary-flux calculations. Numerische Mathematik, 63(4):483–501, December 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Plemmons:1976:RSD

[Ple76] R. J. Plemmons. Regular splittings and the discrete Neumann problem. Numerische Mathematik, 25(2):153–161, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Polya:1968:GME

[Pól68] George Pólya. Graeffe’s method for eigenvalues. Numerische Mathematik, 11(4):315–319, May 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pomentale:1968:DRL

[Pom68] Tomaso Pomentale. On discrete rational least squares approximation. Numerische Mathematik, 12(1):40–46, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pomentale:1971:CIM

[Pom71] Tomaso Pomentale. A class of iterative methods for holomorphic functions. Numerische Mathematik, 18(3):193–203, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ponstein:1966:SCE

[Pon66] J. Ponstein. Splitting certain eigenvalue/eigenvector problems. Numerische Mathematik, 8(5):412–421, August 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Porsching:1966:DST

[Por66] T. A. Porsching. Diagonal similarity transformations which isolate Gerschgorin disks. Numerische Mathematik, 8(5):437– 443, August 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Porsching:1978:EEM

[Por78] T. A. Porsching. Error estimates for MAC-like approximations to the linear Navier–Stokes equations. Numerische Mathematik, 29(3):291–306, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Potempa:1978:MRQ

[Pot78] H. Potempa. Modifizierte Restgliedabschätzungen für Quadraturformeln. (German) [Modified estimates for quadrature errors]. Numerische Mathematik, 30(1): 39–50, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Potra:1982:EAS

[Pot82] Florian A. Potra. An error analysis for the secant method. Numerische Mathematik, 38(3):

427–445, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pothen:1989:SNB

[Pot89] Alex Pothen. Sparse null basis computations in structural optimization. Numerische Mathematik, 55(5):501–519, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Powell:1994:UCT

[Pow94] M. J. D. Powell. The uniform convergence of thin plate spline interpolation in two dimensions. Numerische Mathematik, 68(1): 107–128, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680107.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680107.
htm.

Potra:1980:SEB

[PP80] Florian-A. Potra and Vlastimil Pták. Sharp error bounds for Newton’s process. Numerische Mathematik, 34(1):63– 72, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Potra:1981:GRF

[PP81] Florian A. Potra and Vlastimil Pták. A generalization of regula falsi. Numerische Mathematik, 36(3):333–346, April 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Papamichael:1997:ACM

[PPSS97] N. Papamichael, I. E. Pritsker, E. B. Saff, and N. S. Stylianopoulos. Approximation of conformal mappings of annular regions. Numerische Mathematik, 76(4): 489–513, June 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076004/70760489.
htm;   http://science.springer.
de/nmee/bibs/7076004/70760489.
htm.

Perrin:1969:HON

[PPV69] F. M. Perrin, H. S. Price, and R. S. Varga. On higher-order numerical methods for nonlinear two-point boundary value problems. Numerische Mathematik, 13(2):180–198, May 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Potra:1998:SMS

[PQS98] Florian A. Potra, Liqun Qi, and Defeng Sun. Secant methods for semismooth equations. Numerische Mathematik, 80(2): 305–324, August 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080002/80800305.
htm;   http://link.springer.
de/link/service/journals/

00211/papers/8080002/80800305.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080002/80800305.
ps.gz.

Parlett:1969:BMC

[PR69] B. N. Parlett and C. Reinsch. Balancing a matrix for calculation of eigenvalues and eigenvectors. Numerische Mathematik, 13(4):293–304, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 315–326].

Prossdorf:1986:SGM

[PR86] S. Prössdorf and A. Rathsfeld. On spline Galerkin methods for singular integral equations with piecewise continuous coefficients. Numerische Mathematik, 48(1):99–118, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pierre:1993:NST

[PR93] Michel Pierre and Jean-R. R.

Roche. Numerical simulation of tridimensional electromagnetic shaping of liquid metals. Numerische Mathematik, 65(2): 203–217, June 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pousin:1994:CSP

[PR94] J. Pousin and J. Rappaz. Consistency, stability, a priori and a posteriori errors for Petrov– Galerkin methods applied to nonlinear problems. Numerische Mathematik, 69(2):213–231, December 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/4069002/40690213.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4069002/40690213.
                                        htm.

Prenter:1971:PS

[Pre71] P. M. Prenter. Piecewise Lsplines. Numerische Mathematik, 18(3):243–253, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Prossdorf:1991:SAP

[Prö91] S. Prössdorf. On the superapproximation property of Galerkin’s method with finite elements. Numerische Mathematik, 59(7):711–722, October 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pruess:1975:HOA

[Pru75] S. Pruess. High order approximations to Sturm–Liouville eigenvalues. Numerische Mathematik, 24(3):241–247, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Perlis:1959:RAL

[PS59] A. J. Perlis and K. Samelson. Report on the algorithmic language ALGOL by the ACM committee on programming languages and the GAMM committee on programming. Numerische Mathematik, 1:41–60, December 1959. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Penkov:1966:HMA

[PS66] B. Penkov and Bl. Sendov.

Hausdorffsche Metrik und Approximationen. (German) [Hausdorff metric and approximations]. Numerische Mathematik, 9(3):214–226, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pereyra:1975:MSD

[PS75] V. Pereyra and E. G. Sewell.

Mesh selection for discrete solution of boundary problems in ordinary differential equations. Numerische Mathematik, 23(3): 261–268, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ponisch:1982:LUK

[PS82] Gerd Pönisch and Hubert Schwetlick. Ein lokal überlinear konvergentes Verfahren zur Bestimmung von Rückkehrpunkten implizit definierter Raumkurven. (German) [A locally superlinearly convergent method for computing turning points of implicitly defined space curves]. Numerische Mathematik, 38(3): 455–465, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Petkovic:1984:COA

[PS84] Miodrag Petković and Lidija Stefanović. On the convergence order of accelerated root iterations. Numerische Mathematik, 44(3):463–476, September 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Papamichael:1992:DDM

[PS92a] N. Papamichael and N. S. Stylianopoulos. A domain decomposition method for approximating the conformal modules of long quadrilaterals. Numerische Mathematik, 62(2): 213–234, July 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Prossdorf:1992:QMS

[PS92b] S. Prössdorf and I. H. Sloan. Quadrature method for singular integral equations on closed curves. Numerische Mathematik, 61(4):543–559, May 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Perthame:1996:PPF

[PS96a] Benoit Perthame and Chi-Wang Shu. On positivity preserving finite volume schemes for Euler equations. Numerische Mathematik, 73(1): 119–130, March 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073001/60730119.
htm;   http://science.springer.
de/nmee/bibs/6073001/60730119.
htm.

Pitkaranta:1996:DPE

[PS96b] Juhani Pitkäranta and Manil Suri. Design principles and error analysis for reduced-shear plate-bending finite elements. Numerische Mathematik, 75(2): 223–266, December 1996. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/6075002/
60750223.htm;   http://science.
springer.de/nmee/bibs/6075002/
60750223.htm.

Palencia:1984:ELR

[PSS84] C. Palencia and J. M. Sanz-Serna. An extension of the Lax– Richtmyer theory. Numerische Mathematik, 44(2):279–283, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pitonyak:1990:IMV

[PSS90] A. Pitonyak, P. Shi, and M. Shillor. On an iterative method for variational inequalities. Numerische Mathematik, 58(3):231–242, October 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ptak:1976:RCN

[Ptá76] Vlastimil Pták. The rate of convergence of Newton’s process. Numerische Mathematik, 25(3): 279–285, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pierce:1972:HOC

[PV72] J. G. Pierce and R. S. Varga. Higher order convergence results for the Rayleigh–Ritz method applied to eigenvalue problems. II. improved error bounds for eigenfunctions. Numerische Mathematik, 19(2): 155–169, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Plato:1990:RPM

[PV90] R. Plato and G. Vainikko. On the regularization of projection methods for solving ill-posed problems. Numerische Mathematik, 57(1):63–79, March 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Peters:1970:ERC

[PW70] G. Peters and J. H. Wilkinson. Eigenvectors of real and complex matrices by LR and QR triangularization. Numerische Mathematik, 16(3): 181–204, December 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 372– 395].

Parlett:1984:EMS

[PW84] B. N. Parlett and W. D.

Wu. Eigenvector matrices of symmetric tridiagonals. Numerische Mathematik, 44(1): 103–110, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Papamichael:1986:SCP

[PW86] N. Papamichael and M. K.

Warby. Stability and convergence properties of Bergman kernel methods for numerical conformal mapping. Numerische Mathematik, 48(6): 639–669, June 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Paige:1993:AGT

[PW93] Christopher C. Paige and Mu Sheng Wei. Analysis of the generalized total least squares problem AX ≈ B when some columns of A are free of error. Numerische Mathematik, 65(2): 177–202, June 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pyle:1967:GIA

[Pyl67]  L. Duane Pyle.  A gener-

alized inverse ϵ-algorithm for constructing intersection projection matrices, with applications. Numerische Mathematik, 10(1): 86–102, May 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Pytlak:2002:CGC

[Pyt02]  R. Pytlak.  Comments on

“Global convergence of the method of shortest residuals” by Y. Dai and Y. Yuan. Numerische Mathematik, 91(2): 319–321, April 2002. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [hDxY99].

Pflaum:1999:EAC

[PZ99] C. Pflaum and A. Zhou. Error analysis of the combination technique. Numerische Mathematik, 84(2):327–350, December 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9084002/90840327.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9084002/90840327.pdf.

Qiao:1988:HAF

[Qia88] San Zheng Qiao. Hybrid algorithm for fast Toeplitz orthogonalization. Numerische Mathematik, 53(3):351–366, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Quarteroni:1991:CVI

[QLV91] A. Quarteroni, G. Sacchi Landriani, and A. Valli. Coupling of viscous and inviscid Stokes equations via a domain decomposition method for finite elements. Numerische Mathematik, 59(8):831–859, October 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Qun:1991:FOE

[Qun91] Lin Qun. Fourth order eigenvalue approximation by extrapolation on domains with reentrant corners. Numerische Mathematik, 58(6):631–640, January 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Qiu:1999:NRM

[QYK99] Lin Qiu, Biao Yang, and Jiaoxun Kuang. The NGPstability of Runge–Kutta methods for systems of neutral delay differential equations. Numerische Mathematik, 81(3): 451–459, January 1999. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/9081003/
90810451.htm;   http://link.
springer.de/link/service/
journals/00211/papers/9081003/
90810451.pdf.

Rainey:1969:CST

[Rai69] John W. Rainey. On comparatively stable tridiagonalization methods. Numerische Mathematik, 13(4):316–322, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rakotch:1973:NSV

[Rak73] E. Rakotch. Numerical solution of Volterra integral equations. Numerische Mathematik, 20(4): 271–279, August 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ralston:1965:RCA

[Ral65] Anthony Ralston. Rational Chebyshev approximation by Remes’ algorithms. Numerische Mathematik, 7(4):322–330, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rall:1966:CNP

[Ral66] L. B. Rall. Convergence of the Newton process to multiple solutions. Numerische Mathematik, 9(1):23–37, November 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rannacher:1979:NFE

[Ran79] Rolf Rannacher. Nonconforming finite element methods for eigenvalue problems in linear plate theory. Numerische Mathematik, 33(1):23–42, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rannacher:1984:FES

[Ran84] Rolf Rannacher. Finite element solution of diffusion problems with irregular data. Numerische Mathematik, 43(2):

309–327, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rappaz:1977:ASN

[Rap77] J. Rappaz. Approximation of the spectrum of a non-compact operator given by the magnetohydrodynamic stability of a plasma. Numerische Mathematik, 28(1):15–24, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rappaz:1984:ANN

[Rap84] Jacques Rappaz. Approximation of a nondifferentiable nonlinear problem related to MHD equilibria. Numerische Mathematik, 45(1):117–133, October 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rath:1982:FGR

[Rat82] Wolfgang Rath. Fast Givens rotations for orthogonal similarity transformations. Numerische Mathematik, 40(1):47– 56, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rathsfeld:1993:QMD

[Rat93] A. Rathsfeld. On quadrature methods for the double layer potential equation over the boundary of a polyhedron. Numerische Mathematik, 66(1):67–95, October 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Raugel:1986:FDA

[Rau86] G. Raugel. Finite-dimensional approximation of bifurcation problems in presence of symmetries. Numerische Mathematik, 48(2):137–198, February 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Razelos:1969:AAS

[Raz69] P. Razelos. On the accuracy of the analog solution of heat conduction problems involving Newtonian cooling. Numerische Mathematik, 13(4):323–333, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Razzaghi:1979:CSM

[Raz79] M. Razzaghi. A computational solution for a matrix Riccati differential equation. Numerische Mathematik, 32(3): 271–279, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reinsch:1968:RTN

[RB68] C. Reinsch and F. L. Bauer. Rational QR transformation with Newton shift for symmetric tridiagonal matrices. Numerische Mathematik, 11(3): 264–272, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 257–272].

Remar:1982:AFS

[RBS82] Jaroslav Remar, John C. Bruch, Jr., and James M. Sloss. Axisymmetric free surface seepage. Numerische Mathematik, 40(2): 143–168, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Richtmyer:1962:CFE

[RDM62] R. D. Richtmyer, Marjorie Devaney, and N. Metropolis. Continued fraction expansions of algebraic numbers. Numerische Mathematik, 4:68–84, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Redheffer:1967:BUD

[Red67] R. M. Redheffer. Bemerkungen über Differentialungleichungen bei abzählbaren Ausnahmemengen. (German) [Remark on differential equations for countable sets]. Numerische Mathematik, 9(5):437–445, April 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reddien:1973:PMS

[Red73] G. W. Reddien. Projection methods and singular two point boundary value problems. Numerische Mathematik, 21(3): 193–205, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Redheffer:1977:FND

[Red77] Ray M. Redheffer. Fehlerabschätzung bei nichtlinearer Differentialgleichungen mit Hilfe linearer Differentialgleichungen. (German) [Error estimation in nonlinear problems by use of a family of linear differential inequalities]. Numerische Mathematik, 28(4):393–405, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reddy:1988:CMF

[Red88] B. D. Reddy. Convergence of mixed finite element approximations for the shallow arch problem. Numerische Mathematik, 53(6):687–699, September 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reimer:1965:FLD

[Rei65] Manfred Reimer. Eine Fehlerabschätzung für lineare Differenzenverfahren. (German) [An error bound for linear difference methods]. Numerische Mathematik, 7(4):277–285, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reinsch:1967:SSF

[Rei67] Christian H. Reinsch. Smoothing by spline functions. Numerische Mathematik, 10(3): 177–183, October 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reimer:1971:UWP

[Rei71a] Manfred Reimer. Untersuchungen über das Wachstrum von Potenzreihenabschnitten mit Anwendungen für das Hornerschema. (German) [Investigations of the growth of segments of polynomials with applications to Horner’s scheme]. Numerische Mathematik, 17(2): 172–178, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reinsch:1971:SSF

[Rei71b] Christian H. Reinsch. Smoothing by spline functions. II. Numerische Mathematik, 16(5): 451–454, February 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reimer:1975:ESP

[Rei75]  Manfred Reimer.  Auswer-

tungsverfahren für Polynome in mehreren Variablen. (German) [Evaluation schemes for polynomials in more than one variable]. Numerische Mathematik, 23(4): 321–336, August 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reinhart:1982:NAK

[Rei82] Laure Reinhart. On the numerical analysis of the von Kármán equations: mixed finite element approximation and continuation techniques. Numerische Mathematik, 39(3):371–404, October 1982. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Reimer:1984:BCO

[Rei84a] Manfred Reimer. Best constants occurring with the modulus of continuity in the error estimate for spline interpolants of odd degree on equidistant grids. Numerische Mathematik, 44(3): 407–415, September 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reimer:1984:EES

[Rei84b] Manfred Reimer. Error estimates for spline interpolants on equidistant grids. Numerische Mathematik, 44(3): 417–424, September 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reimer:1988:OIR

[Rei88] Manfred Reimer. Optimal interval refinement near singularities. Numerische Mathematik, 53(4): 411–422, August 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reichel:1989:OTM

[Rei89a] Lothar Reichel. The ordering of tridiagonal matrices in the cyclic reduction method for Poisson’s equation. Numerische Mathematik, 56(2/3):215–227, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reimer:1989:CHS

[Rei89b]  Manfred Reimer.  Cardinal

Hermite-spline-interpolation on the equidistant lattice. Numerische Mathematik, 56(4): 345–357, October 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reichel:1990:FSM

[Rei90] Lothar Reichel. Fast solution methods for Fredholm integral equations of the second kind. Numerische Mathematik, 57(8): 719–736, August 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reitich:1991:SST

[Rei91] Fernando Reitich. Singular solutions of a transmission problem in plane linear elasticity for wedge-shaped regions. Numerische Mathematik, 59(2): 179–216, May 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reich:1997:HOS

[Rei97] Sebastian Reich. On higherorder semi-explicit symplectic partitioned Runge–Kutta methods for constrained Hamiltonian systems. Numerische Mathematik, 76(2):231–247, April 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.de/link/service/
journals/00211/bibs/7076002/

70760231.htm;   http://science.
springer.de/nmee/bibs/7076002/
70760231.htm.

Rentrop:1979:TSM

[Ren79] P. Rentrop. A Taylor series method for the numerical solution of two-point boundary value problems. Numerische Mathematik, 31(4): 359–375, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rentrop:1980:ACE

[Ren80] P. Rentrop. An algorithm for the computation of the exponential spline. Numerische Mathematik, 35(1):81–93, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rentrop:1985:PRK

[Ren85]  Peter Rentrop.  Partitioned

Runge–Kutta methods with stiffness detection and stepsize control. Numerische Mathematik, 47(4):545–564, December 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reusken:1988:CMFa

[Reu88a] Arnold Reusken. Convergence of the multigrid full approximation scheme for a class of elliptic mildly nonlinear boundary value problems. Numerische Mathematik, 52(3):251–277, March 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reusken:1988:CMFb

[Reu88b] Arnold Reusken. Convergence of the multilevel full approximation scheme including the V - cycle. Numerische Mathematik, 53(6):663–686, September 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reusken:1991:SOL

[Reu91] Arnold Reusken. Steplength optimization and linear multigrid methods. Numerische Mathematik, 58(8):819–838, March 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reusken:1995:FAR

[Reu95]  Arnold Reusken.  Fourier

analysis of a robust multigrid method for convectiondiffusion equations. Numerische Mathematik, 71(3):365–397, September 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071003/50710365.
htm;   http://science.springer.
de/nmee/bibs/5071003/50710365.
htm.

Robinson:1994:OSC

[RF94] Mark P. Robinson and Graeme Fairweather. Orthogonal spline collocation methods for Schrödingertype equations in one space variable. Numerische Mathematik, 68(3):355–375, September 1994.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4068003/
40680355.htm;   http://science.
springer.de/nmee/bibs/4068003/
40680355.htm.

Reddy:1988:VPC

[RG88] B. D. Reddy and T. B. Griffin. Variational principles and convergence of finite element approximations of a holonomic elastic-plastic problem. Numerische Mathematik, 52(1): 101–117, January 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rhee:1993:GCC

[RH93] Noah H. Rhee and Vjeran Hari. On the global and cubic convergence of a quasi-cyclic Jacobi method. Numerische Mathematik, 66(1):97–122, October 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rheinboldt:1988:CMS

[Rhe88] Werner C. Rheinboldt. On the computation of multidimensional solution manifolds of parameterized equations. Numerische Mathematik, 53(1/2): 165–181, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rice:1965:CPR

[Ric65] John R. Rice. On the conditioning of polynomial and rational forms. Numerische Mathematik, 7(5):426–435, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Richert:1974:HEP

[Ric74a] Walter R. Richert. On the handling of eigenvalue problems with the parameter of nonlinear form. Numerische Mathematik, 22(4):275–287, August 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Richert:1974:FET

[Ric74b] Walter R. Richert. Zur Fehlerabschätzung für Eigenwertaufgaben vom Typ (λ^{2}I − λA − B)x = 0. (German) [Error bounds construction for eigenvalue distances of the problem (λ^{2}I − λA − B)x = 0]. Numerische Mathematik, 22(3): 233–239, June 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Richter:1978:SPP

[Ric78] Gerard R. Richter. Superconvergence of piecewise polynomial Galerkin approximations, for Fredholm integral equations of the second kind. Numerische Mathematik, 31(1):63– 70, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rieder:1991:WTS

[Rie91] Andreas Rieder. The wavelet transform on Sobolev spaces and its approximation properties. Numerische Mathematik, 58(8):875–894, March 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rieder:1997:WMM

[Rie97] Andreas Rieder. A wavelet multilevel method for ill-posed problems stabilized by Tikhonov regularization. Numerische Mathematik, 75(4):501–522, February 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7075004/70750501.
htm;   http://science.springer.
de/nmee/bibs/7075004/70750501.
htm.

Rissanen:1974:SLE

[Ris74] Jorma Rissanen. Solution of linear equations with Hankel and Toeplitz matrices. Numerische Mathematik, 22(5):361–366, October 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ritter:1980:RSC

[Rit80] Klaus Ritter. On the rate of superlinear convergence of a class of variable metric methods. Numerische Mathematik, 35(3): 293–313, October 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rivlin:1962:PBU

[Riv62] T. J. Rivlin. Polynomials of best uniform approximation to certain rational functions. Numerische Mathematik, 4:345– 349, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Riess:1972:EEC

[RJ72] R. D. Riess and L. W. Johnson.

Error estimates for Clenshaw– Curtis quadrature. Numerische Mathematik, 18(4):345–353, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reemtsen:1984:MNS

[RK84] R. Reemtsen and A. Kirsch. A method for the numerical solution of the one-dimensional inverse Stefan problem. Numerische Mathematik, 45(2): 253–273, November 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rodrigue:1989:CCA

[RKL89] Garry Rodrigue, LiShan Kang, and Yu-Hui Liu. Convergence and comparison analysis of some numerical Schwarz methods. Numerische Mathematik, 56(2/3):123–138, October 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rudert:1966:UPL

[RL66] W. S. Rudert and H. G. Lill.

Über Partitionen und ein lineares diophantisches Problem. (German) [On partitions and a linear Diophantine problem]. Numerische Mathematik, 8(5): 407–411, August 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reemtsen:1981:ATN

[RL81] R. Reemtsen and C. J. Lozano. An approximation technique for the numerical solution of a Stefan problem. Numerische Mathematik, 38(1):141–154, October 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Robert:1972:NAA

[RM72] F. Robert and J. F. Maitre. Normes et algorithmes associés à une découpe de matrice (itération du principe de Gauss– Seidel). (French) [Norms and algorithms associated with a splitting of matrix (iteration of the Gauss–Seidel principle)]. Numerische Mathematik, 19(4): 303–325, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ryoo:1998:NVS

[RN98] Cheon Seoung Ryoo and Mitsuhiro T. Nakao. Numerical verification of solutions for variational inequalities. Numerische Mathematik, 81(2):305–320, December 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.

de/link/service/journals/
00211/bibs/8081002/80810305.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081002/80810305.
pdf.

Rowland:1970:SBD

[RO70] John H. Rowland and James A.

O’Donnell. Some bounds for the deviation and interpolation points in Chebyshev approximation. Numerische Mathematik, 16(1):16–21, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Roark:1971:ECI

[Roa71] A. L. Roark. On the eigenproblem for convolution integral equations. Numerische Mathematik, 17(1):54–61, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Robert:1966:RMP

[Rob66] F. Robert. Recherche d’une M-matrice parmi les minorantes d’un opérateur linéaire. (French) [Search for an M-matrix among the minorants of a linear operator]. Numerische Mathematik, 9(3):189–199, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Robert:1971:AEO

[Rob71]  J. Robert.  Approximations

des espace d’Orlicz et applications. (French) [Approximations of Orlicz spaces and their applications]. Numerische Mathematik, 17(4):338–356, August 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Robinson:1972:ENM

[Rob72] Stephen M. Robinson. Extension of Newton’s method to nonlinear functions with values in a cone. Numerische Mathematik, 19(4):341–347, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Robert:1976:ADT

[Rob76] François Robert. Autour du théorème de Stein–Rosenberg. (French) [More about the Stein– Rosenberg theorem]. Numerische Mathematik, 27(1): 133–141, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Roche:1988:RMD

[Roc88]  Michel Roche.  Rosenbrock

methods for differential algebraic equations. Numerische Mathematik, 52(1):45–63, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rodrigue:1973:GMM

[Rod73] Garry Rodrigue. A gradient method for the matrix eigenvalue problem Ax = λBx. Numerische Mathematik, 22(1):1– 16, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rohde:1998:UFV

[Roh98] Christian Rohde. Upwind finite volume schemes for weakly coupled hyperbolic systems of conservation laws in 2D. Numerische Mathematik, 81(1):85– 123, November 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8081001/80810085.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081001/80810085.
pdf.

Rokne:1972:NMU

[Rok72] Jon Rokne. Newton’s method under mild differentiability conditions with error analysis. Numerische Mathematik, 18(5): 401–412, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Roman:1985:CCR

[Rom85] J. Roman. Calculs de complexité relatifs à une méthode de dissection emboî tée. (French) [Relative complexity bounds for a nested dissection method]. Numerische Mathematik, 47(2): 175–190, September 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rosen:1964:PPP

[Ros64] J. B. Rosen. Primal partition programming for block diagonal matrices. Numerische Mathematik, 6(1):250–260, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rose:1975:WAE

[Ros75]  Milton E. Rose.  Weak-

element approximations to elliptic differential equations. Numerische Mathematik, 24(3): 185–204, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rosier:1995:BDR

[Ros95]  Michael Rosier.  Biorthog-

onal decompositions of the Radon transform. Numerische Mathematik, 72(2):263–283, December 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/5072002/50720263.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/5072002/50720263.
                                        htm.

Roubicek:1987:OCS

[Rou87]  Tomáš Roubíček.  Optimal

control of a Stefan problem with state-space constraints. numerical approximation. Numerische Mathematik, 50(6): 723–744, April 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rowland:1971:IIP

[Row71] John H. Rowland. Inequalities for the interpolation points in Chebyshev approximation by polynomials. Numerische Mathematik, 17(1):40–44, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Roose:1985:NCN

[RP85] Dirk Roose and Robert Piessens.

Numerical computation of nonsimple turning points and cusps. Numerische Mathematik, 46(2): 189–211, May 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rabier:1990:CMS

[RR90] Patrick J. Rabier and Werner C.

Rheinboldt. On a computational method for the second fundamental tensor and its application to bifurcation problems. Numerische Mathematik, 57(6/7):681–694, July 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Paper presented at the Conference on Approximation Theory and Numerical Linear Algebra, 30 March – 1 April, 1989, Kent, Ohio.

Rentrop:1989:ARW

[RRS89] P. Rentrop, M. Roche, and G. Steinebach. The application of Rosenbrock-Wanner type methods with stepsize control in differential-algebraic equations. Numerische Mathematik, 55(5):

545–563, August 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rutishauser:1963:TMS

[RS63] H. Rutishauser and H. R. Schwarz. The LR transformation method for symmetric matrices. Numerische Mathematik, 5(1):273–289, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Roark:1965:PRW

[RS65] A. L. Roark and L. F. Shampine. On paper of roark and wing. Numerische Mathematik, 7(5): 394–395, October 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Roark:1968:EDI

[RS68] A. L. Roark and L. F. Shampine. On the eigenproblem for displacement integral equations. Numerische Mathematik, 12(3): 170–179, October 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rakotch:1970:EEN

[RS70] E. Rakotch and J. Steinberg. Error estimate for the numerical solution of Fredholm’s integral equation. Numerische Mathematik, 15(4):320–328, August 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Russell:1972:CMB

[RS72] R. D. Russell and L. F.

Shampine. A collocation method for boundary value problems. Numerische Mathematik, 19(1):1–28, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reddien:1976:CMS

[RS76] G. W. Reddien and L. L.

Schumaker. On a collocation method for singular two point boundary value problems. Numerische Mathematik, 25(4): 427–432, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reimer:1986:EAR

[RS86] Manfred Reimer and D. Siepmann. An elementary algebraic representation of polynomial spline interpolants for equidistant lattices and its condition. Numerische Mathematik, 49(1):55–65, January 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ruotsalainen:1988:CGM

[RS88] K. Ruotsalainen and J. Saranen. On the convergence of the Galerkin method for nonsmooth solutions of integral equations. Numerische Mathematik, 54(3): 295–302, December 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rahman:1990:CSC

[RS90] Qazi I. Rahman and Gerhard Schmeisser. Characterization of the speed of convergence of the trapezoidal rule. Numerische Mathematik, 57(2): 123–138, April 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Riemenschneider:1991:CHI

[RS91] Sherman Riemenschneider and Karl Scherer. Cardinal Hermite interpolation with box splines. II. Numerische Mathematik, 58 (6):591–602, January 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Riemenschneider:1995:GIL

[RS95] Sherman D. Riemenschneider and Zuowei Shen. General interpolation on the lattice hz^{s}: compactly supported fundamental solutions. Numerische Mathematik, 70(3): 331–351, May 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070003/50700331.
htm;   http://science.springer.
de/nmee/bibs/5070003/50700331.
htm.

Raviart:1996:HAM

[RS96] Pierre-Arnaud Raviart and Eric Sonnendrücker. A hierarchy of approximate models for the Maxwell equations. Numerische Mathematik, 73(3):

329–372, May 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073003/60730329.
htm;   http://science.springer.
de/nmee/bibs/6073003/60730329.
htm.

Rappaz:1997:SPF

[RSSV97] J. Rappaz, J. Sanchez Hubert, E. Sanchez Palencia, and D. Vassiliev. On spectral pollution in the finite element approximation of thin elastic “membrane” shells. Numerische Mathematik, 75(4):473–500, February 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7075004/70750473.
htm;   http://science.springer.
de/nmee/bibs/7075004/70750473.
htm.

Reddy:1992:SML

[RT92] Satish C. Reddy and Lloyd N.

Trefethen. Stability of the method of lines. Numerische Mathematik, 62(2):235–267, July 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ruas:1985:QVP

[Rua85] Vitoriano Ruas. Quasisolenoidal velocity-pressure finite element methods for the threedimensional Stokes problem. Numerische Mathematik, 46(2): 237–253, May 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ruas:1988:QFE

[Rua88] Vitoriano Ruas. A quadratic finite element method for solving biharmonic problems in r^{n}. Numerische Mathematik, 52(1): 33–43, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ruhe:1970:PMV

[Ruh70] Axel Ruhe. Properties of a matrix with a very ill-conditioned eigenproblem. Numerische Mathematik, 15(1):57–60, May 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rumpf:1996:VAO

[Rum96] M. Rumpf. A variational approach to optimal meshes. Numerische Mathematik, 72(4): 523–540, February 1996. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/6072004/
                                        60720523.htm;   http://science.
                                        springer.de/nmee/bibs/6072004/
                                        60720523.htm.

Rund:1970:RCB

[Run70]  Hanno Rund.  The reduc-

tion of certain boundary problems to variational problems by means of transversality conditions. Numerische Mathematik, 15(1):49–56, May 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Russell:1974:CSB

[Rus74] Robert D. Russell. Collocation for systems of boundary value problems. Numerische Mathematik, 23(2):119–133, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ruscheweyh:1984:GDM

[Rus84] Stephan Ruscheweyh. On a global descent method for polynomials. Numerische Mathematik, 45(2):227–240, November 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rutishauser:1960:BNI

[Rut60] Heinz Rutishauser. Bemerkungen zur numerischen Integration gewöhnlicher Differentialgleichungen n-ter Ordnung. (German) [Remark on the numerical integration of n-th order ordinary differential equations]. Numerische Mathematik, 2(1): 263–279, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rutishauser:1963:ARP

[Rut63a] H. Rutishauser. Ausdehnung des Rombergschen Prinzips. (German) [Extension of the Romberg principle]. Numerische Mathematik, 5(1):48– 54, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rutishauser:1963:SSQ

[Rut63b]  Heinz Rutishauser.  Stabile

Sonderfälle des Quotienten-Differenzen-Algorithmus. (German) [Stable special cases of the quotient-difference algorithm]. Numerische Mathematik, 5(1): 95–112, December 1963. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rutishauser:1966:BEO

[Rut66a] H. Rutishauser. Bestimmung der Eigenwerte orthogonaler Matrizen. (German) [Determination of the Eigenvalues of Orthogonal Matrices]. Numerische Mathematik, 9(2):104– 108, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rutishauser:1966:JMR

[Rut66b] H. Rutishauser. The Jacobi method for real symmetric matrices. Numerische Mathematik, 9(1):1–10, November 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 202– 211].

Rutishauser:1969:CAF

[Rut69] Heinz Rutishauser. Computational aspects of F. L. Bauer’s simultaneous iteration method. Numerische Mathematik, 13(1): 4–13, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rutishauser:1970:SIM

[Rut70] H. Rutishauser. Simultaneous iteration method for symmetric matrices. Numerische Mathematik, 16(3):205–223, December 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 284–302].

Ramm:1990:CSI

[RvdS90] A. G. Ramm and A. van der Sluis. Calculating singular integrals as an ill-posed problem. Numerische Mathematik, 57(2): 139–145, April 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Roark:1965:MCE

[RW65] A. L. Roark and G. M. Wing. A method for computing the eigenvalues of certain integral equations. Numerische Mathematik, 7(2):159–170, April 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Robert:1985:EEI

[RW85a] R. Robert and P. Witomski.

Sur une équation d’évolution intégro-différentielle singulière linéaire. (French) [A singular integro-differential evolution equation]. Numerische Mathematik, 46(4):493–498, August 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rye:1985:RVR

[RW85b] Egil Rye and Haakon Waadeland. Reflections on value regions, limit regions and truncation errors for continued fractions. Numerische Mathematik, 47(2):191–215, September 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ruotsalainen:1988:BEM

[RW88] K. Ruotsalainen and W. Wendland. On the boundary element method for some nonlinear boundary value problems. Numerische Mathematik, 53(3): 299–314, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ryan:1967:SMC

[Rya67] Joseph T. Ryan. A sieve method for classifying composite numbers. Numerische Mathematik, 10(4):324–326, November 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sanchez:1984:EEM

[SA84] A. M. Sanchez and R. Arcangeli. Estimations des erreurs de meilleure approximation polynômiale et d’interpolation de Lagrange dans les espaces de Sobolev d’ordre non entier. (French) [Estimates of the best polynomial approximation errors and the Lagrange interpolation error in fractionalorder Sobolev spaces]. Numerische Mathematik, 45(2):

301–321, November 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saad:1986:CNS

[Saa86] Youcef Saad. On the condition number of some Gram matrices arising from least squares approximation in the complex plane. Numerische Mathematik, 48(3):337–347, March 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sablonniere:1978:BTS

[Sab78] Paul Sablonnière. Bases de type Schauder de C[0, 1] et formules de quadrature associèes. (French) [Bases of Schauder type in C[0, 1] and associated quadrature formulas]. Numerische Mathematik, 30(2): 227–240, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sack:1972:FSR

[Sac72] R. A. Sack. A fully stable rational version of the QR algorithm for tridiagonal matrices. Numerische Mathematik, 18(5): 432–441, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sachs:1973:IMN

[Sac73] A. Sachs. Iterationsverfahren für monotone, nicht notwendig Lipschitz-beschränkte Operatoren im Hilbert-Raum. (German) [Iteration method for monotone not-necessarily Lipschitzlimited operators in Hilbert space]. Numerische Mathematik, 20(5):356–363, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

SacchiLandriani:1988:STA

[Sac88] Giovanni Sacchi Landriani. Spectral tau approximation of the two-dimensional Stokes problem. Numerische Mathematik, 52(6):683–699, June 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sadkane:1993:BAD

[Sad93a] Miloud Sadkane. Block-Arnoldi and Davidson methods for unsymmetric large eigenvalue problems. Numerische Mathematik, 64(2):195–211, February 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sadkane:1993:BAC

[Sad93b]  Miloud Sadkane.  A block

Arnoldi-Chebyshev method for computing the leading eigenpairs of large sparse unsymmetric matrices. Numerische Mathematik, 64(2):181–193, February 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sahba:1985:SNI

[Sah85]  M. Sahba.  On the solu-

tion of nonlinear inequalities in a finite number of iterations. Numerische Mathematik, 46(2): 229–236, May 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sakai:1997:ISP

[Sak97]  Manabu Sakai.  Inflections

and singularity on parametric rational cubic curves. Numerische Mathematik, 76(3): 403–417, May 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076003/70760403.
htm;   http://science.springer.
de/nmee/bibs/7076003/70760403.
htm.

Salzer:1960:OPN

[Sal60] Herbert E. Salzer. Optimal points for numerical differentiation. Numerische Mathematik, 2(1):214–227, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Salzer:1961:SEB

[Sal61] Herbert E. Salzer. Some extensions of Bairstow’s method. Numerische Mathematik, 3:120– 124, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Salzer:1963:EWF

[Sal63] Herbert E. Salzer. Equallyweighted formulas for numerical differentiation. Numerische Mathematik, 4:381–392, December 1963. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Salzer:1964:DDF

[Sal64] H. E. Salzer. Divided differences for functions of two variables for irregularly spaces arguments. Numerische Mathematik, 6(1): 68–77, December 1964. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Salzer:1971:FBE

[Sal71] Herbert E. Salzer. Formulas for best extrapolation. Numerische Mathematik, 18(2): 144–153, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Salkauskas:1975:WFC

[Sal75] K. Salkauskas. On weight functions for Chebyshev quadrature. Numerische Mathematik, 24(1): 13–18, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sanchez:1975:ISB

[Sán75] David A. Sánchez. An iteration scheme for boundary value/alternative problems. Numerische Mathematik, 23(3): 223–230, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saridakis:1986:AUS

[Sar86] Yiannis G. Saridakis. On the analysis of the unsymmetric successive overrelaxation method when applied to p-cyclic matrices. Numerische Mathematik, 49(5):461–473, September 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saranen:1988:CED

[Sar88] J. Saranen. The convergence of even degree spline collocation solution for potential problems in smooth domains of the plane. Numerische Mathematik, 53(5): 499–512, August 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saranen:1989:EMS

[Sar89]  Jukka Saranen.  Extrapola-

tion methods for spline collocation solutions of pseudodifferential equations on curves. Numerische Mathematik, 56(4): 385–407, October 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saridakis:1990:DDU

[Sar90] Yiannis G. Saridakis. Domains of divergence of the USSOR method applied on p-cyclic matrices. Numerische Mathematik, 57(4):405–412, June 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saridakis:1996:REC

[Sar96] Yiannis G. Saridakis. Regions of exact convergence of the USSOR method applied on generalized consistently ordered matrices. Numerische Mathematik, 74(2): 203–219, August 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074002/60740203.
htm;   http://science.springer.
de/nmee/bibs/6074002/60740203.
htm.

Sarkis:1997:NCS

[Sar97] Marcus Sarkis. Nonstandard coarse spaces and Schwarz methods for elliptic problems with discontinuous coefficients using non-conforming elements. Numerische Mathematik, 77(3): 383–406, September 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7077003/
70770383.htm;   http://science.
springer.de/nmee/bibs/7077003/
70770383.htm.

Sauer:1963:DHA

[Sau63]  Robert Sauer.  Differen-

zenverfahren für hyperbolische Anfangswertprobleme bei mehr als zwei unabhängigen Veränderlichen mit Hilfe von Nebencharakteristiken. (German) [Difference method for hyperbolic initial value problems in more than two independent variables with the aid of secondary characteristics]. Numerische Mathematik, 5(1):55– 67, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sauer:1967:AWE

[Sau67] R. Sauer. Alwin Watther em.

ord. Professor an der Technischen Hochschule Darmstadt. Numerische Mathematik, 9(4): 379, February 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sautter:1973:SLO

[Sau73] W. Sautter. Schwankungsbereiche linearer Optimierungsaufgaben. (German) [Perturbation analysis of linear programming problems]. Numerische Mathematik, 21(3):264–269, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sautter:1978:FGE

[Sau78] Werner Sautter. Fehleranalyse für die Gauß-Elimination zur Berechnung der Lösung minimaler Länge. (German) [Error Analysis of Gauss Elimination for the Best Least Squares Problem]. Numerische Mathematik, 30(2):165–184, June 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sautter:1984:KLA

[Sau84] Werner Sautter. Zur Kondition des linearen Ausgleichsproblems mit linearen Gleichungen als Nebenbedingungen. (German) [On the condition of the linear least squares problem with linear equations as side conditions]. Numerische Mathematik, 44(1):139–152, June 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sauer:1997:PIM

[Sau97] Thomas Sauer. Polynomial interpolation of minimal degree. Numerische Mathematik, 78(1): 59–85, November 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078001/70780059.
htm;   http://science.springer.
de/nmee/bibs/7078001/70780059.
htm.

Savare:1993:SAA

[Sav93]  Giuseppe Savaré.  A(Θ)-

stable approximations of abstract Cauchy problems. Numerische Mathematik, 65(3): 319–335, July 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saworin:1982:EEN

[Saw82] A. N. Saworin. Über die Effektivität einiger nichtlinearer Verfahren bei der numerischen Behandlung steifer Differentialgleichungssysteme. (German) [On the efficiency of some nonlinear methods for the numerical solution of systems of stiff differential equations]. Numerische Mathematik, 40(2): 169–177, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stiefel:1969:SCM

[SB69] E. Stiefel and D. G. Bettis. Stabilization of Cowell’s method. Numerische Mathematik, 13(2): 154–175, May 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shampine:1983:SEM

[SB83] Lawrence F. Shampine and Lorraine S. Baca. Smoothing the extrapolated midpoint rule. Numerische Mathematik, 41(2): 165–175, June 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shanahan:1990:TFE

[SB90] Robert M. Shanahan and John W. Barrett. Total flux estimates for a finite element approximation of the Dirichlet problem using the boundary penalty method. Numerische Mathematik, 57(4): 351–363, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Smarzewski:1984:QS

[SBH84] R. Smarzewski and A. Bujalska-Horbowicz. Quartic X-splines. Numerische Mathematik, 45(1): 135–148, October 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schopf:1960:KI

[Sch60] Andreas H. Schopf. On the Kantorovich inequality. Numerische Mathematik, 2(1):344– 346, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=2&issue=1&spage=
344.

Schafke:1961:VBC

[Sch61a] Friedrich Wilhelm Schäfke. Ein Verfahren zur Berechnung des charakteristischen Exponenten der Mathieuschen Differentialgleichung. I. (German) [A method for calculation of characteristic exponents of the Mathiew differential equation. I]. Numerische Mathematik, 3: 30–38, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schonhage:1961:FIG

[Sch61b] A. Schönhage. Fehlerfortpflanzung bei Interpolation. (German) [Error transmission in interpolation]. Numerische Mathematik, 3:62–71, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schonhage:1961:KJV

[Sch61c] A. Schönhage. Zur Konvergenz des Jacobi–Verfahrens. (German) [On the convergence of the Jacobi method]. Numerische Mathematik, 3(1):374– 380, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
374.

Schroder:1961:FRG

[Sch61d]  Johann Schröder.  Fehlera-

bschätzung mit Rechenanlagen bei gewönlichen Differentialgleichungen erster Ordnung. (German) [Computational error bounds for first-order ordinary differential equations]. Numerische Mathematik, 3:39– 61, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schroder:1961:VFG

[Sch61e] Johann Schröder. Verbesserung einer Fehlerabschätzung für gewöhnliche Differentialgleichungen erster Ordnung. (German) [Improvement of error bounds for first-order ordinary differential equations]. Numerische Mathematik, 3:125–130, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schonhage:1963:OPD

[Sch63] A. Schönhage. Optimale Punkte für Differentiation und Integration. (German) [Optimal points for differentiation and integration]. Numerische Mathematik, 5(1):303–331, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1964:AMI

[Sch64a]  Jochen W. Schmidt.  Aus-

gangsvektoren für monotone Iterationen bei linearen Gleichungsystemen. (German) [Starting vectors for monotone iterations of linear equation systems]. Numerische Mathematik, 6(1): 78–88, December 1964. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneider:1964:BND

[Sch64b] A. Schneider. Bemerkungen zur numerischen Differentiation. (German) [Remark on numerical differentiation]. Numerische Mathematik, 6(1):332– 334, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneider:1965:POI

[Sch65] Hans Schneider. Positive operators and an inertia theorem. Numerische Mathematik, 7(1):11– 17, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1966:AEK

[Sch66a] Jochen W. Schmidt. Asymptotische Einschließung bei konvergenzbeschleunigenden Verfahren. (German) [Asymptotic bounds for convergence accelerating methods]. Numerische Mathematik, 8(2):105– 113, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schroder:1966:OUI

[Sch66b] Johann Schröder. Operator-Ungleichungen und ihre numerische Anwendung bei Randwertaufgaben. (German) [Operator inequalities and their numerical use in boundary value problems]. Numerische Mathematik, 9(2):149–162, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schafke:1968:ZGL

[Sch68a] F. W. Schäfke. Zum Zeilensummenkriterium. (german) [On the line sum criterion]. Numerische Mathematik, 12(5): 448–453, December 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1968:AEK

[Sch68b] Jochen W. Schmidt. Asymptotische Einschließung bei konvergenzbeschleunigenden Verfahren. II. (German) [Asymptotic enclosure with convergence acceleration method. II]. Numerische Mathematik, 11(1):53– 56, January 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schwarz:1968:TSB

[Sch68c] H. R. Schwarz. Tridiagonalization of a symmetric band matrix. Numerische Mathematik, 12(4):231–241, November 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). Also in [WR71, pp. 273– 283].

Schaefer:1970:ANE

[Sch70a]  Helmut H. Schaefer.  Eine

Abschätzung der nichttrivialen Eigenwerte stochastischer Matrizen. (German) [An approximation of nontrivial eigenvalues of stochastic matrices]. Numerische Mathematik, 15(3): 219–223, July 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schafke:1970:AEI

[Sch70b] F. W. Schäfke. Zum Anwendungsbereich einiger Iterationsverfahren. (German) [On the application of an iteration scheme]. Numerische Mathematik, 15(2):109–121, June 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schonhage:1970:MRI

[Sch70c] Arnold Schönhage. Mehrdimensionale Romberg-Integration. (German) [Multidimensional Romberg integration]. Numerische Mathematik, 14(3): 299–304, February 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schroder:1970:PIP

[Sch70d]  Johann Schröder.  Proving

inverse-positivity of linear operators by reduction. Numerische Mathematik, 15(2): 100–108, June 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schafer:1971:KAL

[Sch71a] Eugen Schäfer. Ein Konstruktionsverfahren bei allgemeiner linearer Approximation. (German) [A construction method for general linear approximation]. Numerische Mathematik, 18(2):113–126, April 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schryer:1971:NMC

[Sch71b] N. L. Schryer. Newton’s method for convex nonlinear elliptic boundary value problems. Numerische Mathematik, 17(4): 284–300, August 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmeisser:1972:OQS

[Sch72a] Gerhard Schmeisser. Optimale Quadraturformeln mit semidefiniten Kernen. (German) [Optimal quadrature formulas with semidefinite kernels]. Numerische Mathematik, 20(1):32– 53, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schryer:1972:SMN

[Sch72b] N. L. Schryer. Solution of monotone nonlinear elliptic boundary value problems. Numerische Mathematik, 18(4):336–344, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schaback:1973:KAE

[Sch73a] Robert Schaback. Konstruktion und algebraische Eigenschaften von M-Spline-Interpolierenden. (German) [Construction and algebraic properties of M spline interpolation]. Numerische Mathematik, 21(2): 166–180, April 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schiop:1973:SRP

[Şch73b] Alexandru I. Şchiop. Stability of Ritz procedure for nonlinear two-point boundary value problem. Numerische Mathematik, 20(3):208–212, June 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schonhage:1973:UTG

[Sch73c] A. Schönhage. Unitäre Transformationen großer Matrizen. (German) [Unitary transformation of large matrices]. Numerische Mathematik, 20(5): 409–417, October 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schuller:1974:OCC

[Sch74a] Günther Schuller. On the order of convergence of certain quasi-Newton methods. Numerische Mathematik, 23(2): 181–192, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schwarz:1974:MCO

[Sch74b] H. R. Schwarz. The method of coordinate overrelaxation for (A − λB)x = 0. Numerische Mathematik, 23(2): 135–151, April 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schiop:1975:SRP

[Şch75a] Alexandru I. Şchiop. On the stability of the Ritz procedure for nonlinear problems. Numerische Mathematik, 23(4):305–310, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneider:1975:VRR

[Sch75b] Claus Schneider. Vereinfachte Rekursionen zur Richardson-Extrapolation in Spezialfällen. (German) [Simplified recursions for Richardson extrapolation in special cases]. Numerische Mathematik, 24(2): 177–184, April 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schroder:1975:ULB

[Sch75c] Johan Schröder. Upper and lower bounds for solutions of generalized two-point boundary value problems. Numerische Mathematik, 23(5):433–457, October 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schumacher:1975:IOS

[Sch75d] Konrad Schumacher. Iterative Optimierungsverfahren, die unter schwachen Voraussetzungen konvergien. (German) [Iterative optimisation, convergence under weak assumptions]. Numerische Mathematik, 24(5): 443–456, October 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Scheurle:1977:SPI

[Sch77a] Jürgen Scheurle. Ein selektives Projektions-Iterationsverfahren und Anwendungen auf Verzweigungsprobleme. (German) [A selective projection-iteration method and bifurcation]. Numerische Mathematik, 29(1): 11–35, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmeisser:1977:RRM

[Sch77b] Gerhard Schmeisser. A representation for the remainder of the Maclaurin quadrature formula. Numerische Mathematik, 27(3):355–358, September 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmid:1978:CFM

[Sch78a] Hans Joachim Schmid. On cubature formulae with a minimal number of knots. Numerische Mathematik, 31(3): 281–297, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1978:EIE

[Sch78b]  Jochen W. Schmidt.  Ein-

schließung inverser Elemente durch Fixpunktverfahren. (German) [Enclosing of inverse elements by fixed-points methods]. Numerische Mathematik, 31(3): 313–320, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schultz:1978:AAN

[Sch78c] Reinhart Schultz. Ein Abstiegsverfahren für Approximationsaufgaben in normierten Räumen. (German) [A descent method for approximation problems in normed spaces]. Numerische Mathematik, 31(1): 77–95, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schafer:1979:FEN

[Sch79a]  Eugen Schäfer.  Fehlerab-

schätzungen für Eigenwertnäherungen nach der Ersatzkernmethode bei Integralgleichungen. (German) [Error estimates for eigenvalues of integral equations computed by approximate kernel methods]. Numerische Mathematik, 32(3):281–290, September 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schomberg:1979:MCI

[Sch79b] Hermann Schomberg. Monotonically convergent iterative methods for nonlinear systems of equations. Numerische Mathematik, 32(1):97–104, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneider:1980:PAS

[Sch80a] Claus Schneider. Produktintegration mit nicht-äquidistanten Stützstellen. (German) [Product integration with non-equidistant knots]. Numerische Mathematik, 35(1):35–43, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schulzendorff:1980:RZB

[Sch80b] Bernd Schulzendorff. Reguläre Zerlegungen und Berechnung des Spektralradius nichtnegativer Matrizen. (German) [Regular splittings and computation of the spectral radius of nonnegative matrices]. Numerische Mathematik, 34(2): 201–216, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Scherer:1981:OGE

[Sch81a]  Karl Scherer.  On optimal

global error bounds obtained by scaled local error estimates.

Numerische Mathematik, 36(2): 151–176, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schittkowski:1981:NPMa

[Sch81b] Klaus Schittkowski. The nonlinear programming method of Wilson, Han, and Powell with an augmented Lagrangian type line search function. I. convergence analysis. Numerische Mathematik, 38(1):83–114, October 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schittkowski:1981:NPMb

[Sch81c] Klaus Schittkowski. The nonlinear programming method of Wilson, Han, and Powell with an augmented Lagrangian type line search function. II. an efficient implementation with linear least squares subproblems. Numerische Mathematik, 38(1): 115–127, October 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schaback:1982:FFK

[Sch82] R. Schaback. Fehlerabschätzungen für Koeffizienten von Exponentialsummen und Polynomen. (German) [Error estimates for coefficients of exponential sums and polynomials]. Numerische Mathematik, 39(2):293–307, August 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmitt:1983:NBR

[Sch83]  Bernhard Schmitt.  Norm

bounds for rational matrix functions. Numerische Mathematik, 42(3):379–389, December 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schroder:1984:NEB

[Sch84] Johann Schröder. Numerical error bounds for fourth order boundary value problems, simultaneous estimation of u(x) and u′′(x). Numerische Mathematik, 44(2):233–245, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schaback:1985:CAG

[Sch85a]  Robert Schaback.  Conver-

gence analysis of the general Gauss–Newton algorithm. Numerische Mathematik, 46(2): 281–309, May 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schippers:1985:MMB

[Sch85b] H. Schippers. Multigrid methods for boundary integral equations. Numerische Mathematik, 46(3):351–363, June 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Scholz:1986:NSO

[Sch86] Reinhard Scholz. Numerical solution of the obstacle problem by the penalty method. II. time-dependent problems. Numerische Mathematik, 49(2/3): 255–268, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1987:SCS

[Sch87a] Günther Schmidt. Spline collocation for singular integrodifferential equations over (0, 1). Numerische Mathematik, 50(3): 337–352, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1987:TLC

[Sch87b] Jochen W. Schmidt. On tridiagonal linear complementarity problems. Numerische Mathematik, 51(1):11–21, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneid:1987:CLI

[Sch87c] Josef Schneid. B-convergence of Lobatto iiic formulas. Numerische Mathematik, 51(2): 229–235, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schock:1987:SIM

[Sch87d] Eberhard Schock. Semi-iterative methods for the approximate solution of ill-posed problems. Numerische Mathematik, 50(3): 263–271, May 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneider:1988:ARQ

[Sch88a] Claus Schneider. An analysis of Ralston’s quadrature. Numerische Mathematik, 52(5): 597–603, May 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schock:1988:PRA

[Sch88b] Eberhard Schock. Pointwise rational approximation and iterative methods for ill-posed problems. Numerische Mathematik, 54(1):91–103, October 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schild:1990:GCD

[Sch90] K. H. Schild. Gaussian collocation via defect correction. Numerische Mathematik, 58(4): 369–386, November 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schneider:1991:SSC

[Sch91] R. Schneider. Stability of a spline collocation method for strongly elliptic multidimensional singular integral equations. Numerische Mathematik, 58(8):855–873, March 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Scherzer:1993:CRI

[Sch93] O. Scherzer. Convergence rates of iterated Tikhonov regularized solutions of nonlinear ill-posed problems. Numerische Mathematik, 66(2):259–279, November 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schittkowski:1994:PES

[Sch94a]  K. Schittkowski.  Parame-

ter estimation in systems of nonlinear equations. Numerische Mathematik, 68(1): 129–142, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680129.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680129.
htm.

Schmidt:1994:BED

[Sch94b]  Günther Schmidt.  Bound-

ary element discretization of Poincaré–Steklov operators. Numerische Mathematik, 69(1):83– 101, November 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069001/40690083.
htm;   http://science.springer.
de/nmee/bibs/4069001/40690083.
htm.

Schmidt:1995:AVS

[Sch95a] Wolfgang M. Schmidt. Appendix to “Approximation of viscosity solutions of elliptic partial differential equations on minimal grids” by M. Kocan: Approximation to orthogonal bases in R^{n} by orthogonal bases with integer coordinates. Numerische Mathematik, 72(1):117–122, November 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5072001/50720117.
htm;   http://science.springer.
de/nmee/bibs/5072001/50720117.
htm.

Schneider:1995:CRM

[Sch95b] Stefan Schneider. Convergence results for multistep Runge– Kutta methods on stiff mechanical systems. Numerische Mathematik, 69(4):495–508, February 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069004/50690495.
htm;   http://science.springer.
de/nmee/bibs/5069004/50690495.
htm.

Schwab:1996:PME

[Sch96] C. Schwab. A-posteriori modeling error estimation for hierarchic plate models. Numerische Mathematik, 74(2): 221–259, August 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074002/60740221.
htm;   http://science.springer.
de/nmee/bibs/6074002/60740221.
htm.

Schropp:1997:NMP

[Sch97]  J. Schropp.  A note on

minimization problems and multistep methods. Numerische Mathematik, 78(1):87– 101, November 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078001/70780087.
htm;   http://science.springer.
de/nmee/bibs/7078001/70780087.
htm.

Scherzer:1998:IML

[Sch98] Otmar Scherzer. An iterative multi level algorithm for solving nonlinear ill-posed problems. Numerische Mathematik, 80(4): 579–600, October 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8080004/
80800579.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8080004/
80800579.pdf.

Schoberl:1999:MMP

[Sch99] Joachim Schöberl. Multigrid methods for a parameter dependent problem in primal variables. Numerische Mathematik, 84(1):97–119, November 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9084001/90840097.htm;   http:

                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9084001/90840097.pdf.

Scott:1973:FEC

[Sco73] Ridgway Scott. Finite element convergence for singular data. Numerische Mathematik, 21(4): 317–327, August 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1967:FPF

[SD67] Jochen W. Schmidt and Hartmut Dressel. Fehlerabschätzungen bei Polynomgleichungen mit dem Fixpunktsatz von Brouwer. (German) [Error estimates by polynomial equations with Brouwer’s fixed point estimate]. Numerische Mathematik, 10(1): 42–50, May 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sack:1972:AGQ

[SD72] R. A. Sack and A. F. Donovan. An algorithm for Gaussian quadrature given modified moments. Numerische Mathematik, 18(5):465–478, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1994:PAM

[SD94] Mechthild Stoer and Geir Dahl. A polyhedral approach to multicommodity survivable network design. Numerische Mathematik, 68(1): 149–167, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680149.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680149.
htm.

Saha:1991:SOT

[SDK91] S. Saha, Purna C. Das, and N. N. Kishore. A shapeoptimization technique for the capillary surface problem. Numerische Mathematik, 58(7): 737–757, February 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Seewald:1982:QDA

[See82] Wolfgang Seewald. Quotienten-Differenzen-Algorithmus: Beweis der Regeln von Rutishauser. (German) [Quotient-difference algorithm: proof of Rutishauser’s rule]. Numerische Mathematik, 40(1):93–98, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schafke:1962:VBC

[SEG62] Friedrich Wilhelm Schäfke, Rudolf Ebert, and Helmut Groh. Ein Verfahren zur Berechnung des charakteristischen Exponenten der Mathieuschen Differentialgleichung. II. (German) [A method for calculation of characteristic exponents of the Mathiew differential equation. II]. Numerische Mathematik, 4:

1–7, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Segeth:1999:PEE

[Seg99] K. Segeth. A posteriori error estimation with the finite element method of lines for a nonlinear parabolic equation in one space dimension. Numerische Mathematik, 83(3):455–475, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083003/90830455.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083003/90830455.pdf.

Seidman:1963:SCD

[Sei63] Thomas I. Seidman. On the stability of certain difference schemes. Numerische Mathematik, 5(1):201–210, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Seoane:1995:FDA

[Seo95] M. L. Seoane. Finite-dimensional approximation of two-parameter nonlinear problems. Numerische Mathematik, 70(3): 353–377, May 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/5070003/50700353.
                                        htm;   http://science.springer.

de/nmee/bibs/5070003/50700353.
htm.

Seppala:1972:NTL

[Sep72] Veikko Seppälä. A note on the theory of limit processes for solving finite equations. Numerische Mathematik, 20(1):80– 86, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Serra:1999:KTT

[Ser99a] Stefano Serra. A Korovkintype theory for finite Toeplitz operators via matrix algebras. Numerische Mathematik, 82(1):117–142, March 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082001/90820117.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082001/90820117.pdf.

Serra:1999:RCT

[Ser99b]  Stefano Serra.  The rate

of convergence of Toeplitz based PCG methods for second order nonlinear boundary value problems. Numerische Mathematik, 81(3):461–495, January 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/9081003/90810461.
htm;   http://link.springer.
de/link/service/journals/

                                        00211/papers/9081003/90810461.
                                        pdf.

Seydel:1979:NCBb

[Sey79a]  Rüdiger Seydel.  Numerical

computation of branch points in nonlinear equations. Numerische Mathematik, 33(3): 339–352, September 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Seydel:1979:NCBa

[Sey79b] Rüdiger Seydel. Numerical computation of branch points in ordinary differential equations. Numerische Mathematik, 32(1): 51–68, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Seydel:1983:BSB

[Sey83] R. Seydel. Branch switching in bifurcation problems for ordinary differential equations. Numerische Mathematik, 41(1): 93–116, April 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Suzuki:1986:RBR

[SF86] Takashi Suzuki and Hiroshi Fujita. A remark on the L^{∞} bounds of the Ritz operator associated with a finite element approximation. Numerische Mathematik, 49(5): 529–544, September 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Simeon:1993:DIM

[SFR93] B. Simeon, C. Führer, and P. Rentrop. The Drazin inverse in multibody system dynamics. Numerische Mathematik, 64(4): 521–539, April 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schafke:1962:BEM

[SG62a] F. W. Schäfke and H. Groh.

Zur Berechnung der Eigenwerte der Mathieuschen Differentialgleichung. (German) [Toward computation of eigenvalues of the Mathieu differential equation]. Numerische Mathematik, 4:64–67, December 1962. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schafke:1962:BES

[SG62b] Friedrich Wilhelm Schäfke and Helmut Groh. Zur Berechnung der Eigenwerte der Sphäroiddifferentialgleichung. (German) [On the calculation of the eigenvalues of the spheroidal differential equation]. Numerische Mathematik, 4:310– 312, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spedicato:1978:SCV

[SG78] E. Spedicato and J. Greenstadt. On some classes of variationally derived quasi-Newton methods for systems of nonlinear algebraic equations. Numerische Mathematik, 29(4):

363–380, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1995:ASM

[SH95] Jochen W. Schmidt and Walter Heß. An always successful method in univariate convex C^{2} interpolation. Numerische Mathematik, 71(2): 237–252, August 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/5071002/50710237.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/5071002/50710237.
                                        htm.

Shaw:1967:SMF

[Sha67] B. Shaw. Some multistep formulae for special high order ordinary differential equations. Numerische Mathematik, 9(4):367– 378, February 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shampine:1968:EBV

[Sha68] L. F. Shampine. Error bounds and variational methods for nonlinear boundary value problems. Numerische Mathematik, 12(5):410–415, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shah:1970:TDP

[Sha70]  Jayant M. Shah.  Two-

dimensional polynomial splines. Numerische Mathematik, 15(1):

1–14, May 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shapiro:1982:OSM

[Sha82]  Alexander Shapiro.  Opti-

mally scaled matrices, necessary and sufficient conditions. Numerische Mathematik, 39(2): 239–245, August 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shampine:1984:ABE

[Sha84] L. F. Shampine. Asymptotic bounds on the errors of one-step methods. Numerische Mathematik, 45(2):201–206, November 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sherman:1977:CGA

[She77] B. F. Sherman. A counterexample to Greenberg’s algorithm for solving linear inequalities. Numerische Mathematik, 27(4): 491–492, December 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shen:1992:EES

[She92] Jie Shen. On error estimates of some higher order projection and penalty-projection methods for Navier–Stokes equations. Numerische Mathematik, 62(1): 49–73, June 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shen:1994:RPE

[She94] Jie Shen. Remarks on the pressure error estimates for the projection methods. Numerische Mathematik, 67(4): 513–520, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067004/40670513.
htm;   http://science.springer.
de/nmee/bibs/4067004/40670513.
htm.

Schmidt:1985:CAN

[SHH85] Jochen W. Schmidt, Wolfgang Hoyer, and Christian Haufe. Consistent approximations in Newton-type decomposition methods. Numerische Mathematik, 47(3): 413–425, November 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shieh:1978:CCG

[Shi78] A. S. L. Shieh. On the convergence of the conjugate gradient method for singular capacitance matrix equations from the Neumann problem of the Poisson equation. Numerische Mathematik, 29(3):307–327, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shieh:1979:FPS

[Shi79] A. S. L. Shieh. Fast Poisson solvers on general twodimensional regions for the Dirichlet problem. Numerische Mathematik, 31(4):405–429, December 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shirey:1982:MEQ

[Shi82] James E. Shirey. Minimization of extended quadratic functions. Numerische Mathematik, 39(2): 157–161, August 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shroff:1991:PAE

[Shr91] Gautam M. Shroff. A parallel algorithm for the eigenvalues and eigenvectors of a general complex matrix. Numerische Mathematik, 58(8):779–805, March 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sidi:1982:ASC

[Sid82] Avram Sidi. An algorithm for a special case of a generalization of the Richardson extrapolation process. Numerische Mathematik, 38(3):299–307, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sidi:1990:GRE

[Sid90] Avram Sidi. On a generalization of the Richardson extrapolation process. Numerische Mathematik, 57(4):365–377, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Siebert:1996:PEE

[Sie96] Kunibert G. Siebert. An a posteriori error estimator for anisotropic refinement. Numerische Mathematik, 73(3): 373–398, May 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6073003/60730373.
htm;   http://science.springer.
de/nmee/bibs/6073003/60730373.
htm.

Sikkema:1959:GAB

[Sik59] P. C. Sikkema. Über den Grad der Approximation mit Bernstein–Polynomen. (German) [On the degree of approximation with Bernstein polynomials]. Numerische Mathematik, 1(1):221–239, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
221.

Sikkema:1961:WEK

[Sik61] P. C. Sikkema. Der Wert einiger Konstanten in der Theorie der Approximation mit Bernstein-Polynomen.. (German) [The value of some constants in the approximation theory for Bernstein polynomials]. Numerische Mathematik, 3:107–116, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sikorski:1982:BO

[Sik82a] K. Sikorski. Bisection is optimal. Numerische Mathematik, 40(1):111–117, November 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sikorski:1982:OQA

[Sik82b] K. Sikorski. Optimal quadrature algorithms in H_{p} spaces. Numerische Mathematik, 39 (3):405–410, October 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/
10.1007/BF01407871.

Sikorski:1984:OSN

[Sik84] K. Sikorski. Optimal solution of nonlinear equations satisfying a Lipschitz condition. Numerische Mathematik, 43(2): 225–240, March 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Simpson:1978:NIS

[Sim78] I. C. Simpson. Numerical integration over a semi-infinite interval, using the lognormal distribution. Numerische Mathematik, 31(1):71–76, March 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Simoncini:1998:MAA

[Sim98] V. Simoncini. A matrix analysis of Arnoldi and Lanczos methods. Numerische Mathematik, 81(1):125–141, November 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8081001/80810125.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081001/80810125.
pdf.

Shi:1997:NSP

[SJX97] Zhong-Ci Shi, Bin Jiang, and Weimin Xue. A new superconvergence property of Wilson nonconforming finite element. Numerische Mathematik, 78(2): 259–268, December 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7078002/
70780259.htm;   http://science.
springer.de/nmee/bibs/7078002/
70780259.htm.

Shimasaki:1973:NSI

[SK73] Masaaki Shimasaki and Takeshi Kiyono. Numerical solution of integral equations in Chebyshev series. Numerische Mathematik, 21(5):373–380, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sauter:1996:ENI

[SK96] S. A. Sauter and A. Krapp. On the effect of numerical integration in the Galerkin boundary element method. Numerische Mathematik, 74(3):337–359, September 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074003/60740337.
htm;   http://science.springer.
de/nmee/bibs/6074003/60740337.
htm.

Skeel:1986:TWE

[Ske86] Robert D. Skeel. Thirteen ways to estimate global error. Numerische Mathematik, 48(1):1– 20, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sapidis:1988:MCT

[SKL88] N. S. Sapidis, P. D. Kaklis, and T. A. Loukakis. A method for computing the tension parameters in convexitypreserving spline-in-tension interpolation. Numerische Mathematik, 54(2):179–192, November 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Skrzipek:1998:PEA

[Skr98] Michael-Ralf Skrzipek. Polynomial evaluation and associated polynomials. Numerische Mathematik, 79(4): 601–613, June 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079004/80790601.
htm;   http://link.springer.

                                        de/link/service/journals/
                                        00211/papers/8079004/80790601.
                                        pdf;   http://link.springer.
                                        de/link/service/journals/
                                        00211/papers/8079004/80790601.
                                        ps.gz.

Sakai:1986:SRS

[SL86] Manabu Sakai and M. C. López de Silanes. A simple rational spline and its application to monotonic interpolation to monotonic data. Numerische Mathematik, 50(2): 171–182, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sloan:1976:EAC

[Slo76] Ian H. Sloan. Error analysis for a class of degenerate-kernel methods. Numerische Mathematik, 25(3):231–238, September 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sloboda:1980:GCG

[Slo80] Fridrich Sloboda. A generalized conjugate gradient algorithm for minimization. Numerische Mathematik, 35(2): 223–230, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sloan:1981:AGQ

[Slo81] Ian H. Sloan. Analysis of general quadrature methods for integral equations of the second kind. Numerische Mathematik, 38(2):263–278, December 1981.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sloan:1988:QBA

[Slo88] Ian H. Sloan. A quadraturebased approach to improving the collocation method. Numerische Mathematik, 54(1):41– 56, October 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Smith:1965:QMB

[Smi65] Francis J. Smith. Quadrature methods based on the Clenshaw–Curtis method of integration. Numerische Mathematik, 7(5):406–411, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Smith:1967:CNM

[Smi67] R. A. Smith. The condition numbers of the matrix eigenvalue problem. Numerische Mathematik, 10(3):232–240, October 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Smith:1991:DDA

[Smi91] Barry F. Smith. A domain decomposition algorithm for elliptic problems in three dimensions. Numerische Mathematik, 60(2):219–234, November 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Smoch:1990:AEM

[Smo90] Martin Smoch. Anmerkungen zu einem Mehrgitterverfahren für lineare Komplementaritätsprobleme. (German) [Remarks on a multigrid method for linear complementarity problems]. Numerische Mathematik, 57(1):80–83, March 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stynes:1986:FEM

[SO86] Martin Stynes and Eugene O’Riordan. A finite element method for a singularly perturbed boundary value problem. Numerische Mathematik, 50(1): 1–15, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stynes:1987:UCD

[SO87] Martin Stynes and Eugene O’Riordan. L^{1} and L^{∞} uniform convergence of a difference scheme for a semilinear singular perturbation problem. Numerische Mathematik, 50(5): 519–531, March 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stynes:1989:UCD

[SO89] Martin Stynes and Eugene O’Riordan. Uniformly convergent difference schemes for singularly perturbed parabolic diffusion-convection problems without turning points. Numerische Mathematik, 55(5):

521–544, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Socolovsky:1988:NAF

[Soc88] E. A. Socolovsky. On the numerical approximation of finite speed diffusion problems. Numerische Mathematik, 53(1/2): 97–105, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Soderlind:1986:BNO

[Söd86] Gustaf Söderlind. Bounds on nonlinear operators in finitedimensional Banach spaces. Numerische Mathematik, 50(1):27– 44, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sonar:1992:EPS

[Son92] T. Sonar. Entropy production in second-order three-point schemes. Numerische Mathematik, 62(3):371–390, August 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Song:1993:SCT

[Son93] Yongzhong Song. Some comparison theorems for nonnegative splittings of matrices. Numerische Mathematik, 65(2): 245–252, June 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Soukup:1969:MFR

[Sou69] Jiří Soukup. A method for finding the roots of a polynomial.

Numerische Mathematik, 13(4): 349–353, August 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saint-Pierre:1972:CND

[SP72]  Patrick Saint-Pierre.  Sur

la convergence numérique de différentes méthodes de décomposition pour une class de problèmes elliptiques. (French) [On the numerical convergence of different methods of decomposition for a class of elliptic problems]. Numerische Mathematik, 19(5): 439–449, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sweldens:1994:AEE

[SP94] Wim Sweldens and Robert Piessens. Asymptotic error expansion of wavelet approximations of smooth functions. II. Numerische Mathematik, 68(3): 377–401, September 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4068003/
40680377.htm;   http://science.
springer.de/nmee/bibs/4068003/
40680377.htm.

Spath:1986:OLS

[Spä86] H. Späth. Orthogonal least squares fitting with linear manifolds. Numerische Mathematik, 48(4):441–445, April 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spann:1993:BEM

[Spa93] W. Spann. On the boundary element method for the Signorini problem of the Laplacian. Numerische Mathematik, 65(3): 337–356, July 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spann:1994:EEA

[Spa94]  W. Spann.  Error esti-

mates for the approximation of semicoercive variational inequalities. Numerische Mathematik, 69(1):103–116, November 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069001/40690103.
htm;   http://science.springer.
de/nmee/bibs/4069001/40690103.
htm.

Spellucci:1971:DPA

[Spe71] P. Spellucci. Double precision approximations to the elementary functions using Jacobifractions. Numerische Mathematik, 18(2):127–143, April 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spence:1975:CNM

[Spe75]  Alastair Spence.  On the

convergence of the Nyström method for the integral equation eigenvalue problem. Numerische Mathematik, 25(1):

57–66, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spence:1978:EBE

[Spe78] Alastair Spence. Error bounds and estimates for eigenvalues of integral equations. Numerische Mathematik, 29(2): 133–147, January 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spence:1979:EBE

[Spe79] Alastair Spence. Error bounds and estimates for eigenvalues of integral equations. II. Numerische Mathematik, 32(2): 139–146, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spijker:1966:CSS

[Spi66] M. N. Spijker. Convergence and stability of step-by-step methods for the numerical solution of initial-value problems. Numerische Mathematik, 8(2):161– 177, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spijker:1971:SEE

[Spi71] M. N. Spijker. On the structure of error estimates for finite-difference methods. Numerische Mathematik, 18(1):73– 100, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spijker:1976:PTE

[Spi76] M. N. Spijker. On the possibility of two-sided error bounds in the numerical solution of initial value problems. Numerische Mathematik, 26(3): 271–300, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spijker:1983:CNS

[Spi83] M. N. Spijker. Contractivity in the numerical solution of initial value problems. Numerische Mathematik, 42(3): 271–290, December 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spijker:1986:MBI

[Spi86] M. N. Spijker. Monotonicity and boundedness in implicit Runge– Kutta methods. Numerische Mathematik, 50(1):97–109, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sprekels:1976:EEE

[Spr76]  Jürgen Sprekels.  Existenz

und erste Einschließung positiver Lösungen bei superlinearen Randwertaufgaben zweiter Ordnung. (German) [Existence and first bounds of positive solutions of superlinear second-order boundary value problems]. Numerische Mathematik, 26(4): 421–428, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sprekels:1980:EBS

[Spr80] Jürgen Sprekels. Exact bounds for the solution branches of nonlinear eigenvalue problems. Numerische Mathematik, 34(1): 29–40, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Squier:1969:OSM

[Squ69] D. P. Squier. One-step methods for ordinary differential equations. Numerische Mathematik, 13(2):176–179, May 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sreedharan:1969:SOL

[Sre69] V. P. Sreedharan. Solutions of overdetermined linear equations which minimize error in an abstract norm. Numerische Mathematik, 13(2):146–151, May 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sreedharan:1971:LSA

[Sre71] V. P. Sreedharan. Least squares algorithms for finding solutions of overdetermined linear equations which minimize error in an abstract norm. Numerische Mathematik, 17(5):387–401, October 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

SriRanga:1994:AQR

[Sri94] A. Sri Ranga. Another quadrature rule of highest algebraic degree of precision. Numerische Mathematik, 68(2): 283–294, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068002/40680283.
htm;   http://science.springer.
de/nmee/bibs/4068002/40680283.
htm.

Schneider:1962:CTS

[SS62] Hans Schneider and W. Gilbert Strang. Comparison theorems for supremum norms. Numerische Mathematik, 4:15– 20, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shah:1965:ACF

[SS65] S. M. Shah and P. Suryanarayana. Approximation of continuous functions by power series with polynomial coefficients. Numerische Mathematik, 7(5):420–425, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schaefke:1966:VBC

[SS66] Friedrich W. Schaefke and Dieter Schmidt. Ein Verfahren zur Berechnung des charakteristischen Exponenten der Mathieuschen Differentialgleichung. III. (German) [A method for computation of the characteristic exponent of Mathieu’ differential equation. III]. Numerische Mathematik, 8(1):68–71, March 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saunders:1976:SNR

[SS76] B. David Saunders and Hans Schneider. A symmetric numerical range for matrices. Numerische Mathematik, 26(1):99– 105, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sloan:1978:PCR

[SS78] Ian H. Sloan and W. E. Smith. Product-integration with the Clenshaw–Curtis and related points-convergence properties. Numerische Mathematik, 30(4): 415–428, December 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schittkowski:1979:FMS

[SS79] K. Schittkowski and J. Stoer. A factorization method for the solution of constrained linear least squares problems allowing subsequent data changes. Numerische Mathematik, 31(4): 431–463, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sloan:1980:PIC

[SS80] Ian H. Sloan and William E. Smith. Product integration with the Clenshaw–Curtis points: implementation and error estimates. Numerische Mathematik, 34(4):387–401, December 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sanz-Serna:1984:CLM

[SS84] J. M. Sanz-Serna. Convergence of the Lambert-McLeod trajectory solver and of the CELF method. Numerische Mathematik, 45(2):173–182, November 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1988:CDB

[SS88] G. Schmidt and H. Strese. The convergence of a direct BEM for the plane mixed boundary value problem of the Laplacian. Numerische Mathematik, 54(2): 145–165, November 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1990:DAC

[SS90] Jochen W. Schmidt and Isa Scholz. A dual algorithm for convex-concave data smoothing by cubic C^{2}-splines. Numerische Mathematik, 57(4): 333–350, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sun:1995:AFO

[SS95] Guang Fu Sun and Martin Stynes. An almost fourth order uniformly convergent difference scheme for a semilinear singularly perturbed reaction-diffusion problem. Numerische Mathematik, 70(4): 487–500, June 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070004/50700487.
htm;   http://science.springer.
de/nmee/bibs/5070004/50700487.
htm.

Stenberg:1996:MFE

[SS96] Rolf Stenberg and Manil Suri. Mixed h-p finite element methods for problems in elasticity and Stokes flow. Numerische Mathematik, 72(3):367–389, January 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6072003/60720367.
htm;   http://science.springer.
de/nmee/bibs/6072003/60720367.
htm.

Sauter:1997:QGB

[SS97] Stefan A. Sauter and Christoph Schwab. Quadrature for hp-Galerkin BEM in R^{3}. Numerische Mathematik, 78(2): 211–258, December 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7078002/
70780211.htm;   http://science.
springer.de/nmee/bibs/7078002/
70780211.htm.

Sonar:1998:DGN

[SS98] Thomas Sonar and Endre Süli.

A dual graph-norm refinement indicator for finite volume approximations of the Euler equations. Numerische Mathematik, 78(4):619–658, February 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078004/70780619.
htm;   http://science.springer.
de/nmee/bibs/7078004/70780619.
htm.

Szymczak:1988:MDS

[SSBB88] W. G. Szymczak, J. M.

Solomon, A. E. Berger, and J. B. Bell. Multiple discrete solutions of the incompressible steady-state Navier–Stokes equations. Numerische Mathematik, 53(1/2):31–50, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sanz-Serna:1986:RSE

[SSS86] J. M. Sanz-Serna and M. N. Spijker. Regions of stability, equivalence theorems and the Courant-Friedrichs-Lewy condition. Numerische Mathematik, 49(2/3): 319–329, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schotzau:1999:MHF

[SSS99] Dominik Schötzau, Christoph Schwab, and Rolf Stenberg. Mixed hp-FEM on anisotropic meshes II: Hanging nodes and tensor products of boundary layer meshes. Numerische Mathematik, 83(4):667–697, October 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083004/90830667.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083004/90830667.pdf.

Saff:1976:EGC

[SSV76a] E. B. Saff, A. Schönhage, and R. S. Varga. Errata: “Geometric convergence to e^{−z} by rational functions with real poles”. Numerische Mathematik, 26(2): 226, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [SSV76b].

Saff:1976:GCRa

[SSV76b] E. B. Saff, A. Schönhage, and R. S. Varga. Geometric convergence to e^{−z} by rational functions with real poles. Numerische Mathematik, 25(3): 307–322, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See errata [SSV76a].

Sanz-Serna:1987:COR

[SSVH87] J. M. Sanz-Serna, J. G. Verwer, and W. H. Hundsdorfer. Convergence and order reduction of Runge–Kutta schemes applied to evolutionary problems in partial differential equations. Numerische Mathematik, 50(4):405–418, July 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schroder:1973:RDB

[ST73] Johann Schröder and Ulrich Trottenberg. Reduktionsverfahren für DIfferenzengleichungen bie Randwertaufgaben. I. (German) [Reduction method for difference equations for boundary value problems. I]. Numerische Mathematik, 22(1): 37–68, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Steidl:1989:ITM

[ST89] Gabriele Steidl and Manfred Tasche. Index transforms for multidimensional DFTs and convolutions. Numerische Mathematik, 56(5):513–528, November 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sikorski:1990:ANO

[ST90] K. Sikorski and G. M. Trojan. Asymptotic near optimality of the bisection method. Numerische Mathematik, 57(5): 421–433, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schumaker:1991:FSD

[ST91] Larry L. Schumaker and Cornelis Traas. Fitting scattered data on spherelike surfaces using tensor products of trigonometric and polynomial splines. Numerische Mathematik, 60(1): 133–144, November 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stallmann:1970:NSI

[Sta70] Friedmann W. Stallmann. Numerical solution of integral equations. Numerische Mathematik, 15(4):297–305, August 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Starius:1977:CMD

[Sta77a] Gøran Starius. Composite mesh difference methods for elliptic boundary value problems. Numerische Mathematik, 28(2): 243–258, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Starius:1977:COC

[Sta77b] Göran Starius. Constructing orthogonal curvilinear meshes by solving initial value problems. Numerische Mathematik, 28(1): 25–48, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Starius:1980:CMD

[Sta80] Göran Starius. On composite mesh difference methods for hyperbolic differential equations. Numerische Mathematik, 35(3): 241–255, October 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stachurski:1981:SCS

[Sta81] Andrzej Stachurski. Superlinear convergence of symmetric Huang’s class of methods. Numerische Mathematik, 38(2): 209–218, December 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Starke:1997:FVA

[Sta97] Gerhard Starke. Field-of-values analysis of preconditioned iterative methods for nonsymmetric elliptic problems. Numerische Mathematik, 78(1):103–117, November 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078001/70780103.
htm;   http://science.springer.
de/nmee/bibs/7078001/70780103.
htm.

Stetter:1961:CCF

[Ste61] Hans J. Stetter. On the convergence of characteristic finitedifference methods of high accuracy for quasi-linear hyperbolic equations. Numerische Mathematik, 3(1):321–344, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
321.

Stetter:1963:MBS

[Ste63] Hans J. Stetter. Maximum bounds for the solutions of initial value problems for partial difference equations. Numerische Mathematik, 5(1):399– 424, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stetter:1965:AEE

[Ste65] Hans J. Stetter. Asymptotic expansions for the error of discretization algorithms for nonlinear functional equations. Numerische Mathematik, 7(1):18– 31, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenger:1966:BEG

[Ste66a] Frank Stenger. Bounds on the error of Gauss-type quadratures. Numerische Mathematik, 8(2): 150–160, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenger:1966:EBE

[Ste66b] Frank Stenger. Error bounds for the evaluation of integrals by repeated Gauss-type formulae. Numerische Mathematik, 9(3): 200–213, December 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stetter:1966:NAF

[Ste66c]  Hans J. Stetter.  Numeri-

cal approximation of Fouriertransforms. Numerische Mathematik, 8(3):235–249, May 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1967:GTF

[Ste67] G. W. Stewart III. A generalization of a theorem of Fan on Gershgorin disks. Numerische Mathematik, 10:162, 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://zbmath.
org/?q=an:0158.03504. See

original paper by Fan at https:

//doi.org/10.1215/S0012-7094-
58-02538-9, https://zbmath.

org/?qan:0081.25202=.

Stetter:1968:INM

[Ste68] Hans J. Stetter. Instability and non-monotonicity phenomena in discretizations to boundary value problems. Numerische Mathematik, 12(2):139–145, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1969:AOI

[Ste69a] G. W. Stewart. Accelerating the orthogonal iteration for the eigenvectors of a Hermitian matrix. Numerische Mathematik, 13(4):362–376, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1969:SIF

[Ste69b] G. W. Stewart. Some iterations for factoring a polynomial. Numerische Mathematik, 13(5): 458–470, October 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1970:SIF

[Ste70] G. W. Stewart. On Samelson’s iteration for factoring polynomials. Numerische Mathematik, 15 (4):306–314, August 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1971:COA

[Ste71] G. W. Stewart. On a companion operator for analytic functions. Numerische Mathematik, 18(1): 26–43, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Steinberg:1972:NSV

[Ste72] J. Steinberg. Numerical solution of Volterra integral equation. Numerische Mathematik, 19(3):212–217, June 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1973:CDM

[Ste73a] G. W. Stewart. Conjugate direction methods for solving systems of linear equations. Numerische Mathematik, 21(4):285–297, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1973:SIF

[Ste73b] G. W. Stewart. Some iterations for factoring a polynomial. II: a generalization of the secant method. Numerische Mathematik, 22(1):33–36, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenger:1975:CTD

[Ste75a] Frank Stenger. Computing the topological degree of a mapping in R^{n}. Numerische Mathematik, 25(1):23–38, March 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1975:CMC

[Ste75b] G. W. Stewart. The convergence of the method of conjugate gradients at isolated extreme points of the spectrum. Numerische Mathematik, 24(2): 85–93, April 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1976:ESP

[Ste76a] G. W. Stewart. The economical storage of plane rotations. Numerische Mathematik, 25(2): 137–138, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1976:SIC

[Ste76b] G. W. Stewart. Simultaneous iteration for computing invariant subspaces of non-Hermitian matrices. Numerische Mathematik, 25(2):123–136, June 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenger:1978:OCM

[Ste78a] Frank Stenger. Optimal convergence of minimum norm approximations in H_{p}. Numerische Mathematik, 29(4): 345–362, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stetter:1978:DCP

[Ste78b]  Hans J. Stetter.  The de-

fect correction principle and discretization methods. Numerische Mathematik, 29(4): 425–443, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1982:CDP

[Ste82] G. W. Stewart. Computing the CS decomposition of a partitioned orthonormal matrix. Numerische Mathematik, 40(3): 297–306, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1984:IPN

[Ste84]  G. W. Stewart.  On the

invariance of perturbed null vectors under column scaling. Numerische Mathematik, 44(1): 61–65, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenberg:1986:COM

[Ste86] Rolf Stenberg. On the construction of optimal mixed finite element methods for the linear elasticity problem. Numerische Mathematik, 48(4): 447–462, April 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenberg:1988:FMF

[Ste88] Rolf Stenberg. A family of mixed finite elements for the elasticity problem. Numerische Mathematik, 53(5):513–538, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stenberg:1990:SNF

[Ste90a]  Rolf Stenberg.  Some new

families of finite elements for the Stokes equations. Numerische Mathematik, 56(8): 827–838, February 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1990:HAM

[Ste90b] David Stewart. A high accuracy method for solving ODEs with discontinuous right-hand side. Numerische Mathematik, 58(3): 299–328, October 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stevenson:1993:RMG

[Ste93a] Rob Stevenson. Robustness of multi-grid applied to anisotropic equations on convex domains and on domains with re-entrant corners. Numerische Mathematik, 66(3):373–398, December 1993. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1993:PMC

[Ste93b]  G. W. Stewart.  On the

perturbation of Markov chains with nearly transient states. Numerische Mathematik, 65(1): 135–141, May 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stevenson:1994:MIS

[Ste94a]  Rob Stevenson.  Modified

ILU as a smoother. Numerische Mathematik, 68(2): 295–309, July 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/4068002/40680295.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4068002/40680295.
                                        htm.

Stewart:1994:CMI

[Ste94b] G. W. Stewart. On the convergence of multipoint iterations. Numerische Mathematik, 68(1): 143–147, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/4068001/40680143.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/4068001/40680143.
                                        htm.

Stengle:1995:EAR

[Ste95] Gilbert Stengle. Error analysis of a randomized numerical method. Numerische Mathematik, 70(1):119–128, March 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer.de/link/service/
journals/00211/bibs/5070001/
50700119.htm;   http://science.
springer.de/nmee/bibs/5070001/
50700119.htm.

Stevenson:1997:RHB

[Ste97] Rob Stevenson. A robust hierarchical basis preconditioner on general meshes. Numerische Mathematik, 78(2):269–303, December 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078002/70780269.
htm;   http://science.springer.
de/nmee/bibs/7078002/70780269.
htm.

Stevenson:1998:STW

[Ste98] Rob Stevenson. Stable threepoint wavelet bases on general meshes. Numerische Mathematik, 80(1):131–158, July 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8080001/
80800131.htm;   http://link.
springer.de/link/service/
journals/00211/papers/8080001/
80800131.pdf;   http://link.
springer.de/link/service/
journals/00211/papers/8080001/
80800131.ps.gz.

Stewart:1999:FAE

[Ste99] G. W. Stewart. Four algorithms for the efficient computation of truncated pivoted QR approximations to a sparse matrix. Numerische Mathematik, 83(2):313–323, August 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9083002/90830313.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9083002/90830313.pdf.

Stiefel:1959:DLT

[Sti59]  E. Stiefel.  Über diskrete

und lineare Tschebyscheff-Approximationen. (German) [On discrete and linear Chebyshev approximations]. Numerische Mathematik, 1:1–28, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stiefel:1960:NJE

[Sti60] E. Stiefel. Note on Jordan elimination, linear programming and Tchebyscheff approximation. Numerische Mathematik, 2(1):1–17, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Still:1988:CBE

[Sti88] Georg Still. Computable bounds for eigenvalues and eigenfunctions of elliptic differential operators. Numerische Mathematik, 54(2):201–223, November 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1961:ZAI

[Sto61] Josef Stoer. Über zwei Algorithmen zur Interpolation mit rationalen Funktionen. (German) [On two algorithms for interpolation with rational functions]. Numerische Mathematik, 3(1): 285–304, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
285.

Stone:1962:BPR

[Sto62] Betty Jane Stone. Best possible ratios of certain matrix norms. Numerische Mathematik, 4:114– 116, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1963:DNP

[Sto63] J. Stoer. Duality in nonlinear programming and the minimax theorem. Numerische Mathematik, 5(1):371–379, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1964:CLU

[Sto64a] J. Stoer. On the characterization of least upper bound norms in matrix space. Numerische Mathematik, 6(1):302– 314, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1964:UDN

[Sto64b] J. Stoer. Über einen Dualitätssatz der nichtlinearen Programmierung. (German) [On a duality theorem of nonlinear programming]. Numerische Mathematik, 6(1):55– 58, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoss:1967:VBC

[Sto67] Hanns-Jörg Stoss. Ein Verfahren zur Berechnung des charakteristischen Exponenten der Differentialgleichung y′′ + (λ + 2λ1cosχ + 2λ2cos2χ)y = 0. (German) [A procedure for the computation of the characteristic exponent of the differential equation y′′ + (λ + 2λ1cosχ + 2λ2cos2χ)y = 0]. Numerische Mathematik, 10(5): 423–436, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1968:LBM

[Sto68] Josef Stoer. Lower bounds of matrices. Numerische Mathematik, 12(2):146–158, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoss:1970:MED

[Sto70] Hanns-Jörg Stoß. Monotonie-Eigenschaften bei Differentialungleichungen über nichtkompaktem Grundbereich. (German) [Monotone eigenvalues of differential equations over noncompact fields]. Numerische Mathematik, 15(1):61–73, May 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1977:RBQ

[Sto77] J. Stoer. On the relation between quadratic termination and convergence properties of minimization algorithms. I. theory. Numerische Mathematik, 28(3):343–366, September 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stokes:1983:ESW

[Sto83] A. N. Stokes. Efficient stable ways to calculate continued fraction coefficients from some series. Numerische Mathematik, 42(2):237–245, November 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1984:CMG

[Sto84] Josef Stoer. The convergence of matrices generated by rank-2 methods from the restricted βclass of Broyden. Numerische Mathematik, 44(1):37–52, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoffer:1993:GLM

[Sto93] Daniel Stoffer. General linear methods: connection to one step methods and invariant curves. Numerische Mathematik, 64(3): 395–408, March 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoffer:1997:QBS

[Sto97] Daniel Stoffer. On the qualitative behaviour of symplectic integrators. Part I: Perturbed linear systems. Numerische Mathematik, 77(4):535–547, October 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/7077004/70770535.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/7077004/70770535.
                                        htm.

Struble:1963:OPV

[Str63] George W. Struble. Orthogonal polynomials: Variablesigned weight functions. Numerische Mathematik, 5(1):88– 94, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strang:1964:APD

[Str64] G. Strang. Accurate partial difference methods. II: Non-linear problems. Numerische Mathematik, 6(1):37–46, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stroud:1966:SAI

[Str66] A. H. Stroud. Some approximative integration formulas of degree 3 for an n-dimensional simplex. Numerische Mathematik, 9 (1):38–45, November 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stroud:1967:SFD

[Str67] A. H. Stroud. Some fifth degree integration formulas for symmetric regions II. Numerische Mathematik, 9(5):460– 468, April 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stroud:1968:SSD

[Str68]  A. H. Stroud.  Some sev-

enth degree integration formulas for the surface of an n-sphere. Numerische Mathematik, 11(3): 273–276, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strassen:1969:GEO

[Str69] Volker Strassen. Gaussian elimination is not optimal. Numerische Mathematik, 13(4): 354–356, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). The author shows that matrix multiplication can be done with fewer than 4.7 × n^{log_{2} 7} ≈ 4.7 × n^{2.807} arithmetic operations, and that Gaussian elimination and matrix inversion can be done with O(n^{log_{2} 7} arithmetic operations.

Strang:1972:AFE

[Str72] Gilbert Strang. Approximation in the finite element method. Numerische Mathematik, 19(1): 81–98, February 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strassen:1973:BEF

[Str73a] Volker Strassen. Die Berechnungskomplexität von elementarsymmetrischen Funktionen und von Interpolationskoeffizienten. (German) [The computational complexity of elementary symmetric functions and interpolation coefficients]. Numerische Mathematik, 20(3): 238–251, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strom:1973:MEB

[Str73b] Torsten Ström. Monotonicity and error bounds in schemes of Romberg type for the computation of f^{(n)}(0) by central differences and extrapolation. Numerische Mathematik, 21(1):14– 21, February 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schroder:1976:RDB

[STR76] Johann Schröder, Ulrich Trottenberg, and Heinz Reutersberg. Reduktionsverfahren für DIfferenzengleichungen bie Randwertaufgaben. II. (German) [Solving difference equations for boundary value problems by reduction methods. II]. Numerische Mathematik, 26(4):

429–459, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strauss:1982:UBO

[Str82] Hans Strauß. Unicity of best one-sided L_{1}-approximations. Numerische Mathematik, 40(2): 229–243, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strauss:1984:MQF

[Str84] H. Strauss. Monotonicity of quadrature formulae of Gauss type and comparison theorems for monosplines. Numerische Mathematik, 44(3): 337–347, September 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strack:1986:SAI

[Str86] Klaus-Günther G. Strack. Stability and accuracy for implicit semidiscretizations of hyperbolic problems. Numerische Mathematik, 48(2): 221–237, February 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stummel:1976:WSW

[Stu76] Friedrich Stummel. Weak stability and weak discrete convergence of continuous mappings. Numerische Mathematik, 26(3): 301–315, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stummel:1985:FEAa

[Stu85a] Friedrich Stummel. Forward error analysis of Gaussian elimination. I. error and residual estimates. Numerische Mathematik, 46(3):365–395, June 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stummel:1985:FEAb

[Stu85b] Friedrich Stummel. Forward error analysis of Gaussian elimination. II. stability theorems. Numerische Mathematik, 46(3): 397–415, June 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Studniarski:1989:ACO

[Stu89] Marcin Studniarski. An algorithm for calculating one subgradient of a convex function of two variables. Numerische Mathematik, 55(6):685–693, August 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stynes:1979:SST

[Sty79] Martin Stynes. A simplification of Stenger’s topological degree formula. Numerische Mathematik, 33(2):147–155, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stynes:1981:CSR

[Sty81] Martin Stynes. On the construction of sufficient refinements for computation of topological degree. Numerische Mathematik, 37(3):453–462, July 1981. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sakai:1989:CSE

[SU89] Manabu Sakai and Riaz A. Usmani. A class of simple exponential B-splines and their application to numerical solution to singular perturbation problems. Numerische Mathematik, 55(5): 493–500, August 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sugihara:1997:ODE

[Sug97] Masaaki Sugihara. Optimality of the double exponential formula — functional analysis approach. Numerische Mathematik, 75(3):379–395, January 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7075003/70750379.
htm;   http://science.springer.
de/nmee/bibs/7075003/70750379.
htm.

Suli:1988:CNS

[Sül88]  Endre Süli.  Convergence

and nonlinear stability of the Lagrange–Galerkin method for the Navier–Stokes equations.

Numerische Mathematik, 53(4): 459–483, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sunouchi:1968:PTD

[Sun68] Haruo Sunouchi. Perturbation theory of difference schemes. Numerische Mathematik, 12(5): 454–458, December 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sun:1995:HOD

[Sun95] W. Sun. A high order direct method for solving Poisson’s equation in a disc. Numerische Mathematik, 70(4): 501–506, June 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070004/50700501.
htm;   http://science.springer.
de/nmee/bibs/5070004/50700501.
htm.

Sun:1998:FAH

[Sun98]  W. Sun.  Fast algorithms

for high-order spline collocation systems. Numerische Mathematik, 81(1):143–160, November 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8081001/80810143.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8081001/80810143.
pdf.

Suschowk:1964:DAG

[Sus64] D. Suschowk. Difference analogues of Green’s identities for grids in R^{n}. Numerische Mathematik, 6(1):200–210, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schultz:1967:S

[SV67] M. H. Schultz and R. S. Varga.

L-splines. Numerische Mathematik, 10(4):345–369, November 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmidt:1970:LMN

[SV70] Jochen W. Schmidt and Klaus Vetters. Ableitsungfreie Verfahren für nichtlineare Optimierungsprobleme. (German) [Lossless method for nonlinear optimization problem]. Numerische Mathematik, 15(4): 263–282, August 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Saff:1975:ZPP

[SV75] E. B. Saff and R. S. Varga. On the zeros and poles of Padé approximants to e^{z}. Numerische Mathematik, 25(1):1–14, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [SV76a].

Saff:1976:EZP

[SV76a] E. B. Saff and R. S. Varga.

Erratum: “On the zeros and poles of Padé approximants to e^{z}”. Numerische Mathematik, 26(2):226, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [SV75].

Saff:1976:STC

[SV76b] E. B. Saff and R. S. Varga. On the sharpness of theorems concerning zero-free regions for certain sequences of polynomials. Numerische Mathematik, 26(4): 345–354, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [SV77].

Saff:1977:EST

[SV77] E. B. Saff and R. S. Varga. Erratum: “On the sharpness of theorems concerning zerofree regions for certain sequences of polynomials” (Numer. Math. 26 (1976), no. 4, 345–354). Numerische Mathematik, 28(1): 119, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [SV76b].

Saff:1978:ZPP

[SV78] E. B. Saff and R. S. Varga. On the zeros and poles of Padé approximants to e^{z}. III. Numerische Mathematik, 30(3): 241–266, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sprekels:1979:PIF

[SV79] Jürgen Sprekels and Heinrich Voss. Pointwise inclusions of fixed points by finitedimensional iteration schemes. Numerische Mathematik, 32(4): 381–392, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

SacchiLandriani:1990:MSC

[SV90] G. Sacchi Landriani and H. Vandeven. A multidomain spectral collocation method for the Stokes problem. Numerische Mathematik, 58(4): 441–464, November 1990. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Starke:1993:HAF

[SV93] Gerhard Starke and Richard S.

Varga. A hybrid Arnoldi-Faber iterative method for nonsymmetric systems of linear equations. Numerische Mathematik, 64(2):213–240, February 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Svecoa:1970:BSC

[Šve70] H. Švecoá. On the Bauer’s scaled condition number of matrices arising from approximate conformal mapping. Numerische Mathematik, 14(5): 495–507, April 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=14&issue=5&spage=
495.

Saff:1976:GCRb

[SVN76] E. B. Saff, R. S. Varga, and W.-C. Ni. Geometric convergence of rational approximations to e^{−z} in infinite sectors. Numerische Mathematik, 26(2): 211–225, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stoer:1962:TDM

[SW62] J. Stoer and C. Witzgall. Transformations by diagonal matrices in a normed space. Numerische Mathematik, 4:158–171, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Symm:1980:REB

[SW80] H. J. Symm and J. H. Wilkinson. Realistic error bounds for a simple eigenvalue and its associated eigenvector. Numerische Mathematik, 35(2): 113–126, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Strehmel:1984:PAR

[SW84] K. Strehmel and R. Weiner. Partitioned adaptive Runge–Kutta methods and their stability. Numerische Mathematik, 45(2): 283–300, November 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spath:1987:OLA

[SW87] H. Späth and G. A. Watson. On orthogonal linear ℓ_{1} approximation. Numerische Mathematik, 51(5):531–543, October 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schwab:1992:NCS

[SW92] C. Schwab and W. L. Wendland. On numerical cubatures of singular surface integrals in boundary element methods. Numerische Mathematik, 62(3):343–369, August 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sivakumar:1993:LSF

[SW93] N. Sivakumar and J. D. Ward.

On the least squares fit by radial functions to multidimensional scattered data. Numerische Mathematik, 65(2): 219–243, June 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Shaw:1996:DGM

[SW96] S. Shaw and J. R. Whiteman.

Discontinuous Galerkin method with a posteriori L_{p}(0, t_{i}) error estimate for second-kind Volterra problems. Numerische Mathematik, 74(3):361–383, September 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/

00211/bibs/6074003/60740361.
htm;   http://science.springer.
de/nmee/bibs/6074003/60740361.
htm.

Sloan:1998:QME

[SW98] Ian H. Sloan and W. L. Wendland. Qualocation methods for elliptic boundary integral equations. Numerische Mathematik, 79(3):451–483, May 1998. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/8079003/
80790451.htm;   http://science.
springer.de/nmee/bibs/8079003/
80790451.htm.

Sloan:1999:SQM

[SW99a] Ian H. Sloan and Wolfgang L. Wendland. Spline qualocation methods for variable-coefficient elliptic equations on curves. Numerische Mathematik, 83 (3):497–533, September 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083003/90830497.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083003/90830497.pdf.

Stoer:1999:API

[SW99b] Josef Stoer and Martin Wechs. On the analyticity properties of infeasible-interior-point paths for monotone linear complementarity problems. Numerische Mathematik, 81(4):631–645, February 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer-
ny.com/link/service/journals/
00211/bibs/9081004/90810631.
htm;   http://link.springer-
ny.com/link/service/journals/
00211/papers/9081004/90810631.
pdf.

Strehmel:1991:ORR

[SWB91] Karl Strehmel, Rüdiger Weiner, and Mario Büttner. Order results for Rosenbrock type methods on classes of stiff equations. Numerische Mathematik, 59(7): 723–737, October 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sweet:1984:FTO

[Swe84] D. R. Sweet. Fast Toeplitz orthogonalization. Numerische Mathematik, 43(1):1–21, January 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Sweet:1991:FBT

[Swe91]  D. R. Sweet.  Fast block

Toeplitz orthogonalization. Numerische Mathematik, 58(6): 613–629, January 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Symm:1966:IEM

[Sym66] George T. Symm. An integral equation method in conformal mapping. Numerische Mathematik, 9(3):250–258, December 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Symm:1967:NME

[Sym67] George T. Symm. Numerical mapping of exterior domains. Numerische Mathematik, 10(5): 437–445, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Symm:1969:CMD

[Sym69] George T. Symm. Conformal mapping of doubly-connected domains. Numerische Mathematik, 13(5):448–457, October 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Salzer:1961:SEBb

[SZ61] Herbert E. Salzer and R. Zurmühl. Some extensions of Bairstow’s method. Numerische Mathematik, 3(1):320, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
320.

Stewart:1991:EGM

[SZ91a] G. W. Stewart and G. Zhang. Eigenvalues of graded matrices and the condition numbers of a multiple eigenvalue. Numerische Mathematik, 58(7):

703–712, February 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Stewart:1991:DMS

[SZ91b] G. W. Stewart and G. Zhang.

On a direct method for the solution of nearly uncoupled Markov chains. Numerische Mathematik, 59(1):1–11, April 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tanabe:1971:PMS

[Tan71]  Kunio Tanabe.  Projection

method for solving a singular system of linear equations and its applications. Numerische Mathematik, 17(3): 203–214, June 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tanabe:1974:CLS

[Tan74] Kunio Tanabe. Characterization of linear stationary iterative processes for solving a singular system of linear equations. Numerische Mathematik, 22(5): 349–359, October 1974. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tang:1992:SNS

[Tan92]  Tao Tang.  Superconver-

gence of numerical solutions to weakly singular Volterra integro-differential equations. Numerische Mathematik, 61(3): 373–382, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tao:1984:ACD

[Tao84] Pham Dinh Tao. Algorithmes de calcul du maximum des formes quadratiques sur la boule unité de la norme du maximum. (French) [Some methods for computing the maximum of quadratic forms on the unit ball of the maximum norm]. Numerische Mathematik, 45(3): 377–401, December 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Taubert:1976:DST

[Tau76] K. Taubert. Differenzverfahren für Schwingungen mit trockener und zäher Reibung und für Regelungssysteme. (German) [A finite difference method in forced vibrations with combined viscous and Coulomb damping and in feedback control systems]. Numerische Mathematik, 26(4):379–395, December 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Taylor:1969:AFH

[Tay69a] G. D. Taylor. Approximation by functions having restricted ranges: Equality case. Numerische Mathematik, 14(1):71– 78, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Taylor:1969:GSR

[Tay69b] P. J. Taylor. A generalisation of systematic relaxation methods for consistently ordered matrices. Numerische Mathematik, 13 (5):377–395, October 1969. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tee:1964:ACM

[Tee64] G. J. Tee. An application of pcyclic matrices, for solving periodic parabolic problems. Numerische Mathematik, 6(1):142– 159, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Temme:1983:NCC

[Tem83] N. M. Temme. The numerical computation of the confluent hypergeometric function U(a, b, z). Numerische Mathematik, 41(1):63–82, April 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Thatcher:1976:UIG

[Tha76] R. W. Thatcher. The use of infinite grid refinements at singularities in the solution of Laplace’s equation. Numerische Mathematik, 25(2):163–178, June 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Thakur:1990:DAO

[Tha90] Lakshman S. Thakur. A direct algorithm for optimal quadratic splines. Numerische Mathematik, 57(4):313–332, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Thomee:1971:CNP

[Tho71] Vidar Thomée. Convergence near plane boundaries of some elliptic difference schemes. Numerische Mathematik, 17(1):45– 53, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Thomas:1975:ASO

[Tho75] K. S. Thomas. On the approximate solution of operator equations. Numerische Mathematik, 23(3):231–239, June 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Thomann:1990:RSF

[Tho90]  Jean Thomann.  Resomma-

tion des series formelles. solutions d’équations différentielles linéaires ordinaires du second ordre dans le champ complexe au voisinage de singularités irrégulières. (French) [Resummation of formal power series. solutions of second-order linear complex ordinary differential equations near irregular singularities]. Numerische Mathematik, 58(5):503–535, December 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsuda:1967:MKP

[TIK67] Takao Tsuda, Kozo Ichida, and Takeshi Kiyono. Monte Carlo path-integral calculations for two-point boundary-value problems. Numerische Mathematik, 10(2):110–116, July 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Timlake:1965:EP

[Tim65] W. P. Timlake. An extension of the Σ^{2} procedure. Numerische Mathematik, 7(3):261–268, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tissier:1972:IPV

[Tis72] Gilles Tissier. Interpolation à plusieurs variables sur la sphère. (French) [Interpolation with several variables on the sphere]. Numerische Mathematik, 19(2):136–145, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsuda:1964:AMC

[TK64] T. Tsuda and T. Kiyono. Application of the Monte Carlo method to systems of nonlinear algebraic equations. Numerische Mathematik, 6(1):59– 67, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tobiska:1991:MSD

[TL91] Lutz Tobiska and Gert Lube.

A modified streamline diffusion method for solving the stationary Navier–Stokes equation. Numerische Mathematik, 59(1): 13–29, April 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Takahasi:1973:QFO

[TM73] Hidetosi Takahasi and Masatake Mori. Quadrature formulas obtained by variable transformation. Numerische Mathematik, 21(3):206–219, June 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

terMaten:1986:CAH

[tMS86] E. Jan W. ter Maten and Gerard L. G. Sleijpen. A convergence analysis of hopscotch methods for fourth order parabolic equations. Numerische Mathematik, 49(2/3):275–290, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Todd:1965:SIG

[Tod65] John Todd. On smallest isolated Gerschgorin disks for eigenvalues. Numerische Mathematik, 7 (2):170–175, April 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Todd:1988:MLM

[Tod88] John Todd. The many limits of mixed means. II. the Carlson sequences. Numerische Mathematik, 54(1):1–18, October 1988. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Todd:1990:WMP

[Tod90]  John Todd.  The Weier-

strass mean. I. the periods of ℘(z|e_{1}, e_{2}, e_{3}). Numerische Mathematik, 57(8):737–746, August 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tomarelli:1984:RTO

[Tom84]  Franco Tomarelli.  Regular-

ity theorems and optimal error estimates for linear parabolic Cauchy problems. Numerische Mathematik, 45(1):23–50, October 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tornig:1963:DRN

[Tör63] Willi Törnig. Über Differenzverfahren in Rechteckgittern zur numerischen Lösung quasilinearer hyperbolischer Differentialgleichungen. (German) [On the difference method in rectangular grid for the numerical solution of quasilinear hyperbolic differential equations]. Numerische Mathematik, 5(1):353– 370, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Torelli:1991:SCG

[Tor91] Lucio Torelli. A sufficient condition for GPN-stability for delay differential equations. Numerische Mathematik, 59(3): 311–320, June 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Torrens:1997:EAP

[Tor97] Juan José Torrens. Sur l’erreur d’approximation par éléments finis en utilisant la méthode des plaquettes splines. (French) [On the finite element approximation error using the “plaquette splines” method]. Numerische Mathematik, 76(1): 69–85, March 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076001/70760069.
htm;   http://science.springer.
de/nmee/bibs/7076001/70760069.
htm.

Tourin:1992:CTP

[Tou92] Agnès Tourin. A comparison theorem for a piecewise Lipschitz continuous Hamiltonian and application to shape-from-shading problems. Numerische Mathematik, 62(1):75–85, June 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

teRiele:1991:NLB

[tR91]  H. J. J. te Riele.  A

new lower bound for the de Bruijn-Newman constant. Numerische Mathematik, 58(6): 661–667, January 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Treuenfels:1962:SLR

[Tre62] P. Treuenfels. Schranken für Lösungen von Randwertaufgaben. (German) [Bounds for solution of boundary value problems]. Numerische Mathematik, 4:21–23, December 1962. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Trefethen:1981:RCA

[Tre81] Lloyd N. Trefethen. Rational Chebyshev approximation on the unit disk. Numerische Mathematik, 37(2):297–320, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Troch:1971:ISR

[Tro71] Inge Troch. Über die Stabilität der Ruhelage bei Differentialgleichungen mit quasihomogenen rechten Seiten. (German) [On the steady-state stability of differential equations with quasi-homogeneous right sides]. Numerische Mathematik, 18(3):204–212, June 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Trottenberg:1976:LLG

[Tro76] Ulrich Trottenberg. Lösung linearer gewönlicher Randwertaufgaben 4. Ordnung mit Hilfe von Aufspaltungen. (German) [Solving linear ordinary boundary value problems of the fourth order by means of splittings].

Numerische Mathematik, 25(3): 297–306, September 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Trojan:1986:OTS

[Tro86] G. M. Trojan. Optimal transformations for scalar sequences. Numerische Mathematik, 48(1): 21–32, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Trouve:1990:CME

[Tro90] P. Trouvé. Sur la convergence des méthodes d’éléments finis nonconformes pour des problèmes linéaires de coques minces. (French) [On the convergence of nonconformal finiteelement methods for linear thin-shell problems]. Numerische Mathematik, 57(5): 481–524, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tourigny:1991:FEM

[TS91] Y. Tourigny and E. Süli. The finite element method with nodes moving along the characteristics for convection-diffusion equations. Numerische Mathematik, 59(4):399–412, July 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tong:1998:IMS

[TS98] Zhanye Tong and Ahmed Sameh. On an iterative method for saddle point problems. Numerische Mathematik, 79(4): 643–646, June 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8079004/80790643.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790643.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8079004/80790643.
ps.gz.

Tsao:1977:NIE

[Tsa77] Nai-Kuan Tsao. Newton interpolation is efficient for approximation of linear functionals. Numerische Mathematik, 29 (1):115–122, March 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsao:1979:EAS

[Tsa79] Nai-Kuan Tsao. Error analysis of splitting algorithms for polynomials. Numerische Mathematik, 32(4):409–421, December 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsuda:1972:NDF

[Tsu72] Takao Tsuda. Numerical differentiation of functions of very many variables. Numerische Mathematik, 18(4):327–335, August 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsuda:1973:NIF

[Tsu73] Takao Tsuda. Numerical integration of functions of very many variables. Numerische Mathematik, 20(5):377–391, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsuda:1975:NIF

[Tsu75] Takao Tsuda. Nonlinear interpolation of functions of very many variables. Numerische Mathematik, 24(5):395–414, October 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsuchiya:1988:EPR

[Tsu88] Takuya Tsuchiya. Enlargement procedure for resolution of singularities at simple singular solutions of nonlinear equations. Numerische Mathematik, 52(4): 401–411, April 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tsuchiya:1999:AKT

[Tsu99] Takuya Tsuchiya. An application of the Kantorovich theorem to nonlinear finite element analysis. Numerische Mathematik, 84(1):121–141, November 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9084001/90840121.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9084001/90840121.pdf.

Thiran:1987:SCA

[TT87] J. P. Thiran and S. Thiry.

Strict Chebyshev approximation for general systems of linear equations. Numerische Mathematik, 51(6):701–725, November 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Toh:1994:PPP

[TT94] Kim-Chuan Toh and Lloyd N.

Trefethen. Pseudozeros of polynomials and pseudospectra of companion matrices. Numerische Mathematik, 68(3): 403–425, September 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4068003/
40680403.htm;   http://science.
springer.de/nmee/bibs/4068003/
40680403.htm.

Turkel:1978:EMD

[Tur78]  Eli Turkel.  Extrapolation

methods for dynamic partial differential equations. Numerische Mathematik, 29(3): 269–285, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tuomela:1998:ADD

[TV98] Jukka Tuomela and Olivier Vacus. Algebraic derivation of discrete absorbing boundary conditions for the wave equation. Numerische Mathematik, 80(3):419–459, September 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080003/80800419.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800419.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800419.
ps.gz.

Thomee:1968:EDE

[TW68] Vidar Thomée and Bertil Westergren. Elliptic difference equations and interior regularity. Numerische Mathematik, 11(3): 196–210, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Thron:1980:ACL

[TW80] W. J. Thron and H. Waadeland. Accelerating convergence of limit periodic continued fractions K(a_{n}/1). Numerische Mathematik, 34(2): 155–170, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Thomee:1983:MNS

[TW83] Vidar Thomée and Lars B. Wahlbin. Maximum-norm stability and error estimates in Galerkin methods for parabolic equations in one space variable. Numerische Mathematik, 41(3): 345–371, July 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tyrtyshnikov:1994:HBH

[Tyr94] Evgenij E. Tyrtyshnikov. How bad are Hankel matrices? Numerische Mathematik, 67(2): 261–269, March 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067002/40670261.
htm;   http://science.springer.
de/nmee/bibs/4067002/40670261.
htm.

Uhlmann:1962:UFD

[Uhl62] W. Uhlmann. Über den Fehler bei Differenzenverfahren für die Poissonsche Differentialgleichung. (German) [On the bound of difference method for Poisson differential equation]. Numerische Mathematik, 4:226– 237, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Uhlig:1992:CPW

[Uhl92] Frank Uhlig. Are the coefficients of a polynomial well-conditioned functions of its roots? Numerische Mathematik, 61(3): 383–393, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Uhlig:1997:ABT

[Uhl97] Frank Uhlig. The DQR algorithm, basic theory, convergence, and conditional stability. Numerische Mathematik, 76(4): 515–553, June 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076004/70760515.
htm;   http://science.springer.
de/nmee/bibs/7076004/70760515.
htm.

Urabe:1967:NSM

[Ura67] Minoru Urabe. Numerical solution of multi-point boundary value problems in Chebyshev series. theory of the method. Numerische Mathematik, 9(4):341– 366, February 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Urabe:1970:IOS

[Ura70] Minoru Urabe. An implicit onestep method of high-order accuracy for the numerical integration of ordinary differential equations. Numerische Mathematik, 15(2):151–164, June 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Utreras:1983:NSF

[Utr83]  Florencio Utreras.  Natural

spline functions, their associated eigenvalue problem. Numerische Mathematik, 42(1): 107–117, September 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Utreras:1985:SND

[Utr85] Florencio I. Utreras. Smoothing noisy data under monotonicity constraint: existence, characterization and convergence rates.

Numerische Mathematik, 47(4): 611–625, December 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VanderCruyssen:1979:ROA

[Van79] P. Van der Cruyssen. A reformulation of Olver’s algorithm for the numerical solution of second-order linear difference equations. Numerische Mathematik, 32(2):159–166, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VanderCruyssen:1981:CFA

[Van81]  P. Van der Cruyssen.  A

continued fraction algorithm. Numerische Mathematik, 37(1): 149–156, May 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VandeVel:1982:HCM

[Van82] H. Van de Vel. The Haar condition and multiplicity of zeros. Numerische Mathematik, 39 (1):139–153, June 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VanLoan:1985:CGS

[Van85] Charles F. Van Loan. Computing the CS and the Generalized singular value decompositions. Numerische Mathematik, 46(4): 479–491, August 1985. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VanIseghem:1994:CVS

[Van94] Jeannette Van Iseghem. Convergence of vectorial sequences. applications. Numerische Mathematik, 68(4):549–562, October 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068004/40680549.
htm;   http://science.springer.
de/nmee/bibs/4068004/40680549.
htm.

Varga:1964:SIG

[Var64] R. S. Varga. On smallest isolated Gerschgorin disks for eigenvalues. Numerische Mathematik, 6(1):366–376, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vasilopoulos:1988:DHO

[Vas88] D. Vasilopoulos. On the determination of higher order terms of singular elastic stress fields near corners. Numerische Mathematik, 53(1/2):51–95, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VanBarel:1992:NFA

[VB92a] Marc Van Barel and Adhemar Bultheel. A new formal approach to the rational interpolation problem. Numerische Mathematik, 62(1):87–122, June 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VanBarel:1992:PAD

[VB92b] Marc Van Barel and Adhemar Bultheel. A parallel algorithm for discrete least squares rational approximation. Numerische Mathematik, 63(1):99–121, October 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Varga:1981:FM

[VC81] Richard S. Varga and Da Yong Cai. On the LU factorization of M-matrices. Numerische Mathematik, 38(2):179–192, December 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verlinden:1992:CFD

[VC92a] Pierre Verlinden and Ronald Cools. On cubature formulae of degree 4k + 1 attaining Möller’s lower bound for integrals with circular symmetry. Numerische Mathematik, 61(3): 395–407, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verschelde:1992:NRS

[VC92b] Jan Verschelde and Ronald Cools. Nonlinear reduction for solving deficient polynomial systems by continuation methods. Numerische Mathematik, 63(2): 263–282, November 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Varga:1994:AZP

[VC94] Richard S. Varga and Amos J. Carpenter. Asymptotics for the zeros and poles of normalized Padé approximants to e^{z}. Numerische Mathematik, 68(1): 169–185, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680169.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680169.
htm.

vanDooren:1973:NCF

[vD73] René van Dooren. Numerical computation of forced oscillations in coupled Duffing equations. Numerische Mathematik, 20(4):300–311, August 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VandenBerghe:1991:ACH

[VD91] G. Vanden Berghe and H. De Meyer. Accurate computation of higher Sturm–Liouville eigenvalues. Numerische Mathematik, 59(3):243–254, June 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vandeGriend:1986:AMR

[vdGK86] J. A. van de Griend and J. F. B. M. Kraaijevanger. Absolute monotonicity of rational functions occurring in the numerical solution of initial value problems. Numerische Mathematik, 49(4):413–424, July 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderHouwen:1972:ERF

[vdH72] P. J. van der Houwen. Explicit Runge–Kutta formulas with increased stability boundaries. Numerische Mathematik, 20(2):149–164, April 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderHouwen:1981:BDT

[vdHtR81] P. J. van der Houwen and H. J. J. te Riele. Backward differentiation type formulas for Volterra integral equations of the second kind. Numerische Mathematik, 37(2): 205–217, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderSluis:1969:CNE

[vdS69] A. van der Sluis. Condition numbers and equilibration of matrices. Numerische Mathematik, 14(1):14–23, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderSluis:1970:CEP

[vdS70a]  A. van der Sluis.  Condi-

tion, equilibration and pivoting in linear algebraic systems. Numerische Mathematik, 15(1): 74–86, May 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderSluis:1970:SSL

[vdS70b] A. van der Sluis. Stability of solutions of linear algebraic systems. Numerische Mathematik, 14(3):246–251, February 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderSluis:1970:URP

[vdS70c]  A. van der Sluis.  Upper-

bounds for roots of polynomials. Numerische Mathematik, 15 (3):250–262, July 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderSluis:1972:RTQ

[vdS72] A. van der Sluis. The remainder term in quadrature formulae. Numerische Mathematik, 19 (1):49–55, February 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderSluis:1975:SSL

[vdS75] A. van der Sluis. Stability of the solutions of linear least squares problems. Numerische Mathematik, 23(3):241–254, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderSluis:1986:RCC

[vdSvdV86] A. van der Sluis and H. A. van der Vorst. The rate of convergence of conjugate gradients. Numerische Mathematik, 48(5):543–560, May 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Veidinger:1960:NDB

[Vei60] L. Veidinger. On the numerical determination of the best approximations in the Chebyshev sense. Numerische Mathematik, 2(1):99–105, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verner:1969:OSI

[Ver69] J. H. Verner. The order of some implicit Runge–Kutta methods. Numerische Mathematik, 13(1): 14–23, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verwer:1977:GLM

[Ver77a]  J. G. Verwer.  On gener-

alized linear multistep methods with zero-parasitic roots and an adaptive principal root. Numerische Mathematik, 27(2): 143–155, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verwer:1977:SPG

[Ver77b] J. G. Verwer. S-stability properties for generalized Runge– Kutta methods. Numerische Mathematik, 27(4):359–370, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verwer:1984:CLO

[Ver84] J. G. Verwer. Contractivity of locally one-dimensional splitting methods. Numerische Mathematik, 44(2):247–259, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verfurth:1987:FEA

[Ver87a] Rüdiger Verfürth. Finite element approximation of incompressible Navier–Stokes equations with slip boundary condition. Numerische Mathematik, 50(6):697–721, April 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verfurth:1987:MFE

[Ver87b] Rüdiger Verfürth. Mixed finite element approximation of the vector potential. Numerische Mathematik, 50(6): 685–695, April 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vermiglio:1988:NCE

[Ver88] R. Vermiglio. Natural continuous extensions of Runge–Kutta methods for Volterra integrodifferential equations. Numerische Mathematik, 53(4):439–458, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verfurth:1989:PEE

[Ver89] R. Verfürth. A posteriori error estimators for the Stokes equations. Numerische Mathematik, 55(3):309–325, June 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verfurth:1991:FEA

[Ver91a] R. Verfürth. Finite element approximation of incompressible Navier–Stokes equations with slip boundary condition. II. Numerische Mathematik, 59(6): 615–636, September 1991. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verfurth:1991:PEE

[Ver91b] R. Verfürth. A posteriori error estimators for the Stokes equations. II. non-conforming discretizations. Numerische Mathematik, 60(2):235–249, November 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vermiglio:1992:SRK

[Ver92] R. Vermiglio. On the stability of Runge–Kutta methods for delay integral equations. Numerische Mathematik, 61(4): 561–577, May 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verfurth:1998:PEE

[Ver98a] R. Verfürth. A posteriori error estimators for convectiondiffusion equations. Numerische Mathematik, 80(4):641–663, October 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080004/80800641.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080004/80800641.
pdf.

Verfurth:1998:RPE

[Ver98b] R. Verfürth. Robust a posteriori error estimators for a singularly perturbed reactiondiffusion equation. Numerische Mathematik, 78(3):479–493, January 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078003/70780479.
htm;   http://science.springer.
de/nmee/bibs/7078003/70780479.
htm.

Veselic:1976:CJM

[Ves76a] K. Veselić. A convergent Jacobi method for solving the eigenproblem of arbitrary real matrices. Numerische Mathematik, 25 (2):179–184, June 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Veselic:1976:SCJ

[Ves76b] K. Veselić. Some convergent Jacobi-like procedures for diagonalising J-symmetric matrices. Numerische Mathematik, 27(1): 67–75, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Veselic:1979:CJL

[Ves79a] K. Veselić. On a class of Jacobi-like procedures for diagonalizing arbitrary real matrices. Numerische Mathematik, 33 (2):157–172, June 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Veselic:1979:NCE

[Ves79b] K. Veselić. On a new class of elementary matrices. Numerische Mathematik, 33(2): 173–180, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Veselic:1993:JEA

[Ves93] K. Veselić. A Jacobi eigenreduction algorithm for definite matrix pairs. Numerische Mathematik, 64(2):241–269, February 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Veselic:1999:NPS

[Ves99] K. Veselić. A note on the pivoted symmetric LR iteration. Numerische Mathematik, 83(4):699–702, October 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9083004/90830699.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9083004/90830699.pdf.

vonFinckenstein:1970:PWM

[vF70] Karl Graf Finck von Finckenstein. Potenzbeschränktheit und Wertebereich einer Matrix. (German) [Potency limitation and ranges of a matrix]. Numerische Mathematik, 15(4): 329–332, August 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vonFinckenstein:1973:KDQ

[vFvH73] Karl Graf Finck von Finckenstein and Karl-Ulrich von Hagenow. Konvergenz eines Differenzenverfahrens für quasilineare parabolische Anfangs-Randwertprobleme in Zylindersymmetrie. (German) [Convergence of a difference method for quasilinear parabolic initialboundary value problems in cylindrical symmetry]. Numerische Mathematik, 20(5): 372–376, October 1973. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Golitschek:1972:CIP

[vG72] M. v. Golitschek. On the convergence of interpolating periodic spline functions of high degree. Numerische Mathematik, 19(2):146–154, April 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vonGolitschek:1980:ASM

[vG80] Manfred von Golitschek. An algorithm for scaling matrices and computing the minimum cycle mean in a digraph. Numerische Mathematik, 35(1):45– 55, March 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vonGolitschek:1982:OCD

[vG82] M. von Golitschek. Optimal cycles in double weighted graphs and approximation of bivariate functions by univariate ones. Numerische Mathematik, 39(1): 65–84, June 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vrscay:1988:EFP

[VG88] Edward R. Vrscay and William J.

Gilbert. Extraneous fixed points, basin boundaries and chaotic dynamics for Schröder and König rational iteration functions. Numerische Mathematik, 52(1):1–16, January 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vonGolitschek:1984:ASP

[vGS84] M. von Golitschek and H. Schneider. Applications of shortest path algorithms to matrix scalings. Numerische Mathematik, 44(1):111–126, June 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Veselic:1989:NOS

[VH89] K. Veselić and V. Hari. A note on a one-sided Jacobi algorithm. Numerische Mathematik, 56(6):

627–633, December 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verlinden:1993:EEC

[VH93] Pierre Verlinden and Ann Haegemans. An error expansion for cubature with an integrand with homogeneous boundary singularities. Numerische Mathematik, 65(3):383–406, July 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verwer:1990:CPR

[VHS90] J. G. Verwer, W. H. Hundsdorfer, and B. P. Sommeijer. Convergence properties of the Runge–Kutta–Chebyshev method. Numerische Mathematik, 57(2):157–178, April 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanIseghem:1984:PTA

[vI84] Jeannette van Iseghem. Padétype approximations of exp(−z) whose denominators are (1 + z/n)^{n}. Numerische Mathematik, 43(2):283–292, March 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vrahatis:1986:RGM

[VI86] M. N. Vrahatis and K. I. Iordanidis. A rapid generalized method of bisection for solving systems of non-linear equations. Numerische Mathematik, 49(2/3):123–138, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vitek:1975:BCP

[Vit75] Yehoshua Vitek. A bound connected with primitive matrices. Numerische Mathematik, 23(3): 255–260, June 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kempen:1966:CCJ

[vK66a] H. P. M. van Kempen. On the convergence of the classical Jacobi method for real symmetric matrices with non-distinct eigenvalues. Numerische Mathematik, 9(1):11–18, November 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kempen:1966:QCS

[vK66b] H. P. M. van Kempen. On the quadratic convergence of the special cyclic Jacobi method. Numerische Mathematik, 9(1): 19–22, November 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Varga:1973:MGS

[VL73] Richard S. Varga and Bernard W.

Levinger. On minimal Gerschgorin sets for families of norms. Numerische Mathematik, 20(3):252–256, June 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VandeVelde:1992:ADD

[VL92] Eric F. Van de Velde and Jens Lorenz. Adaptive data distribution for concurrent continuation. Numerische Mathematik, 62(2):269–294, July 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Varga:1993:ADI

[VN93] Richard S. Varga and Reinhard Nabben. An algorithm for determining if the inverse of a strictly diagonally dominant Stieltjes matrix is strictly ultrametric. Numerische Mathematik, 65(4):493–501, August 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vogelius:1983:AVF

[Vog83a] Michael Vogelius. An analysis of the p-version of the finite element method for nearly incompressible materials. uniformly valid, optimal error estimates. Numerische Mathematik, 41(1):39–53, April 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vogelius:1983:RID

[Vog83b] Michael Vogelius. A rightinverse for the divergence operator in spaces of piecewise polynomials. application to the p-version of the finite element method. Numerische Mathematik, 41(1):19–37, April 1983.

CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Voigt:1974:LMA

[Voi74] A. Voigt. Line method approximations to the Cauchy problem for nonlinear parabolic differential equations. Numerische Mathematik, 23(1):23– 36, February 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Voigt:1979:MLN

[Voi79] Alexander Voigt. The method of lines for nonlinear parabolic differential equations with mixed derivatives. Numerische Mathematik, 32(2):197–207, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Volk:1986:GSL

[Vol86] Wolfgang Volk. Globale Superkonvergenz bei der Lösung von linearen Randwertaufgaben. (German) [Global superconvergence in the solution of linear boundary value problems]. Numerische Mathematik, 48(6): 617–625, June 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Voss:1975:PRN

[Vos75]  H. Voss.  Projectionsver-

fahren für Randwertaufgaben mit nichtlinearen Randbedingungen. (German) [Projection methods for boundary value problems with nonlinear boundary conditions]. Numerische Mathematik, 24(4):317–329, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vonPetersdorff:1990:CMM

[vPS90] T. von Petersdorff and E. P. Stephan. On the convergence of the multigrid method for a hypersingular integral equation of the first kind. Numerische Mathematik, 57(4): 379–391, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vonPetersdorff:1996:WAF

[vPS96] Tobias von Petersdorff and Christoph Schwab. Wavelet approximations for first kind boundary integral equations on polygons. Numerische Mathematik, 74(4):479–516, October 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6074004/60740479.
htm;   http://science.springer.
de/nmee/bibs/6074004/60740479.
htm.

vonSydow:1977:EEG

[vS77] Björn von Sydow. Error estimates for Gaussian quadrature formulae. Numerische Mathematik, 29(1):59–64, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vuik:1990:EEA

[Vui90] C. Vuik. An L^{2}-error estimate for an approximation of the solution of a parabolic variational inequality. Numerische Mathematik, 57(5):453–471, June 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanVeldhuizen:1976:RPC

[vV76] M. van Veldhuizen. A refinement process for collocation approximations. Numerische Mathematik, 26(4): 397–407, December 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanVeldhuizen:1977:NPP

[vV77] M. van Veldhuizen. A note on partial pivoting and Gaussian elimination. Numerische Mathematik, 29(1):1–10, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanVeldhuizen:1978:HOM

[vV78] M. van Veldhuizen. Higher order methods for a singularly perturbed problem. Numerische Mathematik, 30(3): 267–279, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanVeldhuizen:1983:SS

[vV83] M. van Veldhuizen. On Dstability and B-stability. Numerische Mathematik, 42(3): 349–357, December 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vanninathan:1984:ATP

[VV84] M. Vanninathan and G. D. Veerappa Gowda. Approximation of Tricomi problem with Neumann boundary condition. Numerische Mathematik, 44(3): 371–391, September 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verdi:1988:EES

[VV88] Claudio Verdi and Augusto Visintin. Error estimates for a semi-explicit numerical scheme for Stefan-type problems. Numerische Mathematik, 52(2): 165–185, January 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

VanHuffel:1989:ACB

[VV89] Sabine Van Huffel and Joos Vandewalle. Algebraic connections between the least squares and total least squares problems. Numerische Mathematik, 55(4):431–449, July 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verdiere:1998:NTS

[VV98] S. Verdière and M. H. Vignal. Numerical and theoretical study of a Dual Mesh Method using finite volume schemes for two phase flow problems in porous media. Numerische Mathematik, 80(4):601–639, October 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080004/80800601.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080004/80800601.
pdf.

vonWelck:1977:SSA

[vW77a] U. von Welck. Ein stabiles Schichtenverfahren für allgemeine lineare Evolutionsgleichungen. (German) [A stable line method for general linear evolution equations]. Numerische Mathematik, 27(2): 171–178, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vonWelck:1977:SMP

[vW77b]  U. von Welck.  Stabile

Merhschichtverfahren für parabolische Evolutionsgleichungen. (German) [Stable multi-line methods for evolution equations of parabolic type]. Numerische Mathematik, 29(1): 65–82, March 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Veselic:1979:QCJ

[VW79] K. Veselić and H. J. Wenzel. A quadratically convergent Jacobi-like method for real matrices with complex eigenvalues. Numerische Mathematik, 33(4): 425–435, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Vassilevski:1992:MIM

[VW92] Panayot S. Vassilevski and Junping Wang. Multilevel iterative methods for mixed finite element discretizations of elliptic problems. Numerische Mathematik, 63(4):503–520, December 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanWijngaarden:1969:RALa

[vWMPK69] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Report on the algorithmic language AL-GOL 68. Numerische Mathematik, 14(2):79–218, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

                                        //www.fh-jena.de/~kleine/
                                        history/languages/Algol68-
                                        Report.pdf;   http://www.
                                        fh-jena.de/~kleine/history/
                                        languages/Algol68-RevisedReport.
                                        pdf.

Vanmaele:1995:CEN

[VŽ95] Michèle Vanmaele and Alexander Ženíšek. The combined effect of numerical integration and approximation of the boundary in the finite element method for eigenvalue problems. Numerische Mathematik, 71(2):

253–273, August 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071002/50710253.
htm;   http://science.springer.
de/nmee/bibs/5071002/50710253.
htm.

Weiss:1972:PIM

[WA72] R. Weiss and R. S. Anderssen.

A product integration method for a class of singular first kind Volterra equations. Numerische Mathematik, 18(5):442–456, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wagenfuhrer:1976:VHK

[Wag76] E. Wagenführer. Ein Verfahren höherer Konvergenzordnung zur Berechnung des charakteristischen Exponenten der Mathieuschen Differentialgleichung. (German) [A method of high order convergence for calculating the characteristic exponent of Mathieu’s differential equation]. Numerische Mathematik, 27(1):53–65, March 1976. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wagenfuhrer:1980:DBC

[Wag80] E. Wagenführer. Die Determinantenmethode zur Berechnung des charakteristischen Exponenten der endlichen Hillschen Differentialgleichung. (German) [On the determinantal method for calculating the characteristic exponent of the finite Hill differential equation]. Numerische Mathematik, 35(4): 405–420, November 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wagatha:1983:APO

[Wag83a] Ludwig Wagatha. Approximation of pseudodifferential operators in absorbing boundary conditions for hyperbolic equations. Numerische Mathematik, 42(1): 51–64, September 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wagenfuhrer:1983:CEF

[Wag83b] Ekkehard Wagenführer. On calculating the eigenvalues of the finite Hill’s differential equation. Numerische Mathematik, 41(2): 255–279, June 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wagner:1997:TFFa

[Wag97a] Christian Wagner. Tangential frequency filtering decompositions for symmetric matrices. Numerische Mathematik, 78(1): 119–142, November 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/7078001/
                                        70780119.htm;   http://science.
                                        springer.de/nmee/bibs/7078001/
                                        70780119.htm.

Wagner:1997:TFFb

[Wag97b] Christian Wagner. Tangential frequency filtering decompositions for unsymmetric matrices. Numerische Mathematik, 78(1):143–163, November 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7078001/70780143.
htm;   http://science.springer.
de/nmee/bibs/7078001/70780143.
htm.

Wahba:1975:SND

[Wah75a] Grace Wahba. Smoothing noisy data with spline functions. Numerische Mathematik, 24(5): 383–393, October 1975. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wahlbin:1975:EEG

[Wah75b] Lars Wahlbin. Error estimates for a Galerkin method for a class of model equations for long waves. Numerische Mathematik, 23(4):289–303, August 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Walter:1967:BIL

[Wal67] Wolfgang Walter. Bemerkungen zu Iterationsverfahren bei linearen Gleichungssystemen. (German) [Remark on iteration method for linear equation systems]. Numerische Mathematik, 10(1):80–85, May 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Walter:1968:LNP

[Wal68] Wolfgang Walter. Die Linienmethode bei nichtlinearen parabolischen Differentialgleichungen. (German) [The line method for nonlinear parabolic differential equations]. Numerische Mathematik, 12(4): 307–321, November 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Walz:1989:CIR

[Wal89] Guido Walz. A contour integral representation of linear extrapolation methods. Numerische Mathematik, 55(4): 477–480, July 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Walz:1991:AEA

[Wal91] Guido Walz. Asymptotic expansions and acceleration of convergence for higher order iteration processes. Numerische Mathematik, 59(5):529–540, August 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Waldron:1997:IEF

[Wal97]  Shayne Waldron.  Integral

error formulae for the scale of mean value interpolations which includes Kergin and Hakopian interpolation. Numerische Mathematik, 77(1): 105–122, July 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077001/70770105.
htm;   http://science.springer.
de/nmee/bibs/7077001/70770105.
htm.

Waldron:1998:MTF

[Wal98] Shayne Waldron. Multipoint Taylor formulæ. Numerische Mathematik, 80(3):461–494, September 1998. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/8080003/80800461.
htm;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800461.
pdf;   http://link.springer.
de/link/service/journals/
00211/papers/8080003/80800461.
ps.gz.

Wampler:1994:ESS

[Wam94] Charles W. Wampler. An efficient start system for multihomogeneous polynomial continuation. Numerische Mathematik, 66(4):517–523, January 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wang:1989:AEE

[Wan89] Junping Wang. Asymptotic expansions and L^{∞}-error estimates for mixed finite element methods for second order elliptic problems. Numerische Mathematik, 55(4):401–430, July 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Watt:1968:CCS

[Wat68] J. M. Watt. Consistency, convergence and stability of general discretizations of the initial value problem. Numerische Mathematik, 12(1):11–22, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1969:UZI

[WB69] H. Werner and Dietrich Braess. Über den Zusammenhang von Interpolation und diskreter T - Approximation mit rationalen Funktionen. (German) [On the coherence of interpolation and discrete T approximation with rational functions]. Numerische Mathematik, 13(2): 112–128, May 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weber:1979:NBV

[Web79] Helmut Weber. Numerische Behandlung von Verzweigungsproblemen bei gewöhnlichen Differentialgleichungen. (German) [Numerical treatment of bifurcation problems for ordinary differential equations]. Numerische Mathematik, 32(1): 17–29, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weber:1981:NCF

[Web81] H. Weber. Numerical computation of the Fourier transform using Laguerre functions and the fast Fourier transform. Numerische Mathematik, 36(2): 197–209, April 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weba:1991:IRF

[Web91] Michael Weba. Interpolation of random functions. Numerische Mathematik, 59(7):739–746, October 1991. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weber:1994:MMP

[Web94] Hermann Weber. A multigrid method for a Petrov–Galerkin discretization of the Stokes equations. Numerische Mathematik, 66(4):525–541, January 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chu:1986:GBF

[wEC86] King wah Eric Chu. Generalization of the Bauer–Fike Theorem. Numerische Mathematik, 49(6):685–691, November 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wegmann:1978:IKA

[Weg78] Rudolf Wegmann. Ein Iterationsverfahren zur konformen Abbildung. (German) [An iterative method in conformal mapping]. Numerische Mathematik, 30(4):453–466, December 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wegmann:1984:CPE

[Weg84] Rudolf Wegmann. Convergence proofs and error estimates for an iterative method for conformal mapping. Numerische Mathematik, 44(3):435–461, September 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wegmann:1989:CMM

[Weg89] Rudolf Wegmann. Conformal mapping by the method of alternating projections. Numerische Mathematik, 56(2/3): 291–307, October 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weisfeld:1959:OPS

[Wei59] M. Weisfeld. Orthogonal polynomials in several variables. Numerische Mathematik, 1:38– 40, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weinstein:1963:SLT

[Wei63]  Alexander Weinstein.  On

the Sturm–Liouville theory and the eigenvalues of intermediate problems. Numerische Mathematik, 5(1):238–245, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weinitschke:1964:UKI

[Wei64]  H. J. Weinitschke.  Über

eine Klasse von Iterationsverfahren. (German) [On a class of iteration methods]. Numerische Mathematik, 6(1):395– 404, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weisel:1980:NEQ

[Wei80] Jutta Weisel. Numerische Ermittlung quasikonformer Abbildungen mit finiten Elementen. (German) [Numerical determination of quasiconformal mappings by finite elements]. Numerische Mathematik, 35(2): 201–222, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wei:1992:ARB

[Wei92a] Mu Sheng Wei. Algebraic relations between the total least squares and least squares problems with more than one solution. Numerische Mathematik, 62(1):123–148, June 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Weideman:1992:EHR

[Wei92b] J. A. C. Weideman. The eigenvalues of Hermite and rational spectral differentiation matrices. Numerische Mathematik, 61(3): 409–432, April 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wei:1995:UBS

[Wei95] Musheng Wei. Upper bound and stability of scaled pseudoinverses. Numerische Mathematik, 72(2):285–293, December 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5072002/
50720285.htm;   http://science.
springer.de/nmee/bibs/5072002/
50720285.htm.

Wendland:1968:BRI

[Wen68] Wolfgang Wendland. Die Behandlung von Randwertaufgaben im R_{3} mit Hilfe von Einfach- und Doppelschichtpotentialen. (German) [The treatment of boundary-value problems in R_{3} by simple and double blocking layer potentials]. Numerische Mathematik, 11(5): 380–404, June 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1962:SDT

[Wer62]  Helmut Werner.  Ein Satz

über diskete Tschebyscheff-Approximation bei gebrochen linearer Funktionen. (German) [A proposition on the discrete Chebyshev approximation of broken linear functions]. Numerische Mathematik, 4:154– 157, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1968:NSS

[Wer68] Wolfgang Werner. Numerical solution of systems of quasilinear hyperbolic differential equations by means of the method of nebencharacteristics in combination with extrapolation methods. Numerische Mathematik, 11(2):151–169, February 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1969:ENG

[Wer69] Jochen Werner. Einschließungssaätze bei nichtlinearen gewönlichen Randwertaufgaben und erzwungene Schwingungen. (German) [Enclosure principles for nonlinear ordinary boundary value problems and forced vibrations]. Numerische Mathematik, 13(1): 24–38, March 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1972:FRI

[Wer72] Helmut Werner. Eine faktorisierung der bei der rationalen Interpolation auftretenden Matrizen. (German) [A factorization of matrices occurring in rational interpolation]. Numerische Mathematik, 18(5): 423–431, October 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1978:GKV

[Wer78] Jochen Werner. Über die globale Konvergenz von Variable-Metrik-Verfahren mit nichtexakter Schrittweitenbestimmung. (German) [On the global convergence of variable metric algorithms without exact line searches]. Numerische Mathematik, 31(3):321–334, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1979:VON

[Wer79] Wilhelm Werner. Über ein Verfahren der Ordnung 1 + \sqrt{2} zur Nullstellenbestimmung. (German) [On a method of order 1 + \sqrt{2} for determining zeros]. Numerische Mathematik, 32(3):333–342, September 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1982:SSR

[Wer82] Wilhelm Werner. Some supplementary results on the 1 + \sqrt{2} order method for the solution of nonlinear equations. Numerische Mathematik, 38(3): 383–392, March 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1986:MPA

[Wer86]  Helmut Werner.  Multivari-

ate Padé approximation. Numerische Mathematik, 48(4): 429–440, April 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wesseling:1975:AEP

[Wes75] P. Wesseling. An asymptotic expansion for product integration applied to Cauchy principal value integrals. Numerische Mathematik, 24(5):435–442, October 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wetterling:1960:EKB

[Wet60]  W. Wetterling.  Zum Ein-

schließungssatz von Kryloff-Bogoliubov für allgemeine Eigenwertaufgaben bei gewöhnlichen Differentialgleichungen. (German) [An enclosure theorem of Kryloff-Bogoliubov for ordinary eigenvalue problems in ordinary differential equations]. Numerische Mathematik, 2(1): 18–21, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wetterling:1970:DRE

[Wet70] W. Wetterling. Definitheitsbedingungen für relative Extrema bei Optimierungs- und Approximationsaufgaben. (German) [Definiteness conditions for relative extrema in optimization and approximation problems]. Numerische Mathematik, 15(2):122–136, June 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wathen:1995:CRM

[WFS95] Andrew Wathen, Bernd Fischer, and David Silvester. The convergence rate of the minimal residual method for the Stokes problem. Numerische Mathematik, 71(1): 121–134, July 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/5071001/50710121.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/5071001/50710121.
                                        htm.

Werner:1986:LDS

[WH86] Helmut Werner and H. Hilgers. Lösung von Differentialgleichungen mit Splinefunktionen. Eine Störungstheorie. I. Divergenzaussagen. (German) [Solution of differential equations by splines. A perturbation theory. I. statements about divergence]. Numerische Mathematik, 48(3): 323–336, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

White:1977:OEP

[Whi77] Bran E. White. On optimal extreme-discrepancy point sets in the square. Numerische Mathematik, 27(2): 157–164, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wendland:1988:ABE

[WhY88] W. L. Wendland and De hao Yu. Adaptive boundary element methods for strongly elliptic integral equations. Numerische Mathematik, 53(5):539–558, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Widlund:1966:SPD

[Wid66] Olaf B. Widlund. Stability of parabolic difference schemes in the maximum norm. Numerische Mathematik, 8(2):186– 202, April 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Widlund:1977:BEB

[Wid77] Olof Widlund. On best error bounds for approximation by piecewise polynomial functions. Numerische Mathematik, 27(3): 327–338, September 1977. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1959:EZIa

[Wil59a] J. H. Wilkinson. The evaluation of the zeros of ill-conditioned polynomials. I. Numerische Mathematik, 1:150–166, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1959:EZIb

[Wil59b] J. H. Wilkinson. The evaluation of the zeros of ill-conditioned polynomials. II. Numerische Mathematik, 1:167–180, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1960:EAF

[Wil60] J. H. Wilkinson. Error analysis of floating-point computation. Numerische Mathematik, 2(1):319–340, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL https://

link.springer.com/article/
10.1007/BF01386233.

Wilson:1961:DDP

[Wil61] Walter L. Wilson. On discrete Dirichlet and plateau problems. Numerische Mathematik, 3(1): 359–373, December 1961. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=3&issue=1&spage=
359.

Wilkinson:1962:CESa

[Wil62a] J. H. Wilkinson. Calculation of the eigenvalues of a symmetric tridiagonal matrix by the method of bisection. Numerische Mathematik, 4:362– 367, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1962:CESb

[Wil62b] J. H. Wilkinson. Calculation of the eigenvectors of a symmetric tridiagonal matrix by inverse iteration. Numerische Mathematik, 4:368–376, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1962:HMS

[Wil62c] J. H. Wilkinson. Householder’s method for symmetric matrices. Numerische Mathematik, 4:354– 361, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1962:NQC

[Wil62d] J. H. Wilkinson. Note on the quadratic convergence of the cyclic Jacobi process. Numerische Mathematik, 4:296– 300, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilson:1962:MSS

[Wil62e] Walter L. Wilson, Jr. d-minimal surfaces spanning polygons in E_{n} by solving a linear system. Numerische Mathematik, 4:328– 329, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilf:1964:ECN

[Wil64] H. Wilf. Exactness conditions in numerical quadrature. Numerische Mathematik, 6(1):315– 319, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Willson:1971:UGM

[Wil71] A. N. Willson, Jr. A useful generalization of the P_{0} matrix concept. Numerische Mathematik, 17(1):62–70, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1972:NMV

[Wil72] J. H. Wilkinson. Note on matrices with a very illconditioned eigenproblem. Numerische Mathematik, 19(2):

176–178, April 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1978:ASN

[Wil78] J. H. Wilkinson. The Alston S. Householder Award. Numerische Mathematik, 29(4): 463, April 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1984:NMQ

[Wil84] J. H. Wilkinson. On neighbouring matrices with quadratic elementary divisors. Numerische Mathematik, 44(1):1–21, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). In honour of the 60th birthday of F. L. Bauer.

Wild:1987:ACP

[Wil87] Paul Wild. Accelerating the convergence of power series of certain entire functions. Numerische Mathematik, 51(5): 583–595, October 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilderotter:1997:AOA

[Wil97] Klaus Wilderotter. An asymptotically optimal algorithm for approximating bounded analytic functions. Numerische Mathematik, 75(3):397–404, January 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/

00211/bibs/7075003/70750397.
htm;   http://science.springer.
de/nmee/bibs/7075003/70750397.
htm.

Wimp:1974:TAL

[Wim74] J. Wimp. Toeplitz arrays, linear sequence transformations and orthogonal polynomials. Numerische Mathematik, 23(1):1– 17, February 1974. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Witsch:1977:KPL

[Wit77] Kristian Witsch. Konvergenzaussagen für Projectionsverfahren bei linearen Operatoren. (German) [Convergence of projection methods for linear operators]. Numerische Mathematik, 27(3):339–354, September 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Witsch:1978:NQP

[Wit78a] Kristian Witsch. Numerische Quadratur bei Projektionsverfahren. (German) [Numerical quadrature in projection methods]. Numerische Mathematik, 30(2):185–206, June 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Witsch:1978:PNVa

[Wit78b] Kristian Witsch. Projektive Newton-Verfahren bei elliptischen Randwertaufgaben. (German) [Projective Newton methods for elliptic boundary value problems]. Numerische Mathematik, 30(3):333–347, September 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Witsch:1978:PNVb

[Wit78c] Kristian Witsch. Projektive Newton-Verfahren und Anwendungen auf nichtlineare Randwertaufgaben. (German) [Projective Newton methods and applications to nonlinear boundary value problems]. Numerische Mathematik, 31(2): 209–230, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wittum:1989:MGM

[Wit89] G. Wittum. Multi-grid methods for Stokes and Navier– Stokes equations. transforming smoothers: algorithms and numerical results. Numerische Mathematik, 54(5): 543–563, February 1989. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wittum:1990:CMG

[Wit90] Gabriel Wittum. On the convergence of multi-grid methods with transforming smoothers — theory with applications to the Navier–Stokes equations. Numerische Mathematik, 57(1):15– 38, March 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wittum:1997:SI

[Wit97] G. Wittum. Shifted iterations.

Numerische Mathematik, 76(2): 265–278, April 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076002/70760265.
htm;   http://science.springer.
de/nmee/bibs/7076002/70760265.
htm.

Wagner:1997:SCM

[WKW97] C. Wagner, W. Kinzelbach, and G. Wittum. Schur-complement multigrid. A robust method for groundwater flow and transport problems. Numerische Mathematik, 75(4):523–545, February 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7075004/70750523.
htm;   http://science.springer.
de/nmee/bibs/7075004/70750523.
htm.

Wagenfuhrer:1979:BCE

[WL79] E. Wagenführer and H. Lang.

Berechnung des charakteristischen Exponenten der endlichen Hillschen Differentialgleichung durch numerische Integration. (German) [On calculating the characteristic exponent of the finite Hill differential equation by numerical integration]. Numerische Mathematik, 32(1): 31–50, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wohlfahrt:1963:UBP

[Woh63] Klaus Wohlfahrt. Über die Berechnung der Periodenverhältnisse elliptischer Gebilde. (German) [On the calculation of the relationship of the periods of elliptic structures]. Numerische Mathematik, 4:393–397, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wohlmuth:1999:RBE

[Woh99]  Barbara I. Wohlmuth.  A

residual based error estimator for mortar finite element discretizations. Numerische Mathematik, 84(1):143–171, November 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/journals/00211/bibs/
                                        9084001/90840143.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/journals/00211/papers/
                                        9084001/90840143.pdf.

Wolf:1978:AEE

[Wol78a] R. Wolf. Asymptotische Entwicklungen für Eigenwerte und Eigenvektoren bei der Approximation parameternichtlinearer Eigenwertaufgaben. (German) [Asymptotic expansions for eigenvalues and eigenvectors in the approximation of eigenvalue problems depending nonlinearly on a parameter]. Numerische Mathematik, 30(2):

207–226, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wolfe:1978:EIM

[Wol78b] M. A. Wolfe. Extended iterative methods for the solution of operator equations. Numerische Mathematik, 31(2): 153–174, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wolfe:1979:CAD

[Wol79]  Jerry M. Wolfe.  On the

convergence of an algorithm for discrete L_{p} approximation. Numerische Mathematik, 32(4): 439–459, December 1979. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wong:1982:QFO

[Won82] R. Wong. Quadrature formulas for oscillatory integral transforms. Numerische Mathematik, 39(3):351–360, October 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wong:1984:CMS

[Won84] Wing Hung Wong. On constrained multivariate splines and their approximations. Numerische Mathematik, 43(1): 141–152, January 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wong:1992:PNFa

[Won92a] Sze-Ping Wong. Preconditioning nonconforming finite element methods for treating Dirichlet boundary conditions. I. Numerische Mathematik, 62(3): 391–411, August 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wong:1992:PNFb

[Won92b] Sze-Ping Wong. Preconditioning nonconforming finite element methods for treating Dirichlet boundary conditions. II. Numerische Mathematik, 62(3): 413–438, August 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wong:1993:PNF

[Won93]  Sze-Ping Wong.  Precondi-

tioning P_{1} nonconforming finite elements: condition numbers and singular value distributions. Numerische Mathematik, 65(2): 253–271, June 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wood:1968:ENM

[Woo68] W. L. Wood. On an explicit numerical method for solving the biharmonic equation. Numerische Mathematik, 11(5): 413–426, June 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1977:NSS

[Woź77a] H. Woźniakowski. Numerical stability for solving nonlinear equations. Numerische Mathematik, 27(4):373–390, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1977:NSC

[Woź77b]  H. Woźniakowski.  Numer-

ical stability of the Chebyshev method for the solution of large linear systems. Numerische Mathematik, 28(2): 191–209, June 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wozniakowski:1978:REA

[Woź78]  H. Woźniakowski.  Round-

off error analysis of iterations for large linear systems. Numerische Mathematik, 30(3): 301–314, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation, Editors: F. L. Bauer, A. S. Householder, F. W. J. Olver, H. Rutishauser, K. Samelson and E. Stiefel. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6. viii + 439 pp. LCCN QA251 .W67.

Wright:1992:DEA

[Wri92] K. Wright. Differential equations for the analytic singular value decomposition of a matrix. Numerische Mathematik, 63(2): 283–295, November 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wright:1994:SPE

[Wri94]  Stephen J. Wright.  Sta-

ble parallel elimination for boundary value ODEs. Numerische Mathematik, 67(4): 521–535, May 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067004/40670521.
htm;   http://science.springer.
de/nmee/bibs/4067004/40670521.
htm.

Walther:1959:NBG

[WS59] A. Walther and H. Schappert. Numerische Behandlung des Gelenkvierecks. (German) [Numerical treatment of general quadrilaterals]. Numerische Mathematik, 1:110–120, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Werner:1967:RCA

[WSB67] H. Werner, J. Stoer, and W. Bommas. Rational Chebyshev approximation. Numerische Mathematik, 10(4): 289–306, November 1967. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wulbert:1973:NPS

[Wul73] Daniel Wulbert. A note on polynomial splines with free knots. Numerische Mathematik, 21(2): 181–184, April 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wuytack:1971:NTR

[Wuy71] L. Wuytack. A new technique for rational extrapolation to the limit. Numerische Mathematik, 17(3):215–221, June 1971. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wuytack:1973:ARI

[Wuy73] Luc Wuytack. An algorithm for rational interpolation similar to the qd-algorithm. Numerische Mathematik, 20(5):418–424, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wets:1968:TAC

[WW68] R. J.-B. Wets and Christoph Witzgall. Towards an algebraic characterization of convex polyhedral cones. Numerische Mathematik, 12(2):134–138, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wasilkowski:1984:CAH

[WW84] G. W. Wasilkowski and H. Woźniakowski.

Can adaption help on the average? Numerische Mathematik, 44(2):169–190, August 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wagner:1997:AF

[WW97] Christian Wagner and Gabriel Wittum. Adaptive filtering. Numerische Mathematik, 78(2): 305–328, December 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/7078002/
70780305.htm;   http://science.
springer.de/nmee/bibs/7078002/
70780305.htm.

Wynn:1959:CFCa

[Wyn59a] P. Wynn. Converging factors for continued fractions. Part I. Numerische Mathematik, 1(1): 272–307, December 1959. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
272.

Wynn:1959:CFCb

[Wyn59b] P. Wynn. Converging factors for continued fractions. Part II. Numerische Mathematik, 1(1): 308–320, December 1959. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
308.

Wynn:1959:PEC

[Wyn59c] P. Wynn. On the propagation of error in certain nonlinear algorithms. Numerische Mathematik, 1:142–149, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wynn:1959:SCI

[Wyn59d] P. Wynn. A sufficient condition for the instability of the q-d algorithm. Numerische Mathematik, 1(1):203–207, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=article&issn=0029-
599X&volume=1&issue=1&spage=
203.

Wynn:1960:UIA

[Wyn60]  Peter Wynn.  Über einen

Interpolations-Algorithmus und gewisse andere Formeln, die in der Theorie der Interpolation durch rationale Funktionen bestehen. (German) [On interpolation algorithms and certain other formulas existing in the interpolations theory by rational functions]. Numerische Mathematik, 2(1):151–182, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wynn:1962:CTN

[Wyn62] P. Wynn. A comparison technique for the numerical transformation of slowly convergent series based on the use of rational functions. Numerische Mathematik, 4:8–14, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wynn:1963:NCF

[Wyn63] P. Wynn. Note on a converging factor for a certain continued fraction. Numerische Mathematik, 5(1):332–352, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wynn:1964:GPV

[Wyn64] P. Wynn. General purpose vector epsilon algorithm ALGOL procedures. Numerische Mathematik, 6(1):22–36, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wynn:1966:USR

[Wyn66] P. Wynn. Upon systems of recursions which obtain among the quotients of the Padé table. Numerische Mathematik, 8(3):264–269, May 1966. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Wells:1995:WSD

[WZ95] Raymond O. Wells, Jr. and Xiaodong Zhou. Wavelet solutions for the Dirichlet problem. Numerische Mathematik, 70(3): 379–396, May 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5070003/50700379.
htm;   http://science.springer.
de/nmee/bibs/5070003/50700379.
htm.

Xu:1992:GWM

[XS92] Jin-Chao Xu and Wei-Chang Shann. Galerkin-wavelet methods for two-point boundary value problems. Numerische Mathematik, 63(1):123–144, October 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Xu:1994:CBO

[Xu94] Yuan Xu. A class of bivariate orthogonal polynomials and cubature formula. Numerische Mathematik, 69(2):233–241, December 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4069002/40690233.
htm;   http://science.springer.
de/nmee/bibs/4069002/40690233.
htm.

Yuan:1995:CNT

[xY95] Ya xiang Yuan. On the convergence of a new trust region algorithm. Numerische Mathematik, 70(4):515–539, June 1995. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/5070004/
50700515.htm;   http://science.

                                        springer.de/nmee/bibs/5070004/
                                        50700515.htm.

Yamamoto:1966:CMD

[Yam66] Tetsuro Yamamoto. A computational method for the dominant root of a nonnegative irreducible matrix. Numerische Mathematik, 8(4):324–333, June 1966. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1968:ECO

[Yam68] Tetsuro Yamamoto. On the eigenvalues of compact operators in a Hilbert space. Numerische Mathematik, 11(3): 211–219, March 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1975:DEI

[Yam75] Tetsuro Yamamoto. On the distribution of errors in the iterative solution of a system of linear equations. Numerische Mathematik, 24(1):71–79, February 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1976:DEI

[Yam76] Tetsuro Yamamoto. On the distribution of errors in the iterative solution of a system of linear equations. II. Numerische Mathematik, 25(4):461–463, December 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1980:EBC

[Yam80]  Tetsuro Yamamoto.  Error

bounds for computed eigenvalues and eigenvectors. Numerische Mathematik, 34(2): 189–199, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1982:EBC

[Yam82]  Tetsuro Yamamoto.  Error

bounds for computed eigenvalues and eigenvectors. II. Numerische Mathematik, 40(2): 201–206, December 1982. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1986:EBN

[Yam86a] Tetsuro Yamamoto. Error bounds for Newton’s iterates derived from the Kantorovich theorem. Numerische Mathematik, 48(1):91–98, January 1986. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1986:MFS

[Yam86b] Tetsuro Yamamoto. A method for finding sharp error bounds for Newton’s method under the Kantorovich assumptions. Numerische Mathematik, 49(2/3): 203–220, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1987:CTN

[Yam87] Tetsuro Yamamoto. A convergence theorem for Newtonlike methods in Banach spaces.

Numerische Mathematik, 51(5): 545–557, October 1987. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yang:1994:GMW

[Yan94]  Daoqi Yang.  Grid modi-

fication for the wave equation with attenuation. Numerische Mathematik, 67(3): 391–401, April 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4067003/40670391.
htm;   http://science.springer.
de/nmee/bibs/4067003/40670391.
htm.

Yao:1999:HI

[Yao99] Qingchuan Yao. On Halley iteration. Numerische Mathematik, 81(4):647–677, February 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9081004/90810647.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9081004/90810647.pdf.

Ye:1995:CET

[Ye95] Qiang Ye. On close eigenvalues of tridiagonal matrices. Numerische Mathematik, 70(4): 507–514, June 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.

de/link/service/journals/
00211/bibs/5070004/50700507.
htm;   http://science.springer.
de/nmee/bibs/5070004/50700507.
htm.

Guo:1993:FCP

[yGL93] Ben yu Guo and Jian Li.

Fourier–Chebyshev pseudospectral method for two-dimensional vorticity equation. Numerische Mathematik, 66(3):329–346, December 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ying:1988:RAP

[YK88] Xingren Ying and I. Norman Katz. A reliable argument principle algorithm to find the number of zeros of an analytic function in a bounded domain. Numerische Mathematik, 53(1/2): 143–163, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yu:1986:AND

[YM86] Guo Ben Yu and A. R. Mitchell.

Analysis of a nonlinear difference scheme in reactiondiffusion. Numerische Mathematik, 49(5):511–527, September 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yamamoto:1993:NVS

[YN93] Nobito Yamamoto and Mitsuhiro T. Nakao. Numerical verifications of solutions for elliptic equations in nonconvex polygonal domains. Numerische Mathematik, 65(4):503–521, August 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Young:1971:BOR

[You71] David M. Young. A bound for the optimum relaxation factor for the successive overrelaxation method. Numerische Mathematik, 16(5):408–413, February 1971. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ypma:1984:CNL

[Ypm84] T. J. Ypma. Convergence of Newton-like-iterative methods. Numerische Mathematik, 45(2): 241–251, November 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yserentant:1980:MLE

[Yse80] Harry Yserentant. Die Mehrstellenformeln für den Laplaceoperator. (German) [The Hermitian formulas for Laplace’s operator]. Numerische Mathematik, 34(2):171–187, June 1980. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yserentant:1983:MKD

[Yse83] Harry Yserentant. Die maximale Konsistenzordnung von Differenzenapproximationen nichtnegativer Art. (German) [The maximal consistency order of difference approximations of nonnegative kind]. Numerische Mathematik, 42(1): 119–123, September 1983. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yserentant:1986:EMS

[Yse86a] Harry Yserentant. Erratum:

“On the multilevel splitting of finite element spaces”. Numerische Mathematik, 50(1): 123, January 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See [Yse86b].

Yserentant:1986:MLS

[Yse86b] Harry Yserentant. On the multilevel splitting of finite element spaces. Numerische Mathematik, 49(4):379–412, July 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Yse86a].

Yserentant:1989:PID

[Yse89] Harry Yserentant. Preconditioning indefinite discretization matrices. Numerische Mathematik, 54(6):719–734, March 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Yserentant:1990:TPB

[Yse90] Harry Yserentant. Two preconditioners based on the multilevel splitting of finite element spaces. Numerische Mathematik, 58(2):163–184, September 1990. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Yserentant:1997:NCP

[Yse97a]  Harry Yserentant.  A new

class of particle methods. Numerische Mathematik, 76(1):87– 109, March 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076001/70760087.
htm;   http://science.springer.
de/nmee/bibs/7076001/70760087.
htm.

Yserentant:1997:PMC

[Yse97b] Harry Yserentant. A particle model of compressible fluids. Numerische Mathematik, 76(1): 111–142, March 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7076001/70760111.
htm;   http://science.springer.
de/nmee/bibs/7076001/70760111.
htm.

Yserentant:1999:EGS

[Yse99a] Harry Yserentant. Entropy generation and shock resolution in the particle model of compressible fluids. Numerische Mathematik, 82(1):161–177, March 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082001/90820161.htm;   http:
//link.springer-ny.com/link/

service/journals/00211/papers/
9082001/90820161.pdf.

Yserentant:1999:PVS

[Yse99b] Harry Yserentant. Particles of variable size. Numerische Mathematik, 82(1):143–159, March 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00211/bibs/
9082001/90820143.htm;   http:
//link.springer-ny.com/link/
service/journals/00211/papers/
9082001/90820143.pdf.

Zadunaisky:1976:EEP

[Zad76] Pedro E. Zadunaisky. On the estimation of errors propagated in the numerical integration of ordinary differential equations. Numerische Mathematik, 27(1): 21–39, March 1976. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zahar:1977:MAM

[Zah77] R. V. M. Zahar. A mathematical analysis of Miller’s algorithm. Numerische Mathematik, 27(4):427–447, December 1977. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zawilski:1991:NSC

[Zaw91] Adam T. Zawilski. Numerical stability of the cyclic Richardson iteration. Numerische Mathematik, 60(2):251–290, November 1991. CODEN NUMMA7.

ISSN 0029-599X (print), 0945-3245 (electronic).

Zenger:1968:CPB

[Zen68] Chr. Zenger. On convexity properties of the Bauer field of values of a matrix. Numerische Mathematik, 12(2):96–105, August 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zenisek:1970:IPT

[Žen70] Alexander Ženíšek. Interpolation polynomials on the triangle. Numerische Mathematik, 15 (4):283–296, August 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zenger:1972:CSB

[Zen72] Christoph Zenger. A comparison of some bounds for the nontrivial eigenvalues of stochastic matrices. Numerische Mathematik, 19(3):209–211, June 1972. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zenger:1984:PCS

[Zen84a] C. Zenger. Positivity in complex spaces and its application to Gershgorin discs. Numerische Mathematik, 44(1):67–73, June 1984. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zennaro:1984:MPL

[Zen84b] Marino Zennaro. Maximum principles for linear differencedifferential operators in periodic function spaces. Numerische Mathematik, 43(1): 121–139, January 1984. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zennaro:1986:SPR

[Zen86] Marino Zennaro. P -stability properties of Runge–Kutta methods for delay differential equations. Numerische Mathematik, 49(2/3):305–318, March 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zennaro:1988:NRK

[Zen88]  Marino Zennaro.  Natural

Runge–Kutta and projection methods. Numerische Mathematik, 53(4):423–438, August 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zenisek:1989:FEV

[Žen89] Alexander Ženíšek. Finite element variational crimes in parabolic-elliptic problems. I. nonlinear schemes. Numerische Mathematik, 55(3):343–376, June 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zenisek:1990:FEM

[Žen90] Alexander Ženíšek. The finite element method for nonlinear elliptic equations with discontinuous coefficients. Numerische Mathematik, 58(1):51– 77, September 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zenisek:1995:MAC

[Žen95] Alexander Ženíšek. The maximum angle condition in the finite element method for monotone problems with applications in magnetostatics. Numerische Mathematik, 71(3):399–417, September 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5071003/50710399.
htm;   http://science.springer.
de/nmee/bibs/5071003/50710399.
htm.

Zennaro:1997:ASA

[Zen97]  M. Zennaro.  Asymptotic

stability analysis of Runge– Kutta methods for nonlinear systems of delay differential equations. Numerische Mathematik, 77(4):549–563, October 1997. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/7077004/70770549.
htm;   http://science.springer.
de/nmee/bibs/7077004/70770549.
htm.

Zerner:1994:AOF

[Zer94] Martin Zerner. An asymptotically optimal finite element scheme for the arch problem. Numerische Mathematik, 69(1):

117–123, November 1994. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://link.
springer.de/link/service/
journals/00211/bibs/4069001/
40690117.htm;   http://science.
springer.de/nmee/bibs/4069001/
40690117.htm.

Zettl:1970:VMF

[Zet70]  G. Zettl.  Ein Verfahren

zum Minimieren einer Funktion bei eingeschränktem Variationsbereich der Parameter. (German) [A method for the minimization of a function by a limited variation of parameters]. Numerische Mathematik, 15(5): 415–432, September 1970. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zenger:1978:IDS

[ZG78] C. Zenger and H. Gietl. Improved difference schemes for the Dirichlet problem of Poisson’s equation in the neighbourhood of corners. Numerische Mathematik, 30(3): 315–332, September 1978. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:1993:OHM

[ZH93] L.-Q. Zhang and G.-Q. Han.

Optimal homotopy methods for solving nonlinear systems. I. nonsingular homotopy paths. Numerische Mathematik, 65(4): 523–538, August 1993. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:1989:OOM

[Zha89] Shangyou Zhang. An optimal order multigrid method for biharmonic, C^{1} finite element equations. Numerische Mathematik, 56(6):613–624, December 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:1992:MSM

[Zha92a]  Xuejun Zhang.  Multilevel

Schwarz methods. Numerische Mathematik, 63(4):521–539, December 1992. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:1992:ABM

[Zha92b] Zhimin Zhang. Arch beam models: finite element analysis and superconvergence. Numerische Mathematik, 61(1): 117–143, February 1992. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:1993:TSA

[Zha93] Liqing Zhang. Two-sided approximation to periodic solutions of ordinary differential equations. Numerische Mathematik, 66(3):399–409, December 1993. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhang:1995:LRF

[Zha95] Zhimin Zhang. Locking and robustness in the finite element method for circular arch problem. Numerische Mathematik, 69(4):509–522, February 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/5069004/50690509.
htm;   http://science.springer.
de/nmee/bibs/5069004/50690509.
htm.

Zha:1996:CGS

[Zha96] Hongyuan Zha. Computing the generalized singular values/ vectors of large sparse or structured matrix pairs. Numerische Mathematik, 72(3):391–417, January 1996. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/6072003/60720391.
htm;   http://science.springer.
de/nmee/bibs/6072003/60720391.
htm.

Zheng:1989:EUV

[Zhe89] Q. Zheng. Über die Existenz von unendlich vielen periodischen Lösungen einer einfachen Temperaturregelung. (German) [On the existence of an infinite number of periodic solutions for a simple temperature control]. Numerische Mathematik, 55(3): 327–341, June 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zietak:1985:CSL

[Zie85] K. Zietak. The Chebyshev solution of the linear matrix equation AX + Y B = C. Numerische Mathematik, 46(3): 455–478, June 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ziegler:1997:GBB

[Zie97] Markus Ziegler. Generalized biorthogonal bases and tridiagonalisation of matrices. Numerische Mathematik, 77(3): 407–421, September 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/7077003/
                                        70770407.htm;   http://science.
                                        springer.de/nmee/bibs/7077003/
                                        70770407.htm.

Zitnan:1989:PEE

[Žit89] Peter Žitňan. An a posteriori error estimation of the Rayleigh–Ritz eigenvalues. Numerische Mathematik, 54(6): 639–654, March 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Ziv:1983:SME

[Ziv83] Abraham Ziv. A stable method for the evaluation of a polynomial and of a rational function of one variable. Numerische Mathematik, 41(3): 309–319, July 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zhou:1994:FAA

[ZL94] Aihui Zhou and Jichun Li. The full approximation accuracy for the stream function-vorticity-pressure method. Numerische Mathematik, 68(3):427–435, September 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068003/40680427.
htm;   http://science.springer.
de/nmee/bibs/4068003/40680427.
htm.

Zlamal:1968:FEM

[Zlá68]  Miloš Zlámal.  On the fi-

nite element method. Numerische Mathematik, 12(5): 394–409, December 1968. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zlamal:1969:SFE

[Zlá69] Miloš Zlámal. On some finite element procedures for solving second order boundary value problems. Numerische Mathematik, 14(1):42–48, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zlamal:1970:FEP

[Zlá70] Miloš Zlámal. A finite element procedure of the second order of accuracy. Numerische Mathematik, 14(4):394–402, March 1970. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zlatev:1978:SPV

[Zla78] Zahari Zlatev. Stability properties of variable stepsize variable formula methods. Numerische Mathematik, 31(2): 175–182, June 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zlatev:1981:ZSP

[Zla81] Zahari Zlatev. Zero-stability properties of the three-ordinate variable stepsize variable formula methods. Numerische Mathematik, 37(2):157–166, July 1981. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zouraris:1997:CRK

[Zou97] Georgios E. Zouraris. Convergence of Runge–Kutta approximations for parabolic problems with Neumann boundary conditions. Numerische Mathematik, 77(1):123–142, July 1997. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

                                        tronic). URL http://link.
                                        springer.de/link/service/
                                        journals/00211/bibs/7077001/
                                        70770123.htm;   http://science.
                                        springer.de/nmee/bibs/7077001/
                                        70770123.htm.

Zou:1999:AQL

[Zou99] Xiulin Zou. Analysis of the quasi-Laguerre method. Numerische Mathematik, 82(3):

491–519, May 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer-
ny.com/link/service/journals/
00211/bibs/9082003/90820491.
htm;   http://link.springer-
ny.com/link/service/journals/
00211/papers/9082003/90820491.
pdf.

Zulehner:1988:SPS

[Zul88] Walter Zulehner. On the solutions to polynomial systems obtained by homotopy methods. Numerische Mathematik, 54(3): 303–317, December 1988. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zurmuhl:1963:BVH

[Zur63] R. Zurmühl. Bemerkungen zum Verfahren von Hessenberg. (German) [Remark on a method of Hessenberg]. Numerische Mathematik, 4:377–380, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zurmuhl:1964:AGT

[Zur64] R. Zurmühl. Zur angenäherten ganzrationalen Tschebyscheff-Approximation mit Hilfe trigonometrischer Interpolation. (German) [Toward approximate fairly rational Chebyshev approximation with the aid of trigonometric interpolation]. Numerische Mathematik, 6(1):1–5, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Zenisek:1995:ITN

[ŽV95] Alexander Ženíšek and Michèle Vanmaele. The interpolation theorem for narrow quadrilateral isoparametric finite elements. Numerische Mathematik, 72(1):123–141, November 1995. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

                                        URL http://link.springer.
                                        de/link/service/journals/
                                        00211/bibs/5072001/50720123.
                                        htm;   http://science.springer.
                                        de/nmee/bibs/5072001/50720123.
                                        htm.

Zwas:1973:TSL

[Zwa73] Gideon Zwas. On two step Lax–Wendroff methods in several dimensions. Numerische Mathematik, 20(5):350–355, October 1973. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).